A Complete Bibliography of Publications in Computers and Mathematics with Applications (1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 May 2024 Version 1.03

Title word cross-reference

(2 + 1) [DK97c]. (C_{0})m [ZA95]. (F, ρ) [Liu96a, Liu96b]. (h, k) [LFP97]. (n, n) [AP94]. (n, p) [WA96a]. (n ≤ 9) [GF90]. (S)_{+}^{1} [CY95]. 0 [Ano97-85]. 1 [Ano97-85, AB93, BBO99, BGL97, MM93a, O’R98b, SJB97]. 1/f [Ano99-162]. $10.00 [Ano96-30]. $10.95 [Ano97-158]. $103.00 [Ano98-59]. $105.00 [Ano98-71, Ano99z, Ano99-27, Ano99-87, Ano99-114, Ano99-153, Ano99-197, Ano99-83, Ano99-131]. $106.00 [Ano99-181]. $106.50 [Ano99-86]. $109.00 [Ano97-108, Ano96e]. $110.00 [Ano98-101, Ano99-57, Ano99-209, Ano96-123]. $112.00 [Ano97-157, Ano99-117]. $114.00 [Ano98-83]. $115.00 [Ano98-89, Ano98-110, Ano98-137, Ano98-144, Ano99-50, Ano99-94, Ano99-93]. $118.00 [Ano99-121]. $119.00 [Ano98-102]. $12.50 [Ano96-47, Ano96-97, Ano97-156, Ano97-160]. $120.00 [Ano97-56, Ano98-37, Ano98-100, Ano99-88]. $122.00 [Ano99-229]. $124.00 [Ano98i]. $125.00 [Ano97-105, Ano98-133, Ano98-156, Ano99o, Ano99-99, Ano99-100, Ano99-169]. $126.00 [Ano98-50]. $129.00 [Ano99-142, Ano99-171, Ano99-183, Ano99-184]. $129.25 [Ano97-164]. $13.75 [Ano95-75]. $13.95 [Ano96-63, Ano95n, Ano95-39, Ano95-40, Ano96-76, Ano96-77, Ano97-120]. $130.00 [Ano99-234]. $133.00 [Ano98-153, Ano99-137]. $135.00 [Ano97-36, Ano98g, Ano98l, Ano99-213, Ano99-211, Ano96v]. $14.50 [Ano97-153]. $14.95 [Ano96-60, Ano95-71, Ano95-54, Ano96-55, Ano97-146]. $140.00 [Ano99-84, Ano99-85, Ano99-80, Ano99-96, Ano99-97, Ano99-165, Ano99-206, Ano99-207]. $145.00 [Ano98p, Ano98-56, Ano98-132]. $149.00 [Ano98-116, Ano99r, Ano99-52]. $15.00 [Ano99-224, Ano95-27, Ano96-87, Ano96-128, Ano96-162, Ano97-80]. $15.95 [Ano96a, Ano96-143, Ano97-40, Ano98-39, Ano95-44]. $150.00 [Ano96-80, Ano97-64]. $157.00 [Ano99-204]. $159.00 [Ano97-66, Ano99-49, Ano99-196, Ano99-212]. $16.00 [Ano96l]. $16.50 [Ano96-154, Ano97-150]. $16.95 [Ano95-29, Ano96-163, Ano96-105]. $160.00 [Ano98-72]. $162.50 [Ano96-81]. $164.00 [Ano99-113]. $165.00 [Ano98o, Ano98-84]. $168.00 [Ano99-170, Ano99-182]. $169.00 [Ano98-57, Ano98-74, Ano98-155]. $17.00 [Ano99h, Ano99i, Ano96-121]. $17.50 [Ano96-115, Ano97s]. $17.95 [Ano96-50, Ano96-108, Ano96-130, Ano96-141, Ano97-154, Ano95-70, Ano96u, Ano96-116]. $170.00 [Ano99-136, Ano99-172]. $173.00 [Ano99u, Ano99-174]. $174.00 [Ano98-157]. $175.00 [Ano98-66]. $179.00 [Ano99-167, Ano99-168]. $18.00 [Ano99d, Ano99-233, Ano96-44]. $18.50 [Ano96-160]. $18.95 [Ano95d, Ano95y, Ano97-39, Ano97-127, Ano98-145, Ano98-146, Ano96-29, Ano97-63]. $182.00 [Ano98j, Ano98-108, Ano98-126]. $183.00 [Ano98r]. $187.50 [Ano96-83, Ano97-65]. $189.00 [Ano99-152]. $19.00 [Ano99-144]. $19.95 [Ano97-82, Ano97-151, Ano98-60, Ano95-58, Ano95-59, Ano96-92, Ano97-155, Ano97-169, Ano98-42, Ano98-79, Ano98-103, Ano98-129, Ano98-130, Ano99-143]. $190.00 [Ano98-127]. $194.00 [Ano98-92, Ano98-111]. $195.00 [Ano97-58, Ano99-120]. $197.00 [Ano98w]. $198.00 [Ano98-121]. $199.00 [Ano98n]. 2 [Ano99-138, Bel97a, BS95, BRGJ95, Bri97, CC97, Com92, HTT92, HGG99, LM96a, RD99, SJB97]. $20.00 [Ano95-32, Ano96-32, Ano96-100, Ano97v, Ano99-75, Ano99-76]. $20.95 [Ano96-54]. $204.00 [Ano99-101]. $205.75 [Ano96-79]. $208.00 [Ano99-95]. $21.00 [Ano99-51]. $21.95 [Ano98-52, Ano98-120]. $210.00 [Ano97-29, Ano99-128]. $214.00 [Ano98-32]. $214.25 [Ano95-65]. $22.00 [Ano97-79, Ano95-43, Ano95-66]. $22.33 [Ano96-102]. $22.50 [Ano95x, Ano95-69, Ano97-43, Ano99-47, Ano99-48]. $22.95 [Ano97n, Ano97-117, Ano95v, Ano95-76, Ano96-57, Ano99j]. $225.00 [Ano98-135, Ano99-33]. $229.50 [Ano96-82]. $238.00 [Ano99-115, Ano99-116]. $24.00 [Ano99-140, Ano95-67, Ano96-106]. $24.50 [Ano99-78, Ano99-79]. $24.95 [Ano96-136, Ano96-144, Ano97-143, Ano97-94, Ano97-101, Ano97-103, Ano97-135, Ano97-162, Ano97-129, Ano98-70, Ano98-128, Ano99-89, Ano99-215, Ano95-77, Ano95-63, Ano95-82, Ano96g, Ano96-42, Ano96-152, Ano96-51, Ano96-167, Ano99-31, Ano99-62, Ano99-173, Ano99-187, Ano99-210, Ano99-237]. $245.00 [Ano97e]. $248.00 [Ano98-65]. $25.00 [Ano96-149, Ano97-141, Ano97-128, Ano99-39, Ano99-225, Ano99-202, Ano99-239, Ano93f, Ano96c, Ano96-140, Ano96-139, Ano97-110, Ano97-133, Ano98-28, Ano98f, Ano98-151, Ano99-216, Ano99-217, Ano99-222, Ano99-223]. $25.95 [Ano96-98]. $252.00 [Ano99l]. $254.00 [Ano99s]. $26.25 [Ano90q]. $26.95 [Ano96-62]. $260.00 [Ano98-123]. $268.00 [Ano99t]. $27.50 [Ano90u, Ano96-118, Ano97-68]. $27.95 [Ano96r, Ano96-88, Ano97-148, Ano99m]. $28.00 [Ano96-89]. $28.95 [Ano95-31, Ano95u]. $29.00 [Ano96-156]. $29.95 [Ano95-74, Ano95-80, Ano96-64, Ano97-90, Ano97-98, Ano97-99, Ano97-97, Ano97-159, Ano98-73, Ano98-122, Ano99-111, Ano99-112, Ano90s, Ano95o, Ano95i, Ano95-33, Ano95-46, Ano96w, Ano96-28, Ano96-49, Ano96-69, Ano96-70, Ano96-93, Ano96-137, Ano96-142, Ano97-38, Ano97-35, Ano97-75, Ano97-77, Ano97-89, Ano97-125, Ano97-166, Ano97-168, Ano98-87, Ano98-86, Ano98-97, Ano98-152, Ano99-61, Ano99-73, Ano99-138]. $295.00 [Ano98-49]. 2n [AP94]. 3 [BMN96, Bel94a, Com92, Dea98, GCLT98, KKS98, Kle93, Kor97, Li98d, Mah90, Mol96, MZ95, RW91, ZIKM93, ZM96]. $3.00 [Ano97-141]. $30.00 [Ano98-27, Ano98-134, Ano90p, Ano93k, Ano95-68, Ano96s, Ano96-66, Ano96-85, Ano96n, Ano96-158, Ano97-50, Ano98-48, Ano99k, Ano99-232]. $31.00 [Ano96-127, Ano96-65]. $31.95 [Ano95-30, Ano95g]. $32.00 [Ano96k, Ano99-106, Ano99-107]. $32.50 [Ano96x, Ano96-58, Ano96-114, Ano97-32]. $32.95 [Ano97-78, Ano97-95, Ano99-180, Ano95-42, Ano97-167, Ano98-33, Ano98-35, Ano98-54, Ano98-88, Ano98-99, Ano99-72, Ano99-139, Ano99-141, Ano99-195, Ano99-188, Ano99-235, Ano99-236]. $32.95haotic [Ano96-56]. $34.00 [Ano96-107]. $34.38 [Ano95-79, Ano97-107]. $34.95 [Ano97-102, Ano99y, Ano99-218, Ano99-156, Ano99-219, Ano99-157, Ano99-147, Ano95-62, Ano97-76, Ano98-51, Ano99-164, Ano99-148, Ano99-166, Ano99-192, Ano99-238]. $35.00 [Ano95-29, Ano99-135, Ano93l, Ano95q, Ano95-28, Ano96-36, Ano96-52, Ano96-94, Ano96-99, Ano96-125, Ano96-148, Ano97-31, Ano97-37, Ano97-55, Ano97-69, Ano97-93, Ano97-109, Ano97-139, Ano98-38, Ano98-82, Ano98-112, Ano99f, Ano99g, Ano99-122, Ano99-158, Ano99-160, Ano99-159, Ano99-161, Ano99-220]. $35.45 [Ano96-119]. $35.95 [Ano96-161, Ano96-110]. $350.00 [Ano97-45, Ano99-105, Ano99-103, Ano99-102, Ano99-104]. $36.00 [Ano96-71].

$36.53 [Ano95f]. $36.95

[Ano96-59, Ano95-72, Ano95-56, Ano99-176]. $37.50 [Ano97-154, Ano98-148, Ano93u, Ano93i, Ano96-90, Ano97-144, Ano98-77, Ano99-38, Ano99-163, Ano99-226]. $37.95 [Ano96-147, Ano98-145, Ano98-146]. $38.00 [Ano96-143, Ano99-146]. $39.00 [Ano97-128]. $39.95 [Ano95r, Ano95s, Ano96-60, Ano97-94, Ano97-124, Ano97-131, Ano98q, Ano98k, Ano98-149, Ano98-62, Ano99-35, Ano99-36, Ano99-108, Ano99-154, Ano99-155, Ano99-177, Ano99-193, Ano99-194, Ano93t, Ano93w, Ano95-37, Ano95-47, Ano95-55, Ano96-35, Ano96-96, Ano96-133, Ano96-166, Ano97-28, Ano97-41, Ano97-60, Ano97-74, Ano98s, Ano98-40, Ano98-75, Ano98-85, Ano98-98, Ano98-119, Ano98-158, Ano99-150]. $395.00 [Ano98-67]. 3x + 1 [LV92]. 4 [Kun95]. $40.00 [Ano95d, Ano99d, Ano99-203, Ano99-91, Ano90o, Ano90x, Ano96-38, Ano96-95, Ano96-84, Ano96-145, Ano97l, Ano97r, Ano97i, Ano97d, Ano97-115, Ano97-161, Ano98u, Ano98-34]. $40.95 [Ano96-109]. $41.95 [Ano95-50, Ano99e]. $42.00 [Ano93o]. $42.95 [Ano99-162, Ano95-34, Ano95-48, Ano96z, Ano97-47]. $44.06 [Ano95-57]. $44.50 [Ano90w]. $44.95 [Ano98-41, Ano99-82, Ano99-132, Ano99-230, Ano96-124, Ano97-123, Ano98-150]. $45 [Ano96-50]. $45.00 [Ano96-112, Ano97-59, Ano97-82, Ano97-130, Ano98-91, Ano99-77, Ano93j, Ano93s, Ano95h, Ano95-38, Ano95-78, Ano95-60, Ano95-81, Ano96-39, Ano96-33, Ano96-31, Ano96-131, Ano96-138, Ano97u, Ano97k, Ano97-134, Ano98-58, Ano99-40, Ano99-123, Ano99-124]. $45.14 [Ano96-86]. $45.95 [Ano95-61]. $46.95 [Ano99-179, Ano96-27]. $47.29 [Ano95e]. $47.50 [Ano95-73]. $48.00 [Ano99-32, Ano99-231]. $48.50 [Ano99-151]. $48.95 [Ano96-45]. $49.00 [Ano98-44, Ano99h, Ano99i, Ano99-224, Ano99-189, Ano96-165]. $49.43 [Ano95-64].

$49.44 [Ano96-151, Ano96-48]. $49.50 [Ano95t, Ano96-111]. $49.95 [Ano95r, Ano95s, Ano97-121, Ano97-162, Ano98t, Ano98-80, Ano99v, Ano99w, Ano99-55, Ano99-56, Ano99-129, Ano99-130, Ano93r, Ano95j, Ano96m, Ano96-134, Ano97p, Ano97-83, Ano97-87, Ano98-29, Ano98-147, Ano98-163, Ano99-228]. $50.00 [Ano96-130, Ano98v, Ano99-233, Ano96h, Ano96-153, Ano96-40, Ano96-46, Ano96-53, Ano96-132, Ano97f, Ano97-49, Ano97-81, Ano97-149, Ano97-100, Ano98-61, Ano98-63, Ano99-81]. $50.95 [Ano95-35]. $51.59 [Ano99-65]. $51.95 [Ano99-34]. $52.00 [Ano96-141]. $53.00 [Ano99-140]. $54.00 [Ano96-91]. $54.95 [Ano96b, Ano96r, Ano99-60, Ano99-109, Ano99-110, Ano96t, Ano96-155, Ano96-146, Ano98-106]. $55.00 [Ano95y, Ano96-108, Ano96-122, Ano97-39, Ano97-127, Ano90u, Ano93v, Ano95-36, Ano95-41, Ano95-52, Ano96-34, Ano96-61, Ano96-78, Ano96p, Ano96o, Ano97-42, Ano97-44, Ano97-137, Ano97-165, Ano99p]. $55.95 [Ano97-27]. $56.00 [Ano98-131]. $57.00 [Ano97-151]. $58.95 [Ano97-145]. $59.00 [Ano98-136, Ano93n]. $59.95 [Ano95-51, Ano95-49, Ano97n, Ano97q, Ano97-72, Ano97-90, Ano97-159, Ano98-105, Ano99-41, Ano99-42, Ano99-66, Ano99-67, Ano99-69, Ano99-208, Ano95-53, Ano96-72, Ano96-157, Ano96-126, Ano96d, Ano97-62, Ano98-30, Ano98-81, Ano99-28]. $60.00 [Ano99-225, Ano95k, Ano96i, Ano96-74, Ano96-73, Ano97g, Ano97-136, Ano97-52, Ano98m, Ano98-104, Ano98-113, Ano98-114, Ano99-92, Ano99-118, Ano99-119]. $61.95 [Ano96y]. $62.00 [Ano97-116]. $64.50 [Ano96-129]. $64.95 [Ano97-98, Ano97-99, Ano97-97, Ano97-117, Ano98-70, Ano98-128, Ano98-125, Ano99-44, Ano99-46, Ano95c, Ano95z, Ano97-142, Ano97-54, Ano98-141]. $65.00 [Ano95-69, Ano96-144, Ano97h, Ano97-135, Ano93m, Ano96j, Ano96-41, Ano96-67, Ano97-53, Ano97-91, Ano97-118, Ano98y, Ano98-55, Ano99-29, Ano99-30, Ano99-90, Ano99q, Ano99-98]. $67.95 [Ano96-150, Ano96-101]. $68.00 [Ano95x, Ano99-214, Ano97-119]. $69.00 [Ano96-159]. $69.95 [Ano97-30, Ano97-129, Ano98-46, Ano99x, Ano99-43, Ano99-125, Ano99-126, Ano95w, Ano96f, Ano97-106, Ano98-142, Ano98-159]. $699.00 [Ano98-162]. $7.95 [Ano99-178]. $70.00 [Ano96-135, Ano98h]. $72.50 [Ano97-126, Ano95p]. $72.95 [Ano96-75, Ano96-117]. $74.00 [Ano96-64]. $74.67 [Ano99-185]. $74.95 [Ano98-73, Ano98-78, Ano99-68, Ano99-133, Ano99-134]. $75.00 [Ano95-30, Ano98-161, Ano95-45, Ano96-103, Ano97-57, Ano98-118, Ano99-205]. $75.95 [Ano98-143]. $77.95 [Ano95m, Ano95l]. $79.00 [Ano99-58, Ano99-135]. $79.95 [Ano98-43, Ano98-164, Ano99-37, Ano99-145, Ano99-221, Ano99-186, Ano96-164, Ano99-199, Ano99-200]. $80.00 [Ano97-95, Ano96-104]. $84.95 [Ano98-107]. $85.00 [Ano97x, Ano97y, Ano97z, Ano98-138, Ano99-154, Ano99-155, Ano97-88, Ano98-96, Ano98-94, Ano98-95, Ano98-124]. $87.00 [Ano93g]. $87.50 [Ano97-104, Ano98-68, Ano98-139]. $87.67 [Ano97-113]. $87.75 [Ano98-109]. $88.50 [Ano98x]. $89.50 [Ano98-160, Ano95-83, Ano96-113]. $89.95 [Ano99-149, Ano99-198, Ano96-37, Ano97-132, Ano98-117]. $90.00 [Ano97-59, Ano98-53]. $90.95 [Ano99-190, Ano99-191]. $93.95 [Ano99-53, Ano99-54]. $94.00 [Ano98-47, Ano99-201]. $95.00 [Ano96-161, Ano97-84, Ano98k, Ano98-27, Ano98-76, Ano93q, Ano96-43, Ano97-48]. $96.00 [Ano97-163]. $97.00 [Ano99-63, Ano99-64, Ano93h]. $97.50 [Ano98-31, Ano98-45]. $98.00 [Ano98z, Ano98-69, Ano99n, Ano99-227, Ano93p].

$98.50 [Ano97-85]. $99.00 [Ano97o, Ano97-130, Ano98-36, Ano98-90, Ano98-115, Ano98-140, Ano99-70, Ano98-93]. 1 [KM97b]. ^{nd} [Ano96-60]. ^{r}d [Ano99-221]. ^{st} [Iri98]. _{1} [CK97b]. _{2} [CK97b]. _{G} [YQM99]. A [AO99b, Wu98]. α [MC97]. B [Maz90, Shv95]. c [Yan93b]. C^{1} [DO91, xLhW96, Luo98, PR91, Wac91a]. C^{2} [Luo98]. C^{3} [Sch96]. C^{k} [Rou92, RZ94]. cosh(x) [QZ93c]. c → ∞ [Bax92]. E(XZ|Y + ϵX) [Sza90a]. ϵ [Ano95f, LY99b]. G [Ver99b]. g_{ij} [Mur91a]. H [Gla93a, HS95, RP97, SH95b, Ver99b, WY99, IESAH95, IB95, RM98b]. H_{2} [Gre92a]. H_{∞} [HG98]. hp [Ano99-205, BBO99, PSS99]. I [BN92]. ∫ e^{−x^{2}} dx [Bas98]. J [Ano99-245, FLW97]. K [Ano98-111, Bad95, CK97a, CL99c, Dwy93]. κ [DKW98]. L [CKAS96, KM97b, LK99, MHBD96]. L^{2} [ST95]. L^{2}R^{2} [KV95]. L^{∞} [GW91]. L_{0} [CV90]. L_{2} [RT94]. L_{p} [Lub97]. L_{p}[0, 1] [Hu95]. LU [DHW98, LC97b, Tsa90]. m [KH92]. m^{2} + y^{2}n^{2} [AB93]. N [RSC93, WSNS98, APB97, CTG96, CK97a, Hin95, Lin91b, WSDT94]. n^{m} [DWC95]. O(2) [Köh96a, Köh96b]. ω [GSST97]. ω_{n}^{k} [BIZ97]. ω_{n}^{k} = n^{k2} ∫_{−∞}^{∞}[S_{n}(x) − F (x)]^{k} dF (x) [IZ97]. P [ZL99c, Bro98, GK98, RR97a]. p(n) [Bar91a]. P_{n}(cos θ)/p(n) [Bar91a]. ∂^{2}∂T^{2}u(x, t) + ∂∂T α(u(x, t)) = ∂^{2}∂X^{2}β(u(x, t)) [Blo91]. Φ [Din97b, Hua98c, Yos90]. Π_{n} [CP95a]. Ψ [Chi91]. Q [Maz90, KS96, RSS94b]. QR [LVV96]. R [AI97, HK96, LWL96]. R^{2} [CF91a]. R^{m} [PLH95]. R^{n} [Gur91, Hua97]. ρ [DBM95]. S [Ano99-193, Ano99-194, Com92]. Σ [Sni90]. Σ1(a_{i} log a_{i}) [Cla98, EZ93]. sinh(x) [QZ93c]. \sqrt{6} [AB93]. T [Li96c, VAB91, WJ90]. T^{−1}T [Li96c]. tanh(x) [QZ93c]. φ(r) = (r^{2} + c^{2})^{12} [Bax92]. W_{3} [Ano97-140]. X [Raq98]. x^{2} + ny^{2} [Ano97-121]. x_{n+1} = (βx_{n}^{2})□(1 + x_{n−1}^{2})

[CLRN94]. y [AB93]. Y_{0}(z) [ZB95, Zha96]. Y_{1}(z) [Zha95, Zha96].

-Adams [DBM95]. -adic [RR97a]. -analogue [KS96]. -Approximation [KM97b]. -ary [CK97a, KH92]. -convex [Liu96a, Liu96b]. -convexity [Ver99b]. -criteria [BIZ97]. -cube [CK97a]. -D [CC97, GCLT98, RD99, SJB97, BBO99, BS95, Bri97, Dea98, Kle93, Kor97, LM96a, Mah90, SJB97, ZIKM93]. -dense [MC97]. -dimensional [Bro98, DK97c, Hin95]. -disconjugacy [AP94]. -domains [LK99]. -equivalence [GSST97]. -error [WJ90]. -faces [Dwy93]. -function [SH95b]. -functionals [Bad95]. -functions [HS95, RP97]. -fuzzy [LWL96]. -generating [RSS94b]. -hemicontractive [Hua98c]. -hypertournament [RM98b]. -interpolation [xLhW96]. -irreducibility [ZL99c]. -Lanczos [Ano99-245, FLW97]. -loops [Ano98-111]. -means [Yan93b]. -minimum [DKW98]. -nephron [WSDT94]. -NN [CL99c]. -norm [RT94]. -norms [Lub97]. -notation [Sni90]. -Optimality [LY99b]. -parameter [ZA95]. -person [Bel94a, Bel97a]. -point [RSC93]. -proper [AO99b]. -rational [DO91, PR91]. -set [Lin91b, O’R98b]. -smooth [Luo98]. -spaces [Ver99b, WY99]. -stability [YQM99]. -stable [CV90, CKAS96, MHBD96, Wu98]. -Step [IESAH95]. -strongly [Din97b]. -sum [Com92]. -system [Gla93a]. -th [AP94]. -transformation [Chi91]. -trees [Shv95]. -trichotomies [LFP97]. -uniform [ST95]. -version [IB95, PSS99].

/1/N [CTG96]. /G/1 [CK97b].

1 [AO93b, CK97b, Dem94a, Gla96b, Toy92, TT92, YFF92b, Ano95-37, Ano96-81, Ano96-80, Ano96-127]. 1-2-3 [Ano95-39]. 1.4 [Ano99-69]. 10/11 [BM90c]. 101.50 [Ano99-52]. 103.00 [Ano98-155]. 107.00 [Ano98-66, Ano99-172]. 108.00 [Ano99-136]. 108.80 [Ano98-72]. 109.00 [Ano98j, Ano98r, Ano98-108, Ano98-126, Ano99u, Ano99-174]. 11-th [Ano99-212]. 11.25 [Ano96-63]. 110.00 [Ano98-127]. 113.00 [Ano97-51, Ano99-167, Ano99-168]. 116.00 [Ano98-92, Ano98-111]. 118.00 [Ano98w]. 119.00 [Ano99-101, Ano99-152, Ano99-165]. 12.75 [Ano96a, Ano97-40, Ano98-39]. 120.00 [Ano97-58, Ano98-121]. 122.00 [Ano98n]. 126.00 [Ano97-29]. 128.00 [Ano98-32]. 13-15 [Ano99-152]. 13.50 [Ano96-163]. 130.00 [Ano99-95]. 132.00 [Ano99-120]. 135.00 [Ano98-135, Ano99-33]. 13th [Ano90a]. 143.00 [Ano99-128]. 146.00 [Ano97e]. 147.00 [Ano99l]. 149.00 [Ano98-65, Ano99-115, Ano99-116]. 15.25 [Ano99-144]. 15.95 [Ano98-60]. 157 [Ano98-93]. 159.00 [Ano98-123]. 16.75 [Ano99-51]. 160.00 [Ano99s]. 169.00 [Ano99t]. 17 [Kou90b]. 17-21 [Ano99-136]. 17.50 [Ano99-202]. 175.00 [Ano97-84]. 179.00 [Ano98-49]. 18 [BM90c, Zad92]. 18.95 [Ano99-215]. 185.00 [Ano97-85]. 1859 [Ano97h]. 1859-1940 [Ano97h]. 1892-1950 [Ano99-239]. 19 [MG90b]. 19.00 [Ano99-89]. 19.50 [Ano99-111, Ano99-112, Ano99-218, Ano99-219]. 19.95 [Ano97-143, Ano99-39, Ano99-239]. 1922-1996 [Eso99c]. 1940 [Ano97h]. 1946 [Ano99-146]. 1946-1972 [Ano99-146]. 1968-1988 [Har90]. 1972 [Ano99-146]. 1989 [Zad92]. 1991 [Ado91a, Sal92b, Ano95-61]. 1991-1994 [Ano95-61]. 1993 [Ano94a]. 1994 [Ano95-61]. 1995 [Ano97-47, Ano98-111]. 1996 [Ara99b, Ano96e, Ano97-57, Ano98-84]. 1997 [Ano97-51, Ano98-83, Ano99-136, Ano99-137, Ano99-152]. 1998 [Ano99t, Ano99-120].

2 [YFF92a, Ano93v, Ano97f]. 2.0 [Ano93z]. 20 [Ano96-136]. 20-22 [Ano98-155]. 21 [Ado91a, Moo93b]. 21-25 [Ano99-120]. 21-st [Ano98-110]. 22 [Sal92b]. 22.50 [Ano96-89]. 23.25 [Ano95-31]. 23.95 [Ano95-74]. 24.50 [Ano98q]. 240.00 [Ano97-164]. 249.00 [Ano98-67]. 25 [MM94]. 25.00 [Ano99-59]. 25.50 [Ano95-30]. 25.95 [Ano96x, Ano98-134]. 26.00 [Ano98-91, Ano99-132, Ano99-156, Ano99-157]. 27.00 [Ano99-108]. 29.50 [Ano99-109, Ano99-110]. 29.95 [Ano97-121, Ano98-148, Ano98-154].

3 [Ano95-36, Ano96-159]. 3.0 [Ano98-162]. 30.50 [Ano99v, Ano99w, Ano99-82, Ano99-146]. 300.00 [Ano96-83, Ano97-65]. 31.95 [Ano99-91]. 32.50 [Ano98-62, Ano99-77, Ano99-193, Ano99-194]. 34.00 [Ano97-33, Ano99-66, Ano99-55, Ano99-67, Ano99-56, Ano99-162]. 35 [Ano95t, Li10]. 350-page [Ano98-162]. 37.50 [Ano96b, Ano99-44, Ano99-43, Ano99-189, Ano99-208]. 375.00 [Ano95-65]. 38.00 [Ano99-60]. 39.95 [Ano98v].

4 [Ano96-156]. 44.75 [Ano98-131]. 47.25 [Ano98-136]. 47.50 [Ano99-41, Ano99-42]. 48.00 [Ano96-68, Ano97m, Ano97-126]. 48.50 [Ano99-127, Ano99-175]. 49.50 [Ano99x, Ano99-46].

51.00 [Ano98-161]. 51.50

[Ano97t, Ano97-46, Ano97-152]. 52.00 [Ano99-125, Ano99-126]. 53.00 [Ano98-164, Ano99-145]. 53.50 [Ano99-198]. 53.75 [Ano99-58]. 54.50 [Ano99-214]. 55.00 [Ano97-140, Ano98-53, Ano99-71]. 56.25 [Ano98-138]. 56.50 [Ano97-70]. 57.75 [Ano98-68, Ano98-109, Ano98-139]. 58.50 [Ano98x]. 59.00 [Ano98-36, Ano98-90, Ano98-115, Ano98-140, Ano99n]. 5th [Ano90b].

6 [Szk95]. 6/7 [Zad92]. 60.00

[Ano97-71, Ano99-74]. 61.00 [Ano99-186].

62.00 [Ano98-59, Ano99-63, Ano99-64, Ano99-153]. 62.25 [Ano97-104, Ano98-47]. 62.75 [Ano98-76]. 64.00 [Ano98-43, Ano99-201]. 64.35 [Ano98-31]. 64.50 [Ano98-45]. 64.70 [Ano98z, Ano98-69]. 65.00 [Ano97-108, Ano98-100, Ano99-45]. 650.00 [Ano97-45]. 66.00 [Ano99z, Ano99-27]. 67.00 [Ano98-101, Ano99-181, Ano99-197]. 67.00.95 [Ano97-157]. 68.00 [Ano98-83]. 68.25 [Ano97-163]. 69.00 [Ano97-34, Ano97-86, Ano97-138, Ano99-149].

7 [Ano96i]. 70.00 [Ano98-89]. 71.50 [Ano98-71]. 72.00 [Ano97-56, Ano99-87, Ano99-114]. 72.50 [Ano99-86]. 75.00 [Ano98-64, Ano99-57, Ano99-209]. 75.90 [Ano98-144]. 76.00 [Ano98i]. 76.20 [Ano99-117]. 77.50 [Ano97j]. 78.20 [Ano98-137, Ano99-93]. 78.25 [Ano99-50, Ano99-94]. 79.50 [Ano98-37].

80.00 [Ano98-110]. 80.25 [Ano99-121]. 81.00 [Ano98-153]. 81.75 [Ano99-88]. 82.00 [Ano99-171, Ano99-183, Ano99-184]. 82.50 [Ano98-156]. 83.00 [Ano99-229]. 83.20 [Ano98-50]. 84.00 [Ano99-137, Ano99-142]. 85.00 [Ano98-133, Ano99o, Ano99-99, Ano99-100, Ano99-169]. 86.50 [Ano97-122]. 88.75 [Ano97-105]. 89.00 [Ano98-116, Ano99-84, Ano99-85, Ano99-96, Ano99-97, Ano99-206, Ano99-207]. 89.95 [Ano99-212].

9.50 [Ano95-75]. 91.80 [Ano98g, Ano98l, Ano99-213]. 91m [Moo93b]. 92.00 [Ano99-211]. 95 [Ano97-34]. 95.00 [Ano97-66]. 95.25 [Ano99-80]. 95.70 [Ano98-56]. 95.75 [Ano98-132]. 95.85 [Ano97-36]. ’96 [Ano97-88]. 96.00 [Ano99r]. 99.00 [Ano98o, Ano98-84, Ano98-74, Ano99-49, Ano99-170, Ano99-182, Ano99-196, Ano99-204].

= [Ano99-99, Ano99-100, Maz90].

A. [Los92]. AAAI [Ano95-47, Ano96h, Ano96-44, Ano96-36, Ano96-53, Ano96-135, Ano97r, Ano98-77, Ano99-40]. AAI [Ano98-112]. Aaron [Ano95-72]. Abbate [Ano96-148]. Abbott [Ano96-153]. Abdelsalam [Ano97-163, Ano98-164]. Abdugafur [Ano98-89]. Abel [MGM92]. Abelson [Ano97-137]. Abhiram [Ano99x]. Abiteboul [Ano95-35]. Ablex [Ano96j]. Abramov [Ano98-36]. Abrial [Ano97-142]. ABS [Gal99a]. Absolute [AKK^{+}97, CL94a, Gre90c]. Absorption [Dor99]. Abstract [Yao93, Ano96-55, Ano97-136, Ano99-77, GS96a, LX98]. abstraction [Ano99-44, Neu91, tC95]. Abundance [BGL97, Ano96-121]. AC [Kuu94]. AC-driven [Kuu94]. Academic [Ano90r, Ano90t, Ano90v, Ano93m, Ano93q, Ano97e, Ano97-29, Ano97-36, Ano97-45, Ano97-58, Ano97-56, Ano97-66, Ano97-85, Ano97-84, Ano97-108, Ano97-105, Ano97-104, Ano97-126, Ano97-157, Ano97-164, Ano97-163, Ano98g, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98w, Ano98x, Ano98r, Ano98z, Ano98-37, Ano98-31, Ano98-32, Ano98-36, Ano98-45, Ano98-49, Ano98-53, Ano98-50, Ano98-47, Ano98-59, Ano98-56, Ano98-57, Ano98-65, Ano98-66, Ano98-68, Ano98-67, Ano98-69, Ano98-71, Ano98-72, Ano98-76, Ano98-84, Ano98-74, Ano98-83, Ano98-89, Ano98-90, Ano98-92, Ano98-109, Ano98-108, Ano98-100, Ano98-101, Ano98-110, Ano98-115, Ano98-111, Ano98-116, Ano98-121, Ano98-123, Ano98-126, Ano98-127, Ano98-133, Ano98-137, Ano98-132, Ano98-135, Ano98-144, Ano98-138, Ano98-139, Ano98-140, Ano98-153, Ano98-157, Ano98-155, Ano98-156, Ano98-160, Ano98-161, Ano98-164]. Academic [Ano99z, Ano99-27, Ano99o, Ano99r, Ano99s, Ano99n, Ano99t, Ano99u, Ano99-33, Ano99-50, Ano99-213, Ano99-49, Ano99-63, Ano99-64, Ano99-52, Ano99-57, Ano99-58, Ano99-84, Ano99-85, Ano99-86, Ano99-83, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-87, Ano99-105, Ano99-103, Ano99-99, Ano99-96, Ano99-102, Ano99-104, Ano99-100, Ano99-97, Ano99-95, Ano99-101, Ano99-115, Ano99-116, Ano99-117, Ano99-120, Ano99-114, Ano99-128, Ano99-121, Ano99-136, Ano99-137, Ano99-131, Ano99-145, Ano99-142, Ano99-152, Ano99-153, Ano99-167, Ano99-168, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-170, Ano99-174, Ano99-183, Ano99-184, Ano99-181, Ano99-182, Ano99-196, Ano99-197, Ano99-201, Ano99l, Ano99-206, Ano99-207, Ano99-204, Ano99-209, Ano99-229, Ano99-212, Ano99-211, Ano99-227, Ano99-234, Ano97-55]. accelerate [BG91b]. accelerating [BDT95, WW95]. acceleration [EDD91]. Access [CB93a, CFLC95, JW93, LH97, WCY94, YL93b, Ano99m]. accounts [Ueb92]. accretive [Cha99b, Din97b, HB99]. accuracy [CHSS95, CC94a, FY93, FK98a, HWL94, KC92a, Mac93, Mat99, Moo91, Moo93b, Moo94, RD99, Tsa90]. Accurate [JA96, Doh90, Gla93e]. Achen [Ano95y]. achieving [Ano96e, Ano96e, Ano96f]. Achille [Ano96-87]. acid [BD96]. acids [KM96]. ACM [Ano95v]. acoustic [Alk94, Ano96-60]. acquisition [Ano97v]. action [Ano99-117, Ber96, Ano96c]. actions [Ano99-181]. active [KKB97]. activities [Ano99-89]. Acton [Ano96-137]. actor [Ano97-32]. actor-based [Ano97-32]. actors [MR99]. actuated [Alf95, Ano96-171]. acyclic [Nab98]. Ada [Ano97-34, RSC93, Ano98-35]. Adam [Ano95-59]. Adami [Ano99-29, Ano99-30]. Adamowicz [Ano97-87]. Adams [Ano95-68, Ano98-61, Ano99l, DBM95, MBV91]. Adams-type [MBV91]. Adamson [Ano97-33]. Adaptation [McK96]. adapted [MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b]. Adapting [Noa91]. adaption [Gol94, SAS92]. Adaptive [De 99, HTT92, Kar96a, KH92, AAII92, Ano95-36, Ano97-53, Ano99-90, CS96, Cas91, CL99b, EM94, Eso99b, FG93, KU91, Mas92, SVS98, WT90a, Ano99n]. add [Ano97-90]. add-on [Ano97-90]. Addison [Ano95-72, Ano95e, Ano95-35, Ano95-34, Ano95f, Ano95-79, Ano95-42, Ano95-50, Ano95-48, Ano95-56, Ano95-57, Ano95-61, Ano95-64, Ano96-151, Ano96-48, Ano96-86, Ano96-119, Ano97-107]. Addison-Wesley [Ano95-72, Ano95v, Ano95e, Ano95-35, Ano95-34, Ano95f, Ano95-79, Ano95-42, Ano95-50, Ano95-48, Ano95-56, Ano95-57, Ano95-61, Ano95-64, Ano96-151, Ano96-48, Ano96-86, Ano96-119, Ano97-107]. additive [CP98b, RS92]. addressing [BP94a]. ADI [Ano97-141]. adic [RR97a]. adjacency [CF91b]. adjoint [Che99b, Coo98, Il’97, Kre91, Tse95, ZQ93b]. administration [Ano99-176, Ano98-51]. administrators [Ano97-169]. admixture [Dor98a, Dor98b]. Adomian [AC94, AR92b, Eso99c, Gab94b, GGC97, VR95a, VR95b]. Adrian [Ano95-46]. adult [KL96b]. Advanced [Ano97y, Ano99-35, Ano99-36, Ano99-169, Ano95f, Ano96-165, Ano98-76, ZL98b, Ano96g, Ano97e, Ano98g, Ano99o]. Advances [EM94, Ano99-74, Bir98, Ano98-132, Ano95h, Ano96i, Ano96h, Ano96j, Ano97g, Ano97f, Ano98h, Ano99r, Ano99s, Ano99t, Ano99q, Ano99p]. advection [GN99, NN91]. advection-diffusion [GN99, NN91]. AEI [Ano97-139, Ano97-161]. Æleen [Ano98-51]. aerodynamic [Bak92, PC96a]. aerodynamics [WF90, WF91]. aerosol [MS99]. aesthetic [Mur93b, Mur91a]. aesthetics [Mur93a]. affected [AV98]. Affine [DM95]. affinity [Ano99-162]. Afsarmanesh [Ano99-128]. Agarwal [Ano96-33, Ano97e, Ano99-84, Ano99-85]. age [Ano96-52, Bar91b, KP95b, Kos90, SKI92, SOI92, Ano96-45]. age-dependent [Kos90]. age-structured [Bar91b, KP95b]. agency [Ano96-40]. Agent [Ano99-128, Ano96-135, Ano98-50]. agents [Ano97-102, Ano99-88]. aggregation [MM96b]. aggregation/disaggregation [MM96b]. Agre [Ano96-40]. Ahamed [Ano97-36]. ahead [EG98]. Ahmed [Ano97-163]. AHP [Nis97b]. AI [Ano96-128]. aided [Cho95, Lu96, YL93a, Ano95t, Ano96-41]. AIE [Ano97-37]. Air [NWL91, Gil93, RA92a]. aircraft [II93]. Airfoil [CPC93, WF90, WF91]. airfoils [RR90]. Akira [Ano98-115]. Al [Ano95-59, Ano96-163]. Al-kindi [Ano96-163]. Alan [Ano93p, Ano96y, Ano97-49, Ano97-93, Ano97-132, Ano98-35, Ano98-97, Ano98-141]. Alberto [Ano97-118]. Albitz [Ano99-221]. Alexander [Ano97-145, Ano97-152, Ano98-29, Ano98-37, Ano98-36, Ano98-154, Ano99-234]. Alexandros [Ano98-161]. Alf [Ano96-124]. Alfred [Ano98-107, Ano99n]. algebra [Ano95u, Ano96-102, Ano96-120, Ano97-140, Ano97-67, Ano99-45, Ano99-77, Ano99-129, Ano99-130, Che91, Cor95, GT92, HP96, Lei90, MRWM94, Pag90, Pat98, TG96, Ano95-44, Ano95-49, Ano96-101]. Algebraic [Ano98-66, Ano99-198, PYS97, Ano96-79, Ano98y, Ano98-153, Ano99d, Ano99-245, Bal96a, BG95, Bur92, CC97, Car94, FLW97, Fli95, Has91, HG98, KT92, KCB90, MT94a, Mär96, MPR98, NR96b, SAS92, Sol97, SH94, Wac91b, WT93, WW93, Ano96k, Ano98i, Ano99u]. algebraic-code [SH94]. algebraically [BS92, Rhe97]. algebras [Ano95-43, Ano98w, Ano98-92, Ano98-121, HJ91, PM95b, Ano96d, Ano97-140, Ano98-89]. Algorithm [Din99a, JY93, Pal97, RG92, pri94, AGM99, Ano93z, Ano99-245, AS95, AS98, AK90b, BV99, BH91, Bog99, BG95, Cab97, CDH91, CCCY94, CHCC95, CLD97, Chr90, CLY91, CHW99, DJM98, Din97a, DL99, EG97, EG98, EB91, ER98, FLW97, FG93, FWI99, Gal91a, Gal93a, Gal93b, Gal93c, Gla91b, Gla96a, GB96, GMR90, HG98, HLS91, HH93, HGG99, JS95, JCL93, JL94, JLDL92, KG98, KKS98, Koç91, Koç92, KO93, KHI94, LC97a, LLP94, Lew91b, Li98c, LL91b, Lin94, LC94, LOP95, Lin96, LS94, LC97b, MZ97, MV95a, MM96a, May99, MM95b, MJK92, MG90c, MZ95, MM95d, NB95, PN98, Pic91, RR97b, Sad96a, SMT98, Sch96, Sha90c, SWF95, SCBFC93, SCB^{+}93, SCBFC98, SY96, Sto92, SBD99, Sym91, TG95, TWIF98, Tor90, Trz92, TTC96, Tsa90, WSDT94, WF96, WT90b, Xue94, Zha93a]. algorithm [Zho90, ZZT99, dJF91, Ano99-226]. algorithmic [Ano99-70, Ano99-189, AR99b, KIAM98, Ano98j, Ano98k]. Algorithms [Ano98-33, Ano98-43, Ano98-108, Ano98-116, Ano99-209, CB93b, DT93, GS91, LKE97, MS99, ARM91, AW94, Ano96-112, Ano97-33, Ano97-34, Ano97-111, Ano97-112, Ano97-115, Ano98-55, Ano98-64, Ano98-68, Ano98-78, Ano98-124, Ano99-74, Ano99-148, Ano99-147, Ano99-166, AC93b, Bai96b, BM90a, BS95, BS96b, Bog98, Bou93, CY96, CF91c, CF94, EA90, Eso91, FL91, FT97b, FLW99, Gal99b, GS99, HTT92, Hua98a, JS97a, JS98, KK97b, KH92, Liu98, LO93, MV95b, MNB93, NNB97, Nel96, Oll90, Pan96, Pan98c, Par91a, QA96, Rhe96, Rhe97, Sar90, Sir97, Sir99, SMTS97, TCS93, Yan93b, vdV95, Ano96n, Ano96-129, Ano97i, Ano97-118, Ano98-91, Ano99-58, Ano99-133, Ano99-134, Ano99v, Ano99w, Ano99x]. Alice [Ano96-140, Ano97-64]. Alistair [Ano98i]. Alle [Ano99-229]. Alle-Jan [Ano99-229]. Allen [Ano96-75, Ano98-114, Ano99-77]. Allenby [Ano95-44]. allergen [GPCSL90]. allergen-triggered [GPCSL90]. allocation [CCK97, CK97a, KM92, MCG98c, MN91, PS93b]. allowing [Pan98a]. alloys [WSNS98]. almost [BG90, Fau94, Kar96a, KS91, Sol90]. ALPA [AA93]. alpha [Gal93a, HSH90a, HSH90b]. Alshawi [Ano93u]. Alternating [MG95, CS90, LW91, Wac94]. alternating-direction-implicit [Wac94]. alternative [Bec96, Cha94, Coo98, DO92, LLR^{+}90]. Altmann [Ano96-32]. Amadio [Ano99-68]. Amateur [Rea97]. America [Ano95-67, Ano96-29, Ano97-68, Ano97-79, Ano99e, Ano99-73, Ano96-142]. American [Ano97-161, Ano98-145, Ano96l, Ano97h]. amicable [Ano96-170, Ano97-119, YJ94, Yan94]. amino [KM96, BD96]. among [NB98]. AMS [Ano96-111]. AMS-LaTEX [Ano96-111]. Amsterdam [Ano93h, Ano95-65, Ano96-79, Ano96-82, Ano96-81, Ano96-80, Ano96-83, Ano97-64, Ano97-65, Ano99-113]. Amsterdam/Cambridge [Ano97-64]. Ana [Ano99-63, Ano99-64]. Analog [LURL98, Ano98l]. analogue [KS96, LCC99]. analogy [Hen97, RV94]. analysing [Bou93, DK97a]. Analysis [Ano99s, BKN97, Boy96, CTG96, CCY93, Coh94, DNJS98, GHM96, Koç95, Oli98, PBG90, PPC99, Rac99, RSC93, TNTR91, WF91, Abe90, AD93, AA93, AO93b, APV97, Ana95b, Ano96-41, Ano96-83, Ano97w, Ano97-65, Ano97-66, Ano97-91, Ano97-105, Ano98-57, Ano98-67, Ano99-71, Ano99-115, Ano99-116, Ano99-113, Ano00, ASW98, AF90, ADFP93, BBO99, BSS98, BR95b, BKK99, Ben99, BIZ97, BRGJ95, CMMR92, CJ90, CP95b, CCH97, Che91, CL99b, CHW99, Edw92, Ege91, EDO93, FT97a, GRZ96, Gan92, GN99, Gla94a, Gla95a, Gla97a, Gla98, GWL96, HM96, II93, IK95, KT90, Kat92, Kel94, Kla98, Kou89, Kou90b, KPS96, KLMP97, Kwo92, Li98c, LS94, LW99b, MT94b, MP91, MR91, MPS92, MB94b, MG90a, MG90b, NO93, NSS92, Pad92b, PC96a, PC94a, Pet92, RT96, SJB97]. analysis [Sen92, Sha90d, Sni90, SCBGN94, Sza90a, TSLW98, TM91, TG96, TTC96, Tsi91, Wei99, Zha97, ZZ99b, Zhe91, ZK99, vdVVdB99a, Ano90x, Ano95-33, Ano96-131, Ano96-143, Ano98-78, Ano98-125, Ano99z, Ano99-27, Ano99-121]. Analytic [JG95, JD97, NFJ96, Sha96b, SLC97, Zad89, Zad92, AIS95b, Ano95-27, BLP95, CFM91, CIS98, Ioa91, JG96, NPJ95, PM91, PT96, RS96a, RS99a, WF91]. Analytic-numerical [JD97, NPJ95]. Analytical [DGV94, KU99, KESH90, CP95b, KHI94, MBMS92, WF90, Ano97l]. Ananda [Ano99-129, Ano99-130]. Anaphora [Ano95-30]. Anastassiou [Ano98-161]. Anastosios [Ano99-105, Ano99-103, Ano99-102, Ano99-104]. Anatoly [Ano98-32]. and/or [VHH95]. Anderson [Ano96p, Ano97-30]. André [Ano97-71, Ano98-142]. Andreae [Ano99-32]. Andreas [Ano99-69]. Andrew [Ano90p, Ano95-40, Ano97-98, Ano97-99, Ano97-97, Ano98-106, Ano99-160, Ano99-161, Ano99-230]. Andries [Ano95-57]. Andrzej [Ano98-164]. Andy [Ano97-124, Ano98-35, Ano98-40]. Angelika [Ano98-91]. Angeline [Ano97f, Ano99p]. angiograms [GCLT98]. angle [BS99b, MK98, Ram91c]. anharmonic [Gur91]. Anil [Ano97-86]. animals [Ano95-36, Ano97-53, Ano99-90]. Animating [Ano97m]. animation [Rot96]. animats [Ano95-36, Ano97-53, Ano99-90]. Anne [Ano98-136]. annealing [BT94, Fab97, Fox94, Pic91, RI94]. annihilating [BKK97]. Annihilation [KL95, Han92, KLMY97]. annotated [Ano99-117, Ano98m]. Announcement [Ano90d]. annoyances [Ano97-168, Ano98-52, Ano98-120]. annual [Ano96-67, Ano97-57]. anomaly [Coo97, PCR90]. Anselmet [Ano98-84]. ANSI [Ano98-61]. answer [LL90b]. answering [GGB92]. Anthony [Ano95-82, Ano99-81]. Anthropology [Ano99-140]. antibody [CJ90]. Anticausal [HI94]. antifolate [SHT^{+}90]. antiproton [KLMY97]. antiproton-nucleon [KLMY97]. antisymmetry [Ano95-70]. Antoniou [Ano97-109]. Anupam [Ano97-163]. AOR [Bai96e, Bai96f, Wan99]. Apache [Ano99-238, Ano99y]. apertures [MK91]. Apostolico [Ano97-118]. appearance [Byk97]. Appel [Ano97-98, Ano97-99, Ano97-97, Ano98-106]. Appl [Li10, Moo93b]. appliances [Ano99-216, Ano99-217]. Applic [Ado91a, BM90c, Kou90b, MG90b, Sal92b]. Application [Alt90a, BFF^{+}97, Chu99, GPZ93, GOWR^{+}93, Gla97b, Ioa91, IA94, Kou98, KLMY97, PB95, Sán98, SFT99, Shu93a, TCJG91, WT90a, AD96, Ano96f, Ano97-68, BGZ92, BJH90, CJS90, Erm92, EGHS98, FL96, Fli95, Gal93a, Gal93b, Gla93d, Gla94a, HGK93, HP94c, Ike91, Kis91, Kra90, KNS99, LD90, MCG98a, PW95, Pom92, RSS94a, RZ91, Sza90b, Tan93a, WCHL96, ZC91]. Applications [SCG^{+}97, Yav93, Zad92, ZQ93b, AD98, AR92e, Ano92a, Ano96-91, Ano97-47, Ano98-110, Ano98-116, Ano98-140, Ano99-88, Ano99-151, Ano99-171, BW92, CC92, DW97, DDP99, Dra99, EHK92, Eso91, Gir92, Gyö90a, Har90, IZ97, Kan90a, Kan90b, KPS96, Lai97, LC97a, Lew94, LL91b, LL90c, MM94, Mil98, NP96, Nis97b, O’R95, OR97b, Pan92a, PT91, RT96, RFRM99, Róz99, Sni91a, SA95, Sza90a, Tan94, TL94, TM91, Tre95b, TZ93, Ver99a, WSNS98, WU99a, YL93a, Ano98-43, Ano99-93, Ano96-43, Ano96-47, Ano96-123, Ano96-136, Ano97-48, Ano97-66, Ano97-86, Ano97-106, Ano97-116, Ano97-126, Ano97-163, Ano98z, Ano98-33, Ano98-65, Ano98-66, Ano98-68, Ano98-92, Ano98-108, Ano98-117, Ano98-121, Ano98-124, Ano98-150, Ano98-156, Ano99v, Ano99w, Ano99-74, Ano99-103, Ano99-104, Ano99-150, Ano99-198, Ano99-234, Ano98n, Ano99z]. Applications [Ano99-27]. applications-Binary [Nis97b]. applications. [Ano97j]. Applied [Ano90a, KPB97, AC94, AJ95, CMMR92, ES97, Gla94f, Gol94, HPH^{+}94, KHI94, Mat99, VDLM95, Woo91, Ano98o, Ano99-28]. Applying [CL99b, Fog94, tC95]. Approach [Eke99, AO98b, AO99b, AAL99, Ano93k, Ano95t, Ano95c, Ano96b, Ano96c, Ano96m, Ano96-161, Ano97-32, Ano97-34, Ano99-105, Ano99-103, Ano99-102, Ano99-104, Ano99-171, BMR96, BV95, BDT95, Bec96, BMD92, Cha98, Cha94, Cha99a, CM93, CT90, CCY95a, CCH97, Che90b, CHW99, DG91, DJV96, EDOS93, GF97, Hau92, HM91, Hsu91, JK96, KU91, Köh93b, KK91, Kou89, Kou90b, LCP98, LZ92, Lee96, LTC98, LDD94, LURL98, MP99, Mit92a, NPJ95, Oll90, Özy96, Pad92a, PSLT93, QA98, RDK92, RM98a, Ram90b, RA92b, SAS92, SGS93, Sen92, Shu93b, Sti93a, UK99, Ung90, Ung91, dCVC97, VPK99, Wu90b, YZM94, YZM95, Ano90o, Ano95-57, Ano96-150, Ano96-110, Ano97-46, Ano97-119, Ano98u, Ano98-151, Ano99-70, Ano99-98, Ano99-132, Ano99-170, Ano99-189]. approaches [Ano96v, Ano96-34, Ano97v, KK94, PLH95]. approach^{T M} [Ano96-159]. approximants [Kov95, PB95]. Approximate [AJ95, JI94, RR90, RT94, Riz95, RZ95, Sar94, TAD91, AdMD99, Ame97a, CN94, Dat90, Dat93, EDNO92, GGS97, Gla93a, Gla94a, Gla94f, Gla95c, HP94c, JCL93, KSM94, QA98, Rei98, Rei99b, Sar90, SS91a, SS94, Sha96b, Shk97, Tan92]. Approximating [BBKN97, Hua98c, BBI97, Gre93b, JGB95, LLP94, Pan96, RW98]. Approximation [Bar91b, Bro94, HS98, KM97b, Nis95, O’R98a, OR97b, Pou90a, TD99, AV98, AZ96, Ano98-59, Ano98-91, BM91, BL95, BF97, BA96, BA98, BH91, BS96d, CMP95, Cha96, Cha99b, CW95, CG94, DNJS98, EDO93, EGEHSEK95, EGHS98, FO91, GP90, Gyö90a, Hu95, JPC99, JP93, KK97a, Kan90a, Kan90b, Kan92a, KP95a, KO97, KO99, KS91, LW96, Li98c, LW99a, LR96b, LZ98, MC93, MO95, Mei97, Mol93, NO93, Pav97, Rab92, RS95, Sad96b, Sal91, Sal92b, SH95a, Shi98, Sug94, Tch98, Wu90a, Zal95]. approximations [Ano99-123, Ano99-124, BR94, Bro98, DG99, Fab95, Fli94b, GHS91, Han92, Kan90a, KP98, KP99a, KP96, KM97c, Wan91, ZB95, Zha95, Zha96]. apps [Ano98-149]. April [Ano98-101]. Apt [Ano96-114]. Arató [Rod94f]. Arbib [Ano98-113]. Arbitrary [MG91, Mar94, AS95, BLR96, Din97b, Heg91a, Heg91b, RR90, SMT98]. Arcady [Ano95-36, Ano96-34, Ano97-53, Ano99-90]. arch [Kel91]. architecture [Ano96-150, DC94, KCD^{+}97]. architectures [Ano97-115, CDLW95, Ano96-86, Ano98g]. arcs [Say99, YLC96, YL99]. area [dCVC97]. arguments [AG99b, CG94, KAD99, QA98, VPK99, WA98]. Argyris [Ano98-57]. Ariel [Ano95d]. Ariew [Ano96-50]. arising [AH92, BCGP99, Boy96, JR94, KOT99, MC99, MRWM94, SS91b, Tan96, VMA99, XL97]. Aristides [Ano98-121]. arithmetic [Gla94a, Gla94b, Gla94d, Gla95a, Lia99b, Nam96b]. arithmetical [Mor93]. arithmetico [Sib98]. arithmetico-geometric [Sib98]. Arjeh [Ano95u]. arm [Yos93]. ARMA [Sou90]. Armstrong [DT93]. Arney [Ano97-68]. Arnold [Ano95-44, Ano97-50]. Arnoldi [RM98a]. Arpad [Ano98-126]. arrangement [Köh94]. arrangements [Ano98-136]. array [Lin91a]. arrays [MMMS97, MMMS96]. Arrival [Alf95, AK93]. arrivals [NJ92]. Arrow [Ano97n]. arrowhead [Gal99a]. Arrowsmith [Ano96-56]. art [Ano90w, Bes96]. arterial [MPM93]. arteriovenous [PCR90]. Arthur [Ano96-115, Ano99-185]. article [Lei96]. articles [Jay92, Los92]. Artificial [KL96a, Ano93f, Ano95-47, Ano96-61, Ano96-74, Ano96-113, Ano96-145, Ano97-44, Ano97-63, Ano99-29, Ano99-30, BV95, Gir92, RA92a, SS90, Ano97-153, Ano90o, Ano95k, Ano95m, Ano95l, Ano96s, Ano99-29, Ano99-30]. artistic [ALL^{+}96]. artistic-scientific [ALL^{+}96]. arts [Las96, Ano95-75, Ano97s]. Arturo [Ano99-215]. ary [CK97a, KH92]. Ascher [Ano96-65]. Ashwin [Ano96-78]. ASI [Ano98-93]. Askar [Ano99-204]. ASP [Ano99-61, Ano99-31]. aspect [Ano96-140]. aspects [Ano95-81, Bur92, Eso91, Gyö90b, MPM93, Par91b, Tem93]. assembly [BD96, DB96, Fil93, FYS99, Pou90b, YF92a]. assessment [Ano96-127, KYO92]. asset [Nai92]. assets [Nak99]. Assigning [Ano97-142]. assignment [CHCC95, Hwa99, LWL93]. assimilation [DF98]. assisted [Mro96]. associated [AB93, GG92, GHS91, LD95, LOD96, MW99, Pan95, RS96a, RS99a, RSS94b, TCS93, Wen99]. Associates [Ano95-76, Ano95-52, Ano95-46, Ano95-80, Ano95-58, Ano95-63, Ano96w, Ano96u, Ano96-42, Ano96-28, Ano96-49, Ano96-51, Ano96-70, Ano96-88, Ano96-92, Ano96-133, Ano96-166, Ano97-101, Ano97-102]. Association [Ano95-67, Ano96-29, Ano97-68, Ano97-79, Ano99e, Ano99-73]. Associative [KKK91, KNU99, MKS^{+}91, Ano99-32]. associativity [Lev92]. assumptions [Arg99]. assuring [BQN97]. astronomy [Ano97h]. asymmetric [CH99, Luo93a, Luo93c]. Asymptotic [Ahm90, Bal98, CP98a, CL96a, DH98b, Dem94a, Dem94b, Den91a, FLR92, FK97, FBL99, Hal90, KT96, Kel94, Kla98, Kou90a, LC98a, Mic98, MP91, PZ99b, PC94b, PXHG99, SWY96a, Tem93, TP98, Tre94, Tre95a, Yu98, ZKÖ98, Zha98, Ano98o, Bax92, CCVV90, GB90, Ise94, Med97, Mic94, MK93, SWY97, Tra97, Tre98]. Asymptotically [JM91, Hua99, Tre95a]. asymptotics [LFP97, PT91, Ric95]. asynchronous [Mav93]. Athanasios [Ano98-108, Ano98-123]. Atlas [Ano97o, Ano98p]. ATM [HY98]. atmosphere [Mat99]. atom [DR96, KKB97, TNS98]. ATOMFT [CC94b]. Atomic [Mav93, Ber91]. Atoms [Ano95n]. attachment [GK91]. attenuation [WSS92]. attitudes [Ano96-139]. attraction [Gre93b, Gre99, XLPG99]. attractive [VL92]. attractivity [Goh94, ZYW98, ZZ99e]. Attractors [Rom98, BWK97, RZ96, ST97]. attributes [HGWL96, Ita93]. audio [Ano98-109]. Auerbach [Ano97-49]. August [Ano98-111, Ano99-128, Ano99-212]. Augustin [Ano97-157]. Augusto [Ano98-140]. Aumann [Ano95-62]. Austin [Ano94a]. Austria [Ano99-120]. authenticate [CH93]. authentication [WCY94, YL93b]. authentication-combined [WCY94]. authentications [Lia95]. Author [Ano95a, Ano97a, Ano97b, Ano98a, Ano98b, Ano99a, Ano90-29, Ano90-30, Ano91l, Ano92h, Ano93-29, Ano95-87, Ano96-172, Ano96-173, Ano97-141, Ano99-248, Ano90e]. authoring [Ano98-47]. automata [Ano95t, Ano97-96, Ano99-166, CCD99, GB90, KCD^{+}97, KCDV99, MG93, Ano98q]. automata-theoretic [Ano95t]. Automated [MN91, PM95a, SFS95, YTLK94, Cha94, KT90, Pou90b, Wos95b]. automath [Ano95-65]. Automatic [Ano95-38, Abe90, Ano99-99, Ano99-100, LP99, MMZ98]. automating [BGO95]. automation [Ano98-164, Ano99-128]. automaton [BFF^{+}97]. autonomous [BBKN97, KS99]. autoregressive [Faz94a, Kor90a, Kor91]. availability [SKI92]. Average [Li98c, GHWL96, Kur91, KYNY99, Ohn92, Res91, She98, TTC96]. Average-case [Li98c]. averaging [CGH92, Gla94a, Gla94b, Gla94d, Gla95a, Gla95b, Gla96a, Gla97a, Gla98, Gur91]. avoid [Paz90]. Avoidance [ZIKM93, IM93]. Avoiding [Zed90]. Avresky [Ano98-56]. AweSim [Ano97-132]. AWT [Ano97-74]. AX [Maz90]. axially [Gla93d]. Axiomatic [AA93, Nam99a, Ano99-103, Ano99-102, Ano99-104, Ano96t]. Axiomatic-like [AA93]. axioms [Kun95]. axisymmetric [Pas92]. Axtell [Ano97-63]. Ayupov [Ano98-89]. Azer [Ano98-132].

B [Ano90o, Ano95-73, Ano95-44, Ano95-48, Ano95-62, Ano96-35, Ano96-47, Ano96-53, Ano96-67, Ano96-78, Ano96-77, Ano96-107, Ano96-120, Ano96-121, Ano96-140, Ano97q, Ano97r, Ano97-36, Ano97-57, Ano97-59, Ano97-132, Ano98-65, Ano98-92, Ano98-149, Ano98-134, Ano98-162, Ano98-62, Ano99z, Ano99-27, Ano99p, Ano99-83, Ano99-99, Ano99-100, Ano99-117, Ano99-162, Ano99-203, Bak90, Ano97-142, EL92]. b-book [Ano97-142]. B-spline-Bézier [EL92]. Baader [Ano98-147]. Babak [Ano99-131]. Babette [Ano96-76]. background [WU94]. backup [SKI92]. backward [BTT90, MBV91, VBM90]. bacteriophage [CCH^{+}90]. Bagaria [Ano98-140]. Bagchi [KM92]. Bahder [Ano95-48]. Bailey [Ano97-69]. Bakker [Ano96-46]. balance [GW99]. balanced [Ano99-128, PN99]. balances [Ye99]. balancing [JL94]. Baldi [Ano98u]. Baldwin [Ano97-54]. Ballard [Ano97k]. ballistics [SV97]. BAM [Chu99]. Banach [Arg98, Din96, Din97a, Din97b, Hua98c, HB99, LG99, O’R95, O’R99, Yao92]. band [BLR90, Kal90, MMMS96, Nyí99]. bandit [Yos93]. Bandits [Ano96u]. bang [YFF92b, YFF92a, YFF93]. bang-bang [YFF92b, YFF92a, YFF93]. bank [MPS93]. bank-to-turn [MPS93]. banks [KKM96]. Banyaga [Ano97-157]. bar [Ano96-142, Ara99a]. Barbara [Ano95-39]. Barbosa [Ano97i]. Barbu [Ano99-183, Ano99-184]. Barcelona [Ano98-140]. Barlow [Ano99-47, Ano99-48]. Barnes [Ano96-143, HS95]. Barnett [Ano99-47, Ano99-48]. Barnsley [Ano90r]. Barnwell [Ano96-147]. Baron [Ano96-127]. barotropic [Gla94f, dCVC97]. Barrett [Ano96u, Ano96-27]. barrier [Pan90b, TNS98]. Barriers [Ano98r]. Barros [Ano99-63, Ano99-64]. Barry [Ano96-89, Ano96-143, Ano97-41]. Barth [Ano97-85]. Barwise [Ano97-162]. based [AE90, Ahm95, Ano90o, Ano97-32, Ano98-109, Arg99, BMR96, BMB94, BIZ97, BPLS96, CL94c, CSOL96, CT96, DBM92, DBM94, DBM95, DC94, De 99, Fli95, FGN90, FK98a, GPK99, Gup94, HM96, HAO91, Hsu96, HL90, HLS91, HWL94, HGG99, JW93, LCSN98, Lyt92b, MG93, MBV91, MR92, MS99, MM95d, NP96, Oza94, Pad92a, Pan95, PW97, Pet92, RSC93, Ram90a, SM98, SHK99, SH96, Sok91, SI94, SS96, SS98b, Tho92, VBM90, VL95, WY93, Ano96-127, Ano97r, Ano97-85, Ano98-37]. bases [Cie95, IA94, KV95, Wac91b]. Basic [Ano97-98, Ano97-99, Ano97-97, Ano96-112, Ano98-53, KS96, RSS94b, Yao93, Ano97p, Ano98s, Ano99-33]. basis [BN92, CF92, Gir92, HSH90b, Il’97, KO97, Mye92, Pan98a, RK98, RHRZ90, Sal92a, vNvN90a, vNvN90b]. basis-deficiency-allowing [Pan98a]. basophil [GPCSL90]. Basu [Ano97l]. BASYS’98 [Ano99-128]. BASYS’98-Third [Ano99-128]. batch [SMU98]. Batya [Ano98-70]. Baumann [Ano97-92]. Bayes [AJ95]. Bayesian [BBM95]. Be [Ano98t, Ano99-35, Ano99-36, Ano97-141, Ano97-79, Ano99-192, HL99, Tan93b, Ano98t, Ano96-167]. Be&TM [Ano99-36]. beam [MDVK97, RS93, Tad94]. Beans [Ano97-38]. beauty [Ano99j]. Beck [Ano96a]. becoming [BGM97]. Begehr [Ano99-182]. beginners [Ano95f]. Beginning [Ano99-34]. Beguelin [Ano95-59]. Behavior [PC94b, Ano95-36, Bal98, CL96a, CCVV90, CL96b, DH98b, GB96, KP99b, KC91, Med97, PXHG99, She94, TP98, Tra97, Tre94, Tre95a, ZKÖ98, Ano97-53, Ano97-43, Ano98-114, Ano99-90]. behaviour [Bin91, FLR95, Ise94, KP96, Nas92, RSP94, Ano96-56, Ano96-56]. Beidler [Ano97-34]. Beijing [Ano99-137]. being [Ano99-215, BGM97]. Béla [Ano99-156, Ano99-157]. Belew [Ano99-29, Ano99-30]. Belgium [Ano98-155]. Bellman [Ano90b, Bel94b, Eso90, KNUX99]. Beloslav [Ano98-74]. Ben [Ano95p, Ano99y]. Bendat [Ano98-117]. Benders [ZC91]. Beng [Ano98-31]. Benjamin [Ano95-60]. Benoit [Ano98-62, Ano99-162]. Benson [Ano95-73]. Benthem [Ano97-64]. Berezansky [Ano99-115, Ano99-116]. Berg [Ano98-33]. Bergadano [Ano96-90]. Bergé [Ano98l]. Bergeron [Ano97-40]. berghei [BJH90]. Berghmans [Ano98-155]. Bergin [Ano99-55, Ano99-56]. Berkane [Ano99-117]. Berlin [Ano93o, Ano96-68, Ano97t, Ano97-46, Ano97-51, Ano97-152, Ano98-33, Ano98-91, Ano98-102, Ano98-62, Ano99v, Ano99w, Ano99-44, Ano99-43, Ano99-66, Ano99-67, Ano99-69, Ano99-70, Ano99-82, Ano99-123, Ano99-124, Ano99-132, Ano99-198, Ano99-208]. Berman [WHC90]. Bernd [Ano98-138]. Bernstein [AdlC95, GZ95, GlZ95, He95b]. Bertil [Ano96-164]. Bertrand [Mor93]. Bessel [BL96, DMC^{+}95, DTLM98, KAS99, LD93, LOD96, Mic94, ZB95, Zha95, Zha96]. Best [LW99a, Ano98-59, BL95, CW95, LW96].

Bestavros [Ano98-132]. bestselling [Ano99-221]. beta [Ano96-93, FF92, Nar92]. Beth [Ano96-151]. Be^{T M} [Ano99-35]. Betounes [Ano98-125]. Between [Ano96-105, Ano95s, Ano99-215, ASW98, BT98, BH90, DPD98, Gla96c, HSH90a, II93, IGM98, Kim92, KM93, LYL95, OPD97, PSD96, Ste90, ZFM96, vdVVdB99b]. Beverly [Ano97q]. Beyne [Ano98-155]. Beyond [Ano97z, Ano97-39, MNE92]. Bézier [AK98, He95b, KKM96, KM97b, Las94, Las95, PC94a, EL92]. Bhandari [Ano99-209]. bi [Kar96b]. bi-Hamiltonian [Kar96b]. bibliographies [Ano98m]. bibliography [LL90c]. BiCGSTAB [SFT99]. bicharacteristic [Hua95]. bidiagonal [APP96]. bifurcation [Abe90, CST97, Cho92, Cho95, KCB90]. Bifurcations [ACL93, CKM99, GSST97, CC98a, MPM93, ST97]. biharmonic [Har90, Ike91, Jek94]. bilateral [CS94a, JV96]. bilevel [AD93]. Bilinear [Ter91, AK90a, ARS98, CS94a, RZ91, XLG95]. Bill [Ano96-151, Ano96-88, Ano99-160, Ano99-161, Ano99-178, Ano99-179]. billiards [WAL97]. billion [TNS98]. Bimal [Ano98-142]. bimatrix [Bai96a, BJPT93]. Bin [RI94, MCG98a]. Bin-packing [RI94]. Binary [Nis97b, Pan93a, Ban93, Fje91c, KH92, LSH95, Leb98, Tel91, Zha93a]. BIND [Ano99-221]. binding [HSH90a, HSH90b, MCG98c, ZL99c]. Bingham [Sán98]. Bini [Ano96-129]. Binmore [Ano95-37, Ano98-63]. binomial [Fje91a]. biochemical [Hol90]. bioenergetics [RGBL96]. Biographical [Ano92b]. biography [Ano99-146]. Bioinformatics [Ano98u]. biological [Ano98-105, BD96, DB96, Sie90]. biology [Ano97-107, Boy96, KOT99]. Biorheological [MPM93]. Biorthogonal [IN97, KV94b, KV95]. biosynthesis [SHT^{+}90]. bipartite [Che99a, SY96, TYC93]. Biplanar [Bei97]. biquadratic [GF94a]. biquintic [WA94a]. Birgit [Ano98-102]. Birkhäuser [Ano96-111, Ano96-129]. birth [Ano99-239, PL99]. birthday [US91]. bisection [Bar94, Oza94, Woo91]. bistable [Mol93]. bit [PR90]. bit-complexity [PR90]. bits [Ano96-30]. Bjørner [Ano93h]. Bjørstad [Ano97-41]. Blachman [Ano93k]. Black [Ano96-69, Ano99-169, Ano99-91, Ano98v]. black-box [Ano99-169]. blades [CM93]. Blaquière [Ano94a]. Blaschke [Kai94]. Blay [Ano98-134]. blend [Fol92]. Bloch [MAM91]. Bloch-equation [MAM91]. Block [AAD98, AS96a, RG92, VA97, APP96, BG90, Cab97, CKL91, CTY95, Con98, EG98, Gal99a, JLDL92, Nyí99, Pap94b, Tor90, Xu97, Yan99a]. Block-cyclic [AAD98]. block-Lanczos [Yan99a]. block-tridiagonal [CTY95]. blocks [Dri97, NB98]. blood [GCLT98, MPS92]. Bloom [Ano97-81]. Bloor [Ano96-143]. Blum [Ano93r]. Blumberg [Ano99-90]. blunder [Ano96-57]. Board [Ano98-162, Ano93y, Ano90z, Ano90-27, Ano91j, Ano91k, Ano92e, Ano92f, Ano93x, Ano94b, Ano95-84, Ano95-85, Ano96-168, Ano96-169, Ano97-172, Ano97-173, Ano98-166, Ano98-167, Ano99-240, Ano99-241]. Bob [Ano97-131]. Boca [Ano95w, Ano95z, Ano95g, Ano98-43, Ano98-107, Ano99-108, Ano99-125, Ano99-126]. Boden [Ano95-28]. bodies [Ano99-137]. body [DKR97, Gla91b, GRC^{+}96, LQ95a, OHI91, YFF92a]. Bogoyavlenskii [GT97b]. Bohlman [Ano97-40]. Boissonnat [Ano98k]. Bollobás. [Ano99-156, Ano99-157]. Bologna [Ano96e]. Boltyanski [Ano99-101]. Boltzmann [BLP95, BCD98, RW98, Str98, TSLW98, Tch98]. Boltzmann-type [RW98]. bond [BGZ92]. bonded [DPD98]. bonding [Gre99]. bone [MR91, MBMS92]. bonus [Ano99-221]. Book [Ano90o, Ano90p, Ano90q, Ano90r, Ano90s, Ano90t, Ano90u, Ano90w, Ano90v, Ano90x, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92c, Ano92d, Ano93f, Ano93g, Ano93t, Ano93u, Ano93h, Ano93i, Ano93j, Ano93l, Ano93k, Ano93n, Ano93m, Ano93o, Ano93p, Ano93q, Ano93v, Ano93r, Ano93s, Ano93w, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano95b, Ano95k, Ano95-36, Ano95h, Ano95m, Ano95l, Ano95-70, Ano95n, Ano95o, Ano95r, Ano95s, Ano95d, Ano95q, Ano95p, Ano95t, Ano95-72, Ano95v, Ano95y, Ano95u, Ano95x, Ano95c, Ano95w, Ano95-71, Ano95z, Ano95-28, Ano95-27, Ano95-29, Ano95-74, Ano95-73, Ano95-30, Ano95e, Ano95i, Ano95-31, Ano95-75]. Book [Ano95-35, Ano95-34, Ano95-77, Ano95-32, Ano95-76, Ano95-33, Ano95-37, Ano95-38, Ano95-78, Ano95f, Ano95-39, Ano95-40, Ano95j, Ano95-79, Ano95-45, Ano95-41, Ano95-42, Ano95-44, Ano95-43, Ano95-52, Ano95-51, Ano95-47, Ano95-46, Ano95-80, Ano95-50, Ano95-48, Ano95-49, Ano95-53, Ano95-55, Ano95-56, Ano95-57, Ano95-58, Ano95-59, Ano95-60, Ano95-61, Ano95-63, Ano95-62, Ano95-65, Ano95-66, Ano95-64, Ano95-67, Ano95-81, Ano95-69, Ano95g, Ano95-82, Ano95-83, Ano95-54, Ano95-68, Ano96g, Ano96s, Ano96i, Ano96-150, Ano96e, Ano96h, Ano96k, Ano96f, Ano96t, Ano96j, Ano96l, Ano96w, Ano96y, Ano96u, Ano96x, Ano96v, Ano96-32, Ano96-37, Ano96-39, Ano96-27, Ano96-29, Ano96-42, Ano96-45, Ano96z, Ano96-44, Ano96-36, Ano96-151, Ano96-43, Ano96-33, Ano96-34, Ano96-31, Ano96-35, Ano96-28, Ano96-41, Ano96-153, Ano96-152]. Book [Ano96a, Ano96b, Ano96-38, Ano96-30, Ano96-40, Ano96-47, Ano96-46, Ano96-48, Ano96-50, Ano96-54, Ano96-55, Ano96-49, Ano96-53, Ano96-52, Ano96-51, Ano96c, Ano96-56, Ano96-69, Ano96-67, Ano96-65, Ano96-61, Ano96-57, Ano96-63, Ano96-66, Ano96-62, Ano96-60, Ano96-68, Ano96-71, Ano96-58, Ano96-72, Ano96-70, Ano96-59, Ano96-64, Ano96-74, Ano96-73, Ano96-75, Ano96-76, Ano96-78, Ano96-77, Ano96-86, Ano96-79, Ano96-87, Ano96-82, Ano96-81, Ano96-80, Ano96-88, Ano96-83, Ano96-85, Ano96p, Ano96-89, Ano96-91, Ano96q, Ano96-90, Ano96r, Ano96n, Ano96m, Ano96o, Ano96-92, Ano96-93, Ano96-95, Ano96-94, Ano96-96, Ano96-104, Ano96-100, Ano96-98, Ano96-99, Ano96-103, Ano96-154, Ano96-101, Ano96-102, Ano96-97, Ano96-119, Ano96-109, Ano96-114, Ano96-105, Ano96-116, Ano96-113, Ano96-111, Ano96-118, Ano96-106, Ano96-107, Ano96-112, Ano96-156, Ano96-108]. Book [Ano96-117, Ano96-120, Ano96-115, Ano96-110, Ano96-155, Ano96-122, Ano96-125, Ano96-124, Ano96-157, Ano96-123, Ano96-121, Ano96-126, Ano96-158, Ano96-134, Ano96-128, Ano96-129, Ano96-131, Ano96-136, Ano96-132, Ano96-159, Ano96-127, Ano96-133, Ano96-130, Ano96-138, Ano96-140, Ano96-137, Ano96-139, Ano96-148, Ano96-144, Ano96-141, Ano96-142, Ano96-161, Ano96-146, Ano96-147, Ano96-149, Ano96-143, Ano96-164, Ano96-163, Ano96-165, Ano96-162, Ano96d, Ano96-167, Ano96-84, Ano96-135, Ano96-145, Ano96-160, Ano96-166, Ano97c, Ano97g, Ano97-140, Ano97f, Ano97-141, Ano97-53, Ano97m, Ano97n, Ano97o, Ano97h, Ano97l, Ano97e, Ano97p, Ano97-142, Ano97q, Ano97r, Ano97-31, Ano97s, Ano97-28, Ano97w, Ano97-32, Ano97u, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97t, Ano97-30, Ano97v, Ano97-29, Ano97-40, Ano97-41, Ano97-33, Ano97-39, Ano97-34, Ano97-43, Ano97-37]. Book [Ano97-36, Ano97-38, Ano97-42, Ano97-35, Ano97-45, Ano97-44, Ano97-143, Ano97-46, Ano97-54, Ano97-48, Ano97-47, Ano97-51, Ano97-50, Ano97-49, Ano97-62, Ano97-57, Ano97-63, Ano97-61, Ano97-60, Ano97-55, Ano97-59, Ano97-58, Ano97-56, Ano97-144, Ano97-64, Ano97-65, Ano97-72, Ano97i, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97-68, Ano97k, Ano97-69, Ano97j, Ano97-66, Ano97-75, Ano97-77, Ano97-76, Ano97-74, Ano97-78, Ano97-145, Ano97-83, Ano97-81, Ano97-85, Ano97-84, Ano97-79, Ano97-82, Ano97-88, Ano97-86, Ano97-80, Ano97d, Ano97-146, Ano97-87, Ano97-149, Ano97-92, Ano97-95, Ano97-93, Ano97-148, Ano97-89, Ano97-96, Ano97-90, Ano97-147, Ano97-98, Ano97-99, Ano97-97, Ano97-100, Ano97-91, Ano97-94, Ano97-111, Ano97-101, Ano97-112, Ano97-102, Ano97-150, Ano97-106, Ano97-109, Ano97-103, Ano97-110, Ano97-108, Ano97-107, Ano97-105, Ano97-104, Ano97-113]. Book [Ano97-114, Ano97-116, Ano97-115, Ano97-117, Ano97-122, Ano97-123, Ano97-119, Ano97-120, Ano97-124, Ano97-151, Ano97-118, Ano97-121, Ano97-125, Ano97-152, Ano97-126, Ano97-137, Ano97-135, Ano97-156, Ano97-158, Ano97-153, Ano97-155, Ano97-128, Ano97-138, Ano97-154, Ano97-131, Ano97-136, Ano97-134, Ano97-132, Ano97-133, Ano97-127, Ano97-157, Ano97-160, Ano97-139, Ano97-161, Ano97-164, Ano97-162, Ano97-163, Ano97-165, Ano97-159, Ano97-166, Ano97-169, Ano97-167, Ano97-168, Ano97-52, Ano97-129, Ano97-130, Ano98g, Ano98m, Ano98h, Ano98q, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98k, Ano98p, Ano98s, Ano98u, Ano98w, Ano98x, Ano98c, Ano98r, Ano98v, Ano98-34, Ano98-29, Ano98-28, Ano98-38, Ano98z, Ano98y, Ano98-27, Ano98-37, Ano98-33, Ano98-31, Ano98-32, Ano98-35, Ano98-36, Ano98-30, Ano98t, Ano98-39, Ano98-44, Ano98-42, Ano98-41, Ano98-40]. Book [Ano98-46, Ano98-45, Ano98-43, Ano98-52, Ano98-54, Ano98-51, Ano98-48, Ano98-49, Ano98-148, Ano98-53, Ano98-50, Ano98-47, Ano98-60, Ano98-61, Ano98-58, Ano98-59, Ano98-56, Ano98-55, Ano98-57, Ano98-64, Ano98-65, Ano98-66, Ano98-63, Ano98-68, Ano98-67, Ano98-69, Ano98-70, Ano98-71, Ano98-73, Ano98-72, Ano98-78, Ano98-77, Ano98-82, Ano98-79, Ano98-76, Ano98-84, Ano98-74, Ano98-75, Ano98-81, Ano98-80, Ano98-83, Ano98-89, Ano98-88, Ano98-87, Ano98-86, Ano98-85, Ano98-96, Ano98-94, Ano98-93, Ano98f, Ano98-90, Ano98-95, Ano98-91, Ano98-92, Ano98-109, Ano98-102, Ano98-105, Ano98-99, Ano98-106, Ano98-97, Ano98-98, Ano98-108, Ano98-100, Ano98-101, Ano98-103, Ano98-110, Ano98-107, Ano98-104, Ano98-115, Ano98-113, Ano98-117, Ano98-149, Ano98-111, Ano98-118, Ano98-112, Ano98-114, Ano98-116, Ano98-121, Ano98-120, Ano98-119, Ano98-123, Ano98-124, Ano98-126, Ano98-127]. Book [Ano98-128, Ano98-129, Ano98-150, Ano98-122, Ano98-125, Ano98-130, Ano98-133, Ano98-137, Ano98-136, Ano98-131, Ano98-132, Ano98-135, Ano98-134, Ano98-151, Ano98-141, Ano98-142, Ano98-144, Ano98-138, Ano98-143, Ano98-139, Ano98-140, Ano98-153, Ano98-145, Ano98-157, Ano98-146, Ano98-155, Ano98-152, Ano98-147, Ano98-154, Ano98-156, Ano98-159, Ano98-158, Ano98-160, Ano98-161, Ano98-163, Ano98-162, Ano98-164, Ano98d, Ano98e, Ano98-62, Ano99-29, Ano99z, Ano99v, Ano99-30, Ano99-27, Ano99w, Ano99-90, Ano99o, Ano99r, Ano99s, Ano99n, Ano99t, Ano99y, Ano99x, Ano99-31, Ano99u, Ano99-28, Ano99-32, Ano99q, Ano99m, Ano99p, Ano99-35, Ano99-36, Ano99-33, Ano99-34, Ano99-37, Ano99-41, Ano99-47, Ano99-42, Ano99-48, Ano99-50, Ano99-213, Ano99-44, Ano99-51, Ano99-49, Ano99e, Ano99-39, Ano99-40, Ano99d, Ano99-43, Ano99-46, Ano99-45, Ano99-38, Ano99-63]. Book [Ano99-66, Ano99-53, Ano99-55, Ano99-64, Ano99-67, Ano99-54, Ano99-56, Ano99-69, Ano99-52, Ano99-65, Ano99-60, Ano99-57, Ano99-61, Ano99-62, Ano99-59, Ano99-58, Ano99-68, Ano99-75, Ano99-78, Ano99-76, Ano99-79, Ano99-73, Ano99-71, Ano99-74, Ano99-77, Ano99-72, Ano99-70, Ano99-84, Ano99-85, Ano99-86, Ano99-92, Ano99-83, Ano99-81, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-214, Ano99-89, Ano99-82, Ano99-87, Ano99-105, Ano99-103, Ano99-99, Ano99-106, Ano99-96, Ano99-109, Ano99-102, Ano99-104, Ano99-100, Ano99-107, Ano99-97, Ano99-110, Ano99-95, Ano99-98, Ano99-101, Ano99-108, Ano99-115, Ano99f, Ano99-111, Ano99-118, Ano99h, Ano99-116, Ano99g, Ano99-112, Ano99-119, Ano99i, Ano99-113, Ano99-117, Ano99-120, Ano99-114, Ano99-216, Ano99-123, Ano99-125, Ano99-129, Ano99-133, Ano99-217, Ano99-124, Ano99-126, Ano99-130, Ano99-134, Ano99-215, Ano99-135, Ano99-128]. Book [Ano99-121, Ano99-136, Ano99-137, Ano99-132, Ano99-127, Ano99-131, Ano99-122, Ano99-139, Ano99-138, Ano99-140, Ano99-145, Ano99-143, Ano99-141, Ano99-144, Ano99-142, Ano99-158, Ano99-160, Ano99-218, Ano99-156, Ano99-154, Ano99-159, Ano99-161, Ano99-219, Ano99-157, Ano99-155, Ano99-152, Ano99j, Ano99-151, Ano99-163, Ano99-149, Ano99-146, Ano99-162, Ano99-153, Ano99-164, Ano99-148, Ano99-150, Ano99-147, Ano99-167, Ano99-222, Ano99-168, Ano99-223, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-166, Ano99-220, Ano99-221, Ano99-170, Ano99-173, Ano99-174, Ano99-176, Ano99-177, Ano99-178, Ano99-179, Ano99-175, Ano99-183, Ano99-193, Ano99-190, Ano99-184, Ano99-194, Ano99-191, Ano99-180, Ano99-185, Ano99k, Ano99-181, Ano99-195, Ano99-187, Ano99-188, Ano99-224, Ano99-192, Ano99-186, Ano99-182, Ano99-189, Ano99-196, Ano99-199, Ano99-200, Ano99-197, Ano99-201, Ano99l, Ano99-198, Ano99-206, Ano99-207, Ano99-204]. Book [Ano99-225, Ano99-203, Ano99-209, Ano99-202, Ano99-205, Ano99-226, Ano99-208, Ano99-229, Ano99-212, Ano99-230, Ano99-210, Ano99-211, Ano99-231, Ano99-228, Ano99-233, Ano99-232, Ano99-227, Ano99-234, Ano99-235, Ano99-236, Ano99-91, Ano99-238, Ano99-237, Ano99-239, Ano99b, Ano99c, Rod93a, Rod94a, Rod94b, Rod94c, Rod94d, Rod94e, Ano95-79, Ano97-147, Ano99-213]. Books [Ano99-81, Ano97x, Ano97y, Ano97z, Ano99-221]. bookshelf [Ano99-221]. boolean [Ano98-133, BS94, Mit90, PM95b, VDLM95]. Bootle [Ano95-71]. border [Ano99-215]. bordered [KG95b]. Boris [Ano97-58]. Borko [Ano98-110]. Bornholdt [Ano99-166]. Bosković [Ano98-151]. Bossel [Ano95-51]. Boston [Ano93m, Ano93q, Ano96-111, Ano96-129, Ano97-36, Ano97-105, Ano97-104, Ano97-126, Ano97-163, Ano98g, Ano98l, Ano98x, Ano98z, Ano98-37, Ano98-31, Ano98-45, Ano98-50, Ano98-47, Ano98-56, Ano98-68, Ano98-69, Ano98-71, Ano98-72, Ano98-76, Ano98-109, Ano98-110, Ano98-133, Ano98-137, Ano98-132, Ano98-144, Ano98-138, Ano98-139, Ano98-156, Ano98-160, Ano98-161, Ano98-164, Ano99o, Ano99-50, Ano99-213, Ano99-52, Ano99-57, Ano99-58, Ano99-86, Ano99-88, Ano99-80, Ano99-94, Ano99-87, Ano99-99, Ano99-100, Ano99-117, Ano99-120, Ano99-128, Ano99-121, Ano99-131, Ano99-145, Ano99-142, Ano99-167, Ano99-168, Ano99-165, Ano99-169, Ano99-201, Ano99-209, Ano99-229, Ano99-212, Ano99-211, Ano99-227, Ano99-93]. both [Ram91a, Tan92]. Botsford [Ano95-74]. bottom [Ano97-63, HW90]. Bouchard [Ano96-161]. bound [BH91, Cha96, Cha99a, Cla98, EZ93, Eso91, Li96a, PvZ95, Pat98, Tse95, dJPvZ97]. bound-outer [BH91]. boundaries [Ano99-51]. Boundary [AO97, Bag90, Bal96a, BK91, BT98, Chi93, DELE99, LS96a, LDD94, Ram91a, XS90, AL93, AHY98, AO98b, ABW99, AO99f, AO98c, Ano90-28, Ano99-84, Ano99-85, Ano99-183, Ano99-184, Ano99-198, AP98b, BACS98, Bed95b, BR95c, Bog98, CC97, CKM99, CK91, CK94a, CK94b, DBM92, DBM94, Dav96, DK97a, Dem94a, Dem94b, EDNO92, EK94, EGLN95, Eti91, Far96a, GS93, Gás95, Gol90, Gol94, GBB91, GHM96, HM99, HK98, HP99, HM93b, HNS^{+}90, Hun95, HS98, Ike91, JPF96, Joh99, Kar96a, Kar94, KG98, KL95, Li97, Li98a, LN98, Li10, LM91, Lop93, MN91, MM93a, Mei93, Moi97, Mur92a, O’R98c, PA94, PR96, QA98, RSS94a, Ram90a, Sau95, SD95, SS94, SM90, Sod92, Som90, Str98, Sun91, SB99, SAS99, WCC99, Wan98a, Wan98b, Wan99, WJ99, WA96a]. boundary [WL96, Won98b, YP98b, YLC96, ZCC96]. Boundary-Value [DELE99, DBM92, DBM94, Moi97, QA98, SS94]. Bounded [EDO98, TR99, AI97, Hua97, LL99, LM96a, Pin98, Pro95]. boundedness [Li99a, SWY98]. Bounding [FZ93, RS96b, Ano96v]. Bounds [AP98b, HW93a, RSR98, AW91, AL93, BP90, DW97, EGLN95, FLR95, JP95b, JD97, JPC99, Käm94b, KT97, Mad92, NFJ96, RT94, RJ94, Roj99, RRS99, Sto99, WA94b, WA96c, Yan99a]. Bouwknegt [Ano97-140]. box [Ano99-169, FLM94].

Boyce [NGNM97]. Boyd [Ano90o]. Boyer [BKM95]. Boykoff [Ano96-127]. BR992 [AS93]. bracketing [GH90]. Bradd [Ano98i]. Bradford [Ano98-148]. Brady [Ano96-77]. Brain [Ano98-61, Ano90v, Ano97-43, Ano97-149, GCLT98, ZZ94, Ano93t, Ano95-41]. Brainerd [Ano98-61]. Brakhage [Ber93]. Brakhage-like [Ber93]. Branch [PL90, BH91, Cha99a, Eso91, RR97b]. branch-and-cut [RR97b]. branches [KCB90]. Branscomb [Ano98-82]. Bratislava [Ano98-74]. Brauer [Ano98w]. Breakdowns [BRZS97, SG90]. Breaking [TNS98, GT95, GT97b]. breathing [YTLK94]. Brendan [Ano99-106, Ano99-107]. Brent [Ano96w, Ano97v]. Brian [Ano90w, Ano96-54, Ano96-58, Ano96r, Ano96-148, Ano97x, Ano97y, Ano97z, Ano98-42, Ano99-176, Ano99-233]. Bricaud [Ano99-201]. bridge [PT91, Sem93]. Bridger [Ano97-139]. Brief [Zel92]. brightness [DGV94]. Brill [LD90]. Brisbane [Ano93p]. Britain [Ano99-225]. Broadcasting [Lia99a]. Brody [Ano96-100]. Bronstein [Ano97-138]. Brookhaven [Ano99-146]. Brooks [Ano95k]. Brookshire [Ano95-57]. Browder [WX97]. Brownian [Ara97, PT91]. Broy [Ano98-102]. Bruce [Ano93r, Ano95-73, Ano98-59, Ano98-105, Ano99-90, Ano99-62, Ano99-143]. Bruijn [SG95a]. Brunak [Ano98u]. Brusilovsky [Ano99n]. Brusselator [Ado95]. Bryan [Ano95-46]. bubble [Dor97, DIR97, Dor99]. bubbles [KG97]. Bucci [Ano96e]. Buchwald [Ano96-144]. Buehrer [LH97]. Buekenhout [Ano96-82]. Buffering [Ano98x]. build [Ano98-149]. building [Ano96-155, HJ91, Ano95o, Ano96w, Ano97q]. Buildings [Ano96-82]. bulb [AÉ96]. bulk [AD93]. bump [Ste90]. bundles [Ano99-234]. Burdea [Ano96-43]. Burgers [DG91, LDD94, vNvN90b]. Burleson [Ano99-125, Ano99-126]. Burns [Ano98-35]. Burr [Ahm94, Ahm95, ED91a, ED91b, Raq98]. bus [BLR96, Szt92]. business [Ano98-41, Ano98-40, Ano96y, Ano96x, Ano99-37]. Buttazzo [Ano98-68, Ano99-58]. Butzer [GGK97]. Buus [Ano95-46]. BV [Ano97-140]. BVP [RT99b]. BVP’s [PS96]. Byron [Ano97-148].

C [Ano90w, Ano93o, Ano95-61, Ano95-65, Ano95-81, Ano96-43, Ano96-56, Ano96-68, Ano96-73, Ano96-75, Ano96-87, Ano96o, Ano96-98, Ano96-109, Ano96-112, Ano96-121, Ano96-163, Ano97g, Ano97o, Ano97-27, Ano97-68, Ano97-95, Ano97-97, Ano97-116, Ano98h, Ano98q, Ano98-61, Ano98-64, Ano98-65, Ano98-68, Ano98-73, Ano98-101, Ano98-118, Ano98-127, Ano98-155, Ano99-41, Ano99-47, Ano99-42, Ano99-48, Ano99-55, Ano99-56, Ano99-58, Ano99-77, Ano99-109, Ano99-110, Ano99-133, Ano99-134, Ano99-176, Ano99-228, GS96a, CKL98, Wu90a, Ano97-73, Ano97-145, Ano97-111, Ano98-80, Ano99-147, Ano99-238, Ano96-27, Ano96z]. C. [Jay92, Zel92]. C.G. [DFS93]. CA [Ano95-76, Ano95-52, Ano95-47, Ano95-46, Ano95-80, Ano95-58, Ano95-63, Ano96h, Ano96w, Ano96u, Ano96-42, Ano96-44, Ano96-36, Ano96-28, Ano96-49, Ano96-53, Ano96-51, Ano96-70, Ano96-59, Ano96-88, Ano96-92, Ano96-93, Ano96-133, Ano96-135, Ano96-166, Ano97-141, Ano97n, Ano97q, Ano97r, Ano97-38, Ano97-35, Ano97-75, Ano97-77, Ano97-76, Ano97-74, Ano97-78, Ano97-148, Ano97-89, Ano97-101, Ano97-102, Ano97-103, Ano97-117, Ano97-123, Ano97-124, Ano97-131, Ano97-162, Ano97-166, Ano97-169, Ano97-167, Ano97-168, Ano97-129, Ano98t, Ano98-52, Ano98-54, Ano98-51, Ano98-70, Ano98-77, Ano98-79, Ano98-88, Ano98-86, Ano98-85, Ano98-99, Ano98-97, Ano98-98, Ano98-103, Ano98-112, Ano98-120, Ano98-129, Ano98-122, Ano98-130, Ano98-163, Ano98-162, Ano99y, Ano99-31, Ano99m, Ano99-35, Ano99-36, Ano99-40, Ano99-61, Ano99-62, Ano99-72, Ano99-139, Ano99-138, Ano99-143]. CA [Ano99-141, Ano99-164, Ano99-148, Ano99-147, Ano99-221, Ano99-173, Ano99-176, Ano99-177, Ano99-178, Ano99-179, Ano99-180, Ano99-195, Ano99-187, Ano99-188, Ano99-192, Ano99-230, Ano99-210, Ano99-235, Ano99-236, Ano99-238, Ano99-237, Ano97q]. CA/Cambridge [Ano95-47, Ano96h, Ano97r, Ano97-148, Ano97-129, Ano99-40]. CA/MIT [Ano96-44, Ano96-36, Ano96-53, Ano96-135]. CA/New [Ano98-70]. CAD [YT98]. Caenepeel [Ano98w]. Cahan [Ano96-141]. Cahn [CC98b]. calculate [KCB90]. Calculating [Joh99, AO93a, Ano99-244, BB99, Ano99-38]. Calculation [MSB90, Feu90, Feu91, Fre98, KO93, Mil98]. calculations [Gre91a, MS99, Ano96-137]. calculi [Zad99, Ano99-68]. calculus [AIS95a, AIS95b, Ano97m, Ano97-125, CIS98, GT92, KNS99, LS95, RS96a, RS99a, SA95, Ano97-152]. California [Ano96-135]. call [Saw92]. Callaway [Ano97-83]. Callender [Ano96-71]. calls [CK97b]. Calogero [Gre90a, MG91]. Calugareanu [Ano98-53]. Camarinha [Ano99-128]. Camarinha-Matos [Ano99-128]. CAmbridge [Ano96l, Ano90w, Ano93f, Ano93t, Ano93u, Ano93i, Ano93j, Ano93l, Ano93v, Ano93s, Ano93w, Ano95k, Ano95-36, Ano95h, Ano95-70, Ano95o, Ano95r, Ano95s, Ano95d, Ano95q, Ano95-71, Ano95-28, Ano95-29, Ano95-77, Ano95-37, Ano95-38, Ano95-78, Ano95-40, Ano95-45, Ano95-41, Ano95-47, Ano95-59, Ano95-60, Ano95-62, Ano95-81, Ano95-82, Ano96s, Ano96i, Ano96h, Ano96k, Ano96t, Ano96-32, Ano96-39, Ano96-44, Ano96-36, Ano96-43, Ano96-34, Ano96-31, Ano96-35, Ano96-153, Ano96b, Ano96-30, Ano96-40, Ano96-47, Ano96-46, Ano96-55, Ano96-53, Ano96-52, Ano96c, Ano96-69, Ano96-67, Ano96-61, Ano96-57, Ano96-66, Ano96-62, Ano96-58, Ano96-72, Ano96-74, Ano96-73, Ano96-78, Ano96-87, Ano96-81, Ano96-80, Ano96-85, Ano96p, Ano96-90, Ano96r, Ano96n, Ano96o, Ano96-95, Ano96-94, Ano96-96, Ano96-104, Ano96-100]. Cambridge [Ano96-99, Ano96-103, Ano96-154, Ano96-97, Ano96-114, Ano96-105, Ano96-118, Ano96-115, Ano96-155, Ano96-122, Ano96-121, Ano96-158, Ano96-134, Ano96-128, Ano96-132, Ano96-138, Ano96-140, Ano96-139, Ano96-148, Ano96-142, Ano96-149, Ano96-162, Ano96-84, Ano96-135, Ano96-145, Ano96-160, Ano97g, Ano97f, Ano97-53, Ano97l, Ano97p, Ano97-142, Ano97r, Ano97-31, Ano97s, Ano97-28, Ano97-32, Ano97u, Ano97x, Ano97y, Ano97z, Ano97v, Ano97-41, Ano97-43, Ano97-37, Ano97-44, Ano97-48, Ano97-47, Ano97-49, Ano97-57, Ano97-63, Ano97-55, Ano97-144, Ano97-64, Ano97i, Ano97k, Ano97-69, Ano97-83, Ano97-81, Ano97-88, Ano97-80, Ano97d, Ano97-149, Ano97-95, Ano97-148, Ano97-90, Ano97-98, Ano97-99, Ano97-97, Ano97-91, Ano97-94, Ano97-150, Ano97-109, Ano97-110, Ano97-115, Ano97-137, Ano97-156, Ano97-153, Ano97-134, Ano97-133, Ano97-160, Ano97-159, Ano97-52, Ano97-129]. Cambridge [Ano98h, Ano98k, Ano98s, Ano98u, Ano98-34, Ano98-28, Ano98-38, Ano98-35, Ano98-41, Ano98-40, Ano98-48, Ano98-61, Ano98-58, Ano98-63, Ano98-70, Ano98-73, Ano98-82, Ano98-87, Ano98-96, Ano98-94, Ano98f, Ano98-95, Ano98-106, Ano98-104, Ano98-113, Ano98-149, Ano98-114, Ano98-119, Ano98-128, Ano98-150, Ano98-151, Ano98-152, Ano98-147, Ano99-29, Ano99-30, Ano99-90, Ano99q, Ano99p, Ano99-40, Ano99-38, Ano99-68, Ano99-75, Ano99-76, Ano99-92, Ano99-106, Ano99-107, Ano99f, Ano99-118, Ano99g, Ano99-119, Ano99-216, Ano99-217, Ano99-122, Ano99-158, Ano99-160, Ano99-154, Ano99-159, Ano99-161, Ano99-155, Ano99-163, Ano99-220, Ano99k, Ano99-226, Ano99-228, Ano99-232]. camera [Lee96, LT94]. Campbell [Ano99-218, Ano99-219, GS95]. can [Ano99-216, Ano99-217, HL99]. cancer [Ush94b]. Canonical [Mär96, QZ91, CKL91, KH92, ZQ93d]. canons [Ano97-40]. cantilever [NAKA98]. Cantor [Nas95]. Cap [Ano99t]. capability [Ano96-151]. capacities [Käm94b]. Capacity [Bro90, CCY95a, CCY95b, NSS92, TBA99, Tsi91]. capillary [Jon96, TNTR91]. capillary-gravity [Jon96]. capturing [Gla94e, Gla95d, Gla95e]. carbon [KG97, WT97]. Cardinal [CWJ92, Sun92, Bax92]. cardinalised [LSH95]. cards [CH93]. careers [Ano97h]. Carl [Ano95-81, Ano97-45]. Carlitz [GW97]. Carlo [GHM96, JI94, Kou98, Mic90, SC98b]. Carlos [Ano98-140, Ano99o]. Carlton [MM96e]. Carmichael [Ano98-40]. Carol [Ano96-126]. Carolyn [Ano96o]. Cartan [Shu93a]. Carter [Ano99u]. Cartesian [AAD98, Ano99-39]. Cartography [Ano96-108]. Cary [Ano99-28]. Casati [Ano96-87]. cascade [Ano99-242, Chu99, MATA98, WF91]. cascadic [Sha96a]. case [CC92, DIT97, Gal93a, GMW96, Kel94, Ker91a, Li98c, LO99, Nak99, PvZ95, TTC96, VA94, Ano97r]. Case-based [Ano97r]. Castano [Ano96-48]. Casti [Ano96-167]. casualty [Ano98-148]. catastrophe [Ano99-214]. categorical [Yoo93]. categories [Ano96t]. Caterpillar [LW97]. Caterpillar-Spannable [LW97]. Cauchy [ES97, Kim99, LOD96, PTC^{+}98, SZ91, UWT99]. causality [Ano93s]. causation [Ano99-40]. cavities [CHW93]. cavity [BR95a, CR90a, KKG97, KMD99, LT96]. Cay [Ano96-109, Ano97-27]. CD [Ano96-59, Ano96-120, Ano97o, Ano97q, Ano97-111, Ano97-101, Ano97-112, Ano97-102, Ano97-103, Ano97-124, Ano97-131, Ano98p, Ano98t, Ano98-41, Ano98-105, Ano98-149, Ano98-122, Ano98-125, Ano98-162, Ano99y, Ano99-35, Ano99-36, Ano99-37, Ano99-69, Ano99-78, Ano99-79, Ano99-70, Ano99-129, Ano99-130, Ano99-154, Ano99-155, Ano99-221, Ano99-180, Ano99-230, Ano99-227, Ano96-59, Ano99-221]. CD-ROM [Ano96-59, Ano96-120, Ano97-111, Ano97-101, Ano97-112, Ano98p, Ano98t, Ano98-41, Ano98-125, Ano99y, Ano99-35, Ano99-36, Ano99-37, Ano99-69, Ano99-78, Ano99-79, Ano99-70, Ano99-129, Ano99-130, Ano99-154, Ano99-155, Ano99-180, Ano99-230, Ano99-227, Ano96-59, Ano99-221]. Cebysev [Ano99-244, BB99]. Celia [Ano93j]. Cell [Aug90, CGH92, BCB90, BM89, BM90c, DC94, Gyö90b, LZ98, ML90b, Mic90, MH93a, Mol96, Pou90b, YF92b]. cell-based [DC94]. cell-centered [Mol96]. cell-to-cell [MH93a]. cells [ZK99]. Cellular [Ano97-96, Ano97-46, Ano99-166, BP90, BFF^{+}97, CCD99, GB90, KCD^{+}97, KCDV99, Sie90]. Censor [Ano98-124]. censored [JCL93]. centennial [OR97b, OR97a]. center [CCGL99]. centered [LR93b, Mol96, Ano90x]. Central [Kor90b, Sha90a]. centrifuge [Ano99-242, MATA98]. centroids [Köh97]. century [Ano98-110, Iri98, Ano97-151]. Cerny [Ano99-117]. certain [AG99a, AG99c, AIS95a, AIS95b, Azb99, Bal98, Bar91a, CIS98, Cim90, DBM92, Dat93, Gra94, HS95, Har96, Haw91, Hen97, Ise94, JS99, Lag91, LS95, MM95c, PXHG99, RS96a, RS99b, SA95, TMGS98, US91, ZC91, ZY98, ZL99b]. Ceruzzi [Ano99f, Ano99g]. CFD [Ano99-205]. CG [CP98b]. CGI [Ano96-28]. CGS [SBS99]. Ch [Ano93o]. chain [HP99, KMŚ98, Tan96]. Chair [Ano96-135]. Chakravarthy [Ano98r]. Challenger [Ano96-152]. Challenges [Ano96-127, Ano98r, Bir98]. Chambers [Ano99-193, Ano99-194]. Champaign [Ano97-90]. chance [Liu97]. chances [Ano92a]. Chandler [Ano97-147]. Chandrasekaran [Ano96-53]. Chandrasekhar [Ano97-130, Ano99-91]. Chandru [Ano99-175]. Chang [Ano98-161]. change [PBG90, Sou90, LH97, Ano96-94]. changed [Ano95i]. changes [Mur92b]. changing [Ano96-167]. channel [Gla95d, Gla95e, GS96b, KLMY97, Sau92, SC98b, WHC90, WW95, YSM92, YFZM95a, YFZM95b]. channels [Gla92c]. Chaos [ABB97, Mur93a, AR90c]. Chaotic [ABB97, ADFP93, Ano96-56, BKN97, Cor94, DSKP96, GB92, GB96, Mag97, RZ96, Tso91, WAL97, Ano95p]. Chapman [Ano95m, Ano95l, Ano95-53, Ano96e, Ano96f, Ano96w, Ano96-56, Ano96-65, Ano97-62, Ano97-93, Ano97-165]. character [Ano99-140]. Characterisation [CCD99]. characteristic [BPLS96]. characteristics [Moi97, NMUY92, SHK99]. Characterization [Bal99, CW95, GGK97, MC97, Kur91, LX98, WW93]. Characterizing [Mil91, KBLS90]. charge [WW98]. charged [KESH90]. Charles [Ano90x, Ano96-151, Ano97-156, Ano98-141, Ano99-29, Ano99-30, Ano99-65, Ano99-177]. chart [Oda91]. Chartrand [Ano97-62]. Chebyshev [BS99a, CGH92, EDOS93, EGEHSEK95, Kar94, MSB90, Pan98c, RM98a, Riv97]. Chebyshev-type [MSB90]. checking [Ano99-86]. checkpoint [FKO92]. chemical [Aug90, Gre99, JR94, KG94, Tau90a, TM91]. chemistry [Edw92, Gla90a]. chemotherapy [Ush94b]. Chen [Ano99-94].

Cheng [Ano99-86]. Chesser [Ano96-130]. Chi [Ano99-78, Ano99-79, Ano98x, Tem93]. chi-square [Tem93]. Chicago [Ano95n, Ano95y, Ano95x, Ano95-74, Ano95-73, Ano95-30, Ano95i, Ano95-31, Ano95-75, Ano95-32, Ano95-43, Ano95-66, Ano95-69, Ano95-54, Ano95-68, Ano96x, Ano96-152, Ano96a, Ano96-50, Ano96-63, Ano96-60, Ano96-64, Ano96-89, Ano96-108, Ano96-127, Ano96-130, Ano96-144, Ano96-141, Ano96-161, Ano96-143, Ano96-163, Ano97h, Ano97-40, Ano97-39, Ano97-143, Ano97-50, Ano97-59, Ano97-82, Ano97-146, Ano97-100, Ano97-151, Ano97-135, Ano97-158, Ano97-155, Ano97-128, Ano97-154, Ano97-127, Ano97-130, Ano98v, Ano98-27, Ano98-39, Ano98-148, Ano98-60, Ano98-136, Ano98-131, Ano98-134, Ano98-145, Ano98-146, Ano99-51, Ano99-39, Ano99d, Ano99-214, Ano99h, Ano99i, Ano99-140, Ano99-144, Ano99-146, Ano99-224, Ano99-225, Ano99-203, Ano99-202, Ano99-233, Ano99-91, Ano99-239]. Chicagp [Ano99-202]. Chicester [Ano99-45, Ano99-74]. Chichester [Ano93p, Ano95u, Ano95-33, Ano96g, Ano96-120, Ano96-131, Ano97-125, Ano97-136, Ano98m, Ano98-64, Ano98-154, Ano99-47, Ano99-48, Ano99-59, Ano99-71, Ano99-98]. Chierchia [Ano95-30]. Chikrii [Ano97-29]. Chilukuri [Ano97-44]. China [Ano99-137]. Chip [Ano99-178, Ano99-201]. Chirality [Mez97]. choice [FH99, Gün99, Nai92, NO93, Sod90, Ano93-27]. choices [Ano98-159]. Cholesky [Més96, Res91, Tor90]. Chomicki [Ano99-145]. chopping [KG98]. Chordal [KM97a]. choreography [ALL^{+}96]. Chris [Ano93g, Ano99-173]. Chrissis [Ano96-151]. Christian [Ano97-83]. Christof [Ano95-41, Ano98-104]. Christoffel [Has97]. Christoph [Ano99-29, Ano99v, Ano99-30, Ano99w]. Christopher [Ano95y, Ano97-164]. Chuck [Ano96-88]. Chun [Ano99-182].

Churchland [Ano93t, Ano97-149]. Ciarlet [Ano96-83, Ano97-65, Ano99-113]. Cichocki [Ano98-164]. Cieslik [Ano99-206, Ano99-207]. Cihan [Ano95l]. Cios [Ano99-52]. circle [PCR90]. circles [Ano97-162, Ano96-29]. circuit [Ano99-132, MRWM94, Ano95q]. circuitry [NR96b]. circuits [GHPLZ90]. circulant [CTY95, ESIP98, Roj90]. circular [He95a, PSD96]. circumstance [DO95]. citizens [Ano97-101]. City [Ano96-30]. civilization [Ano96-108, Ano99-222, Ano99-223]. Clan [BMR96]. Clan-based [BMR96]. Clarendon [Ano93g, Ano98y, Ano98-55, Ano99-135]. Clarissa [Ano95-74]. Clark [Ano99-40]. Clarke [Ano99-172]. class [Ame99, Ano97-121, BW96a, Bai96c, BKM91, CCGL99, CAZAS95, CZ94, Coh94, CSW99, CK94a, CH99, DH98b, DH98a, DNJS98, DSKP96, DKR97, DLS99, EDO97, EK94, EA90, GOWR^{+}93, GS96a, Har96, Hua98b, Jha90, Kat92, KG98, Kij92, KNS99, LTR98, LW96, LC98a, Liu97, LW99b, Luo93c, Mak90, MM95c, NFJ96, OH93, PA94, PA98, PM91, RTJ96, SSBG91, Tor90, Ver99a, Wan98a, Yan99b, ZC99, ZT99]. classes [GS96a, Kov95, NF92, SHA99b, TMGS98]. Classical [BNR96, CCC^{+}98, AD98, APV97, Ano96-146, Ano97-87, Ano97-92, Ano97-157, PP97]. classification [FL96, Gal91b, GMBS94, HW90]. classify [Mey99]. Classifying [CF91b, PS93a]. classroom [Ano99-89]. clause [BRGJ95]. client [Ano97-166]. Cliff [Ano95-36]. Clifford [Ano96-27, Ano97-148, Ano97-125]. Cliffs [Ano90p, Ano90q, Ano90s, Ano90x, Ano93k, Ano93n]. climate [DF98]. clinical [Ano96-43]. Clinton [Ano97-166]. clique [BS91b]. clique-partitioning [BS91b]. Close [Ano98-141, BF97, BLR94, Moo93a].

Closed [DH91, Sjo92, AO99a, AO99b, Ano98-134]. Closed-form [Sjo92]. closeness [Gar96]. Cloth [Ano90u, Ano95d, Ano95t, Ano95y, Ano95x, Ano95-29, Ano95-73, Ano95-30, Ano95-31, Ano95-69, Ano96-50, Ano96-60, Ano96-64, Ano96r, Ano96-108, Ano96-130, Ano96-144, Ano96-141, Ano96-161, Ano96-143, Ano97-39, Ano97-59, Ano97-82, Ano97-151, Ano97-135, Ano97-128, Ano97-154, Ano97-127, Ano97-130, Ano99d, Ano99h, Ano99i, Ano99-135, Ano99-140, Ano99-224, Ano99-225, Ano99-203, Ano99-233]. cloud [LCSN98]. Cluster [EL99, CT90, CHW99, SSN94]. clustering [EL99, Yan93b]. CMOS [NR96b]. co [Ano98-69, Cha93]. co-design [Ano98-69]. co-integrated [Cha93]. coaxial [ASW98, PSD96]. COBOL [Ano98-143]. cock [Ano97-79]. Codd [NGNM97]. code [Ano98-137, BP94b, HSH90a, KM96, LSH98, SH94]. Codes [ABVB92, BG90, AG91, Cas98, Jha90, Ngu97, Ram95, VAB91, WJ90, Yan93c, Ano98-81, Ano98y]. Codesign [Ano98z]. coding [Ano93g, Ano96-147, Lia99b, SHK99, Ano96r]. coefficient [AAII92, HM93b, JCM96, JPF96, Kwo92, MM93b, MM95a, NN91, SS94, TG96, Tre95a, Var98, Wan96, ZZ99c]. coefficients [AIS95a, CLY95, CS93b, Doh91, EDNO92, EDO96, FHWW99, GW91, HM93a, Hin95, KP97a, KNS99, RS99b, Sch92b, SA95, VHH95, YCZ98, ZL98a, dJF91]. cofactor [SHT^{+}90]. cognate [HSH90a]. cognition [Ano96-47, Ano96c, IK95, Ano95-31, Ano96-31, Ano96-84, Ano99-220, Ano96-31, Ano97s]. Cognitive [Ano96-53, Ano96-34, Ano97-143, Ano97-120, Ano99-92, Ano96-32, Ano98-28, Ano98-27]. cographs [LOP95]. Cohen [Ano95u, Ano96-61, Ano97-47]. coherent [BKK97]. Cohn [Ano99q]. cohomology [Ano97-140]. cokriging [Mye92]. Cole [Ano96-105]. Coletti [Ano96-113]. Colin [Ano99-218, Ano99-219]. collaboration [Ano95-62]. Collected [Ano96-36]. collection [CGS99, OR97b, OR97a]. collective [FWI99, ZZ94]. College [Ano99-32]. colliding [Gre90b]. collision [Kou98, RW98, Tch98]. Collisions [Wya94]. Collocation [SAS99, Ahm90, Ano98-165, BR95c, Bon97, Fun97, Kar94, KO96, LTB96, MMP98, Rob97, RT99b, SZ91, ZS93]. Collocation-continuation [SAS99]. collocation-interpolation [BR95c, Bon97, LTB96]. colloquim [Ano99-152]. colloquium [Ano98-101]. Coloring [Cim90, BP90, DdW93, JL94]. coloured [Ben99, Kis91]. Colouring [SS91b]. COMAP [Ano97-122]. combat [RA92a]. combination [TDF^{+}92]. combinations [Arn90, Sto92]. Combinatorial [Kos93, Ano98m, Ano99-127, FWI99, Ano96-33, Ano98-29, Ano98-30]. combinatorics [Ano96-81, Ano96-80]. combined [MMP98, Pic91, WCY94]. combining [Sir99]. combustion [ZSDB99]. command [YFZM95a, YFZM95b]. commemorate [OR97b, OR97a]. Comment [McL95, ZQ95, Zel92]. Commentary [Jay92, Ano96a]. Comments [Har92a, Ter00, Los92]. Commerce [Ano96-116, Ano97-167]. Commins [Ano96-128]. commodity [BM90b]. common [Ano97-134, Los89b]. communicating [Ano99-212]. communication [Ano97-36, BGM95, YNK92, Ano99-106, Ano99-107, Ano97t]. communications [Ano96y]. Community [Ano97h]. commutative [Ano98-115]. Commuting [Xu97]. Compact [LK99, PR90, Ano95-43, AH97, Bad95, CY94, PR93, PR96, Sun92, Tan93a, Tse95]. compactness [Bri97]. companion [CKL91]. Company [Ano95-72, Ano95v, Ano95-35, Ano95-34, Ano95-42, Ano95-50, Ano95-48, Ano95-56, Ano95-61]. comparative [CMMR92, HNS^{+}90, Ano96-34]. Comparison [Ana95a, HP99, JLDL92, SB99, SBS99, BW96b, Bai96d, BT94, Boy99, CS96, CQ98, Cas98, CR90c, DGV94, Gla95b, Gla95c, MK98, NL93]. comparisons [HS94, KO97, KO99]. compartmental [BM89, BM90c]. compartmentalized [Par91b]. compensator [Tad94]. compete [Ano96-118]. competing [LL96, VMA99]. Competition [Ano97-161, Ano97-139, BT95, BTED97, PC94b]. competitive [KL96b, TC99]. Compiled [Ano95-54, Ano97-150]. compiler [Ano97-98, Ano97-99, Ano97-97, Ano98-106, Ano98-31]. complementarity [Bai96b, BW96a, Bai96c, Bai96d, Cub96, Gup94, Yao93]. complementation [Ano98-151]. Complete [Ano98-61, Ano99-213, Ano99-158, Ano99-160, Ano99-159, Ano99-161, AK90b, PDK96, TYC93]. Completely [Gre95, Hua98b]. Completing [Coh92]. completion [TPF98]. Complex [Sib98, Sti93b, Ano95p, Ano95-66, Ano97-121, Ano99-216, Ano99-217, ABI99, Bak92, BG91b, CCC^{+}98, Gil91, KBLS90, Pan95, PC91, Riz95, RZ95, RSR93, Sie90, TNS98, YX95, Yas98b]. Complex/generalized [Sib98]. complexes [KMŚ98]. Complexity [Ano97-108, MCG98a, Ano95q, Ano97u, Ano97t, Ano97j, Ano98-32, Ano99-132, Ano99-233, Bro94, CD98, DKW98, EG97, Eso99a, Li96a, MCG98c, PR90, Res91, Wit96b, Ano99-60]. complicated [BS99b]. component [Ano99-121, Gla92e, HW96, KESH90, Zar92]. component-consistent [HW96]. components [Leo90, Los89b, Ano99-61]. componentware [Ano98-149]. composed [WW98]. Composers [SH96]. composing [Opp96]. Composite [Kar96b, Ano98-57, FCP95, Flo97, Köh94, MV95a, Sez95, SD93]. composition [Hel96, Leo96, ML90b, Rot96, Sco96]. compositional [Ano96b, Cho96]. comprehension [Ano96-158]. compressed [Ano98x]. compressible [Gla91c, Gla91d, Gla92b, Gla94b, Gla95f, Gla97a, Gla99, Rac99, YX95]. compression [BLR90, KCDV99, Sad96a, Ano98-72]. Comput [Li10, Moo93b]. computability [Ano98q, Ano96-35, Ano97u]. Computation [Ano90a, Cho96, CK91, EK92, YFZM95a, YFZM95b, Ano93w, Ano96j, Ano98-93, Ano98-118, Ano99-81, Ano99-166, Bak92, BL96, Bar93, BP94a, BR94, Bro98, Che99a, DGV94, Ech96, EKC91, FWI99, GT97a, GT97b, GM99, JCM96, KMŚ98, ND95, PYS97, Rei93a, Rou92, WF99a, YL93a, Ano93l, Ano95j, Ano96-165, Ano97k, Ano99-232, Ano96-37, Ano96-36, Ano99-40]. Computational [Ano95-31, Ano97-73, Ano98-80, Ano99-118, Ano99-119, Bal90, EA90, Eso91, KM90, Pom91, AAL99, Ano90-28, Ano93t, Ano96j, Ano96-32, Ano96-53, Ano96-126, Ano97-51, Ano97-119, Ano98r, Ano98-125, Ano99-70, Ano99-111, Ano99-112, BLP95, Cse95, DKW98, Eso99a, Gla95b, JTY97, Kan90a, Kan90b, LR93a, MH99, Wea92, Ano95r, Ano95s, Ano96-39, Ano96-38, Ano96-40, Ano97w, Ano97v, Ano98-33, Ano98-32, Ano99-41, Ano99-42, Ano99-44, Ano99-43]. computations [Ano96-129, Dea98, KO94, LK97, Pan92a, RM98a, Rhe96, ZM99, Ano98-32, Ano99-229]. compute [Ano95-76, TPF98]. computed [Bou99, BS92]. Computer [Ano98-93, Cho95, DC94, Gre93a, KV94a, Lei90, LH94, Lu96, Pag90, RSV98, TG96, TNS98, Wan96, Ano95-71, Ano96-150, Ano96-153, Ano96-147, Ano97-28, Ano97-137, Ano97-133, Ano97-129, Ano98j, Ano98n, Ano98-64, Ano98-70, Ano99-74, Ano99-216, Ano99-217, Ano99-150, Ano99-190, Ano99-191, Ano99-242, ANV94, Bec96, CC92, Che91, CG95, Cho99, EA92, HBB90, Han92, Leo96, LH99, Lia99a, MATA98, Mro96, NAKA98, Opp96, Pat98, Rot96, RG90c, SKI92, SK91, SZ98a, Tso91, hCPW98, Ano93i, Ano95-40, Ano90p, Ano90q, Ano95t, Ano95v, Ano95u, Ano95w, Ano96-42, Ano96-43, Ano96-41, Ano97x, Ano97y, Ano97z, Ano99-46, Ano99-45]. Computer-aided [Cho95, Ano95t, Ano96-41]. Computer-integrated [Ano96-43]. Computer-related [Ano95v]. computerized [TM91]. Computers [Ado91a, Ano99-222, Ano99-223, BM90c, Kou90b, MM94, MG90b, Sal92b, SL99, Zad92, Amo93, APP96, AL94, Ano96-58, Ano96-122, Ano96-162, Ano97-148, BP94b, CSW99, Kal90, PC92, Wit90, hC96b, Ano95-64, Ano99-81, Ano99-215, Ano99-166, Ano96-44, Ano98-34]. Computing [GNR95, JP95a, KO99, LPR96, RS99c, SBD99, Win91, AGM99, Ano96-86, Ano96-137, Ano97o, Ano97-65, Ano98p, Ano98-68, Ano98-123, Ano98-139, Ano99-66, Ano99-67, Bak90, CS93a, CHCC95, CCY95a, CCY95b, CCH97, KO97, LL91b, MM96b, Ter00, Trz92, WCL99, ZM94, Ano90w, Ano93m, Ano95-34, Ano95-59, Ano96-62, Ano96-75, Ano96m, Ano96o, Ano97t, Ano99f, Ano99g, Ano96-45, Ano97-28, Ano97-27, Ano99-47, Ano99-48]. computing. [Ano97x]. concave [BH91]. concavity [Elo98]. concentrating [MT98]. concentration [Ano98-62, Dor98a, Dor98b]. concentrations [ZFM96]. Concept [Wil92, BR95b, CL94c, Köh91, Luo93a, Pol90]. conception [Ano99-140, KM93]. Concepts [Ano98-164, Mar92, Ano97-27, Cas94, DHW98, Gar96]. Conceptual [Ano96-85, Lyt92a, Sow92, Chu99, Hau92, Wil92]. Concerning [CLM92, KC91, dJPvZ97]. concise [Ano99d]. Concrete [Ana95b, Ano95c]. concurrency [NJ92, Ano98-35]. concurrent [MR99, Sti97, Yoo93]. condensation [CC94c, Fli94b, Fli94a, Fli94c, Fli95, TSLW98]. condensing [O’R95]. condition [CC97, GBB91, LM91, MM95a, Roj99, Tan93b, Zaf98b]. Conditional [Yos90, BSP94]. conditioned [Sha90c]. conditioning [JPF96]. Conditions [Ker91a, Pap94b, AR98, Ano98-36, BJH99, Bal96a, Bed95b, Car92, CKM99, Dor90, Far96a, GS93, HS98, Il’97, Kar94, Kla98, Kur91, Lar90, Lay99, Li99a, LPR96, LOD96, MN97, Mei93, Mur92a, O’R98c, RS92, SD95, ST95, YP98b, Yav91b, ZL99a, ZQ93d]. conducting [CR90b, EL96, LLP94, MMZ91]. conduction [Far99, Fra96, GMR90, Kha97, MG90c, Mur92b, MH93b, She99]. conductive [YP98a]. conductivity [Ram91b]. cone [dKSG94]. cones [Sza91]. Conference [Ano99t, Ano90-28, Ano95-36, Ano95-45, Ano96e, Ano96-67, Ano96-104, Ano96-135, Ano97-53, Ano97-47, Ano97-51, Ano97-57, Ano97-88, Ano98-96, Ano98-95, Ano98-111, Ano99-29, Ano99-30, Ano99-90, Ano99r, Ano99-120, Ano99-128, WY93, Ano96-166, Ano97g, Ano98h, Ano99q]. conferences [Ano97q, Ano98-140]. Confidence [FF92]. confidentiality [DFMS99]. configuration [LY99a]. configurations [CCH^{+}90, Son92]. Conflict [Ano97-29]. Conflict-controlled [Ano97-29]. conflicting [CL94a]. Confluent [Sal91, Sal92b]. conformal [HGK93, Ano97-30]. conforming [Sau95]. Congress [Ano90a]. congruence [Koç92]. conic [BC97]. conics [CG95, Qu96]. conjecture [Gyi96]. Conjugate [CLM99, PS92, ABW99, Ara96, Ara99a, Bro96, Heg91a, Heg91b, Luo93b, PZDH95, Sha96a, Sha95, Won98b]. conjugate-gradient [Sha96a]. connected [AS95, Bel97a]. Connectedness [Ano98-36]. connection [ALL^{+}96, EGLN95, Ste90]. connectionism [Ano97-144, Ano97-31]. connectionist [Sha92]. Conner [Ano95-57]. Connor [Ano90q]. consecutive [GMW96]. consensus [FLNW98]. consequences [US91]. Conservation [HSH90a, Sha99a, Cha96, Che93a, Gla93c, Gla93e, Gla96c, Hua95, KOT99, Oli93, Prn97, Sha98, Wan91]. Conservative [CC98b, FG98, Gre90a, Gre94, Hua95, Nke94, Söv99, Tch98, Ber96, FG97, Gre92b, Gre95, Kan92a]. conservativity [Tan93c]. conserving [MG91, Mar94]. consideration [AR92a, YNK92]. consistency [GWL96, LRLR95]. Consistent [NWB92, Bou93, Dor98a, Dor98b, HW96, KO97, KO99, MAM91, Zhu92]. conspiracy [LS94]. constancy [SWY96a, Zha98]. constant [AAII92, Ano96-57, BM97, CG94, Dem94b, Gal93a, KAD99, Kwo92, NN91, Tre95a, Wan96]. constant-coefficient [AAII92, NN91]. constants [BL95, Sto99]. Constitutive [Fli94a]. Constitutive-operator [Fli94a]. Constrained [PN99, AK90a, Ano97-144, CW95, ER98, GRZ96, JTY97, KNUX99, LTR98, Par91a, Sha95, ZZ99d, dKSG94]. constraint [Ano96-134, Ano97-85, GHPLZ90, KHC92, SZ98a, Sny90, Ano98-37]. Constraint-based [Ano98-37]. Constraints [Ano95r, Ano99-94, Ano99-171, BQN97, BS99b, CCY95a, CCY95b, CF98, FHWW99, Gue93, Har92c, Jan96, Köh96b, Med90, MPR98, SFY92, Ste90, WW98, WT90b, WF99b]. construct [AS95]. constructed [Zhu92]. Constructing [POH99, ZQ93c, DK97c, RA92b, VAB91]. Construction [JCM96, Mic94, QZ93c, TDF^{+}92, Ano99-80, ANV96, BSW96, KV95, LD90, LKC95, Ngu97, PR91, QA98, Sch96, SS96]. constructions [SS99]. Constructive [Ame98, JPC99]. contact [Bri97, ESKG98, PSS99, SSZ91, Shu93a, Shu93b, Tan93c]. containing [Ano99-172, VR95a]. contemporaries [Ano96-50]. Content [Ano98-109]. Content-based [Ano98-109]. contents [Ano90-29, Ano90-30, Ano91l, Ano92h, Ano93-29, Ano95-86, Ano95-87, Ano96-172, Ano96-173, Ano97-174, Ano97-175, Ano98-168, Ano98-169, Ano99-247, Ano99-248]. context [BCGP99, Ano98-34]. contextual [Ano98-100]. Contingency [KM97a, DFMS99]. continua [KPS96]. Continuation [AO99a, AC93a, BK91, CCCY94, GP91, MM95b, SAS99]. Continuation-minimization [AC93a]. continued [AE90, EKC91, LL91b, Mac93]. continuity [Ana95a, LX98, RFRM99, Sch96]. Continuous [BM91, BL94, DKO93, JP95b, Ano98n, Ano99l, BL95, BM97, Boh96, Bon99, Cha91, CSOL96, DOS93, EDNO92, Gan92, JP95a, JA96, Kis91, KLMY97, Lew94, LK99, Pad92b, Pic91, Pro95, Rom98, Sar90, Sar94, SZC91, Tra97, TH95, ZYC98, ZL99a, Ano99-49]. continuous-time [SZC91, Tra97]. Continuously [Chi93, CT96, YC92]. Continuum [Ano90b, Ano97-149, Eso90, Nam97, Nam99a]. contour [HS95]. contoured [Sza90a, Sza90b]. contours [Byk97]. contract [Ano95-37, Ano98-63]. contraction [CH99]. contractive [HB99, O’R98b]. contractor [Alt90a]. contradictionary [Ano95-71]. contribution [Har92b]. contributions [Ano99-172, Eso90]. Control [Eis92, FYS99, Gue93, Oda91, WR93, Alf95, Ame99, Ano96-122, Ano96-171, Ano99-93, Ano99-98, ABB97, ARS98, BMR96, BT95, BK91, Bes96, BS96d, Bou98, BNV90, Cha98, CB93a, CFLC95, CT96, Cra98, DBA98, DG91, DNJS98, DKO93, EGEHSEK95, Erm92, EM94, Eso99b, Fog94, Gal91b, GZ93, GOWR^{+}93, GB96, GM96, GG92, GHS91, HM91, JM91, JW93, KMR90, Kra92, KS97b, KS98b, LH97, LTC98, LTR98, LDD94, LB94, LL90c, MJK92, NJ92, OHI91, Pou90b, RG90b, RHG91, SFY92, Sti96, SI94, Tad93, Tra97, TT92, WW93, WAL97, WT90b, WCY94, YNK92, Yav91b, YF92a, YZM94, YZM95, Yav96, YL93b, Ano99s, Ano96-46]. controllability [DBA98]. controlled [AD93, Ano97-29]. controllers [HP96, LCSN98, SZC91]. controls [Gal93d, RTJ96]. convection [BGL90, Che92, Gol90, Ha90, HH98b, MS96, RM98a, SS98a, ST95, WF99a, vNvN90a]. convective [AKK^{+}97, DF94, Kwo92]. convective-diffusion [Kwo92]. Convergence [AC94, AP98a, DH98a, FT97b, HK96, KK97b, KG95b, SMU98, Yan93b, YC94, AC95, Alt90b, AB94, AR99a, Arg98, Arg99, Bai96b, Bai96c, Bai96e, Bai96f, BC98, BG91b, CC96, Dem94a, Dem94b, Far99, HS95, Hua97, KAD99, KT95, Kun90, KS91, KN94, LA96, LGP98, Sha96a, SCBFC98, ST95, SCC97, Tan96, VPK99, Yan99a]. convergent [CMP95, Car92, Li97, Li98a, LN98, LN99, Li10, Oza94, Pas97, Pin98, Tre95b, Xue94]. converging [Ker91a, MP93]. converging-diverging [MP93]. Conversations [Ano95-69]. converse [GlZ95]. conversion [Koç91, Lew91b]. convex [BH91, Bör90, CF98, Che99a, Din99c, GMBS94, Kai94, Kai96, KKS98, Kra90, Liu96a, Liu96b, Ola92, PLH95, Sch96, SH95a, SY96, WF99b]. Convexification [WCHL96]. convexity [Ano99-95, BL99, Ver99b, Yas94]. Convolution [Ano98-43, Has93a, RB97]. convolutional [AG91]. Cook [Ano98-29, Ano99-199, Ano99-200]. Cooke [Ano97-80]. cooling [RD96b]. Coombes [Ano96-54, Ano98-42]. Cooper [Ano96-35, Ano99-40]. Cooperative [BKK97, Ano98-156, SJS92, ZC96]. coordinate [Gla93d, Gla95d]. coordinates [Gla93a, Has93b, Has95, Wad95]. coordinating [Ano95t, Ano97-32, Ano98-38]. Coordination [Ano95x]. Copies [Ano97-141]. copositive [Hu95]. coprime [CGS99]. corank [ACL93]. CORBA [Ano98-41]. core [Ano93u, Ano97-156, Ano98-103, DHW98, Ano99-158, Ano99-159]. Corless [Ano93o]. Cornelius [OR97b, OR97a]. corporate [Ano90s]. Corporation [Ano96j, Ano97-132]. correct [MR94]. corrected [Rok90]. correcting [ABVB92, Ano98-81, Jha90, WJ90, Yan93c]. correcting/all [WJ90]. correcting/unidirectional [Jha90]. correction [BACS98, CP98b, HW96]. corrections [PC93, Zha98]. corrector [VA97, VK99]. correlated [YI97]. correlations [WAL97]. corresponding [CCVV90, Sok91]. Corrigendum [Ano98-165, Ano00, Li10]. Cosine [Mak90]. cosmological [Ano96-57]. cosmology [Ano97-61, Ano98-65]. cosmos [Ano99j]. Cost [HO97, Kur91, Ohn92, SOI92]. Coulomb [Mic98]. Council [Ano99-89]. counterexample [PR91]. counting [DWC95]. country [Özy96]. coupled [Ado96, BR95b, BG95, Boy96, JG95, JG96, JG98, NFJ96, SM90, TNS98]. Coupling [CLD97, HM99, Kla98, LCP98]. course [Ano95d, Ano95-54, Ano97-67, Ano97-87, Ano99e, Ano99d, Ano97-33]. covariant [Gre95]. cover [LOP95]. coverage [FF92]. Cowles [HH93]. Cox [Ano97-121]. Coyne [Ano96-52]. crack [DPD98, DD99, Ioa91, MB92]. Craig [Ano93v, Ano96-147]. Crandall [Ano96-121, Ano96-165]. Crank [CCVV90, EGLN95, Sir99]. CRC [Ano95w, Ano95z, Ano95g, Ano98-43, Ano98-107, Ano99-108, Ano99-125, Ano99-126]. Crease [Ano99-146]. Creating [Ano98-82]. Creative [Ano96-47]. creativity [Ano95-28]. Credibility [Ano99-51]. credit [Kra90]. creeping [Kro99]. Cricket [Ano95-46, Ano99-221]. crime [Ano96-42]. crimefighter [Ano96-42]. Crisp [Köh94]. Cristian [Ano99-98]. Criteria [AdMD99, Ano99-243, BIZ97, He95b, Ill91, LL90a, LC98b, LY98, LF99, LC96a, LF98, ST98, WA96b, Zha98, ZT98b, ZL99b, ZZ99c, ZT99]. criterion [EL99, LO99, Ohn92, SKI92]. Critical [FWI99, Ano99-140, Kai96, Mol93, Ano98-138]. critique [Ano96a]. Crosbie [Ano99-225]. Cross [BW93, Ano95-72, WY93, Ano95y, Ano99-50]. cross-GUI [Ano95-72]. Cross-Language [Ano99-50]. Cross-level [Ano95y]. cross-product [WY93]. crossing [KKD^{+}97]. crossover [LKE97]. Crouzeix [Ano99-95]. Cryptanalysis [CFLC95]. Cryptographer [Rea97]. cryptographic [Ano96-136, Hwa99, LWL93, YL93a]. cryptography [Ano98-86, Ano93g, Ano95z]. cryptoschemes [LK97]. cryptosystem [Lia99a]. cryptosystems [SH94]. crystals [BGZ92]. CSLI [Ano97n, Ano97-148, Ano97-117, Ano97-162, Ano97-129, Ano98-70]. CSP [TT92]. CSP-1 [TT92]. Ct [Ano97-141]. Cubature [Web95]. cube [CK97a, LM96a, ZM96]. cubic [CCGL99, FG93, Gal91a, Gal93a, Gal93b, Gal93c, LPR96, MN97, MMP98, MTLGFG95a, PM95a, RSS94a, Ung90, Ung91]. cuboidal [Kar94]. cuboids [NB95, NB98]. Cucker [Ano97-51]. cultural [Ano99-225, Ano99-51]. culture [Ano96-152, Ano96-108, Ano96-141]. Cummins [Ano96-139, Ano98-41]. Cunningham [Ano96-59, Ano98-29]. Curacao [Ano98-140]. cure [Ano97-146]. Curien [Ano99-68]. curl [HS97]. Current [Fli94b, Bir98, BS95]. Curtis [Ano96-151, Ano96-153]. curvature [FM96, HGK93, Mil91, NG91]. Curvatures [AK98]. curve [Kou90a, Qu96]. curved [Ano97-59, BM91, Has95, Luo98, RHRZ90]. curves [AK98, Ano95-55, Ano98-44, Ano98-107, EL92, ISM93, KKM96, KM97b, KS95, LK97, MC97, PM95a, PC94a, YJ99, Ano98y, Ano99-153]. curvilinear [Gla93d, Gla95e]. customers [Kat92]. cut [RR97b]. Cutcliffe [Ano98-145, Ano98-146]. cutting [WF99b]. cyberspace [Ano96-85]. cycle [BBKN97, BM89, BM90c]. Cycles [Pal97, Ano96x, GF90, Kri97, MN97, Nis97b, Ran93, Ves98]. cyclic [AAD98, Amo93, JLDL92, Kat92, SK91]. cyclic-service [Kat92]. cylinders [ASW98, PSD96]. cylindrical [KMD99, LYL^{+}91, LCL94, LT94, MP94]. Cylindrically [Gla93b]. Czech [Ano99-128].

D [Ano90u, Ano93h, Ano93o, Ano95-57, Ano96w, Ano96-56, Ano96o, Ano97u, Ano97-30, Ano97-41, Ano98r, Ano98-28, Ano98-64, Ano98-85, Ano98-140, Ano99e, Ano99-83, Ano99-136, Ano99-138, Ano99-190, Ano99-191, Ano99l, Ano99-226, CC97, GCLT98, RD99, SJB97, BMN96, BBO99, BGL97, BS95, BRGJ95, Bri97, Dea98, HTT92, HGG99, KKS98, Kle93, Kor97, Li98d, LM96a, Mah90, MM93a, Mol96, MZ95, Pou90a, RW91, SJB97, ZIKM93, ZM96]. D-IHCP [MZ95, ZM96]. D/G/K [Pou90a]. DAEs [Bal96b, CC94b, MC99, Rhe97, RS99c]. Dagli [Ano95l]. Dale [Ano95-80]. d’Alembert [JG98]. Dalheimer [Ano99-195]. Daly [Ano95f]. Dam [Ano95-57]. damped [Kuu94, LF99]. damping [GL93, KL96a]. Dan [Ano96-29, Ano99-192]. Dana [Ano97k]. dance [MM96d]. Daniel [Ano95x, Ano96u, Ano96-154, Ano96-160, Ano98k, Ano98-48, Ano98f, Ano99-92, Ano99-224]. Daniele [Ano96-90]. Dao [Ano98o]. Darboux [Leb98]. Darcy [KP99a]. Dario [Ano96-129]. Darrell [Ano98-164]. Darren [Ano96-156, Ano96-159, Ano97-147, Ano99-218, Ano99-219]. data [Ano96h, Ano96y, Ano96-120, Ano97-136, Ano98-32, Ano98-72, Ano98-139, Ano99v, Ano99w, Ano99-80, Ano99-94, Ano99-99, Ano99-100, Ano99-193, Ano99-194, BKM91, BIZ97, Bro94, CF92, Cha91, CF94, DF98, FBL99, Fol92, FLNW98, GRZ96, GB92, JCL93, Kan90a, Kan90b, Kou95, LD90, Los89a, MCG98c, Mil91, Pio94, QA96, Ram91c, Sad96a, Sen92, YNK92, Ano97-33, Ano97-34, Ano99-53, Ano99-55, Ano99-54, Ano99-56, Ano99-52]. database [Ano95-61, Ano97-163, Ano98-132, Ano99m, Ano99-94, Ano99-125, Ano99-126, Ano99-176, FO93, FKO92, NJ92, ZO92, Ano96-48, Ano97-35, Ano99-57]. databases [Ano98-76, Ano98-158, Ano99-88, Ano99-145, Gün99, MS91, Yan91, Ano95-35, Ano98-45, Ano98-109]. Datalog [BHK95]. Datamodel [DT93]. Date [Ano95-61]. dates [Che90b]. Dave [Ano95-36]. David [Ano95-52, Ano95-57, Ano96g, Ano96i, Ano96f, Ano96l, Ano96-42, Ano96-67, Ano96-76, Ano96-78, Ano96-92, Ano96-93, Ano96-98, Ano96-102, Ano96-118, Ano96-155, Ano96-141, Ano96-149, Ano96-143, Ano96-163, Ano97g, Ano97r, Ano97-37, Ano97-57, Ano97-72, Ano97-67, Ano97-68, Ano97-151, Ano97-121, Ano97-132, Ano97-168, Ano98-148, Ano98-56, Ano98-99, Ano98-122, Ano98-125, Ano98-152, Ano99q, Ano99-46, Ano99-158, Ano99-159, Ano99-176, Ano99-190, Ano99-191, Ano99-185, Ano99-231]. Davida [HH93]. Davidson [Ano97-50]. Davis [Ano95-41]. Davison [Ano95-40]. Dawes [Ano99-178]. Day [Con95]. Dayanand [Ano99-28]. DC [Ano95-67, Ano96-29, Ano97-37, Ano97-68, Ano97-79, Ano97-139, Ano97-161, Ano99e]. DCE [Ano96-49]. DCT [SHK99]. Deadline [Ano99-58]. deadlines [CD98]. deadlock [MG93]. Dealing [Ano98-39]. Dear [Ano97-154]. death [Ano99-239, PL99]. DeBevoise [Ano98-134]. Deborah [Ano96-64]. Debra [Ano99-92]. debugging [Ano99-86]. Decay [Duc95a, Duc95b, WAL97, KKD^{+}97]. decaying [BGM97, KKD^{+}97]. decentralized [KS99, WW93]. Decision [AK93, Ano96-152, Ano98-27, Ano99-53, Ano99-54, CL99a, Eke99, Eso91, HGWL96, HL90, Jay90, KSK90, Kur91, KK94, Lew91a, Lew91b, LH94, Li99c, LL90c, MC93, Ohn92, TDF^{+}92, Ano95-27]. decision-support [HL90, TDF^{+}92]. decisions [Mil92]. Declarative [Bol92, Yan91]. Deco [Ano96m]. decoding [HH93, WJ90]. decomposable [Sjo92]. Decomposed [Tor90]. decomposing [Luo93d]. decomposition [AC95, AR90a, Ado91a, ARM91, Ado91b, AR92a, AR92c, Ado94, Ado96, Ano97-41, BS93, BS95, BS96b, BS96c, Bog98, Bog99, Bon99, CP95b, CSOL96, DG99, Flo95, Gab94a, GPK99, JT99, LCP98, LC97b, LT99, LF91, Luo93c, LURL98, NGNM97, PC92, RAM92, Sir97, Tsa90, VR95a, WHL^{+}99, Yan97, ZC91]. decompositional [RB97]. decompositions [Ano95-33, CGS99, JAK96, KP98, PS95, Sto99]. deconvolution [CL99a]. decoupling [LS99]. decreasing [CD98, RJ94, Won98a]. Dedicated [Ano93h]. Dedication [Ano90y, Ano94a]. Deeba [Ano99-129, Ano99-130]. Defeasible [Pad92a]. defect [LCP98]. defects [Ano99-137]. deferred [BACS98]. deficiency [Pan98a]. deficient [MPR98]. defined [AIS95b, Bad95, BR94, Bro98, FK98b, LS96a, LS98b, Rhe96, Yos90]. Definite [BRGJ95, LURL98]. Definite-clause [BRGJ95]. definiteness [LW99b, Mye92]. definitive [Ano96-88, Ano96-93, Ano98-85, Ano99y, Ano99-188]. deflection [MB94b]. deformation [ARI96, Ano99-136]. Deformations [Köh95, Köh96a, Köh96b, Köh93a, WW91]. Degree [KKM96, KM97b, EDO93, PC94a]. delamination [FCP95]. Delaunay [Bör90, LS96a, LS98b, Wea92]. delay [AD93, AdMD99, CYS94, CLY95, CL96a, CG94, CT96, HY98, Hen94, IESAH95, KV94a, KP99b, LY98, Pao98, ST98, SZ98b, SWY98, SC98a, Sta95, TY99a, TY99b, TSL96, YCC96, YQM99, Yan99b, Yu98, Zaf98b, Zha97, ZY98, ZT98a, ZC98, ZT98b, Zha99, ZT99, ZZ99a, ZZ99e]. delay-difference [Hen94]. delayed [Ame99]. delays [BGM95, DBA98, Ise94, SWY96b, SWY97, TC99, WA96b, XLPG99, ZC96]. deletion [MCG98b]. delivery [Ano98x, RR97b]. Dell’Amico [Ano98m]. Delta [Ano99-59]. demand [Ano98x, Ano98-159, Ano98-161, Ano99-131, DKO93]. demands [ZG97]. democracy [Ano95-75]. Dempster [LZ92]. Denis [Ano90q, Ano96-161]. Dennie [Ano96-127]. Dennis [Ano96y, Ano97-37, Ano99-199, Ano99-200]. dense [Fre98, Lin96, LC97b, MC97]. density [Ano98-84, GNR95, NR96a]. departure [Sie90]. dependability [Cse95]. Dependence [EDO93, NO93]. Dependencies [DT93, BKM91, DT95, ZO92]. Dependency [DRS91, Lyt92a]. Dependent [Liu97, Ame97a, Ame97b, Ano96-164, BP97, Bra92, CP96, CR90a, CR90b, Du94, Eti91, FK97, IIP^{+}97, JP95b, JA96, JD97, KNU99, KV94a, Kos90, Leo90, Mat99, SOI92, She92, TG96, VK95, WHL^{+}99, YOO92]. Dependent-chance [Liu97]. depending [Gal93d, Prn97, SKI92]. depletion [SHT^{+}90]. depth [YZM95]. Derek [Ano95-77, Ano98-31]. derivation [DOS93, LOS99, Nam99a]. derivative [Arg99, EDO96, GH90, Kan90a, VDLM95]. derivative-free [GH90]. derivatives [AW91, Doh91, Feu90, Gal93d, GS99, Has93c, WA94b, WA96c, Zad89, Zad92]. DERIVE [Ano96-98]. Deriving [MN97]. DeRusso [Ano98-141]. Descartes [Ano97-61, Ano96-50]. descendants [Lyt92a]. descent [Alt90c, CH99, dKSG94]. described [LZ98]. describing [SMU98]. Description [YX95, Ano98l, Ano99-227, Hol90]. descriptions [DT95, FMS92, TCS94]. descriptive [Zar92, Ano99-60]. descriptor [Tra97]. Desfray [Ano95-56]. Design [Ano97-126, CPC93, DH97, Ano95-52, Ano97-84, Ano98-34, Ano98-37, Ano98-69, Ano98-70, Ano98-119, Ano99v, Ano99w, Ano99m, Ano99-86, Bed95a, BF95, BS99b, CKK97a, Cho99, Chu99, FO93, GSM91, GCT98, HP96, Lia99b, LY92, MMMS97, NMUY92, PZ99a, RA92b, TCS94, YT98, Yav93, YL93b, ZO92, tC95, Ano95-72, Ano95-81, Ano96-27, Ano97-91, Ano98-127, Ano99-94, Ano99-122, Ano97-36]. Designing [Ano97-131, MMMS96, Ano96-52, Ano96-51, Ano97-37]. designs [SYZ94, Ano99-201]. desion [Ano99o]. desktop [Ano96-92, Ano97-169, Ano98-79, Ano98-103, Ano98-130, Ano99-31, Ano99-62, Ano99-143, Ano99-187, Ano99-210]. Desrochers [Ano98-141]. destruction [Ano96-96]. desynhcronized [ADFP93]. detecting [ABVB92, Jha90, WJ90]. detection [HP94c, Kle93, MG93, PBG90, Sou90, ZIKM93]. deterioration [KK92]. determinant [De 99, GM99, Mur91a]. determinants [DJM98, PYS97]. Determination [GP93, LCL94, BMD92, DJV96, LYL^{+}91, LKE97, UK99]. determine [Lia99b]. determined [Sun91]. determining [Bai96a, KKS98, LL92, Mil91]. determinism [Ano95-29]. Deterministic [Pan95, Sug94]. Deugo [Ano98-87]. developed [Ano97l]. developer [Ano98t, Ano98-163, Ano99-180]. Developing [Ano97-38, Ano98-41, Ano98-40, Ano99-61]. Development [Ame99, Ano98t, Ano99-35, Ano99-36, FCP95, Ano95-73, Ano95-34, Ano96c, Ano97l, Azb99, SD93, Ano90q, Ano98-102, Ano99-154, Ano99-155]. developments [DKG98]. deviance [Ano96-152]. Deviant [Ano97-39]. deviating [AG99b, KAD99, QA98, WA98]. deviation [ED91a, Kou90a, PT91]. deviations [CL94a]. device [Che93b, Sie90, Szm90]. devices [Che90a, Che96, Ano90t]. Deville [Ano97-110]. Devlin [Ano97-61]. Dewdney [Ano99j]. Dewilde [Ano99-229]. Dexter [Ano98q]. Dfl [Ano95-65, Ano96-83, Ano97e, Ano97-29, Ano97-36, Ano97-45, Ano97-58, Ano97-56, Ano97-66, Ano97-85, Ano97-84, Ano97-104, Ano97-126, Ano97-157, Ano97-164, Ano97-163, Ano98-71, Ano98-115, Ano98-121, Ano98-123, Ano98-137, Ano98-153, Ano98-160]. Diagnosing [YS90]. diagnosis [Eso99b, KKOS96, LYH98]. diagnostic [FGN90, LR90, NSS92]. diagonal [AS97, BG90, DBM92, EG97, EG98, ESIP98]. diagonalization [OB97]. diagram [Dwy93]. Diagrammatic [Ano96-53]. diagrams [Ano99-117]. diameter [She98]. Diane [Ano96-152]. diaphysial [MBMS92]. DiBona [Ano99-173]. dichotomically [DELE99]. dichotomies [Pin94]. dichotomy [AHO99, AG94]. Dictionaries [Ano96-58]. dictionary [Ano97-150]. Didier [Ano96-113]. Dieter [Ano97-136]. Dietmar [Ano99-89, Ano99-206, Ano99-207]. diffentiation [MG90b]. diffeomorphism [Ano97-157]. Difference [BK98, BGL90, KN94, PY94, AH98, AO99e, AP94, AHO99, AHL98, Ano96-41, Ano96-164, Ano97e, Ano99-84, Ano99-85, Ano99-243, AMZ98, Bal98, Bog98, Cas94, CHSS95, CP98a, CS93a, CE94, CLY95, CL96a, Che99b, CZ94, CL98, CL99d, CCVV90, CS94b, CC98b, CKRV98, DBM92, DH98b, DH98a, Dor90, Doš98, EGLN95, Far99, FT98, Gla90a, Gla91a, Gla91e, Gla92b, Gla93d, Gla93c, Gla93e, Goh94, GT99, Gre90a, Gre94, Gre98a, HP94a, Hen94, HK98, HNS^{+}90, HS98, Ise94, Ism97, JV96, JPF96, KP97a, KS97b, KS98b, Kwo92, Lal94, LY94, LC98a, LC98b, LY98, LF99, Li99a, LC96a, LFP97, MT94b, Med97, MR94, Pac94, PA98, Pao98, Pap94b, PXHG99, Pin98, Pop94, QZ93b, Rom98, SK91, Sha90a, SZ98b, SA94, SWY96a, SWY96b, SWY97, SWY98, Sta95, ST95, TY99a, TY99b]. difference [Tha94, TSL96, TP98, TMGS98, TR99, Tre95b, Tre98, Wen99, WA96a, WA96b, Won98a, WA98, Yan93d, Yu98, Zaf98a, Zaf98b, Zha97, ZY98, Zha98, ZT98a, ZC98, ZT98b, ZYC98, ZL98a, ZL98b, Zha99, ZL99a, ZL99b, ZZ99c, ZT99, ZZ99a, ZY96, ZYW98, ZZ99e, ZQ93d]. differences [CWJ92, Elo98, GMR90, MG90c]. different [ASW98, Che90b, Gla95c, Gla96a, OSLC98, RD96a]. differentiability [EB93]. differentiable [AD96, KK91]. Differential [Ano96-56, EDNO92, EDO96, AC94, AR90a, AR90b, Ado91d, Ado96, AWY97, AG99a, AG99b, AG99c, ACHP94, Ahm91, Ano96-41, Ano96-91, Ano97-41, Ano97-116, Ano98-107, Ano98-126, Ano98-125, Ano99-33, Ano99-34, Ano99-43, Ano99-71, Ano99-84, Ano99-85, Ano99-105, Ano99-103, Ano99-102, Ano99-104, Ano99-183, Ano99-184, Ano99-185, Ano99-182, Ano99-197, AG94, Azb99, Bal96a, BA96, CC97, CS96, Cas94, CP96, CJS90, CYS94, CS93b, CG94, CT96, CS90, CRV98, EDO97, Fab95, Fil93, GPZ93, GS91, GSC98, GL93, Gra94, GMS98, GPK99, GP90, Gyö90a, HBB90, HP94b, IESAH95, Il’97, JP95a, JP95b, JA96, JCM96, JD97, JPC99, JLDL92, Kan90b, KV94a, Ker91a, Ker91b, KS98a, KSM94, KP99b, KB90, Lee93, LYL95, LX98, LP99, LG99, LPR96, LOD96, LM91, LS95, LNG92, MT94a, Mär96, MPS94, Med90]. differential [ML90b, MP91, MK93, MB90, MW99, MT98, Moo94, MPR98, MRJ97, NPJ95, NFJ96, O’R98c, O’R99, PR90, QZ91, QZ93b, Riz95, RZ95, Rua91, Sar90, Sar94, Sha90d, ST98, Shi98, SLC97, SCB^{+}93, Sym91, Tör94, Tre95a, WCC99, Wan96, YQM99, Yan93d, Yav93, ZZ99b, ZQ93b, Ano96-54, Ano98-42]. differential-algebraic [Bal96a, MPR98]. differential-difference [Yan93d]. differentiated [Doh91]. Differentiation [Mar95, Ano99-244, BB99, CFM91, MBV91, MG90a, MMZ98, VBM90]. diffpack [Ano99-43]. diffuse [WU94]. diffusion [Ado95, BBO99, BGL90, Boy96, BM89, BM90c, CAZ99, Che92, CCVV90, CC98a, Cho92, CC98c, DKSI94, GN99, Gol90, GW91, Har96, JG95, JG96, KP95a, Kwo92, LN98, Li10, MS96, MB94a, Mil90, NN91, Pao98, Par91b, RM98a, Ram91a, SS98a, Sha90d, SC98a, ST95, VK99, Wu90a, vNvN90a, Ano96-91]. diffusion-Brusselator [Ado95]. diffusion-reaction [CCVV90, MB94a, Ram91a]. diffusive [YCC96]. diffusivity [MM95a]. digamma [KS96]. digital [Ano98-137, Ano99-106, Ano99-107, SZC91, Ano96-55, Ano98-43]. Digraphs [CT97, Ano97-62]. dihedral [Arn90, MTLGFG95b]. dihydrofolate [SHT^{+}90]. Dijkstra [Ano99-83]. dilemma [Ano95-29]. Dillon [Ano98-50]. dimension [Ano95-56, CHSS95, MAM91]. dimensional [Ano99-231, Ano00, ANV94, BSS98, BBI97, BR94, Bro98, CC94a, CST97, CCD99, DKR97, DK97c, Dri97, FG93, Fin99, GF94a, Gla91a, Gla92b, Gla94d, Gla95a, Gla95d, Gla95f, Gla98, Gla99, GB90, GB92, GB96, GM96, GT99, GL96, Gre91b, GC99, GK91, HM93a, Hin95, HNS^{+}90, IK97, JKK98, KCDV99, Kha97, KS95, Köh95, Köh96a, Köh96b, Li96b, Lin95, MMZ91, MRZ96, MM95b, NN91, SSZ91, ST95, Tad94, TE93, Wan91, WF99a, Wen99, XLG95, ZM98]. dimensionality [Lew91b, Sok91]. Dimensions [DO91, McK96, Ano99-38, DM95, EA90, GG97, GGG^{+}98, Kel91, Kre91, PW97, Ano95-28]. Dimiter [Ano98-56]. Dimitris [Ano98-161]. Dionysios [Ano96-119]. Diophantine [TCS93]. dioxide [KG97]. Dirac [AR99a, Fau94]. Direct [CKRV98, KP98, PDK96, APP96, Ano99-246, ARS98, Cas91, CO99, Gla96c, Kou98, KB90, Nov90, Nyí99, OH93, PN98]. directed [Nis97b]. direction [CS90, LW91, MG95, Sny90, Wac94].

Directions [Ano96-35, Alt90a, Ano97-28, Heg91a, Heg91b, SS90, Ano97r]. Director [Ano99-62]. Dirichlet [AO99f, CC97, She93, YL99]. disaggregation [MM96b]. disagree [Ano98-39]. discipline [AD93, Ano95e, HY98]. Disciplining [Ano97-40]. Disconjugacy [HP94a, MW99, AP94, Che99b]. disconjugate [Doš98]. discontinuities [Ech95, Ech98]. Discontinuity [Ano98-62, Zhe93, HNS^{+}90, Ioa91]. Discontinuous [Cub95a, BBO99, ZZ99d]. discounted [KK94]. discounting [VMA99]. discourses [Ano95-74]. Discovering [Ano99j, Ano98-44]. discovery [Ano95-38, Ano96h, Ano98-139, Har90, Har92a, Hen97, Ano99-40, Ano99-52]. Discrete [AL93, AO98a, BMM94, Boh96, Boh98, CK97b, EDNO92, MG90a, MG90b, MMZ98, Pin94, PA99, SD95, Yam94, AO97, AO98b, ABW99, ACHP94, Ano97-70, Ano99-135, Ano99l, BB93, BG95, Cha91, CTG96, Fli95, FG97, GMS98, HI94, HI98, Har96, Hsu96, Huj99, JS97b, KM92, KK92, KSK90, KP99a, KSM94, Lee94, LA96, LTR98, Lew94, LB94, Li98d, LF98, LCC99, MM95a, MM96c, Mic94, Mic98, Mil95, OCS96, OSLC98, Oll90, Özy96, PC96a, PR90, PA94, Sar94, Sto99, Tch98, TD99, Wan91, Wan98a, Wan98b, Wu90b, XLPG99, ZM99, ZCC96, Ano99-63, Ano99-64, Ano99-65]. Discrete-time [CK97b, BG95, CTG96, KM92, LB94]. Discretization [AG94, Bab98, LTB96, Bal99, BRV98, Bon99, DK93, Fai98, KHI94, Lay93, Rac99, Shi94, SS91b]. discretizations [DSKP96, EL96, HS97]. discretized [Dia91]. discretum [Nas95]. discriminant [Ahm94, Ahm95]. discriminate [BH90]. discrimination [BJH90]. disk [CC94c, GM99, SKI92]. diskette [Ano95-51, Ano95-80, Ano95-50, Ano95-49, Ano96q, Ano96-112, Ano96-122, Ano96-159, Ano96-149, Ano97w, Ano97-30, Ano97-72, Ano97-73, Ano97-78, Ano97-145, Ano97-79, Ano97-92, Ano97-96, Ano97-147, Ano97-113, Ano98-44, Ano98-46, Ano98-57, Ano98-80, Ano98-143, Ano98-160, Ano99-147, Ano99-177, Ano99-179]. diskettes [Ano96-147, Ano97-67, Ano99-53, Ano99-54]. disks [Ano96-120]. dispersion [KKK^{+}97]. dispersive [Sha90a]. displacements [KMD99]. dissimilar [DPD98]. dissipation [Alk94, XL97]. dissipative [XLG95]. Distance [CT97, Jar97, LYL95, MPS93, PL90]. distinct [CD98]. DISTO [BFF^{+}97]. Distributed [MR99, APP96, AGM99, Ano97-32, Ano97i, Ano97-126, Ano98-56, Ano99-163, Ano99-177, AS95, AS98, BGM95, BP94b, CHCC95, CCY95a, CCY95b, CCH97, CDLW95, FO93, KKOS96, Lag91, Nke94, Pou90b, Rei94, Rei99a, Ano98-45]. distributed-memory [CDLW95]. Distribution [Pan93b, AKS96, BGM95, CKK97a, ED91a, ED91b, GW97, GS96c, Kou90a, Kov95, MM96e, MH93b, Nar92, Niw96, Pap94a, RR90, SV91a, SV91b, Tem93, WY93, Yan93c]. distributions [Ahm94, Dev93, Fau94, Gab94a, Pan93b, RL91, ZG97]. disturbance [SC98b, WSS92]. disturbances [LL99]. divergence [HS97]. diverging [MP93]. diverse [Ano99-209]. diversity [Ano99-144]. divided [HH98b]. divisible [BLR96]. division [BCB90, Chr90, FL91, HP94c, SRCB98, Tan92]. Djurdjica [Ano98-126]. DKT [AO99a, AO99b]. Dlf [Ano97-108, Ano97-105]. DM [Ano96-68, Ano96q, Ano97-140, Ano97m, Ano97w, Ano97t, Ano97-33, Ano97-34, Ano97-46, Ano97-51, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97j, Ano97-86, Ano97-92, Ano97-96, Ano97-147, Ano97-111, Ano97-112, Ano97-122, Ano97-152, Ano97-138, Ano98-44, Ano98-43, Ano98-91, Ano98-125, Ano98-62, Ano99v, Ano99w, Ano99x, Ano99-37, Ano99-41, Ano99-42, Ano99-44, Ano99-43, Ano99-46, Ano99-66, Ano99-55, Ano99-67, Ano99-56, Ano99-69, Ano99-60, Ano99-78, Ano99-79, Ano99-77, Ano99-70, Ano99-89, Ano99-82, Ano99-109, Ano99-110, Ano99-108, Ano99-111, Ano99-112, Ano99-125, Ano99-129, Ano99-126, Ano99-130, Ano99-132, Ano99-218, Ano99-156, Ano99-219, Ano99-157, Ano99-162, Ano99-193, Ano99-194, Ano99-186, Ano99-189, Ano99-198, Ano99-208]. DMIX [Opp96]. DNA [HSH90a, HSH90b, Jos90, Ano99-66, Ano99-67]. DNS [Ano99-221]. Document [Ano95f, Ano95-42]. Documentation [Ano98-162]. Dodson [Ano96-68]. Doeblin [Kur91]. does [Ano95-76, Ano95-67, Ano95-29]. Doherty [Ano97-117]. doing [Ano96-144]. Domain [Bon99, DG99, XLPG99, Ano96t, BS93, BS95, BS96b, BS96c, Bog98, Bog99, Che97, GPK99, KP98, KL96a, LCP98, Riz95, RZ95, Sir97, YOO92, Yan97, Ano97-41]. domains [Bör90, EDO98, HSH90b, JV96, Kar94, Kri97, LK99, LM96a, Ano99-68]. Domik [Ano96o]. dominant [DF94]. domination [SMT98]. Dominik [Ano98-133]. Donald [Ano96-57, Ano96-94, Ano97-129, Ano99-53, Ano99-54, Ano99-216, Ano99-125, Ano99-217, Ano99-126, Ano99-227, Ano99-239]. Dongarra [Ano95-59, Ano96-118, Ano99-158, Ano99-159]. Donnie [Ano98-130, Ano99-187]. door [CB93a]. Dordrecht [Ano93m, Ano93q, Ano97e, Ano97-29, Ano97-45, Ano97-58, Ano97-56, Ano97-66, Ano97-85, Ano97-84, Ano97-108, Ano97-157, Ano97-164, Ano98o, Ano98j, Ano98i, Ano98n, Ano98w, Ano98r, Ano98-32, Ano98-36, Ano98-49, Ano98-53, Ano98-59, Ano98-57, Ano98-66, Ano98-67, Ano98-84, Ano98-74, Ano98-83, Ano98-89, Ano98-90, Ano98-92, Ano98-108, Ano98-100, Ano98-101, Ano98-115, Ano98-111, Ano98-116, Ano98-121, Ano98-123, Ano98-126, Ano98-127, Ano98-135, Ano98-140, Ano98-153, Ano98-157, Ano98-155, Ano99z, Ano99-27, Ano99r, Ano99s, Ano99n, Ano99t, Ano99u, Ano99-33, Ano99-49, Ano99-63, Ano99-64, Ano99-84, Ano99-85, Ano99-83, Ano99-105, Ano99-103, Ano99-96, Ano99-102, Ano99-104, Ano99-97, Ano99-95, Ano99-101, Ano99-115, Ano99-116, Ano99-114, Ano99-136, Ano99-137, Ano99-152, Ano99-153, Ano99-171, Ano99-172, Ano99-170, Ano99-174, Ano99-183, Ano99-184, Ano99-181]. Dordrecht [Ano99-182, Ano99-196, Ano99-197, Ano99l, Ano99-206, Ano99-207, Ano99-204, Ano99-234]. Dordrecht/Bratislava [Ano98-74]. Dordrect [Ano98-65]. Doru [Ano99-189]. Double [WA96a, YL99, CKAS96, Gil91, Kel94]. double-stride [CKAS96]. Doubly [CGS99, Che99a]. Doug [Ano99-238]. Dougherty [Ano95-80]. Douglas [Ano95z, Ano96r, Ano98-114, Ano99e, Ano99-222, Ano99-223]. Dowla [Ano96-145]. Dowler [Ano97-55]. Downing [Ano97-78]. downwash [RR90]. Dragan [Ano96m]. Drastal [Ano95r, Ano95s]. Drawing [HIIN99]. Drescher [Ano96-120]. dressed [KKD^{+}97]. Drew [Ano98-152]. Drift [Har96]. drive [Ano95-29]. driven [CHW93, HW90, Kuu94, Ano96-78]. drop [Kor97]. droplet [RS97]. drops [ARI96]. drugs [KM93]. drum [Log96]. DSMC [Bir98]. Dual [LZ94, RG90a, Bar94, LKT99, Pan98b, SWF95]. Duality [BJH99, Kim92, Lin91b, Liu96a, BGO95, Che97, Liu96b, LY99b, Ano99-151]. Dubinsky [Ano97-70]. Dubois [Ano96-113]. Dubrovin [DKR97]. duct [Che91, Gla92d]. Due [Chi93, Che90b, KL96b, MBMS92, PR91]. due-dates [Che90b]. Dukas [Ano98-27]. Dunham [Ano99-73]. Dupuit [Ano99-203]. Durban [Ano99-136]. during [Gil93].

Durnbaugh [Ano96-77]. Durrmeyer

[AdlC95, GZ95]. Dusan [Ano99-49]. dwarfs [Ano99-91]. Dwight [Ano98-87]. Dye [Ano99-177]. Dykstra [ER98]. Dym [DK97c]. Dynamic [Bec90, Che90b, Cho99, HM94, KNU99, LH94, LH99, LO99, Nai92, Oda94, VV99, Yan97, AH92, AL94, Ano96c, Ano96-140, Ano98-105, CK97a, CL96b, EA90, Eso99a, Eso99b, Fli94b, Flo95, HJ91, Hwa99, ISM93, Iwa91, JW93, KSK90, KPB97, KYNY99, LRLR95, LK94, LS98a, LH91, LB94, LWL93, Lin94, Mar94, ML90b, MR99, Özy96, RG90a, RT94, She94, SFY92, TG95, Yas94, YL93b, Yos96, Yos99, Zad89, Zad92, Ano97-43, Ano97-42, Ano99-69]. Dynamical [Gre98a, Gre98b, Mur94, PS90, Abe90, BAS97, Boh96, Cor94, DSKP96, Gre93a, Shi94, Wit96b, Yav96, Ano96-56]. dynamically [BW92]. Dynamics [AKK^{+}97, IIP^{+}97, RS93, APV97, Ano96r, Ano96-130, Ano97-108, Ano97-107, ADKM97, AH97, BJ90, CC95, CCW96, FT97a, Fre98, Gil91, Gla90b, Gyö90b, HPH^{+}94, Kan90a, Kan90b, KP95b, KC91, LQ95a, MPP97, Mar91, MPS94, MDVK97, Mil90, Mro96, Nas92, PP97, RSP94, RS94, Wea92, Wen99, ZZ94, Ano98r, Ano98-113, Ano99-83, Ano95-30, Ano98-46].

E. [Los92]. earliness [Cha99a, WF96]. earliness/tardiness [WF96]. early [Ano96-158, Ano97-154]. earth [Ano98-60, Har92b]. easy [WJ90]. Eberlein [Ano97-59]. ECd [VAB91]. Eckstein [Ano99-237]. ECO [CGHS92]. ecological [Ano96-130, Ano99-233, CC99, LL94a, LL99]. ecology [Ano98-27]. econometric [Ano98-131, Gan92, Pad92b]. economic [Ano95-27, Ano97-49, Erm92, GG92, Pan92b, Pet92]. economics [AH92, Ano97l, Ano97w, Ano98-148, Ano97-69]. economies [Ano97-42]. economy [Ano97l, Ano98-151]. ecosystem [Ano96-96]. Ed [Ano97m, Ano97-70]. EDAUED [VAB91].

Eddy [BS95]. EDF [Ano99-58]. Edgar [Ano97-147, Ano98-75]. Edge [Jar97, JAK96, FCP95, MB92]. Edited [Ano90t, Ano93u, Ano93h, Ano93j, Ano93p, Ano93q, Ano93v, Ano93w, Ano95k, Ano95-36, Ano95h, Ano95m, Ano95l, Ano95r, Ano95s, Ano95u, Ano95x, Ano95w, Ano95-28, Ano95-29, Ano95-77, Ano95-40, Ano95-45, Ano95-41, Ano95-65, Ano95-67, Ano95-81, Ano95-69, Ano95-83, Ano96i, Ano96e, Ano96h, Ano96j, Ano96v, Ano96-32, Ano96-39, Ano96-44, Ano96-36, Ano96-151, Ano96-43, Ano96-34, Ano96-31, Ano96-35, Ano96-38, Ano96-40, Ano96-50, Ano96-53, Ano96-67, Ano96-78, Ano96-86, Ano96-79, Ano96-82, Ano96-81, Ano96-80, Ano96-83, Ano96-89, Ano96-104, Ano96-103, Ano96-114, Ano96-113, Ano96-115, Ano96-123, Ano96-121, Ano96-134, Ano96-128, Ano96-127, Ano96-130, Ano96-148, Ano96-144, Ano96-141, Ano97g, Ano97f, Ano97-53, Ano97n, Ano97r, Ano97-28, Ano97w, Ano97v, Ano97-40, Ano97-45, Ano97-54, Ano97-47, Ano97-51, Ano97-50]. Edited [Ano97-49, Ano97-57, Ano97-55, Ano97-56, Ano97-64, Ano97-65, Ano97-68, Ano97-69, Ano97-81, Ano97-82, Ano97-88, Ano97-149, Ano97-90, Ano97-100, Ano97-106, Ano97-117, Ano97-118, Ano97-135, Ano97-128, Ano97-154, Ano98g, Ano98m, Ano98h, Ano98i, Ano98l, Ano98r, Ano98v, Ano98-34, Ano98-28, Ano98-27, Ano98-40, Ano98-45, Ano98-49, Ano98-148, Ano98-47, Ano98-58, Ano98-56, Ano98-64, Ano98-65, Ano98-66, Ano98-70, Ano98-77, Ano98-82, Ano98-84, Ano98-83, Ano98-87, Ano98-96, Ano98-94, Ano98-93, Ano98-95, Ano98-91, Ano98-92, Ano98-102, Ano98-108, Ano98-101, Ano98-110, Ano98-104, Ano98-111, Ano98-112, Ano98-114, Ano98-121, Ano98-123, Ano98-127, Ano98-150, Ano98-136, Ano98-132, Ano98-135, Ano98-138, Ano98-139, Ano98-140, Ano98-145, Ano98-157, Ano98-146, Ano98-155, Ano98-156, Ano98-161, Ano99-29, Ano99-30, Ano99-90, Ano99o, Ano99r, Ano99s, Ano99n, Ano99t]. Edited [Ano99x, Ano99u, Ano99q, Ano99p, Ano99-50, Ano99-40, Ano99-45, Ano99-74, Ano99-92, Ano99-83, Ano99-81, Ano99-80, Ano99-214, Ano99-95, Ano99-118, Ano99-119, Ano99-113, Ano99-120, Ano99-128, Ano99-136, Ano99-137, Ano99-122, Ano99-145, Ano99-152, Ano99-163, Ano99-165, Ano99-169, Ano99-172, Ano99-220, Ano99-173, Ano99-186, Ano99-182, Ano99-212, Ano99-211, Ano99-232, Ano99-91]. Edition [Ano96-93, Ano98-46, Ano98-143, Ano93n, Ano95-71, Ano95-32, Ano95f, Ano95-54, Ano96f, Ano96y, Ano96-45, Ano96-60, Ano96q, Ano96-154, Ano96-102, Ano96-109, Ano96-133, Ano97x, Ano97y, Ano97z, Ano97-62, Ano97-71, Ano97j, Ano97-86, Ano97-95, Ano97-93, Ano97-150, Ano97-137, Ano97-158, Ano97-153, Ano98-35, Ano98-30, Ano98-54, Ano98-81, Ano98-85, Ano98-106, Ano98-107, Ano98-104, Ano98-141, Ano99y, Ano99-28, Ano99m, Ano99-108, Ano99-158, Ano99-159, Ano99-150, Ano99-221, Ano99-179, Ano99-187, Ano99-224, Ano99-227, Ano99-237]. Editor [Wit90]. Editorial [Ano91i, Ano93y, Ano97-170, Ano97-171, Rod93b, Ano92k, Ano92l, Ano92m, Ano93-35, Rod91a, Rod91b, Rod92a, Rod92b, Rod92c, Rod92d, Rod92e, Rod92f, Rod92g, Rod93c, Rod93d, Rod93e, Rod93f, Rod93g, Rod93h, Rod93i, Rod93j, Rod93k, Ano90z, Ano90-27, Ano91j, Ano91k, Ano92e, Ano92f, Ano93x, Ano94b, Ano95-84, Ano95-85, Ano96-168, Ano96-169, Ano97-172, Ano97-173, Ano98-166, Ano98-167, Ano99-240, Ano99-241]. Editors [RGBK91]. EDs [VAB91]. educated [Ano97-143]. Education [Ano96-127, BGEZ90, Gal99c, Ano96-89]. Edward [Ano95-44, Ano96-44, Ano97-146, Ano97-101].

Effect [MB92, CC98c, GL94, IK97, MPS93, Tso91, ZO92]. Effective [KLMP97, TPF98]. Effects [KKK^{+}97, Lag91, ML90b, CC99, KKD^{+}97, KL96b, MDVK97, SHT^{+}90, YC92]. Efficiency [HS94, Mav93, Ahm95, CC94a, Gla95b, Gla96a, JMPS94, Rán92]. Efficient [AR99b, CPC93, CK97a, Käm94a, LYH98, LC97b, NNB97, Rei93b, Rei98, Rei99b, SM98, BGM95, BS96a, DdW93, FGN90, Gla91a, Gla92a, Gla94b, Gla94d, Gla94c, Gla94e, Gla95d, Gla96b, Gla97b, HL96, KKS98, LT96, LL92, SS96, TG95, WSDT94, Ano99-70]. Egan [Ano97-143]. Ehrich [Ano97-136]. eigen [LURL98]. eigen-decomposition [LURL98]. eigenfunctions [ABW99, BK98]. eigenpair [Zha93c]. eigenpolynomials [BB93]. eigenvalue [AH98, Ame98, BKK99, BMD92, Boh98, Bou99, BG91a, BG91b, CCCY94, CY93, GC99, Mey99, Nan90, Pom91, Pom92, RS92, RSR93, Zha93d, vdVVdB99a]. Eigenvalues [ABW99, LLR^{+}90, CS93a, Con95, DR96, GSM91, HP99, Pan90a, RSAR95, RRS99, SBD99]. eight [MN97]. Eighth [Ano98-143, Ano99-120]. Einstein [Ano95i, Ano96-57]. Ekeland [ABK97]. Ekelund [Ano99-203]. elastic [Ara99a, BR95a, Ber98, CR90b, DD99, Dem94b, ESKG98, Kel91, MB92, MB94b, NSD98, Pok97, RS99c, She92, TAD91, XL97]. electric [ARI96, Ano96-58]. electrical [Ano90u, Ano99-190, Ano99-191]. electrically [EL96, LLP94]. electroacoustic [Sco96]. electrodynamics [KKK^{+}97]. Electromagnetic [BU90, Duc95a, KKB97, MRZ96]. Electron [Gre93b, Gre99, BKK97]. electron-hole [BKK97]. electron-positron [BKK97]. Electronic [Ano96-116, Dre90, Ano95-75, Ano96-85, Ano98-155, Ano96-59, Ano98-47]. electronics [Pen96]. Eleftheriadis [Ano98-161]. element [Ano96-83, Ano98-57, Ano99-205, BBO99, Bed95a, Bla95, BS96d, Che93b, Che96, Che91, Coo97, Dav96, DG99, DK97a, Dem94a, Dem94b, Du94, Fai98, Far91a, Far91b, GN99, GHS91, HM99, IB95, Joh99, Kar96a, KP99a, KPS96, Lai97, LLP94, Lay99, Li97, Li98a, Li98b, LN98, LN99, Li10, Luo98, OPD97, PSS99, PR91, Ram90a, SS98a, Sem93, SFT99, Shi98, Sol97, Ye99, ZS93, vNvN90a]. element-by-element [SFT99]. Elementary [Gyi96, Leb98, Pas97, Ano99-71]. elements [Arn90, BL94, Che90a, DO91, GF97, HSH90b, HAO91, Nis97b, PW97, RHRZ90, Sok91, Tad93, Wac91b, Wac91a, Ano96-60, Ano97-44, Ano98-48]. Eles [Ano98-144]. eleventh [Ano95-45]. Elias [Ano99-129, Ano99-130]. Elimination [MM93c, MM94, GHWL96, Mil95, MM95d, ND95, RZ91, Zou91]. elimination-based [MM95d]. ELISA [AGM99]. Elizabeth [Ano96w, Ano96o]. Ellen [Ano97-79]. Elliotte [Ano97-76]. Elliptic [Has97, AO93b, AO93a, Ano97-41, ASW98, Bed95b, CLM99, Cho94, Ech95, Ech98, FS96, HM93a, Hin95, JLDL92, KT96, Kan90b, Kor96, LK97, Li96b, Li97, Li98a, Li98b, LN98, LN99, Li10, PDK96, Pom91, Pom92, RT99a, RSAR95, Sib98, Sir97, VPK99, Wan99, YJ99]. elliptic-type [KT96]. Elliptically [Sza90b, Sza90a]. Ellis [Ano96-125]. Elmagarmid [Ano97-163]. Elmér [Ano99-181]. Elsevier [Ano95-65, Ano96-31, Ano96-79, Ano96-82, Ano96-81, Ano96-80, Ano96-83, Ano97-64, Ano97-65, Ano99-113]. Elsevier/MIT [Ano96-81, Ano96-80, Ano97-64]. EM- [Sch92b]. Embedded [AS96b, Ano97-126, Ano99o, Ben99, CMR90, CT96, Sar94]. Emerson [Ano95o]. Emilio [Ano96-63]. Emily [Ano97-90]. emission [KKB97, ZFM96]. Emmanuel [Ano98-58]. emphasis [Ano96-91]. empirical [CC92, PT91, Yos90, Ano96-61]. empty [NB98]. Enabling [Ano98-47, Ano96-62]. enclosing [NB95]. enclosure [Ha90, HH98b, Wan98a]. encoding [WJ90]. encoding/decoding [WJ90]. Encyclopaedia [Ano98-49]. Encyclopedia [Ano97-48, Ano97-45]. End [Rok90, Ano97-61, FG98]. End-point [Rok90]. endemic [Voi90]. endpoint [KM97b]. Endre [Ano98-126]. Energy [DR96, Pol90, Ano99-225, CC97, Duc95a, Gue93, ISM93, Oli93, Tan94, TL94]. Energy-information [Pol90]. Engelbrecht [Ano97-108]. Engeln [Ano97-111, Ano97-112]. engine [Ano93u]. engineering [Ano96-137, Ano96-145, Ano97-36, Ano98-64, Ano98-116, Ano99-74, Kan92a, LC97a, RA92b, TM91, Ano93n, Ano93r, Ano95e, Ano95-56, Ano96-90, Ano97-52, Ano98-118, Ano98-50]. engineers [Ano97-93, Ano95-48, Ano96-112, Ano97-125, Ano98-141, Ano99-190, Ano99-191, Ano99-203]. England [Ano93f, Ano93g, Ano93t, Ano93u, Ano93i, Ano93j, Ano93l, Ano93v, Ano93s, Ano93w, Ano95-34, Ano95-50, Ano95-56, Ano96-48, Ano96-119, Ano96-120, Ano97-42, Ano97-125, Ano98m, Ano98-64, Ano99-71]. England/Stuttgart [Ano96-120]. Englander [Ano96-150, Ano97-38]. Englewood [Ano90p, Ano90q, Ano90s, Ano90x, Ano93k, Ano93n]. English [Ano98-73]. enhanced [LC97a]. enhancement [BKM95, GK98]. enhancements [MS96]. Enlarged [Ano98-46]. Enriched [ZS93]. Enrico [Ano95n, Ano95-54, Ano96-63]. Enrique [Ano99-95]. ensembles [Köh95, Köh96a, Köh96b, Köh99, PS93b]. Enskog [Fre98]. enterprise [Ano97-154, Ano95x, Ano99-72]. entropy [PS93a, RZ96]. enumerability [Ano96-35]. enumeration [AKV91, BV95, GMBS94]. envelopment [Sen92]. environment [Eke99, Opp96, SH96, ZC96]. environmental [Ano96-145, Ano98-39]. environments [Ano95-27, Ano96-119, KSK90, Szt92]. enzyme [BH90]. enzyme-kinetic [BH90]. Epcot [DdW93]. epidemic [ZC99]. Epstein [Ano97-63, Ano97-107, Ano99-62, Ano99-143, AB93]. Equality [AR90a, CF98, WT90b]. Equation [BW93, Ado95, AM95, AP94, AHL98, Ano95i, Ano97-148, Ano99-243, Ara95, AS96b, BM96, BL94, Bal99, BK91, BLP95, BGL97, BA96, CC97, CAZ99, CL99d, CCVV90, CC98b, CT96, DG91, DK97c, Dor90, Duc95b, ESKG98, EGLN95, FLM94, FH99, Far99, GT95, GT97a, GT97b, GN99, Goh94, GPCGV92, GMS98, GGC97, GHS91, GP90, HI94, HP94a, HH98a, HM93b, HNS^{+}90, HG98, IB95, JA97, KP95a, KC92b, Kre91, Kro99, Kwo92, LCP98, Lee94, LA96, LDD94, LQ95b, LC98b, LYL95, LCC99, LD93, LD95, LOD96, LM91, MMP98, MMZ91, MAM91, Maz90, Med97, Mol93, MGM92, NN91, O’R98a, Oli98, Pao98, Pap94b, PY94, Ram93, RV94, RS90, SVS98, Say99, Sch99, Sch94, SD95, Sha90b, Sha90d, Sha96b, She92, SWY97, Shi98, SLC97, SM94, SM95, Sim97, Sim98]. equation [SMU98, Str98, Tah91, TSLW98, TVZPG96, TY99b, TG96, Tse95, Ush94a, WHC90, Wan98b, Wu90b, Yu98, ZZ99e, vNvN90b]. equational [Jay90]. equations [AC94, AR90a, AR90b, Ado91a, Ado91b, Ado91d, Ado91c, Ado96, AP98a, AH98, AO99c, AO99e, AG99a, AG99b, AG99c, Ahm91, AHO99, Alt90a, Alt90b, Alt90c, Ame97a, Ame97b, Ano96-54, Ano96-56, Ano96-91, Ano97-116, Ano98-42, Ano98-126, Ano98-125, Ano98-165, Ano99-43, Ano99-183, Ano99-184, Ano99-185, Ano99-245, AF90, AG94, AMZ98, Azb99, BMN96, Bab98, BSW96, Bal96a, Bal98, Ber93, BGL90, BMB94, BG95, Boy96, BPLS96, BCC95, CS96, Cas94, CP96, CJS90, Che93a, Che93b, CE94, CYS94, CLY95, CL96a, Che99b, CZ94, CKM99, CS93b, CS94b, Cho94, Cho92, CL96b, CG94, CS90, CKRV98, CRV98, DH98b, DH98a, Dat90, DJV96, Din97b, DF94, DKR97, Doh90, DHV91, EDD91, EDNO92, EDO96, EDO97, ESIP98, ES97, EGHS98, Fab95, Fai98, Far91b, FLW97, FT98, GPZ93, GS91, GS93, Gás95, Gla93d]. equations [Gla94d, Gla94c, Gla94e, Gla95a, Gla95b, Gla96a, Gla98, GSC98, GL93, Gra94, Gre94, Gre98a, GP91, GPK99, GW91, Gyö90a, HW93b, HBB90, Heg91a, Heg91b, Hen94, HK98, HP94b, HM91, HW96, Huj96, IESAH95, Ise94, JP95a, JP95b, Jon96, JLDL92, Kan90b, Kan92a, KV94a, Ker91a, Ker91b, KS98a, KT92, Kim98, Kim99, KAS99, Kla98, Koç92, KSM94, KHI94, KP97a, KP99b, Kos90, KS97b, KS98b, KLMP97, KN94, Lai97, Lal94, Lay93, Lee93, LZ94, LY94, Li96b, LC98a, LY98, LF99, Li99a, LX98, LC96a, LP99, LG99, LW91, LS95, MS93, Mak90, MH99, Mar94, MT94a, Mär96, MT94b, MPS94, MO98b, MO98a, Mei93, MZMQ90, Mic94, Mic98, MP91, MK93, MB90, MT98, Moo94, MPR98, MRJ97, MM95d, NFJ96, Nov90, O’R95, O’R98c, O’R99, Ola90, Oli93, Oza94, PR90]. equations [PXHG99, Pop94, QZ91, QZ93c, QZ93b, Rac99, RL91, RSP94, RT99a, Riz95, RZ95, RW98, Rob97, Roj90, Rom98, Rua91, Sak90, Sar90, Sar94, Sch92b, Scr91, Sha90c, Sha95, ST98, SA94, SWY96a, SWY96b, Sol90, Sol97, SCBFC98, SZ91, Sta95, Szm90, TY99a, Tha94, TSL96, TP98, TMGS98, TR99, Tör94, Tre95a, Tre95b, TCS93, Ung90, Ung91, VR95a, VHH95, VK99, WCC99, WT93, Wan93b, Wan96, WA96a, WA96b, Won98a, WA98, Ye99, YCZ98, ZKÖ98, Zaf98a, Zaf98b, Zha91, ZY98, Zha98, ZT98a, ZC98, ZT98b, ZYC98, ZL98a, ZL98b, ZL99a, ZL99b, ZZ99c, ZZ99b, ZT99, ZZ99a, ZY96, ZYW98, ZQ93b, vNvN90a, Ano96-41, Ano97e, Ano97-41, Ano99-33, Ano99-34, Ano99-71, Ano99-84, Ano99-85, Ano99-182, Ano99-197]. equiconvergence [Il’97]. equidistant [GGK96]. equilateral [Köh96c]. equilibrating [AO93a]. equilibria [AP98a, BJPT93, YY98]. equilibrium [Ano95-83, Ano99-75, Ano99-76, Ano99-171, Bai96a, Cub95b, Din99c, Gla90a, Gla91a, Liu98, Tre98, Ano96-69, Ano98-159]. equipped [Ano96-155]. Equivalence [GT92, Zed99, Ano99-86, GS95, GSST97, Gup94, Van92]. equivalent [DT95]. equivariant [Ano99-208]. Érdi [Ano98-113]. Erdos [Ara99b, BL97, BT97, Lub97]. erf [Ano97q]. ergodic [AD93, BS96d, Nas92]. Eric [Ano97-150, Ano97-169, Ano98-48, Ano96v]. Erik [Ano96-46]. Erlang [Bak90, Gup94]. Erlang-B [Bak90]. Ernst [Ano98-91]. Errata [Ado91a, Ano96-171, Ano96-170, Ano99-244, Ano99-242, Ano99-243, Ano99-245, BM90c, Kou90b, MM94, MG90b, Rei99b, Ung91, Zad92]. Erratum [Ano99-246, Moo93b, Sal92b]. Error [AO91, BL95, CT96, EDO93, NO93, AW91, AL93, AO93b, AO93a, Ano98-81, Cha96, DW97, EDO96, EDO97, EL96, EGLN95, FLR92, FLR95, FK97, FO91, Gre91a, Jha90, JP95b, JD97, JPC99, KO91, Kor96, Li98a, Mad92, NFJ96, PC94a, WJ90, WA94a, WA94b, WA96c, Yan93c, Yan99a, YNK92, Ano96-64]. error-correcting [Ano98-81, Yan93c]. errors [ABVB92, Ano96-137, Ano99-244, BB99, BM90a, Din96, Din97b, FK97, FBL99, Hua98c, Hua99]. errors-in-variables [BM90a, FK97, FBL99]. Ershov [Ano93h]. Essays [Ano96-94, Ano96-97, Ano96-128, Ano96-141]. essential [AO99b, Ano99-176, Ano98-51]. essentials [Ano97-27]. estate [Ano97-102]. Esther [Ano96c]. Estimate [Zha99, Ara95, BM96, BS94]. estimated [Ahm94, AGM99]. estimates [AI97, BL95, De 99, DK97b, EDO96, Hal90, KK97b, Kan90a, KAD99, Kor96, Li98a, Pap94a, RSR93, Sha96a, Toy92, WA94a]. Estimating [DIT97, EA92, Pan90a, JCL93].

Estimation [Bag90, CKK97a, ED91a, Li99b, AO91, Ara97, BBM95, BGZ92, CL95, CKK97b, CT96, DW97, EDNO92, EL96, FO91, GW91, Har96, KNU99, KO91, Kun90, ML90c, Nar92, Pap92, PPH98, ST90, Zad89, Zad92]. Estimations [Kri97]. estimator [FK97, FBL99, Kou90a, MCG98b, Tad94, Zhu92]. estimators [AJ95, AO93b, AO93a, AKS96, Faz94b]. Ethnomathematics [Ano96-65]. Étienne [Ano99-152]. Euclidean [TCS93]. Eugene [Ano95j]. Euler [Lei96, BMN96, Gla93d, Gla94c, Gla95a, Gla96a, Gla98, GMS98, HM96, MSR95, Wu90b, Ano99-73]. Euler-like [GMS98]. Eulerian [GN99]. EUROMECH [Ano98-101, Ano99-152]. Europe [Ano97-154]. European [Ano99t]. EUROTHERM [Ano98-155]. evader [GOWR^{+}93, Kle93]. evaluate [CB93b]. Evaluating [McK96, Toy92, DJM98]. Evaluation [Cab97, BS99a, BN92, Dat93, KS99, PSLT93, Pan98c, Tch98, Ano96-155, Ano97-105]. Evan [Ano99-89]. Evans [Ano96-131]. Evaporation [TSLW98]. evasion [Gal93b, Gue93, II93, Yav91a]. Eve [Ano97-135]. Eventual [Che99b]. Eventually [Won98a, AHL98]. everywhere [Ano90r]. Evidence [Ano96-158, Ano98-131]. evolution [Ano98-64, Ano99-74, Bon97, Den91a, Fil93, Har92a, Ano97-46]. evolutionary [Ano96-67, Ano97-165, Ano98-27, Ano99-74, Ano99-166, BMR96, Fog94, Ano96-67, Ano96-66, Ano99-75, Ano99-76, Ano99-74]. evolutions [Kor90b]. Evolver^{T M} [Ano93z]. Evseev [Ano98-153]. Ewing [Ano95-82, Ano99-160, Ano99-161]. Exact [AKS96, GM99, JG96, PT91, Tsi91, VHH95, Ech95, Hun95, Koç92, LS95, WT90b]. exactly [GT97b]. exactness [Zed90]. example [Los89a]. Examples [Ano99-103, Ano97-92, Ano98-126, Ano99-104, AR99b, Ano93o]. exams [Ano98-103]. excel [Ano96-110, Ano99-53, Ano99-54, Ano98-52]. exceptional [Ano97-82]. excess [ZFM96]. exchange [CF91c, CF94, PC92, pri94]. exchangers [DK97a]. exchanges [JAK96]. excited [RS97]. exclusion [FLNW98]. executable [Sni90, Sni91a]. execution [Rei94, Rei99a]. exercise [Saw92]. exercises [Ano98-126, Ano98-53]. exhaustive [SFS95]. Exhibiting [AR90c]. exiomatic [Ano99-105]. Existence [AO99c, CHW93, Cub95b, DL99, Hak99, HW90, KK94, LZ98, Maz90, MO98b, MO98a, WHC90, WJ99, Wu99b, AHY98, AO99d, BM96, Blo91, CL94c, CTY99, Dor90, Ker91a, MN97, SFS95, TC99, WW93, Zaf98a, ZT98a]. existential [Rob92, Ano95-31]. existentialist [Ano97-127]. existing [Gla95c]. Exotic [CL96b]. expanded [Ano98-106]. expansion [Doh90, EDOS93, FLR92, LR93a, MK93, SZ98a]. Expansions [Sza90a, AI97, BDT95, Doh91, HGG99, Il’97, IK97]. expectations [HM95, SGS93]. expected [Dwy93, FF92]. experience [Cse95]. Experiences [Ano97r, dSB96, Ano97-79]. experiment [Ano95s]. Experimental [Ha90, Ano96-64, Ano96-155, Ano98-101, Noa91]. experimentation [Ano93i]. Experiments [QZ93a, Cho95, EA90, FG93, LM96a, hCPW98, Ano98-37, Ano96-68]. expert [Bal90, Chu99, HL90, Lew91a, LR90, YS90, Ano93-27]. experts [Ano98-39, Ano98-110]. Explanatory [RGBK91]. Explicit [JV96, KAS99, NN91, WA94a, ZQ93a, hC96a, hC96b, Bog99, CJS90, CSW99, FLM94, GL96, Hua91, Rhe97, Som90, Tsi99, Wan99, Wu98, hCPW98, vNvN90a]. exploitation [BD96, DB96]. Exploiting [RC92]. exploration [Ano96-155, Wit96b]. Explorations [Ano97-160, Ano98-46]. Exploring [Ano99-81, SHT^{+}90, Ano97-79, Ano96-69, Ano96-71, Ano96-70, Ano98-54, Ano99-78, Ano99-79, Ano99-77]. exponent [Kun95]. Exponential [AHO99, AG94, Car94, CLM90, CLM92, Far99, KYO92, LL91a, MG95, MBV91, SS98a, SM94, Tre98]. exponential-integral [CLM92]. exponential-S-shaped [KYO92]. Exponentially [Car93, Pou90b]. exponentiation [Koç95, LSH95, PPC99]. express [Il’97]. expression [KTO97]. expressions [Ano97-89, MB94a]. Extended [CKAS96, DK97b, DKR97, Hua93, Luo93a, SJS92, CAZAS95, CAZ99, DK97a, Fje91a, Sim97, VDLM95]. Extending [BGM97, Fje91b]. Extension [LH91, SG95a, ADKM97, BHK95, GF90, KLMY97, Lai97, PP97, Sza90b]. Extensions [RT96, AI97, Coo97, Gir92, Gla95c, Pan90b, Ano99-160, Ano99-161, Ano96-72]. extensively [KT98]. exterior [HM99]. external [IIP^{+}97, SV97]. externally [RS97]. extracting [PN98]. extraction [Ano99-80, Sch92a]. extragradient [Noo99]. extrapolated [FLM94]. extrapolation [BS96c]. extrema [EDO97]. extremal [BT97, Pan90a]. Extreme [Gry94, KKS98, RRS99, Zha93c]. extremum [Ano98-36]. extrinsic [Lee96].

F [Ano90x, Ano93o, Ano95-73, Ano95-83, Ano95-68, Ano96-36, Ano96-75, Ano96-82, Ano96-120, Ano96-123, Ano97-54, Ano97-81, Ano97-89, Ano97-115, Ano97-151, Ano98-45, Ano98-148, Ano98-93, Ano99-51, Ano99-40, Ano99-53, Ano99-54, Ano99-65, Ano99-59, Ano99-172, Ano99-203]. F.E.M [Ano98-116]. Faber [He95a]. Fabien [Ano98-84]. Fabrikant [RF91]. Fabry [KKG97]. face [Ano99-231]. Faces [RR97b, Dwy93]. facility [BC98, Chu99, Xue94]. fact [Ano93f, Ano99h, Ano99i]. factoring [Tan92]. Factorization [Doš98, NP96, Ara96, DHW98, LVV96, Més96, NWB92, Pan95, Res91, Tor90]. factors [Ano96-57, DPD98, Los89b]. Facts [dJPvZ97, Mar92]. Fagin [Ano96-138]. fail [Ano99-216, Ano99-217, NMUY92]. fail-safe [NMUY92]. failure [Dei93, KKOS96, YF93]. fair [Ano95-37]. Families [KV94b, VBM90, CS94a, SG95b]. Family [DT95, Ano95n, Bog96, CGHS92, GT95, JS99, LR93b, Pat98, PIT97, PPH98, SR92, SM94, WF92, WS92, Xu97]. Fan [WX97, WY99, YY98]. Farid [Ano96-145]. Farin [Ano95-55]. Farrell [Ano95-39]. Fast [Ano98-43, BN92, BB93, EKC91, FS96, JS95, JS97a, JS98, KO93, McK96, Més96, PS92, Pas97, YJ99, CL90, DJM98, Gal93c, GG90, HLS91, LYL^{+}91, Lee96, LJ98, MZ97, PSLT93, PZDH95, Pan98c, RSC93, Ano98-55]. Fauconnier [Ano97-135]. Fault [DH97, HTT92, LYH98, Rei93b, Ano98-56]. Fault-Tolerant [DH97, HTT92, Ano98-56]. Faurre [BM90a]. Fay [Ano98-112, Ano98-132]. Fay-Wolfe [Ano98-132]. Fayolle [Ano99-198]. Fayyad [Ano96h]. FDE [WJ99]. FDMS [NN91]. feasibility [Ano98-32, YZM94]. feasible [Sny90, YFZM95a, YFZM95b, dKSG94]. Feature [SB97, FT97b, Ano99-80]. feed [MKS^{+}91]. feed-forward [MKS^{+}91]. feedback [Ame99, CKL98, EK97, Gal93d, LTR98, RTJ96, SHT^{+}90]. feeding [YF92b]. Fehlberg [Sim93b]. Feigenbaum [Ano96-44]. Feldman [Ano96-44]. Felicitas [Ano99-140]. Felipe [Ano97-51]. Felleisen [Ano96-154, Ano96-160, Ano98f]. Feller [AZ96]. FEM [IB95]. Feng [Ano98x]. Fenton [Ano97-70]. Fermat [Ano96-124, Ano97-121, BL99]. Fermi [Ano95n, Ano96-63, Ano95n, Ano95-54]. Ferrat [Ano99t]. fertility [KL96b]. Fetzer [Ano97-120]. Feuerstein [Ano99-178, Ano99-179]. Few [EK92, Ano98f]. Feynman [DR96, Ano99-81]. FFT [HLS91, HWL94, HGG99]. FFT-based [HLS91, HWL94, HGG99]. Field [Mur91a, Ano96-99, Ano97-121, ASW98, BK91, BKK97, Dor98a, Dor98b, EGK96, Jek94, KK97a, LD95, Mur93b, UW94, Wos95b, Ano97-128]. fields [ARI96, Ano97-48, Ano97-47, Ano98-65, Ano99-197, Ara90, Leo90, RP97, TCJG91, Web95, Ano99-186, Ano95-32]. fifth [Ano96y, Ano96-127, Ano97-93, Ano99-90, Ano99-181]. figures [IK95, Köh94, Köh96b, Köh96c, Köh97]. file [Ano98-163, CHCC95, CCH97]. Filippov [Ano99-33]. Fillia [Ano98-47]. Fillmore [Ano96d]. film [CC94c]. filter [BBM95, KKM96, NL91, OCS96, Ram90b, SGS93, Sza90b]. filtering [Ano99-174, CSOL96, FK95, Gyö90a, MG90c, Sch92b, Ter00, ZM94]. filters [OSLC98]. filtration [Zed99]. finance [Ano97w, Ano98-62]. financing [Ano99-214]. find [Nis97b]. Finding [CST97, LM96b, Ram91b, TYC93, Ano99-123, Ano99-124, NB95, PT96, PIT97, PPH98]. Findler [Ano93f]. Finite [Ano96-83, Bog98, Du94, Far91a, Far91b, GHS91, Gup94, IB95, MR94, SS98a, Sem93, Shi98, Tad93, Alm96, BBO99, BL94, Bed95a, BS96d, CHSS95, CP98a, CS93a, Che90a, Che93b, Che96, Che91, CCVV90, DBM92, DG99, Dem94a, Dem94b, DO91, Elo98, Fai98, GN99, Gla91a, Gla91e, Gla92b, Gro95, GMR90, HM99, HNS^{+}90, HS98, KP99a, KPS96, Kwo92, LLP94, Lay99, Li97, Li98a, Li98b, LN98, LN99, Li10, MPS93, MG90c, OPD97, Pac94, PSS99, Pao98, PR91, RD96a, RV94, SK91, Sha90a, She92, She93, SFT99, SWY98, Sol97, VDLM95, Wac91a, WF99a, YZM95, Ye99, ZS93, Zha91, ZM96, vNvN90a, Ano95-33, Ano97-48, Ano97-47, Ano98-58, Ano98-57]. Finite-difference [MR94, CHSS95, HS98]. Finite-element [Shi98, Fai98, GN99, Ye99]. Finite-state [Ano98-58]. finitely [CR90b]. Finiteness [Ano99-82]. Finke [Ano96-47]. Fiore [Ano96t]. fire [RG90b]. Firebaugh [Ano90o]. firewalls [Ano96w]. firings [Fau94]. First [Ano90-28, Ano96-135, MK93, Mol99, Ano96-72, Ano97-33, Ano97-57, BL96, CL96a, Cho94, De 99, GC99, HBB90, Sán98, WCC99, WSNS98]. First-order [MK93, De 99, Sán98, WCC99]. Fiscal [Ano97-49]. Fischer [Ano96-69]. fish [Niw96]. fission [BCC95]. Fitted [Car93, Gla91b, SM94]. fitting [MBMS92, Nyí99, ZM99]. fitz [Cho92]. fitz-hugh-nagumo [Cho92]. Fitzgerald [Ano96y, Ano99-154, Ano99-155]. FitzHugh [Cho95, CL96b]. Fiume [Ano95j]. Five [DBM92, ESIP98]. Five-diagonal [DBM92, ESIP98]. fixed [AO98b, HK96, Hua97, Hua98c, Kle93, LK94, MM93c, MM94, O’R95, O’R98b, WW98, WT90a, WW93, YX95, YY98, Ano98-59]. fixed-charge [WW98]. fixed-point [AO98b, Hua97]. fixed-points [O’R98b]. FL [Ano98-43, Ano98-107, Ano99-108, Ano99-125, Ano99-126]. Flanagan [Ano95-52, Ano96-92, Ano96-93, Ano98-85]. Flannery [Ano90w]. Flaschel [Ano97-42]. Flat [Chi93, WF90, YC92]. Flattening [Ano98-60]. flexible [FYS99, YF92a, YF92b]. Flight [RG90b, Edw92]. Floating [CL91a, CL91b, CHW93, JP93, Ano99-83]. Floating-point [CL91a, CL91b]. Flow [Chi93, Ano96-46, Ano99-70, Ano99-136, Ano99-242, ASW98, Ber92, Bla95, CC94a, Cha91, Che91, Gla91b, Gla91c, Gla91d, Gla92b, Gla92c, Gla94e, GK91, HR99, Iri98, Kha90, KESH90, LT96, LMW95, MATA98, MPS92, MPM93, MP93, Mol93, PSD96, Pio94, SM98, Sán98, SJB97, Son92, SC98b, Ste90, ZSDB99]. flow-shop [Ste90]. flowfields [Bak92]. flowing [Ano99-83]. flows [Ano98-84, BM90b, CHW93, EL96, Eti91, Gla91a, Gla92a, Gla92d, Gla92e, Gla93b, Gla93d, Gla94a, Gla94b, Gla94f, Gla95c, Gla95d, Gla95e, Gla95f, Gla96b, Gla97a, Gla97b, Gla99, Hak99, HW90, Iri98, IGM98, KP99a, KS95, Kro99, LK94, LLP94, LSH98, LS99, Pas92, Sak90, Tan93a, WHC90, WW95, Yas98b]. fluctuation [Ano97-42]. fluctuations [AKK^{+}97]. Fluid [Edw92, Ano98r, Ano99-83, CH91, Gre91b, Kan90a, Kan90b, LLP94, LS99, Mah90, MP93, MP94, PSD96, Pas92, RD96a, Sán98, Sau92, TH95, Wea92, YX95]. fluid-dynamics [Kan90a, Kan90b]. Fluid/chemistry [Edw92]. fluids [Ano99-113, BMN96, EL96, Hak99, Kup91]. Flux [Gla90a, Gla93d, Gla93c, Prn97]. Flux-difference [Gla90a]. fluxes [AO93a, GRC^{+}96]. flying [Ano99-83]. Flynn [Ano97-127]. Focal [HK98, AP98b, Ano99-84, Ano99-85]. Fogel [Ano96-67, Ano97-57]. Foliated [Mar92]. FoLLI [Ano97n]. followers [Liu98]. following [BS99b]. Fomenko [Ano98j]. Fomin [Ano99-174]. force [MBMS92]. force-fitting [MBMS92]. forced [BF97, CL96b, Ira93, NAKA98, TR99]. forcing [Zaf98a]. Forecasting [Cha91, Mit92a, Cha93, JS97a, VA94, Ano96x]. forest [JT99]. forest-like [JT99]. Foreward [Dey90]. Foreword [Lee91, LL94b, Mai96a]. Form [Kup91, LC96b, Ano96-100, Ano97-121, BLR94, DC94, Den91a, Fai98, HV92, Köh93a, Köh95, Köh96a, Köh96b, Köh96c, Men91, NGNM97, Sjo92, Var98]. Form-invariance [Kup91]. Formal [Tan94, TL94, Ano97-159, Ano99-82, Hol90, LRLR95, ZQ93c, tC95, Ano99-86, Ano99-87]. formalism [Tan93a, Ano97-39]. formally [Sun91]. Forman [Ano96-137]. formation [Bar93, Kor97, Mai96b]. forms [AF90, CKL91, Köh91, Köh93a, Köh95, Köh96a, Köh96b, LR93b, Mit90]. formula [Bak90, CAZ99, FK98a, JG98]. formulae [DELE99, Sar94, Zed90, ZA95]. Formulas [And91, BS90, CMR90, CV90, DOS93, EG97, FLR95, KT96, VL95]. formulation [EDO98, ESKG98, GN99, Gla94f, Gla95c, Gla95f, She92]. formulations [Gla96c, Gre90a]. Fortran [Ano98p, Ano99-48, Ano99-47, SZM98, Ano96g, Ano97-112, Ano98-61]. Forward [BTT90, MKS^{+}91]. Fosdick [Ano96o]. Foucault [Ano97-127, Ano97-50]. foundation [Gab94b]. Foundations [Ano99v, Ano99w, Ano99-232, AO93b, Ano96-82, Ano99-149, Ano95-35, Ano95-34, Ano96-73, Ano97-51, Ano97-52, Ano99-88]. four [CST97, GF96, SM94, Köh96a]. four-dimensional [CST97]. four-step [SM94]. Fourier [Ano98-43, RSC93, Oli98, Wan91]. fourteenth [Ano98-96]. Fourth [Ano96-166, Ano95k, Ano95-56, Ano96-67, Ano96-154, Ano97-53, Ano99-227, CHSS95, DBM92, DBM94, MHBD96, PS96, Shi98]. fourth-order [DBM92, DBM94, MHBD96, PS96, Shi98]. Fox [MM95c]. Fractal [Cie95, Ano98-75, Nel96]. fractals [AR90c, Ano97-92, Ano90r, Ano98-62, Ano99-89]. fraction [LR93a, Mac93, TCS94]. Fractional [AIS95a, Ism97, SH95b, AIS95b, Ano99-63, Ano99-64, AK90b, CIS98, KNS99, Liu96b, LY99b, LKT99, LS95, RS96a, RS99a, SA95]. fractions [AE90, EKC91, LL91b]. fragmentable [MCG98a]. fragmentation [CC99, FO93]. frame [Ano97-134, Gla91d, Gla93b]. frames [NF92]. framework [BSW96, Yoo93]. frameworks [Ano96-119]. France [Ano98-84, Ano99t, Ano99-152]. Francesco [Ano96-90, Ano98m]. Francisco [Ano96-135]. Franck [Ano96-31]. Franco [Ano96-114, Ano99-165]. Francois [Ano97-56]. Frank [Ano97-111, Ano97-112]. Franke [Ano97-42]. Franklin [Ano96s]. Franz [Ano98-147]. Fraser [Ano90o, Ano96-89]. Fréchet [Arg99]. Fred [Ano98-41]. Fredholm [BAS97, MH99, Ola90]. Free [Has98, AO98c, Ano90-28, Eti91, EGHS98, FG98, GH90, Gla95e, HJ91, Hun95, MM93a]. free-boundary [Hun95]. Free-streamlines [Has98]. free-surface [Gla95e]. Freire [Ano96-149]. Freiss [Ano98-129]. frequency [Fli95, KL96a]. Fresnel [SD95]. Frey [Ano99-106, Ano99-107]. Friedl [Ano97-89]. Friedman [Ano96-154, Ano96-160, Ano98-70, Ano98f]. friendly [Ano99-190, Ano99-191]. Frieze [Ano93p]. Frisch [Ano98-51, Ano99t]. Fritzsch [Ano95i]. Frobenius [dOF91]. Frølund [Ano97-32]. Front [GGG^{+}98, SCG^{+}97]. frontier [Ado91d, Ano96-85]. frontiers [Ano96-167]. Froot [Ano99-214]. FSMs [LYH98]. Fu [Ano98-161]. Fuchssteiner [Ano96-120]. Fudenberg [Ano98-152]. fudge [Ano96-57]. Fugini [Ano96-48]. Fukushima [CH99]. Fulachier [Ano98-84]. full [Sak90]. Fully [KP99a, Lee94, LA96]. function [Ahm94, Ahm95, Ano98-113, Ano99-98, AB93, BL95, Bax92, Bin91, CB93a, Che93a, CL98, CLM92, EDO96, EDO97, EK94, Fai98, Ike91, KS96, LR93a, LL91a, LS99, Mat99, Pan90b, PDK96, Pan93b, PN98, RS95, SFY92, SH95b, UWT99, Wei99, WT90b, WCY94, ZB95, Zha95]. Functional [DT93, Kur91, AG99a, AG99c, DT95, Fab95, FK97, GL93, Gra94, JTY97, LH99, MW99, Rua91, SLC97, Tse95, ZL98b, ZZ99c]. Functionals [ABI99, Ara90, Bad95, DLS99, Kou95, Ruz90, Sza91, Yan93d]. functions [AR92b, AL93, AD96, AIS95a, AIS95b, Ano96-146, Ano97o, Ano98p, Ano99-59, Ano99-153, AH97, Bad95, BL96, BBI97, BN92, BM97, Bro94, BS94, BFR94, Cab97, CFM91, CS94a, CIS98, CWJ92, Cie95, DCL^{+}93, DMC^{+}95, DTLM98, FT97a, GG97, GS99, Gir92, GZ95, GBB91, HS95, Haw91, Ioa91, JS99, Kar96a, KL95, Kov95, KNS99, LOS99, LW99a, Lin91b, LR93b, LR96b, Liu96a, Liu96b, xLhW96, Mar95, MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b, Men91, Mil91, MM95c, Mur92a, Mye92, Noa91, PDK96, Pas97, PM91, PT96, Pro95, QZ93c, RS96a, RS99a, RS99b, RK98, RSS94b, RHRZ90, Rok90, RFRM99, RP97, Rou92, RZ94, Sal92a, She95, Shu93b, SSH92, SA95, SG95b, Sri95, SHA99b, Sun92, VPK99, WCHL96, XC92, Zha96, Zhe93, ZZ99d, vNvN90a, vNvN90b, Ano96-37, Ano97-138]. functions [Ano99-59]. Fundamental [Ano96-129, KO96]. Fundamentals [Ano96-74, Ano96-75, Ano99-92]. Furht [Ano98-110]. Further [HM96, Roj99, AR92a, TY99a]. fuselage [BR95b]. Future [Ano97-28, Ano95-73, Ano95-77, Ano95-76, Ano97r, MM96e]. future-oriented [Ano95-73]. Fuzzy [CL94a, CL94b, CL99a, Che99c, Eso99b, GHWL96, HGWL96, Köh99, Lew91a, Li99c, MS91, Zad99, Ano96-122, Ano97-39, Ano97-52, Ano99-215, CL94c, CL95, CL99b, CL99c, CLY97, CR90c, DK97b, Din99a, Din99b, Eke99, Eso91, EM94, Eso99a, EL99, FHWW99, Köh91, Köh93b, Köh94, KYNY99, LC97a, LZ94, LL90a, LL91a, LWL96, LO99, ML90a, NL93, SNU98, Sen92, Sya97, Sya99, TI96, WW98, Yan93b, YC94, Yas94, Yos96, Yos99, Ano97-54, Ano99-94, Ano99-93]. fuzzy-logic [Ano96-122].

G [Ano93o, Ano95v, Ano95x, Ano95-27, Ano95-64, Ano95g, Ano96-41, Ano96-67, Ano96-64, Ano96-83, Ano96-91, Ano96-116, Ano96-111, Ano96-107, Ano96-117, Ano96-120, Ano96-123, Ano96-140, Ano97-106, Ano97-105, Ano98-37, Ano98-64, Ano98-92, Ano98-114, Ano98-116, Ano98-135, Ano99-66, Ano99-67, Ano99-83, Ano99-81, Ano99-113, Ano99-140, Ano99-182, Ano99-196, Ano99-211, Ano99-231, Ano99-91, CK97b, Kim92, Pou90a, Ver99a, Wu90a, Wu90a]. G-H-KKM [Ver99a]. G/G/c [Wu90a]. Gabriel [Ano96-95]. Gaile [Ano99-231]. Galerkin [Bed95a, Bla95, DNJS98, EDO98, EY91, KS91, Say99, vNvN90b]. Galil [Ano97-118]. Gallagher [Ano95w]. Galois [Ano98w]. game [Ano95d, Ano96-66, Bel94a, Bel97b, Bel97a, Gue93, Nak99, Özy96, Ano95-37, Ano98-63]. games [AK93, Ano95-62, Ano99-75, Ano99-76, Bai96a, BJPT93, GZ91, Gal93b, GP93, KK91, Kos93, Yav93, YY98, Ano98-152]. Gammerman [Ano96-38]. Ganzah [Ano96-41]. Garfinkel [Ano95-58, Ano96-133, Ano97-167]. Garg [Ano98-43]. Garrett [Ano96-54, Ano97-81, Ano98-42]. Gary [Ano96-86, Ano97-62, Ano99-65]. gas [ASW98, Fre98, Gla90b, Gla91d, Gla92d, Gla94b, Gla97a, Gla99, KESH90, KLMP97, SC98b, Wen99]. gases [Gla90a, Gla92a, Gla92b, KKK^{+}97]. Gass [Ano97-45]. Gastel [Ano95u]. Gateways [Ano97-55]. Gauge [Ano97-56]. Gauss [Ano99-244, BB99, BRV98, GPGCM94, KNUX99, KT95, Mac93, WT93, Zou91]. Gaussian [AJ95, Ara90, Faz94a, Faz94b, HGK93, Isp98, Kim98, Kor91, Leo90, MM95d]. Gaylord [Ano96-76, Ano96q, Ano97-96]. GBP [Ano98q, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98w, Ano98x, Ano98r, Ano98z, Ano98-27, Ano98-37, Ano98-31, Ano98-32, Ano98-36, Ano98-44, Ano98-45, Ano98-43, Ano98-49, Ano98-53, Ano98-50, Ano98-47, Ano98-59, Ano98-56, Ano98-57, Ano98-64, Ano98-65, Ano98-66, Ano98-68, Ano98-67, Ano98-69, Ano98-72, Ano98-76, Ano98-84, Ano98-74, Ano98-83, Ano98-89, Ano98-90, Ano98-91, Ano98-92, Ano98-109, Ano98-108, Ano98-100, Ano98-101, Ano98-110, Ano98-111, Ano98-116, Ano98-126, Ano98-127, Ano98-136, Ano98-131, Ano98-132, Ano98-135, Ano98-144, Ano98-138, Ano98-139, Ano98-140, Ano98-155, Ano98-156, Ano98-161, Ano98-164, Ano98-62, Ano99z, Ano99v, Ano99-27, Ano99w, Ano99o, Ano99r, Ano99s, Ano99n, Ano99t, Ano99x, Ano99u, Ano99-33, Ano99-37, Ano99-41, Ano99-42, Ano99-50, Ano99-213, Ano99-44, Ano99-49, Ano99d, Ano99-43, Ano99-46]. GBP [Ano99-63, Ano99-66, Ano99-55, Ano99-64, Ano99-67, Ano99-56, Ano99-69, Ano99-52, Ano99-60, Ano99-57, Ano99-58, Ano99-78, Ano99-79, Ano99-77, Ano99-70, Ano99-84, Ano99-85, Ano99-86, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-214, Ano99-89, Ano99-82, Ano99-87, Ano99-105, Ano99-103, Ano99-99, Ano99-96, Ano99-109, Ano99-102, Ano99-104, Ano99-100, Ano99-97, Ano99-110, Ano99-95, Ano99-101, Ano99-108, Ano99-115, Ano99-111, Ano99-116, Ano99-112, Ano99-117, Ano99-120, Ano99-114, Ano99-125, Ano99-129, Ano99-126, Ano99-130, Ano99-128, Ano99-121, Ano99-136, Ano99-137, Ano99-132, Ano99-145, Ano99-144, Ano99-142, Ano99-218, Ano99-156, Ano99-219, Ano99-157, Ano99-152, Ano99-149, Ano99-146, Ano99-162, Ano99-153, Ano99-167, Ano99-168, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-170, Ano99-174, Ano99-175, Ano99-183, Ano99-193, Ano99-184, Ano99-194, Ano99-181, Ano99-186, Ano99-182, Ano99-189]. GBP [Ano99-196, Ano99-197, Ano99-201, Ano99l, Ano99-198, Ano99-206, Ano99-207, Ano99-204, Ano99-209, Ano99-208, Ano99-229, Ano99-212, Ano99-211, Ano99-227]. GDM [Ano98q, Ano98-105]. Geddes [Ano96-107]. Gee [Ano98-73]. Gegenbauer [GKT99]. Geisselhardt [Ano99o]. Geist [Ano95-59]. Gellman [DR96]. gems [Ano98-87]. Gena [Ano98-66]. Gene [Ano96-133]. General [Ano96-135, Ano97-130, Ano99-105, But96, CL99d, Din97a, MPP97, MTGFG93, AO97, Ame97a, Ano96-69, Ano98-159, Ano99-102, BKM91, CSW99, Cub96, Eso99b, Gla93d, Hua98b, Ill91, LB94, MT94a, Noo99, PA94, PA98, RSC93, She98, SH96, SFY92, SZ91, TP98, Ush94a, ZG97, Pol90]. generalisations [Ano98-92]. generalised [Ano95-68, Ano99-59, Ano99-181, EGK96, Gla95d, Mai96b]. Generalization [AR92b, Elo98, Shv95]. Generalizations [MM96e, WX97]. Generalized [Bör90, CLY97, DTLM98, Din99b, Din99c, GT97a, GS96c, Noo95, Noo96, Noo98, PR93, TCS93, Ver99b, YLY96, ZZT99, ACHP94, Ano99-95, Ara90, BGL97, BL99, CTY99, CLM90, CLM92, CY94, CY95, DCL^{+}94, DMC^{+}95, DL99, Din99d, GT97b, Gla93a, Hua93, Hua98a, HLTB99, JA97, JY93, KT96, KS98a, Kim99, KO99, LCL94, Liu96a, Liu96b, MM96c, NAS97, PS93a, Qu96, RB97, RW98, Shi95b, Sib98, Sni90, Sni91a, Sri95, UK99, Yan93b, YC94, YQM99, ZC91, Ano99-96, Ano99-97, Ano99-95, Ano99-98]. generally [Ter00, ZM94]. generate [Bis91, BG91a]. generated [CC97, Cha91, HBB90, Han92, JGB95, WA98]. Generating [DT93, FGN90, Heg91a, Heg91b, Lin91a, SYZ94, CS94a, CCH97, LOS99, NR96b, RSS94b, Shu93b, SSH92, SG95b, Sto92, Zho90]. Generation [ISM93, Ano98-137, Ano99-246, AÉ96, Bak92, BV95, BV99, Cas91, CO99, DC94, Eis92, KKK91, MC97, Sau95, Son92, SH94, WT90a, Ano99-142]. generations [FKO92]. Generator [SL99, LKC95, LHC96, Sez95, SD93, Tan92]. Generators [MZ93, Jab96, KT97, KW98, KT98, PW95]. genesis [Shi95b]. Genetic [Ano99-99, Ano99-100, AC93b, HSH90b, AW94, Ano93z, Ano95h, Ano96n, Ano97f, Ano99p, Ano99-74, Ano99-226, CLD97, CHW99, HSH90a, Jan96, KM96, Liu98, May99, MJK92, TWIF98, WF96, Ano95-38, Ano97-57, Ano98-64, Ano99-99, Ano99-100]. Gennaro [Ano95-30]. genomic [Sch90]. Geo [CK97b]. Geodesic [Tan93a]. Geoffrey [Ano99-232]. Geom [CTG96].

Geometric [GMBS94, Ano99-46, CKL98, EL92, Ill91, Köh94, Köh96b, Köh97, KCB90, MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b, Sib98, Sti93a, Ano99-101]. Geometrical [MP96, Lee96]. geometries [Gla95e]. Geometry [LY92, Ano95-55, Ano96-82, Ano98-33, Ano98-107, Ano99-105, Ano99-103, Ano99-102, Ano99-104, CP95a, Huj99, II93, Nas95, Ram91e, Ano96-33, Ano97-56, Ano98k, Ano97-60, Ano97-59, Ano97-58, Ano98-65, Ano99-105, Ano99-103, Ano99-102, Ano99-104]. geophysics [ZR91]. George [Ano95r, Ano95s, Ano95-69, Ano96-36, Ano96-125, Ano97-35, Ano99-127, Ano99-164, Ano99-205, Eso99c]. geosciences [Ano96-145]. Gerald [Ano95-79, Ano95-55, Ano96i, Ano97-67, Ano97-137, Ano98-71, Ano98-75]. Gerd [Ano97-92, Ano99-88]. Gerhard [Ano98-45, Ano98-111]. German [Ano97-151]. Germany [Ano96-120, Ano96-165, Ano97-140, Ano97-92, Ano97-111, Ano97-112, Ano97-136, Ano98-93, Ano98-111, Ano99-166]. Gerry [Ano96-32]. Gert [Ano97-56, Ano98-66]. Getting [Ano95-39, Ano96-76, Ano96-77]. Geurts [Ano99-83]. Geuvers [Ano95-65]. GFLOPS [DF98]. Gheorghe [Ano98-100, Ano99-114]. Ghetie [Ano97-105]. GI [CTG96, Kim92]. GI/Geom [CTG96]. GI/M/1/s/s [Kim92]. Giacomo [Ano96e]. Giancarlo [Ano96-48]. Giannessi [Ano95-83, Ano96-123, Ano99-165]. Gibbons [Ano95c]. Gibbs [Ano96-119, GG97]. Giblin [Ano99-231]. Giere [Ano99-202]. Gieryn [Ano99-51]. Gilbert [Ano99-182]. Gilles [Ano97-135]. Ginzburg [AM95, Du94, GHS91]. Giorgio [Ano98-68, Ano99-58]. Gips [Ano96-110]. Girodias [Ano99-117]. Gisela [Ano97-111, Ano97-112]. Gitta [Ano96o]. Giulianella [Ano96-113]. given [AR99a, Ano95-54]. Givens [LLR^{+}90].

Glasgow [Ano96-53, Ano97-47]. Glassman [RSC93]. Glen [Ano97-30]. Glenys [Ano99-234]. Glesner [Ano95-53]. Global [AO99d, Flo97, Goh94, KB90, Li98a, Moo91, Moo93b, PZ99a, Par91a, Par91b, SWY97, ZZ99d, ZYW98, ZZ99e, Ano96-121, Ano97-110, Ano99-181, AKV91, BS91a, CC96, Dei91, Dei93, FZ93, GZ91, GPZ93, Gal93b, GZ93, GW99, GS99, KVCA^{+}99, LN99, Lu96, LW99b, Men91, Pic91, RR91, SC98a, TZ93, Woo91, ZC96, Zhe91, Zhe93]. globally [Oza94, Xue94]. globin [LNM90]. Glymour [Ano99-40]. GMRES [Rac99]. GNU [Ano97-124]. Goal [Ano96-78]. Goal-driven [Ano96-78]. Gödel [Ano97d, Ano95-78, Nam96b]. Gödel’s [Ano97-94]. Godunov [Wen99]. Godunov-type [Wen99]. Goertzel [Ano95p]. Goguen [Ano96k]. Goldberg [Ano97-57]. Golden [Ano97-91]. Goldschlager [Ano90p]. Goldsmith [Ano96-57]. Golinkoff [Ano96-158]. Golshtein [Ano96-117]. Golub [Ano97-146]. Gong [Ano98-76]. Gonzalez [Ano96-68]. good [Ano95-58, Ano96-39, Ano99-216, Ano99-217, Cor94]. Goodbye [Ano97-61]. Goodman [Ano99-190, Ano99-191]. Gordon [Ano96-153, Ano99-186, Ano99-231, Ado91c]. Gossain [Ano98-119]. Gould [Ano99-230]. governed [Cho92]. governing [Kan92a]. GP [YQM99]. GQVI [Cub96]. graceful [JLL^{+}93, LJL^{+}93]. graded [DKW98]. Gradient [PS92, Ara96, Ara99a, Bro96, CLM99, HS97, Luo93b, PZDH95, Sha96a, Sha95, ZM99]. Graeffe [JS98]. Graf [DCL^{+}93]. Graf-type [DCL^{+}93]. Graham [Ano95-50, Ano96-81, Ano96-80]. Gram [MNB93]. grammar [Ano96-158, Ano95-30, Ano96-161, Ano97-135, Ano98-100]. grammars [BRGJ95, EHK92]. Gramß [Ano99-166]. Grand [Ano97-75]. Granino [Ano96-122]. Grant [Ano96k]. Granular [Hel96, BCGP99]. Granularity [MGBP92]. Graph [BP90, HO97, KM97a, Mez97, PJ91, Rea97, Ano98-55, Ano99-156, Ano99-157, AS95, AS98, BGZ92, CF91c, EHK92, HM94, JL94, LS96b, MR92, Nis97b, Reg94, SG95a, SS96, SS98b, Ano98-66]. graph-based [SS96, SS98b]. graphical [Ano95j, Ano95-57, Ano99-106, Ano99-107]. Graphics [Ano99-44, Ano99-46, Ano99-150, Ano99-199, Ano99-200, Ano99-138, Ano99-199, Ano99-200, Ano99-109, Ano99-110, Ano99-108]. Graphs [Bei97, Hev97, Jar97, LW97, AAD98, Ano93p, BV95, Che99a, Cim90, CF91b, CF94, GF94b, JAK96, LO93, NNB97, ND95, Rob92, RR97a, SMT98, Sow92, SSBG91, SY96, SS91b, SS93, TYC93, Zad99, Ano97-62]. Grätzer [Ano96-111]. gravitating [RD96a]. gravity [ARI96, Jek94, Jon96]. Gray [Ano98-107]. Grazia [Ano96-48]. Grazyna [Ano98-44]. great [Gil93]. greater [ACL93]. greatest [Ano96-57, Ric95]. greedy [LO93, TTC96]. Green [AL93, CL98, EK94, Haw91, Ike91, KL95]. Greenberg [Ano99-224]. Greenspun [Ano96-153]. Grefenstette [Ano99-50]. Greg [Ano99-188]. Gregory [Ano96h, Ano96-132, Ano97-100, Ano97-133, Ano99-50, Ano99-40]. Grene [Ano96-50]. Grid [LD90, Son92, Ano99-246, Bed95b, Cas91, CO99, CLD97, Eis92, Fun97, Gol94, PC96a, PR93, SAS92, Sau95, SCC97, WT90a, Yan93a, Yan97]. grid-generation [Ano99-246, CO99]. grids [Boy99, CHSS95, HS97, Mas92, PR96, YX95]. Griffith [Ano99-163, DD99]. Grigore [Ano96-43, Ano98-53]. Grigoris [Ano97-109]. Griliches [Ano98-131]. Grillner [Ano98-114]. grinding [SMU98]. Gröbner [IA94]. Gronwall [PA99]. groove [RSV98]. Gropp [Ano95-82, Ano97-41, Ano99-160, Ano99-161]. Groß [Ano99-166]. Grötschel [Ano96-81, Ano96-80]. ground [Gre92a]. groundwater [SM98]. group [Ano97-79, Ano99-71, Ano99-181, JK96, Jon96, Li99c, MTLGFG93, MTLGFG95a, YJ99]. groups [Ano98w, Ano98-92, Ano98-112, Ano99u, AH97, ARS98, AR99b, Kun95, MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b, Sok91, Ano95-43, Ano97-58, Ano97-79, Ano97-82, Ano97-157]. Growing [Ano97-63]. growth [Ano96-64, Ano97-42, CC98c, KYO92, Mit92a, Pan92b]. Grüss’ [DW97]. Guedes [Ano98-127]. GUI [Ano95-72, Ano95-52]. guidance [FY93, YFF92b, YFF92a, YFF93]. Guide [Ano96-116, Ano95f, Ano95-42, Ano95-59, Ano96-99, Ano96d, Ano97o, Ano98p, Ano98-149, Ano99-69, Ano99-193, Ano99-194, Ano99k, Ano99-246, CO99, FG98, Ano96-88, Ano96-93, Ano96-106, Ano96-107, Ano98t, Ano98-85, Ano98-122, Ano98-163, Ano99y, Ano99-45, Ano99-180, Ano99-188]. Guidelines [Ano96-151, Ano90e]. guiding [Ano93i]. Guillemin [Ano99-208]. GUIs [Ano99-108]. Gunawardena [Ano99-129, Ano99-130]. Gundavaram [Ano96-28]. Gunetti [Ano96-90]. Gunter [Ano95-81, Ano99-145]. Guo [Ano99-182]. Guo-Chun [Ano99-182]. Guoquin [Ano99-94]. Gürlebeck [Ano97-125]. Gurtin [FLM94]. gust [WF90]. Gustafsson [Ano96-164]. Gustavo [Ano96m]. Guthrie [Ano96-58]. Guy [Ano99-198]. Gwynne [Ano96-131]. gyroscopic [FLW99].

H [Ano90w, Ano93n, Ano95l, Ano95y, Ano95-65, Ano96-27, Ano96-43, Ano96-62, Ano96-74, Ano96-120, Ano96-130, Ano97p, Ano97-47, Ano97k, Ano97-145, Ano97-120, Ano98i, Ano98-29, Ano98-39, Ano98-82, Ano98-83, Ano98-101, Ano98-145, Ano98-146, Ano99-32, Ano99-57, Ano99-71, Ano99-83, Ano99-101, Ver99a]. Haack [Ano97-39]. Habermas [Ano97-80]. habitat [CC99]. hacker [Ano97-150]. Hadamard [Coo97, Gro95, Zha91]. Haefel [Ano99-72]. Haifa [Ano99-136]. Haitao [Ano97-163]. Haith [Ano95-73]. Hal [Ano97w]. half [CR90b, CK94a, DPD98, Duc95a, LYL95, MB92, O’R98a, WA98]. half-linear [LYL95, WA98]. half-plane [DPD98, MB92]. half-space [CR90b, Duc95a]. Hall [Ano90p, Ano90q, Ano90s, Ano90x, Ano93k, Ano93n, Ano96e, Ano96-71, Ano96-98, Ano96-102, Ano97-93, Ano97-113, Ano97-114, Ano97-116, Ano97-165, Ano98-30, Ano99-185, Ano95m, Ano95l, Ano95-53, Ano96f, Ano96-56, Ano96-65, Ano97-62]. Halley [GH98]. Hallinan [Ano99-231]. Halpern [Ano96-138]. Halton [KO93]. Hamburg [Ano98-53, Ano98-111]. Hamdy [Ano97-113, Ano97-114]. Hamideh [Ano99-128]. Hamidzadeh [Ano99-131]. Hamilton [Ano99-38, LW97, LQ95b]. Hamiltonian [McL95, ZQ95, APV97, Boh98, Gre90a, HI98, IK97, JAK96, Kar96b, Kup91, MG91, MZMQ90, Pal97, PP97, Reg94, Shi94, Tan93a, Tan93b, Tan94, TL94, Wan91, ZQ94]. Hammerstein [Ame97b, EGHS98, O’R95]. Hammerstein-type [EGHS98]. handbook [Ano95-72, Ano95-80, Ano96-156, Ano98-61, Ano98-150, Ano96-42, Ano99-218, Ano99-219, Ano96-79, Ano96-82, Ano96-81, Ano96-80, Ano96-83, Ano97-64, Ano97-65, Ano98-67, Ano99-111, Ano99-112, Ano99-113, Ano99-114]. handling [Ano96-113]. hanging [NAKA98]. Hanif [Ano99l]. Hankel [CB93b, Feu91]. Hannon [Ano98-105]. Hans [Ano97-136, Ano98-91, Ano99-43]. Hans-Dieter [Ano97-136]. Hansen [Ano96-106]. Hanson [Ano95r, Ano95s, Ano96-39]. Hao [Ano97d]. Harald [Ano95i, Ano97-48]. Harasim [Ano96-99]. hard [Fre98, SKI92, TSLW98, Ano98-68]. hard-sphere [TSLW98]. hardback [Ano97n, Ano97-95, Ano97-90, Ano97-98, Ano97-99, Ano97-97, Ano97-94, Ano97-117, Ano97-162, Ano97-159, Ano97-129, Ano98k, Ano98-27, Ano98-70, Ano98-73, Ano98-128, Ano98-145, Ano98-146, Ano99-154, Ano99-155]. hardware [Ano96-150, Ano98l, Ano99-227, Pfi98, Ano98-69]. Hardware/software [Ano98-69]. Hardy [Ano92b, Jek94]. Hari [Ano96-53, Ano98-43]. Harmen [Ano99-120]. Harmonic [BR95b, GBB91, Gla90c, Jon96, MRZ96, Ran93, Ano99-115, Ano99-116]. Harnessing [Ano90s]. Harnett [Ano99-53, Ano99-54]. Harold [Ano97-76, Ano97-137]. Harrington [Ano96z]. Harris [Ano97-45, Ano99-111, Ano99-112]. Harry [DK97c]. Hart [Ano98i]. Harten [Ano99-172]. Hartley [HLS91]. Hartmut [Ano95-51, Ano99-89]. Haruo [Ano98-115]. Harvard [Ano96-57, Ano96-96, Ano96-105, Ano96-142]. harvest [LL99]. harvesting [LL94a]. Harvey [Ano97x, Ano97y, Ano97z]. Hassani [Ano99-149]. Hasselmo [Ano97g]. Hassoun [Ano96-74]. Hausdorff [CR90c]. having [MR94, Sun92]. Hawaii [LH94]. Hayashi [AD96]. hazardous [Ano99-38]. Hazewinkel [Ano98-49]. head [BU90]. Heal [Ano96-106, Ano96-107]. Heat [CH91, Hon99, Che91, DK97a, EGLN95, FH99, Far99, Fra96, GMR90, HM93b, KG94, Kha97, KM90, MG90c, Mur92a, Mur92b, MH93b, Rèp91, SM90, She99, SBS99]. heat-conduction [Far99]. Heath [Ano99x]. heating [Ha90]. Hebbian [Mar90]. Hébert [Ano99-203]. Heck [Ano97-71]. Heidelberg [Ano93o, Ano96-165, Ano97-140, Ano97-92, Ano97-111, Ano97-112, Ano98-93]. Heinrich [Ano99-182]. Heinz [Ano96-164, Ano99-89]. Heinz-Otto [Ano96-164, Ano99-89]. Helal [Ano97-163, Ano98-164]. held [Ano97-51, Ano98-84, Ano98-83, Ano98-101, Ano99t, Ano99-136, Ano99-137, Ano99-152, Ano99-172]. Hele [HR99, Hun95]. Helena [Ano98-46]. Hélène [Ano96v]. helices [HSH90a, HSH90b]. helicopter [CM93]. Hellinger [Faz94a, FL96]. Helmholtz [ABV97, IB95, MMZ91, Ano96-141]. Helmut [Ano95f, Ano98-93]. hemicontractive [Hua98c]. Hendler [Ano95-47]. Hendrik [Ano98-73]. Henrika [Ano97-128]. Henry [Ano96-76, Ano96-143, Ano97-77, Ano98-99]. Hentenryck [Ano95-45, Ano96-134, Ano97-110]. Herbert [Ano96-34, Ano96-89, Ano97-153]. hereditary [Ano98-115]. Heribert [Ano99-132]. Hermann [Ano96-141, Ano97-126]. Hermida [Ano99o]. Hermite [AW91, DCL^{+}94, Feu90, KTO97, NR96a, Ric95]. heteroclinic [Byk97]. heterogeneous [Alm96, Bla95, HM95, ML90b, Mic90]. heuristic [CHCC95, CCY95a, CCH97]. heuristics [Ano98-116]. Heusch [Ano95i]. hexadecahedrane [Kou90b]. hexagonal [BV99]. hexcadecahedrane [Kou89]. Hey [Ano99-81]. Hibbard [Ano99-77]. Hidden [Pal97]. Hidetoshi [Ano98-115]. Hierarchical [BP94a, HGWL96, Ano99-117]. Hierarchies [GWL96, Coh92, MGBP92, MR92]. hierarchy [Ano95-27, CB93a, CFLC95, DIT97, Hwa99, LH97, LL92, Lia99b, Sok91]. Hietaniemi [Ano99-148]. High [EDO96, Mac93, SM90, Ano96o, Ano98-139, Ano99-120, BS90, CS93b, CS94b, Cho94, HTT92, IB95, KO94, KM90, MR99, PC96b, Tsi99, Zha93b, SZM98, Ano96-86, Ano96-85, Ano96-84, Ano99-118, Ano99-119, Ano99-120]. High-accuracy [Mac93]. high-level [MR99, Ano96-84]. high-order [BS90, Tsi99]. high-performance [Ano96o]. Higher [BF97, GB90, MRJ97, AO99e, AG99b, Doš98, EA90, Gal93d, GB92, JP93, Köh95, Köh96a, Köh96b, Kre91, LR96b, SA94, TSL96, WA96a, ZKÖ98, Zaf98a, Zaf98b, ZY96, ZQ93b, ZQ93c]. Higher-dimensional [GB90, Köh95].

Higher-order [BF97, AO99e, AG99b, Doš98, SA94, ZKÖ98, Zaf98a]. highly [Ran93]. Hilbert [Ano99-181, Ahm91, Ame97a, ADKM97, AMZ98, BAS97, Dor90, GS96a, JY93, MO95, Sza91, Tse95]. Hill [Ano95-78, Ano96-102, Ano96-112, Ano97-67, Ano97-137, Kha90, AI97]. Hilliard [CC98b]. Hillston [Ano96b]. Hiltz [Ano96-99]. Hindley [Ano98s]. Hintikka [Ano98-90]. Hinton [Ano99-232]. HiQ(R) [Ano93-28]. Hiroaki [Ano95-47, Ano99-29, Ano99-30]. Hiroshi [Ano99-80]. histone [LNM90]. Historical [Ano95x, Ano97-127]. history [Ano95-29, Ano96x, Ano99f, Ano99h, Ano99g, Ano99i, Ano99-225, Ano97-127, Ano98-145]. Hiyan [Ano93u]. HLLE [Wen99]. Hoeve [Ano98-73]. Hohenberg [BGL97]. hole [BKK97]. holes [Ano98v, Ano99-91, Ano96-87]. holistic [Ano93r]. Holland [Ano93h]. Holtzman [Ano96-55]. Homage [Ano99-165]. Homoclinic [Shi97, GSST97, GST97]. homogeneous [AS93, ES97, Ker91a, LYH98, MB94b, SS94, Tan96]. homogenization [Shk97]. Homology [KMŚ98, Ano95-68]. homomorphisms [MR92]. homotopy [Ano95-68, KK91]. honest [YL93a]. Hooker [Ano99-175]. hope [Ano97-146]. Hopf [Ano98w, CST97, Cho92, Cho95]. Hopkins [Ano95-79]. Horgan [Ano97-31]. horizontal [CC94c]. hormone [HSH90b]. Horrell [Ano99-53, Ano99-54]. Horst [Ano99-111, Ano99-112]. Horstmann [Ano96-109, Ano97-27]. Horwood [Ano96-125, Ano99-59]. Hoskins [Ano99-59]. host [MP99]. host-parasite [MP99]. hot [RD96b]. hot-rolled [RD96b]. House [Ano97-120]. Hoyles [Ano93j]. HPN’98 [Ano99-120]. Hromkovic [Ano97t]. HTML [Ano96-88]. Hu [Ano96-49, Ano98-67]. Huan [Ano99-80]. Huang [Ano99-86]. Hudspeth [Ano98-52, Ano98-120]. huge [BS93]. hugh [Cho92]. Hughes [Ano99-41, Ano99-42]. Hull [Ano95-35, PLH95]. human [Ano97-144, GRC^{+}96, GPCSL90, Ano98-70]. Humour [Ano95-40]. Humphrey [Ano95e]. Hung [Ano98n]. Hunt [Ano96-54, Ano98-42]. Hunter [Ano99-139]. Hurtubise [Ano97-56]. Husbands [Ano95-36]. Huss [Ano96-118, Ano99-158, Ano99-160, Ano99-159, Ano99-161]. Huss-Lederman [Ano96-118, Ano99-158, Ano99-160, Ano99-159, Ano99-161]. Husty [Ano99s]. Hybrid [Che90a, Che93a, APV97, Ano98-50, CCD99, De 96, FMS92, GHPLZ90, PC91, Sim98, TWIF98, Zar92]. hydrides [Gre98b]. hydrocodes [HL99]. hydrodynamic [BCGP99, Che96, Dor97, DIR97, Kla98]. hydrodynamics [Dor98a, Dor98b]. hydrogen [DR96, Kor92]. Hydromagnetic [PSD96, RD96a]. Hyper [LW97, KAS99]. hyper-Bessel [KAS99]. Hyper-Hamilton [LW97]. hyperbolas [RHRZ90]. Hyperbolic [CCW96, CGH92, Chu92, EY91, Gla93c, Gla90c, Haw91, Hua95, Kan90b, KHI94, LM91, MM93b, Mur91b, QZ93c, Scr91, ST97, Zha91]. hypercomplex [Ano99-115, Ano99-116]. Hypercube [FLNW98, GH97, PC92, PS93b, LNG92]. hypercubes [CDH91, Chu92, CY93, LC97b, Mur91b]. hypergeometric [KS96, RB97, RSS94b]. hypermedia [Ano99n]. hypersonic [Edw92]. hypertext [Ano99n]. hypertournament [RM98b]. hypo [Alt90b]. hypo-quadratic [Alt90b]. hypoquadratic [AB94]. Hypothesis [Kor90a, Kor91, AAL99, Nam97, Nam99a, Sch90]. hysteretic [TBA99].

Ian [Ano95q, Ano95-50, Ano96-96, Ano97-28, Ano99-93]. Iasnogorodski [Ano99-198]. IBM [Ano95o, Ano99-37]. Ibragimov [Ano99-71]. Icosahedral [KC92b, MTLGFG93]. Icove [Ano96-42]. Idan [Ano98-104]. ideal [Gla92a]. Ideas [Ano99-199, Ano99-200, AC95, Ano96-65]. Idempotent [Ano97-66]. identical [CD98]. Identifiability [CKK97a]. Identification [BM90a, HM93a, Hin95, Los89a, ZM98, Ano98-101, Bag90, KU91, MM93b, MM95a, Mur92a, MH93b, SI94, Ano96v]. identities [JKS94, PM95b]. identity [dJF91]. IEEE [Ano99-128]. IEEE/IFIP [Ano99-128]. IFIP [Ano99-120, Ano99-128, Ano99-212]. Ignition [WT97]. IHCP [LM96a, MM96c, MZ95, ZM98, ZM96]. II [Ano95r, Ano95s, Ano98-63, Ano98-155, Ano99-103, Ano99-104, Noo98, TL94, AR92d, BD96, DIR97, Dor98b, Gla97b, Has95, Heg91b, Hin95, Hua95, Kan90b, Köh96a, MC99, MTLGFG93, Noo95, Sha99a, WU94, Wit96b, WA94a, WA96c, Ano93g, Ano95-38]. III [Ano96-39, Ano96-147, Dor99, Köh96b, MTLGFG95a, MR94, Ano99p]. IL/Cambridge [Ano97-90]. ILAPs [Ano97-68]. ill [DEMS94, Sha90c]. ill-conditioned [Sha90c]. ill-posed [DEMS94]. illustrated [Ano97o, Ano98p]. IMACS [Ano90a]. image [Ano98-76, DGU90, GCLT98, KU99, Li96c, Li98d, SHK99]. images [Sad96b, Ano93h]. imaginaries [Ano95-69]. imagination [Ano97-151]. imaging [OOWD98]. imbedded [Sim98]. imbedding [NL91, SV91a, SV91b, UW91, UW94, UWT99, Wan90, WU99a]. Immerman [Ano99-60]. immersed [Li96b]. Impact [Azb99]. impedance [MRZ96]. imperative [Ano96k]. imperfect [Köh91]. Imperial [Ano99-32]. Implementation [BR95b, Ano97-98, Ano97-99, Ano97-97, Ano98-106, Ber93, BRZS97, Con95, DBM95, MNB93, Mis90, TCS94, Wan96, Ano98-31, Ano98-71]. implementations [Dav96]. Implementing [BP94b]. implication [BKM91]. implications [Gup94]. Implicit [Bar93, AKV91, Bog99, CC94a, CS90, Cub96, Din99a, Din99b, GPZ93, Hua98a, HLTB99, JV96, LW91, MG95, Mey99, MRJ97, SK91, VK99, Wac94, WY99]. implicit-explicit [Bog99]. implicitly [BR94, Bro98, FK98b, Rhe96]. importance [Ano99-215]. important [CCC^{+}98]. imposition [Lay99]. improve [SK91]. Improved [APB97, BW96b, HLS91, KC92a, Kou98, Pan94, BM92, HH93, Koç91, PC93, Sha90c, Ye99, Zha93a]. Improvement [FK98a, BW92, Hwa99, LS96a, LS98b, Pan95]. Improving [Huj96, Rán92, Ric95, vDHO^{+}95, Ano96-151, NMUY92, Ano96-89]. impulse [Bou98, GL94]. impulsive [CYS94, KB90, LZ98, LG99, MB90, RT99b, TCS94]. Inc./John [Ano99-28]. incentive [Ano97-37]. incidence [Ano96-82]. included [Ano95-51, Ano95-50, Ano95-49, Ano96-59, Ano96q, Ano96-120, Ano96-147, Ano96-149, Ano97o, Ano97q, Ano97w, Ano97-30, Ano97-73, Ano97-67, Ano97-78, Ano97-145, Ano97-79, Ano97-92, Ano97-96, Ano97-147, Ano97-111, Ano97-101, Ano97-112, Ano97-102, Ano97-103, Ano97-113, Ano97-124, Ano97-131, Ano98p, Ano98t, Ano98-44, Ano98-46, Ano98-57, Ano98-80, Ano98-122, Ano98-125, Ano98-143, Ano98-160, Ano99y, Ano99-35, Ano99-36, Ano99-37, Ano99-53, Ano99-54, Ano99-69, Ano99-70, Ano99-129, Ano99-130, Ano99-154, Ano99-155, Ano99-147, Ano99-177, Ano99-179, Ano99-180, Ano99-230, Ano99-227, Ano98-105]. Includes [Ano98-41, Ano98-149, Ano99-221, Ano98-162, Ano99-78, Ano99-79]. Including [Ano96f, Ano99-181, Ano96-112, Ano96-159, Ano97-72, PC96a]. inclusion [Fil93, Med90, Men91, MSR95, PC93, ZO92]. inclusions [AO99d, Cha99b, Din99b, Hua98a, Hua98b, Lei96]. income [Ueb92]. incomplete [Ano95-60, Ano95-62, Ara96, CDH91, Luo93b, Mil92, Nis97b]. incompressible [Ano99-70, Ber92, EL96, HW96, LLP94, LS99, LMW95, Mei93, SCBFC93, YX95]. Increasing [Som90, Gol94, HWL94]. indefinite [EB91, Luo93c, Par91a, YK95]. independence [Ves98, Yos93]. Independent [Nak99, APB97, LS96b, Ano99-121]. Index [Ano92g, Ano90-29, Ano90-30, Ano91l, Ano92h, Ano93-29, Ano95a, Ano95-87, Ano96-172, Ano96-173, Ano97a, Ano97b, Ano98a, Ano98b, Ano99a, Ano99-248, Bal96a, Bal96b, Gün99, MT94a, MPR98]. index-1 [MPR98]. index-2 [MT94a]. indicators [Ano96x, Mit92a]. Indies [Ano98-83]. indifference [LO93]. individual [Ano98-159, She95]. induced [Ahm91, ASW98, MBMS92, WK94]. Inductive [Ano96-90]. Industrial [Ano99-150, Ano99-74, Ano99-93]. industrially [CCC^{+}98]. industrially-important [CCC^{+}98]. industry [Ano95o, Ano95u, Ano97-37]. inelasticity [Ano99-41, Ano99-42]. inequalities [APB97, AWY97, AD98, Ano95-83, Ano97-30, BS96d, Bou98, CTY99, Cub95a, CY95, CLY97, CH99, Din97a, Din99a, DL99, Din99c, Din99d, GY93, Hua93, HLTB99, JY93, Kor96, Lei96, MSB90, MSR95, Noo95, Noo96, NAS97, Noo98, Noo99, Pac94, PA99, Ver99a, WX97, Wu99b, WY99, Yao92, YLY96, Ano98-135]. inequality [AD96, Ano99-170, CY94, DW97, DDP99, Dra99, GlZ95, JTY97, WW98, WX97, WY99, Yao93, YY98]. inertia [Ano97-134]. inetgro [GSC98]. inetgro-differential [GSC98]. inexact [Alt90c, Arg99, KN94, Los89a]. inf [Sto99]. inf-sup [Sto99]. infeasible [SWF95]. infeasible-interior-point [SWF95]. infections [Voi90]. inference [Ano95y, Ano99-175]. Inferring [DT93]. infiltration [CKK97b]. infinite [AO97, AO98a, Ano97-140, Ano98-134, Bal99, Ber98, CG94, CK94b, DBA98, FW96, Fin99, GM96, JG95, JG96, KMD99, LW99a, Mat99, Pok97, SWF95, Tad94, Tre95b, Tse95, Wan91, WF96, WK99, XLG95, YSM92]. infinite-dimensional [GM96, Tad94, XLG95]. infinitely [WF99b]. infinity [Sim93b]. inflexible [Mil92]. influence [BS99b]. infobahn [Ano96-30]. information [AGM99, Ano93f, Ano95x, Ano95-46, Ano96i, Ano96-150, Ano96u, Ano96-45, Ano96-52, Ano96m, Ano96-148, Ano97g, Ano98h, Ano98-27, Ano98-77, Ano98-109, Ano98-132, Ano99q, Ano99-50, Ano99-216, Ano99-217, Ano99-128, Ano99-145, Cho99, DDP99, HSH90a, LL92, MM96a, Mil92, Pol90, Ter00, ZM94, Ano95-62, Ano98-71, Ano98-73, Ano98-72, Ano99-122]. information-theoretic [Ano96m]. infrastructure [Ano96-148]. infusion [Oll90]. Ingo [Ano97-139]. inheritance [Pad92a, Tho92]. Inhibited [KKB97]. inhibition [SHT^{+}90]. inhomogeneous [Bed95b, EGLN95]. inhomogeneously [Kan92a]. Initial [But94, Car93, Szm90, Ahm95, AS96a, BR95c, BA98, Car92, EDNO92, EGLN95, HNS^{+}90, JG98, JPC99, LD90, LNG92, MM95a, MG90c, MH93b, Riz95, Sim93b, Sim93a, Tsi99, vdHS94, vdV95, Ano99-123, Ano99-124]. initial-boundary [BR95c, HNS^{+}90]. Initial-Value [Car93, AS96a, Car92, Riz95, Sim93b, Sim93a]. injection [XLJS98]. inner [MT98, Sni91a, WSDT94]. innovation [Ano98-82, Kor91]. innovator [Ano99-83]. input [MW99, Ano90t]. input-output [MW99]. inquiries [Pag90]. inquiry [Ano98-48]. insights [Ano99-215]. inspection [IN92]. inspirator [Ano99-83]. instabilities [KKK^{+}97]. Instability [Ano97-42, Dea98, KKB97, KKG97, KG95b, RD96a]. installation [Bes96]. Instantons [Mur93b]. instigator [Ano99-83]. Institute [Ano99-28]. institution [Ano99-239]. instruction [CCK97]. instructions [BD96, DB96]. instrument [Bec96, Far96b]. insulated [HW90]. insulin [Oll90]. insurance [Ano98-148]. integer [Ano97-84, Ano98-154, CCK97, CLD97, EB91, Gal91a, LTC98, TI96, Ano99-127]. Integrability [BMN96, LPR96]. integrable [AH97, Lee93, YCZ98]. Integral [GZ91, Kro99, TZ93, APB97, AO99c, Ano99-182, Ber93, CLM92, Dra99, ES97, ESKG98, EGHS98, Gás95, HM99, IZ97, Iwa91, Kim98, Kim99, KAS99, Lar90, LM91, LC96b, MS93, MH99, MO98b, MO98a, MGM92, O’R95, Ola90, Ram91a, Ram93, RL91, Ram91e, RS90, Sch99, She92, Shi95b, Shu93a, SZ91, Tch98, YLC96, Zha91, Zhe93, Ano98-74, Ano98-75]. integrals [CLM90, CLM92, Dat93, EGK96, Feu91, GKT99, Gro95, HS95, Has97, KT96, KK97a, Mar94, Mil98, PDK96, Sib98, Sri95, Zha91]. integrands [DR97]. integrated [Cha93, Lyt92b, Ano96-43]. Integrating [Pen96, BCC95, HS94, Li96c]. integration [And91, And93, Ano97-138, Ber93, DBM94, DR97, Gro95, KS92, Kar96a, Leo96, LQ95a, Sim93a, SM94, Sim98, SH95b, dCVC97, Web95]. integrator [LQ95b]. integrators [Kar96b]. integro [AWY97, CRV98, GS91, KP99b, LG99, LM91, RZ95]. integro-differential [AWY97, CRV98, GS91, KP99b, LG99, LM91, RZ95]. integrodifferential [LCC99, SMU98]. intel [LSH98, OOWD98, CCC^{+}98, LNG92]. intelligence [Ano90o, Ano93f, Ano96-36, Ano99-37, BV95, Gir92, SS90, Ano95-47, Ano96-61, Ano96-113, Ano99-32]. Intelligent [Far96b, LR90, Ano95l, Ano97-36, Ano98-50, Dre90, Jay90, Jos90, Sti96, Ano98-77, Ano98-76, Ano99-128]. intelligently [Ano93i]. intensity [DPD98]. interacting [Kos90]. Interaction [DD99, Ano96-40, Gre91b, LR90, RS94]. interactions [KM93]. Interactive [Dei91, SNU98, Sco96, Bec96, BKM95, KBLS90, MM96d, Nan90, Pen96, Ano97-67, Ano99-129, Ano99-130, Ano99-131]. interactivity [Ano99-44]. interarrival [EA92]. interative [Gás95]. interchangeable [YI97]. interconnection [Ano97-161, HTT92]. Interdisciplinary [Ano97-68]. interface [Ano95-72, Ano99-117, Ioa91, Li96b, Mol96, Mur92a, Sau95, YX95, Ano95-82, Ano96-128]. interface-oriented [Ano99-117]. Interfacial [ARI96, DPD98, DD99]. Interior [LS98b, LM96a, Més96, SWF95, Sny90, Ano98-78]. Interlacing [MO95]. interlocutors [Ano97-50]. internal [Son92]. internals [Ano98-163]. International [Ano90b, Ano96-166, Ano99-128, Ano99-212, Ano90-28, Ano95k, Ano95-36, Ano95-45, Ano96e, Ano96-104, Ano96-103, Ano96-135, Ano97-53, Ano97-47, Ano97-88, Ano98-38, Ano98-96, Ano98-94, Ano98-95, Ano99-29, Ano99-30, Ano99-90, Ano99r, Ano99-120, Mit92a]. Internet [Ano95-46, Ano96w, Ano96-133, Ano96-77, Ano97-101, Ano97-102, Ano97-103, Ano98-112, Ano97-69, Ano98-79]. interoperability [Ano98-161]. interplay [Sch92b]. interpolant [Fol92]. interpolants [LS96a, LS98b, Luo98, Sch96, Sim93c, Sim93a]. Interpolating [BS90, Den91b]. Interpolation [CF92, CQ98, Fli95, Har92b, NG91, She95, XC92, AW91, AL93, BC97, BR95c, BM97, BMB94, Bon97, Boy99, BT97, BM92, CF91a, CN94, CWJ92, DBM92, DBM94, DBM95, FS95, Feu90, Gem93, ISM93, JS95, KC92a, KTO97, KM97b, LTB96, xLhW96, MN93, MBV91, PSLT93, Pan98c, PDK96, Qu96, QA96, RK98, Ram95, Sal92a, Sau92, Sun92, VBM90, WA94a, WA94b, WA96c]. Interpolation-based [Fli95]. interpolators [Mad92]. interpolatory [DR97, FLR92, QA96]. interpretation [ACHP94, Ano96-140, Ano97-137]. Interpretational [Neu91]. interpreter [Abe90]. intersection [Alf95, Ano96-171]. Intersections [Ano95s, POH99]. Intertemporal [GG92]. Interval [LR96b, RR91, Ano98-32, BM96, BKK99, BM97, Boy99, CG94, CC96, Dia91, EDO93, Hua97, LZ92, Mar90]. intervals [AO97, AO98a, CK94b, FF92, Ioa91]. intracellular [Hol90]. intramolluscan [KC91]. intrinsic [AS97, Sok91]. Introduction [Ano96-109, EHK92, Fus90, Ano95j, Ano96p, Ano96q, Ano96r, Ano96n, Ano96o, Ano96-95, Ano96-111, Ano96-146, Ano97i, Ano97k, Ano97j, Ano99-59, Ano99-149, Ano99-190, Ano99-191, Rab92, SH96, Ano90p, Ano97-113, Ano97-114, Ano97-72, Ano97-71, Ano97-70, Ano97-73, Ano98-81, Ano98-80, Ano99-133, Ano99-134, Ano99-132]. Introductory [Ano96-91]. invariance [Kup91]. Invariant [JGB95, NL91, RZ96, UW91, UW94, UWT99, Wan90, WU99a, AH97, Jon96, LS98a, Shu93a, SV91a, SV91b, Ter00, ZM94]. invariants [Ano97-30, Mro96, Win91]. invasion [CC99]. inventory [CJ90, DKO93, DO95, KS97a, Oda94, RG90a]. Inverse [Eti91, AJ95, BK91, BG91b, EGHS98, Fra96, GMR90, Gur91, Ike91, Luo93a, Luo93b, MMZ91, MRZ96, MG90c, Mur92b, MH93b, PDK96, PL99, RW91, Rèp91, RS92, She99, Sun91, Sym91, VPK99, Wan93a, WU99a, YP98a, YP98b, ZR91]. inverse-free [EGHS98]. inverses [BLR90, Kim99, UK99]. Inversion [EG97, Ram91c, Usm94, FK98b, GS96a, HLS91, HWL94, Mat99, MGM92]. inverted [Ano99-245, FLW97]. Invertibly [Tre95b]. inverting [FK98a, Luo93c]. Investigation [Tel91, Ano97-134]. investigations [Ker91b]. Investing [Ano98-82]. investment [HL90, KM92, Mil92]. invexity [Cra98]. inviscid [Gla90a, Gla94e]. invisible [Ano99-216, Ano99-217]. Invitation [Ano99-135]. involving [Arg99, Byk97, CIS98, KAD99, KS96, Liu96a, Liu96b, LD93, Mil98, Mis90, Sri95]. inward [AO99b]. Ionescu [Ano99-98]. ionized [KKK^{+}97]. ions [Ano98-104, RS93].

Iosif [Ano97-105]. IRb [Szk95]. IRb-6 [Szk95]. irreducibility [ZL99c]. irregular [Kan92a, Ruz90]. Irregularity [HO97]. irreversibility [AS97, HPH^{+}94]. Irv [Ano96-150]. Irving [Ano95-32, Ano95-43]. is-a [MNE92]. Isabel [Ano99-63, Ano99-64]. ISDN [Ano95-79]. isentropic [Gla90b, Gla91b]. ISETL [Ano97-70]. Ishikawa [Cha99b, Hua98c, Hua98a, Hua99]. Isi [Ano98-128]. Isis [Ano97-154]. island [Ano96-96]. ISO [Ano98-61]. ISO/ANSI [Ano98-61]. isochronous [CCGL99]. isodiastemic [Köh96c]. isolation [JI94]. Isosymmetric [Köh93a, Köh95, Köh96a, Köh96b]. Israel [Ano99-136]. issue [Ano90c, CCK97]. Issues [Ano97-163, Pag90, Ano96-85, Erm92, GG92, KSM94, Ano98-39, Ano98-47]. Ister [Ano98-74]. Iterated [Dia91]. Iteration [BWK97, Kun90, Bai96b, BW92, Bin91, Hua97, Hua98c, LW91, Ohn92, Oza94, Pan92a, Wac94]. iterations [HK96, Hua99, MM93c, MM94, Mil95]. Iterative [Ame97a, Ame97b, BD90, BS95, BS96b, BP97, Din96, Din97b, GPGCM94, Iwa91, MM96b, NAS97, PW97, Sol97, BSW96, Cha99b, Fin99, Gol90, Hua98a, HB99, LM96b, MV95b, MG97, MRWM94, Oza94, PA94, Pom92, RG90c, Scr91, Ser93, SFT99, SLC97, Sir97, Zha93c, dOF91]. Ito [Ano98-136]. IUTAM [Ano98-84, Ano98-83, Ano99-136, Ano99-137, Ano98-84, Ano98-83, Ano99-136, Ano99-137]. IV [Ano96-67, Ano96-83, MTLGFG95b, Ano95k, Ano97-107]. IVPs [Söv99]. IWTCS’98 [Ano99-212]. Iyengar [Ano99-133, Ano99-134]. Iyer [Ano98-65].

J [Ano90v, Ano93t, Ano93v, Ano93w, Ano95m, Ano95d, Ano95-73, Ano95-44, Ano95-61, Ano95-62, Ano95-65, Ano95-68, Ano96f, Ano96u, Ano96-30, Ano96-40, Ano96-54, Ano96-68, Ano96-71, Ano96-83, Ano96-89, Ano96o, Ano96-108, Ano97-53, Ano97o, Ano97e, Ano97-142, Ano97-43, Ano97-54, Ano97-49, Ano97-65, Ano97-72, Ano97-106, Ano97-132, Ano97-133, Ano97-165, Ano98p, Ano98s, Ano98-29, Ano98-42, Ano98-52, Ano98-64, Ano98-101, Ano98-120, Ano98-141, Ano98-157, Ano98-155, Ano99u, Ano99p, Ano99-41, Ano99-47, Ano99-42, Ano99-48, Ano99e, Ano99d, Ano99-45, Ano99-52, Ano99-74, Ano99-83, Ano99-81, Ano99-106, Ano99-107, Ano99-113, Ano99-136, Ano99-172, Ano99-220, Ano99-190, Ano99-191, Ano99-205, Ano99-211, Ano99-232]. J.-R [Ano97-142]. Jaakko [Ano98-90]. Jack [Ano95-59, Ano96-118, Ano99-158, Ano99-159]. Jackendoff [Ano96-97]. Jackson [Ano96-71]. Jacky [Ano97-159]. Jaco [Ano96-46]. Jacobi [LQ95b, MG97, PDK96, SSH92, SG95b, Sri95, Usm94, VL95, Xu97]. Jacobi-type [MG97]. Jacobian [PDK96, VPK99, Zed90]. Jacobson [Ano99-122]. Jacques [Ano96-31, Ano97-56, Ano97-136, Ano98l]. Jadran [Ano99s]. Jaffar [Ano98-95]. Jajodia [Ano98g]. James [Ano95-47, Ano96-79, Ano96p, Ano96-110, Ano97-120, Ano98-46, Ano98-82, Ano98-160, Ano99-38, Ano99-53, Ano99-54]. jams [Ano97-160]. Jan [Ano96z, Ano98-73, Ano98-94, Ano99-145, Ano99-229]. Jane [Ano96b]. Janeiro [Ano97-51]. Janet [Ano96-115, Ano96-148]. Janette [Ano95-73]. Janice [Ano96-53]. János [Ano98-113, Ano96-33]. January [Ano97-51, Ano98-83]. Japanese [VA94]. Jari [Ano98-156]. Jarkko [Ano99-148]. Jaromír [Ano95-33]. Jaroslav [Ano99-135]. Jarratt [Arg98, EGHS98]. Jarratt-type [EGHS98]. Jasanoff [Ano96-142]. Jason [Ano99-139]. jaundice [YS90]. Java [Ano96-92, Ano97-38, Ano97-98, Ano98f, Ano98-87, Ano97-35, Ano96-70, Ano98-54, Ano98-106, Ano96-92, Ano97-75, Ano97-77, Ano97-76, Ano97-74, Ano97-78].

Javabeans^{T M} [Ano99-72]. JavaScript [Ano96-93, Ano98-85]. Java^{T M} [Ano98-88, Ano98-87, Ano98-86, Ano99-139, Ano99-138]. Jay [Ano97-137, Ano98-162, Ano99-164]. Jaynes [Los92]. JDBC [Ano97-35]. Jean [Ano95-36, Ano96-34, Ano97-53, Ano97-132, Ano98l, Ano98k, Ano98-140, Ano99-90, Ano99t, Ano99-39, Ano99-95]. Jean-Arcady [Ano95-36, Ano96-34, Ano97-53, Ano99-90]. Jean-Daniel [Ano98k]. Jean-Luc [Ano99-39]. Jean-Michel [Ano98l]. Jean-Pierre [Ano99-95]. Jeanne [Ano98-61]. Jed [Ano96-144]. Jeff [Ano99-210]. Jeffrey [Ano97-93, Ano97-89, Ano98-45]. Jennifer [Ano96-51]. Jerrold [Ano98-61]. Jerry [Ano95-46, Ano96y, Ano97-60]. Jersey [Ano93k, Ano93n, Ano97-116]. Jerzy [Ano98-72]. Jessup [Ano96o]. jet [ZSDB99]. jets [PB95]. Jewels [Ano98-87]. Jiang [Ano95-59, Ano97-163]. Jianming [Ano96-37]. Jim [Ano97-140]. Jimmy [Ano99-131]. Jin [Ano96-37]. Jirf [Ano99-135]. Jiri [Ano99-171]. JiríRohn [Ano98-32]. Joan [Ano96-77, Ano96-127, Ano98-140]. Job [ML90a]. jobs [LS96b]. Jochem [Ano99-171]. Joel [Ano95-41, Ano95-49]. Johan [Ano97-64, Ano99-169]. Johanna [Ano99-144]. John [Ano93n, Ano93p, Ano93v, Ano95u, Ano95c, Ano95-33, Ano95-38, Ano95-78, Ano95-39, Ano95-81, Ano96g, Ano96-150, Ano96y, Ano96v, Ano96-37, Ano96-27, Ano96-45, Ano96z, Ano96-33, Ano96-41, Ano96-153, Ano96-38, Ano96-54, Ano96-67, Ano96-73, Ano96-75, Ano96-76, Ano96-77, Ano96-91, Ano96-103, Ano96-101, Ano96-109, Ano96-116, Ano96-117, Ano96-120, Ano96-110, Ano96-125, Ano96-124, Ano96-126, Ano96-128, Ano96-131, Ano96-146, Ano96-147, Ano96-143, Ano96-164, Ano96d, Ano96-167, Ano97o, Ano97h, Ano97-31, Ano97-27, Ano97-30, Ano97-34, Ano97-54, Ano97-57, Ano97-61, Ano97-60, Ano97-72, Ano97-74, Ano97-145, Ano97-87, Ano97-106, Ano97-121, Ano97-125, Ano97-132, Ano98m, Ano98p, Ano98-29, Ano98-30, Ano98-42, Ano98-60, Ano98-57, Ano98-64, Ano98-78, Ano98-84, Ano98-81, Ano98-117, Ano98-141, Ano98-142, Ano98-143, Ano98-154, Ano98-159, Ano98-161, Ano99-28]. John [Ano99-32, Ano99-34, Ano99-47, Ano99-48, Ano99-45, Ano99-53, Ano99-54, Ano99-65, Ano99-58, Ano99-71, Ano99-74, Ano99-98, Ano99-111, Ano99-112, Ano99-133, Ano99-134, Ano99-127, Ano99-154, Ano99-155, Ano99j, Ano99-151, Ano99-148, Ano99-175, Ano99-193, Ano99-190, Ano99-194, Ano99-191, Ano99-199, Ano99-200, Ano99-228]. Johs [Ano98-111]. join [Rei94, Rei99a]. joint [Ano97-88, Ano98-95, BS99b, JLL^{+}93, Tsi91]. Jointly [AK90a]. joints [TCJG91]. Jon [Ano97-78, Ano97-162, Ano99-148]. Jonathan [Ano97-159, Ano98-86, Ano99-78, Ano99-79, Ano99-138]. Jones [Ano95-46, Ano97u, Ano97-164]. Joos [Ano99-169]. Jordan [Ano97-53, Ano98h, Dri97, HV92, Var98, Ano98-89]. Jörg [Ano98z]. Jörgen [Ano96-66, Ano98-69]. José [Ano95r, Ano95s, Ano96-39]. Joseph [Ano90s, Ano96k, Ano96-138, Ano96-164, Ano97-73, Ano97-69, Ano98-80, Ano99-55, Ano99-56]. Joshi [Ano97-163]. Joshua [Ano96-70, Ano97-63, Ano97-107, Ano98-54]. Journal [Ano96-166, Ano97-128]. journey [Ano97d]. Joxan [Ano98-95]. joy [Ano97-145]. Jr [Ano95h, Ano95-73, Ano96-75, Ano96-91, Ano96-130, Ano97f, Ano97-161, Ano98-30, Ano99-203]. Juan [Ano99o, Ano99-95]. Juan-Enrique [Ano99-95]. Jude [Ano96-39]. judgment [Ano99-140]. Judson [Ano96-59]. Julian [Ano96-44]. Juliana [Ano96-149]. Julie [Ano97-137, Ano99-180]. Julita [Ano99n]. Julius [Ano98-117]. July [Ano97-47, Ano98-84, Ano98-111, Ano99t]. jump [BG95]. June [Ano96-135, Ano99t]. Juraj [Ano97t]. Jürgen [Ano98-91]. Jürgens [Ano99-89]. Jüri [Ano97-108]. Just [Ano98-63]. Justification [Nam97, RF91, ES97]. juvenile [KL96b].

Kaeli [Ano98-56]. Kahin [Ano96-148]. Kahl [Ano98-32]. Kalle [Ano99-195]. Kalman [Ter00, BBM95, Ram90b, SGS93, Sch92b, Sza90b, ZM94]. Kalman-filtering [Ter00, ZM94]. Kalman-tracking [Ram90b]. Kalmbach [Ano99-196]. Kalyuzhnyi [Ano99-115, Ano99-116]. Kam [Ano93i, Ano93i]. Kambiz [Ano96-147]. Kamin [Ano96q]. Kant [Ano96a, Ano99-140]. Kantorovich [GlZ95, Han92]. Kaplan [Ano99-151]. Kaplansky [Ano95-32, Ano95-43]. Kapteyn [DTLM98]. Karin [Ano95i]. Karl [Ano96-42]. Kármán [CKM99]. Karmarkar [Tor90]. Karniadakis [Ano99-205]. Karp [Ano97-168]. Karpinski [Ano98-55]. Kasabov [Ano97-52]. Katavolos [Ano98-121]. Katherine [Ano97-40]. Kathleen [Ano97-83]. Kathy [Ano96-158]. Kaufman [Ano95-63]. Kaushik [Ano97l]. Kautz [SG95a]. Kayne [Ano95-70]. Kazume [Ano97-96]. KBSs [LRLR95]. KdV [VHH95]. Kearns [Ano99q]. Keating [Ano99-201]. Kee [Ano98-31]. Keener [Ano96j]. Keller [Ano98-82]. Kelly [Ano95-71]. Kelly-Bootle [Ano95-71]. Kelso [Ano97-43]. Kemper [Ano96-120]. Ken [Ano95-37, Ano98-63]. Kennedy [Ano96-88]. Kenneth [Ano93i, Ano95h, Ano96-126, Ano97f, Ano99-77, Ano99-214]. Kepler [Ano96-163]. Kerkrade [Ano98-101]. Kernel [KHI94, CL99c, ES97, FLR95]. kernels [DIT97, Mil98]. Kerschberg [Ano98g]. Kersnar [Ano97-102]. Kevin [Ano96-54, Ano98-42, Ano99-230]. Key [DHW98, SH94, Hwa99, LWL93, WY93]. Keyton [Ano99-31]. Kharazishvili [Ano99z, Ano99-27]. Khatib [Ano93v]. Khattree [Ano99-28]. Khordoc [Ano99-117]. Khosla [Ano98-50]. Khuri [Ano98-142]. Kichoon [Ano99-153]. Kieran [Ano97-143]. kill [ML90b]. Killelea [Ano99-235, Ano99-236]. Kimball [MM96e]. kind [BL96, KAS99, KS97b, PDK96, Tse95, Wu99b]. kindi [Ano96-163]. kinds [MZMQ90, ZQ93d]. kinematical [Pap94b]. kinematics [Ano99s, Szk95]. kinetic [Bab98, BJH90, BH90, BCC95, DKG98, Gre92b, Kla98]. Kinetics [KM93, Aug90, Ber96]. King [Ano99-164]. Kink [Bas98]. Kinnear [Ano95h, Ano97f]. kinodynamic [RT94]. Kiong [Ano98-31]. Kip [Ano99-38]. Kirchhoff [BM96]. Kirsh [Ano96-158]. Kirsh-Pasek [Ano96-158]. Kishan [Ano97-44]. kit [Rot96]. Kitano [Ano95-47, Ano99-29, Ano99-30]. KKM [Ano98-59, Ver99a]. KKM-map [Ano98-59]. KL [WS92]. Klas [Ano98-156]. Klatte [Ano93o]. Klaus [Ano97-125]. Klein [Ado91c]. Klimek [Ano98-44]. Kline [Ano99-230]. Kluge [Ano96-120]. Kluwer [Ano93m, Ano93q, Ano97e, Ano97-29, Ano97-36, Ano97-45, Ano97-58, Ano97-56, Ano97-66, Ano97-85, Ano97-84, Ano97-108, Ano97-105, Ano97-104, Ano97-126, Ano97-157, Ano97-164, Ano97-163, Ano98g, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98w, Ano98x, Ano98r, Ano98z, Ano98-37, Ano98-31, Ano98-32, Ano98-36, Ano98-45, Ano98-49, Ano98-53, Ano98-50, Ano98-47, Ano98-59, Ano98-56, Ano98-57, Ano98-65, Ano98-66, Ano98-68, Ano98-67, Ano98-69, Ano98-71, Ano98-72, Ano98-76, Ano98-84, Ano98-74, Ano98-83, Ano98-89, Ano98-90, Ano98-92, Ano98-109, Ano98-108, Ano98-100, Ano98-101, Ano98-110, Ano98-115, Ano98-111, Ano98-116, Ano98-121, Ano98-123, Ano98-126, Ano98-127, Ano98-133, Ano98-137, Ano98-132, Ano98-135, Ano98-144, Ano98-138, Ano98-139, Ano98-140, Ano98-153, Ano98-157, Ano98-155, Ano98-156, Ano98-160, Ano98-161, Ano98-164, Ano99z, Ano99-27, Ano99o]. Kluwer [Ano99r, Ano99s, Ano99n, Ano99t, Ano99u, Ano99-33, Ano99-50, Ano99-213, Ano99-49, Ano99-63, Ano99-64, Ano99-52, Ano99-57, Ano99-58, Ano99-84, Ano99-85, Ano99-86, Ano99-83, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-87, Ano99-105, Ano99-103, Ano99-99, Ano99-96, Ano99-102, Ano99-104, Ano99-100, Ano99-97, Ano99-95, Ano99-101, Ano99-115, Ano99-116, Ano99-117, Ano99-120, Ano99-114, Ano99-128, Ano99-121, Ano99-136, Ano99-137, Ano99-131, Ano99-145, Ano99-142, Ano99-152, Ano99-153, Ano99-167, Ano99-168, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-170, Ano99-174, Ano99-183, Ano99-184, Ano99-181, Ano99-182, Ano99-196, Ano99-197, Ano99-201, Ano99l, Ano99-206, Ano99-207, Ano99-204, Ano99-209, Ano99-229, Ano99-212, Ano99-211, Ano99-227, Ano99-234]. knapsack [JW93]. Knowing [Ano96-115, Ano96-115, Ano96-94]. Knowledge [Las96, Ano90o, Ano95-60, Ano96h, Ano96-113, Ano96-143, Ano97-55, Ano97-52, Ano98-139, Ano99-52, Ano99-88, Ano99h, Ano99i, Ano99-142, FGN90, Pet92, Sow92, Wil92, Zar92, Ano96-64, Ano96-138, Ano96-95]. knowledge-based [Ano90o, Pet92]. knowledgebase [BNR96]. known [BG91a]. Knudsen [Ano98-86, Ano99-138]. Knuth [Ano97-129]. Ko [BL97]. Kobsa [Ano99n]. Koch [Ano95-41, Ano98-104]. Kocvara [Ano99-171]. Kohler [Ano97-128]. Kohonen [FT97b]. Kolmogorov [Ano97j]. Kolokoltsov [Ano97-66]. Koman [Ano95-80]. Komlósi [Ano99-165]. Komrakov [Ano98-92]. KOND [KHI94]. Kopetz [Ano97-126]. Kopka [Ano95f]. Korn [Ano96-122]. Korteweg [Ado91c, SVS98, Sch94]. Kostantinos [Ano96-149]. Kothanda [Ano99-87]. Kotov [Ano93h]. Kou [Ano97-104]. Kowalski [Ano98-71]. Koza [Ano95-38, Ano97-57]. Kozen [Ano98q]. Krakatau [Ano96-96]. Kramer [GS95, Ano98-138]. Krantz [Ano95g]. Krasil’shchik [Ano98-92]. Krayola [SH96]. Kreines [Ano99-176]. Kreinovich [Ano98n, Ano98-32]. Kreiss [Ano96-164]. Kreveld [Ano98-33]. Kriging [Mye92]. Krishna [Ano98-43]. Krithi [Ano99-58]. Kronecker [JAK96, Róz99]. Kronecker-product [JAK96]. Kronstadt [Ano96-76]. Krueger [Ano98-136]. Krylov [SB99]. Krzysztof [Ano96-114, Ano97-140, Ano98-144, Ano99-52]. Kuchcinski [Ano98-144]. Kuerten [Ano99-83]. Kuhn [Ano97-158]. Kuiken [Ano99-83]. Kuipers [Ano95-60]. Kuklick [Ano97-128]. Kulisch [Ano93o]. Kumar [Ano99-57]. Kunz [Ano98-133]. Kuramoto [MMP98]. Kurshan [Ano95t]. Kutta [CT96, Pat98, QZ91, Sim93a, hC96a, hC96b, AS96a, BS90, BS96a, Bog96, BT98, CMR90, Coh94, DOS93, Gol94, Sar94, Sim93b, Sim93c, Tan93b, Tsi99, YQM99, vdHS94]. Kuzmin [Mac93]. Kwang [Ano99-86]. Kwang-Ting [Ano99-86]. Ky [WX97, WY99, YY98]. Kyle [Ano99-147]. Kyng [Ano98-34]. Kyurkchiev [Ano99-123, Ano99-124].

L [Ano90s, Ano90u, Ano92b, Ano95r, Ano95s, Ano95-27, Ano95-76, Ano95-79, Ano95-41, Ano96j, Ano96-27, Ano96z, Ano96-34, Ano96-54, Ano96-75, Ano96-81, Ano96-80, Ano96-83, Ano96-91, Ano96-106, Ano96-167, Ano96-145, Ano97s, Ano97-37, Ano97-57, Ano97-65, Ano97-161, Ano98-42, Ano98-61, Ano98-84, Ano98-80, Ano98-92, Ano99s, Ano99-53, Ano99-54, Ano99-65, Ano99-92, Ano99-87, Ano99-113, Ano99-231, GS95]. L. [GS95]. lab [Ano97-147]. Labahn [Ano96-107]. laboratories [BGEZ90]. Laboratory [Ano99-146, KZ95, Ano96-147, Ano97-67, Ano97m, Ano97-79]. labs [Ano96-102]. Laceable [LW97]. Lachlan [Ano98i]. Ladefoged [Ano96-60]. Lafayette [Ano97-132]. lag [Hal90, Lag91, Sim93b, Sim93c, WT90b]. Lagging [Van92]. Lagrange [AB94, CMP95, Con98, Feu90, HW93a, Wu90b]. Lagrange-type [CMP95, Con98]. Lagrangian [GN99, Gla91d, Gla93b]. Lagrangians [Ano96-117]. lags [JCL93]. Laguerre [DCL^{+}93, KS98a, LOS99, NR96a]. Lahanier [Ano96v]. Lakeyev [Ano98-32]. Lakshmikantham [Ano99-96, Ano99-97]. Lalonde [Ano97-56]. Lam [Ano98-149]. Lamb [Ano96-91]. lambda [Den91b, Ano99-68]. lambda-calculi [Ano99-68]. Lambert [HM96]. Laminar [CC94c, WHC90]. Lamoreaux [Ano95x]. Lamport [Ano95-42]. Lanczos [Ano99-245, Cab97, FLW97, OR97b, OR97a, Sha90c, SBD99, SBS99, Yan99a, BRZS97]. Landau [AM95, Du94, GHS91]. Landauer [Ano96-162]. Lang [Ueb92]. Lang- [Ueb92]. Langdon [Ano99p, Ano99-99, Ano99-100]. Langtangen [Ano99-43]. Language [Ano93o, Ano95p, Ano93u, Ano95-81, Ano96-158, Ano96-140, Ano97v, Ano97-144, Ano97-75, Ano97-110, Ano98-31, Ano98-58, Ano99-178, BD96, DB96, FMS92, KHC92, Lyt92b, Zar92, Ano95-78, Ano97-64, Ano98l, Ano99-193, Ano99-194, Ano99-227, Ano97-81, Ano97-80, Ano98-90, Ano99-50]. Languages [Sti93b, Ano96-55, MG93, Sti94, Sti96, Sti97, Ano96-73, Ano99-82, Ano99-228, Ano96-97]. Lanham [Ano95-75]. Lankford [Ano97h]. Laplace [BD90, BRV98, Che93a, FK98a, FK98b, HLS91, HWL94, Mat99]. Laplacians [HM91, Joh99]. Lar [Ano95-63]. Large [WW91, Ano96-170, Ano98-109, Ano99-245, BW96a, BSW96, Cho94, FLW97, GRKM96, GL94, KH92, LT99, MB94b, Pet92, PT91, RG90c, SM98, Sie90, Tor90, YJ94, Yan94, Yan95, Ano95-41]. large-scale [GRKM96, LT99, Pet92, Sie90, Ano95-41]. largest [RSR98]. Larry [Ano98g, Ano99-75, Ano99-76]. Lars [Ano98-34]. Larsen [Ano99-154, Ano99-155]. Larson [Ano96-95, Ano96-149, Ano98-140]. Lasance [Ano98-155]. laser [RS93, Shi97]. lasers [KM90, MDVK97]. Laskey [Ano99-176]. last [Ano96-124, Ano97-139]. László [Ano97n]. Latin [Ano99-65]. lattice [Kuu94, Mur94, Pad92a]. lattices [Pro95, RC92, Wil92]. launch [Ano96-152, Yav91b]. Laura [Ano95n]. Laurence [Ano99-127]. Laurent [Ano97-110]. Laurie [Ano99y]. LaVal [Ano97-132]. Lavalléee [Ano96-43]. Law [Ano96-142, Ano97-134, Pas92, YFF92b, YFF92a, YFF93]. Lawrence [Ano97-36, Ano97-55, Ano97-145, Ano97-95, Ano97-162, Ano98-99, Ano99k]. laws [Cha96, CC92, Gla93c, Gla93e, Gla96c, Hua95, Prn97, Sha98, Sha99a, Wan91, Ano99-202]. Lawson [Ano96-101]. Lax [Pok97]. Layer [Chi93, McK96, DPD98, Dev93, GS96b, Li97, Li98a, LMW95, NNB97, RD96a, Sod92, YL99]. layered [JS97b]. layers [Bog98]. layout [Chu99]. Laywine [Ano99-65]. Lazaro [Leo96]. Lazarus [Ano98-57]. LBB [Dem94b]. lead [KO90, LY92]. leading [Mit92a, ST97]. Leah [Ano96-145]. leakage [Jab96]. Leake [Ano96-78, Ano97r]. Learning [Ano96-106, Paz90, Ano95r, Ano95s, Ano96-39, Ano96-38, Ano96-90, Ano96-99, Ano97-55, Ano98u, Ano98-152, Ano99-32, Ano99-106, Ano99-107, Ano99-232, AR99b, Bel94b, KU91, KPB97, Mar90, Ano93q, Ano96-78, Ano93j, Ano96-98, Ano96-99, Ano99-141]. Least [BF95, BA96, CL94a, Che92, DK97b, ER98, Fra96, ISM93, Li96a, Los89b, Sal92a, Ye99]. least-squares [ER98, Fra96, Li96a, Ye99]. Lebesgue [BM97]. Lecture [Ano97-107, Ano99-172]. lectures [Ano95-32, Ano95-43, Ano95-66, Ano95-68, Ano90u, Ano97-82, Ano98-91]. Lederman [Ano96-118, Ano99-158, Ano99-160, Ano99-159, Ano99-161]. Lee [Ano96j, Ano96-155, Ano97-69, Ano98-52, Ano98-96, Ano98-120, Ano98-160, Ano99p, Ano99-121, Ano99-230]. Leebaert [Ano95-77]. Leen [Ano96i]. Leendert [Ano95u]. left [SRCB98, TCS94]. Legacy [Leo96, Ano99-91]. Legaz [Ano99-95]. Legendre [DCL^{+}93, Fun97]. Leitsch [Ano97-152]. Lenarcic [Ano99s]. length [EDO93, PPC99]. Leo [Ano95-29]. Leon [Ano96-104]. Leonhard [Ano98-52, Ano98-120]. Leslie [Ano95-42]. Lesniak [Ano97-62]. less [Ano97l]. Lesser [Ano96-135]. Lessons [Ano97-49, Ano97r]. Leupers [Ano98-137]. Leuven [Ano98-155]. Levasseur [Ano99-77]. Level [RFRM99, Ano99-118, Ano99-119, Ano99-222, Ano99-223, CL99d, EL96, Fai98, Lay93, LLP94, MR99, NN91, RS93, Ano95y, Ano96-84]. Level-continuity [RFRM99]. levels [KKD^{+}97]. Leveson [Ano95-64]. Levia [Ano98l]. Levine [Ano98-152, Ano99-92]. Lewin [Ano99-78, Ano99-79]. Lewis [Ano96a, Ano98-82]. lexical [CF91b, Ano99-142]. Lexico [Ano96-100]. Lexico-logical [Ano96-100]. lexicographic [Sto92]. Leyton [Ano93s]. Li [Ano97j]. Liapunov [CKRV98]. libraries [Ano97-55]. library [Ano99-55, Ano99-56, GS96a, Ano95-52]. Lidl [Ano97-48]. Lidstone [HP99, WA94b, WA96c]. Lidstone-spline [WA94b, WA96c]. Lie [Ano97-58, Ano98-92, Ano97-82, Ano98-89, Ano99-71, Ano99-181, ARS98, AH97, LQ95a, LQ95b, Ano95-43, Ano98-92]. life [Ano95k, Ano95n, Ano99-29, Ano99-30, Ano99-83, Dre90, Pol90, Ano99-29, Ano99-30]. lifetime [AJ95, Aug90]. lift [RR90]. like [AA93, Ano97-148, BMD92, Ber93, CTY99, Cub95a, Dat90, Din99d, GMS98, JT99, PZDH95, PT96, Vic90]. likelihood [ED91b, Faz94b]. LIL [Kun90]. Lima [Ano96-149]. Limit [Ran93, Ruz90, Ahm91, Kor90b, MN97]. limited [Ram91c, dCVC97]. limited-angle [Ram91c]. limiting [EK94, Kov95]. limits [Ano97-146, Ano99-81]. Lincoln [Ano99-238]. Lind [Ano96r]. Linda [Ano96c, Ano96-99, Ano97-62, Ano98-150]. Lindberg [Ano96-163]. Lindenfeld [Ano97-151]. Lindfield [Ano96-125]. line [AI97, BWK97, CK94a, GF97, O’R98a, RV94, TTC96, Ano99-51]. Linear [Bal96b, DELE99, FHWW99, Gal99b, GF90, GF97, JT99, MS96, Mas92, MKS^{+}91, PLH95, Ram95, RG92, Sha98, Tre98, AR90a, AP94, Alt90c, Ame99, Amo93, APP96, Ano97-72, Ano97-67, Ano98-142, Ano98-154, Ano99-129, Ano99-130, Ano99-136, AK90b, AMZ98, Bai96b, BSW96, BJH99, BF91, BM97, BGM95, Bin91, Boh98, BG95, BRZS97, BT98, But96, CV90, CL94b, CCK97, Che92, CL98, CLY91, CL90, Cor95, CH99, De 96, DK97b, EG97, ESIP98, FW96, Far96a, FG97, FG98, Gal99a, GSM91, GGK96, GM96, GC99, HP96, HBB90, HP94a, Ita93, JLDL92, Ker91a, Kis91, Koç92, KP99b, Kou95, KPB97, LS96a, LS98b, LL91a, LB94, LN99, LYL95, Lin96, Los89a, LC96b, Mär96, MRWM94, Més96, MR94, MB90, MM95d, NNB97, Nov90, Nyí99, NR96b, O’R98c, OH93, Pan90b, Pan94, Pan98a, Pan98b]. linear [Pap92, Pap94b, PC96b, Ram90a, RG90a, Rei98, Rei99b, RT99a, Roj90, RZ91, RZ94, RG90c, SNU98, Sha90c, Sha95, Sha90d, ST98, SWF95, SWY96a, Sny90, SY96, Tör94, Tre95a, Vos91, WSS92, WT93, Wan93b, WK99, WA98, WF99b, XLJS98, Yas94, Yu98, Zad89, Zad92, Zha98, ZT99, pri94, Ano95-44, Ano96-101, Ano96-102]. linear-quadratic [LB94]. Linear-time [JT99]. linearised [Gla96c]. linearization [Ano99l]. Linearized [ZY98]. linearly [Par91a, VK99, dKSG94]. Lines [HIIN99, JM91, Lia95, SVS98, Sch94, YF92b]. Lingo [Ano99-143]. linguistic [Sti93a, Ano99-144]. link [Ano99-140]. linked [CT90]. links [MNE92]. linux [Ano95-63]. Lions [Ano96-83, Ano97-65, Ano99-113]. Liouville [AHY98, BMD92, CS93a, Con95, Doš98, GS96c, Hua91, She93, Tan93c]. Lipschitz [BL95, Gal93a]. Lipschitzian [Dra99, GS99]. Lipsman [Ano96-54, Ano98-42]. liquid [ARI96, Dor97, DIR97, Kor97, KMD99]. LISA [RD99]. LISP [Win91, SH96, Ano97-83]. Lisp-based [SH96]. List [Ano90-29, Ano90-30, Ano91l, Ano92h, Ano93-29, Ano95-86, Ano95-87, Ano96-172, Ano96-173, Ano97-174, Ano97-175, Ano98-168, Ano98-169, Ano99-247, Ano99-248]. Listening [Wit96b]. Lister [Ano90p]. lists [Ano98-97]. literacy [Ano95-79]. little [Ano96-154, Ano98f]. Littlewood [dJF91]. Litvinov [Ano98-92]. Liu [Ano95-46, Ano99-80, Ano99-221]. live [Pen96]. lively [Ano97-68]. living [Ano95k]. Ljapin [Ano98-153]. Llinás [Ano97-149]. Lloyd [Ano95-78, Ano96o, Ano96-103]. LMMs [ZZ99b]. Load [JL94, AGM99, BGM95]. loading [MR91, MB94b]. loads [BLR96]. Lobatto [Ano99-244, BB99, BACS98]. Local [AH97, Bin91, BL99, DH97, GZ95, DC94, Dei93, KN94, LNM90, LTR98, Sak90, WR93]. Local-Sparing [DH97]. Localization [RSAR95]. localized [BGL97, Ha90]. locally [Ano95-43, Din96, EDO96, WF90, WF91]. location [Ano99-63, Ano99-64, BC98, ED91b, Ioa91, LYL^{+}91, LCL94, MM96e, Xue94, Ano97-84]. locking [RS97]. LOD [VK95]. Loeckx [Ano97-136]. log [LL91b]. logarithmic [Ram93]. logarithms [DDP99]. Logic [Ano98-133, Ano95-34, Ano95-45, Ano96-90, Ano96-104, Ano96-114, Ano96-122, Ano97n, Ano97-39, Ano97-61, Ano97-64, Ano97-88, Ano98-96, Ano98-90, Ano98-95, Ano99-94, AS97, Bou93, Bur92, FGN90, GHPLZ90, LL90c, Paz90, RG90b, Yan91, Zad99, Ano95p, Ano96-72, Ano96-126, Ano97n, Ano97-54, Ano97-87, Ano95-45, Ano96-104, Ano96-103, Ano97-85, Ano97-88, Ano97-86, Ano97-87, Ano98-96, Ano98-94, Ano98-93, Ano98-95]. Logic-based [Ano97-85]. logical [Ano97d, Ano99-175, Ano96-100]. logically [HS97]. Logics [Ano99-145, Ano96-114]. logiform [BGLT90]. logistic [CC98c, GMS98, ZZ99e]. lognormality [Nai92]. logo [Ano97x, Ano97y, Ano97z, Ano93j]. logstable [Pan93b]. London [Ano93f, Ano93t, Ano93u, Ano93h, Ano93i, Ano93j, Ano93l, Ano93m, Ano93q, Ano93v, Ano93s, Ano93w, Ano95m, Ano95l, Ano95-44, Ano95-53, Ano96-71, Ano96-98, Ano97-33, Ano97-42, Ano97-139, Ano97-161, Ano99-32, Ano99-37, Ano99-109, Ano99-110]. London/Washington [Ano97-139, Ano97-161]. Long [Ano99-92, Blo91, RSP94, BT94, KW98, MBMS92]. look [EG98, LR93a]. look-ahead [EG98]. loop [Yav91a]. loops [Ano98-111]. López [Ano99o]. Lorentz [KLMP97]. Lorenz [GGC97]. loss [CKL98, HY98, Pou90a]. lost [KS97a]. lot [ZG97]. lot-sizing [ZG97]. Lotka [Lu96, LW99b, PC94b, TC99]. Lotus [Ano95-39]. Loudon [Ano99-147]. Louis [Ano96-58, Ano98-84, Ano99-68]. Loukides [Ano97-124]. Lovász [Ano96-81, Ano96-80]. low [Ano98-84]. low-speed [Ano98-84]. lower [Cub95b, FO91, KO91, LG99, OCS96]. lower-order [OCS96]. Lozano [Ano93v]. LP [Men91, Tor90]. LP-form [Men91]. LSMS [WSNS98]. Ltd [Ano95-51, Ano99-45, Ano99-74]. Lubbe [Ano98-73]. lubricated [Ano98-83]. Luby [Ano96-136]. Luc [Ano99-39, Ano99-186]. Luca [Ano99-82]. Lucas [WCL99]. Lucio [Ano96-99]. Luczak [Ano93p]. Ludlow [Ano96-85]. Luger [Ano96-36]. Luis [Ano95-27, Ano99-128]. Luke [Ano99-234]. Lumley [Ano98-84]. Lumpkin [Ano96-77]. Lumsdaine [Ano99-160, Ano99-161]. Lunel [Ano95u]. lunes [He95a]. LUP [JT99]. Lusk [Ano95-82, Ano99-160, Ano99-161]. Lutz [Ano97-123]. Lyapunov [BG95, KB90, LW91, Nov90, Yan93d]. Lynn [Ano97-81]. Lynne [Ano95-72].

M [Ano93o, Ano95u, Ano95x, Ano95-73, Ano95-33, Ano96h, Ano96-32, Ano96-47, Ano96-79, Ano96-81, Ano96-80, Ano96-98, Ano96-106, Ano96-107, Ano97-37, Ano97-45, Ano97-63, Ano97-91, Ano97-107, Ano97-139, Ano97-165, Ano98v, Ano98-49, Ano98-82, Ano98-108, Ano98-123, Ano98-134, Ano98-141, Ano98-155, Ano98-160, Ano99-68, Ano99-74, Ano99-77, Ano99-83, Ano99-115, Ano99-116, Ano99-128, Ano99-163, Ano99-150, CKL98, Kim92, Toy92, CKL98, Kim92, Toy92]. M.P.P. [SCBGN94]. M/G/s/s [Kim92]. M/M/1 [Toy92]. M/M/c [CKL98]. MA [Ano95k, Ano95-36, Ano95h, Ano95-70, Ano95o, Ano95r, Ano95s, Ano95d, Ano95q, Ano95-72, Ano95v, Ano95-71, Ano95-28, Ano95-29, Ano95e, Ano95-35, Ano95-77, Ano95-37, Ano95-38, Ano95-78, Ano95f, Ano95-40, Ano95j, Ano95-79, Ano95-45, Ano95-41, Ano95-42, Ano95-51, Ano95-47, Ano95-48, Ano95-49, Ano95-55, Ano95-57, Ano95-59, Ano95-60, Ano95-61, Ano95-62, Ano95-64, Ano95-81, Ano95-82, Ano96s, Ano96i, Ano96h, Ano96k, Ano96l, Ano96-32, Ano96-39, Ano96-44, Ano96-36, Ano96-151, Ano96-43, Ano96-34, Ano96-31, Ano96-153, Ano96-30, Ano96-40, Ano96-47, Ano96-46, Ano96-55, Ano96-53, Ano96-52, Ano96c, Ano96-69, Ano96-67, Ano96-61, Ano96-57, Ano96-66, Ano96-62, Ano96-58, Ano96-74, Ano96-73, Ano96-78, Ano96-86, Ano96-87, Ano96-81, Ano96-80, Ano96-85, Ano96p, Ano96-90, Ano96n, Ano96o, Ano96-95, Ano96-94, Ano96-96]. MA [Ano96-104, Ano96-100, Ano96-99, Ano96-103, Ano96-154, Ano96-97, Ano96-114, Ano96-105, Ano96-118, Ano96-115, Ano96-122, Ano96-121, Ano96-158, Ano96-134, Ano96-128, Ano96-132, Ano96-138, Ano96-140, Ano96-139, Ano96-148, Ano96-142, Ano96-149, Ano96-162, Ano96-84, Ano96-135, Ano96-145, Ano96-160, Ano97g, Ano97f, Ano97-53, Ano97l, Ano97r, Ano97-31, Ano97s, Ano97-32, Ano97u, Ano97x, Ano97y, Ano97z, Ano97v, Ano97-43, Ano97-37, Ano97-44, Ano97-49, Ano97-57, Ano97-63, Ano97-55, Ano97-144, Ano97-64, Ano97i, Ano97k, Ano97-69, Ano97-81, Ano97-88, Ano97-80, Ano97d, Ano97-149, Ano97-91, Ano97-150, Ano97-109, Ano97-110, Ano97-107, Ano97-105, Ano97-104, Ano97-115, Ano97-126, Ano97-137, Ano97-156, Ano97-153, Ano97-134, Ano97-133, Ano97-160, Ano97-163, Ano97-52, Ano98h, Ano98l, Ano98u, Ano98x, Ano98-34, Ano98-28, Ano98-38, Ano98z, Ano98-37, Ano98-31]. MA [Ano98-45, Ano98-48, Ano98-50, Ano98-47, Ano98-61, Ano98-58, Ano98-56, Ano98-63, Ano98-68, Ano98-69, Ano98-82, Ano98-76, Ano98-96, Ano98-94, Ano98f, Ano98-95, Ano98-106, Ano98-110, Ano98-104, Ano98-113, Ano98-114, Ano98-132, Ano98-151, Ano98-144, Ano98-138, Ano98-139, Ano98-152, Ano98-156, Ano98-161, Ano98-164, Ano99-29, Ano99-30, Ano99-90, Ano99o, Ano99q, Ano99p, Ano99-50, Ano99-213, Ano99-40, Ano99-38, Ano99-52, Ano99-57, Ano99-58, Ano99-75, Ano99-76, Ano99-86, Ano99-92, Ano99-81, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-87, Ano99-99, Ano99-106, Ano99-100, Ano99-107, Ano99f, Ano99-118, Ano99g, Ano99-119, Ano99-117, Ano99-120, Ano99-216, Ano99-217, Ano99-128, Ano99-121, Ano99-131, Ano99-122, Ano99-145, Ano99-142, Ano99-158, Ano99-160, Ano99-159, Ano99-161, Ano99-163, Ano99-167, Ano99-168, Ano99-165, Ano99-169, Ano99-220, Ano99k, Ano99-201, Ano99-209]. MA [Ano99-226, Ano99-229, Ano99-212, Ano99-211, Ano99-231, Ano99-232, Ano99-227]. MA/New [Ano97-137]. MA/The [Ano96-31]. MA/Washington [Ano97-37]. Maarten [Ano97n]. Maccamy [FLM94]. MacCormack [Ber92]. Macdonald [Ano99-148]. MacEachern [Ano99-238]. MacGregor [Ano90v]. Mach [IGM98]. Machine [Bin91, Ano95-59, Ano96-90, Ano97-133, Ano98u, Ano99-106, Ano99-107, Cha99a, Che90b, CD98, HC98, KCD^{+}97, May99, Sie90, Sni90, Sni91a, Ano97-78]. machines [Ano95-76, Ano96-94, Ano97-46, Ano97-94, HC98, VDLM95, YF93]. Maciej [Ano98-44]. Macintosh [Ano95-80]. MacKenzie [Ano96-94]. macro [GG92]. macro-economic [GG92]. macroeconometric [Dei91]. macroeconomics [Ano97-42]. macroscopic [Ano99-233, AKK^{+}97]. made [Ano96-137]. Maes [Ano95k, Ano97-53]. Maeve [Ano97-80]. Maffioli [Ano98m]. Magdy [Ano97-163]. magnetic [BKK97, LD95, TCJG91, WSNS98]. Magneto [NSD98, Alk94, Ber98]. magneto-acoustic [Alk94]. magneto-thermo-visco-elastic [Ber98]. Magneto-thermoelastic [NSD98]. Magnetohydrodynamic [Chi93, YC92]. magnetohydrodynamics [Mei93]. magnetothermoelastic [CR90b]. Maher [Ano97-88]. Mahmud [Ano97-82]. mailing [Ano97-141, Ano98-97]. Main [IZ97, SV97]. maintenance [KK92]. Maja [Ano97-53]. majority [GF90]. Makedon [Ano98-47]. Mäkelä [Ano99-74]. Making [Mil92, Ano95-27, Ano97-161, Dei91, Eke99, Erm92, HGWL96, KSK90, Li99c, Ano98-27, Ano96-105, Ano99-146]. Malcolm [Ano96k]. Málek [Ano97-165]. Maletsky [Ano99-89]. Mallios [Ano99-105, Ano99-103, Ano99-102, Ano99-104]. Maluszyński [Ano98-94]. Malyshev [Ano99-198]. Mammone [Ano95m]. Mamoru [Ano97-82]. Man [Ano93i]. Man-Kam [Ano93i]. management [Ano97-45, Ano97-105, Ano98-155, Ano98-156, Ano99-131, FLNW98, VMA99]. Managing [Ano95-46, Ano98-99, Ano98-97, Ano98-98]. Manchek [Ano95-59]. Manchester [Ano99-172]. Mandelbrot [Ano98-62, Ano99-162]. Maneuver [RA92a]. maneuverability [YZM94, YZM95]. Manfred [Ano95-53, Ano97-84, Ano98-102, Ano99s, Ano99-142]. manifold [Rhe97]. manifolds [BR94, Bro98, Köh93a, Köh95, Köh96a, Köh96b, Köh96c, Kou95, McL95, Rhe96, ZQ94, ZQ95, Ano97-59, Ano98j]. manipulating [CCY93]. manipulation [BD96, DB96, TM91]. manipulator [BS99b, Szk95]. Mankiw [Ano97-100]. Mann [Cha99b, Hua98a, Hua99]. MANPAK [Rhe96, Rhe97]. MANPAK-manifold-algorithms [Rhe97]. mantras [Ano96-55]. manual [Ano96-120, Ano97-79, Ano97-114, Ano99-69, Ano99-201, Ano95-42]. Manuel [Ano97-138, Ano98r]. manufacturing [Ano99-128, Cha94, FYS99, Lew94, Tsi91, YF92a, YF92b, YF93, ZK99, Ano95l]. many [PA99, WF99b]. Manzano [Ano96-72]. map [AO99b, Ano96-155, Ano98-60, Ano98-59, AH97, Dri97, Fol92]. map-building [Ano96-155]. Maple [Ano96-156, Ano96-159, Ano97-73, Ano99-129, Ano99-130, Bis91, Yan95, YJ99, Ano96-68, Ano97-71, Ano96-106, Ano96-107]. Maple(R) [Ano97-147, Ano98-125, Ano98-44]. Maple^{T M} [Ano96-159, Ano96-54]. mapping [FT97b, HGK93, IN97, MH93a, TE93, Wu99b]. mappings [BW92, Cha99b, CLY97, DK93, Din99a, Din99b, Dra99, Hua98c, Hua98b, Hua99, HB99, Sya99, Ano99-49]. maps [AO99a, AO99b, Ano96-56, Ano96-117, ABB97, Dia91, Din96, Gla90c, GB96, Mag97, Mro96, O’R98b, Pav97, RZ96, RA92b, Sib98, Ano96t, Ano97-30, Ano96-108]. Marc [Ano96-118, Ano98-33, Ano99-158, Ano99-160, Ano99-159, Ano99-161]. Marcello [Ano95-74]. Marcelo [Ano96t]. March [Ano99-136, Ara99b]. Marchand [Ano99-108]. marching [GMR90, MG90c, MH93b]. Marcia [Ano96-65]. Marco [Ano99-58]. Marcus [Ano95-72, Ano95-69, Ano96r, Ano98-100]. Marek [Ano98-55, Ano98-164]. Margaret [Ano95-28, Ano98-112]. Margolis [Ano98-39]. Maria [Ano96-48, Ano96-72]. Mariette [Ano98k]. Marik [Ano99-128]. Mario [Ano96v]. Marion [Ano99-39]. Marjorie [Ano96-50]. Mark [Ano96-151, Ano97-75, Ano97-123, Ano98-33, Ano98-77, Ano99-37, Ano99-109, Ano99-110, Ano99-173]. marker [Hen92]. marker-passing [Hen92]. market [VA94]. Markov [AK93, Cse95, Dri97, Kor90b, Kur91, KK94, MM96b, Ohn92, Tan96, TPF98]. Markovian [Alf95, BTT90, KK92, Kij92, Nab98, Ruz90, Szt92]. Markus [Ano97-136]. Marla [Ano95-67]. Marsaglia [PW95]. Marseille [Ano98-84]. Marshall [Ano95-73, Ano98-30]. Martella [Ano96-48]. Martello [Ano98m]. Martin [Ano90x, Ano95d, Ano96-27, Ano97-90, Ano98-48, Ano98-61, Ano98-129, Ano99-98]. Martinez [Ano99-95]. Martinez-Legaz [Ano99-95]. Martini [Ano99-101]. Marubayashi [Ano98-115]. Marx [Ano95-29, Ano97n]. Mary [Ano96-151, Ano97-81, Ano99-37, Ano99h, Ano99i]. Maschler [Ano95-62]. Maslov [Ano97-66]. mass [Ber96, YFF93, Ye99].

Massachusetts [Ano93f, Ano93t, Ano93u, Ano93i, Ano93j, Ano93l, Ano93v, Ano93s, Ano93w]. Massively [Hen92, LVV96, AL94, Ano97-160, CK97a, Ano95-47]. Massively-parallel [Hen92]. master [Ano99-73]. Mastering [Ano96-109, Ano96-110, Ano97-89, Ano99-148, Ano99-147]. Masuch [Ano97n]. Mataric [Ano97-53]. matching [Ano97-118, Ano98-55, Hsu91, TYC93]. matchings [BV95, SY96]. material [Nel96, Pol90, Ram91a, vdVVdB99b, Ano98-101]. materials [Ano98-83, Shk97, Ano90u]. Math [Ado91a, BM90c, Kou90b, Li10, Moo93b, MG90b, Sal92b, Ano99-215, Ano95-67, Ano96-111]. MathCAD(R) [Ano97-67]. Mathematica [Ano98-80, Ano97o, Ano98p, Ano98-107, Ano97-90]. Mathematical [Ano95-67, Ano97-68, Ano97-79, Ano99-73, Ano99-242, CC92, MATA98, MPR97, MR91, NAKA98, Niw96, Wit90, Ano95c, Ano95-33, Ano96-29, Ano96-65, Ano96-146, Ano97o, Ano97-107, Ano97-134, Ano98p, Ano98-134, Ano99e, Ano99j, Ano99-165, Bon99, Fau94, Gal99c, GPCSL90, Gyö90b, Hen97, MP99, MPS92, PCR90, She94, Ush94b, Ano96-113, Ano96-112, Ano97-91, Ano98-102, Ano99-149]. Mathematica(R) [Ano97m, Ano96q, Ano97w, Ano97-96, Ano99-77, Ano95-48, Ano97-92]. Mathematica^{T M} [Ano98-42]. Mathematics [Ano90a, Ano95-44, Ano99-89, Zad92, Ano93j, Ano93n, Ano95-32, Ano95-43, Ano95-66, Ano95-68, Ano96-68, Ano97-48, Ano97-51, Ano97-70, Ano97-87, Ano97-162, Ano98n, Ano98-49, Ano99-65, Ano99-78, Ano99-79, Ano99-111, Ano99-112, BGEZ90, But94, Gal99c, MM94, Sch90, Var92, Ano96j, Ano97-122, Ano98-90, Ano99-135, Ano93k, Ano97-93, Ano99-150]. Mathematics/Springer [Ano99-89]. Mathematics/Springer-Verlag [Ano99-89]. Mathew [Ano98-142]. Mathews [Ano93n, Log96]. Mathias [Ano98-105, Ano98-140]. Mathiassen [Ano98-34]. Matlab [Ano96-125]. MATLAB(R) [Ano96-102, Ano99-218, Ano99-219, Ano99-108]. MATlab^{T M} [Ano95-49]. Matos [Ano99-128]. Matousek [Ano99-135]. matrices [Ano99-244, BB99, BLR94, BLR90, BG91a, CGS99, CF91b, Den91b, DFS93, DKW98, EG97, EG98, Heg91a, Heg91b, JT99, Kal90, LR93a, LLR^{+}90, Luo93c, LURL98, Nan90, OB97, Pan92a, PTC^{+}98, RS92, RSR93, RS96b, RM98b, SS94, Szu96, Ter00, Tre95b, Var98, Xu97, ZM94]. Matricial [Car93]. Matrix [BW93, Has93b, Has95, KTO97, AAD98, Ano96-129, Ano00, BSW96, BSS98, BG91b, Cab97, CE94, CDLW95, Coo98, ER98, FL91, GGS97, Han92, Heg91a, Heg91b, JP95a, JG98, Käm94a, KO97, KO99, Li99b, LC97b, Luo93a, Luo93d, MV95a, MV95b, MM96a, MP96, MMMS97, MMMS96, MRJ97, NPJ95, Pan90a, Pan93a, PYS97, PC96b, TCS93, TCS94, Usm94, Wan93b, Zed90, dSB96, Ano95-49]. matroid [Coo97]. Matt [Ano95-63]. Matthew [Ano98i]. Matthias [Ano96-154, Ano96-160, Ano98f, Ano99-195]. Matti [Ano97-30]. maturity [Ano96-151]. Matveev [Ano98j]. Mátyás [Rod94f]. Maugeri [Ano95-83]. Maurer [Ano99-233]. Maurice [Ano99-189]. Mauro [Ano98m]. Mavina [Ano97-30]. Max [Hel96, FW96, GHWL96, WW98, Bes96]. max-min [GHWL96]. max-product [WW98]. Maxima [Ano99-151]. Maximal [NB98, Dev93]. Maximum [ED91b, Faz94b, AL93, Li98c, NB95, SY96, TYC93, WCHL96]. May [Ano97-79, Ano99p, Ano99d, Ano99-152, Ano97-141]. May-cock [Ano97-79]. Maybury [Ano98-77]. Mayo [Ano96-64]. Mayr [Ano98-91]. McAloon [Ano96-126]. McCarthy [Ano97-140]. McClamrock [Ano95-31]. McClellan [KV95]. McDonnell [Ano96-67]. McGraw [Ano96-112]. McGraw-Hill [Ano96-112, Ano97-137]. McKeehan [Ano99-180]. McKeown [Ano96-116]. McKnight [Ano97-69]. McLaughlin [Ano98-112]. McOwen [Ano97-116]. MCSE [Ano98-103]. Mean [Ara97, ED91a, Toy92]. meaning [Ano95-30, Ano96-95]. meanings [Ano96-58, Ano97-142]. means [BB93, DW97, Ram95, Yan93b]. measurability [Zhe93]. measurable [AI97]. Measure [KM96, Ano97-165, Ano98-74, Cse95, GL94, HP99, Nab98]. measure-valued [Ano97-165]. measurements [Noa91, Ram91b]. Measures [Mez97, TXC98, Ahm91, AR99a, Ano99-196, GB90, JGB95, Kai96, KK94, LF98, MC93, RZ96, SB97, Sza90a, Ano98-75]. Measuring [Bri97]. mechanical [Gre91a, Gre92b, Gre93b, KU91, Kor92, VV99]. mechanics [Ano95-54, Ano97-92, Ano98-116, BMM94, BGM97, HPH^{+}94, RV94, Ano98-157, Ano99-152]. mechanism [Bes96, Byk97, FKO92, Lia99b, MT98, RW98]. Media [Ano97-90, Ano97-148, BCGP99, Bla95, Eti91, KG94, KP99a, MDVK97, NSD98]. Media/Cambridge [Ano97-90]. medical [Eso99b, SS90]. medicine [Ano97-146, Wit90]. Meditations [Ano96-50]. medium [BR95a, Ber98, CR90a, JG96, KKB97, KMD99, Pas92]. medulla [WSDT94]. medullary [MBMS92, MT98, SMTS97]. Mehler [Ano96-31]. Mehrotra [Ano93w, Ano97-44]. Meinel [Ano99v, Ano99w]. meiotic [BCB90]. Meixner [BK98]. Melanie [Ano96n, Ano99-166]. Mellin [HS95]. melt [Dor98a, Dor98b, Dor99]. membership [LL91a]. membrane [CMMR92]. memoirs [Ano95-69]. Memorial [Ano99-172]. memoriam [Eso99c]. memories [KKK91, KNU99, MKS^{+}91]. memory [APP96, Ano93h, Ara99b, BP94b, CDLW95, LH99, Man96, Mav93, Tau90b, WR93, hCPW98]. Menlo [Ano95-47, Ano96h, Ano96-44, Ano96-36, Ano96-53, Ano96-135, Ano97r, Ano98-77, Ano98-112, Ano99-40]. mental [Ano96-97]. Meromorphic [Ano99-153]. meromorphically [JS99, SHA99b]. Merrill [Ano97-81]. Merritt [Ano95-29]. Mersenne [Tan92]. Mesh [Bak92, CS96, FH99, Gla91c, KG98, MN91, She94]. Meshes [Ano97-115, Dia91, Gla91b, Gla91e]. message [Ano95-82]. message-passing [Ano95-82]. Messina [Ano96-62]. Mesterton [Ano95c]. Mesterton-Gibbons [Ano95c]. Meta [Ano96-114]. Meta-logics [Ano96-114]. Metacognition [Ano96-115]. metallic [RSV98]. Metamathematics [Ano97-94]. Metamorphosis [Ano96-116]. metaphor [Ano96-52]. metaphysics [Ano99-39]. Metcalfe [Ano96-115]. Method [Ano99-39, EDO93, EDOS93, KO91, LG99, McK96, Sch94, AC94, AO91, AR90a, Ado91a, Ado91b, AR92a, APV97, AO98c, Alt90b, Alt90c, AB94, And93, Ano96-52, Ano00, Ara96, Ara99a, Arg98, AKV91, BBO99, BW96b, Bai96e, Bai96d, Bai96f, BSS98, Bak90, BBI97, Bed95a, Bed95b, BRZS97, BH90, BPLS96, BG91a, BG91b, Cas91, CC94a, CS93a, Che96, CLD97, Chi91, CCY93, CS94b, CDLW95, CP98b, Con95, CKRV98, CH99, Dav96, DK97a, DOS93, EDO98, ESIP98, ESKG98, Far91a, Far91b, Fin99, Flo95, Fra96, FO91, Fun97, Gab94a, Gab94b, GZ91, Gal91b, GT97a, Gil91, Gla94b, GG90, GGC97, GHWL96, GHM96, GPK99, GH98, Han92, Har90, Har92a, Har92b, HM96, HJ91, HAO91, HW96, Huj96, HP94c, HH93, HWL94, IESAH95, JR94, Jek94, JTY97, Joh99, JKK98, KK97a, Kar96a]. method [KO96, KT92, KP96, Kou98, KPS96, Kra90, KB90, KCB90, Lai97, LD90, Lay93, LYL^{+}91, LJ98, LLR^{+}90, Li96b, Li98b, Li99c, LN99, LS99, LGP98, LP99, LT99, LM96b, Luo93b, MS96, MS93, MH99, MMP98, MG95, MG97, MM96b, Mar95, MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b, MM93b, Mey99, MK93, Mil90, MH93a, Mol96, MG90a, MG90b, Mur92b, Nan90, Nis97b, Noo99, Nov90, Nyí99, Oli98, OPD97, Oza94, PSS99, Pan90b, Pan98a, Pan98b, PBG90, PS96, PPC99, PC91, PT96, PR96, Pom92, Ram90a, Ram91a, RF91, RT99a, Roj90, RT99b, RSV98, RG90c, Sán98, SVS98, Sch92b, Scr91, Sha96a, Sha95, ST94, Sim93b, SM95, Sim97, Sny90, SFY92, Sod90, Sol97, SCB^{+}93, SS91b, SI94, Sug94, SCC97, SBD99, TPF98, TZ93, VR95a, VK99, WT93, WSNS98, WCL99, WK99]. method [WF90, WF91, Wu98, WF99b, Yan93a, Ye99, YK95, YLC96, ZS93, Zha93c, Zha93d, ZB95, Zha95, Zha96, ZS97, dKSG94, vNvN90a, vNvN90b, NO93, Ano99-117]. Methodology [DH97, Ano99-201, Gre92b, Gre93b, Gre95, Jan96, KVCA^{+}99]. Methods [Ano97-116, Ano99-52, AC95, AE90, AO99a, AC93a, APP96, Ano93n, Ano96-61, Ano96-83, Ano96-125, Ano97-41, Ano97-91, Ano98o, Ano98j, Ano98-66, Ano98-102, Ano98-101, Ano98-127, Ano99-43, Ano99-101, Ano99-113, Ano99-172, Ano99-175, Ano99-198, Ano99-205, Arg99, AS96a, AS96b, BW96a, Bai96c, BSW96, Bar91b, BF95, BLP95, BR95c, BGL90, BM90b, Bon97, Bro96, BT98, But94, But96, CGH92, CF92, CQ98, CJS90, CAZAS95, CKAS96, Che92, Che93b, CLM99, CS93b, Cho94, Coh94, Con98, CSW99, CT96, CRV98, CC96, CH99, DBM92, DBM95, DG99, Dem94a, Dem94b, DEMS94, DLS99, Du94, EDO97, EY91, FZ93, Fli94b, FM96, FK98a, Gal99a, GRZ96, GZ93, GS93, Gol90, Gol94, GPGCM94, GP91, HM99, HNS^{+}90, HS94, HS98, JMPS94, Kar94, KP98, KS98a, KT95, KG95b, Kro99, KN94]. methods [LCP98, Lay99, Lee94, Li96c, Li97, Li98a, LN98, Li10, LTB96, Lop93, LF91, Mak90, Mar94, MM93a, Mas92, MT92, MRWM94, Més96, MBV91, MHBD96, MRJ97, NAS97, Ohn92, PA94, Pao98, Pat98, PW97, PM91, PC93, PIT97, PPH98, PS90, Pom91, QZ91, SS98a, Say99, Sem93, Ser93, SA94, SM94, Sir99, Som90, Söv99, SMU98, SB99, Tan93d, TVZPG96, Ter00, VBM90, VK95, Wan98b, Wan99, WT90a, YQM99, Yan99a, YY98, ZM94, Zha93b, ZCC96, dOF91, hC96a, hC96b, hCPW98, vdHS94, Ano96-164, Ano97-159, Ano99-123, Ano99-124, Ano98-104]. metric [Bad95, Sya97]. metrics [EA92]. Metzner [Hak99]. Meulen [Ano96-140]. Meyer [Ano95-36, Ano96-34, Ano97-53, Ano97-78, Ano99-90]. Michael [Ano90r, Ano93s, Ano95c, Ano95-62, Ano96-100, Ano96-136, Ano96-130, Ano97g, Ano97n, Ano97v, Ano97-51, Ano97-88, Ano97-103, Ano98h, Ano98-28, Ano98-103, Ano98-113, Ano99x, Ano99q, Ano99-45, Ano99-166, Ano99-170, Ano99-201, Ano99-226]. Michal [Ano99-171]. Michel [Ano97-110, Ano98l, Ano98-134, Ano99-95]. Michnick [Ano96-158]. microcirculatory [XS90]. Microcomputer [BGEZ90, AR90c, Lei90]. microcomputers [Wag90]. microdrop [Gre91b]. microdrops [Gre90b, Wya94]. microeconomics [Ano99-203]. micropolar [GK91, NSD98]. Microscopic [RP97]. microturbulence [KG95a]. microworlds [Ano97-160]. Micula [Ano99-114]. MIDI [SH96]. midst [Ano95-76]. Miettinen [Ano99-74]. Migdalas [Ano98-108, Ano98-123]. Mignotte [Ano99-189]. Migrating [Ano95-50]. Mike [Ano97-124]. Milanese [Ano96v]. miles [Ano97-139]. Miller [Ano97-145, Ano97-115]. Milner [Ano97-28]. Milovanović [Ano98-135]. Milton [Ano97-161]. MIMD [CDLW95]. mimetic [HS98]. Mimura [Ano97-82]. min [FW96, GHWL96]. min-max [FW96]. mind [Ano96-97, Ano97-143, Ano97-149, Ano93s, Ano97-61]. mind-brain [Ano97-149].

Mindlin [Tad93]. minds [Ano95-31, Ano96s]. Minendra [Ano97-84]. Ming [Ano97j]. minima [Ano99-151, Moo93a]. MINImal [Ano95-49, MM96a, Mit90, Ano99-206, Ano99-207, AS98, CP95a, SOI92, Sim93c]. minimal-repair [SOI92]. minimalist [Ano96-100, Ano96-161]. Minimax [BNV90, JTY97, LKT99, Ver99b, Ver99a, WX97, YY98]. Minimization [Sza91, AC93a, BH91, Fra96, GH90, Gre90c, LYH98, LT99, Sni91b, Zhe93, ZZ99d, dKSG94, dOF91]. minimized [NG91]. Minimizing [Med90, BM97, GS99]. Minimum [FM96, DKW98, NB95, WCHL96]. mining [Ano99-52, Ano99-80, Ano96h, Ano98-139]. minisum [BC98]. MINQUE [Sch92b]. MINQUE-method [Sch92b]. mirror [Oli93]. Miscellaneous [Mad92]. Mishchenko [Ano99-234]. miss [MPS93]. miss-distance [MPS93]. missile [MPS93]. missiles [II93]. missing [KT97]. misspecification [Lag91]. Mistakidis [Ano98-116]. Mitchel [Ano97-160]. Mitchell [Ano93g, Ano95-81, Ano96-30, Ano96-73, Ano96n, Ano97-139, Ano99-166]. Mitrani [Ano98-128]. Mitropolskii [Ano98o]. Mixed [JP95b, KP95a, Ano98l, Ano98-101, Ano98-142, BMB94, CLD97, DBM92, DBM94, DBM95, DG99, Din99a, DL99, HH98b, JA96, JD97, JA97, KP99a, Koç91, LTC98, Mei93, NPJ95, NFJ96, RF91, She92, She93, VBM90]. mixed-radix [Koç91]. mixed-signal [Ano98l]. mixed-type [DBM92, DBM94, DBM95]. mixing [Yos90]. mixture [Ahm94, Ahm95]. Miyamoto [Ano98-115]. ML [Ano97-99, Ano97-95]. Möbius [BWK97]. modal [Ano97n, RSV98]. modality [Ano97-117]. Mode [KKG97, RS97, YF93]. Model [SI94, AJ95, AS93, Ano96-151, Ano97-141, Ano98i, Ano00, Bab98, BSS98, BKM91, BCGP99, BJH90, BM89, BM90c, CMMR92, Che96, Cho95, Den91a, Dor97, DIR97, Du94, Ege91, Fli94b, Fli94c, FYS99, FGN90, FF92, FG97, GGB92, Gre92a, Gre93a, GPCSL90, Iri98, Jon96, JS97b, KS99, KM92, KG94, KP95b, Kor91, KC91, KPB97, LRLR95, LL94a, LZ98, LO99, LKT99, MG93, MR92, Mil90, MW99, Nas92, NIS97a, Niw96, Pan92b, PBG90, PCR90, RR90, Raq98, SK91, She94, SFT99, Sod92, SCBGN94, SG90, Tau90a, Tau90b, Ueb92, dCVC97, VMA99, WF96, WW98, XLJS98, Yas98b, YFF92a, YFF93, Yav96, ZS93, pri94, Ano99-125, Ano99-126]. model-based [FGN90]. Modeling [Ano95-60, Ano97w, BJ90, Ano97-110, Ano98-104, Ano98-119, Ano98-159, Ano99-92, Ano99-94, Ano99-169, Ano99-242, AÉ96, CC92, Edw92, Kor97, MATA98, MP99, ML90c, NWL91, NL91, NAKA98, Tad93, Wit90, ZK99, Ano90v, Ano99e, Ano99-46, Ano95-51, Ano97-96, Ano98-105]. modelled [Ado91d]. Modelling [Alf95, Ano96-171, BCB90, GB92, KSK90, Szk95, Szt92, Ano90-28, Ano96-98, Gyö90b, Kel91, Kos90, Mah90, MPR97, PY94, Voi90, Ano95c, Ano96b, Ano98-128, Ano99-154, Ano99-155]. Models [KM97a, AR92e, Ano96-32, Ano96-113, Ano98g, Ano99-53, Ano99-54, Ano99-106, Ano99-107, Bal90, Bar91b, BM90a, BTT90, Bon99, Bra92, BH90, Cse95, DK97b, FK97, FBL99, Fli95, FG98, Gan92, GMS98, Kor90a, Lag91, LCSN98, LMW95, LTB96, MT98, MK98, Nab98, NJ92, Pad92b, PC94b, Pio94, RHG91, SGS93, Sem93, Sha92, Shi97, Sjo92, SJS92, SMTS97, TPF98, Ter00, Tra97, US91, Ush94b, Var98, Vos91, VA94, WSDT94, YOO92, Yam94, YCC96, Zad89, Zad92, ZM94, ZC99, Ano96-39, Ano98-142, Ano99-63, Ano99-64, Ano99-87]. modern [Ano90p, Ano97-87, Ano97-156, Ano97-154, Ano99f, Ano99h, Ano99g, Ano99i, Ano99-149, Ano99-203, Mar91, Ano97-98, Ano97-99, Ano97-97, Ano98-106, Ano98-107, Ano99-156, Ano99-157]. modes [TCS94, WW93]. Modification [RT99b, VR95a, DC94, De 99, ESIP98, MG90b, VR95b, Yan97]. Modified [MBV91, BMD92, Fun97, MJK92, MH93a, Noo99, Pan90b, RAM92, YLC96, Ano96-117]. Modular [Ano95-44, And91]. module [SH96]. Modules [Ano97-140, Rot96, Ano99-69, Ano99-204, Ano99-238]. moduli [Ana95a, Bad95, GSST97]. modulus [Tan92]. Mohamad [Ano96-74]. Mohan [Ano97-44]. MOL [CP96]. molding [SCG^{+}97]. Molecular [Ano96-121, Ano99-166, Fre98, HPH^{+}94, Ber91, CCC^{+}98, MPS94]. molecule [Gre93a, Kor92]. molecules [TSLW98]. mollification [HM93a, Hin95, MM95a, MM96c, MG90a, Mur92b, MMZ98, ZM99]. Mollified [MM93b, GMR90]. Moment [AR92e, AR98, MB94a, DIT97]. Moments [Gau97, YSM92, AR99a, BFR94, GNR95, NR96a, TPF98]. momentum [YC92]. Momme [Ano98-111]. Monagan [Ano96-107]. Moncur [Ano98-103]. monetary [Ano97-42, Ano97-100]. Monkey [MZ93, PW95]. monograph [Ano97-141]. monopoly [Ano97-161]. Monotone [CZ94, LK94, Mil95, PA94, Pao98, SA94, TMGS98, Wan98a, Wan98b, ZCC96, Ano96-117, Bai96c, Bai96e, Bro94, Dev93, Noo99, TR99, Wu99b]. monotonely [Won98a]. monotonic [SFT99]. monotonicity [Ano99-95]. Monson [Ano99-72]. Monson-Haefel [Ano99-72]. Monte [GHM96, JI94, Kou98, Mic90, SC98b]. Monte-Carlo [JI94]. Montego [Ano97-141]. Mooney [Ano99e]. Moorby [Ano99-227]. Moore [BKM95]. moral [Ano99-140]. morality [Ano99-140]. Morisse [Ano96-120]. morphology [Ola92, Ano93l]. Morris [Ano90o]. Morten [Ano98-34]. Mortenson [Ano99-150]. mosaic [Ano95-80]. Moses [Ano96-138].

Mösges [Ano96-43]. Moshe

[Ano96-138, Ano97-46]. Moss [Ano97-162]. motif [Köh94, Ano95-52]. Motifs [Köh96b]. motion [Ano99-118, Ano99-119, AB93, KNUX99, KPS96, RR90, Sak90, VV99]. motions [SJS92]. Motoda [Ano99-80]. motor [Ano98-114]. motors [Yav93]. mountain [HH98a]. movement [IA94]. Moving [Chi93, Pfi98, Sal92a, AO98c, Ano90-28, KKS98, Kle93, LS98a, SM90, TH95, YC92, Yav92]. Mozer [Ano97g, Ano98h]. MPI [Ano95-82, Ano99-158, Ano99-160, Ano99-159, Ano99-161, Ano96-118, Ano99-158, Ano99-160, Ano99-159, Ano99-161]. MPI-2 [Ano99-160, Ano99-161]. MPI-The [Ano99-160, Ano99-161]. MR1096131 [Moo93b]. mRNAs [LNM90]. mSQL [Ano99-164]. Muelen [Ano97-64]. Mueller [Ano97-161]. Mullen [Ano99-65]. Müllges [Ano97-111, Ano97-112]. Multi [Ano96-120, Ano99-128, LD90, LL90a, Lin95, MZMQ90, Ser93, And93, Ano96-135, Ano97n, Ano98-50, ANV94, BM90b, CCK97, Coh94, FK95, Gla92e, GK98, Kat92, PR93, Tan93d]. multi- [GK98]. Multi-Agent [Ano99-128, Ano96-135, Ano98-50]. multi-class [Kat92]. multi-commodity [BM90b]. multi-component [Gla92e]. Multi-criteria [LL90a]. Multi-dimensional [Lin95, ANV94]. Multi-Grid [LD90, PR93]. multi-issue [CCK97]. Multi-iterative [Ser93]. multi-modal [Ano97n]. multi-oscillator [FK95]. Multi-processing [Ano96-120]. multi-rate [And93]. Multi-stage [MZMQ90, Coh94]. multi-step [Tan93d]. multiagent [Sti97]. multiarmed [Yos93]. multibody [RS99c]. Multicompartmental [AÉ96]. multicriteria [Shi95a]. multicultural [Ano96-65]. multidatabases [BQN97]. multidelay [ZZ99b]. multidelay-differential [ZZ99b]. Multidimensional [AL94, Bar94, RK98, Woo91, Che93a, Faz94b, KP97b, Leo90, Pap94a, RZ96]. multidisciplinary [CM93, CP95b]. multidiscipline [Bes96]. multidomain [Yan93a]. multifaceted [Opp96]. Multifractals [Ano99-162]. multigrid [Gás95, KK97b, Mol96, Oli98, PR90, SM98, SS91b, WF99a, YK95]. multigroup [CCY93]. multilateral [Ano98-136]. multilayer [BIZ97]. Multilevel [LB94, Ano97-41, CP95b, IK95, Liu98, SNU98, Ano98-108]. multilingual [Ano99-142]. multimedia [Ano97-131, Ano98-47, Ano98-77, Rot96, Ano96-119, Ano98-109, Ano98-110]. multinephron [SMTS97]. multinomial [Lon95]. Multiobjective [XLJS98, BJH99, EB93, FCP95, GW99, Liu96a, LY99b, YY98]. Multiparameter [Gur91, Ame98]. multiparty [LY99a]. multiplatform [Ano95-72]. Multiple [AO99f, AO99e, CC98a, DELE99, CCY93, Gün99, HS95, HGWL96, Iri98, Ita93, JKS94, KT97, KT98, KS92, Kha97, LL91a, LB94, Liu98, PN98, PBG90, Rán92, RSS94b, SMT98, Tan92, ZG97]. multiplication [CDLW95, KH92, MMMS97, MMMS96, Nis95, PC96b, Ano97-121]. multiplicities [NP96]. multiplicity [SV91a, SV91b]. multiplier [Tan92]. multipliers [AB94, HW93a]. multipoint [PSLT93]. Multipole [McK96, GG90, LJ98]. multiprocessor [LC94, PS93b, Szt92]. multiprocessors [Ano93w]. Multiprocessorsa [DH97]. multiquadratic [Bax92]. Multiquadric [BM92, Sau92, CF91a, CN94, Har90, Har92a, Har92b, Jek94, KC92a, Mad92, WHL^{+}99]. multiquadric-biharmonic [Har90, Jek94]. Multiquadrics [Kan90a, Kan90b]. multirate [De 96]. multisectoral [Pan92b]. multisegment [NAKA98]. multisplitting [Ano00, BW96a, BW96b, Bai96d, BSW96, BSS98, Wan99]. multistage [Eso91]. multistart [Pic91]. multistart-annealing [Pic91]. multistep [AE90, BT98, CV90, Coh94, FM96, MRJ97]. Multistride [MHBD96]. multivalent [JS99, KNS99, SHA99b]. Multivalued [CY95, O’R99, AO99b, Ano98-67, Ano99-49, HB99, JY93, Mro96, Noo95, Noo96, Noo98, Wu99b]. multivariable [CKL91, Trz92]. Multivariate [BIZ97, PS95, Ano99-28, MSB90, Nar92, PZ99b, Riv97]. multiwavelets [ANV96]. Multizone [WHL^{+}99]. Mumford [Ano99-231]. Munzel [Ano99-140]. MuPAD [Ano99-69, Ano96-120, Ano96-120]. Murray [Ano96-99, Ano97q, Ano97-134]. Musciano [Ano96-88]. Music [Rot96, Ano96-153, Ano97-40, Bec96, LL90b, MM96d, Opp96]. musical [Far96b, Nel96, Wit96b, Ano99-163]. Musick [Ano98-139]. Musicology [Ano97-40]. mutual [FLNW98, Har92c]. My [Ano95n, Los92]. MySQL [Ano99-164]. mystery [Ano99j].

N [Ano96-53, Ano96-117, Ano97-66, Ano97-100, Ano97-94, Ano98-157, Ano98-159, Ano99-202, LL91b, CTG96]. N.J [Ano90p, Ano90q, Ano90s, Ano90x]. N.Y [Ano98-118]. Nadel [Ano97-81]. Nagar [Ano98-163]. nagumo [Cho92, Cho95, CL96b]. Naik [Ano99-28]. Nail [Ano99-71]. Naish [Ano98-96]. Namir [Ano96-112]. Nancy [Ano93k, Ano95-64, Ano96-45, Ano98-143, Ano99-141]. nanometer [TNS98]. nanometer-scale [TNS98]. Nanotechnology [Ano96-121]. Naomi [Ano95x]. Narasimhan [Ano95-66]. Narayanan [Ano96-53]. narrow [MPS92, Sto99, YFZM95a, YFZM95b]. NASA [Ano96-152]. Nash [Kos93, Liu98]. Nass [Ano97-148]. Natick [Ano99-231]. National [Ano96-127, Ano99-89, Ano99-146, Ueb92]. Nato [Ano98-93]. Natural [HS97, Sok91, Ano95r, Ano95s, Ano96-39, Ano96-140, Ano97k, Ha90, KHC92, Lyt92b, MB93, Sau92, SV91a, SV91b]. nature [Ano96-105, Ano97-96, Ano99-220]. Naughton [Ano98-45]. Naum [Ano99-167, Ano99-168]. Naundorf [Ano96-120]. Navier [AF90, CS94b, Fai98, GP91, HM91, HW96, Lay93, Rac99, Sak90, SCBFC93, SCBFC98]. Navigation [Yav96, YFZM95a, YFZM95b]. Navy [Kra90]. Nayebi [Ano96-147]. NC [Ano99-28]. NC/New [Ano99-28]. Neaga [Ano93o]. Near [Gün99, Ano99-239, GK91, KKB97, KCB90, MPM93]. nearest [YC94]. nearfields [Ano98-111]. Nearly [KP97b, Var98, KHI94, Kor90a, Kor91, Pan96, PZ99b]. nearly-analytical [KHI94]. nearrings [Ano98-111, Ano98-111]. Necas [Ano97-165]. Necessary [Dor90, ST95, Zaf98b, ZL99a, Ano98-36, Tan93b]. neck [CCH^{+}90]. Nederpelt [Ano95-65]. negative [AIS95a, KNS99, RS99b, SA95]. neighbor [YC94]. Neil [Ano97u, Ano99-60, Ano99-180]. Neittaanmäki [Ano99-74]. Nemhauser [Ano99-127]. nephron [WSDT94]. Nerode [Ano97-86]. nerve [Cho95, Tau90b]. Nesetril [Ano99-135]. nested [BHK95, GT92, JS97b]. Net [SJS92, KS99]. NetActivism [Ano97-101]. Netherlands [Ano96-31, Ano96-83, Ano97e, Ano97-29, Ano97-45, Ano97-56, Ano97-66, Ano97-85, Ano97-84, Ano97-108, Ano97-157, Ano97-164, Ano98o, Ano98j, Ano98i, Ano98w, Ano98-59, Ano98-65, Ano98-89, Ano98-90, Ano98-108, Ano98-100, Ano98-101, Ano98-111, Ano98-126, Ano98-127, Ano98-157]. NETL [Tho92]. netplay [Ano98-112]. nets [Ben99, Bol92, GS96b, KKOS96, MR99, Nis97b, Nke94, ZK99]. netscape [Ano98-149]. NetSuccess [Ano97-102]. NetTravel [Ano97-103]. Network [Sti93b, Sti96, Sti97, Ano93m, Ano95-83, Ano97-76, Ano97-91, Ano99-92, Ano99-242, BGM95, CT90, CL99b, Cho99, HTT92, Iri98, IIP^{+}97, JS97b, KS99, MATA98, MPP97, MNE92, Mar90, May99, SHT^{+}90, SZ98a, SSBG91, Sti94, Ano98-112]. network-flow [Iri98]. networked [Ano95-59]. networking [Ano99-120, Ano99-221, Ano96y, Ano97-104]. networks [Ano95m, Ano95l, Ano95-53, Ano96-99, Ano96-122, Ano97-106, Ano97-52, Ano98-114, Ano98-129, Ano98-133, Ano99-163, Ano99-209, BLR96, CL99a, CT90, CF98, EHK92, Fau94, GCT98, Hau92, Hen92, KHC92, LS96b, LJ98, Leh92, Lev92, Lia99a, Mar92, MKS^{+}91, RA92a, Sha92, Tan96, Wan93b, WR93, YF93, Ano95q, Ano96-74, Ano96p, Ano96-145, Ano97-44, Ano98-104, Ano97-105]. Neubrunn [Ano98-74]. Neumann [Ano95v]. Neural [Fau94, Ano95m, Ano95l, Ano95q, Ano95-53, Ano96i, Ano96-74, Ano96p, Ano96m, Ano96-145, Ano97g, Ano97-44, Ano97-91, Ano97-52, Ano98h, Ano99-92, Ano99-232, CF98, Chu99, FT97a, GCT98, IIP^{+}97, KBLS90, MPP97, Mar90, MKS^{+}91, RA92a, SI94, Wan93b, WR93, Ano90v, Ano93m, Ano96-122, Ano97-106, Ano98-113]. neuro [Ano99q, Eso99b]. neurobiological [Ano99-118, Ano99-119]. Neurocomputers [Ano95-53]. neuronal [Ano95-41, Ano98-104]. Neurons [Ano98-114]. neurophysiological [RT96]. Neuropsychology [Ano99-92]. neuroscience [Ano98-28, Ano99-92]. neutral [AG99c, CLY95, DH98a, DBA98, KP97a, Lal94, LC98a, LY98, YCZ98, ZZ99b, ZYW98]. Newport [Ano96-134]. Newton [Ano95i, LLP94, Arg99, EDD91, FM96, Gil91, Han92, KT92, KT95, KN94, Mil95, Ohn92, Pan92a, She93, ST94, WT93]. Newtonian [Gre93b, TH95]. next [Ano99-222, Ano99-223]. NEXTSTEP [Ano96-149]. Nguyen [Ano98o, Ano98n]. Niall [Ano99-163]. Niche [GF96, GF94b]. Nicholas [Ano93f]. Nichols [Ano99-144]. Nick [Ano95-72]. Nico [Ano96-146].

Nicolson [CCVV90, EGLN95, Sir99]. Niederreiter [Ano97-48, Ano97-47]. Niederst [Ano96-51]. Niemeyer [Ano96-70, Ano98-54]. Niguidula [Ano95-57]. Nikola [Ano97-52]. Nikolas [Ano98-67]. Nikolay [Ano99-123, Ano99-124]. Nina [Ano99-212]. nineteenth [Ano97-151]. Ninety [Ano96-127]. Ninety-fifth [Ano96-127]. Nipkow [Ano98-147]. Nishidate [Ano97-96]. nitrogen [Kou98]. Nitzberg [Ano99-160, Ano99-161]. NJ [Ano95t, Ano96j, Ano96-102, Ano96-136, Ano96-137, Ano97-113, Ano97-114, Ano99-215, Ano99-185]. NLG [Ano97-65, Ano98g, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98w, Ano98x, Ano98r, Ano98z, Ano98-37, Ano98-31, Ano98-32, Ano98-36, Ano98-45, Ano98-49, Ano98-53, Ano98-50, Ano98-47, Ano98-59, Ano98-56, Ano98-57, Ano98-65, Ano98-66, Ano98-68, Ano98-67, Ano98-69, Ano98-72, Ano98-76, Ano98-84, Ano98-74, Ano98-83, Ano98-89, Ano98-90, Ano98-92, Ano98-109, Ano98-108, Ano98-100, Ano98-101, Ano98-110, Ano98-111, Ano98-116, Ano98-126, Ano98-127, Ano98-133, Ano98-132, Ano98-135, Ano98-144, Ano98-138, Ano98-139, Ano98-140, Ano98-155, Ano98-156, Ano98-161, Ano98-164, Ano99z, Ano99-27, Ano99o, Ano99r, Ano99s, Ano99n, Ano99t, Ano99u, Ano99-33, Ano99-50, Ano99-213, Ano99-49, Ano99-63, Ano99-64, Ano99-52, Ano99-57, Ano99-58, Ano99-84, Ano99-85, Ano99-86, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-87, Ano99-105, Ano99-103, Ano99-99]. NLG [Ano99-96, Ano99-102, Ano99-104, Ano99-100, Ano99-97, Ano99-95, Ano99-101, Ano99-115, Ano99-116, Ano99-117, Ano99-120, Ano99-114, Ano99-128, Ano99-121, Ano99-136, Ano99-137, Ano99-145, Ano99-142, Ano99-152, Ano99-149, Ano99-153, Ano99-167, Ano99-168, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-170, Ano99-174, Ano99-183, Ano99-184, Ano99-181, Ano99-182, Ano99-196, Ano99-197, Ano99-201, Ano99l, Ano99-206, Ano99-207, Ano99-204, Ano99-209, Ano99-229, Ano99-212, Ano99-211, Ano99-227]. NN [CL99c]. No [Ado91a, Sal92b, Zad92, Lay99, MM94, Moo93b, MG90c]. no-slip [Lay99]. nodal [GK91]. node [Pan98c]. nodes [GGK96, Kan92a, SZ91]. Noise [AR92c, AV98, Ano99-162, Kis91]. noisy [Ram95]. Non [Tan93c, AS93, Ano97-162, Ano97-130, Ano99-136, Bör90, CY94, Gla90a, Gla91a, Gla91c, Gla91e, Gla92a, HAO91, Il’97, KP99a, Nov90, Ram90a, Sun92, Tan92, TH95, WT93, Ano98-115, Ano99-166]. Non-commutative [Ano98-115]. non-compact [CY94, Sun92]. Non-conservativity [Tan93c]. non-convex [Bör90]. non-Darcy [KP99a]. non-equilibrium [Gla90a, Gla91a]. non-homogeneous [AS93]. non-ideal [Gla92a]. non-linear [Ano99-136, Nov90, Ram90a, WT93]. non-Mersenne [Tan92]. non-Newtonian [TH95]. non-radial [Ano97-130]. non-self-adjoint [Il’97]. Non-standard [Ano99-166]. non-uniform [Gla91c, Gla91e, HAO91]. non-wellfounded [Ano97-162]. nonadiabatic [KKD^{+}97]. nonassociated [vdVVdB99a]. nonautonomous [LFP97, PC94b, SWY97, SC98a, TC99, YCC96, ZZ99e]. noncentral [Tem93]. nonchaotic [Nas92]. noncoercive [Bou98]. noncompact [Hua98b]. nonconcave [ZC91]. Nonconforming [YK95, KP98]. nonconvex [Ano99l, BF91, Chi91, WCHL96, Ano98-116]. noncooperative [Bel97b, KK91, Nak99]. nondegenerate [CST97]. nondifferentiable [Liu96a, Ano99-167, Ano99-168]. nonequilibrium [RP97]. Nonexistence [ZT98a]. nonexpansive [Hua99]. nonfinite [Ano98-151]. nonhomogeneous [BL94, Far96a]. Nonhydrostatic [LMW95]. nonhyperbolic [PS93a]. nonintegrable [Mur93b]. Nonlinear [AR92d, CK94b, HW93b, MB94b, NIS97a, OHI91, Pad92b, RGBL96, WK94, WY99, Yao92, AR90a, AR90b, Ado91a, Ado91b, Ado91d, Ado96, AO98b, AH98, AO99d, Ahm94, Ahm95, Alt90a, Alt90b, Ame97a, Ame97b, Ame98, Ano98o, Ano98-165, Ano99r, Ano99-96, Ano99-97, Ano99-181, Ano99-243, BKN97, BBKN97, BW96a, BW96b, Bai96c, Bai96e, Bai96d, Bai96f, BM96, BLP95, BR95c, BA96, Bon97, BPLS96, Cha98, CCCY94, CL96a, CZ94, CLD97, Chi91, CC98b, CK91, CK94a, DH98b, DBA98, DNJS98, Din97a, Din97b, Dor97, DIR97, Ech95, Ech98, EK97, EGHS98, Far91a, Far91b, FK97, FK95, FT98, Gal99a, GZ91, GS93, Goh94, GRKM96, GCT98, GL93, GL94, Gyö90a, Han92, HH98a, HM93b, HMZ99, Hua98a, HLTB99, Isp98, JPC99, JI94, KS98a, KT92, KT95, KP95b, KKG97, KPS96, KS97b]. nonlinear [Lee94, LTC98, LTR98, LY94, LC98a, LC98b, LF99, Li99a, LD93, LD95, LOD96, LTB96, LM91, MO98a, MDVK97, MH93a, Mur92a, NAS97, O’R98a, O’R98c, OCS96, Oza94, PXHG99, Pom91, Pom92, Prn97, Ram90b, RV94, Ran93, RTJ96, Rèp91, RZ95, Rob97, Scr91, Sha90a, Sha90b, Sha96b, SWY96b, SWY97, SAS99, Tah91, TAD91, TVZPG96, TWIF98, Tha94, TSL96, TP98, TG96, Ush94a, VR95b, WA96b, XLPG99, Yan99b, YLY96, ZKÖ98, Zaf98b, Zha93b, ZY98, ZL99b, ZY96, ZS97, Ano96-123, Ano97-108, Ano97-107, Ano98-117, Ano99-169, Ano99-170]. nonlinearities [Ram91a, Wan98a]. nonlinearity [EY91]. nonlinearly [Kos90]. Nonlocal [KT90, DLS99, EY91, Moi97]. nonmodal [dCVC97]. nonmonic [Trz92]. Nonmonotonic [Ano97-109]. nonmonotonicity [Ano97-117]. nonnegative [RS96b]. nonnegativity [SMU98]. Nonoscillation [LY94, YCZ98]. nonoscillations [WA98]. nonoscillatory [CYS94]. Nonparametric [CL99c, IZ97, Zhu92]. nonparaxial [MDVK97]. nonperiodic [Shk97]. nonpositively [Ano97-59]. nonpredictive [Wan90]. nonpreemptive [LC94]. nonreliable [Alm96, AS99]. nonresonant [MO98b]. nonscalar [Ech98]. nonseparable [LH91]. nonsingular [AO99f]. nonsmooth [Liu96b, Ano99-171]. Nonstationary [KKD^{+}97, BGM97, Kor90a, Kor91, VA94]. nonsymmetric [Rac99, YK95]. nontransitive [GWL96]. nonuniform [CHSS95, PR96]. nonuniformities [VHH95]. nonzero [PPC99]. noon [Ano96-85]. Norbert [Ano98-138]. Norm [Lei96, MSR95, LX98, RT94, dOF91]. Normal [TNTR91, EA92, Köh93a, Mil91, NGNM97, Var98]. normality [Hal90]. normalized [CL95, PT91]. Norman [Ano96-108, Ano99-216, Ano99-217]. norms [Ahm90, Lub97]. North [Ano93h]. North-Holland [Ano93h]. Norton [Ano96v]. Norwood [Ano96j]. Noss [Ano93j]. Notation [Has95, Has93b, Has93c, Sni90]. note [Bai96a, BC98, Car94, Cha93, CP98b, Fil93, HM95, Kov95, LQ95b, Li96b, MK91, Nar92, QZ93b, RSS94a, RS95, RGBK91, SCC97, Zha96, Zhu92]. notebook [Ano99-78, Ano99-79]. notebooks [Ano97m]. Notes [Bel94b, Zha98, Ano97-107, Ano95-54, Ano96-124]. nothing [KM96]. notions [CR90c]. novel [BJH90]. Novo [LL90a, Shi95a]. NT [Ano97-169, Ano98-51, Ano98-98, Ano98-163]. nth [Rua91]. nth-order [Rua91]. NuAlgebra [Nam96a]. nucleation [Gre91b]. nucleon [KLMY97]. Nucleotide [DB96]. Null [DBA98, Kou98]. null-collision [Kou98]. NuMachine [Nam96a]. Number [Ano98-43, MZ93, SL99, AG91, Ano92a, Ano96-157, CL91a, CL91b, Chr90, Dwy93, EDO97, GF94b, GF96, GM99, Has93a, IB95, JP93, KT97, KW98, KT98, Pav97, PW95, Roj99, Sez95, SD93, Ste90, Ves98, Wag90]. numbers [Ano97-119, Ban93, BGZ92, CL91a, CL91b, HP94c, Koç91, KH92, KO93, LS94, MB93, NL93, Yan95]. Numerica [Ano97-110]. Numerical [ARM91, AO98c, Ano98-46, Ano99-43, ASW98, Ber93, BMB94, CFM91, Car93, CG94, DBM94, FLW99, FG93, Feu91, GS93, GGC97, Has91, HH98a, Kim99, KG95a, Kos90, LM96a, MT92, MM95a, MM96c, MDVK97, Mil98, MT98, Moo94, Mur92a, QZ93a, Ram93, RL91, Sha90d, SSZ91, SCBGN94, TSLW98, Tör94, VL92, ZSDB99, ZR91, hCPW98, AE90, AO91, APV97, And93, Ano93i, Ano96-75, Ano96-83, Ano96-131, Ano97-65, Ano98-101, Ano99-83, Ano99-113, Ano99-171, Ano99-245, AS96b, Bab98, Boy99, BPLS96, But94, CS96, Cas94, CMMR92, CAZAS95, CKAS96, CCVV90, Cor94, CKRV98, DGV94, DG91, DJV96, DW97, Ech95, Ech96, Ech98, Far96a, FH99, FLW97, Feu90, FK98a, FK98b, Gar96, Gla94b, Gla94d, Gla94c, Gla95b, GRKM96, GSC98, Gre95, Ha90, HR99, HP94b, HAO91, Huj99, HLS91]. numerical [HWL94, JR94, JG95, JP95a, JP95b, JA96, JG96, JD97, KU99, KG94, Lar90, LDD94, LT96, LZ98, Mac93, MG91, Mar94, MP96, Mat99, MT94b, MM95b, MHBD96, Mey99, Mil90, MG90a, MG90b, MGM92, MMZ98, NPJ95, Oli98, PYS97, PS96, Pav97, PW97, SJB97, Say99, Sch99, SD95, Sha90b, Sha99a, SM90, She99, Sim93a, SM94, SM95, Sim97, Sim98, Sod92, SMU98, Sug94, SBS99, Tah91, TPF98, Tem93, Wan99, Yav92, ZM99, tC95, vDHO^{+}95, vdVVdB99b, Ano90w, Ano93n, Ano96-125, Ano97-111, Ano97-112, Ano98-118, Ano99-172]. numerically [Bou99, CB93b]. numerov [Sim97, BMD92].

Numerov-like [BMD92]. numerov-type [Sim97]. NURB [Ano95-55]. Nusse [Ano98-46]. nutritional [TNTR91]. nutshell [Ano96-92, Ano97-169, Ano98-79, Ano98-103, Ano98-130, Ano99-31, Ano99-62, Ano99-143, Ano99-187, Ano99-210, Ano99-237]. NY [Ano98-41, Ano98-158]. Nye [Ano95-46, Ano96l]. Nyström [CJS90, Pat98, QZ91, Sim93a, hC96a, hC96b].

O [Ano96-107, Ano96-120, Ano98l, Ano98-136, Köh95, LL91b, WSNS98]. O.D.E [Ano97-147]. O29 [CCH^{+}90]. Oaks [Ano97-77, Ano98-88]. Oara [Ano99-98]. OBDD [Ano99v, Ano99w]. OBDD-Foundations [Ano99v, Ano99w]. Object [Ano96-84, Ano95-50, Ano95-81, Ano96z, Ano96-109, Ano97-34, Ano99-125, Ano99-126, LYL^{+}91, LCL94, MCG98a, MR99, NB95, Yav92, Yoo93, Ano95-56, Ano95-57, Ano98-119]. object-oriented [Ano96z, Ano96-109, Ano97-34, Yoo93, Ano95-57]. object-orientedness [MR99]. objections [Ano96-50]. objective [LL91a]. Objects [Ano96-119, Ano97-32, LT94, Ano98-40, Ano98-158]. oblique [MK91]. Obradovic [Ano96m]. Observability [ARS98, AH97, De 96]. observation [CSOL96, CKK97a, Har96, Zou91]. Observations [Ush94a, OSLC98]. observatory [Ano99-239]. observed [Kra92]. observer [XLG95]. obsolescent [DO95]. obstacle [MMZ91, MRZ96, YP98b, YZM94, YZM95, ZZT99]. obtain [KM96]. obtention [MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b]. occupation [KK94]. Ockman [Ano99-173]. odd [LC98a]. ode [And91, AO91, BP94b, Han92, ZS97, dSB96]. ODE’s [And93, Gar96, CC94b, CAZAS95, CKAS96, HS94, Lop93, MHBD96, Sha98, Sha99a, Söv99, VA97]. Oevel [Ano96-120]. off [Tso91]. offering [Ano92a]. Office [Ano98-120]. oil [SJB97]. oilfield [XLJS98]. old [Gal99c]. OLE [Ano98-149]. olfactory [AÉ96]. Oliger [Ano96-164]. Olin [Ano96-130]. Oliver [Ano98y]. on-line [GF97, TTC96]. One [AAL99, ABK97, Byk97, HM93a, AG91, AR99a, Ano97-127, Ano00, BSS98, BBI97, CHSS95, CB93a, CAZAS95, DM95, DKR97, Dri97, FG93, FG98, GF94a, GG97, Gla95a, GB96, GC99, HNS^{+}90, Hun95, IK97, KESH90, MAM91, May99, NN91, PJ91, SSZ91, WCY94, Wu98, Yan93a, ZM98, WS92]. One-dimensional [HM93a, Ano00, BSS98, BBI97, DKR97, Dri97, FG93, Gla95a, GB96, GC99, HNS^{+}90, IK97, NN91, SSZ91, ZM98]. one-grid-overlapped [Yan93a]. one-phase [Hun95]. one-step [CAZAS95, Wu98]. one-to-one [GF94a]. one-way [CB93a, WCY94]. O’Neil [Ano99-34]. online [Ano96-99]. onto [CP95a, TSLW98]. oocytes [BCB90]. Oomens [Ano98-101]. OOP [Hen97]. Open [Yav91a, Ano99-173, Gla92c, Gla95d, GM99, KKG97, Say99, YLC96, YL99]. Open-loop [Yav91a]. OpenSources [Ano99-173]. operating [Ano99-192, Pfi98, Szt92]. operational [LS95]. operations [Ano97-45, Bol92, LO99, Pan93a, SYZ94, YJ99, Ano98-153, Ano97-113, Ano97-114]. Operator [Gus97, AO99d, Ame97a, AZ96, Ano96d, Ano98-89, AI97, Din97b, Fli94a, Gla91a, HSH90a, Ike91, Ism97, LD93, Maz90, MO98b, MO98a, O’R98a, Sán98, Tse95, WW98, ZA95, Ano98-121]. operators [AdlC95, AIS95b, AC93b, BAS97, BK98, BRV98, BCD98, CIS98, Coo98, CY95, Doš98, GGK96, GZ95, GlZ95, Il’97, JOW98, LC96b, MM96b, Nis95, O’R95, RS96a, RS99a, RZ94, SA95, Wan90]. Oppenheim [Ano98-159]. Optimal [AW91, BA98, Cha98, Che99a, CF94, Cra98, DO92, DO95, IN92, KO90, Kou95, KS97b, KS98b, LO93, ML90c, NMUY92, Pan96, Pan92b, Pap92, PS93b, RTJ96, Rei94, Sad96a, Saw92, SOI92, SZC91, Tad94, Yav91b, YF92a, Yos93, AO93a, Bed95a, BF95, DNJS98, EGEHSEK95, FH99, GM96, Gün99, GHS91, HW93a, HM91, KK92, KK94, LTC98, Lew91b, LOP95, MJK92, MMMS97, MMMS96, NR96b, Oda94, OHI91, RDK92, SKI92, SFY92, SSBG91, YT98, Yan93d, Yan99b, YF92b, Yos96, Ano99-174]. Optimality [Ill91, Liu96b, LY99b]. optimally [DdW93]. optimisation [SFY92, Woo91]. Optimising [Rei99a]. Optimization [AS99, CPC93, SHK99, Ano96-123, Ano98-108, Ano98-116, Ano98-124, Ano99-101, Ano99-151, Ano99-167, Ano99-168, Ano99-171, Ano99-246, AKV91, BF91, BS91a, BCB90, CO99, CM93, CP95b, CT90, CCY95a, CCY95b, Chi91, CC96, Dei93, FZ93, FCP95, FWI99, GZ91, Gal91b, GPZ93, GZ93, GW99, GS96a, GPK99, HW93b, HM94, JTY97, Kat92, KK91, Kos93, KVCA^{+}99, LH91, Liu97, Men91, Moo91, Moo93b, Oll90, Par91a, Pic91, PL90, PN99, RR91, SZ98a, TZ93, Zhe91, Ano96-117, Ano97-110, Ano97-164, Ano98m, Ano98-29, Ano98-123, Ano99-127, Ano96-126, Ano99-175]. optimizations [YTLK94]. Optimized [Amo93]. optimizing [Ano99-87, Ano99-242, MATA98]. Optimum [GMW96, BS99b, Dei91, KHI94, PZ99a, Shi95a]. optimum-path [Shi95a]. option [HL90]. option-based [HL90]. options [DO92, Saw92, TDF^{+}92]. O’Quinn [Ano98-130, Ano99-187]. Oracle [Ano99-176, Ano99-177, Ano99-178, Ano99-179]. Oram [Ano97-124]. orbit [SSN94]. orbits [FS95, GSST97, Mag97]. Order [Gol94, Pat98, Raq98, ZQ93d, AH98, AO99f, AO99e, AG99b, AP94, AO93b, AO93a, Ano96-72, BF97, BS90, BTT90, Car92, CHSS95, CP98a, Cha91, CJS90, CE94, CL96a, CS93b, CS94b, Cho94, DBM92, DBM94, De 99, Doš98, EDO96, ED91a, Gal93d, GS91, Gla93e, GL93, HBB90, HP94a, Isp98, JK96, JV96, JPF96, Ker91b, LC98a, LY98, LF99, LX98, LR96b, LG99, LS95, Mak90, MMP98, MG91, Mar94, MT92, MHBD96, MK93, MRJ97, OCS96, PS96, Paz90, PXHG99, PR96, Prn97, QZ91, Rua91, Sán98, Sar90, Sar94, SA94, Shi98, Sim93b, Sim93a, Söv99, Sto92, Tha94, TSL96, TP98, TMGS98, TR99, Trz92, Tsi99, UW94, VK95, WCC99, Wan98b, WJ99, WA96a, Wu98, ZKÖ98, Zaf98a, Zaf98b, Zha93b, ZL98b, ZZ99c, ZY96, ZQ93b, ZQ93c]. Order-increasing [Gol94]. order-of-scattering [UW94]. ordered [Ita93]. ordering [CF91b, KO90, Ano98-74]. orders [CMR90, GF90, Ano98-115]. Ordinary [MPS94, Ano99-33, Ano99-71, BA96, LP99, LNG92, MP91, MK93, QZ91, QZ93b, Sar90, Sar94, SAS99, Wan96]. ordinates [Tch98]. Ore [Pal97]. O’Reilley [Ano95-52]. O’Reilly [Ano95-76, Ano95-46, Ano95-80, Ano95-58, Ano95-63, Ano96w, Ano96u, Ano96-42, Ano96-28, Ano96-49, Ano96-51, Ano96-70, Ano96-59, Ano96-88, Ano96-92, Ano96-93, Ano96-133, Ano96-166, Ano97q, Ano97-38, Ano97-35, Ano97-75, Ano97-77, Ano97-76, Ano97-74, Ano97-78, Ano97-89, Ano97-101, Ano97-102, Ano97-103, Ano97-123, Ano97-124, Ano97-131, Ano97-132, Ano97-166, Ano97-169, Ano97-167, Ano97-168, Ano98t, Ano98-52, Ano98-54, Ano98-51, Ano98-79, Ano98-88, Ano98-86, Ano98-85, Ano98-99, Ano98-97, Ano98-98, Ano98-103, Ano98-120, Ano98-129, Ano98-122, Ano98-130, Ano98-163, Ano98-162, Ano99y, Ano99-31, Ano99m, Ano99p, Ano99-35, Ano99-36, Ano99-61, Ano99-62, Ano99-72, Ano99-139, Ano99-138, Ano99-143, Ano99-141, Ano99-164, Ano99-148, Ano99-147, Ano99-221, Ano99-173, Ano99-176, Ano99-177, Ano99-178, Ano99-179, Ano99-180, Ano99-195, Ano99-187, Ano99-188, Ano99-192, Ano99-230]. O’Reilly [Ano99-210, Ano99-235, Ano99-236, Ano99-238, Ano99-237]. organization [Ano95x, Ano97-43, Ano98-113, Hol90]. organizations [Ano99-128]. organizing [NIS97a]. orientation [Mis90]. oriented [Ano95-73, Ano95-81, Ano96z, Ano96-109, Ano97-34, Ano99-117, CHCC95, Yoo93, Ano95-57]. orientedness [MR99]. original [Ano93o]. origins [Ano96-158, Ano99-203]. Ornstein [AKS96, ABI99]. orthogonal [Ano98-165, DIT97, Gem93, GNR95, Gla93d, Has93b, KO97, MMP98, MO95, Rob97, WK94]. orthogonalisation [MNB93]. orthogonally [NAKA98]. orthonormal [Bis91, Lub97]. orthotropic [DPD98, MP93, MB92, MB94b]. Orwant [Ano99-148]. Osborn [Ano96-54, Ano98-42]. Osborne [Ano95d]. oscillating [PS96, Sim93b, Tsi99]. Oscillation [AG99a, AG99c, Ano99-243, CE94, CLY95, FT98, GL93, Gra94, GT99, KP97a, Lal94, LC98b, LY98, LF99, SWY96b, TY99b, Tha94, TSL96, WA96b, YCZ98, ZT98b, ZYC98, ZL98a, ZL98b, ZL99b, ZZ99c, AG99b, Pop94, ST98, TY99a, Zaf98a, Zaf98b, ZY98]. Oscillations [KL96b, Rua91, Sta95, WA98, Ano97-130, ZL99a, Ano98o]. oscillator [FK95, Sha96b]. oscillators [BF97, Gur91]. Oscillatory [MP93, ZY96, Cas98, KC91, LYL95, Mil98, PXHG99, TP98, TXC98, VR95b]. osculatory [RK98]. Oseen [HM99]. Osherson [Ano98-48]. Osiris [Ano97-128]. Osterbrock [Ano99-239]. Ostrowski [DW97, Dra99, Fin99]. Ostrowski-Grüss’ [DW97]. Otfried [Ano98-33]. other [Ano96-57, Ano96-87, Ano97-89, Ano97-130, Ano99-215, Tse95]. otolith [CMMR92]. Otto [Ano96-118, Ano96-164, Ano99-89, Ano99-158, Ano99-159]. OUP [Ano90u]. our [Ano95-76, Ano97-143, Ano97-146]. Oussama [Ano93v]. out-of-core [DHW98]. outer [BH91]. outline [Pol90]. output [AH97, MW99]. Outrata [Ano99-171]. overall [Bes96]. overdetermined [Kim98]. overflow [Pou90a]. overlapped [Yan93a]. overlapping [Sir99]. Overmars [Ano98-33]. overseas [Ano97-141]. overturning [GT95]. overview [AK90a, Ano95-53, KZ95, SH96]. own [Ano97q]. Oxford [Ano90u, Ano93g, Ano93r, Ano97-118, Ano98y, Ano98-55, Ano98-118, Ano98-124, Ano99-135, Ano99-222, Ano99-223, Ano99-205]. ozone [ZFM96].

P [Ano90w, Ano93h, Ano93w, Ano95t, Ano95-65, Ano96t, Ano96-83, Ano96-154, Ano96-120, Ano96-115, Ano96-160, Ano97e, Ano97-65, Ano97-69, Ano97-66, Ano97-84, Ano98-36, Ano98-60, Ano98f, Ano98-92, Ano99d, Ano99-74, Ano99-84, Ano99-85, Ano99-117, Ano99-131, Ano99-146, Ano99-182, Ano99l, Ano99-211]. PA [Ano95-27]. Pach [Ano96-33]. package [DF98, SH96]. packages [Ano97-90]. Packel [Ano97m]. packing [GMW96, MCG98a, RI94]. Padberg [Ano97-84]. Padé [Kov95, PB95]. Padhraic [Ano96h]. page [Ano98-162]. pages [Ano90o, Ano90p, Ano90q, Ano90s, Ano90u, Ano90w, Ano90x, Ano93f, Ano93g, Ano93t, Ano93u, Ano93h, Ano93i, Ano93j, Ano93l, Ano93k, Ano93n, Ano93m, Ano93o, Ano93p, Ano93v, Ano93r, Ano93s, Ano93w, Ano95k, Ano95-36, Ano95h, Ano95m, Ano95l, Ano95-70, Ano95n, Ano95o, Ano95r, Ano95s, Ano95d, Ano95q, Ano95p, Ano95t, Ano95-72, Ano95v, Ano95y, Ano95u, Ano95x, Ano95c, Ano95w, Ano95-71, Ano95z, Ano95-28, Ano95-27, Ano95-29, Ano95-74, Ano95-73, Ano95-30, Ano95e, Ano95i, Ano95-31, Ano95-75, Ano95-35, Ano95-34, Ano95-77, Ano95-32, Ano95-76, Ano95-33, Ano95-37, Ano95-38, Ano95-78, Ano95-39, Ano95-40, Ano95j, Ano95-79, Ano95-45, Ano95-41, Ano95-42, Ano95-44, Ano95-43, Ano95-52, Ano95-51, Ano95-47, Ano95-46, Ano95-80, Ano95-50, Ano95-48, Ano95-49, Ano95-53]. pages [Ano95-55, Ano95-56, Ano95-57, Ano95-58, Ano95-59, Ano95-60, Ano95-61, Ano95-63, Ano95-62, Ano95-65, Ano95-66, Ano95-64, Ano95-67, Ano95-81, Ano95-69, Ano95g, Ano95-82, Ano95-83, Ano95-54, Ano95-68, Ano96g, Ano96s, Ano96i, Ano96-150, Ano96e, Ano96h, Ano96k, Ano96f, Ano96t, Ano96j, Ano96l, Ano96w, Ano96y, Ano96u, Ano96x, Ano96v, Ano96-32, Ano96-37, Ano96-39, Ano96-27, Ano96-29, Ano96-42, Ano96-45, Ano96z, Ano96-44, Ano96-36, Ano96-151, Ano96-43, Ano96-33, Ano96-34, Ano96-31, Ano96-35, Ano96-28, Ano96-41, Ano96-153, Ano96-152, Ano96a, Ano96b, Ano96-38, Ano96-30, Ano96-40, Ano96-47, Ano96-46, Ano96-48, Ano96-50, Ano96-54, Ano96-55, Ano96-49, Ano96-53, Ano96-52, Ano96-51, Ano96c, Ano96-56, Ano96-69, Ano96-67, Ano96-65, Ano96-61, Ano96-57, Ano96-63, Ano96-66, Ano96-62, Ano96-60, Ano96-68, Ano96-71]. pages [Ano96-58, Ano96-72, Ano96-70, Ano96-59, Ano96-64, Ano96-74, Ano96-73, Ano96-75, Ano96-76, Ano96-78, Ano96-77, Ano96-86, Ano96-79, Ano96-87, Ano96-82, Ano96-81, Ano96-80, Ano96-88, Ano96-83, Ano96-85, Ano96p, Ano96-89, Ano96-91, Ano96q, Ano96-90, Ano96r, Ano96n, Ano96m, Ano96o, Ano96-92, Ano96-93, Ano96-95, Ano96-94, Ano96-96, Ano96-104, Ano96-100, Ano96-98, Ano96-99, Ano96-103, Ano96-154, Ano96-101, Ano96-97, Ano96-119, Ano96-109, Ano96-114, Ano96-105, Ano96-116, Ano96-113, Ano96-111, Ano96-118, Ano96-106, Ano96-107, Ano96-112, Ano96-156, Ano96-108, Ano96-117, Ano96-120, Ano96-115, Ano96-110, Ano96-155, Ano96-122, Ano96-125, Ano96-124, Ano96-157, Ano96-123, Ano96-121, Ano96-126, Ano96-158, Ano96-134, Ano96-128, Ano96-129, Ano96-131, Ano96-136, Ano96-132, Ano96-159, Ano96-127, Ano96-133, Ano96-130, Ano96-138, Ano96-140, Ano96-137, Ano96-139, Ano96-148, Ano96-144]. pages [Ano96-141, Ano96-142, Ano96-161, Ano96-146, Ano96-147, Ano96-149, Ano96-143, Ano96-164, Ano96-163, Ano96-165, Ano96-162, Ano96d, Ano96-167, Ano96-84, Ano96-135, Ano96-145, Ano96-160, Ano96-166, Ano97g, Ano97-140, Ano97f, Ano97-53, Ano97m, Ano97n, Ano97o, Ano97h, Ano97l, Ano97e, Ano97p, Ano97-142, Ano97q, Ano97r, Ano97-31, Ano97s, Ano97-28, Ano97w, Ano97-32, Ano97u, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97t, Ano97-30, Ano97v, Ano97-29, Ano97-40, Ano97-41, Ano97-33, Ano97-39, Ano97-34, Ano97-43, Ano97-37, Ano97-36, Ano97-38, Ano97-42, Ano97-35, Ano97-45, Ano97-44, Ano97-143, Ano97-46, Ano97-54, Ano97-48, Ano97-47, Ano97-51, Ano97-50, Ano97-49, Ano97-62, Ano97-57, Ano97-63, Ano97-60, Ano97-55, Ano97-59, Ano97-58, Ano97-56, Ano97-144, Ano97-64, Ano97-65, Ano97-72, Ano97i, Ano97-71, Ano97-70, Ano97-73, Ano97-68]. pages [Ano97k, Ano97-69, Ano97j, Ano97-66, Ano97-75, Ano97-77, Ano97-76, Ano97-74, Ano97-78, Ano97-145, Ano97-83, Ano97-81, Ano97-85, Ano97-84, Ano97-79, Ano97-82, Ano97-88, Ano97-86, Ano97-80, Ano97d, Ano97-146, Ano97-87, Ano97-149, Ano97-92, Ano97-95, Ano97-93, Ano97-148, Ano97-89, Ano97-96, Ano97-90, Ano97-147, Ano97-98, Ano97-99, Ano97-97, Ano97-100, Ano97-91, Ano97-94, Ano97-111, Ano97-101, Ano97-112, Ano97-102, Ano97-150, Ano97-106, Ano97-109, Ano97-103, Ano97-110, Ano97-108, Ano97-107, Ano97-105, Ano97-104, Ano97-113, Ano97-116, Ano97-115, Ano97-117, Ano97-122, Ano97-123, Ano97-119, Ano97-120, Ano97-124, Ano97-151, Ano97-118, Ano97-121, Ano97-125, Ano97-152, Ano97-126, Ano97-137, Ano97-135, Ano97-156, Ano97-158, Ano97-153, Ano97-155, Ano97-128, Ano97-138, Ano97-154, Ano97-131, Ano97-136, Ano97-134, Ano97-132, Ano97-133, Ano97-127, Ano97-157, Ano97-160, Ano97-139, Ano97-161]. pages [Ano97-164, Ano97-162, Ano97-163, Ano97-165, Ano97-159, Ano97-166, Ano97-169, Ano97-167, Ano97-168, Ano97-52, Ano97-129, Ano97-130, Ano98g, Ano98m, Ano98h, Ano98q, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98k, Ano98s, Ano98u, Ano98w, Ano98x, Ano98r, Ano98v, Ano98-34, Ano98-29, Ano98-28, Ano98-38, Ano98z, Ano98y, Ano98-27, Ano98-37, Ano98-33, Ano98-31, Ano98-32, Ano98-35, Ano98-36, Ano98-30, Ano98t, Ano98-39, Ano98-44, Ano98-42, Ano98-41, Ano98-40, Ano98-46, Ano98-45, Ano98-43, Ano98-52, Ano98-54, Ano98-51, Ano98-48, Ano98-49, Ano98-148, Ano98-53, Ano98-50, Ano98-47, Ano98-60, Ano98-61, Ano98-58, Ano98-59, Ano98-56, Ano98-55, Ano98-57, Ano98-64, Ano98-65, Ano98-66, Ano98-63, Ano98-68, Ano98-67, Ano98-69, Ano98-70, Ano98-71, Ano98-73, Ano98-72, Ano98-78, Ano98-77, Ano98-82, Ano98-79, Ano98-76, Ano98-84]. pages [Ano98-74, Ano98-75, Ano98-81, Ano98-83, Ano98-89, Ano98-88, Ano98-87, Ano98-86, Ano98-85, Ano98-93, Ano98-90, Ano98-95, Ano98-91, Ano98-92, Ano98-109, Ano98-102, Ano98-105, Ano98-99, Ano98-106, Ano98-97, Ano98-98, Ano98-108, Ano98-100, Ano98-101, Ano98-103, Ano98-110, Ano98-107, Ano98-104, Ano98-115, Ano98-113, Ano98-117, Ano98-149, Ano98-111, Ano98-118, Ano98-112, Ano98-114, Ano98-116, Ano98-121, Ano98-120, Ano98-119, Ano98-123, Ano98-126, Ano98-127, Ano98-128, Ano98-129, Ano98-150, Ano98-122, Ano98-130, Ano98-133, Ano98-137, Ano98-136, Ano98-131, Ano98-132, Ano98-135, Ano98-134, Ano98-151, Ano98-141, Ano98-142, Ano98-144, Ano98-138, Ano98-143, Ano98-139, Ano98-140, Ano98-153, Ano98-145, Ano98-157, Ano98-146, Ano98-155, Ano98-152, Ano98-147, Ano98-154, Ano98-156, Ano98-159, Ano98-158, Ano98-160, Ano98-161, Ano98-163, Ano98-164, Ano98-62, Ano99-29, Ano99z, Ano99v, Ano99-30, Ano99-27]. pages [Ano99w, Ano99-90, Ano99o, Ano99r, Ano99s, Ano99n, Ano99t, Ano99y, Ano99x, Ano99-31, Ano99u, Ano99-28, Ano99-32, Ano99q, Ano99m, Ano99p, Ano99-35, Ano99-36, Ano99-33, Ano99-34, Ano99-37, Ano99-41, Ano99-47, Ano99-42, Ano99-48, Ano99-50, Ano99-213, Ano99-44, Ano99-51, Ano99-49, Ano99e, Ano99-39, Ano99-40, Ano99-43, Ano99-46, Ano99-45, Ano99-38, Ano99-63, Ano99-66, Ano99-53, Ano99-55, Ano99-64, Ano99-67, Ano99-54, Ano99-56, Ano99-69, Ano99-52, Ano99-65, Ano99-60, Ano99-57, Ano99-61, Ano99-62, Ano99-59, Ano99-58, Ano99-68, Ano99-75, Ano99-78, Ano99-76, Ano99-79, Ano99-73, Ano99-71, Ano99-77, Ano99-72, Ano99-70, Ano99-84, Ano99-85, Ano99-86, Ano99-92, Ano99-83, Ano99-81, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-214, Ano99-89, Ano99-82, Ano99-87, Ano99-105, Ano99-103, Ano99-99, Ano99-106, Ano99-96, Ano99-109]. pages [Ano99-102, Ano99-104, Ano99-100, Ano99-107, Ano99-97, Ano99-110, Ano99-95, Ano99-98, Ano99-101, Ano99-108, Ano99-115, Ano99f, Ano99-111, Ano99-118, Ano99h, Ano99-116, Ano99g, Ano99-112, Ano99-119, Ano99i, Ano99-113, Ano99-117, Ano99-120, Ano99-114, Ano99-216, Ano99-123, Ano99-125, Ano99-129, Ano99-133, Ano99-217, Ano99-124, Ano99-126, Ano99-130, Ano99-134, Ano99-215, Ano99-135, Ano99-128, Ano99-121, Ano99-136, Ano99-137, Ano99-132, Ano99-127, Ano99-131, Ano99-122, Ano99-139, Ano99-138, Ano99-140, Ano99-145, Ano99-143, Ano99-141, Ano99-144, Ano99-142, Ano99-158, Ano99-160, Ano99-218, Ano99-156, Ano99-154, Ano99-159, Ano99-161, Ano99-219, Ano99-157, Ano99-155, Ano99-152, Ano99j, Ano99-151, Ano99-163, Ano99-149, Ano99-146, Ano99-162, Ano99-153, Ano99-164, Ano99-148, Ano99-150, Ano99-147, Ano99-167, Ano99-222, Ano99-168, Ano99-223, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-166, Ano99-220]. pages [Ano99-170, Ano99-173, Ano99-174, Ano99-176, Ano99-177, Ano99-178, Ano99-179, Ano99-183, Ano99-193, Ano99-190, Ano99-184, Ano99-194, Ano99-191, Ano99-180, Ano99-185, Ano99k, Ano99-181, Ano99-195, Ano99-188, Ano99-224, Ano99-192, Ano99-186, Ano99-182, Ano99-189, Ano99-196, Ano99-199, Ano99-200, Ano99-197, Ano99-201, Ano99l, Ano99-198, Ano99-206, Ano99-207, Ano99-204, Ano99-225, Ano99-203, Ano99-209, Ano99-202, Ano99-205, Ano99-226, Ano99-208, Ano99-229, Ano99-212, Ano99-230, Ano99-210, Ano99-211, Ano99-231, Ano99-228, Ano99-233, Ano99-232, Ano99-227, Ano99-234, Ano99-235, Ano99-236, Ano99-91, Ano99-238, Ano99-237, Ano99-239, Ano95-49, Ano97-61, Ano97-114, Ano99-221]. Painlevé [TG96]. paint [SH96]. pair [BS96a, CMR90, Tsi99, YJ94]. paired [GHWL96]. pairs [Ano96-170, Bog96, Ech96, Yan94]. pairwise [KO97, KO99]. Palm [Ano99-180]. Palmer [Ano96-127]. PalmPilot [Ano98-122]. Pan [Ano96-129]. Pandey [Ano99-87]. panel [Lai97]. Pankaj [Ano96-33]. Pannell [Ano97-72]. Panos [Ano98-108, Ano98-123]. Pap [Ano98-126]. Papageorgiou [Ano98-67]. Paper [Ano90u, Ano95d, Ano95y, Ano95x, Ano95-29, Ano95-69, Ano96-50, Ano96-60, Ano96-64, Ano96r, Ano96-108, Ano96-130, Ano96-144, Ano96-141, Ano96-161, Ano96-143, Ano97-39, Ano97-59, Ano97-82, Ano97-95, Ano97-98, Ano97-99, Ano97-97, Ano97-117, Ano97-151, Ano97-135, Ano97-128, Ano97-154, Ano97-127, Ano97-130, Ano99d, Ano99h, Ano99i, Ano99-135, Ano99-140, Ano99-224, Ano99-225, Ano99-233]. paperback [Ano97n, Ano97-90, Ano97-94, Ano97-162, Ano97-159, Ano97-129, Ano98k, Ano98-70, Ano98-73, Ano98-128, Ano98-145, Ano98-146, Ano99-154, Ano99-155, Ano98-27]. papers [Ano95-65, Ano97-51, Ano97-129, Ano97-130, OR97b, OR97a, Zel92, Ano96-134]. parabolic [AAII92, Bal99, BS96b, BS96c, Bog98, Bog99, Boy96, CS96, CV90, CS90, DNJS98, Doh90, Far91a, Far91b, Far96a, GW91, HM93a, Kan90b, KHI94, Lar90, LD93, LD95, SWY96a, Sir99, Yan97, Zha98]. paradigm [Ano96z, Bec96, CDH91, LY99a]. paradigms [Ano99-66, Ano99-67]. paradox [Ano96-76]. Paragon [Ano97-120, LSH98, OOWD98, CCC^{+}98]. Paragon^{T M} [Pfi98]. Parallel [AAII92, Ano95-59, Ano97-41, Ano99-163, Bai96f, BS99a, BG95, Dav96, Dea98, DEMS94, EK92, FL91, Gal99a, Hen97, HIIN99, HC98, LH94, ND95, RG92, SZM98, SSN94, SMTS97, VK95, Wan99, WF99a, Amo93, APP96, AL94, Ano93m, Ano95-47, Ano95-82, Ano96-86, Ano96-132, Ano97t, Ano97-46, Ano97-160, Ano98-45, Ano98-56, Ano98-55, Ano99x, Ano99-133, Ano99-134, Bai96b, BW96a, Bai96c, Bai96e, BMD92, Bog96, CCCY94, CK97a, CLY91, CTY95, Con98, CSW99, Cor95, DC94, DKSI94, DF98, DHW98, FS96, GS99, GG90, HL96, Hen92, JS95, JS97a, JS98, JS97b, JLDL92, KG98, Koç92, LJ98, LC97b, LHC96, LVV96, May99, MNB93, MM95d, NNB97, Pan94, Pat98, Rei93a, Res91, SM98, SK91, Sir99, SCBFC93, SCBFC98, SJS92, TG95, Tsa90, VA97, WSDT94, YI97, Zha93c, dSB96, hC96a, hC96b, vDHO^{+}95]. parallel [vdHS94, vdV95, Ano97-115, Ano98-123, Ano98-124]. Parallelism [CL90]. Parallelization [LSH98, SCB^{+}93]. parallelized [Bon97]. Parallelizing [CDLW95]. Parameter [CKK97b, ST90, Ame97a, Ame97b, CF91a, Gla99, KIAM98, Nar92, ZA95]. parameter-dependent [Ame97b]. parameters [AKS96, CL94a, Faz94b, KNU99, KC92a, Lee96, LL90a, Li98b, MM95c, Pap94a, Pap92, Sou90, ZM98]. parametric [CST97, EB93, RS95, Ano99-181]. Parametrization [Pan92a]. parasite [MP99]. Parberry [Ano95q]. Pardalos [Ano98-108, Ano98-123]. Pareto [BJPT93, YY98]. Park [Ano95-47, Ano96h, Ano96-44, Ano96-36, Ano96-53, Ano96-135, Ano97r, Ano98-77, Ano98-112]. Parker [Ano95-67, Ano97-79]. Parks [Ano99-92, Ano99-192]. Part [Ano97-65, Ano99-113, YFF93, AR92d, AO93a, Ano96-83, Ano96-127, Dem94b, Gro95, MNE92, Zha91, AO93b, Dem94a, DIR97, Dor98a, Dor98b, Dor99, Gla96b, Gla97b, Has93b, Has95, HM93a, Hin95, IB95, LM96a, YFF92b, YFF92a]. part-whole [MNE92]. Partial [Ker91b, AR90a, Ado91d, AR92a, Ado96, Ano96t, Ano96-41, Ano96-91, Ano96-113, Ano97-41, Ano98-153, Ano99-34, Ano99-43, Ano99-197, CC97, CS96, CP96, CL98, CL99d, CS90, GPZ93, GF90, GPK99, Gyö90a, Has93c, JP95b, JA96, JD97, JLDL92, Kan90a, Kan90b, Lee93, LR93a, LC96a, NPJ95, NFJ96, PR90, SA94, SWY96b, WA96b, Won98a, ZY98, ZL98a, ZL99a, ZL99b, ZT99, ZQ93b, Ano97-116, Ano98-126, Ano98-125, Ano99-183, Ano99-184, Ano99-185, Ano99-182]. Partiality [Ano97-117]. partially [HH98b, Kra92]. Particle [Kel91, Kor97, KG97, Mah90, SV91a, SV91b, Gre92a, HPH^{+}94, IA94, Kro99, Mar91, MP94, RW98, WT97]. particle-fluid [MP94]. particles [Dor98a, Dor98b, Dor99, Ano99-186]. particular [CCD99, Seb91]. partition [ND95]. partitioning [BS91b, SCB^{+}93, SBD99]. parts [Cha94]. Pascal [Ano95-45, Ano95-57, Ano96-109, Ano96-134, Ano97-110, CG95, Fje91a, Fje91b, Fje91c, Ano93o]. Pascal-XSC [Ano93o]. Pasek [Ano96-158]. pass [HH98a]. passing [Ano95-82, Hen92]. passive [KKK91, KKK93]. Password [Lia95]. passwords [CH93]. past [CH91, Kha90]. patch [BM91, ZC96]. patches [BC97]. path [BS99b, DC94, GMW96, LOP95, MCG98c, Shi95a, Tho92, Zha93b]. path-based [Tho92]. paths [Che99a, Reg94]. Patricia [Ano93t, Ano95-78, Ano97e, Ano97-149]. Patrick [Ano95f, Ano96-70, Ano96-116, Ano97-59, Ano97-117, Ano98-32, Ano98-54, Ano99-108, Ano99-229, Ano99-235, Ano99-236]. Patriksson [Ano99-170]. Pattern [Lev92, Mai96b, MDVK97, YTLK94, Ano97-118]. patterns [Ano97-43, Ano98-150, Ano99-231, Ram95, Ano98f]. Pattie [Ano95k, Ano97-53]. Paul [Ano96-61, Ano96-62, Ano96q, Ano97j, Ano97-81, Ano98-98, Ano98-114, Ano98-141, Ano98-139, Ano99f, Ano99g, Ano99-221, Ano99-210, Ara99b]. Paulk [Ano96-151]. Paulo [Ano96-157]. Paulson [Ano97-95]. Paun [Ano98-100, Ano99-66, Ano99-67]. Pavel [Ano99-49]. Pavlik [Ano98-161]. Pawel [Ano97-87]. PC [Con98, Ano90s]. PCG [KG95b]. PDE [Chu92, GZ93]. PDEs [Ano99-181, AAII92, Mur91b, CV90, HL96, LZ98, MC99, VK95, Yan93a, Ano97-165]. Peano [FLR95, Kou95, Nas92]. Pearce [Ano97-169]. Pearson [Ano99-136]. Peck [Ano96-70, Ano97q, Ano98-54, Ano98-162]. Pedoe [Ano96-29]. Pedrycz [Ano99-52]. Peek [Ano95-46]. Peircean [Bur92]. Peitgen [Ano99-89]. Pellizzari [Ano99-166]. penalties [Cha99a]. Penalty [SFY92, WT90b, ZZ99d]. penetration [LD95]. Peng [Ano98-144]. Pennington [Ano95-73]. Penny [Ano96-125]. people [Ano97-148]. Pera [Ano95-74]. perception [Ano99-163]. perceptron [BIZ97]. Perciante [Ano99-89]. Pérez. [Ano93v]. perfect [BV95, Ngu97, ND95, SS96, SS99, Ano97-119]. perfectly [MMZ91]. Performability [Nab98]. Performance [GK98, Kat92, SZM98, AA93, AS99, Ano96b, Ano96-86, Ano96o, Ano98-139, Ano99-120, Ano99-235, Ano99-236, Ben99, EA92, HTT92, JLDL92, KS99, Leo96, Li98c, PC96b, Pen96, SK91, Sco96, WSS92, YFF92b, YFF92a, YFF93, YF93, vDHO^{+}95, Ano99-163, Ano96-127]. Performance-based [Ano96-127]. performing [Opp96, RR90]. Périaux [Ano98-64, Ano99-74]. Perimeter [Man96, Bri97]. Period [Yan93c, KW98]. Periodic [LD93, LD95, AHL98, AS96a, Dor90, HH98a, KP97a, LL91b, Mag97, Ran93, RT99b, Sim93a, Ter00, Vic90, WCC99, XC92, YCC96, ZM94, ZC99]. periodical [ABV97, Fau94]. periodicity [Den91a]. Peristaltic [MP94]. perl [Ano97-89, Ano99-238, Ano99-141, Ano97-166, Ano99-148]. Perl/Tk [Ano99-141]. Permanence [YCC96, ZC96]. permeable [Kha90]. permutation [AR99b, LKC95, Zho90]. permutation-generating [Zho90]. permutations [Arn90, Bha94]. Perot [KKG97]. Perron [RJ94, RS96b]. Perseus [Ano99-81]. Persistence [SC98a, CYS94, TC99]. person [Bel94a, Bel97a, Nak99]. Personal [SL99, Ano96-122, Ano99-216, Ano99-217, Las96]. perspective [Ano95p, Ano96z, Ano97u, Ano99-80, Ano99-233]. perspectives [Ano95x, Ano96-53, Ano96-32, Ano99-118, Ano99-119]. persymmetric [RS96b]. PERT [SS91a]. perturbation [JR94, MT92, NO93, WK99]. perturbations [CYS94]. perturbed [BS93, BS96b, BS96c, Bog99, Duc95a, EDO96, EY91, GPCGV92, HG98, Hua98a, HB99, Li98b, LN98, LN99, Li10, MP91, MK93, Sir97, Sir99, Tha94]. petaflops [Ano96-62]. Peter [Ano95v, Ano96-60, Ano96-85, Ano96d, Ano97-140, Ano97f, Ano97-42, Ano97-85, Ano97-154, Ano98-53, Ano98-109, Ano98-108, Ano99n, Ano99y, Ano99p, Ano99-34, Ano99-154, Ano99-155, Ano99-163, Ano99-231, Ano98-113]. Peters [Ano95j, Ano95-51, Ano95-49, Ano95-55, Ano99-231]. Peterson [Ano97-81]. Petrenko [Ano99-212]. Petri [Ben99, KS99, KKOS96, MR99, Nis97b, Nke94, SJS92, ZK99]. Petrov [Bla95]. Petru [Ano98-144]. Petsche [Ano96-39, Ano98h]. Petter [Ano97-41, Ano99-43]. Pfeifer [Ano99-90]. PGP [Ano95-58]. phantasies [dJPvZ97]. Phase [Zha93b, CLY91, Hun95, Mur92b, PS93a, Sch90, Sim93b, Sim93c]. phase-lag [Sim93b, Sim93c]. phase-state [Sch90]. phenomena [ADFP93, BKK97, FWI99, Gre93b, KKG97, Mah90, NIS97a, Shi97, Ano97-162]. phenomenon [GG97, KYO92]. Philip [Ano95-36, Ano96-153, Ano96-40, Ano97-40, Ano99-87, Ano99-227]. Philippe [Ano95-56]. Phillips [Ano95y, Pok97]. philosophical [Ano96-141]. philosophy [Ano97-31, Ano97d, Ano96-128, Ano97-120]. phoenetics [Ano96-60]. phosphofructokinase [BJH90]. photodimerization [MPR97]. Photoshop [Ano99-187]. phychophysical [Ano99-118, Ano99-119]. physical [Ado94]. physicist [Ano96-63]. physicists [Ano97-125]. physics [Ano96-57, Ano97-92, Ano99-149, Ano99-146, Ano99-162, Ano99-225, Kan92a, Mar91, Ano96-144, Ano96-146, Ano97-156]. physiology [XS90]. Piatetsky [Ano96h]. Piatetsky-Shapiro [Ano96h]. pickup [RR97b]. pieces [Ano97-83]. Piecewise [Vos91, BF91, Bro94, CG94, EDNO92, LS96a, LS98b]. Pierangela [Ano96-48]. Pierre [Ano98u, Ano99-68, Ano99-95, Ano99-201]. Pierre-Louis [Ano99-68]. Piet [Ano96v]. Piet-Lahanier [Ano96v]. Pieter [Ano99-83]. Pilch [Ano97-140]. Pillo [Ano96-123]. pin [MBMS92]. pipe [Ano99-242, MATA98]. piping [YT98]. pirimidine [MPR97]. pitch [LY92]. Pittsburgh [Ano95-27]. pivoting [LC97b]. PL [Ano99-178, Ano99-179]. PL/SQL [Ano99-178, Ano99-179]. Place [Ano96-56, Ano96-30]. placement [GS96b, NNB97]. places [Ano97-148]. planar [Ano99-246, BS99b, CO99, Gue93, Kai94, Kai96, MMMS97, Pav97]. plane [ABV97, Ano99-198, DPD98, KKS98, Kle93, MB92, Pas92, PB95, RSV98, WF99b, YC92, YLC96]. planes [Pok97]. planning [RT94, WF96]. plans [Cha94]. plant [YT98]. plasmas [BKK97]. Plasmodium [BJH90]. plasticity [vdVVdB99a]. Plate [Chi93, Sal92a, Tad93, TH95, WF90]. plates [FCP95, MB94b]. platform [Pfi98]. Platforms [Ano97-105]. players [GP93]. Playing [Ano95-37, Ano98-63]. Plenum [Ano95p, Ano95-83, Ano96v, Ano96-113, Ano96-123]. Pless [Ano98-81]. plus [Ano97-141, EG97, EG98]. PNG [Ano99-188]. PNTBLSIM [Lew94]. Pöchmüller. [Ano95-53]. pocket [Ano99-178]. Podolsky [Ano90s]. Pogue [Ano98-122]. Poincaré [Tre94, Shu93a]. Poincarè-Cartan [Shu93a]. Point [PPH98, Sie90, AW91, AO98b, Ano98-59, Ano98-78, Ano99z, Ano99-27, BACS98, BCC95, CL91a, CL91b, CRV98, DBM92, DMC^{+}95, Eis92, FG98, GS93, Gol94, GB90, HK96, Hua97, JP93, Kai94, Kai96, KG98, Köh95, Köh96a, Köh96b, Köh99, KCB90, Kwo92, LKE97, Lon95, MN91, MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b, Més96, MM93c, MM94, NN91, O’R95, RSC93, RSS94a, RT96, Rok90, SWF95, Sou90, VL92, YFF93, YY98]. point-mass [YFF93]. points [Ano99-244, Bai96a, BB99, CST97, GK91, HW93a, Hua98c, KO96, Köh95, Köh96a, LK94, MN91, NB98, O’R98b]. Pointwise [DG91, Li98a]. Poiseuille [Kha90]. Poisson [McL95, ZQ95, LJ98, LQ95a, LQ95b, LHC96, LVV96, NJ92, ZQ94]. polar [Wad95]. poles [PN98]. policies [DO95, IN92, KO90, KK94, Pap92, Saw92, YNK92]. Policy [Ohn92, Ano96-148, Ano97-49, Ano98-82, Dei91, Erm92, KK92, SKI92, YF92b, Ano97-100, Ano99-38]. policy-making [Dei91, Erm92]. political [Ano95-27, Ano99-38]. politics [Ano99-224]. Pollack [Ano97-53]. Pollock [Ano96-128]. pollution [NWL91]. Poloni [Ano98-64]. Pólos [Ano97n]. polygon [GMBS94]. Polygonal [Köh96c, Mol93, Cha96, Köh93a, Köh95, Köh96a, Köh96b, WW91]. polygons [Köh96c, Ola92]. polyhedra [Bro90]. polyhedral [Bel94a, Bel97b, Bel97a]. polyhedron [KKS98]. Polynomial [MV95b, RHRZ90, Sad96b, SRCB98, Wac91b, AL93, Ano99-167, Ano99-168, Boy99, FL91, GM99, Huj96, JS95, JPC99, KO91, KK91, MV95a, Ngu97, PSLT93, Pan95, Pan96, Pan98c, PC91, PC93, Rei99a, Ter91, Ano96-129]. polynomials [AR92b, BV99, Bar91a, BK98, Bis91, Car94, DCL^{+}94, Den91b, Doh90, Doh91, Gem93, GNR95, GW97, He95a, He95b, IN97, KS98a, KTO97, KM97c, LOS99, LW99b, Lub97, MZ97, MP96, MO95, NR96a, NP96, Ric95, Riv97, Róz99, SSH92, SG95b, VL95, WK94, ZL99c, Ano99-189]. polynucleotides [MPR97]. polytope [LM96b]. Poorten [Ano96-124]. Poovey [Ano99h, Ano99i]. Popov [Ano99-98]. popular [Ano96-141]. population [Bar91b, CC98c, CC99, Gyö90b, KP95a, KP95b, LZ98, Mil90, Ano96-130]. populations [Ahm95, Kos90, KC91, Mic90]. porous [Bla95, CH91, Eti91, KG94, KP99a, KMD99, PSD96, Pas92, WHC90]. Portable [Ano95-82, KO93]. Porter [Ano97-67]. portfolio [Nai92]. posed [DEMS94]. Positive [AWY97, AH98, Ber96, GGK96, Hu95, WL96, AHY98, AP98b, LC98a, LW99b, LURL98, Mye92, Nis95, TC99, Tra97, WJ99, WA96a, Won98a, Won98b, YCC96, Zaf98a, ZT98a]. positron [BKK97]. possibilities [Ano96-127]. Possibility [NL93]. possible [PS93a]. possibly [Maz90]. Post [FK98a]. Postel [Ano96-120]. posteriori [AO93b, AO93a, EL96, Kor96]. postmodern [Ano96-52]. postprocess [Joh99]. Postulate [Mor93]. potential [Ano97-144, AI97, Ara95, Eti91, Har92b, KO96, RF91, Sch90, YL99]. Potentials [McK96, BSP94]. Pountney [Ano96-98]. power [Ano97h, BG91b, Gil93, HGG99, KM90, Pas92, YT98, ZQ93c]. Powerful [Ano97-89, Sug94]. Powers [Ano99-61, Ism97]. Pozrikidis [Ano98-118]. pp [Ado91a, MM94, Sal92b, Zad92]. pp. [Ano90r, Ano90t, Ano90v]. Practical [Ano97-159, Ano99-154, Ano99-155, Ano99-151, Ano93k, Ano95-55, Ano96a, Ano96-159, Ano97-72, Ano97-151, Ano99-45, Ano99-246, CO99, Cha94, SAS92, Ano96-131, Ano96-132, Ano96-133]. practice [Ano96-134, Ano96-144, Ano97-122, Cho96, Ano95u, Ano95z, Ano98-69]. practitioners [Ano97o, Ano98p]. pragacz [dJF91]. pragmatics [Ano97-80]. Prague [Ano99-128]. Pramila [Ano98-59]. precedence [Ste90]. Precise [Feu90]. Precision [KKK93, KO94]. precomposition [SH96]. Preconditioned [Gol90, MRWM94, PS92, Ara96, Ara99a, GL96, PZDH95, Rac99]. preconditioner [DF94, GK98, Luo93b]. preconditioners [Ara96]. Preconditioning [vdHS94, DFS93, SB99]. precursor [LNM90]. predator [SC98a, WC97].

Predictable [Ano98-68]. predicted [HSH90b]. Prediction [Gla92c, Gla96b, Gla97b, RA92a, SSN94]. predictions [Jek94]. predictive [SI94]. Predictor [GP91, VA97, VK99]. predictors [Con98]. Preface [Aga94, Aga98, Ana95b, Ano90-31, Ano91m, Ano92i, Ano92j, Ano93-30, Ano93-31, AKP97a, AKP97b, AKP97c, ABF90, AS91, CI98, Eso99d, Gal91c, Gal99d, Kan92b, MU96, Mit92b, OR97a, Ram91d, Róz96, SGS99, Wit96a, Wos95b, Yas98a, OR97b]. preference [Dei93, Sla92, WF92]. Prefix [EK92, LHC96, Rei93a]. Preinvex [Sya99]. prejudices [Los89b]. Prentice [Ano90p, Ano90q, Ano90s, Ano90x, Ano93k, Ano93n, Ano96-71, Ano96-98, Ano96-102, Ano97-113, Ano97-114, Ano97-116, Ano99-185]. preparation [Ano95f, Ano95-42]. preprint [HP99]. prescribed [BS99b, GSM91, Moo91, Moo93b, Moo94, SMT98]. presence [FBL99, LL99, Med90, PCR90, RR90]. presentation [RS99b, SHA99b]. Preserving [QZ93a, CP95a, Fab95, KM97c, LC96b]. Press [Ano90r, Ano90t, Ano90w, Ano90v, Ano93f, Ano93g, Ano93t, Ano93u, Ano93i, Ano93j, Ano93l, Ano93v, Ano93r, Ano93s, Ano93w, Ano95k, Ano95-36, Ano95h, Ano95-70, Ano95n, Ano95o, Ano95r, Ano95s, Ano95d, Ano95q, Ano95p, Ano95t, Ano95v, Ano95y, Ano95x, Ano95w, Ano95-71, Ano95z, Ano95-28, Ano95-29, Ano95-74, Ano95-73, Ano95-30, Ano95i, Ano95-31, Ano95-75, Ano95-77, Ano95-32, Ano95-37, Ano95-38, Ano95-78, Ano95-40, Ano95-45, Ano95-41, Ano95-43, Ano95-47, Ano95-59, Ano95-60, Ano95-62, Ano95-66, Ano95-81, Ano95-69, Ano95g, Ano95-82, Ano95-83, Ano95-54, Ano95-68, Ano96s, Ano96i, Ano96h, Ano96k, Ano96t, Ano96l, Ano96x, Ano96v, Ano96-32, Ano96-39, Ano96-44, Ano96-36, Ano96-43, Ano96-34, Ano96-31, Ano96-35, Ano96-153, Ano96-152, Ano96a, Ano96b, Ano96-30, Ano96-40]. Press [Ano96-47, Ano96-46, Ano96-50, Ano96-55, Ano96-53, Ano96-52, Ano96c, Ano96-69, Ano96-67, Ano96-61, Ano96-57, Ano96-63, Ano96-66, Ano96-62, Ano96-60, Ano96-58, Ano96-72, Ano96-64, Ano96-74, Ano96-73, Ano96-78, Ano96-87, Ano96-81, Ano96-80, Ano96-85, Ano96p, Ano96-89, Ano96-90, Ano96r, Ano96n, Ano96o, Ano96-95, Ano96-94, Ano96-96, Ano96-104, Ano96-100, Ano96-99, Ano96-103, Ano96-154, Ano96-97, Ano96-114, Ano96-105, Ano96-113, Ano96-118, Ano96-108, Ano96-115, Ano96-155, Ano96-122, Ano96-123, Ano96-121, Ano96-158, Ano96-134, Ano96-128, Ano96-136, Ano96-132, Ano96-127, Ano96-130, Ano96-138, Ano96-140, Ano96-137, Ano96-139, Ano96-148, Ano96-144, Ano96-141, Ano96-142, Ano96-161, Ano96-149, Ano96-143, Ano96-163, Ano96-162, Ano96-84, Ano96-135, Ano96-145, Ano96-160, Ano97g, Ano97f, Ano97-53, Ano97h, Ano97l, Ano97p, Ano97-142, Ano97r, Ano97-31, Ano97s]. Press [Ano97-28, Ano97-32, Ano97u, Ano97x, Ano97y, Ano97z, Ano97v, Ano97-40, Ano97-41, Ano97-39, Ano97-43, Ano97-37, Ano97-42, Ano97-44, Ano97-143, Ano97-48, Ano97-47, Ano97-50, Ano97-49, Ano97-57, Ano97-63, Ano97-55, Ano97-59, Ano97-144, Ano97-64, Ano97i, Ano97k, Ano97-69, Ano97-83, Ano97-81, Ano97-82, Ano97-88, Ano97-80, Ano97d, Ano97-146, Ano97-149, Ano97-95, Ano97-148, Ano97-90, Ano97-98, Ano97-99, Ano97-97, Ano97-100, Ano97-91, Ano97-94, Ano97-150, Ano97-109, Ano97-110, Ano97-115, Ano97-151, Ano97-118, Ano97-137, Ano97-135, Ano97-156, Ano97-158, Ano97-153, Ano97-155, Ano97-128, Ano97-154, Ano97-134, Ano97-133, Ano97-127, Ano97-160, Ano97-139, Ano97-161, Ano97-159, Ano97-52, Ano97-129, Ano97-130, Ano98h, Ano98k, Ano98s, Ano98u, Ano98v, Ano98-34, Ano98-28, Ano98-38, Ano98y, Ano98-27, Ano98-35, Ano98-39, Ano98-41, Ano98-40, Ano98-43]. Press [Ano98-48, Ano98-148, Ano98-60, Ano98-61, Ano98-58, Ano98-55, Ano98-63, Ano98-70, Ano98-73, Ano98-77, Ano98-82, Ano98-87, Ano98-96, Ano98-94, Ano98f, Ano98-95, Ano98-106, Ano98-107, Ano98-104, Ano98-113, Ano98-149, Ano98-118, Ano98-112, Ano98-114, Ano98-119, Ano98-124, Ano98-128, Ano98-150, Ano98-136, Ano98-131, Ano98-134, Ano98-151, Ano98-145, Ano98-146, Ano98-152, Ano98-147, Ano99-29, Ano99-30, Ano99-90, Ano99-32, Ano99q, Ano99p, Ano99-51, Ano99-39, Ano99-40, Ano99d, Ano99-38, Ano99-68, Ano99-75, Ano99-76, Ano99-92, Ano99-214, Ano99-106, Ano99-107, Ano99-108, Ano99f, Ano99-118, Ano99h, Ano99g, Ano99-119, Ano99i, Ano99-216, Ano99-125, Ano99-217, Ano99-126, Ano99-215, Ano99-135, Ano99-122, Ano99-140, Ano99-144, Ano99-158, Ano99-160, Ano99-154, Ano99-159, Ano99-161, Ano99-155, Ano99-163, Ano99-146, Ano99-150, Ano99-222, Ano99-223, Ano99-220, Ano99k, Ano99-224]. Press [Ano99-225, Ano99-203, Ano99-202, Ano99-205, Ano99-226, Ano99-228, Ano99-233, Ano99-232, Ano99-91, Ano99-239]. Press/AEI [Ano97-139, Ano97-161]. Press/AIE [Ano97-37]. Press/Elsevier [Ano96-31]. Press/McGraw [Ano97-137]. Press/McGraw-Hill [Ano97-137]. Press/MIT [Ano95-47, Ano96h, Ano97r, Ano98-112, Ano99-40]. Press/Stanford [Ano97-129]. pressure [HW96]. prestarlike [RS99b, SA95]. presupposition [Ano95-30]. Pretty [Ano95-58]. Pretzel [Ano98y]. Preventing [Ano96-137]. prey [SC98a, WC97]. Pribyl [Ano99-178, Ano99-179]. Price [VA94, BTED97, Nai92]. primal [SWF95]. primal-dual [SWF95]. Primality [Yan95]. Prime [KS91, Ano96-157, Tan92]. primer [Ano95g]. primes [Mor93, Ano97-121]. primitive [Cla98, EZ93, FMS92, Gla99]. Princeton [Ano95t, Ano96-136, Ano96-137, Ano99-215]. principal [Los89b, Mil91]. principle [ABK97, DM95, KNUX99, Kis91, Ano98-59]. Principles [Ano98-69, Ano99-47, Ano99-48, AL93, Ano97-126, MO98b, MO98a, WSNS98, WCHL96, Ano96-134, Ano97-122]. priori [JD97, JPC99, NFJ96]. Priority [Mis90]. Prisco [Ano98-140]. Pritsker [Ano97-132]. privacy [Ano95-58]. Prize [Rod94f]. Probabilistic [ABB97, LOS99, Rei93a, Ano96-38, Bou93, MCG98b, RDK92, Ano98-127, Ano98-128]. probabilities [PT91, Zad99]. probability [AR99a, Ano96-71, Ano98-75, FZ93, KMR90, MM96b, Mar95, RZ96, Ano99-190, Ano99-191]. Problem [Ano95u, Ano96-86, Gal91b, AO99f, ABV97, Ano92a, Ano96-110, Ano97-73, Ano97-134, Ano98-80, Ano00, AS98, AP98b, BT95, BSS98, BKK99, Bec90, Bed95a, BF91, BA98, BS91b, BS95, BS96c, Bog98, Bog99, BC98, BL99, BT97, Cas98, Che92, Che90b, CC98a, CKK97b, CY93, CY94, Cub96, DBM92, DBM94, DKR97, DKW98, EY91, Eso99b, EGLN95, FO91, FK98b, GGK97, GS95, Gla90b, GM96, GHS91, GMR90, Gur91, HS95, HM99, HM93b, Hua91, Hun95, HW90, Ike91, Ioa91, IM93, JW93, Jan96, JTY97, JKK98, KM92, KO96, Kha97, KESH90, Kle93, Kra90, Kre91, KS97b, KM90, Li96a, Li97, Li98a, Li98b, LN99, LM96a, Mac93, MCG98b, MMZ91, May99, MG90c, MH93b, Nak99, Nan90, PSS99, PTC^{+}98, PLH95, PL99, Ram91e, Rèp91, Riz95, RS92, RSV98, RR97b, SS91a]. problem [Seb91, SOI92, Sod90, Ste90, SV97, Sym91, TI96, TWIF98, TTC96, TE93, UW91, UWT99, US91, Wan93a, Wan98b, WJ99, YP98a, YP98b, Yav91b, Yav91a, Yos96, ZS93, Zed99, ZG97, ZCC96, Ano97-141, Ano99-181]. Problem-method [Gal91b]. problem-solving [Ano96-110]. Problems [Ano99h, Ano99i, Car93, DELE99, PW97, RD99, Ado91d, Ado94, AL93, AO97, AHY98, AO98b, AH98, ABW99, AK90a, AO98c, AC93a, AS99, Ame98, AR92e, AR98, Ano90-28, Ano96-164, Ano96-145, Ano99-70, Ano99-84, Ano99-85, Ano99-171, Ano99-170, Ano99-198, AKV91, AS96a, BMR96, BBO99, Bai96b, BW96a, Bai96c, Bai96d, BACS98, BF95, Bed95b, BK91, BR95c, BKM91, BMD92, BS93, BS96b, Boh98, Bon97, BS96d, Bou98, BP97, But94, CGH92, Car92, CS93a, CCCY94, CF98, CLD97, Chi91, CLM99, CK91, Cor95, CK94a, CK94b, Cub95b, DG99, DNJS98, Dem94b, Din99c, DEMS94, Ech95, Ech98, EB93, EDNO92, EGEHSEK95, ESKG98, EK94, ER98, Eti91, FW96, Far96a, FL96, FS96, Fog94, FWI99, GOWR^{+}93, Gau97, Gol90, Gol94, GWL96, Haw91, HK98, HP99]. problems [HM93a, Hin95, HMZ99, Hsu91, HM94, Huj99, Ill91, IA94, Ita93, JR94, JG95, JP95b, JA96, JCM96, JG96, JD97, JA97, JG98, JPC99, KNU99, KK97a, KG98, KT95, KC91, KK91, KOT99, KPS96, LTC98, LTR98, LH91, LW96, LN98, Li10, Liu96a, LT99, LNG92, MS96, MN91, MM93a, MT92, MB94a, MRWM94, Med90, MM93b, Mey99, MJK92, Mis90, Moi97, Mol96, Mur92a, Mur92b, PA94, Par91a, PR96, Pom91, Pom92, PN99, QA98, RSS94a, Ram90a, Ram91a, RF91, RT96, SS98a, SNU98, SS94, SM90, She92, She93, She99, SSZ91, Shi95a, ST90, Sim93b, Sim93a, Sir97, Sir99, SFS95, Sni91b, SFY92, SCBFC93, SCB^{+}93, Str98, ST95, Sug94, Sun91, SB99, SBS99, SAS99, Tor90, TCS93, Tsi99, VL92, WCC99, Wan98a, WU99a, Wan99, WK99, WT90b, WA96a, WL96]. problems [Won98b, WHL^{+}99, XS90, Yan97, Yan93d, Yan99b, Yao93, YK95, Yos93, YL99, Zha93c, Zha93d, ZZ99d, ZZT99, ZR91, dKSG94, vdHS94, vdV95, Ano95-83, Ano98-55, Ano99-96, Ano99-97, Ano99-101, Ano99-167, Ano99-168, Ano99-183, Ano99-184, Ano99l]. Procedure [CPC93, AW94, BJH90, BCC95, CP95b, CKRV98, DdW93, FCP95, HB99, MG90c, PC96a, Ram90a, Sim98]. Procedures [KS92, FGN90, Gla95b, Gla96a, Gla97a, Gla98]. Proceedings [Ano95k, Ano95-36, Ano95-45, Ano96-67, Ano96-104, Ano96-103, Ano97g, Ano97-53, Ano97-47, Ano97-57, Ano97-88, Ano98h, Ano98-84, Ano98-83, Ano98-96, Ano98-94, Ano98-95, Ano98-111, Ano98-155, Ano99-29, Ano99-30, Ano99-90, Ano99r, Ano99t, Ano99q, Ano99-128, Ano99-136, Ano99-137, Ano99-152, Ano99-212, Ano96e, Ano98-101, Ano96-135]. process [AO93a, Ano95-27, Ano98-164, BTT90, Cha94, Cho96, Din96, Din97b, DKO93, Hua98c, Pou90a, Ram95, RZ91, Ano95-27, Ano96-151]. Processes [Alf95, AK93, Ano99-190, Ano99-191, AKS96, ABI99, Bag90, BSP94, DKSI94, Eso91, Faz94a, Faz94b, Har96, KSK90, KP97b, Kra92, Kur91, KK94, Lon95, Nis95, Oda91, Oda94, Ohn92, PBG90, Pap94a, PZ99b, PL99, PT91, PS90, RT96, Ruz90, Sar90, Yos93, Yos90, Ano95t, Ano95-73, Ano97-29, Ano97-149, Ano99-174]. Processing [Ano98-43, KHC92, Ano96i, Ano96-32, Ano96-120, Ano97g, Ano98h, Ano98-27, Ano98-32, Ano99q, Ano99-118, Ano99-119, CD98, DKSI94, Hen97, Jan96, KKOS96, Lyt92b, ML90a, Pio94, Pou90b, RGBL96, Ano98-58, Ano99x]. processor [BLR96, CK97a, MMMS96, Vos91]. processor-time [MMMS96]. Processors [EK92, BGM95, CCK97, CK97a, Ano98-137]. Product [DR97, LW97, AAD98, ARS98, Ber93, DO95, Gro95, JAK96, Las94, Las95, LJL^{+}93, Mit90, Sni91a, TYC93, Ves98, WW98, Web95, WY93]. product-type [Web95]. Production [BTED97, BT95, CJ90, JM91, RG90a, Sch92a, Tsi91, TT92, WF96, YF92b]. production/inventory [RG90a]. productivity [Ano98-131, Ano96-162]. products [Ano97-163, Ano99-216, Ano99-217, Tre95b]. profiles [Ano95-69, Fre98]. program [Ano98-102, Bis91]. programmer [Ano98-149, Ano99k, Ano97-95]. programmers [Ano96-92, Ano96-109]. Programming [Ano96-107, SZM98, AL94, Ano95-38, Ano95-78, Ano95-45, Ano95-52, Ano95-57, Ano95-81, Ano95-82, Ano96-28, Ano96-67, Ano96-73, Ano96q, Ano96-90, Ano96-104, Ano96-103, Ano96-119, Ano96-109, Ano96-134, Ano97f, Ano97u, Ano97-35, Ano97-46, Ano97-57, Ano97-73, Ano97-88, Ano97-159, Ano97-166, Ano98-80, Ano98-96, Ano98-94, Ano98-95, Ano99r, Ano99p, Ano99-47, Ano99-48, Ano99-63, Ano99-55, Ano99-64, Ano99-56, Ano99-69, Ano99-74, Ano99-99, Ano99-109, Ano99-100, Ano99-110, Ano99-165, Ano99-170, Ano99-180, Ano99-228, BJH99, Bec90, CCK97, CF98, Che90b, CLD97, EB93, EB91, EA90, Eso99b, FHWW99, Fog94, FGN90, Gal99b, GHPLZ90, GF97, HJ91, HM94, Ill91, ISM93, Iwa91, Jan96, Jay90, KNU99, Kra90, LTC98, Leo96, LH94, LH99, LL90a, LH91, LL91a, LB94, Lin94, LL90c, Liu96b, Liu97, Liu98, LY99b, LKT99, Més96, MR99]. programming [MKS^{+}91, NL93, Oda94, Pan90b, Pan98a, Pan98b, PJ91, PLH95, Paz90, Ram90b, RG90a, RG90b, SNU98, SWF95, Shi95a, Sny90, TG95, TWIF98, VV99, WF96, WK99, XLJS98, ZS97, Ano96-49, Ano93h, Ano95h, Ano95-45, Ano96-67, Ano96-104, Ano96-114, Ano96-132, Ano97-72, Ano97-76, Ano97-85, Ano97-84, Ano97-88, Ano98-96, Ano98-95, Ano98-143, Ano98-154, Ano99r, Ano99m, Ano99-43, Ano99-100, Ano99-139, Ano99-179, Ano99-230, Ano97-123, Ano97-124, Ano99-193, Ano99-194, Ano99-195, Ano99-192]. programming.

[Ano97z]. programs [AK90a, AH92, Ano95-38, Ano97o, Ano97-142, Ano98p, AK90b, Eso99a, FW96, GRZ96, LV92, PL90, WF99b, ZC91, Ano96k, Ano97-137]. progress [Ano98-135]. progressions [Mor93]. prohibited [SS98b]. projectile [FY93, YFF92b, YFF92a, YFF93]. Projection [Che93b, Ahm90, Bar91b, CH99, DEMS94, Gal99a, HW96, LM96b, ST94]. projections [CP95a, POH99, Sny90, Ano98-60]. projective [Ano95-55, Ano99-153, Pan98b, TD99]. Projector [GSC98]. Projector-splines [GSC98]. Projectors [Den91b, Mär96]. projects [Ano97-68]. Prokhorov [AR99a]. prolog [Mit90]. Prömel [Ano98-91]. promising [KIAM98]. Prony [Sal92b, Sal91]. proof [Ano97p, Ano98-91, Ano99-87, Gyi96, Lu96, WW93, Ano97-94]. proofs [Hen97, Mro96]. Propagation [Ber98, ABVB92, Ano96-91, Ano99-172, AÉ96, GHM96, MDVK97, SC98b, Vic90]. propagators [LX98]. proper [AO99b]. Properties [DBM95, Kra92, SSBG91, ZG97, Ahm90, Ano90u, BLR94, DSKP96, Far96a, FL96, FK97, FBL99, GL94, IK97, Isp98, IZ97, KW98, Li99a, MPP97, MP96, MO95, MR94, Mic98, Pan95, PZ99b, Ruz90, SS93, Szu96, TMGS98, TR99, Yan93b, ZA95]. property [AD98, Ano98-148, BSP94, Fil93, Sok91, ZY96]. property-casualty [Ano98-148]. proportional [Ise94]. proposal [dJPvZ97]. prospects [Ano95r]. Protecting [Ano98-129]. protein [HSH90a]. proteins [HSH90b]. protocol [LY99a]. Prototypes [FMS92]. prover [BKM95]. proving [AC95]. proximity [Cim90]. Pruess [Ano96-75, Con95]. PSAS [DF98]. Pseudo [SL99, CSW99, Gla94e, HB99, hCPW98]. pseudo-contractive [HB99]. Pseudo-Random [SL99]. pseudo-unsteady [Gla94e]. pseudocontractive [Din96]. pseudoproper [TCS94]. pseudorandom [LKC95, PW95]. Pseudorandomness [Ano96-136]. Pseudospectra [JOW98]. pseudospectral [Ano99-244, BB99, Sak90]. psi [MM95c]. PSN [Myl92]. Psycholinguistic [Ano96-32]. psychology [Ano97-31]. public [Ano98-39]. publication [Ano92k, Ano92l, Ano92m, Ano93-35, Rod91a, Rod91b, Rod92a, Rod92b, Rod92c, Rod92d, Rod92e, Rod92f, Rod92g, Rod93c, Rod93d, Rod93e, Rod93f, Rod93g, Rod93h, Rod93i, Rod93j, Rod93k]. Publications [Ano95-27, Ano97n, Ano97-148, Ano97-117, Ano97-162, Ano97-129, Ano98-70, Ano98-158]. Publications/Cambridge [Ano97-148, Ano97-129, Ano98-70]. Publications/FoLLI [Ano97n]. Publishers [Ano93m, Ano93q, Ano97e, Ano97-29, Ano97-36, Ano97-45, Ano97-58, Ano97-56, Ano97-66, Ano97-85, Ano97-84, Ano97-108, Ano97-105, Ano97-104, Ano97-126, Ano97-157, Ano97-164, Ano97-163, Ano98g, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98w, Ano98x, Ano98r, Ano98z, Ano98-37, Ano98-31, Ano98-32, Ano98-36, Ano98-45, Ano98-49, Ano98-53, Ano98-50, Ano98-47, Ano98-59, Ano98-56, Ano98-57, Ano98-65, Ano98-66, Ano98-68, Ano98-67, Ano98-69, Ano98-71, Ano98-72, Ano98-76, Ano98-84, Ano98-74, Ano98-83, Ano98-89, Ano98-90, Ano98-92, Ano98-109, Ano98-108, Ano98-100, Ano98-101, Ano98-110, Ano98-115, Ano98-111, Ano98-116, Ano98-121, Ano98-123, Ano98-126, Ano98-127, Ano98-133, Ano98-137, Ano98-132, Ano98-135, Ano98-144, Ano98-138, Ano98-139, Ano98-140, Ano98-153, Ano98-157, Ano98-155, Ano98-156, Ano98-160, Ano98-161, Ano98-164, Ano99z, Ano99-27, Ano99o]. Publishers [Ano99r, Ano99s, Ano99n, Ano99t, Ano99u, Ano99-33, Ano99-50, Ano99-213, Ano99-49, Ano99-63, Ano99-64, Ano99-52, Ano99-57, Ano99-58, Ano99-84, Ano99-85, Ano99-86, Ano99-83, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-87, Ano99-105, Ano99-103, Ano99-99, Ano99-96, Ano99-102, Ano99-104, Ano99-100, Ano99-97, Ano99-95, Ano99-101, Ano99-115, Ano99-116, Ano99-117, Ano99-120, Ano99-114, Ano99-128, Ano99-121, Ano99-136, Ano99-137, Ano99-131, Ano99-145, Ano99-142, Ano99-152, Ano99-153, Ano99-167, Ano99-168, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-170, Ano99-174, Ano99-183, Ano99-184, Ano99-181, Ano99-182, Ano99-196, Ano99-197, Ano99-201, Ano99l, Ano99-206, Ano99-207, Ano99-204, Ano99-209, Ano99-229, Ano99-212, Ano99-211, Ano99-227, Ano99-234]. Publishers/Ister [Ano98-74]. Publishing [Ano95-72, Ano95v, Ano95-35, Ano95-34, Ano95-42, Ano95-50, Ano95-48, Ano95-56, Ano95-61, Ano96j, Ano97-132, Ano99-59, Ano96-59, Ano98-47]. Pugh [Ano95o]. Pulleyblank [Ano98-29]. pulse [MDVK97, Yav93]. punch [DPD98]. Purcell [Lai97]. purchased [Ano97-141]. pure [Ano99-224, Bai96a]. pure-strategy [Bai96a]. Purely [AR90b]. purine [SHT^{+}90]. pursuer [GOWR^{+}93]. pursuer-evader [GOWR^{+}93]. Pursuit [II93, Gal93b, Gue93, Yav91a]. Pursuit-evasion [II93, Gal93b, Yav91a]. PVM [Ano95-59]. pyramids [Ano97-115]. pyrimidine [SHT^{+}90]. Python [Ano97-123].

QAP [BT94]. Qiang [Ano98-37]. Qt [Ano99-195]. quadranguli [Köh97]. Quadratic [Ara90, AK98, Alt90b, Che99c, EDD91, EB91, GRZ96, GSM91, GM96, KAD99, Kou90a, LB94, Par91a, PJ91, SZ98b, Sza91, Ung90, Ung91]. Quadrature [Ola90, DW97, FLR92, FLR95, Gau97, Kim98, PW97, Rok90, SZ91, VL95, Zha91]. quadrature-collocation [SZ91]. quadratures [CFM91, Mil98]. quadric [BM91, BC97]. quadrilaterals [xLhW96]. Quagliarella [Ano98-64]. Qualitative [Far96a, Ano95-60]. qualities [Mar92]. quality [Ano96e, Ano96f, JM91, ML90c, Pou90b, RDK92, Tsi91, TT92]. quantitative [Ano96-155, SZ98b]. quantity [Tsi91]. quantity-quality [Tsi91]. quantization [Ram95, SHK99]. Quantum [AS97, Kor92, KLMP97, Ano95-54, Ano99-166, AKK^{+}97, BMM94, BGM97, Che96, DKR97, Gre91a, Gre92b, Gre93b, RV94, RP97, VV99, Vic90, Ano99-196]. quantum-like [Vic90]. QuarkXPress [Ano98-130]. quarter [Ano99-198]. quarter-plane [Ano99-198]. quartic [Ung90, Ung91]. Quartics [EL92]. Quasi [RT99a, Bou98, Cla98, CY94, Cub95b, CLY97, DR97, Den91a, Din99a, Din99d, EZ93, FM96, Gre91a, Gre92b, Hua98a, KT92, KT95, LN99, Noo96, Noo98, Sha90d, Van92, WT93, Wu99b]. quasi-equilibrium [Cub95b]. quasi-equivalence [Van92]. quasi-Gauss [KT95, WT93]. quasi-interpolatory [DR97]. quasi-linear [LN99]. quasi-Newton [FM96, KT92]. quasi-periodicity [Den91a]. quasi-primitive [Cla98, EZ93]. quasi-quantum [Gre91a, Gre92b]. Quasi-Trefftz [RT99a]. quasi-variational [Bou98, CY94, CLY97, Din99a, Hua98a, Noo96, Noo98, Wu99b]. quasi-variational-like [Din99d]. quasiattractor [Byk97]. Quasiattractors [GST97]. quasiconformal [Ano97-30]. Quasigroup [SFS95]. quasilinear [TR99, VL92]. quasilinearization [Ano99-96, Ano99-97, NWL91]. Quasioptimal [Li97]. quasipower [Pas97]. quasivariational [Cub95a, Din99b]. quasivariational-like [Cub95a]. Quaternionic [Ano97-125]. Queinnec [Ano97-83]. Quentin [Ano97-115]. Quercia [Ano98-79]. queries [Rei94, Rei99a]. QUEST [GGB92]. question [GGB92, LL90b]. questions [Ano99-39]. queue [CTG96, CKL98, Tan96, Toy92, Wu90a, YI97]. Queueing [NJ92, AD93, AS93, Alm96, AR92e, CK97b, Gup94, Kat92, MC93, Mis90, RHG91, SG90, TBA99, TT92, YF93]. queues [Alf95, Ano96-171, Kij92, Kim92, YSM92]. queuing [CT90]. quick [Ano96-92, Ano97-169, Ano98-79, Ano98-103, Ano98-130, Ano99-31, Ano99-62, Ano99-143, Ano99-187, Ano99-210]. quickening [Ano97-133]. quickstart [Ano98-160]. Quilici [Ano97-145, Ano98-37]. Quintic [WA94a, KC92b, Ano97-60]. quintics [EL92]. quorum [FLNW98].

R [Ano93o, Ano93w, Ano95x, Ano95z, Ano95-38, Ano95-44, Ano95-65, Ano96g, Ano96-54, Ano96-55, Ano96-67, Ano96-61, Ano96-71, Ano96-75, Ano96-81, Ano96-80, Ano96o, Ano96-102, Ano96-114, Ano96-165, Ano97-142, Ano97w, Ano97v, Ano97-57, Ano97-127, Ano98r, Ano98-29, Ano98-42, Ano98-56, Ano98-65, Ano98-140, Ano99u, Ano99-41, Ano99-47, Ano99-42, Ano99-48, Ano99-59, Ano99-120, Ano99-136, Ano99-199, Ano99-200, Ano99-211, Ano99-227]. R&D [Ano98-131]. Radial [KMD99, Ano97-130, BN92, CF92, Gir92, Mye92, Pas92, RK98, Sal92a, SJB97, SM94, SM95, XC92]. radiation [EGK96, GBB91, KK97a, UW94]. radiative [WU99a]. radical [LJL^{+}93]. radically [Ano96-100]. radicals [Ung90, Ung91, VR95a]. radio [BP90, Log96]. radius [RSR98]. radix [JP93, Koç91]. Rado [BL97]. radon [Hsu96, MS93]. Rafaeli [Ano98-112]. Raff [Ano95x]. Raghavan [Ano95-66]. rail [LR96a]. raindrops [Gre91b]. Rainer [Ano98-137]. Raines [Ano99-210]. Rajeev [Ano98-163]. Rajiv [Ano98-50]. Rakhimov [Ano98-89]. Ralph [Ano95-73, Ano96-43].

Ram [Ano96-78]. Ramamritham [Ano99-58]. Ramasamy [Ano96h]. Ramesh [Ano99-209]. Ramkissoon [Ano98-83]. Ranade [Ano99x]. Randall [Ano99e]. Randolph [Ano99-92]. Random [Fox94, KW98, MZ93, SL99, AR98, Ano92a, Ara90, Ber98, KT97, KT98, Kor90b, Leo90, PS90, Sez95, SD93, Sod90, Sza90b, TBA99, Web95, Ano93p, Ano99-197, Ano99-198]. Randomization [Cor95, Cse95]. randomized [AC93b, KO90]. randomly [Yav92]. randomness [Gry94]. Randy [Ano99-164]. range [MP96]. ranges [LR96b]. ranging [KKK91, KKK93]. rank [MPR98]. rank-deficient [MPR98]. Ranka [Ano97-44]. Ranking [CL94c, Che99c, JK96]. Rapcsák. [Ano99-165]. Raphson [Ohn92]. rapid [Gre92b]. rarefied [ASW98, KESH90, SC98b]. RAS [Ege91]. Rassias [Ano95-33]. rate [And93, Bai96e, CR90a, CR90b, MPS93, Ram91a, Sha96a, Tan96, YC94, YI97]. Rates [AR99a, CD98, Mit92a]. Ratio [BW93]. Rational [Gem93, HV92, JP93, Las94, Las95, xLhW96, AK98, BM97, CLRN94, CGS99, DO91, FO91, GP93, LR93b, MM95c, PR91, SGS93, Wac91a, vNvN90a, vNvN90b]. rationals [CW95]. rationing [pri94]. ratios [Shi95a]. Raton [Ano95w, Ano95z, Ano95g, Ano98-43, Ano98-107, Ano99-108, Ano99-125, Ano99-126]. Ratz [Ano93o]. Ravi [Ano97e, Ano99-84, Ano99-85]. Ravindra [Ano99-28]. Rawlins [Ano97-133]. Ray [Ano96-97, Oli93]. Rayleigh [Dea98, KESH90]. Raymond [Ano97-150]. Raymund [Ano98-38]. Razumikhin [ZC98]. reachability [De 96]. Reaction [LN98, Li10, Boy96, BM89, BM90c, CCVV90, CC98a, Cho92, MB94a, Pao98, Par91b, Ram91a, VK99]. reaction-diffusion [Boy96, CC98a, Cho92, Par91b, VK99]. reactions [Gre92b]. reactive [LSH98]. reactor [JR94]. reactors [BCC95]. Reading [Ano95-72, Ano95v, Ano95e, Ano95-35, Ano95f, Ano95-79, Ano95-42, Ano95-57, Ano95-61, Ano95-64, Ano96-151, Ano96-86, Ano97-107, Ano99-81, Ano95-48]. Readings [Ano97-154, Ano96-36]. Real [Gla92d, Nam99b, Nel96, Ano96-137, Ano97-148, Ano97-102, Ano98z, Ano98-68, Ano98-89, Ano98-138, Ano99z, Ano99-27, Ano99-58, BWK97, Bro94, Gla90a, Gla91d, Gla92b, Gla94b, Gla97a, Gla99, GM99, GC99, MB93, Som90, Szu96, Ano96-137, Ano97-126, Ano98-132]. real-time [Ano98z, Ano98-68, Ano98-138, Ano99-58, Ano97-126, Ano98-132]. realistic [IIP^{+}97, MPP97]. reality [Ano95p]. Realization [TCS94, BM90a, Ter91]. reason [Ano97-80, Ano97-127, Ano96a]. reasoning [Ano96-53, Ano97r, Bal90, LC97a, LCSN98, PM95a, SFS95, Sti93a, Wos95b, Ano95-60, Ano96-38, Ano97-109, Ano96-138, Ano98-133, Ano98-134]. reassembly [Ano96-96]. Rebelsky [Ano98-47]. receptivity [GPCSL90]. receptor [HSH90b]. receptors [KM93]. recipes [Ano90w]. recoding [KH92]. recognition [Ano96-84, BFF^{+}97, HSH90a, HSH90b, PS90]. Recognizing [LT94]. reconstruction [GCLT98, MRZ96]. Records [Dev93, Ano96-157]. Recovery [HM93b, Ano98-37, FKO92, Kou95, XLJS98, Ano99-57]. recta [MT98]. rectangles [NB95]. rectangular [CHW93, Che91, ER98, HS97, OB97]. Rectilinear [Has93b]. Recurrence [GH98, AP98a, Li98c, Tre94, WK94]. recurrences [CLY91]. Recurrent [Wan93b, MKS^{+}91, Sol90, Yos99]. recursion [Ano96-35]. Recursive [SS99, CLRN94, Cha98, EDO98, KT97, KT98, Kou90a, SBD99, Tan92, UK99]. redesign [SZC91]. Redfern [Ano96-156, Ano96-159, Ano97-147, Ano99-218, Ano99-219]. REDUCE [RZ91]. reduced [AF90, CP98b, Yav91a]. reducing [FCP95]. Reduction [Lew91b, MM95c, Amo93, CM93, Fli94b, Fli94c, JLDL92, KMŚ98, KKM96, KM97b, KPB97, Lee93, PC94a, SK91]. redundant [NMUY92]. reengineering [Ano98-37]. Reese [Ano97-35, Ano99-164]. Reeves [Ano97-148]. reference [Ano93o, Ano95-42, Ano96-92, Ano97-169, Ano99-158, Ano99-160, Ano99-159, Ano99-161, Ano96-118, Ano97-75, Ano97-74, Ano98-61, Ano98-79, Ano98-103, Ano98-130, Ano99-31, Ano99-62, Ano99-143, Ano99-176, Ano99-178, Ano99-187, Ano99-210]. Refining [Wu90a]. reflecting [Oli93]. reflection [Gla90b, IGM98, Mat99, Sym91, Wan93a]. Reflections [Eso90]. reflective [Ano99-140]. reformulation [Ano99l]. reformulation-linearization [Ano99l]. regenerative [AD93]. Regier [Ano97-144]. region [Bou93]. regional [KM92]. Regionalism [Ano98-136]. regions [Kan92a, Köh97, RSAR95]. register [CCK97, SRCB98]. registry [Ano98-98]. regression [CL94a, CL94b, Che99c, CL99b, CL99c, Ita93, Kou90a, Sch92b, Zhu92, Ano99-199, Ano99-200]. regression- [CL99c]. regressions [Ano99-199, Ano99-200]. Regular [Far99, Ano97-89, Ano97-115, Bro90, CF91b, IGM98]. regularity [Ano99-82, BMN96]. Regularization [HMZ99, MS93, MPR98, Ech95]. Regularized [BCD98, Gro95]. regulation [Ano97-37]. regulators [GSM91]. regulatory [HSH90a]. Reid [ACHP94]. Reiner [Ano97-42, PSD96]. reiner-rivlin [PSD96]. Reinforcement [Ano93q]. Related [RRS99, AK90a, Ano96f, Ano99-58, AI97, BS96d, Bou98, CCH^{+}90, DKR97, Feu91, Gol90, Gyi96, Il’97, IK97, JA97, Kai96, KK97a, KO91, KM90, Mad92, Sri95, Yan93d, Yan99b, dOF91, Ano95v]. relatedness [Joh95]. Relation [DT93, Hsu96, DH91, LZ94, vdVVdB99b]. Relation-based [Hsu96]. Relational [DT93, Ano98-158, BKM91, GT92, Gün99, MS91, ZO92, Ano95-61]. relations [BHK95, CH99, GT92, GH98, Li98c, MS91, OPD97, Özy96, WK94, Yas94]. relationship [Gla95c, Gla96c, Pol90]. Relationships [ZFM96, LL92, Lia99b]. Relative [KPS96, GSM91]. relativisation [Kup91]. relativistic [Ano98v, NR96a]. Relativity [Ano97-130, Ano95i]. relaxation [BW96a, BW96b, Bai96c, Bai96d, CRV98, JOW98, JMPS94, Kou98]. relaxed [WF99b]. release [Ano96-156, Ano96-159, BLR96]. Reliability [CCY95b, YNK92, CHCC95, CCY95a, CCH97, KYO92, RD99, SZ98a, YOO92]. reliability-oriented [CHCC95]. reliable [Cho99, JM91, KO93]. Relocation [KOT99]. remainder [Tse95]. remark [ABK97]. Remarks [LH97, Lei96, WCC99, FO91, Fun97, OPD97, Riv97, Sez95, Wit90, Zal95]. Remodelling [MBMS92]. remote [CH93]. Ren [Ano99-137]. renaissance [Ano99-222, Ano99-223]. renal [MT98, SMTS97, WSDT94]. repair [Jos90, SOI92]. repairable [Cse95]. Repeated [Ano95-62]. repetitive [Kra92]. replacement [SOI92]. replicated [FLNW98]. replicating [Ban93]. replies [Ano96-50]. Reply [Los92, ZQ95]. report [Ano94c, Ano95b, Ano98c, Ano99b, Ano98-87]. Reports [Ano98d, Rod94b, Rod94c, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93-32, Ano93-33, Ano93-34, Ano97c, Ano98e, Ano99c, Rod93a, Rod94a, Rod94d, Rod94e]. Repovs [Ano99-49]. Representation [Arn90, Ano99-142, Bha94, Che97, DIT97, EL92, GM99, Har92c, LSH95, MB93, Reg94, Sow92, Tse95, Zar92, ZA95, Ano96-97]. Representations [Mez97, Ano98-92, Ano99u, Ano96-139]. representing [Gla94e, HS95, Ano96-140]. Republic [Ano99-128]. requirements [Ano90x, Köh93b]. Resampling [DFMS99]. Research [Ano98-132, Ano96-47, Ano97-28, Ano97-45, Ano97-113, Ano97-114, Ano97-128, Ano98-82, Ano99-239, Wag90, Ano97-49, Ano97-55]. reservoir [LO99]. Residence [KMR90]. Residual [Fra96, MM96a]. Residual-minimization [Fra96]. residuals [Isp98]. residue [CL91a, CL91b, Chr90, HP94c, Koç91]. resin [SCG^{+}97]. resistive [RD96a]. Resnick [Ano97-160]. resolution [Ano97-152, DEMS94, GG97, Moo93a]. resonance [Wos95a]. Resonances [ADKM97, BAS97, BF97, KLMY97]. resonant [Jon96, MO98a]. Resource [Ano99-131, Sch92a]. resources [NL91]. respect [KKS98, Nis95]. Response [Alm96, Fli95, HSH90b]. rest [Ira93]. restarted [CP98b]. restarting [Fox94]. restoration [DGU90]. restricted [Ano92a, Man96, Tsi91]. result [Hua97, SCBFC98, TY99a]. resultant [Pan95]. results [Ado91a, Ado91b, Ahm94, Ame99, BL99, CE94, CLM90, Dem94a, HK96, MCG98c, Mei97, Mol99, O’R98b, RM98b, Sri95, Tse95, Ano99-95, Ano99-171]. resurrection [Ano99-239]. retarded [Riz95]. Retargetable [Ano98-137]. Retraction [Ano91n]. retrial [CK97b, CKL98]. Retrieval [GW99, Ano98-71, Ano98-109, Lev92, MS99, Ano93f, Ano98-77, Ano98-76, Ano99-50]. returns [Bra92]. reusable [Ano95-38]. Reuse [Ano99-201]. Reuven [Ano98-27]. reverse [HI94]. reverse-time [HI94]. reversible [CCGL99]. Review [Ano90o, Ano90p, Ano90q, Ano90r, Ano90s, Ano90t, Ano90u, Ano90w, Ano90v, Ano90x, Ano93f, Ano93g, Ano93t, Ano93u, Ano93h, Ano93i, Ano93j, Ano93l, Ano93k, Ano93n, Ano93m, Ano93o, Ano93p, Ano93q, Ano93v, Ano93r, Ano93s, Ano93w, Ano95k, Ano95-36, Ano95h, Ano95m, Ano95l, Ano95-70, Ano95n, Ano95o, Ano95r, Ano95s, Ano95d, Ano95q, Ano95p, Ano95t, Ano95-72, Ano95v, Ano95y, Ano95u, Ano95x, Ano95c, Ano95w, Ano95-71, Ano95z, Ano95-28, Ano95-27, Ano95-29, Ano95-74, Ano95-73, Ano95-30, Ano95e, Ano95i, Ano95-31, Ano95-75, Ano95-35, Ano95-34, Ano95-77, Ano95-32, Ano95-76, Ano95-33, Ano95-37, Ano95-38, Ano95-78, Ano95f, Ano95-39, Ano95-40, Ano95j, Ano95-79, Ano95-45, Ano95-41, Ano95-42, Ano95-44, Ano95-43, Ano95-52, Ano95-51, Ano95-47, Ano95-46]. Review [Ano95-80, Ano95-50, Ano95-48, Ano95-49, Ano95-53, Ano95-55, Ano95-56, Ano95-57, Ano95-58, Ano95-59, Ano95-60, Ano95-61, Ano95-63, Ano95-62, Ano95-65, Ano95-66, Ano95-64, Ano95-67, Ano95-81, Ano95-69, Ano95g, Ano95-82, Ano95-83, Ano95-54, Ano95-68, Ano96g, Ano96s, Ano96i, Ano96-150, Ano96e, Ano96h, Ano96k, Ano96f, Ano96t, Ano96j, Ano96l, Ano96w, Ano96y, Ano96u, Ano96x, Ano96v, Ano96-32, Ano96-37, Ano96-39, Ano96-27, Ano96-29, Ano96-42, Ano96-45, Ano96z, Ano96-44, Ano96-36, Ano96-151, Ano96-43, Ano96-33, Ano96-34, Ano96-31, Ano96-35, Ano96-28, Ano96-41, Ano96-153, Ano96-152, Ano96a, Ano96b, Ano96-38, Ano96-30, Ano96-40, Ano96-47, Ano96-46, Ano96-48, Ano96-50, Ano96-54, Ano96-55, Ano96-49, Ano96-53, Ano96-52, Ano96-51, Ano96c, Ano96-56, Ano96-69, Ano96-67, Ano96-65, Ano96-61, Ano96-57, Ano96-63]. Review [Ano96-66, Ano96-62, Ano96-60, Ano96-68, Ano96-71, Ano96-58, Ano96-72, Ano96-70, Ano96-59, Ano96-64, Ano96-74, Ano96-73, Ano96-75, Ano96-76, Ano96-78, Ano96-77, Ano96-86, Ano96-79, Ano96-87, Ano96-82, Ano96-81, Ano96-80, Ano96-88, Ano96-83, Ano96-85, Ano96p, Ano96-89, Ano96-91, Ano96q, Ano96-90, Ano96r, Ano96n, Ano96m, Ano96o, Ano96-92, Ano96-93, Ano96-95, Ano96-94, Ano96-96, Ano96-104, Ano96-100, Ano96-98, Ano96-99, Ano96-103, Ano96-154, Ano96-101, Ano96-102, Ano96-97, Ano96-119, Ano96-109, Ano96-114, Ano96-105, Ano96-116, Ano96-113, Ano96-111, Ano96-118, Ano96-106, Ano96-107, Ano96-112, Ano96-156, Ano96-108, Ano96-117, Ano96-120, Ano96-115, Ano96-110, Ano96-155, Ano96-122, Ano96-125, Ano96-124, Ano96-157, Ano96-123, Ano96-121, Ano96-126, Ano96-158, Ano96-134, Ano96-128, Ano96-129, Ano96-131, Ano96-136, Ano96-132, Ano96-159, Ano96-127, Ano96-133, Ano96-130]. Review [Ano96-138, Ano96-140, Ano96-137, Ano96-139, Ano96-148, Ano96-144, Ano96-141, Ano96-142, Ano96-161, Ano96-146, Ano96-147, Ano96-149, Ano96-143, Ano96-164, Ano96-163, Ano96-165, Ano96-162, Ano96d, Ano96-167, Ano96-84, Ano96-135, Ano96-145, Ano96-160, Ano96-166, Ano97g, Ano97-140, Ano97f, Ano97-141, Ano97-53, Ano97m, Ano97n, Ano97o, Ano97h, Ano97l, Ano97e, Ano97p, Ano97-142, Ano97q, Ano97r, Ano97-31, Ano97s, Ano97-28, Ano97w, Ano97-32, Ano97u, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97t, Ano97-30, Ano97v, Ano97-29, Ano97-40, Ano97-41, Ano97-33, Ano97-39, Ano97-34, Ano97-43, Ano97-37, Ano97-36, Ano97-38, Ano97-42, Ano97-35, Ano97-45, Ano97-44, Ano97-143, Ano97-46, Ano97-54, Ano97-48, Ano97-47, Ano97-51, Ano97-50, Ano97-49, Ano97-62, Ano97-57, Ano97-63, Ano97-61, Ano97-60, Ano97-55, Ano97-59, Ano97-58, Ano97-56, Ano97-144]. Review [Ano97-64, Ano97-65, Ano97-72, Ano97i, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97-68, Ano97k, Ano97-69, Ano97j, Ano97-66, Ano97-75, Ano97-77, Ano97-76, Ano97-74, Ano97-78, Ano97-145, Ano97-83, Ano97-81, Ano97-85, Ano97-84, Ano97-79, Ano97-82, Ano97-88, Ano97-86, Ano97-80, Ano97d, Ano97-146, Ano97-87, Ano97-149, Ano97-92, Ano97-95, Ano97-93, Ano97-148, Ano97-89, Ano97-96, Ano97-90, Ano97-147, Ano97-98, Ano97-99, Ano97-97, Ano97-100, Ano97-91, Ano97-94, Ano97-111, Ano97-101, Ano97-112, Ano97-102, Ano97-150, Ano97-106, Ano97-109, Ano97-103, Ano97-110, Ano97-108, Ano97-107, Ano97-105, Ano97-104, Ano97-113, Ano97-114, Ano97-116, Ano97-115, Ano97-117, Ano97-122, Ano97-123, Ano97-119, Ano97-120, Ano97-124, Ano97-151, Ano97-118, Ano97-121, Ano97-125, Ano97-152, Ano97-126, Ano97-137, Ano97-135, Ano97-156, Ano97-158, Ano97-153, Ano97-155, Ano97-128, Ano97-138, Ano97-154]. Review [Ano97-131, Ano97-136, Ano97-134, Ano97-132, Ano97-133, Ano97-127, Ano97-157, Ano97-160, Ano97-139, Ano97-161, Ano97-164, Ano97-162, Ano97-163, Ano97-165, Ano97-159, Ano97-166, Ano97-169, Ano97-167, Ano97-168, Ano97-52, Ano97-129, Ano97-130, Ano98g, Ano98m, Ano98h, Ano98q, Ano98o, Ano98j, Ano98i, Ano98l, Ano98n, Ano98k, Ano98p, Ano98s, Ano98u, Ano98w, Ano98x, Ano98r, Ano98v, Ano98-34, Ano98-29, Ano98-28, Ano98-38, Ano98z, Ano98y, Ano98-27, Ano98-37, Ano98-33, Ano98-31, Ano98-32, Ano98-35, Ano98-36, Ano98-30, Ano98t, Ano98-39, Ano98-44, Ano98-42, Ano98-41, Ano98-40, Ano98-46, Ano98-45, Ano98-43, Ano98-52, Ano98-54, Ano98-51, Ano98-48, Ano98-49, Ano98-148, Ano98-53, Ano98-50, Ano98-47, Ano98-60, Ano98-61, Ano98-58, Ano98-59, Ano98-56, Ano98-55, Ano98-57, Ano98-64, Ano98-65, Ano98-66, Ano98-63, Ano98-68, Ano98-67]. Review [Ano98-69, Ano98-70, Ano98-71, Ano98-73, Ano98-72, Ano98-78, Ano98-77, Ano98-82, Ano98-79, Ano98-76, Ano98-84, Ano98-74, Ano98-75, Ano98-81, Ano98-80, Ano98-83, Ano98-89, Ano98-88, Ano98-87, Ano98-86, Ano98-85, Ano98-96, Ano98-94, Ano98-93, Ano98f, Ano98-90, Ano98-95, Ano98-91, Ano98-92, Ano98-109, Ano98-102, Ano98-105, Ano98-99, Ano98-106, Ano98-97, Ano98-98, Ano98-108, Ano98-100, Ano98-101, Ano98-103, Ano98-110, Ano98-107, Ano98-104, Ano98-115, Ano98-113, Ano98-117, Ano98-149, Ano98-111, Ano98-118, Ano98-112, Ano98-114, Ano98-116, Ano98-121, Ano98-120, Ano98-119, Ano98-123, Ano98-124, Ano98-126, Ano98-127, Ano98-128, Ano98-129, Ano98-150, Ano98-122, Ano98-125, Ano98-130, Ano98-133, Ano98-137, Ano98-136, Ano98-131, Ano98-132, Ano98-135, Ano98-134, Ano98-151, Ano98-141, Ano98-142, Ano98-144, Ano98-138, Ano98-143, Ano98-139, Ano98-140, Ano98-153, Ano98-145, Ano98-157, Ano98-146]. Review [Ano98-155, Ano98-152, Ano98-147, Ano98-154, Ano98-156, Ano98-159, Ano98-158, Ano98-160, Ano98-161, Ano98-163, Ano98-162, Ano98-164, Ano98-62, Ano99-29, Ano99z, Ano99v, Ano99-30, Ano99-27, Ano99w, Ano99-90, Ano99o, Ano99r, Ano99s, Ano99n, Ano99t, Ano99y, Ano99x, Ano99-31, Ano99u, Ano99-28, Ano99-32, Ano99q, Ano99m, Ano99p, Ano99-35, Ano99-36, Ano99-33, Ano99-34, Ano99-37, Ano99-41, Ano99-47, Ano99-42, Ano99-48, Ano99-50, Ano99-213, Ano99-44, Ano99-51, Ano99-49, Ano99e, Ano99-39, Ano99-40, Ano99d, Ano99-43, Ano99-46, Ano99-45, Ano99-38, Ano99-63, Ano99-66, Ano99-53, Ano99-55, Ano99-64, Ano99-67, Ano99-54, Ano99-56, Ano99-69, Ano99-52, Ano99-65, Ano99-60, Ano99-57, Ano99-61, Ano99-62, Ano99-59, Ano99-58, Ano99-68, Ano99-75, Ano99-78, Ano99-76, Ano99-79, Ano99-73, Ano99-71, Ano99-74, Ano99-77, Ano99-72, Ano99-70]. Review [Ano99-84, Ano99-85, Ano99-86, Ano99-92, Ano99-83, Ano99-81, Ano99-88, Ano99-80, Ano99-94, Ano99-93, Ano99-214, Ano99-89, Ano99-82, Ano99-87, Ano99-105, Ano99-103, Ano99-99, Ano99-106, Ano99-96, Ano99-109, Ano99-102, Ano99-104, Ano99-100, Ano99-107, Ano99-97, Ano99-110, Ano99-95, Ano99-98, Ano99-101, Ano99-108, Ano99-115, Ano99f, Ano99-111, Ano99-118, Ano99h, Ano99-116, Ano99g, Ano99-112, Ano99-119, Ano99i, Ano99-113, Ano99-117, Ano99-120, Ano99-114, Ano99-216, Ano99-123, Ano99-125, Ano99-129, Ano99-133, Ano99-217, Ano99-124, Ano99-126, Ano99-130, Ano99-134, Ano99-215, Ano99-135, Ano99-128, Ano99-121, Ano99-136, Ano99-137, Ano99-132, Ano99-127, Ano99-131, Ano99-122, Ano99-139, Ano99-138, Ano99-140, Ano99-145, Ano99-143, Ano99-141, Ano99-144, Ano99-142, Ano99-158, Ano99-160, Ano99-218, Ano99-156, Ano99-154, Ano99-159, Ano99-161, Ano99-219, Ano99-157, Ano99-155, Ano99-152, Ano99j]. Review [Ano99-151, Ano99-163, Ano99-149, Ano99-146, Ano99-162, Ano99-153, Ano99-164, Ano99-148, Ano99-150, Ano99-147, Ano99-167, Ano99-222, Ano99-168, Ano99-223, Ano99-165, Ano99-169, Ano99-171, Ano99-172, Ano99-166, Ano99-220, Ano99-221, Ano99-170, Ano99-173, Ano99-174, Ano99-176, Ano99-177, Ano99-178, Ano99-179, Ano99-175, Ano99-183, Ano99-193, Ano99-190, Ano99-184, Ano99-194, Ano99-191, Ano99-180, Ano99-185, Ano99k, Ano99-181, Ano99-195, Ano99-187, Ano99-188, Ano99-224, Ano99-192, Ano99-186, Ano99-182, Ano99-189, Ano99-196, Ano99-199, Ano99-200, Ano99-197, Ano99-201, Ano99l, Ano99-198, Ano99-206, Ano99-207, Ano99-204, Ano99-225, Ano99-203, Ano99-209, Ano99-202, Ano99-205, Ano99-226, Ano99-208, Ano99-229, Ano99-212, Ano99-230, Ano99-210, Ano99-211, Ano99-231, Ano99-228, Ano99-233, Ano99-232, Ano99-227, Ano99-234, Ano99-235, Ano99-236, Ano99-91, Ano99-238, Ano99-237, Ano99-239, Ado91a, Ado91b, Con90]. Revised [Ano98-46, Ano98-106, Has93c]. revisited [YFF93, Ano97l]. revolution [Ano97-155, Ano99-173]. revolutions [Ano97-158]. reward [KYNY99, TPF98]. rewards [Yos96]. Rewrite [KZ95]. rewriting [Ano98-147]. Reynolds [Ano96-67, Ano98-145, Ano98-146, Ano99-228]. Rham [Ano99-208]. rheology [Ano99-137]. Rhodes [Ano96-130, Ano99-180]. Ribenboim [Ano96-157]. Riccati [Ano99-245, Ano99-98, Bal98, BW93, FLW97, GPCGV92, HI94, HG98]. Richard [Ano93j, Ano93l, Ano93q, Ano95m, Ano95-70, Ano95w, Ano95-75, Ano95-35, Ano95-80, Ano95-62, Ano96-52, Ano96q, Ano96-95, Ano96-116, Ano96-106, Ano96-149, Ano97-86, Ano97-96, Ano97-91, Ano98-149, Ano99-29, Ano99-30, Ano99-72]. Richardson [Ano96-147, dJF91]. Rick [Ano97-57]. Riecan [Ano98-74]. Riemann [AAL99, Gla91d, Gla92a, Gla93a, Gla94a, Gla94f, Gla95a, Gla95c, Gla96c, Gla97a, Gla98, Gla99, Shi95b, TAD91, Wen99]. Riesz [BAS97]. rigged [ADKM97, BAS97]. right [JCL93]. rigid [Dem94b, LQ95a, OHI91, ZG97]. rigorous [Bou99]. Rijal [Ano97-84]. ring [Ano98-53, MNB93]. rings [Ano98-115, Leb98, Ano95-32, Ano99-204]. Rio [Ano97-51]. Riolo [Ano97-57]. Rising [Ano98-150]. risk [Ano98-39, Ano98-62, Ano99-214]. risks [Ano99-38, VMA99, Ano95v]. Risky [Ano96-152]. River [Ano96-102, Ano97-113, Ano97-114, Ano99-185, Sau92]. Rivest [Ano95r, Ano95s]. rivlin [PSD96]. RK [hCPW98]. RKN [QZ91, Con98, CSW99]. RKN-type [Con98]. RNS [DJM98, HP94c]. Roads [Ano96-153]. Rob [Ano98-141]. Robbin [Ano95-49]. Robert [Ano95t, Ano95-59, Ano95-62, Ano95-54, Ano96-45, Ano96-67, Ano96-128, Ano96-139, Ano97s, Ano97-38, Ano97-63, Ano97-116, Ano97-128, Ano98v, Ano98-143, Ano98-158, Ano99x, Ano99-122, Ano99-146, Ano99-220, Ano99-182, Ano99-203, Ano99-237]. Roberta [Ano96-158]. Roberto [Ano96-87, Ano99-68]. Roberts [Ano95-73]. Robertson [Ano99-222, Ano99-223]. Robichaux [Ano98-98]. Robin [Ano97-28, CC97, CKM99]. Robot [ALL^{+}96, Ano96-155, Ano99s, Ano99-32, BS99b, LCL94, Tau90a]. robotics [Ano93v]. Robust [WSS92, Zhe91, Ano99-98, DJM98, Li98b, Zhe93]. robustness [Pet92]. Roche [Ano98-58]. rod [FG97, FG98]. Rodney [Ano95k]. Rodolfo [Ano97-149]. Roe [Ano95-29, Gla94f, Gla95f]. Roelofs [Ano99-188]. Roger [Ano96-50, Ano98s]. Rogers [Ano96-145]. Roitblat [Ano96-34]. Rojo [ESIP98]. Rokyta [Ano97-165]. role [Ano97-55, MT94b, Mye92, SHT^{+}90]. Rolf [Ano98-156, Ano99-90]. roll [MPS93]. Rolland [Ano92b]. rolled [RD96b]. ROM [Ano96-59, Ano96-120, Ano97-111, Ano97-101, Ano97-112, Ano98p, Ano98t, Ano98-41, Ano98-105, Ano98-149, Ano98-122, Ano98-125, Ano98-162, Ano99y, Ano99-35, Ano99-36, Ano99-37, Ano99-69, Ano99-78, Ano99-79, Ano99-70, Ano99-129, Ano99-130, Ano99-154, Ano99-155, Ano99-180, Ano99-230, Ano99-227, Ano96-59, Ano99-221]. Roman [Ano99m, Ano99-52, Ano99o]. Romano [Ano96-113]. Ron [Ano95-31, Ano98-139].

Ronald [Ano90v, Ano95r, Ano95s, Ano96-47, Ano96-54, Ano96-130, Ano96-138, Ano98-42, Ano99-202]. Roorda [Ano98-157]. root [Ano99-123, Ano99-124, Gil91, JS98, JI94, Kel94, RJ94, RS96b]. rootfinding [Car94]. roots [NP96, PC93]. Rosebush [Ano96-59]. Rosenbrock [Zed90]. Roseneau [LA96]. Rosenfeld [Ano97-58]. Rosenschein [Ano96-40]. Rosinger [Ano99-181]. rotating [BMN96, Ber98, IM93, OHI91, WW91]. Rotation [Jar97, Kra90, Pav97]. rotor [BR95b, CM93]. rotor/fuselage [BR95b]. Roudolf [Ano99-198]. Rough [NSS92, Duc95b]. Rouillard [Ano98l]. round [Tso91]. round-off [Tso91]. roundabout [ACHP94]. rounding [AR98]. route [YT98]. routing [GS96b, HTT92, SSBG91, Ano99-209]. row [TCS94]. row-pseudoproper [TCS94]. Roxanne [Ano96-99]. Roy [Ano98-141, Ano99-190, Ano99-191]. Rozanov [Ano99-197]. Rozenberg [Ano99-66, Ano99-67]. RRL [KZ95]. RSA [Zha93a]. Rubinstein [Ano95d]. Rudolf [Ano97-48]. Rugg [Ano98-28]. rugged [UWT99]. Rule [KZ95, Bar93, HM96, YC94, Zha91]. rules [AR98, DC94, DW97, Rok90, Web95, Zad99]. run [BT94]. Runge [AS96a, BS90, BS96a, Bog96, BT98, CMR90, Coh94, CT96, DOS93, Gol94, Pat98, QZ91, Sar94, Sim93b, Sim93c, Sim93a, Tan93b, Tsi99, YQM99, hC96a, hC96b, vdHS94]. Running [Ano95-63]. Rusinkiewicz [Ano98-164]. Russ [Ano95-46, Ano97-115]. Russell [Ano96-43]. Russia [Ano99-212]. Rusty [Ano97-76]. Ruth [Ano98-105]. Ruzicka [Ano97-165]. RWS [Ano95-27]. Rytter [Ano98-55].

S [Ano93t, Ano93q, Ano95-70, Ano95w, Ano96-35, Ano96-75, Ano96-109, Ano96-107, Ano96-137, Ano96-149, Ano97g, Ano97p, Ano97-27, Ano97-47, Ano97-149, Ano97-150, Ano97-158, Ano97-130, Ano98j, Ano98-38, Ano98-61, Ano98-67, Ano98-92, Ano98-117, Ano98-114, Ano98-116, Ano98-153, Ano98-145, Ano98-146, Ano99x, Ano99q, Ano99-92, Ano99-93, Ano99-96, Ano99-97, Ano99-133, Ano99-134, Ano99-222, Ano99-223, Ano99-224, Ano99-234, Ano99-91, Voi90, Kim92, KYO92, Kim92]. S-system [Voi90]. Saad [Ano98-111]. Saake [Ano99-145]. Saaty [Ano95-27]. Sabidussi [Ano97-56, Ano98-66]. Sabot [Ano96-86]. Sachs [Ano96-76]. Saddle [Ano96-102, Ano97-113, Ano97-114, Ano99-185, GK91]. Sadiq [Ano98-41]. Sadri [Ano99-149]. safe [NMUY92]. safety [Ano95-64, Ano96f, Ano98-138]. Safety-critical [Ano98-138]. safety-related [Ano96f]. Safeware [Ano95-64]. Sagdeo [Ano99-213]. Sagonas [Ano96-149]. Saint [Ano99t, Ano99-152]. Saint-Étienne [Ano99-152]. Saint-Jean [Ano99t]. Salas [Ano98r]. sales [KS97a]. salesman [TE93]. saliency [SB97]. salmon [VA94]. Salomaa [Ano99-66, Ano99-67]. Salomon [Ano99-46]. Saltz [Ano93w]. Sam [Ano99-173]. Samarati [Ano96-48]. sample [Ahm94, KS91]. samples [Ahm95, AV98, Ter00, ZM94]. sampling [AW94, De 96, GS95]. Samuel [Ano96q, Ano98-47]. Samuelson [Ano99-75, Ano99-76]. San [Ano96-135, Leo96]. Sándor [Ano99-165]. Sandra [Ano99-114]. Sandro [Ano96e]. sandwich [Ano99-152, CF91c]. sandwiched [RD96a]. Sang [Ano99-57]. Sangalli [Ano99-215]. Sanjay [Ano97-44]. Sanjiv [Ano98-119]. Sankatha [Ano98-59]. Sanov [BSP94]. Saphir [Ano99-160, Ano99-161]. Sappington [Ano97-37]. Sara [Ano99q]. Sarafyan [CT96]. Saraswat [Ano96-134]. Sartre [Ano97-127]. SAS [Ano99-28]. SATAN [Ano98-129]. satellite [KU99, SSN94]. Satisfaction [Kar94]. satisfying [GBB91]. saturated [FT97a, KMD99]. saturation [ZA95]. Saul [Ano90w, Ano97-45]. Saulyev [EGLN95]. Saupe [Ano99-89]. saving [Shv95]. Sawyer [Ano95g]. Saxl [Ano99u]. Saying [NF92]. scalable [Ano93w, Yas98b, Ano98-139]. scalar [CE94, Fab95, Gla93e, Goh94, GHS91, Isp98, KL96a, Oza94, Prn97]. scalar-valued [Isp98]. scale [GRKM96, GL94, LT99, Pet92, SM98, Sie90, TNS98, Ano95-41]. scales [LK94, LS98a, ZKÖ98]. Scaling [BFR94, Ano98-62, DKSI94, SV91a, SV91b]. scattered [Fol92, Kan90a, Kan90b, Kan92a, QA96, Sun92]. scattered-cardinal [Sun92]. Scattering [RR97a, Dem94b, MRZ96, MW99, MK98, MS99, Pok97, RW91, RSV98, UW94, UWT99, YP98a, YP98b, ZR91]. scenario [RA92b]. Schabes [Ano98-58]. Schauble [Ano96o, Ano98-109]. Schauder [Cie95]. schedules [ZK99]. Scheduling [BLR96, LS96b, LC94, AK93, AGM99, Ano97-84, Ano98-68, Ano99-58, Ano99-131, BTED97, Cha94, CCK97, Cha99a, CD98, HY98, HC98, Seb91, Ste90]. schema [KHC92]. Scheme [Car93, DK97c, Ber92, Che93a, CL98, CCVV90, CC98b, CS90, EGLN95, FLM94, GH90, Gla91a, Gla91c, Gla91e, Gla92b, Gla94d, Gla94c, Gla94e, Gla95b, Gla95d, Gla95f, Hwa99, JW93, Kan90a, Kan90b, Kan92a, KP95b, KS91, LH97, LL92, LWL93, LM91, PS93a, Prn97, QZ93b, Shk97, Sir99, Tah91, Tan93b, Tan94, TL94, Wen99, WHL^{+}99, WCY94, YL93b, YL99]. schemer [Ano96-154, Ano96-160]. Schemes [QZ93a, AAII92, Ano96-41, Bab98, Ber96, Car92, CHSS95, DH91, DKG98, Far99, Gla93e, Gla95c, GPCGV92, GPGCM94, HP94b, Hua95, Kwo92, Los89b, McL95, MZMQ90, MR94, ND95, QZ93c, RW98, Sha90a, ST95, SS96, SS98b, SS99, Tan93c, VA97, ZQ93a, ZQ93b, ZQ93c, ZQ93d, ZQ94, ZQ95]. Scherf [Ano97q]. Scheurer [Ano95-34]. Schieder [Ano98-102]. Schluter [Ano95-54]. Schmidt [Ano98-38, MNB93, Tse95]. Schmitt [Ano97-131]. Schoenberg [Rab92]. schools [Niw96]. Schreiber [Ano99x]. Schrijver [Ano98-29, Ano98-154]. Schrödinger [Ano98-165, Dat90, AI97, AS96b, BSP94, Lee94, RV94, Rob97, Sha90b, SM94, SM95, Sim97, Sim98, Tah91, TVZPG96, TG96]. Schrod̈inger-like [Dat90]. Schrödinger’s [Ara95]. Schulze [Ano96-120]. Schur [EG98, HV92]. Schwartz [Ano97-101, Ano98-97]. Schwarz [Has97, Sir99, ZZT99]. Schwarzkopf [Ano98-33]. Schwichtenberg [Ano97p, Ano98-93]. Science [Ano98-74, Ano90p, Ano93n, Ano95p, Ano96-89, Ano96-105, Ano96-142, Ano97x, Ano97y, Ano97z, Ano97-63, Ano97-146, Ano97-107, Ano98-118, Ano98-125, Ano99-51, Ano99-74, Ano99-225, Wit96b, Ano95-74, Ano96-34, Ano96-167, Ano97-28, Ano97-45, Ano97-120, Ano97-129, Ano98n, Ano98-64, Ano99-111, Ano99-112, Ano99-224, Ano96-141, Ano96-142, Ano97-128, Ano99-202]. Sciences [Ano98-93, Ano97-151, Ano97-153, Ano99h, Ano99i, Yam94]. Scientific [Ano97-119, Ano90w, Ano93w, Ano95j, Ano96g, Ano96o, Ano96-137, Ano96-165, Ano97-65, Ano97-73, Ano97-158, Ano97-155, Ano97-154, Ano98-48, Ano98-80, Ano99-78, Ano99-79, Ano99-239, ALL^{+}96, Ano96-144, Ano96-143]. scientists [Ano95-48, Ano96-112, Ano99-47, Ano99-48, Ano95g, Ano97-93]. scoring [Sch92b]. Scott [Ano97-43, Ano97-77, Ano97-102, Ano98-88]. Scozzafava [Ano96-113]. screws [LY92]. sea [Kra90]. sea/shore [Kra90]. search [Ano97-61, AC93b, ANV94, BT94, CLD97, Dei91, Huj96, KVCA^{+}99, LV92, Man96, MN93, Paz90, SFS95, Yav92, DdW93]. searches [CG95]. Searchlight [WU94, UW91, UWT99]. seasoned [Ano96-160]. Sebastopol [Ano95-76, Ano95-52, Ano95-46, Ano95-80, Ano95-58, Ano95-63, Ano96w, Ano96u, Ano96-42, Ano96-28, Ano96-49, Ano96-51, Ano96-70, Ano96-59, Ano96-88, Ano96-92, Ano96-93, Ano96-133, Ano96-166, Ano97q, Ano97-38, Ano97-35, Ano97-75, Ano97-77, Ano97-76, Ano97-74, Ano97-78, Ano97-89, Ano97-101, Ano97-102, Ano97-103, Ano97-123, Ano97-124, Ano97-131, Ano97-166, Ano97-169, Ano97-167, Ano97-168, Ano98t, Ano98-52, Ano98-54, Ano98-51, Ano98-79, Ano98-88, Ano98-86, Ano98-85, Ano98-99, Ano98-97, Ano98-98, Ano98-103, Ano98-120, Ano98-129, Ano98-122, Ano98-130, Ano98-163, Ano98-162, Ano99y, Ano99-31, Ano99m, Ano99-35, Ano99-36, Ano99-61, Ano99-62, Ano99-72, Ano99-139, Ano99-138, Ano99-143, Ano99-141, Ano99-164, Ano99-148, Ano99-147, Ano99-221, Ano99-173, Ano99-176, Ano99-177, Ano99-178, Ano99-179, Ano99-180, Ano99-195, Ano99-187, Ano99-188, Ano99-192, Ano99-230, Ano99-210]. Sebastopol [Ano99-235, Ano99-236, Ano99-238, Ano99-237]. Second [Ano93n, Ano96-109, Ano97-86, Ano98-46, Gla93e, MMP98, Prn97, AO99f, AO93b, AO93a, Ano95-71, Ano95-32, Ano95f, Ano95-54, Ano96-45, Ano96q, Ano96-102, Ano96-133, Ano97x, Ano97y, Ano97z, Ano97-71, Ano97j, Ano97-95, Ano97-137, Ano98-35, Ano98-30, Ano98-54, Ano98-107, Ano98-104, Ano98-141, Ano99y, Ano99-28, Ano99m, Ano99-108, Ano99-158, Ano99-159, Ano99-150, Ano99-179, Ano99-187, Ano99-237, Arg99, BL96, BTT90, Car92, CP98a, Cha91, CJS90, CE94, CS93b, Eso90, GS91, GL93, Isp98, JV96, JPF96, Jon96, Ker91b, KAS99, Köh93b, KS97b, LY98, LF99, LX98, LG99, Mak90, MT92, MK93, PDK96, PXHG99, PR96, QZ91, Sim93a, Söv99, Tha94, TP98, TMGS98, TR99, Trz92, Tse95, VK95, Wan98b, WJ99, ZL98b, ZZ99c]. Second-order [MMP98, Prn97, AO99f, Car92, CP98a, Isp98, JV96, LY98, LF99, LG99, Mak90, MK93, PXHG99, PR96, Sim93a, Söv99, TMGS98, TR99, Wan98b, WJ99, ZL98b, ZZ99c]. Secret [SS98b, SS96, SS99, Ano99-203]. secretary [Ano92a, Bec90]. section [Ano90-32, Ano90-33, Ano90-34, Ano90-35, Ano90-36, Ano90-37, Ano90-38, Ano90-39, Ano90-40, Ano90-41, Ano90-42, Ano90-43, Ano90-44, Ano91o, Rod90]. sectionally [Ioa91]. sections [RD96b]. security [Ano96-49, Ano97-104, Ano97-167, Ano96-48, Ano96-133, Ano98-88]. seeker [FY93]. seemingly [Sch92b]. Segal [Ano96-95]. Seger [Ano96-42]. Segev [Ano98-104]. segment [IM93]. segmentation [KH92, Pan93a]. segmented [Mei97]. segments [Oll90]. Segrè [Ano96-63]. Seidler [Ano98-72]. Seismic [OOWD98]. seismology [Sym91]. Sejnowski [Ano93t, Ano99-232]. Selecta [Ano98-62]. Selected [Ano97-51, Ano97-92, Ano99-172, Fog94, Ano95-65, Ano97-129, Ano97-130]. Selecting [Ano96-39]. selection [Ano99-80, Gün99, GP91, YF92a, YF92b, Ano99-75, Ano99-76]. selections [Ano99-49]. Selective [DHV91]. Self [Ban93, AO93a, Ano97-43, Ano99-162, AS95, AS98, Che99b, Dor98a, Dor98b, Il’97, Kre91, MAM91, NIS97a, Tse95, ZQ93b]. self-adjoint [Che99b, Kre91, Tse95, ZQ93b]. self-affinity [Ano99-162]. self-consistent [Dor98a, Dor98b, MAM91]. self-equilibrating [AO93a]. self-organization [Ano97-43]. self-organizing [NIS97a]. Self-replicating [Ban93]. self-stabilizing [AS95, AS98]. selfadjoint [AI97]. selling [Nak99]. Selverston [Ano98-114]. Semantic [Joh95, Leh92, Ano96-95, Ano97-144, BV99, EHK92, GOWR^{+}93, Hau92, Hen92, Lev92, MNE92, Mar92, Sha92]. Semantics [Ano99-94, Ano95-81, Ano96k, Ano96-161, Ano99-87, Ano99-142, WF92, Ano96-46]. semblance [Sym91]. Semenoff [Ano99-186]. Semenov [Ano99-49]. Semi [AD93, Wan91, Ano97-140, Ano98-115, CC94a, CP95b, CK94b, FW96, Fli95, GN99, JG95, JG96, Kor90b, Mat99, Mav93, Ruz90, SWF95, WF96, WK99]. semi-analytical [CP95b]. Semi-discrete [Wan91, Fli95]. semi-hereditary [Ano98-115]. semi-implicit [CC94a]. semi-infinite [Ano97-140, CK94b, FW96, JG95, JG96, Mat99, SWF95, WF96, WK99]. semi-Lagrangian [GN99]. semi-Markov [Kor90b]. semi-Markovian [Ruz90]. Semi-regenerative [AD93]. semi-synchronous [Mav93]. semianalytical [PC96a]. semicircle [CFM91]. semiclassical [Gre93a]. semiconductor [Bed95a, Che90a, Che93b, Che96]. semiconductors [Nas95, Szm90]. semicontinuous [Cub95b]. semicycles [ZZ99a]. Semidiscrete [Che92, DNJS98]. Semidistributive [Ano99-204]. Semigroups [XL97, Ano99-82, ZA95]. Semilattices [Pro95]. semilinear [AG94, CLM99, Mak90, Sir97, Wan99]. seminar [Ano98-155]. semiotic [Cho96]. semirigid [YFF92a]. semiseparable [EG97, EG98]. Semmler [Ano97-42]. sending [Ano97-141]. Sense [Sla92, Ano97-134]. Sensitivity [Gan92, BS94, CHSS95, CP95b, Ege91, KT90, PC96a]. Sensory [Ano97-149]. Sentient [Nam96b]. separable [RT99a]. separation [Kis91, PJ91]. September [Ano98-155, Ano99-120, Ano99-137, Ano99-212, Ara99b]. sequence [CLRN94, Lei96, MSR95, RJ94]. Sequences [Sya97, Ban93, BD96, Cla98, DB96, EZ93, HSH90b, JGB95, LL90b, Med90, TXC98, WCL99, Yos90]. sequencing [Che90b, May99, ML90a]. Sequential [GSM91, GS99, Pio94, AW94, Sou90]. Serge [Ano95-35]. Serial [KKOS96, JM91]. serializability [BQN97]. Series [Ano95-44, Ano98-93, AR92c, AR92d, Ano95-27, BS99a, CC94b, DTLM98, EDOS93, HGG99, Il’97, Isp98, JKS94, KS96, Pas97, RSS94b, Tse95, VA94, Wan96, ZQ93c]. series-part [AR92d]. Server [YL93a, TBA99]. Server-aided [YL93a]. service [AD93, Ano97-161, EA92, Kat92, YI97]. services [Ano95-46]. servlet [Ano99-139]. set [Ano95-34, Ano97x, Ano97y, Ano97z, Ano99z, Ano99-27, Bel97a, BH91, Din99b, GW99, GF94a, HLTB99, Kai94, Kai96, Kou89, Kou90b, KLMY97, LW99a, Lin91b, Mol99, Nam99b, O’R98b, Pan98c, Rhe96, Ano99-105, Ano99-103, Ano99-102, Ano99-104, Ano98-140]. set-valued [Din99b, HLTB99]. Sets [BL97, Win91, Bel94a, Bel97b, CL94c, Com92, CY94, DH91, Dev93, KB90, LS98a, LWL96, NSS92, SH95a, WCHL96, Yos99]. seven [Boy99]. Seventh [Ano99t]. Several [KT98, AR92b, HS95, SH95b, YF92b, Ano95-66]. sextet [BV99]. sFr [Ano97-140, Ano97m, Ano97w, Ano97t, Ano97-33, Ano97-34, Ano97-46, Ano97-51, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97j, Ano97-86, Ano97-92, Ano97-96, Ano97-147, Ano97-111, Ano97-112, Ano97-122, Ano97-152, Ano97-138, Ano98-44, Ano98-43, Ano98-91, Ano98-125, Ano98-62, Ano99v, Ano99w, Ano99x, Ano99-41, Ano99-42, Ano99-44, Ano99-43, Ano99-46, Ano99-66, Ano99-55, Ano99-67, Ano99-56, Ano99-69, Ano99-60, Ano99-78, Ano99-79, Ano99-77, Ano99-70, Ano99-89, Ano99-82, Ano99-109, Ano99-110, Ano99-108, Ano99-111, Ano99-112, Ano99-125, Ano99-129, Ano99-126, Ano99-130, Ano99-132, Ano99-218, Ano99-156, Ano99-219, Ano99-157, Ano99-149, Ano99-162, Ano99-193, Ano99-194, Ano99-186, Ano99-189, Ano99-198, Ano99-208]. Shadow [Kra90]. Shafer [LZ92]. Shaking [Ira93]. shallow [CC94a, GT97a, Gla94d, Gla95b, Gla96b, Gla97b]. Shammas [Ano96-112].

Shampine [Ano96-75]. Shanahan [Ano97-134]. Shanjie [Ano96-37]. Shankar [Ano97-94]. Shannon [GS95]. Shape [He95b, KM97c, Ano97-143, Bed95a, BF95, BRGJ95, Bri97, CP95a, HL99, KC92a, Luo93a, xLhW96, MRZ96, PZ99a]. shape-preserving [CP95a]. shape-shifting [HL99]. shaped [KYO92]. shapes [Ano97-146, Bak92, LCL94]. Shapin [Ano97-155]. Shaping [Ano95o]. Shapiro [Ano96h, Ano97-103]. share [HSH90b]. shared [Mav93, hCPW98]. sharing [SS96, SS98b, SS99]. Sharp [EGLN95, WA94b, WA96c, YX95]. Shavkat [Ano98-89]. Shavlik [Ano96-39]. Shaw [Ano99-93, HR99, Hun95]. sheaves [Ano99-105, Ano99-103, Ano99-102, Ano99-104]. sheet [Ara97, CH91]. Sheetanshu [Ano99-87]. sheets [Kha97]. Sheila [Ano96-142]. Sheizaf [Ano98-112]. Shelley [Ano99-61]. Sheppard [Ano99-211]. Sherali [Ano99l]. Sherr [Ano90t]. Sherwin [Ano99-205]. Shi [Ano99-86]. Shi-Yu [Ano99-86]. shift [Ano99-245, BG91b, FLW97, SRCB98]. shift-inverted [Ano99-245, FLW97]. shifted [Zha93d]. shifting [HL99]. Shih [Ano98-161]. Shih-Fu [Ano98-161]. Shimamura [Ano96-115]. Shimon [Ano96-84]. Ship [YZM95, YZM94, YFZM95a, YFZM95b]. Shishir [Ano96-28]. Shively [Ano95y]. Shlomo [Ano99-208]. Shock [Gla95e, Fre98, Gla90b, Gla94e, Gla95d, RSP94, RS94, Sak90]. shock-reflection [Gla90b]. Shockwave [Ano97-131]. Shooting [MM93a, DELE99, Rán92]. shop [Ste90]. Shor [Ano99-167, Ano99-168]. Shore [Ano97-86, Kra90]. Shortening [KS95]. shortest [Che99a, Kun95]. Shouchuan [Ano98-67]. Shrinking [Jab96]. Shub [Ano97-51]. shuffle [PC92]. shuffle-exchange [PC92]. Shukhrat [Ano98-89]. Shurman [Ano97-60]. shutdowns [HC98]. sided [GS96b]. sides [Luo98]. Siek [Ano96-120]. sigmoidal [FT97a]. sign [CL94b, HP94c]. Signal [Ano98-43, AÉ96, Ano98l, Ano98-137, IIP^{+}97, RGBL96]. signalized [Alf95, Ano96-171]. Signed [BL97]. signs [DJM98]. SIGS [Ano98-158]. Silva [Ano96-149]. Silvana [Ano96-48]. Silvano [Ano98m]. SIMD [PC92]. Similarity [Gla92e, Pas92, WW95, CHW93, Gla90b, HSH90b, KM96, LZ94, Pap94b]. simile [Mar90]. Simo [Ano99-41, Ano99-42]. Simon [Ano96-119, Ano97-153]. Simple [ST97, Ano96-155, Ano98s, Ano99-226, ANV94, BW92, BG91a, HSH90a, Mol96, Sto92, Sug94, ZK99]. simpler [CC97]. simplest [Gre98b, LKC95]. simplex [Pan98a, Pan98b]. simplexes [He95b]. simplicial [BR94, Bro98]. simplicity [Ano97-108]. simplified [WW93]. Simsa [Ano95-33]. Simson [Ano95-58, Ano96-133, Ano97-167]. Simulated [BT94, Fab97, Tan92, Fox94, Mar90, RI94, RD96b, Tso91]. Simulating [MN93, GB92, HM95]. Simulation [GHPLZ90, Str98, Ano95k, Ano95-36, Ano95-60, Ano96-167, Ano97-53, Ano99-90, Ano99-242, DKSI94, Gre90b, Gre98b, Hon99, HAO91, IGM98, KU99, KG97, Kor92, Kou98, LSH98, MATA98, MRWM94, Mic90, NAKA98, Nke94, Pad92b, Rei93b, Sán98, SJB97, SHT^{+}90, SCG^{+}97, SC98b, TNS98, US91, WHL^{+}99, ZSDB99, Ano95-51, Ano97-132]. simulations [Ano97-96, CCC^{+}98, Cor94, DC94, Kro99, RD99, SM98, Sie90]. simulator [Lew94]. Simultaneous [AZ96, OB97, SH95a, AV98, BL96, Car94, LW96, LW99a, PM91, PT96, PIT97, PPH98]. Sinc [LTB96, BR95c]. Sine [Ado91c]. Sine-Gordon [Ado91c]. Singapore [Ano93p, Ano97-119, Ano99-78, Ano99-79, Ano99-189]. Singh [Ano98-59]. Single [GS96b, Mar90, PM95b, BC98, Cha99a, Che90b, CD98, Coo97, Kun95, LT94, NNB97, Szt92, YCC96]. single- [NNB97]. Single-interval [Mar90]. Single-layer [GS96b]. single-machine [Che90b]. single-sided [GS96b]. single-species [YCC96]. Singular [O’R98c, PC92, AO99c, Ber93, BSP94, BG91a, CJS90, DR97, EK94, Feu91, GS93, HK98, JR94, Kar96a, Kim98, Kim99, Li99b, MT92, Mil98, Ola90, RSR98, Roj99, Rok90, SZ91, WL96, Zha91]. singularities [Ano99-136, KL95]. singularity [AS97, BS93, Ram93]. singularly [BS96b, BS96c, Bog99, EY91, GPCGV92, HG98, Li98b, LN98, LN99, Li10, MP91, MK93, Sir97, Sir99]. Sinha [Ano98-142]. Sipper [Ano97-46]. site [HSH90a, HSH90b]. site-specific [HSH90a, HSH90b]. sites [HSH90a, LNM90]. Sivashinsky [MMP98]. six [Ano97-156, Kwo92]. six-point [Kwo92]. sixth [Ano99-29, Ano99-30, Wu98]. sixth-order [Wu98]. size [CT96, Dor99, Niw96]. Sized [GH97]. sizes [CC99, MR94]. sizing [ZG97]. Sjoerd [Ano95u]. sketch [Ano92b]. Skew [Has95]. Skjellum [Ano95-82]. slab [GMR90, UW91]. slackness [BJH99]. SLAM [Ano97-132]. Slaman [Ano96-35]. Slator [Ano96-58]. Slaves [Ano97-133]. sleepers [LR96a]. Slide [Jar97, FG98]. sliding [Koç95]. slip [Lay99]. slot [HW90]. Slots [NF92]. slow [Hak99]. slowly [SS94]. slowly-varying [SS94]. Small [Ahm94, GH97, Ano97-83, Ano97-79, CS93b, GF90, Li98b, MK91, ZZ99c]. smallest [GF94b, Roj99]. smart [CH93]. Smilonich [Ano95-72]. Smith [Ano95-29, Ano96f, Ano96-47, Ano96c, Ano96-62, Ano96-130, Ano97-41, Ano98-61, Ano99-225]. Smooth [Qu96, QA96, Hua98c, Luo98, RS99c, Say99, YL99]. smoothed [HPH^{+}94]. smoothed-particle [HPH^{+}94]. smoothing [CL99c, Fli94a]. Smoothness [Hen94, Bad95, GZ95]. Smyth [Ano96h].

SNePS [SR92]. Snider [Ano99-185]. Snir [Ano96-118, Ano99-158, Ano99-160, Ano99-159, Ano99-161]. snRNA [LNM90]. Snyder [Ano98-60, Ano99k]. Soares [Ano98-127]. Sobolev [BK98]. sociable [Ano97-119]. Social [Ano97-63, Ano95-27, Ano95-37, Ano97-107, Ano98-63, BJ90, Yam94]. societies [Ano97-63]. Society [Ano96-127, Ano96-105, Ano96-108, Ano99h, Ano99i]. socio [Den91a]. socio-spatial [Den91a]. sociological [Ano96-143]. Soft [Mol99]. Software [Ano90-32, Ano90-33, Ano90-34, Ano90-35, Ano90-36, Ano90-37, Ano90-38, Ano90-39, Ano90-40, Ano90-41, Ano90-42, Ano90-43, Ano90-44, Ano91o, Ano93-32, Ano93-33, Ano93-34, Ano94c, Ano98-162, Con90, KYO92, Log96, Rod90, Ano95e, Ano96-150, Ano96e, Ano96f, Ano96-27, Ano96-151, Ano96-90, Ano98-37, Ano99-154, Ano99-155, Mis90, YOO92, dSB96, tC95, Ano95-77, Ano97-124, Ano99-28, Ano93r, Ano98-69]. soil [CKK97a]. sojourn [Toy92]. Sojourns [Leo90]. Sol [Ano90t, Ano98-101]. solid [Kor97]. solids [Ano96-83, Ano99-113]. solitary [Pom92]. soliton [Bas98, DK97c, GT95, GT97b]. soliton-type [DK97c]. solitons [HAO91]. Solla [Ano99q]. Solso [Ano97s]. Soltan [Ano99-101]. Solution [Ado94, Ado96, BKK99, BR95c, BS96c, Bon97, Car93, Chi91, Cse95, GZ93, LW91, LNG92, Mur92b, PSS99, Sni91b, AO91, AR92c, Amo93, Ano98-165, Ano99-181, AS96b, BLP95, BMB94, BS93, BS95, BS96b, CC97, CS96, Cas94, CP96, CAZAS95, CKAS96, CCVV90, CL90, Cor95, Doh90, EDD91, Ech95, Ech98, Far96a, FH99, GPZ93, GS91, GGS97, Gás95, Gla90b, GSC98, HL96, HI94, HBB90, Hun95, IESAH95, IB95, Ita93, JG95, JG96, JA97, KO96, Kha97, Kij92, Koç92, KSM94, Lar90, LA96, LYL95, LP99, LD93, Lop93, LS95, Maz90, MM96c, MHBD96, Moo94, MH93b, MM95d, NPJ95, NFJ96, Ola90, Pan94, PS96, PC96b, Ram90a, Ram91a, Ram93, RL91, RSP94, Rei98, Rei99b, Riz95, Rob97, RSV98, SGS93, Sar94, SVS98, Say99, Sch99, Sch94]. solution [SD95, SS91a, SS94, Sha99a, Sha95, Sha96b, Sim93b, SM95, Sim97, Sir97, Sir99, Sjo92, Tör94, VR95b, VA97, VL92, Wan96, WL96, YCC96, Zha98, ZC99, ZR91, Ano99-37, Ano99-216, Ano99-217]. solutions [AR90a, AR92a, AWY97, AHY98, AH98, AO99c, AO99f, AO99e, Ame97a, Ame97b, Ame98, AHL98, Ano97-114, Ano97-165, Ano99-181, Ano99-245, AP98b, BM96, Bal99, BGL97, Blo91, CTY99, Cha99b, CHW93, CYS94, CL96a, CZ94, CL99d, CK91, CG94, Cub95b, DH98b, DH98a, Dat90, Din97a, Din99a, DK97c, Dor90, DHV91, Duc95b, ESKG98, FLW97, Fil93, GT95, GGK97, Gla92e, Hen94, HH98a, Hua91, HW90, JP95a, JP95b, JA96, JV96, JD97, JS97b, KV94a, Kel94, Ker91a, KESH90, KAD99, Kim99, KAS99, KP99b, Kos90, Lee93, LC98a, LG99, LD95, MG91, MC99, Med97, Mic94, Mic98, MT98, O’R98a, PR90, PS96, Pas92, PXHG99, Pin98, Pop94, QA98, RS99c, RZ95, Sak90, Sar90, SSZ91, SWY96a, SLC97, SFY92, Sol90, Szm90, TC99, TP98, TMGS98, TR99, Tre94, Tre95a, Tse95]. solutions [Tsi99, VHH95, WHC90, WW95, Wan99, WJ99, WA96a, Won98a, Won98b, Wu99b, YTLK94, Zaf98a, ZT98a, ZZ99a, Kan90b, Ano99-185]. solve [CLD97, GT97b, SM90, VR95a, ZQ93a]. solved [CCVV90]. solvents [Den91b]. Solver [PS92, CTY95, Gla91d, Gla92a, Gla93a, Gla94f, Gla95a, Gla95c, Gla96c, Gla97a, Gla98, GL96, LJ98, OH93, Rac99, SFT99, Wen99, vDHO^{+}95]. solvers [Ano99-70, CS96, CT96, FS96, Gla94a, Gla99, GRKM96, LVV96, PZDH95, SM98, SBS99, TAD91, dSB96]. Solving [Ado91d, Ara99a, CF98, Chu92, CY93, FW96, GM96, HM91, Kim98, MT94a, Mur91b, Nov90, Rhe97, RG92, SZ91, Sug94, Tor90, WF99b, Alt90a, Ano95u, Ano96-86, Ano96-110, Ano97-73, Ano98-80, Ano99l, BMR96, Bed95b, BGL90, BG95, BRZS97, CC94b, CLY91, CY96, DL99, EGEHSEK95, ESIP98, Far91a, Far91b, HNS^{+}90, HG98, Huj96, JLDL92, KT92, KK91, Lin96, MS93, Mar94, Nyí99, Oza94, RF91, Roj90, Sha90c, Sol97, SCBFC93, SCBFC98, SB99, SBS99, SAS99, TCS93, Ung90, Ung91, WT93, Wan93b, WSDT94, WK99, Zha91, Zha93c, Zha93b, Zha93d, ZS97, pri94, Ano96-145, Ano97-134]. Some [AH92, BW92, BQN97, CS94a, EGK96, Erm92, FL96, FO91, Fun97, Gal99c, GKT99, Gir92, GH97, Gyö90b, Huj99, IK97, JKS94, KSM94, KC91, MCG98c, Mei97, OPD97, Oza94, Pac94, PC93, RS99a, RSS94b, Riv97, RM98b, RHG91, Sez95, Sha96a, SSH92, SA95, SG95b, Sri95, TM91, Ush94b, Wag90, Zal95, ZA95, dOF91, AD93, Ado91a, Ado91b, AD98, BLR94, CS96, Cas98, CLM92, CR90c, DL99, DW97, Ech95, Ech98, Eso99a, KK97a, Kov95, KYNY99, Lee93, LOS99, MN97, MB94a, Mil98, Moi97, Pas97, RL91, SHA99b, Szu96, YF93, ZR91, hCPW98]. Son [Ano99-57]. sonar [Ano96-155, MM96d]. sonar-equipped [Ano96-155]. Song [Ano97-119]. Sonine [GKT99]. Sons [Ano93p, Ano95u, Ano95c, Ano95-33, Ano95-39, Ano96g, Ano96-150, Ano96y, Ano96-37, Ano96-27, Ano96-45, Ano96z, Ano96-33, Ano96-41, Ano96-38, Ano96-54, Ano96-75, Ano96-76, Ano96-77, Ano96-91, Ano96-101, Ano96-109, Ano96-116, Ano96-117, Ano96-110, Ano96-124, Ano96-126, Ano96-131, Ano96-146, Ano96-147, Ano96-164, Ano96d, Ano96-167, Ano97o, Ano97-27, Ano97-30, Ano97-54, Ano97-61, Ano97-60, Ano97-72, Ano97-145, Ano97-87, Ano97-106, Ano97-121, Ano97-125, Ano98m, Ano98p, Ano98-29, Ano98-30, Ano98-42, Ano98-64, Ano98-78, Ano98-81, Ano98-117, Ano98-141, Ano98-142, Ano98-143, Ano98-154, Ano98-159, Ano99-28, Ano99-34, Ano99-47, Ano99-48, Ano99-45, Ano99-53, Ano99-54, Ano99-65, Ano99-71, Ano99-74, Ano99-98, Ano99-133, Ano99-134, Ano99-127, Ano99j, Ano99-151, Ano99-175, Ano99-190, Ano99-191, Ano99-199, Ano99-200]. Sonsg [Ano97-132]. Sonsg/Systems [Ano97-132]. SOR [KP96]. Soren [Ano98u]. Sorgatz [Ano96-120, Ano99-69]. sort [Coh92]. Sossinsky [Ano98-92]. sound [KW98, Pok97, SK91]. sounds [Wit96b]. source [Ano99-173, Cha96, Gla93c, Gla93e, Gup94, Mur92a]. Sources [Ano99-185, DJV96, ZFM96]. Soylmar [Ano90u]. Space [Ano96-30, LT99, Shv95, Ahm91, Ano96-130, Ano99-144, ADKM97, Arg98, AMZ98, BL94, BTT90, CS96, CP98b, CR90b, Din97a, Dor90, Duc95a, GMR90, Har92c, Hau92, HAO91, KU99, Köh95, Köh96a, Köh96b, LG99, MG90c, MH93b, Prn97, RC92, Sya97, Ter91, VA94, WW91, Yav91a, Ano97-81]. Space-decomposition [LT99]. space-time [HAO91]. space/descriptor [Tra97]. spaces [Ame97a, Bad95, BAS97, BL95, CR90c, DO91, Din96, Din97b, Din99c, Fin99, Hua98c, HB99, JY93, Köh93a, Köh96b, Köh96c, Lei96, MSR95, O’R95, O’R99, Sza91, Ver99b, WY99, Yao92, Ano99-196, Ano97-135]. Spafford [Ano96-133]. Spainhour [Ano99-237]. span [AG91]. Spannable [LW97]. Spanning [GH97, AS95, AS98, CCH97, DKW98]. Sparing [DH97]. Sparse [CN94, Ano99-245, BW96a, BSW96, FLW97, MRWM94, NWB92, Rei98, Rei99b, RG90c, dSB96, vDHO^{+}95]. sparsity [Luo93a, Luo93d]. Spatial [Ano97-144, DK93, Mai96b, Ano99-38, Den91a, Kan92a, KKK^{+}97, Mil90]. spatiotemporal [Mai96b]. special [Ano96-37, CY96, DW97, Ike91, Pas97, Sim93a, Ano96-146]. species [LL96, PC94b, YCC96, ZC96]. specific [HSH90a, HSH90b]. Specification [Ben99, Ano99-117, tC95, Ano97-136]. specified [GSM91]. specimen [MR91]. Spector [Ano99p]. spectra [BG91a]. Spectral [Ano00, BLR94, LS99, Moi97, RT96, AI97, BSS98, Bla95, Gur91, Gus97, IK97, JKK98, KT97, KP98, KS98a, KLMP97, PW97, RT99a, RSR98, Roj99, Wan91, Yan93a, Zal95, Ano99-205]. Spectral/ [Ano99-205]. spectrometer [BFF^{+}97]. spectrum [LR96a, PS93a]. speculations [Ano96-121]. speech [Ano95m, Ano96-32, Ano96-147]. Speed [Ano92k, Ano92l, Ano92m, Ano93-35, Rod91a, Rod91b, Rod92a, Rod92b, Rod92c, Rod92d, Rod92e, Rod92f, Rod92g, Rod93c, Rod93d, Rod93e, Rod93f, Rod93g, Rod93h, Rod93i, Rod93j, Rod93k, Ano98-84, CC96, Kun90, SM90]. Speeding [LK97]. speeds [Sau92]. Spencer [Ano98-99, Ano99-205]. SPH [HL99, LGP98]. sphere [Fol92, Fre98, TSLW98, Xue94]. Spherical [CC95, BR95a, CR90a]. Spherically [BR95a, CR90a]. SPICE2 [SHT^{+}90]. spiral [Byk97]. spliced [GG97]. spline [Ano98-165, EL92, Fab95, Hu95, IESAH95, MMP98, RS95, Rob97, Rou92, RZ94, Say99, Sch96, Söv99, WA94a, WA94b, WA96c]. splines [DR97, GSC98, KM97c, RSS94a, SS98a, Sal92a, Zhu92, Ano99-114]. split [CV90]. Splitting [GRZ96, Li96c, Gla90a, Gla91a, Gla93d, Gla93c, MZ97, MMP98, Sán98]. Splitting-integrating [Li96c]. spontaneous [KKB97]. SPPs [JS97b]. spreads [CL94b]. spreadsheets [Ano93z, Ano96-71]. spring [HH98a]. Springer [Ano93o, Ano96-68, Ano96q, Ano96m, Ano96-106, Ano96-107, Ano96-156, Ano96-157, Ano96-159, Ano96-165, Ano97-140, Ano97m, Ano97w, Ano97t, Ano97-33, Ano97-34, Ano97-46, Ano97-51, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97j, Ano97-86, Ano97-92, Ano97-96, Ano97-147, Ano97-111, Ano97-112, Ano97-122, Ano97-152, Ano97-138, Ano98q, Ano98-33, Ano98-44, Ano98-46, Ano98-75, Ano98-80, Ano98-93, Ano98-91, Ano98-102, Ano98-105, Ano98-125, Ano98-62, Ano99v, Ano99w, Ano99x, Ano99-37, Ano99-41, Ano99-42, Ano99-44, Ano99-43, Ano99-46, Ano99-66, Ano99-55, Ano99-67, Ano99-56, Ano99-69, Ano99-60, Ano99-78, Ano99-79, Ano99-77, Ano99-70, Ano99-82, Ano99-109, Ano99-110, Ano99-111, Ano99-112, Ano99-129, Ano99-130, Ano99-132, Ano99-218, Ano99-156, Ano99-219, Ano99-157, Ano99-149, Ano99-162, Ano99-193, Ano99-194, Ano99-186, Ano99-189, Ano99-198, Ano99-208]. Springer-Verlag [Ano93o, Ano96q, Ano96m, Ano96-106, Ano96-107, Ano96-156, Ano96-157, Ano96-165, Ano97-140, Ano97m, Ano97w, Ano97t, Ano97-33, Ano97-34, Ano97-46, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97j, Ano97-92, Ano97-96, Ano97-147, Ano97-111, Ano97-112, Ano97-122, Ano97-138, Ano98q, Ano98-80, Ano98-93, Ano98-102, Ano98-105, Ano99v, Ano99w, Ano99x, Ano99-41, Ano99-42, Ano99-44, Ano99-43, Ano99-46, Ano99-66, Ano99-55, Ano99-67, Ano99-56, Ano99-69, Ano99-60, Ano99-78, Ano99-79, Ano99-77, Ano99-70, Ano99-89, Ano99-82, Ano99-109, Ano99-110, Ano99-111, Ano99-112, Ano99-129, Ano99-130, Ano99-132, Ano99-218, Ano99-156, Ano99-219, Ano99-157, Ano99-149, Ano99-162, Ano99-193, Ano99-194, Ano99-186, Ano99-189, Ano99-198, Ano99-208]. Sproat [Ano93l]. Sprössig. [Ano97-125]. Spuri [Ano99-58]. SQL [Ano99-178, Ano99-179, Ano99-230]. square [AH97, BA96, Tem93]. squares [BF95, Che92, DK97b, ER98, Fra96, Li96a, Los89b, Sal92a, Ye99, Ano99-65]. squaring [JS98]. Sravastava [Ano98-59]. Srinivasan [Ano99-91]. Srivastava [Ano98-157]. st [Ano98-110]. Stability [Ahm91, And93, Cas94, CC94a, CCVV90, CKRV98, Fab95, GN99, HP94b, JMPS94, KAD99, Kwo92, LS98a, Lar90, Li99a, LC96a, LF98, MB90, PA98, SJB97, SWY98, Zha97, ZZ99b, ZT99, AdMD99, AC93a, AR92e, ADFP93, AMZ98, CL98, Den91a, GSM91, Goh94, GCT98, GL94, HM96, Ker91b, KL96b, KB90, LR96a, LNM90, Lu96, LW99b, MR94, MG90a, MG90b, QZ93b, Ran93, SZ98b, SWY97, Sol90, Som90, SC98a, Tre98, YQM99, Yu98, ZC96, Zha99, vDHO^{+}95]. Stabilization [LL96, Mag97, BKN97, De 99, Flo95, Flo97, WSS92]. stabilized [CAZAS95, FY93, HL99]. stabilizing [Ame99, AAII92, AS95, AS98, HI94, LL94a]. Stable [CS93b, Cho94, BGL97, CV90, CKAS96, Che97, DELE99, Hsu91, MZ97, MHBD96, MG90c, MGM92, MZ95, Tsi91, Wu98, Ano95-68]. Stackelberg [Liu98]. stacks [CCY93]. stage [Coh94, MZMQ90, QZ93b, WC97, YI97]. stage-structure [WC97]. stair [Luo93a]. stair-shape [Luo93a]. Staircase [Sch96]. Stan [Ano95-71, Ano96s, Ano97m]. Standard [Ano97-90, Ano99-55, Ano99-56, ED91a, Ano99-166]. standards [Ano97-104, Ano96-148]. standarization [Ano98-38]. Standford [Ano98-70]. Stanford [Ano97n, Ano97-148, Ano97-117, Ano97-162, Ano97-129, Log96]. Stankovic [Ano99-58]. Stanley [Ano95g, Ano96-40]. star [AAD98, CF94, LS96b, SS93]. Stark [IK97]. starlike [AIS95a]. Starr [Ano96-99]. stars [Ano97-130, Ano98v]. started [Ano95-39, Ano96-76, Ano96-77]. State [Bra92, Tra97, Ame99, Ano97-151, BTT90, CS94b, GSM91, Gre92a, KV94a, MM93a, Med90, Sch90, Ter91, VDLM95, VA94, Yav91a, Ano98-141, Ano98-58]. state-dependent [KV94a]. State-space [Tra97, VA94]. State-space/descriptor [Tra97]. states [Gus97, KKD^{+}97, vdVVdB99b]. static [Liu96a]. station [Pou90b, Tsi91]. Stationarity [Isp98]. stationary [BGL97, BGM97, Dor90, Faz94b, KK94, MM96b, Mei93, Pap94a, RT96]. Statistical [DSKP96, DGU90, IGM98, Nas95, DFMS99, Jab96, KW98, MC93, TZ93, Ano98-142]. statistics [Ano96-71, Ano99-28, Ano99-53, Ano99-54, ED91a, IZ97, Raq98]. Staunstrup [Ano98-69]. Stavros [Ano98-124]. Stavroulakis [Ano98-116]. Steady [LT96, Bab98, CS94b, Ech96, Gla92b, Gla94a, Gla94e, Gla95f, Hak99, HM99, IGM98, MM93a, WW95]. steady-state [MM93a]. Stearns [Ano95-62]. Stede [Ano99-142]. steel [RD96b]. steepest [Alt90c]. Stefaan [Ano98w]. Stefan [Ano99-70, Ano99-166, Sod90]. Stefanescu [Ano99-189]. Stefano [Ano99-82]. Steger [Ano98-91]. Stein [Ano98-114, Ano99-238]. Steiner [TTC96, Ano99-206, Ano99-207]. Steinhardt [Kai94]. Sten [Ano98-114]. stenosed [TNTR91]. Step [IESAH95, vdV95, Abe90, CAZAS95, CSW99, CT96, Lop93, MG97, MR94, SM94, SM95, Sim98, Tan93d, Wu98, hC96a, hCPW98]. Step-parallel [vdV95]. step-sizes [MR94]. Stéphane [Ano96-43]. Stephen [Ano95r, Ano95s, Ano95-76, Ano96-39, Ano96-105, Ano98-145, Ano98-146, Ano99-237]. steplength [GP91]. stepping [Sak90]. Sterling [Ano96-62, Ano96-104]. Stern [Ano96-45, Ano98-143]. Sternberg [Ano99-220, Ano99-208]. steroid [HSH90b]. Steve [Ano96-59, Ano96-118, Ano98-73, Ano99-158, Ano99-159]. Steven [Ano95g, Ano96-47, Ano96-55, Ano96-118, Ano97-155, Ano98-37, Ano99m, Ano99-158, Ano99-160, Ano99-159, Ano99-161, Ano99-165, Ano99-178, Ano99-179]. Stevens [Ano97-156]. Stewart [Ano95-36, Ano96-98, Ano97-53, Ano99-90]. Stieltjes [Lin95]. Stiff [Car93, BACS98, Car92, Cas98, MN91, Wu98, vdHS94, vdV95]. Stinson [Ano95z]. Stochastic [AK93, CJ90, Ahm91, Ano99-174, Ano99-190, Ano99-191, Ano99-197, Ara95, BS91a, Dor90, FZ93, Flo95, Flo97, FYS99, GF97, GP90, Gyö90a, HP94b, KM93, KSM94, KS97b, KS98b, Liu97, Oda91, OSLC98, Pap92, RW98, Sjo92, Sol90, SJS92, Sug94, Ter91, Tör94, VMA99, Yav91b, Yav93, YZM94, YZM95]. StochGran [Hel96]. stock [Bra92, Cha91, DO92]. Stocker [Ano99-111, Ano99-112]. Stoffel [Ano98-133]. Stokes [AF90, BP97, CS94b, Fai98, GP91, HM91, HW96, JKK98, Lay93, Rac99, Sak90, SCBFC93, SCBFC98, Ye99]. Stolorz [Ano98-139]. Stone [Ano99-173, Che97]. stopping [Yos93, Yos96]. Storage [BLR90, IN92]. stories [Ano96-144]. Storøy. [Ano98-123]. Stout [Ano97-115]. straight [BFF^{+}97, Lia95]. strange [Nas92]. Strassen [CDLW95, PC96b]. Strategic [Ano99-89]. strategies [Ano96-155, Ano98-119, Ano98-150, Ano99-74, Bel97a, FYS99, GP93, Gue93, HL90, Sni91b, Yav91a, YF92a, YFZM95a, YFZM95b, Ano99-131]. strategy [Ano98-64, Bai96a, BGM95, DO92, LL94a, LL99, Wos95a]. Strawn [Ano96-153]. stream [Fai98, LS99, ML90c]. streamline [MS96]. Streamlined [Ano95-52]. streamlines [Has98]. Stress [DPD98, vdVVdB99b]. stresses [FCP95]. stretching [CH91]. strict [Tre98]. strictly [Din96, Kan92a, TC99]. stride [CKAS96]. string [FG97, FG98, Ira93, TAD91]. stringer [dJPvZ97, PvZ95]. Strings [Gla90c]. strip [Bal99]. strong [BF97, BKK97, GlZ95, Ves98]. strongly [CF91b, Din97b, HLTB99, TNS98]. Stroock [GP90]. Strothotte [Ano99-44]. Structural [AMZ98, Ano98-127, Ano98-157, KU91, OPD97, RD96b]. structurally [GSST97]. Structure [Ano98-113, Ano97-158, Ano97-157, Ano99-55, Ano99-56, AR99b, Gal99a, Köh96c, Lag91, LWL93, PY94, RB97, RP97, SOI92, Tel91, WC97, WSNS98, Win91, Yas94, Ano97-137]. Structured [Sha92, AH92, Bar91b, KP95b, Pan92a, PL90, Ano98-143]. structures [ABV97, Ano97-33, Ano97-34, Ano98-89, Ano99v, Ano99w, Ano99-33, Ano99-99, Ano99-100, Ano99-152, Ara99a, Den91a, Hsu91, Käm94a, Kup91, She98, SS96, SS98b, Vic90, WW91, Win91, Ano98-57]. Stuart [Ano98-114]. Stuck [Ano96-54, Ano98-42]. student [Ano96-127]. Studies [Ano98-38, Gre92b, LNM90, Shi95a, Gre90c, Ha90, KG95a, MDVK97]. studio [Ano97-131]. Study [Ano96-127, CCH^{+}90, AE90, Ano97-80, AB93, CC92, EB93, FKO92, GRKM96, Gre93a, GGC97, HR99, HNS^{+}90, KM90, LO99, LF91, MBMS92, PB95, SZ98b, She99, Yav92, YZM94, YY98]. studying [Ano99-199, Ano99-200]. Sturm [AHY98, BMD92, CS93a, Con95, Doš98, Hua91, She93]. Stuttgart [Ano96-120, Ano97-136]. style [Ano97x, Ano97y, Ano97z, Bel94b]. subclass [AIS95b, CIS98, RS96a]. subclasses [RS99a, RS99b, SA95]. subdivision [QA96, QA98]. subdomains [LCP98]. Subgraphs [GH97]. subharmonics [BF97]. subject [AS93, AR98, CCY95a, CCY95b, LL90c, SG90]. subjected [MR91]. sublime [Ano96l]. Suboptimal [CSOL96, LTR98, OSLC98, OCS96, RTJ96, Yav93]. subsequent [Tho92]. Subset [Jay90, Li98c]. Subset-equational [Jay90]. subsets [Lin91a]. subsidiary [Lar90]. Subspace [ST94, SB99]. subspaces [Bal96b]. substance [LD95]. substitute [Sch92a]. substructuring [Fli94b, PW97]. Subsumption [BS92]. success [FZ93]. suction [CC94c, WHC90]. Sudweeks [Ano98-112]. Sufficient [Car92, Dor90, Zaf98b, ZL99a]. suitable [KL95]. Sukumar [Ano98r]. sum [Com92, JLL^{+}93, Li98c, Mit90]. sum-of-product [Mit90]. Sumi [Ano98-164]. summation [CP98a, ZZ99c]. Summations [KS96]. Summing [Bar91a]. sums [Ano95-33, POH99]. Sunderam [Ano95-59]. sup [Sto99]. super [GH98, LKC95, Tor90]. super-Halley [GH98]. Supercomputer [Gre90c, Gre90b, Ha90]. superconductivity [Du94]. supercritical [Gla92c, Gla94a, Gla95c, Gla95e]. superficialities [Ano96-87]. Superharmonic [Yos99]. superhighway [Ano96u]. supermultiplicity [AD98]. superradiance [AKK^{+}97, KKG97]. supersonic [Gla92b, Gla95f]. Supersymmetry [Ano99-208]. SUPG [Rac99]. Supplement [Ano98-49]. supply [KS97a]. support [AAL99, Ano98-156, GP90, HL90, LL90c, MC93, Sun92, TDF^{+}92]. supported [LR96a]. sure [KS91, Sol90]. Surface [Nyí99, ZM99, CHW93, Dor99, Gla95e, Hun95, HW90, Joh99, Kan90a, KKB97, Kor97, Mol93, NSD98, QA96, Ram91b]. surface-tension [CHW93, HW90]. surfaces [Ano98-44, Ano98-107, BM91, DK97a, Duc95b, Has91, Mil91, MBMS92, Oli93, Tan93a, Ano95-55]. surfaces-as [Tan93a]. surgery [Ano96-43]. Survey [Bei97, APP96, Ano90-32, Ano90-33, Ano90-34, Ano90-35, Ano90-36, Ano90-37, Ano90-38, Ano90-39, Ano90-40, Ano90-41, Ano90-42, Ano90-43, Ano90-44, Ano91o, BLP95, Rod90]. Survivable [Ano99-209]. Susan [Ano96-77, Ano97q, Ano97-39, Ano98-162]. Susana [Ano96-31]. Sushil [Ano98g]. suspended [Dor98a, Dor98b]. suspension [MP94, Sem93, WT97]. Sussman [Ano97-137]. Sutton [Ano93q]. Suykens [Ano99-169]. Svend [Ano97-32]. Sverre [Ano98-123]. Sweetser [Ano97-135]. Swift [Ano99e, BGL97]. Swiniarski [Ano99-52]. switches [HY98]. switchover [RHG91]. Sydow [Ano99-192]. Syed [Ano97-36].

Sylvester [De 99, JCM96]. Symbolic [Ano97x, AF90, Che91, Ano95j, Ano96r, Bar93, BP94a, GT97a, GT97b, HP96, HBB90, Han92, TM91, Ano97-138]. Symm [Say99]. Symmetric [Haw91, AS98, BR95a, CS93a, CR90a, CY93, ESIP98, FLR95, GGS97, Gla93b, Gla93d, Jha90, Kal90, Luo93d, OH93, Pan90a, Roj90, Zha93c, Zha93d, hC96b]. symmetries [CC95, CCW96, Köh91, Köh93b, Köh94, Köh95, Köh96a, Köh96b, Köh99]. Symmetry [Ber91, IK95, LL90b, Var92, Ano98-66, CLM99, GMR90, Kai96, KC92b, Köh91, Köh93b, Köh95, Köh96a, Köh96b, Köh97, Kou89, Kou90b, Lin91c, Mar91, MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b, Par91b, Sok91, Ano93s]. symmetry-adapted [MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b]. Symplectic [TVZPG96, Ano97-56, MZMQ90, QZ93c, Tan93b, Tan94, TL94, ZQ93a]. symplecticity [Tan93d]. symposium [Ano97-88, Ano98-84, Ano98-83, Ano98-95, Ano99-136, Ano99-137, Ano96-103, Ano98-94]. symptoms [NSS92]. Synchronization [LY99a, Ano97-32]. synchronized [SCB^{+}93]. synchronous [Mav93]. synovial [TCJG91]. Syntactica [Ano96-149]. syntax [Ano96-161, Ano98-151, Ano95-70]. synthesis [Ano95k, Ano98-133, Ano98-144, Hel96, MCG98c]. synthetic [Dre90]. System [Ano95-34, Ano95-64, AS93, AS99, Ano93i, Ano95-42, Ano96v, Ano97-169, Ano98-51, Ano98-117, Ano98-163, Ano99-201, BSW96, BR95b, BGLT90, Boy96, Cha91, CHCC95, Cho92, Chr90, Chu99, FYS99, GGS97, Gla93a, Gla93d, Gla95d, GC99, HBB90, HL90, IN92, JM91, Jos90, KS97a, KS99, Kat92, KKOS96, KK92, Kim98, Lar90, LL94a, LL99, LL96, LPR96, Los89a, Lyt92b, MN97, MC93, Mic90, MM96d, Mit92a, Nyí99, Pan94, Pao98, Par91b, PS93a, Pet92, Pfi98, PY94, Pou90a, RSP94, Sau95, SH96, SC98a, SG90, Szt92, TBA99, TDF^{+}92, Tsi91, TT92, UWT99, Voi90, Wad95, WC97, WY93, XLJS98, YS90, YT98, YNK92, YF92a, ZC96, Ano97-161, Ano99-192, Ano98-144]. system-on-a-chip [Ano99-201]. system-theoretic [Mit92a]. Systematic [WJ90, Tau90b]. Systems [Ano97-132, Ano98-93, RG92, Sch92b, Sti93b, Abe90, AD93, AO97, AO98a, ACHP94, AO93b, AO93a, APV97, AdMD99, Alm96, Alt90a, Alt90b, Alt90c, Ame99, Amo93, APP96, AGM99, And91, Ano90q, Ano95k, Ano95r, Ano95s, Ano95p, Ano96i, Ano96-150, Ano96-39, Ano96c, Ano96-56, Ano96-135, Ano97g, Ano97-105, Ano97-126, Ano97-163, Ano97-52, Ano98h, Ano98-41, Ano98-68, Ano98-71, Ano98-72, Ano98-132, Ano98-155, Ano98-161, Ano99o, Ano99q, Ano99-45, Ano99-58, Ano99-88, Ano99-93, Ano99-128, Ano99-131, Ano99-154, Ano99-155, Ano99-229, Ano99-212, AS97, Ara99a, ADFP93, ARS98, BKN97, BBKN97, BV99, Bal90, BAS97, Ben99, Ber91, Boh96, BG95, BG90, BRZS97, BJ90, CCGL99, CP98a, CST97, CKL91, Cha93, Cha98, Cha94, CCY95a, CCY95b, CCH97, CL91a, CL91b, CK97b, CTY95, CY96, CCC^{+}98]. systems [CL90, Cor94, CS90, Cse95, DBA98, De 96, De 99, DSKP96, Dre90, ESIP98, Far91a, Far96b, FLW99, Flo97, FK95, Gal99a, Gla93c, Gla96c, Gla90c, GSST97, GM96, GT99, Gre90a, GL94, GG92, Gup94, HI98, Han92, Jay90, JP95b, JV96, JPF96, JP93, JI94, KU91, Kar96b, KAD99, Kis91, KG95b, KBLS90, KPB97, KB90, KLMY97, KYNY99, LK94, LS98a, Lew91a, Lew94, LB94, LL92, Lin96, LR90, LL90c, LF98, LP99, LFP97, Lu96, LW99b, Mai96b, MG91, Mav93, McL95, MZMQ90, MP91, MK93, MP99, MW99, MH93a, Mur94, NMUY92, NPJ95, Nov90, OCS96, OSLC98, Oli93, OH93, Pag90, PA98, Pap92, PC96b, Pet92, PP97, Pin98, Pol90, Ran93, RG90a, RTJ96, Rei98, Rei99b, RS99c, Roj90, RZ91, RG90c, RS90, Sar94, Scr91, Sen92, SZ98b]. systems [SWY98, Shi94, Shu93a, Shu93b, Sie90, Sol97, Sti97, SJS92, SCC97, Tan93b, Tan93c, Tan94, TL94, TC99, Ter91, TM91, Tra97, Tre94, Tre95a, Tre98, Trz92, TNS98, VR95b, Wan91, WSS92, Wil92, Wit96b, WR93, Wu98, XL97, XLG95, XLPG99, YQM99, Yan93d, Yan99b, Yas94, Yoo93, Yos96, Yos99, ZQ93a, Zha93b, Zha97, Zha99, ZZ99b, ZQ94, ZQ95, Ano96f, Ano97-36, Ano98x, Ano98-50, Ano98-56, Ano98-105, Ano98-138, Ano99-115, Ano99-116, Ano99-145, Ano99-177]. systems-Towards [Wit96b]. systolic [Lin91a, Lin94, Lin96, MMMS97, MMMS96, Sto92]. Széchenyi [Rod94f]. Szentágothai [Ano98-113].

T [Ano90w, Ano95-44, Ano96-32, Ano96-35, Ano96-68, Ano96-71, Ano96-116, Ano96-120, Ano97-144, Ano98j, Ano98i, Ano98n, Ano98-52, Ano98-61, Ano98-77, Ano98-120, Ano99x, Ano99-41, Ano99-42, Ano99-131, Los92]. table [KM96, Lew91b, SHK99]. Tables [KM97a, DFMS99, Lew91a, LH94]. Tabu [BT94, KVCA^{+}99, DdW93]. tabular [MR91]. Taha [Ano97-113, Ano97-114]. taintings [PvZ95]. Takaci [Ano98-126]. Takatoshi [Ano98-136]. Takefuji [Ano93m]. Takeo [Ano99-118, Ano99-119]. Talbott [Ano95-76]. Tamás [Ano99-165]. tandem [Kat92, Tsi91, TT92, YI97]. tangencies [GST97]. tangent [Zha91]. tanh [GT97a]. tardiness [Cha99a, WF96]. target [GP93, MM96e, WU94]. targets [Ano96-139]. task [PS93b]. tasks [BT94, LC94]. Tau [ZB95, Zha95, AO98c, EDO98, LP99, OPD97, Zha96, AO91, EDO93, EDOS93, FO91, HAO91, KK97a, KO91, NO93]. Tau-method [ZB95, Zha95, Zha96]. taxonomy [Bro96]. Taylor [Ano96-43, Ano97-106, Ano99-29, Ano99-30, CC94b, Dea98]. TC [Ano99-120]. TC-6 [Ano99-120]. TC6 [Ano99-212]. Tchebychef [PT96]. Tchebychef-like [PT96]. TCL [Ano99-210]. TCL/TK [Ano99-210]. Te [Ano99-121]. Te-Won [Ano99-121]. Teachers [Ano99-89]. teaching [Ano96-99, Ano97-55]. Team [Ano98t, Ano99-35, Ano99-36]. technical [Ano96-94]. technique [Ano93f, Ano99l, Ber93, BS91a, Bla95, BG91b, Dat90, EDD91, FK95, HWL94, JS98, Koç92, Kou98, NR96b, Rán92, SM90, SAS99]. Techniques [Ano97-65, Ano98-65, Ano98-150, Ano98-140, Ano97-89, Ano98x, Ano98-117, Ano99o, Ano99-63, Ano99-64, Ano99-87, Ano99-154, Ano99-155, Ano99-242, BP97, CMMR92, CL99c, EDOS93, FGN90, HM93a, Hin95, HMZ99, Kar96a, KO91, KL96a, Koç95, LT96, Li98d, MATA98, ND95, PYS97, PJ91, RT96, SB99, VAB91, Ano97-98, Ano97-99, Ano97-97, Ano98-119, Ano98-133, Ano99-169]. techniques. [Ano97y]. technological [Ano95-27, Ano95-29, Ano96l]. technologies [Ano96-62, Ano98-47, Ano98-110]. Technology [Ano96-43, Ano98-161, Ano95-29, Ano95-75, Ano96-150, Ano96-152, Ano96-52, Ano96-142, Ano98-38, Ano98-31, Ano99-128, Gil93, Las96, Ano95o, Ano95-50, Ano97-133, Ano98-70, Ano98-164, Ano98-145, Ano98-146]. Technoscientific [Ano95-69]. telecommunications [Ano97-37, Ano98-38, Ano97-139]. Telephone [Ano97-161]. Telepresence [Ano99-211]. Teles [Ano96-99]. television [Ano97-148]. Temme [Ano96-146]. Temperature [CR90b, ASW98, CR90a, MB92, MH93b]. Temperature-rate [CR90b, CR90a]. temperatures [ASW98]. tempered [CWJ92]. template [Ano99-55, Ano99-56]. temporal [VDLM95]. ten [Ano97-139]. Tenek [Ano98-57]. Teneketzis [Ano98-57]. tense [Ano96-140]. tension [CHW93, Hun95, HW90]. Tensor [Has95, Has93b, Las94, Las95]. ter [Ano96-140, Ano97-64]. Terence [Ano97-31]. term [LF99, NO93, Ano98-147]. Terminal [Seb91, AS93, AS99, SG90, WT90b]. termination [LNM90]. termites [Ano97-160]. terms [AR92c, Cha96, CS93b, Cho94, DF94, EDO98, FK97, Gla93c, Gla93e, KT97, KAD99, LF98, Zaf98a]. ternary [PM95b]. terrain [UWT99]. Terrance [Ano99-232]. Terrence [Ano93t]. Terry [Ano96-101, Ano98-145, Ano98-146, Ano99-89]. ters [Ano95-49]. Tesauro [Ano96i]. test [BG91a, De 99, KT97, PW95, Sou90, Ano99o]. testability [FF92]. tested [KT98]. Testing [Ano99-212, Szu96, YOO92, Ano98-133, Kor90a, Kor91, LY99a, SD93, Yan91, Yan95, Ano99-212]. Testing-domain [YOO92]. Tests [MZ93, Ano98-142, Gry94]. tethered [MP93]. tetrahydrofolate [SHT^{+}90]. Teubner [Ano96-120, Ano97-136]. Teukolsky [Ano90w]. Text [KCDV99, Ano98-109, Ano99-142, BD96, DB96]. textbook [Ano96-147]. th [Ano99-212, AP94]. Tharam [Ano98-50]. Their [Ano98-92, AS97, AR92e, Ano97-106, Ano99u, Ano99-234, AC93b, Bal96a, CH99, Gir92, Gla95c, Gup94, GMBS94, HSH90a, Har92c, Köh96c, Mis90, NP96, Róz99, Sar94, Saw92, Sza90a, Sza90b, Ver99a]. Thelen [Ano96c]. Themistocles [Ano95-33]. Theobald [Ano99v, Ano99w]. Theodorsen [Sch99]. Theorem [BL97, BW93, Pal97, ACHP94, Arg98, Arg99, BW96b, Bai96d, BKM95, Cub96, DCL^{+}93, Fje91a, GP90, Kor90b, O’R95, Yao93, ZC98, Ano96-124]. theorems [AD93, AG99c, Ana95a, CTY99, FT97b, GS95, GL93, Gyi96, Lal94, Lin91b, Nam96b, Tha94, TSL96, Van92, Ver99b, Ver99a, Wu99b, ZY98, ZL98b]. theoretic [Ano95t, Ano96m, Kou89, Kou90b, Mit92a]. Theoretical [AO93b, Yan99a, Ano97-92, Dem94a, Gab94b, Ano95-81]. Theories [Ano96-144, Ano95-41, Ano96-40, Ano97-156, Tho92, Ano96-163, Ano99-228]. Theory [Ano95z, Ano96x, Ano96-47, Ano98-37, Ano98-43, Ano98-71, Ano98-78, Ano98-124, Ano99-93, Ano99-171, Bou93, DCL^{+}94, EK97, Gla96b, Har90, Iri98, PW95, Rea97, Ano95r, Ano95s, Ano95-30, Ano95i, Ano95-34, Ano95-37, Ano96t, Ano96j, Ano96-39, Ano97-56, Ano97-79, Ano97-121, Ano97-127, Ano98-59, Ano98-63, Ano98-81, Ano98-140, Ano98-153, Ano98-152, Ano99z, Ano99-27, Ano99-98, Azb99, BJH99, BGM97, BGZ92, BT97, Cho99, DMC^{+}95, DDP99, Erm92, Eti91, Fau94, Fli94c, Fre98, JR94, KLMY97, MTGFG93, MG93, Mol99, Mur93b, Nam99b, OR97b, Pok97, RF91, RC92, RSP94, RS94, RS90, UW94, Vic90, Wag90, Mur91a, Ano98-67, Ano95d, Ano96-35, Ano96-66, Ano96-95, Ano96-100, Ano97p, Ano98i, Ano98s, Ano98w, Ano98-30, Ano98-53, Ano98-73, Ano99-105, Ano99-102, Ano99-156, Ano99-157]. theory [Ano99-208, Ano98-154]. Thermal [GRC^{+}96, TH95, MB94b, Sod92, WF99a, Ano98-155]. thermo [BR95a, Ber98]. thermo-visco-elastic [BR95a]. thermodynamic [SHT^{+}90]. thermodynamics [Lin91c]. thermoelastic [CR90a, GC99, NSD98, SSZ91]. thermomicropolar [CH91]. Thierry [Ano95-34]. thin [DK97a, MB94b, Pok97, RR90, Sal92a, WF91]. thin-plate [Sal92a]. Thinning [Lon95]. Third [Ano99-128, AH98, Ano95-36, Ano96e, Ano96f, Ano97-47, Ano97-62, Ano97-150, Ano97-158, Ano97-153, Ano98-81, Ano98-85, BT98, HP94a, Jon96, KKK91]. third-generation [KKK91]. third-order [AH98]. Thomas [Ano95-27, Ano95-48, Ano96-39, Ano96-47, Ano96-62, Ano96-147, Ano96-162, Ano97-158, Ano97-127, Ano98h, Ano98-142, Ano99-44, Ano99-51, Ano99-166, Ano99-227].

Thompson [Ano95-72, Ano97o, Ano98p]. Thomsen [Ano98-111]. Thornton [Ano96-96]. Thorsten [Ano99v, Ano99w]. thought [Ano95p, Ano96-44]. thousand [Ano98-60]. threaded [LY92]. threads [Ano97-77]. Three [Gre91b, GK91, LKT99, MRZ96, Özy96, Wac94, AR99a, Ano99-231, CC94a, CL99d, CLY91, DKR97, GF94b, Gla91a, Gla95b, GGG^{+}98, GL96, JLDL92, Kel91, Kha97, KS95, Köh95, Li96b, Los89a, MMZ91, NN91, PW97, QZ93b, WF99a, Ano98j, Ano99-89]. three-body [DKR97]. Three-country [Özy96]. Three-dimensional [Gre91b, GK91, MRZ96, Ano99-231, CC94a, Gla91a, GL96, Kha97, KS95, Li96b, WF99a]. three-level [CL99d, NN91]. three-manifolds [Ano98j]. three-phase [CLY91]. three-point [NN91]. three-stage [QZ93b]. Three-variable [Wac94, Los89a]. throughput [YI97]. Thrower [Ano96-108]. Tibor [Ano98-74]. Tienson [Ano97-31]. Tight [Käm94b]. Tim [Ano99-86, Ano99-164]. Time [CRV98, DKSI94, HI98, Lin91c, Shi94, Alm96, AAII92, Ano96-140, Ano98z, Ano98-68, Ano98-138, Ano99-58, BL94, BMM94, Blo91, BS96c, BG95, BP97, CP96, CKL91, Cha91, CTG96, CK97b, CC98c, DKO93, Du94, Eti91, Gal99b, Gan92, HI94, Har92c, HAO91, Isp98, IIP^{+}97, JT99, JP95b, JA96, JD97, KNU99, KM92, KK92, KYNY99, LK94, LS98a, LTR98, LB94, MRZ96, Mat99, MMMS96, Nel96, NR96b, OHI91, Pad92b, Pao98, Pap92, RC92, RSP94, Rei99a, Sak90, Sch92b, She92, SZC91, SC98a, SY96, TPF98, Ter00, TG96, Toy92, Tra97, dCVC97, VK95, VA94, WT90b, WHL^{+}99, Yan99b, ZM94, ZKÖ98, ZC96, Ano96-130, Ano99-144, Ano96-164, Ano99-229, Ano97-126, Ano98-132]. time-delay [Yan99b]. time-dependent [Du94, Eti91, IIP^{+}97, JD97, KNU99, Mat99, She92, WHL^{+}99]. time-dependent-coefficient [TG96].

Time-discretization [Shi94]. time-integration [dCVC97]. time-invariant [Ter00, ZM94]. time-lag [WT90b]. time-stabilizing [AAII92]. Time-varying [HI98, CKL91, Pap92, Ter00, ZM94, Ano99-229]. Timed [ZK99]. times [BLR96, CD98, EA92, KO90, ML90a, Pou90b, YI97]. timetable [Seb91]. Timoshenko [Tad94]. Ting [Ano99-86]. Tino [Ano99-166]. Tips [Ano98-119]. Tk [Ano99-141, Ano99-210]. Tobago [Ano98-83]. Tobias [Ano98-147]. Toda [Gre90a, Kuu94, MG91]. Todd [Ano96i, Ano99-163]. Toeplitz [BLR94, DFS93, PS92, PZDH95, SCC97]. Toeplitz-like [PZDH95]. Tokyo [Ano93h]. Tolerant [DH97, HTT92, Ano98-56]. Tomás [Ano93v]. Tomasz [Ano93p]. tomographic [Ram91c]. tomorrow [Ano97-28]. Tomsk [Ano99-212]. tool [Ano96-120, Rot96]. toolbox [SH96]. Tools [Ano96-27, Ano98-31, Ano95-52, Ano97-143, Ano99-154, Ano99-155, RR91, Ano97-89]. toolset [KBLS90]. topics [Ano97e, Ano99-35, Ano99-36, Ano96-165]. Topological [Mro96, SZ98a, SS93, Ano99-33, CT90, CR90c, MTGFG93, MTLGFG93, MTLGFG95a, MTLGFG95b]. topology [Ano96-79, Ano99d]. Toro [Ano99-172]. Toronto [Ano93p]. torsional [MR91]. torus [Fje91c, HTT92]. Total [CF91c, Zha99]. tour [Ano99j]. Touretzky [Ano96i, Ano97g]. Townend [Ano96-98]. tracing [Oli93, Zha93b]. track [CF92]. tracking [BPLS96, GGG^{+}98, Ram90b, Sau95, SCG^{+}97]. tracks [BFF^{+}97]. trade [Ano98-136, Özy96]. trading [DO92]. traditional [Tan93c]. traditions [Azb99, Gal99c]. traffic [Alf95, Ano96-171, Ano97-160]. trajectories [Pan92b, Tso91]. Transact [Ano99-230]. Transact-SQL [Ano99-230]. transaction [Ano98g, Ano98-156]. transactional [Sco96]. Transcendental [Ano97-138, BKK99, DHV91]. Transcending [Ano95-76]. transcription [LNM90]. transfer [Bal96a, CH91, Che91, HM93b, KM90, LR93a, Mur92a, PN98, Rèp91, SCG^{+}97, SBS99, WU99a, YC92]. Transfinite [BC97]. transform [Che93a, CB93b, Faz94a, FL96, Feu91, Hsu96, HLS91, MO95, Mat99, RSC93]. transform/weighting [Che93a]. transformation [AR92d, CCD99, Chi91, Has97, HV92, KV95, Mar95, Nel96, RB97, TD99, Zed99]. transformations [BWK97, BNR96, GF94a, GB90, IN97, JGB95, Leb98, Li96c, Li98d, PTC^{+}98]. Transforming [CKL91]. Transforms [Ano98-43, BD90, BB93, FK98a, FK98b, HLS91, HWL94, Lin95]. Transient [HH98b, Ano99-242, Che93a, HM93b, Kha97, Kij92, MATA98]. transients [Szm90]. transition [IGM98]. Transitional [Hev97]. transitions [PS93a]. Translated [Ano93o, Ano97-83, Ano98-73, Ano98-134, Ano95-74, Ano95i]. translation [Ano93o, Ano98-73]. Translations [Sti94, DH91]. translators [Ano98-31]. transmission [MK91, Tau90b, YP98b]. transmissivity [HM93a, Hin95]. transparent [SD95]. transport [Ano98-83, DJV96, HNS^{+}90, KNU99, KG94, MP94, Oli98, TNTR91, TH95]. transportation [TI96]. transputers [MNB93]. Transverse [MDVK97, NAKA98]. Tranter [Ano99-210]. trap [CB93a]. trapezoidal [CAZ99, LY92, Rok90]. travel [Ano98-159]. travelers [Ano97-103]. traveling [MC99, TE93]. treat [Ano97-148]. treatment [AO98c, JCL93, Ola90, Sha90b]. Tree [HIIN99, Reg94, AS95, AS98, BV99, DKW98, Fje91c, Hwa99, LWL93, She98, Sok91, TTC96]. trees [Arn90, Bha94, CCH97, JLL^{+}93, LJL^{+}93, Shv95, Ano99-206, Ano99-207]. Trefftz [RT99a]. Trekoff [Ano96-126]. trematode [KC91]. trends [Ano99-165, CL94a, Ano98-157]. Tretyakov [Ano96-117]. trials [Ano98-161]. triangle [Fje91a, Fje91b, Fje91c]. triangles [BC97, Lia95, SS98a]. Triangular [Wei99, BLR94, LT96, Pan94, Pap94b]. triangularitate [Köh97]. triangulated [ND95]. triangulation [AAD98, Wea92]. triangulations [Bör90, GMBS94, LS96a, LS98b]. tribe [Myl92]. trichotomies [LFP97]. trichotomy [AHO99, LC98a]. Tricomi [GW97]. Tridiagonal [RG92, Amo93, CY93, CTY95, CY96, LLR^{+}90, Roj90, Usm94, Zha93c]. Tridiagonalization [Kal90]. triggered [GPCSL90]. trigonometric [BS99a, BDT95, Car94, Il’97, PS95, RT99b]. Triple [Won98b, Com92]. triple-sum [Com92]. trivalent [SSBG91]. trivial [BW92]. Troelstra [Ano97p]. trolley [Yav96]. Trouble [Ano96-162]. Troy [Ano97-78]. Trummer [PTC^{+}98]. truncated [EA92]. truncation [ES97, Tso91]. truth [Ano98-90, Ano99j]. Tsichritzis [Ano96-119]. tube [MP94]. tubes [MP93]. Tuganbaev [Ano99-204]. tumor [Mic90]. tumors [ML90b]. tuning [Ano99-235, Ano99-236]. turbid [UW91]. Turbine [CPC93]. turbulence [Ano99t, SCBGN94]. turbulent [Ano98-84, ZSDB99]. Turek [Ano99-70]. turing [Mai96b]. Turini [Ano96-114]. turn [MPS93]. turning [VL92]. Turoff [Ano96-99]. Turtles [Ano97-160]. tutorial [Ano95-59, Ano96-153]. TVD [Gla91e]. twelfth [Ano96-104]. Two [Ano98-60, GPCGV92, Köh95, Lop93, McK96, RS94, US91, AW91, Ahm94, Ahm95, ACL93, AR98, ASW98, BD90, BACS98, Bed95b, BR94, CCH^{+}90, CST97, CCD99, CD98, CK97b, CSW99, DBM92, DM95, EL96, Fai98, GS93, GF94a, GG97, GP93, Gla92b, Gla94d, Gla95d, Gla95f, Gla98, Gla99, GGG^{+}98, Gol94, GT99, II93, Jay92, JKK98, KO90, KG98, KCDV99, KS95, Lag91, Lay93, LLP94, Li98b, LC96a, LMW95, LF98, Los92, MN91, MG97, MM95b, MZMQ90, Nak99, NNB97, NAKA98, Oli93, PSD96, PC94b, RSS94a, RS93, SM95, ST95, SCC97, Tan96, Tsi91, TE93, Wen99, YI97, ZC96, ZQ93d, hC96a, hCPW98, Ano99-231]. Two- [Ano99-231]. two-dimensional [BR94, CCD99, GF94a, Gla92b, Gla94d, Gla95d, Gla95f, Gla98, Gla99, GT99, JKK98, KCDV99, KS95, MM95b, ST95, TE93, Wen99]. two-grid [Bed95b, SCC97]. two-layer [LMW95, NNB97]. two-level [Fai98, Lay93, LLP94, RS93]. two-parametric [CST97]. two-patch [ZC96]. two-person [Nak99]. two-point [BACS98, DBM92, Gol94, KG98, MN91]. two-queue [Tan96]. two-species [PC94b, ZC96]. two-stage [YI97]. two-station [Tsi91]. Two-step [Lop93, CSW99, MG97, SM95, hC96a, hCPW98]. two-target [GP93]. two-variable [Lag91]. Type [Ahm94, LN98, AO99f, Ahm95, Ame97b, Ano98s, BK98, BM90a, CMP95, Cha99b, CLM99, Com92, Con98, DBM92, DBM94, DBM95, DCL^{+}93, DO91, DK97c, DW97, EA92, EGHS98, FS95, GKT99, GS93, Hak99, Hua98a, Iri98, KT96, Kou95, KS98b, Li10, MG97, MM96b, MSB90, MBV91, RB97, Raq98, RW98, She93, Sim97, TWIF98, VBM90, Ver99a, Web95, Wen99, Wu99b, WY99, YFF92b, YFF92a, YFF93]. Types [Ano95-81, CK97b, DRS91, KO90, LKT99, OSLC98, Ano97-136]. typical [GT95].

U [Ano93o, LNM90]. U-snRNA [LNM90]. U.K [Ano90w, Ano96t, Ano96-72, Ano96-155, Ano97p, Ano97-142, Ano97-28, Ano97-41, Ano97-48, Ano97-47, Ano97-83, Ano97-95, Ano97-148, Ano97-90, Ano97-98, Ano97-99, Ano97-97, Ano97-136, Ano97-159, Ano97-129, Ano98-154, Ano99-172]. U.K./Stuttgart [Ano97-136]. U.S.A.

[Ano97-141]. UED [VAB91]. Ueda [Ano98-115]. Uhlenbeck [AKS96, ABI99]. Uhlig [Ano97-111, Ano97-112]. UK [Ano98y, Ano98-40, Ano98-87, Ano98-119]. Ullman [Ano96-84]. Ullrich [Ano93o]. ultimate [Ano98-122]. Ultra [KO94]. Ultra-high [KO94]. ultraspherical [Doh90, Doh91]. Umamageswaran [Ano99-87]. Una-May [Ano99p]. unary [ZO92]. unbiased [FO93, HM95]. unbounded [BSP94, EDO98, ESKG98, JV96, KP99b, Maz90, ZT98a, ZT98b]. uncertain [Ame99, LS98a, LL94a, LL99, LL96, WSS92]. Uncertainty [LCSN98, Bal90, DM95, EK97, KS97a]. unclassified [Ahm95]. unconstrained [Ill91]. uncontrollable [Iri98]. undamped [FLW99]. underdetermined [ST90]. understanding [Ano97-143]. underwater [SK91]. undulator [DGV94]. unfoldings [Jon96]. unidirectional [Jha90, WJ90]. unification [Lyt92b]. unification-based [Lyt92b]. unified [Ano99-170, BSW96, DMC^{+}95, EDOS93, NB95, RS99b, SHA99b, UK99, Ung90, Ung91]. Uniform [TC99, AR99a, Ano99-132, AI97, ASW98, Dev93, Gla91c, Gla91e, GPCGV92, Har92c, HAO91, Ola90, PvZ95, ST95, Tre98]. Uniformly [CMP95, LN98, Li10, BMN96, Car92, Hua98c, Li97, Li98a, LN99]. unifying [Hsu91]. union [SYZ94]. unique [CW95, EK94]. Uniqueness [JA97, Sun91, YP98a, YP98b]. unit [Dia91, GM99]. united [BDT95]. Univ [Ano90w]. univariate [GH90]. Universal [Jon96, Kup91, Sow92, Ano97-161]. universe [Ano96-57, Ano98-134]. University [Ano93r, Ano95n, Ano95t, Ano95y, Ano95x, Ano95-74, Ano95-73, Ano95-30, Ano95i, Ano95-31, Ano95-75, Ano95-32, Ano95-43, Ano95-66, Ano95-69, Ano95-54, Ano95-68, Ano96t, Ano96x, Ano96-35, Ano96-152, Ano96a, Ano96b, Ano96-50, Ano96-57, Ano96-63, Ano96-60, Ano96-72, Ano96-64, Ano96-89, Ano96r, Ano96-96, Ano96-105, Ano96-108, Ano96-155, Ano96-136, Ano96-127, Ano96-130, Ano96-137, Ano96-144, Ano96-141, Ano96-142, Ano96-161, Ano96-143, Ano96-163, Ano97h, Ano97p, Ano97-142, Ano97-28, Ano97-40, Ano97-41, Ano97-39, Ano97-143, Ano97-48, Ano97-47, Ano97-50, Ano97-59, Ano97-83, Ano97-82, Ano97-146, Ano97-95, Ano97-148, Ano97-90, Ano97-98, Ano97-99, Ano97-97, Ano97-100, Ano97-94, Ano97-151, Ano97-118, Ano97-135, Ano97-158, Ano97-155, Ano97-128, Ano97-154, Ano97-127, Ano97-159, Ano97-129, Ano97-130, Ano98k, Ano98s, Ano98v, Ano98-27, Ano98-35]. University [Ano98-39, Ano98-41, Ano98-40, Ano98-148, Ano98-60, Ano98-70, Ano98-73, Ano98-87, Ano98-106, Ano98-149, Ano98-118, Ano98-119, Ano98-124, Ano98-128, Ano98-150, Ano98-136, Ano98-131, Ano98-134, Ano98-145, Ano98-146, Ano98-147, Ano99-51, Ano99-39, Ano99d, Ano99-68, Ano99-214, Ano99h, Ano99i, Ano99-215, Ano99-140, Ano99-144, Ano99-154, Ano99-155, Ano99-146, Ano99-222, Ano99-223, Ano99-224, Ano99-225, Ano99-203, Ano99-202, Ano99-205, Ano99-228, Ano99-233, Ano99-91, Ano99-239]. UNIX [Ano96-133, Ano96-27]. unknown [Gal93a]. unlagging [Van92]. Unpredictable [ZZ94]. unrelated [Sch92b]. unresolved [Erm92]. unrestricted [CL94b, Tan92]. unsolvability [Ano96-35]. unstable [BBKN97, GSST97, HR99, KP97b, Mag97, PZ99b, PP97, Var98]. Unsteady [MPS92, Gla92a, Gla94e, Gla95d, Gla95f, KM90, RR90, WF90]. Unstructured [Ano93w]. unsupervised [KPB97, Ano99-232]. Untangling [Ano99-233]. Update [Ano98-143]. upgrading [Sar90, Sar94, Ano98-158]. upon [JW93, Kup91, YC92]. Upper [Ano96-102, Ano97-113, Ano97-114, Ano99-185, EZ93, FO91, KT97, KO91, Cla98, LG99, Pap94b, RJ94]. upstream [Gla91e]. upwind [Gla93e, Hua95, KP95b, LM91]. upwinding [MG95]. Urban [Ano98-159]. Uriel [Ano99t]. usability [Ano96-162]. usage [RDK92]. Usama [Ano96h]. Use [Kha97, Sak90, Ano95-55, Ano97-101, Ano97-102, Ano97-103, GW99, KS98a]. used [Ano97-79]. useful [Ush94a]. Usefulness [Ano96-162]. usenet [Ano98-99]. User [Ano99-69, Ano95-72, Ano95-42, Ano95-59, Ano96-120, Ano96d, YL93b, Ano90x]. User-centered [Ano90x]. users [Ano95f]. uses [Wag90, YFF92b, YFF92a, YFF93]. Using [Ano95-49, Ano98-149, CH93, CCK97, GT97b, PC96b, SK91, SZM98, AIS95b, Ano96-68, Ano96-125, Ano97-34, Ano97-73, Ano97-67, Ano98-80, Ano98-101, Ano98-133, Ano98-165, Ano99-65, BC97, Bar91b, Ber93, Bri97, CFM91, CB93a, CC94b, CP95b, Che91, Cho99, CG94, CF91b, DR97, De 99, Ech95, Feu90, FY93, FGN90, Gla91a, Gla94b, Gla94d, ISM93, KC92a, KCDV99, Kim98, Kim99, Kun90, LCL94, LJ98, Li98c, Lia95, Liu98, LHC96, LS95, MR99, ND95, Nov90, PSS99, QA96, RK98, RS94, RT94, Rob97, RHRZ90, RSV98, SS98a, SRCB98, SYZ94, SZ91, SSN94, WHL^{+}99, WCY94, Yav96, vNvN90a, Ano95-82]. Usmanov [Ano98-89]. Uthurusamy [Ano96h]. utilities [Ano96-159]. utilization [Shv95, Vos91].

V [Ano93f, Ano93h, Ano96-151, Ano96-41, Ano96-106, Ano96-107, Ano96-156, Ano96-117, Ano96-132, Ano96-159, Ano97-40, Ano97-36, Ano97-65, Ano97-164, Ano98j, Ano98r, Ano98-65, Ano98-135, Ano98-161, Ano99-33, Ano99-34, Ano99-96, Ano99-97, Ano99-101, Ano99-129, Ano99-130]. Vachoux [Ano98l]. Vadim [Ano99-198]. Vaidy [Ano95-59]. Vajda [Ano99-165]. Valental [Bur92]. Valerie [Ano98-79].

Valeriote [Ano98i]. Valid [FO93]. validation [FBL99]. Validity [Gal93d, EL99]. Valmir [Ano97i]. valuation [Ano98-115, Saw92]. Value [Car93, DELE99, AL93, AO97, AHY98, AO98b, ABW99, AO99f, AO98c, Ano99-84, Ano99-85, Ano99-183, Ano99-184, Ano99-198, AP98b, AS96a, Bag90, BACS98, BR95c, BA98, BT98, BG91a, But94, Car92, CK91, CK94a, CK94b, DBM92, DBM94, EDNO92, EK94, EGLN95, Gol90, Gol94, HK98, HP99, HM93b, HNS^{+}90, Ike91, JPF96, JG98, JPC99, KG98, LN98, Li10, Lop93, LNG92, MN91, Moi97, PC92, PA94, PR96, QA98, RSS94a, Ram90a, Riz95, RSR98, Roj99, SS94, SM90, Sim93b, Sim93a, Str98, Sun91, SB99, SAS99, Tsi99, WCC99, Wan98a, Wan98b, Wan99, WJ99, WA96a, WL96, Won98b, XS90, vdHS94, vdV95]. valued [Ano97-165, CTY99, Din99b, Hua98b, HLTB99, Isp98, Lin91b, O’R98b]. Values [Hev97, Ano98-70, Gry94, Iri98, Li99b, Mac93, Pan93b]. Vamanamurthy [Ano97-30]. Vandewalle [Ano99-169]. vanishing [BFR94]. Varadhan [GP90]. Värbrand. [Ano98-108]. Vardi [Ano96-138]. Vargas [Ano95-27]. variable [Ano98-84, Che90a, CLY95, Dor99, JCM96, JPF96, KC92a, Lag91, LY92, Los89a, PPC99, PY94, Prn97, Sim98, VHH95, Wac94, Yu98, ZYC98, ZL98a]. variable-step [Sim98]. Variables [BR95b, AR92b, APB97, Ano98-141, BD90, BM90a, FK97, FBL99, HS95, PA99, Ran93, SH95b, Sza90b, Tel91, ZL99a, Ano95-66]. Varian [Ano97w]. variance [Sni91b]. variants [ST94]. Variation [GGK97, HGK93, Pan98a]. Variational [DLS99, Ano99-170, ABK97, BS96d, Bou98, Cas91, CTY99, Cha99b, CY94, CY95, CLY97, CH99, Din97a, Din99a, DL99, Din99c, Din99d, GY93, Hua93, Hua98a, Hua98b, HLTB99, JY93, JKK98, Kor96, Liu96a, Mas92, Noo95, Noo96, NAS97, Noo98, Noo99, WT90a, Wu90b, WX97, Wu99b, WY99, Yao93, YLY96, Ano95-83]. variational-like [CTY99]. variations [Hsu96]. varies [AB93]. Various [Gar96, BSW96]. VARMA [Ter00, ZM94]. Varricchio [Ano99-82]. varying [CKL91, CC98c, HI98, Pap92, Sch92b, SS94, Ter00, ZM94, Ano99-229]. Varzi [Ano96-87]. vasa [MT98]. Vassileva [Ano99n]. Vassili [Ano97-66]. Vatsala [Ano99-96, Ano99-97]. Vaughan [Ano96-152]. Vautrin [Ano99-152]. VCH [Ano99-123, Ano99-124, Ano99-166]. vector [Ano96-86, Ano98-125, Ano99-105, Ano99-103, Ano99-102, Ano99-104, CTY99, CSOL96, Din99d, Kal90, Lai97, Lin91b, MM96a, MMZ91, MMMS96, Pan93a, Ram95, RG90c, Ano99-234]. vector-valued [CTY99, Lin91b]. Vectorized [CY96]. vectors [Gla99, MM96b, Mil91]. Veen [Ano99-229]. vehicle [Alf95, Ano96-171]. vehicle-actuated [Alf95, Ano96-171]. Veijalainen [Ano98-156]. velte [Sto99]. Venkatakrishnan [Ano98r]. Vera [Ano98-81]. Verduyn [Ano95u]. verification [Ano95t, Ano98-91, Ano98-133, Ano99-117, Bou99, LRLR95, VDLM95]. Verilog [Ano99-213]. Verilog(R) [Ano99-227, Ano98-160]. Verlag [Ano93o, Ano96q, Ano96m, Ano96-106, Ano96-107, Ano96-156, Ano96-157, Ano96-165, Ano97-140, Ano97m, Ano97w, Ano97t, Ano97-33, Ano97-34, Ano97-46, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97j, Ano97-92, Ano97-96, Ano97-147, Ano97-111, Ano97-112, Ano97-122, Ano97-138, Ano98q, Ano98-80, Ano98-93, Ano98-102, Ano98-105, Ano99v, Ano99w, Ano99x, Ano99-41, Ano99-42, Ano99-44, Ano99-43, Ano99-46, Ano99-66, Ano99-55, Ano99-67, Ano99-56, Ano99-69, Ano99-60, Ano99-78, Ano99-79, Ano99-77, Ano99-70, Ano99-89, Ano99-82, Ano99-109, Ano99-110, Ano99-111, Ano99-112, Ano99-123, Ano99-129, Ano99-124, Ano99-130, Ano99-132, Ano99-218, Ano99-156, Ano99-219, Ano99-157, Ano99-149, Ano99-162, Ano99-193, Ano99-194, Ano99-186, Ano99-189, Ano99-198, Ano99-208]. Vermeulen [Ano98-158]. versatile [HY98]. version [Ano93-27, Ano93-28, Ano96-120, Ano96-149, GG90, IB95, PSS99, RSC93]. versioning [Käm94a]. versions [JLDL92]. versus [Ano98-136, Fox94, Iwa91, Mav93, SHT^{+}90]. vertex [MCG98b]. vertical [FG98]. vertices [KKS98]. vessels [GCLT98, MPS92]. vessels- [MPS92]. Vetterling [Ano90w]. VHDL [Ano99-87, Ano98-144]. VI [Ano99-29, Ano99-30, Ano99-113]. via [Ame99, BKK99, BS93, GZ93, Gem93, HLS91, KL95, KS95, Koç92, QZ93c, RA92a, SFY92, Sug94]. viable [Fil93]. Vianu [Ano95-35]. vibrating [FG97, FG98, LR96a]. vibration [BR95b, CM93]. vibrational [Kou98]. vibrations [NAKA98]. vibroacoustic [NSS92]. vicinity [YZM94, YZM95]. Vicious [Ano97-162]. Victor [Ano95-35, Ano96x, Ano96-41, Ano96-129, Ano96-135, Ano97-36, Ano97-66, Ano98-132, Ano99-208]. Victorian [Ano99-225]. video [Ano98x, Ano98z, Ano99-131, Pen96, Ano97-163, Ano98-161]. video-on-demand [Ano98x, Ano99-131]. Vienna [Ano99-120]. view [Ano93r, Ano96-65, DMC^{+}95, Las96, Ano96-29]. viewpoint [Bar94, LR93a]. views [LT94]. VII [Ano95-27, Ano99t]. Vijay [Ano96-134, Ano99-57, Ano99-175]. Vinet [Ano99-186]. Vink [Ano96-46]. Viorel [Ano99-183, Ano99-184]. Virtual [Ano98-112, Ano95-59, Ano96-55, Ano97-78, Ano99-128, Bec96]. visco [BR95a, Ber98]. visco-elastic [BR95a]. viscoelastic [Hak99]. Viscous [WF90, Ano98-83, LT96, MP93, WF91]. Viscusi [Ano99-38]. Vishveshwara [Ano98-65]. vision [Ano96-163, Ano96-84, Ano95m]. Visions [Ano98-110]. visual [Ano96-112, Ano96-84, Ano97s, Ano97-132, SHK99]. visualization [Ano99-44, Ano95w, Ano97-164]. visualized [MR99]. Vitányi. [Ano97j]. Vivek [Ano99-213]. Vlad [Ano99-98]. Vladik [Ano98n, Ano98-32]. Vladimir [Ano99-128, Ano99-174]. Vladimirovic [Ano99-49]. VLRPs [FT97b]. VLSI [Ano95-53, Ano99v, Ano99w, KCD^{+}97, Rei93b]. Vogelsang [Ano97-139]. Voices [Ano99-173]. Voigt [Ano93w]. Vol [Ado91a, Sal92b, Zad92, Ano95-27, MM94]. Völker. [Ano98-138]. Volle [Ano99-95]. Vollmer [Ano99-132]. Volterra [Zha98, KS97b, BMB94, CKRV98, CRV98, DIT97, GS91, GS93, KS98b, Lu96, LW99b, MH99, Med97, Ola90, PC94b, SWY96a, TC99]. Volume [Ano95r, Ano95s, Ano97x, Ano97y, Ano97z, Ano97-65, Ano97-127, Ano97-130, Ano98-67, Ano98-93, Ano98-62, Ano99-105, Ano99-103, Ano99-102, Ano99-104, Ano99-113, Ano99-158, Ano99-160, Ano99-159, Ano99-161, QZ93a, Ano90-30, Ano93p, Ano95-37, Ano96-39, Ano96-81, Ano96-80, Ano96-83, Ano96-129, Ano97f, Ano97-107, Ano98-49, Ano99p, Ano99-89, Ano99-174, WF99a]. Volume-Preserving [QZ93a]. volumes [Las94, Las95]. VonStorch [Ano96-42]. Vorkoetter [Ano96-107]. Voronoi [Dwy93]. Vorozhtsov [Ano96-41]. vortex [Ech96, ZQ93a]. vorticity [LS99]. Vose [Ano99-226]. Vries [Ado91c, SVS98, Sch94]. Vrijer [Ano95-65]. Vuorinen [Ano97-30].

W [Ano90o, Ano95-36, Ano95o, Ano95y, Ano95f, Ano95-49, Ano96-39, Ano96-66, Ano96-108, Ano96-120, Ano97-53, Ano97-69, Ano97-98, Ano97-99, Ano97-97, Ano98-91, Ano98-106, Ano98-101, Ano99-90, Ano99u, Ano99-38, Ano99-92, Ano99-99, Ano99-100, Ano99-111, Ano99-112, Ano99-182, Ano99-208, Ano99-231]. Wachspress [Ano97-141, DO91, PR91]. Waerden [Gyi96]. Wagener [Ano98-61]. Wagner [Ano96-27, Ano99-88]. Wagon [Ano97m]. Wainer [Ano96-35]. Walberg [Ano96-89]. Wald [Ano98v]. Walker [Ano96-118, Ano99-158, Ano99-159, Ano99-211]. walks [Ano99-198]. wall [CC94c]. walls [PSD96, WW95]. Walmsley [Ano99-109, Ano99-110]. Walsh [Ano90u, Ano99-141, Wad95]. Walter [Ano96v, Ano98-61, Ano99o]. Wand [Ano97-28]. Wang [Ano97d, Ano99-137]. Waqar [Ano98-41]. war [Gil93]. Ward [Ano96-47]. Warren [Ano96-149, Ano99l]. Washington [Ano95-67, Ano96-29, Ano97-37, Ano97-68, Ano97-79, Ano97-139, Ano97-161, Ano99e]. waste [Ano99-38]. Watanabe [Ano99-118, Ano99-119]. water [CC94a, CKK97a, GT97a, Gla94d, Gla95b, Gla96b, Gla97b, Gre90b, Gre93a, KG94, KG97, KG95a, NL91, Wya94, YZM95]. Watson [Ano96-116, Ano98-59]. Watts [Ano95e]. wave [Ano96-91, Ano97-108, Ano99-172, BL94, BK91, BPLS96, DG99, Duc95b, GT97a, IB95, JA97, JG98, KKK^{+}97, Kre91, Mak90, MC99, MZMQ90, Pom92, QZ93c, Vic90]. waveform [JOW98, JMPS94]. Wavelet [AV98, Lin95, KV95, PS95]. wavelets [ANV94, BFR94, KV94b]. Waves [LD90, Alk94, BR95a, Ber98, BU90, CR90a, CR90b, DD99, Hon99, Jon96, KL96a, MK91, NSD98, Pok97, Sak90, Sha90a, TAD91]. way [Ano97-159, BT98, CB93a, DWC95, WCY94]. Wayne [Ano98-69]. WDKT [AO99a]. Weak [FS95, Lay99, Gla94f, Gla95c, Gla95f, Gla96c, HP94b, JK96, Sha90d, Ano97-165]. Weak-type [FS95]. weakly [AO99a, AO99b, Ber93, KKK^{+}97]. wealth [Ano99h, Ano99i]. Web [Ano96-116, Ano96-166, Ano97q, Ano96-28, Ano96-51, Ano97-131, Ano97-166, Ano97-167, Ano98-162, Ano99-235, Ano99-236]. Weber [Ano96-151, BL99]. Webmaster [Ano99-237]. Wei [Ano96-49, Ano99-78, Ano99-79]. Wei-Chi [Ano99-78, Ano99-79]. Weibull [Ano96-66]. Weicheng [Ano95-59]. Weidong [Ano97-104]. Weierstrass [BRV98, KP96, PC93]. Weighted [Pin98, BNR96, GHWL96, LC97a, Lub97, TYC93]. weighting [Che93a, Gla91e]. weights [CL95, ED91b, Lub97]. Weikum [Ano98-45]. Weinheim [Ano99-166]. Weisman [Ano97-37]. Weiss [Ano99-98]. Weissinger [Ano99-31]. Well [JPF96, SJB97]. Well-conditioning [JPF96]. Wellesley [Ano95j, Ano95-51, Ano95-49, Ano95-55]. wellfounded [Ano97-162]. Wellin [Ano96q]. Wellings [Ano98-35]. Welsh [Ano95-63]. Wen [Ano99-182]. Werle [Ano98-38]. Werner [Ano95-53]. Wesley [Ano95-72, Ano95v, Ano95e, Ano95-35, Ano95-34, Ano95f, Ano95-79, Ano95-42, Ano95-50, Ano95-48, Ano95-56, Ano95-57, Ano95-61, Ano95-64, Ano96-151, Ano96-48, Ano96-86, Ano96-119, Ano97-107]. West [Ano97-132, Ano98-83, Ano98-146]. Wester [Ano99-45]. wheels [GMW96]. which [Jon96, Lei96, MSR95, YFF92b, YFF92a, YFF93]. White [Ano96a, Ano99-91, Hak99]. Whitehorn [Ano99-37]. whole [AI97, MNE92]. Widder [FK98a]. Wide [Ano96-116, Ano96-166, Ano96-28]. Wiener [DKO93]. Wilberg [Ano98z]. Wild [Ano99-162]. Wiley [Ano93p, Ano95u, Ano95c, Ano95-33, Ano95-39, Ano96g, Ano96-150, Ano96y, Ano96-37, Ano96-27, Ano96-45, Ano96z, Ano96-33, Ano96-41, Ano96-38, Ano96-54, Ano96-75, Ano96-76, Ano96-77, Ano96-91, Ano96-101, Ano96-109, Ano96-116, Ano96-117, Ano96-120, Ano96-110, Ano96-124, Ano96-126, Ano96-131, Ano96-146, Ano96-147, Ano96-164, Ano96d, Ano96-167, Ano97o, Ano97-27, Ano97-30, Ano97-54, Ano97-61, Ano97-60, Ano97-72, Ano97-145, Ano97-87, Ano97-106, Ano97-121, Ano97-125, Ano97-136, Ano97-132, Ano98m, Ano98p, Ano98-29, Ano98-30, Ano98-42, Ano98-64, Ano98-78, Ano98-81, Ano98-117, Ano98-141, Ano98-142, Ano98-143, Ano98-154, Ano98-159, Ano99-28, Ano99-34, Ano99-47, Ano99-48, Ano99-45, Ano99-53, Ano99-54, Ano99-65, Ano99-71, Ano99-74, Ano99-98, Ano99-123, Ano99-133, Ano99-124, Ano99-134, Ano99-127, Ano99j, Ano99-151, Ano99-166, Ano99-175, Ano99-190, Ano99-191]. Wiley [Ano99-199, Ano99-200]. Wiley-VCH [Ano99-123, Ano99-124, Ano99-166]. Wiley/B [Ano96-120]. Wiley/Teubner [Ano97-136]. Wilfred [Ano99-151]. Wilks [Ano96-58]. Willé [Ano96g]. Willi [Ano97-42]. William [Ano90w, Ano95-82, Ano96-42, Ano96-30, Ano97o, Ano97-41, Ano97-70, Ano98p, Ano98-29, Ano99p, Ano99-73, Ano99-160, Ano99-161]. Willis [PCR90]. Wilsey [Ano99-87]. Wilson [Ano95-36, Ano96-132, Ano97-53, Ano99-90]. wind [FY93]. window [Hal90, Koç95, PPC99]. Windows [Ano98-51, Ano98-98, Ano95-39, Ano96-76, Ano97-169, Ano97-168, Ano98-163]. Windsor [Ano97-141]. Winter [Ano98-64]. WIP [TT92]. within [HSH90b, Mar90, Pol90]. without [Ano99-202, EB93, Gla91e, Wan98a]. Witold [Ano99-52]. Wiwianka [Ano96-120]. Woelk [Ano98-164]. Wojciech [Ano98-55]. Wokingham [Ano95-34, Ano95-50, Ano95-56, Ano96-119]. Wolf [Ano96-127, Ano97-136, Ano98-69]. Wolfe [Ano98-132]. Wolff [Ano93o]. Wolfgang [Ano97-125, Ano98-133, Ano98-156]. Wolfram [Ano97-90]. Wolsey [Ano99-127]. Won [Ano99-121]. Wong [Ano97e, Ano97-77, Ano97-166]. Woods [Ano98-37]. Woody [Ano98-52, Ano98-120]. word [Ano95-75]. words [Ano96-58]. Workflow [Ano98-164]. working [Ano97-95, DC94]. Workingham [Ano96-48]. works [Ano98-82]. Workshop [Ano99-212, Ano90b, Ano95k, Ano99-172]. workstation [SSN94]. workstations [LJ98, Ano96-122]. World [Ano90a, Ano96-116, Ano96-166, Ano97-119, Ano95-31, Ano96-28, Ano98-110, Ano98-134, Ano95i, Ano96-166]. worlds [Ano96-167, Ano97-135, Ano96-55]. Would [Ano96-167]. Would-be [Ano96-167]. Wright [MM95c]. Writing [Ano99-238]. writings [Ano95-61, Ano97-130]. Wu [Ano98x]. Wu-chi [Ano98x].

Xavier [Ano99-133, Ano99-134]. XB

[Maz90]. xenopus [BCB90]. xiang [Ano99r]. XII [Ahm94, Ahm95]. Xmt [Ano95-52]. XSC [Ano93o].

Ya [Ano99r]. Ya-xiang [Ano99r]. Yair [Ano98-124]. Yan [Ano97-119]. Yang [Ano98-37, Ano99-78, Ano99-79, Ano99-153]. Yarger [Ano99-164]. Yates [Ano99-190, Ano99-191]. Year [Ano98-143]. yearbook [Ano96-127]. years [Ano98-60, Har90]. Yee [Ano98-78]. Yerkes [Ano99-239]. Yevtushenko [Ano99-212]. yield [Lei96, MSR95, ZG97]. Yihong [Ano98-76]. Yinyu [Ano98-78]. Yip [Ano93i]. Yoram [Ano96-138]. Yorick [Ano96-58]. York [Ano90r, Ano90t, Ano90v, Ano93h, Ano93o, Ano93p, Ano93r, Ano95p, Ano95v, Ano95c, Ano95-39, Ano95-83, Ano96-150, Ano96e, Ano96f, Ano96y, Ano96v, Ano96-37, Ano96-27, Ano96-45, Ano96z, Ano96-33, Ano96-41, Ano96-38, Ano96-54, Ano96-56, Ano96-65, Ano96-75, Ano96-76, Ano96-77, Ano96-91, Ano96q, Ano96m, Ano96-101, Ano96-109, Ano96-116, Ano96-113, Ano96-106, Ano96-107, Ano96-112, Ano96-156, Ano96-117, Ano96-110, Ano96-125, Ano96-124, Ano96-157, Ano96-123, Ano96-126, Ano96-159, Ano96-146, Ano96-147, Ano96-164, Ano96d, Ano96-167, Ano97m, Ano97o, Ano97w, Ano97-27, Ano97-30, Ano97-34, Ano97-54, Ano97-62, Ano97-61, Ano97-60, Ano97-72, Ano97-71, Ano97-70, Ano97-73, Ano97-67, Ano97j, Ano97-145, Ano97-86, Ano97-87, Ano97-93, Ano97-96, Ano97-147, Ano97-94, Ano97-106, Ano97-122, Ano97-120, Ano97-118, Ano97-121, Ano97-137, Ano97-138]. York [Ano97-132, Ano97-165, Ano98q, Ano98k, Ano98p, Ano98s, Ano98-29, Ano98-35, Ano98-30, Ano98-44, Ano98-42, Ano98-46, Ano98-70, Ano98-73, Ano98-78, Ano98-75, Ano98-81, Ano98-80, Ano98-105, Ano98-117, Ano98-149, Ano98-124, Ano98-128, Ano98-150, Ano98-125, Ano98-141, Ano98-142, Ano98-143, Ano98-147, Ano98-159, Ano98-162, Ano99x, Ano99-28, Ano99-34, Ano99-41, Ano99-42, Ano99-46, Ano99-53, Ano99-55, Ano99-54, Ano99-56, Ano99-65, Ano99-60, Ano99-77, Ano99-89, Ano99-111, Ano99-112, Ano99-129, Ano99-133, Ano99-130, Ano99-134, Ano99-127, Ano99-218, Ano99-156, Ano99-154, Ano99-219, Ano99-157, Ano99-155, Ano99j, Ano99-151, Ano99-149, Ano99-162, Ano99-150, Ano99-222, Ano99-223, Ano99-175, Ano99-193, Ano99-190, Ano99-194, Ano99-191, Ano99-186, Ano99-199, Ano99-200, Ano99-205]. York/Addison [Ano95v]. York/Addison-Wesley [Ano95v]. York/West [Ano97-132]. Yorke [Ano98-46]. Yoshiyasu [Ano93m]. Yu [Ano98o, Ano99-86, Ano99-115, Ano99-116, Ano99-197]. Yuan [Ano99r]. Yuille [Ano99-231]. Yves [Ano97-110, Ano98-58]. Yvinec [Ano98k].

Z  [Ano96-144,  Ano97-159,

Ano99-167, Ano99-168]. Zachary [Ano97-73, Ano98-80]. Zafer [Ano97-82]. Zandbergen [Ano99-83]. Zanevsky [Ano99-230]. Zarnowitz [Ano96x]. Zbierski [Ano97-87]. Zebo [Ano98-144]. Zeljko [Ano98-151]. Zellner [Los92]. Zenios [Ano98-124]. zero [ARI96, GW97, Hun95, IN97, Ker91a, PJ91, PC91, PIT97, PPH98]. zero-finding [PIT97, PPH98]. zero-mapping [IN97]. zeroone [PJ91]. zero-surface-tension [Hun95]. zerofinding [MV95a, MV95b]. zeros [AB93, GNR95, GM99, Kov95, LYL95, MO95, Mey99, NR96a, Pan96, PN98, PM91, PT96, Ric95, VL95, Zha93b]. zeta [AB93, PDK96]. Zhang [Ano96-37]. zigzag [YFZM95a, YFZM95b]. Zimmermann [Ano96-120]. Zitarelli [Ano96-102]. Zofia [Ano97-87]. zonal [DK97a]. zone [KKB97]. Zowe [Ano99-171]. ZPL [Ano99k]. Zukowski [Ano97-74]. Zvi [Ano97-118, Ano98-131]. Zwicky [Ano96w].

References

Adari:1993:ALP

[AA93] S. Adari and R. A. Ammar. Axiomatic-like performance analysis (ALPA). Computers and Mathematics with Applications, 26 (10):59–81, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                                     //www.sciencedirect.com/
                                                     science/article/pii/089812219390026R.

Al-Ayyoub:1998:BCM

[AAD98] A.-E. Al-Ayyoub and K. Day. Block-cyclic matrix triangulation on the Cartesian product of star graphs. Computers and Mathematics with Applications, 36(5):113–126, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001540.

Amitai:1992:PAT

[AAII92] Dganit Amitai, Amir Averbuch, Samuel Itzikowitz, and Moshe Israeli. Parallel adaptive and time-stabilizing schemes for constant-coefficient parabolic PDE’s. Computers and Mathematics with Applications, 24(10):33–53, November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290018D.

Aizenberg:1999:OCA

[AAL99] L. Aizenberg, V. Adamchik, and V. E. Levit. One computational approach in support of the Riemann hypothesis. Computers and Mathematics with Applications, 37(1):87–94, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002442.

Arenstorf:1993:SMZ

[AB93] R. F. Arenstorf and L. L.

Brewer. A study of the motion of zeros of the Epstein zeta function associated to m^{2} + y^{2}n^{2} as y varies from 1 to \sqrt{6}. Computers and Mathematics with Applications, 26(5):57–69, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193900746.

Altman:1994:HCM

[AB94] T. Altman and P. F. Boulos. A hypoquadratic convergence method for Lagrange multipliers. Computers and Mathematics with Applications, 27(1):25–31, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900035.

Antoniou:1997:PCC

[ABB97] I. Antoniou, V. Basios, and F. Bosco. Probabilistic control of chaos: Chaotic maps under control. Computers and Mathematics with Applications, 34(2–4):373–389, July/August 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700134X.

Abelson:1990:BIS

[Abe90] H. Abelson. The bifurcation interpreter: a step towards the automatic analysis of dynamical systems. Computers and Mathematics with Applications, 20(8):13–35, ????

1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090207Z.

Arato:1990:P

[ABF90] M. Arató, A. Bagchi, and I. Fazekas. Preface. Computers and Mathematics with Applications, 19(1):vii, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090075U.

Arato:1999:FCO

[ABI99] M. Arató, S. Baran, and M. Ispány. Functionals of complex Ornstein–Uhlenbeck processes. Computers and Mathematics with Applications, 37(1):1–13, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002375.

Arutyunov:1997:ORE

[ABK97] A. Arutyunov, N. Bobylev, and S. Korovin. One remark to Ekeland’s variational principle. Computers and Mathematics with Applications, 34 (2–4):267–271, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001284.

Akishin:1997:HPP

[ABV97] P. G. Akishin, F. Bosco, and S. I. Vinitsky. Helmholtz problem for plane periodical structures. Computers and Mathematics with Applications, 34(5–6):613– 625, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001570.

Al-Bassam:1992:CDC

[ABVB92] S. Al-Bassam, R. Venkatesan, and B. Bose. Codes for detecting and correcting propagation errors. Computers and Mathematics with Applications, 24(4):103–109, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192900138.

Agarwal:1999:EED

[ABW99] R. P. Agarwal, M. Bohner, and P. J. Y. Wong. Eigenvalues and eigenfunctions of discrete conjugate boundary value problems. Computers and Mathematics with Applications, 38(3–4):159– 183, August 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001923.

Allgower:1993:CMM

[AC93a] E. L. Allgower and C.- S. Chien. Continuationminimization methods for stability problems. Computers and Mathematics with Applications, 25(7):65–79, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390313K.

Arunkumar:1993:GSA

[AC93b] S. Arunkumar and T. Chockalingam. Genetic search algorithms and their randomized operators. Computers and Mathematics with Applications, 25(5):91–100, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902027.

Abbaoui:1994:CAM

[AC94] K. Abbaoui and Y. Cherruault. Convergence of Adomian’s method applied to differential equations. Computers and Mathematics with Applications, 28(5):103–109, September 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001448.

Abbaoui:1995:NIP

[AC95] K. Abbaoui and Y. Cherruault. New ideas for proving convergence of decomposition methods. Computers and Mathematics with Applications, 29(7):103–108, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500022Q.

Ahlbrandt:1994:DIR

[ACHP94] C. D. Ahlbrandt, S. L. Clark, J. W. Hooker, and W. T. Patula. A discrete interpretation of Reid’s roundabout theorem for generalized differential systems. Computers and Mathematics with Applications, 28(1–3):11–21, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194000891.

Allgower:1993:BCG

[ACL93] E. L. Allgower, C.-S. Chien, and W.-C. Lee. Bifurcations of corank greater than two. Computers and Mathematics with Applications, 25 (7):51–64, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390312J.

Abolnikov:1993:SRA

[AD93] Lev Abolnikov and Jewgeni H. Dshalalow. Semiregenerative analysis of controlled bulk queueing systems with a bilevel service delay discipline and some ergodic theorems. Computers and Mathematics with Applications, 25(3):107–116, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219390149P.

Agarwal:1996:AHI

[AD96] R. P. Agarwal and S. S.

Dragomir. An application of Hayashi’s inequality for differentiable functions. Computers and Mathematics with Applications, 32 (6):95–99, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001460.

Agarwal:1998:PSS

[AD98] R. P. Agarwal and S. S.

Dragomir. The property of supermultiplicity for some classical inequalities and applications. Computers and Mathematics with Applications, 35(6):105–118, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000236.

Asarin:1993:CPD

[ADFP93] E. Asarin, Phil Diamond, I. Fomenko, and A. Pokrovskii.

Chaotic phenomena in desynhcronized systems and stability analysis. Computers and Mathematics with Applications, 25(1):81–87, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390214G.

Antoniou:1997:RED

[ADKM97] I. Antoniou, L. Dmitrieva, Yu. Kuperin, and Yu. Melnikov. Resonances and the extension of dynamics to rigged Hilbert space. Computers and Mathematics with Applications, 34(5–6):399– 425, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700148X.

Adell:1995:BDO

[AdlC95] J. A. Adell and J. de la Cal. Bernstein–Durrmeyer operators. Computers and Mathematics with Applications, 30 (3–6):1–14, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500081X.

Alastruey:1999:CSA

[AdMD99] C. F. Alastruey, J. R. González de Mendívil, and M. De la Sen. Criteria for stability in approximate delay systems. Computers and Mathematics with Applications, 38 (7–8):199–206, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002503.

Adomian:1991:ERD

[Ado91a] G. Adomian. Errata to “A review of the decomposition method and some recent results for nonlinear equations”: Computers Math. Applic., Vol. 21, No. 5, pp. 101– 127, 1991. Computers and Mathematics with Applications, 22(8):95, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190018Y.

See [Ado91b].

Adomian:1991:RDM

[Ado91b] G. Adomian. A review of the decomposition method and some recent results for nonlinear equations. Computers and Mathematics with Applications, 21(5):101–127, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190220X.

See errata [Ado91a].

Adomian:1991:SGK

[Ado91c]  G. Adomian.  The sine-

Gordon, Klein–Gordon, and Korteweg–de Vries equations. Computers and Mathematics with Applications, 21(5): 133–136, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190222P.

Adomian:1991:SFP

[Ado91d] G. Adomian. Solving frontier problems modelled by nonlinear partial differential equations. Computers and Mathematics with Applications, 22(8):91–94, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190017X.

Adomian:1994:SPP

[Ado94] G. Adomian. Solution of physical problems by decomposition. Computers and Mathematics with Applications, 27(9–10):145–154, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901325.

Adomian:1995:DBE

[Ado95] G. Adomian. The diffusion-Brusselator equation. Computers and Mathematics with Applications, 29(5):1–3, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400244F.

Adomian:1996:SCN

[Ado96] G. Adomian. Solution of coupled nonlinear partial differential equations by decomposition. Computers and Mathematics with Applications, 31(6):117–120, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000107.

Abelman:1990:NSM

[AE90] S. Abelman and D. Eyre.

A numerical study of multistep methods based on continued fractions. Computers and Mathematics with Applications, 20(8):51–60, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902093.

Aradi:1996:SGP

[AÉ96] I. Aradi and P. Érdi. Signal generation and propagation in the olfactory bulb: Multicompartmental modeling. Computers and Mathematics with Applications, 32 (11):1–27, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001939.

Armfield:1990:SAR

[AF90] S. W. Armfield and C. A. J. Fletcher. Symbolic analysis of reduced forms of the Navier– Stokes equations. Computers and Mathematics with Applications, 19(5):55–65, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090101O.

Abdel-Ghaffar:1991:NCC

[AG91] Khaled A. S. Abdel-Ghaffar. The number of convolutional codes of span one. Computers and Mathematics with Applications, 21(10):33–38, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190059D.

Aulbach:1994:DSD

[AG94] B. Aulbach and B. M. Garay. Discretization of semilinear differential equations with an exponential dichotomy. Computers and Mathematics with Applications, 28 (1–3):23–35, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194000905.

Agarwal:1999:OCF

[AG99a] R. P. Agarwal and S. R. Grace. Oscillation of certain functional differential equations. Computers and Mathematics with Applications, 38(5–6):143–153, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002217.

Agarwal:1999:OHO

[AG99b] R. P. Agarwal and S. R.

Grace. The oscillation of higher-order differential equations with deviating arguments. Computers and Mathematics with Applications, 38 (3–4):185–199, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001935.

Agarwal:1999:OTC

[AG99c] R. P. Agarwal and S. R.

Grace. Oscillation theorems for certain neutral functional differential equations. Computers and Mathematics with Applications, 38(11– 12):1–11, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002801.

Agarwal:1994:P

[Aga94]  Ravi P. Agarwal.  Pref-

ace. Computers and Mathematics with Applications, 28(1–3):xiii, August 1994.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219490104X.

Agarwal:1998:P

[Aga98] Ravi P. Agarwal. Preface. Computers and Mathematics with Applications, 36(10– 12):xv, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800035.

Anand:1999:EEL

[AGM99] L. Anand, D. Ghose, and V. Mani. ELISA: an estimated load information scheduling algorithm for distributed computing systems. Computers and Mathematics with Applications, 37(8): 57–85, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001017.

Amir:1992:SSD

[AH92] R. Amir and A. Hadim. Some structured dynamic programs arising in economics. Computers and Mathematics with Applications, 24(8–9): 209–218, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290199R.

Ayala:1997:LOI

[AH97] V. Ayala and A. K. Hacibekiroğlu.

Local observability of invariant dynamics on compact Lie groups with square integrable output map functions. Computers and Mathematics with Applications, 34 (12):61–70, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002344.

Agarwal:1998:PSN

[AH98] R. P. Agarwal and J. Henderson. Positive solutions and nonlinear eigenvalue problems for third-order difference equations. Computers and Mathematics with Applications, 36(10–12):347–355, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800357.

Amleh:1998:DEE

[AHL98] A. M. Amleh, J. Hoag, and G. Ladas. A difference equation with eventually periodic solutions. Computers and Mathematics with Applications, 36(10–12):401–404, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800400.

Ahmed:1990:APC

[Ahm90] A. H. A. Ahmed. Asymptotic properties of collocation projection norms. Computers and Mathematics with Applications, 19(4):45–50, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190901368.

Ahmed:1991:SLM

[Ahm91]  N. U. Ahmed.  Stabil-

ity of limit measures induced by stochastic differential equations on Hilbert space. Computers and Mathematics with Applications, 22(8):41–48, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190011R.

Ahmad:1994:SSR

[Ahm94] K. E. Ahmad. Small sample results for a nonlinear discriminant function estimated from a mixture of two Burr Type XII distributions. Computers and Mathematics with Applications, 28 (5):13–20, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001367.

Ahmad:1995:END

[Ahm95] K. E. Ahmad. The efficiency of a nonlinear discriminant function based on unclassified initial samples from a mixture of two Burr type XII populations. Computers and Mathematics with Applications, 30(2):1–7, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000727.

Alonso:1999:EDT

[AHO99] A. I. Alonso, Jialin Hong, and R. Obaya. Exponential dichotomy and trichotomy for difference equations. Computers and Mathematics with Applications, 38(1):41–49, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001674.

Agarwal:1998:EPS

[AHY98] R. P. Agarwal, Huei-Lin Hong, and Cheh-Chih Yeh. The existence of positive solutions for the Sturm–Liouville boundary value problems. Computers and Mathematics with Applications, 35(9): 89–96, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000601.

Antoniou:1997:UWL

[AI97] I. Antoniou and V. A. Il’in. The uniform over the whole line R estimates of spectral expansions related to the selfadjoint extensions of the Hill operator and of the Schrödinger operator with a bounded and measurable potential. Computers and Mathematics with Applications, 34(5–6):627–632, September 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001582.

Altintas:1995:FCC

[AIS95a] O. Altintas, H. Irmak, and H. M. Srivastava. Fractional calculus and certain starlike functions with negative coefficients. Computers and Mathematics with Applications, 30(2):9–15, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000738.

Altintas:1995:SAF

[AIS95b] O. Altintas, H. Irmak, and H. M. Srivastava. A subclass of analytic functions defined by using certain operators of fractional calculus. Computers and Mathematics with Applications, 30(1):1–9, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000635.

Ahmad:1995:ABE

[AJ95] K. E. Ahmad and Z. F. Jaheen. Approximate Bayes estimators applied to the inverse Gaussian lifetime model. Computers and Mathematics with Applications, 29 (12):39–47, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000565.

Al-Khayyal:1990:JCB

[AK90a] F. A. Al-Khayyal. Jointly constrained bilinear programs and related problems: an overview. Computers and Mathematics with Applications, 19(11):53–62, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090148D.

Arsham:1990:CAL

[AK90b] H. Arsham and A. B. Kahn. A complete algorithm for linear fractional programs. Computers and Mathematics with Applications, 20(7):11–23, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090344J.

Altman:1993:SSG

[AK93] E. Altman and G. Koole. Stochastic scheduling games with Markov decision arrival processes. Computers and Mathematics with Applications, 26(6):141–148, September 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390123D.

Ahn:1998:CQR

[AK98] Y. J. Ahn and H. O. Kim. Curvatures of the quadratic rational Bézier curves. Computers and Mathematics with Applications, 36 (9):71–83, November 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800193X.

Antoniou:1997:ACS

[AKK^{+}97] I. E. Antoniou, V. V. Kocharovsky, Vl. V. Kocharovsky, Yu. M. Mironov, and I. A. Shereshevsky. Absolute and convective superradiance: Dynamics and macroscopic quantum fluctuations. Computers and Mathematics with Applications, 34(7– 8):751–771, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001739.

Antoniou:1997:Pa

[AKP97a] I. Antoniou, Yu. Kuperin, and I. Prigogine. Preface. Computers and Mathematics with Applications, 34(2–4): xiii–xvi, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001156.

Antoniou:1997:Pb

[AKP97b] I. Antoniou, Yu. Kuperin, and I. Prigogine. Preface. Computers and Mathematics with Applications, 34(5– 6):xiii–xvi, September 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001478.

Antoniou:1997:Pc

[AKP97c] I. Antoniou, Yu. Kuperin, and I. Prigogine. Preface. Computers and Mathematics with Applications, 34 (7–8):xiii–xvi, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197893340.

Arato:1996:EDE

[AKS96] M. Arató, A. Kuki, and A. Szabó. Exact distribution of estimators of parameters in Ornstein–Uhlenbeck processes. Computers and Mathematics with Applications, 31(11):45–54, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000600.

Asic:1991:IEM

[AKV91] Miroslav D. Ašić and Vera V. Kovačević-Vujčić. An implicit enumeration method for global optimization problems. Computers and Mathematics with Applications, 21 (6–7):191–201, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901732.

Agarwal:1993:DPI

[AL93] Ravi P. Agarwal and Bikkar S. Lalli. Discrete polynomial interpolation, Green’s functions, maximum principles, error bounds and boundary value problems. Computers and Mathematics with Applications, 25(8):3–39, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390169V.

Angel:1994:MDP

[AL94] E. Angel and P. Leong. Multidimensional dynamic programming on massively parallel computers. Computers and Mathematics with Applications, 27(9–10):67–75, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901252.

Alfa:1995:MTQ

[Alf95] A. S. Alfa. Modelling traffic queues at a signalized intersection with vehicle-actuated control and Markovian arrival processes. Computers and Mathematics with Applications, 30(8):105–119, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500142L.

See errata [Ano96-171].

Alkahby:1994:DMA

[Alk94] H. Y. Alkahby. The dissipation of magneto-acoustic waves. Computers and Mathematics with Applications, 27 (5):9–15, March 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900728.

Apostolos:1996:RCA

[ALL^{+}96] M. K. Apostolos, M. Littman, S. Lane, D. Handelman, and J. Gelfand. Robot choreography: an artistic-scientific connection. Computers and Mathematics with Applications, 32(1):1–4, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000818.

Almasi:1996:RTF

[Alm96] B. Almási. Response time for finite heterogeneous nonreliable queueing systems. Computers and Mathematics with Applications, 31(11): 55–59, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000612.

Altman:1990:ACD

[Alt90a]  T. Altman.  Application

of contractor directions for solving systems of nonlinear equations. Computers and Mathematics with Applications, 19(6):9–14, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090043J.

Altman:1990:HQC

[Alt90b] T. Altman. A hypo-quadratic convergence method for systems of nonlinear equations. Computers and Mathematics with Applications, 20(1): 57–60, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090069V.

Altman:1990:MIS

[Alt90c] T. Altman. A method of inexact steepest descent for systems of linear equations. Computers and Mathematics with Applications, 19(12): 65–69, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090252F.

Adomian:1995:GLE

[AM95] G. Adomian and R. E. Meyers. The Ginzburg–Landau equation. Computers and Mathematics with Applications, 29(3):3–4, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002227.

Amer:1997:IAS

[Ame97a]  M. A. Amer.  Iterative

and approximate solutions of general parameter dependent nonlinear operator equations in Hilbert spaces. Computers and Mathematics with Applications, 34 (12):81–88, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002368.

Amer:1997:ISP

[Ame97b] M. A. Amer. Iterative solutions of parameter-dependent nonlinear equations of Hammerstein type. Computers and Mathematics with Applications, 34(9):123–129, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001934.

Amer:1998:CSN

[Ame98] M. A. Amer. Constructive solutions for nonlinear multiparameter eigenvalue problems. Computers and Mathematics with Applications, 35 (11):83–90, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000868.

Amemiya:1999:SCC

[Ame99] T. Amemiya. A stabilizing control for a class of uncertain linear delayed systems via state feedback — development of the results. Computers and Mathematics with Applications, 37 (11–12):159–169, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001546.

Amodio:1993:OCR

[Amo93] P. Amodio. Optimized cyclic reduction for the solution of linear tridiagonal systems on parallel computers. Computers and Mathematics with Applications, 26(3):45–53, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901099.

Aulbach:1998:SSL

[AMZ98] B. Aulbach, N. Van Minh, and P. P. Zabreiko. Structural stability of linear difference equations in Hilbert space. Computers and Mathematics with Applications, 36(10–12): 71–76, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800102.

Anastassiou:1995:CTM

[Ana95a] G. A. Anastassiou. Comparison theorems on moduli of continuity. Computers and Mathematics with Applications, 30(3–6):15–21, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000828.

Anastassiou:1995:PCA

[Ana95b] George A. Anastassiou. Preface: Concrete analysis. Computers and Mathematics with Applications, 30 (3–6):xv, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195900594.

Andrus:1991:FMI

[And91] J. F. Andrus. Formulas for modular integration of systems of ode’s. Computers and Mathematics with Applications, 21(8):47–56, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190050E.

Andrus:1993:SMR

[And93] J. F. Andrus. Stability of a multi-rate method for numerical integration of ODE’s. Computers and Mathematics with Applications, 25(2): 3–14, January 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390218K.

Anonymous:1990:IWC

[Ano90a] Anonymous. 13th IMACS world Congress on computation and applied mathematics. Computers and Mathematics with Applications, 19(10):III–IV, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090365Q.

Anonymous:1990:IWB

[Ano90b] Anonymous. 5th international workshop of the Bellman continuum. Computers and Mathematics with Applications, 19(10):III, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090364P.

Anonymous:1990:AI

[Ano90c] Anonymous. About this issue. Computers and Mathematics with Applications, 20 (4–6):ix, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190903087.

Anonymous:1990:A

[Ano90d] Anonymous. Announcement.

Computers and Mathematics with Applications, 20 (3):III, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090035I.

Anonymous:1990:AG

[Ano90e] Anonymous. Author’s guidelines. Computers and Mathematics with Applications, 20 (4–6):xi, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190903098.

Anonymous:1990:BRa

[Ano90f] Anonymous. Book reports. Computers and Mathematics with Applications, 19(2): 99–100, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090014B.

Anonymous:1990:BRb

[Ano90g] Anonymous. Book reports. Computers and Mathematics with Applications, 19(4): 95–96, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901416.

Anonymous:1990:BRc

[Ano90h] Anonymous. Book reports. Computers and Mathematics with Applications, 19 (6):95, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090049P.

Anonymous:1990:BRd

[Ano90i] Anonymous. Book reports.

Computers and Mathematics with Applications, 19(10): 99–102, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090362N.

Anonymous:1990:BRe

[Ano90j] Anonymous. Book reports.

Computers and Mathematics with Applications, 19(12): 95–96, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090255I.

Anonymous:1990:BRf

[Ano90k] Anonymous. Book reports.

Computers and Mathematics with Applications, 20(2): 81–92, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090243D.

Anonymous:1990:BRg

[Ano90l] Anonymous. Book reports.

Computers and Mathematics with Applications, 20(3): 83–91, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090033G.

Anonymous:1990:BRj

[Ano90m] Anonymous. Book reports. Computers and Mathematics with Applications, 20(7): 75–86, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090351J.

Anonymous:1990:BRk

[Ano90n] Anonymous. Book reports. Computers and Mathematics with Applications, 20(8): 77–88, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190902112.

Anonymous:1990:BRAa

[Ano90o] Anonymous. Book review: Artificial intelligence — A knowledge-based approach: B. Morris W. Firebaugh. Boyd & Fraser. (1988). 740 pages. $40.00. Computers and Mathematics with Applications, 19 (7):100, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090202U.

Anonymous:1990:BRCa

[Ano90p] Anonymous. Book review: Computer science — a modern introduction: By Les Goldschlager and Andrew Lister. Prentice–Hall, Englewood Cliffs, N.J. (1988). 330 pages. $30.00. Computers and Mathematics with Applications, 19(7):99–100, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902004.

Anonymous:1990:BRCb

[Ano90q] Anonymous. Book review:

Computer systems development: By Denis A. Connor. Prentice–Hall, Englewood Cliffs, N.J. (1988). 194 pages. $26.25. Computers and Mathematics with Applications, 19(7):100, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090203V.

Anonymous:1990:BRFa

[Ano90r] Anonymous. Book review:

Fractals everywhere: By Michael Barnsley. Academic Press, New York. 394 pp. Computers and Mathematics with Applications, 20(4– 6):339, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090340P.

Anonymous:1990:BRH

[Ano90s]  Anonymous.  Book re-

view: Harnessing the corporate PC: By Joseph L. Podolsky. Prentice–Hall, Englewood Cliffs, N.J. (1988). 354 pages. $29.95. Computers and Mathematics with Applications, 19(7):99, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090199T.

Anonymous:1990:BRI

[Ano90t] Anonymous. Book review: Input devices: Edited by Sol Sherr. Academic Press, New York. 301 pp. Computers and Mathematics with Applications, 20(4–6):339–340, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090341G.

Anonymous:1990:BRL

[Ano90u] Anonymous. Book review: Lectures on the electrical properties of materials: By L. Soylmar and D. Walsh. OUP, Oxford (1988). 465 pages Paper $27.50; Cloth $55.00. Computers and Mathematics with Applications, 19(7):99, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090198S.

Anonymous:1990:BRNb

[Ano90v] Anonymous. Book review:

Neural and brain modeling: By Ronald J. MacGregor. Academic Press, New York. 643 pp. Computers and Mathematics with Applications, 20(4–6):339, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090339L.

Anonymous:1990:BRNa

[Ano90w] Anonymous. Book review:

Numerical recipes in C — The art of scientific computing: By William H. Press, Brian P. Flannery, Saul A. Teukolsky and William T. Vetterling. Cambridge Univ. Press, U.K. (1988). 735 pages. $44.50. Computers and Mathematics with Applications, 19 (7):100, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090201T.

Anonymous:1990:BRU

[Ano90x] Anonymous. Book review:

User-centered requirements analysis: By Charles F. Martin. Prentice–Hall, Englewood Cliffs, N.J. 305 pages. $40.00. Computers and Mathematics with Applications, 19 (7):100, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090204W.

Anonymous:1990:D

[Ano90y]  Anonymous.  Dedication.

Computers and Mathematics with Applications, 19 (8–9):vii–viii, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090258L.

Anonymous:1990:EBa

[Ano90z] Anonymous. Editorial Board. Computers and Mathematics with Applications, 19 (1):i, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090074T.

Anonymous:1990:EBb

[Ano90-27] Anonymous. Editorial Board. Computers and Mathematics with Applications, 20 (1):i, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090061N.

Anonymous:1990:FIC

[Ano90-28] Anonymous. First international conference on computational modelling of free and moving boundary problems. Computers and Mathematics with Applications, 20(7):III–IV, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090353L.

Anonymous:1990:LCAa

[Ano90-29] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 19(12):i–xii, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090257K.

Anonymous:1990:LCAb

[Ano90-30] Anonymous. List of contents and author index volume 20, 1990. Computers and Mathematics with Applications, 20(12):i–xi, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090169K.

Anonymous:1990:P

[Ano90-31] Anonymous. Preface. Computers and Mathematics with Applications, 20(9–10):vii– viii, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090107U.

Anonymous:1990:SSSa

[Ano90-32] Anonymous. Software survey section. Computers and Mathematics with Applications, 19(2):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090015C.

Anonymous:1990:SSSb

[Ano90-33] Anonymous. Software survey section. Computers and Mathematics with Applications, 19(4):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901427.

Anonymous:1990:SSSc

[Ano90-34] Anonymous. Software survey section. Computers and Mathematics with Applications, 19(6):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090050T.

Anonymous:1990:SSSd

[Ano90-35] Anonymous. Software survey section. Computers and Mathematics with Applications, 19(7):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090205X.

Anonymous:1990:SSSe

[Ano90-36] Anonymous. Software survey section. Computers and Mathematics with Applications, 19(10):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090363O.

Anonymous:1990:SSSf

[Ano90-37] Anonymous. Software survey section. Computers and Mathematics with Applications, 19(12):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090256J.

Anonymous:1990:SSSg

[Ano90-38] Anonymous. Software survey section. Computers and Mathematics with Applications, 20(1):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090073S.

Anonymous:1990:SSSh

[Ano90-39] Anonymous. Software survey section. Computers and Mathematics with Applications, 20(2):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090244E.

Anonymous:1990:SSSi

[Ano90-40] Anonymous. Software survey section. Computers and Mathematics with Applications, 20(3):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090034H.

Anonymous:1990:SSSj

[Ano90-41] Anonymous. Software survey section. Computers and Mathematics with Applications, 20(7):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090352K.

Anonymous:1990:SSSk

[Ano90-42] Anonymous. Software survey section. Computers and Mathematics with Applications, 20(8):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190902123.

Anonymous:1990:SSSl

[Ano90-43] Anonymous. Software survey section. Computers and Mathematics with Applications, 20(11):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190902215.

Anonymous:1990:SSSm

[Ano90-44] Anonymous. Software survey section. Computers and Mathematics with Applications, 20(12):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090168J.

Anonymous:1991:BRa

[Ano91a] Anonymous. Book reports.

Computers and Mathematics with Applications, 21(4): 91–100, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191902494.

Anonymous:1991:BRb

[Ano91b] Anonymous. Book reports.

Computers and Mathematics with Applications, 21(8): 95–100, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190056A.

Anonymous:1991:BRc

[Ano91c] Anonymous. Book reports.

Computers and Mathematics with Applications, 21(9): 73–79, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190126O.

Anonymous:1991:BRd

[Ano91d] Anonymous. Book reports. Computers and Mathematics with Applications, 22(1): 77–89, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900272.

Anonymous:1991:BRe

[Ano91e] Anonymous. Book reports. Computers and Mathematics with Applications, 22(2): 83–92, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900456.

Anonymous:1991:BRf

[Ano91f] Anonymous. Book reports. Computers and Mathematics with Applications, 22(3): 89–92, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190073D.

Anonymous:1991:BRg

[Ano91g] Anonymous. Book reports. Computers and Mathematics with Applications, 22(9): 97–101, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190210U.

Anonymous:1991:BRh

[Ano91h] Anonymous. Book reports.

Computers and Mathematics with Applications, 22(12): 89–92, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190152T.

Anonymous:1991:E

[Ano91i] Anonymous. Editorial. Computers and Mathematics with Applications, 22(10):1–2, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901879.

Anonymous:1991:EBa

[Ano91j] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 21 (1):i, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190224R.

Anonymous:1991:EBb

[Ano91k] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 22 (1):i, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190019Z.

Anonymous:1991:LCA

[Ano91l] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 22(12):i–ix, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190153U.

Anonymous:1991:P

[Ano91m] Anonymous. Preface. Computers and Mathematics with Applications, 21(5):vii, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190211L.

Anonymous:1991:R

[Ano91n]  Anonymous.  Retraction.

Computers and Mathematics with Applications, 22 (11):91, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900375.

Anonymous:1991:SSS

[Ano91o] Anonymous. Software survey section. Computers and Mathematics with Applications, 21(9):I–V, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190127P.

Ano:1992:SPR

[Ano92a] Katsunori Ano. A secretary problem with restricted offering chances and random number of applications. Computers and Mathematics with Applications, 24(1– 2):157–162, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902419.

Anonymous:1992:BSR

[Ano92b]  Anonymous.  Biographical

sketch of Rolland L. Hardy. Computers and Mathematics with Applications, 24 (12):ix–x, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290164D.

Anonymous:1992:BRa

[Ano92c] Anonymous. Book reports.

Computers and Mathematics with Applications, 23(12): 99–102, June 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290096Z.

Anonymous:1992:BRb

[Ano92d] Anonymous. Book reports.

Computers and Mathematics with Applications, 24(10): 117–125, November 1992.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290024C.

Anonymous:1992:EBa

[Ano92e] Anonymous. Editorial Board. Computers and Mathematics with Applications, 23 (1):i, January 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290073Q.

Anonymous:1992:EBb

[Ano92f] Anonymous. Editorial Board. Computers and Mathematics with Applications, 24 (1–2):i, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192902198.

Anonymous:1992:I

[Ano92g] Anonymous. Index. Computers and Mathematics with Applications, 23(6–9):747– 758, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901333.

Anonymous:1992:LCA

[Ano92h] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 23(12):i–viii, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900972.

Anonymous:1992:Pa

[Ano92i] Anonymous. Preface. Computers and Mathematics with Applications, 23(2–5):ix–x, January/March 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901344.

Anonymous:1992:Pb

[Ano92j] Anonymous. Preface. Computers and Mathematics with Applications, 24(1–2):vii– viii, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290220C.

Anonymous:1992:SPEa

[Ano92k] Anonymous. Speed of publication — an editorial. Computers and Mathematics with Applications, 23 (11):1–2, June 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290063N.

Anonymous:1992:SPEb

[Ano92l] Anonymous. Speed of publication — an editorial. Computers and Mathematics with Applications, 24(3): 1–2, August 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290208Y.

Anonymous:1992:SPEc

[Ano92m] Anonymous. Speed of publication — an editorial. Computers and Mathematics with Applications, 24(7): 1–2, October 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290149C.

Anonymous:1993:BRb

[Ano93a] Anonymous. Book reports. Computers and Mathematics with Applications, 25 (4):113–118, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390254S.

Anonymous:1993:BRd

[Ano93b] Anonymous. Book reports. Computers and Mathematics with Applications, 25(7): 101–109, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390315M.

Anonymous:1993:BRe

[Ano93c] Anonymous. Book reports.

Computers and Mathematics with Applications, 25(12): 107–112, June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901907.

Anonymous:1993:BRf

[Ano93d] Anonymous. Book reports.

Computers and Mathematics with Applications, 26(5): 105, September 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390078A.

Anonymous:1993:BRg

[Ano93e] Anonymous. Book reports.

Computers and Mathematics with Applications, 26(5): 107–112, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390079B.

Anonymous:1993:BRA

[Ano93f] Anonymous. Book review:

An artificial intelligence technique for information and fact retrieval: Nicholas V. Findler. The MIT Press, Cambridge, Massachusetts and London, England. (1991). 155 pages. $25.00. Computers and Mathematics with Applications, 25(3):118, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390152L.

Anonymous:1993:BRCa

[Ano93g]  Anonymous.  Book re-

view: Cryptography and coding II: Chris Mitchell. Clarendon Press, Oxford, England. (1992). 301 pages. $87.00. Computers and Mathematics with Applications, 25(3): 118, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390154N.

Anonymous:1993:BRI

[Ano93h] Anonymous. Book review: Images of programming: Dedicated to the memory of A. P. Ershov: Edited by D. Bjørner and V. Kotov. North-Holland, Amsterdam, London, New York and Tokyo. (1991). 270 pages. $97.00. Computers and Mathematics with Applications, 25(3):119, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390157Q.

Anonymous:1993:BRK

[Ano93i] Anonymous. Book review:

KAM: a system for intelligently guiding numerical experimentation by computer: Kenneth Man-Kam Yip. The MIT Press, Cambridge, Massachusetts and London, England. (1991). 242 pages. $37.50. Computers and Mathematics with Applications, 25(3):117, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390151K.

Anonymous:1993:BRL

[Ano93j]  Anonymous.  Book re-

view: Learning mathematics and logo: Edited by Celia Hoyles and Richard Noss. The MIT Press, Cambridge, Massachusetts and London, England. (1992). 483 pages. $45.00. Computers and Mathematics with Applications, 25(3):117, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390150T.

Anonymous:1993:BRMb

[Ano93k] Anonymous. Book review:

Mathematics: a practical approach: Nancy Blachman. Prentice Hall, Englewood Cliffs, New Jersey. (1992). 365 pages. $30.00. Computers and Mathematics with Applications, 25(3):120, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390160W.

Anonymous:1993:BRMa

[Ano93l] Anonymous. Book review: Morphology and computation: Richard Sproat. The MIT Press, Cambridge, Massachusetts and London, England. (1992). 295 pages. $35.00. Computers and Mathematics with Applications, 25(3):118, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390153M.

Anonymous:1993:BRNb

[Ano93m] Anonymous. Book review: Neural network parallel computing: Yoshiyasu Takefuji. Kluwer Academic Publishers, Boston, Dordrecht, and London. (1992). 230 pages. $65.00. Computers and Mathematics with Applications, 25(3):120, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390159S.

Anonymous:1993:BRNa

[Ano93n] Anonymous. Book review: Numerical methods for mathematics, science, and engineering. Second edition: John H. Mathews. Prentice–Hall, Englewood Cliffs, New Jersey. (1992). 646 pages. $59.00. Computers and Mathematics with Applications, 25(3): 118, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390155O.

Anonymous:1993:BRP

[Ano93o] Anonymous. Book review:

Pascal-XSC: Language reference with examples: R. Klatte, U. Kulisch, M. Neaga, D. Ratz and Ch. Ullrich. Translated by G. F. Corless, R. Klatte, U. Kulisch, D. Ratz and C. Wolff. Springer-Verlag, Berlin, Heidelberg and New York. (1991, original; 1992, translation). 344 pages. $42.00. Computers and Mathematics with Applications, 25(3):121, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390162O.

Anonymous:1993:BRRa

[Ano93p]  Anonymous.  Book re-

view: Random graphs (volume 2): Edited by Alan Frieze and Tomasz Luczak. John Wiley & Sons, Inc., New York, Chichester, Brisbane, Toronto, and Singapore. (1992). 285 pages. $98.00. Computers and Mathematics with Applications, 25(3):119, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390156P.

Anonymous:1993:BRRb

[Ano93q]  Anonymous.  Book re-

view: Reinforcement learning: Edited by Richard S. Sutton. Kluwer Academic Publishers, Boston, Dordrecht and London. (1992). $95.00. Computers and Mathematics with Applications, 25(3): 122, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390167T.

Anonymous:1993:BRSa

[Ano93r] Anonymous. Book review: Software engineering: a holistic view: Bruce I. Blum. Oxford University Press, New York and Oxford. (1992). 588 pages. $49.95. Computers and Mathematics with Applications, 25(3):119, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390158R.

Anonymous:1993:BRSb

[Ano93s]  Anonymous.  Book re-

view: Symmetry, causality, mind: Michael Leyton. The MIT Press, Cambridge, Massachusetts and London, England (1992). 630 pages. $45.00. Computers and Mathematics with Applications, 25(3):120, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390161N.

Anonymous:1993:BRCb

[Ano93t] Anonymous. Book review:

The computational brain: Patricia S. Churchland and Terrence J. Sejnowski. The MIT Press, Cambridge, Massachusetts and London, England. (1992). 544 pages. $39.95. Computers and Mathematics with Applications, 25(3):121, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390164Q.

Anonymous:1993:BRCc

[Ano93u] Anonymous. Book review:

The core language engine: Edited by Hiyan Alshawi. The MIT Press, Cambridge, Massachusetts and London, England. (1992). 322 pages. $37.50. Computers and Mathematics with Applications, 25(3):122, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390165R.

Anonymous:1993:BRRc

[Ano93v] Anonymous. Book review: The robotics review 2: Edited by Oussama Khatib, John J. Craig, and Tomás Lozano– Pérez. The MIT Press, Cambridge, Massachusetts and London, England. (1992). 372 pages. $55.00. Computers and Mathematics with Applications, 25(3):122, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390166S.

Anonymous:1993:BRU

[Ano93w] Anonymous. Book review: Unstructured scientific computation on scalable multiprocessors: Edited by P. Mehrotra, J. Saltz and R. Voigt. The MIT Press, Cambridge, Massachusetts and London, England. (1992). 405 pages. $39.95. Computers and Mathematics with Applications, 25(3):121, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390163P.

Anonymous:1993:EBa

[Ano93x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 25 (1):i, January 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390205A.

Anonymous:1993:EBb

[Ano93y] Anonymous. Editorial board.

Computers and Mathematics with Applications, 26 (1):iii–iv, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390080F.

Anonymous:1993:EGA

[Ano93z]  Anonymous.  Evolver^{T M}

2.0: a genetic algorithm for spreadsheets. Computers and Mathematics with Applications, 26(12):94, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900643.

Anonymous:1993:ECV

[Ano93-27] Anonymous. Expert Choice version 8.0. Computers and Mathematics with Applications, 25(8):117, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390179Y.

Anonymous:1993:HV

[Ano93-28] Anonymous. HiQ(R) version 2.0. Computers and Mathematics with Applications, 26(4):99, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390038W.

Anonymous:1993:LCA

[Ano93-29] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 25(12):i–xii, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390191W.

Anonymous:1993:Pa

[Ano93-30] Anonymous. Preface. Computers and Mathematics with Applications, 25(10–11):ix– x, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390274Y.

Anonymous:1993:Pb

[Ano93-31] Anonymous. Preface. Computers and Mathematics with Applications, 26(6):ix, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193901129.

Anonymous:1993:SRa

[Ano93-32] Anonymous. Software reports. Computers and Mathematics with Applications, 25 (8):115, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390178X.

Anonymous:1993:SRb

[Ano93-33] Anonymous. Software reports. Computers and Mathematics with Applications, 26 (4):97, August 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390037V.

Anonymous:1993:SRc

[Ano93-34] Anonymous. Software reports. Computers and Mathematics with Applications, 26(12):93, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900632.

Anonymous:1993:SPE

[Ano93-35] Anonymous. Speed for publication — an editorial. Computers and Mathematics with Applications, 25(8):1–2, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390168U.

Anonymous:1994:DAB

[Ano94a]  Anonymous.  Dedication:

Austin Blaquière 1923–1993. Computers and Mathematics with Applications, 27(9– 10):ix, May 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901155.

Anonymous:1994:EB

[Ano94b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 28(1– 3):i–iii, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901031.

Anonymous:1994:SR

[Ano94c] Anonymous. Software report. Computers and Mathematics with Applications, 28(8):99–101, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900094.

Anonymous:1995:AI

[Ano95a]  Anonymous.  Author in-

dex. Computers and Mathematics with Applications, 30 (12):x–xi, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900551.

Anonymous:1995:BR

[Ano95b] Anonymous. Book report.

Computers and Mathematics with Applications, 29(7): 109–114, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902759.

Anonymous:1995:BRCl

[Ano95c]  Anonymous.  Book re-

view: A concrete approach to mathematical modelling: By Michael Mesterton-Gibbons. John Wiley & Sons, New York. (1995). 597 pages. $64.95. Computers and Mathematics with Applications, 30(9):124, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902139.

Anonymous:1995:BRCc

[Ano95d] Anonymous. Book review:

A course in game theory: By Martin J. Osborne and Ariel Rubinstein. MIT Press, Cambridge, MA. (1994). 352 pages. $40.00 (cloth), $18.95 (paper). Computers and Mathematics with Applications, 29(3):115, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195901019.

Anonymous:1995:BRDg

[Ano95e]  Anonymous.  Book re-

view: A discipline for software engineering: By Watts Humphrey. Addison-Wesley, Reading, MA. (1995). 789 pages. $47.29. Computers and Mathematics with Applications, 30(9):126, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902430.

Anonymous:1995:BRGd

[Ano95f] Anonymous. Book review: A guide to LaTEX2ϵ: Document preparation for beginners and advanced users (second edition): By Helmut Kopka and Patrick W. Daly. Addison-Wesley, Reading, MA. (1995). $36.53. Computers and Mathematics with Applications, 30(9):126, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902376.

Anonymous:1995:BRTc

[Ano95g] Anonymous. Book review: A TEX primer for scientists: By Stanley A. Sawyer and Steven G. Krantz. CRC Press, Boca Raton. (1995). 399 pages. $31.95. Computers and Mathematics with Applications, 30(9):125, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902279.

Anonymous:1995:BRAc

[Ano95h] Anonymous. Book review:

Advances in genetic programming: Edited by Kenneth E. Kinnear, Jr. MIT Press, Cambridge, MA. (1994). 518 pages. $45.00. Computers and Mathematics with Applications, 29(3):114, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900918.

Anonymous:1995:BREa

[Ano95i] Anonymous. Book review:

An equation that changed the world: Newton, Einstein, and the theory of relativity: By Harald Fritzsch, translated by Karin Heusch. University of Chicago Press, Chicago. (1994). 279 pages. $29.95. Computers and Mathematics with Applications, 29(3):110, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219590073X.

Anonymous:1995:BRIa

[Ano95j] Anonymous. Book review:

An introduction to scientific, symbolic, and graphical computation: By Eugene Fiume. A. K. Peters, Wellesley, MA. (1995). 306 pages. $49.95. Computers and Mathematics with Applications, 29(11):110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902732.

Anonymous:1995:BRAa

[Ano95k] Anonymous. Book review: Artificial life IV: Proceedings of the fourth international workshop on the synthesis and simulation of living systems: Edited by Rodney A. Brooks and Pattie Maes. MIT Press, Cambridge, MA. (1994). 444 pages. $60.00. Computers and Mathematics with Applications, 29(3):110, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900713.

Anonymous:1995:BRAe

[Ano95l]  Anonymous.  Book re-

view: Artificial neural networks for intelligent manufacturing: Edited by Cihan H. Dagli. Chapman & Hall, London. (1994). 469 pages. $77.95. Computers and Mathematics with Applications, 29(11):107, June 1995. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902546.

Anonymous:1995:BRAd

[Ano95m] Anonymous. Book review:

Artificial neural networks for speech and vision: Edited by Richard J. Mammone. Chapman & Hall, London. (1994). 586 pages. $77.95. Computers and Mathematics with Applications, 29(11):106, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219590249X.

Anonymous:1995:BRAg

[Ano95n] Anonymous. Book review:

Atoms in the family: My life with Enrico Fermi: By Laura Fermi. The University of Chicago Press, Chicago. (1954). 267 pages. $13.95. Computers and Mathematics with Applications, 30(9): 124, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219590221X.

Anonymous:1995:BRB

[Ano95o] Anonymous. Book review:

Building IBM: Shaping an industry and its technology: By Emerson W. Pugh. MIT Press, Cambridge, MA. (1995). 405 pages. $29.95.

Computers and Mathematics with Applications, 30(9): 123, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902074.

Anonymous:1995:BRCe

[Ano95p] Anonymous. Book review: Chaotic logic: Language, thought, and reality from the perspective of complex systems science: By Ben Goertzel. Plenum Press, New York. (1994). 278 pages. $72.50. Computers and Mathematics with Applications, 29(3):116, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195901116.

Anonymous:1995:BRCd

[Ano95q] Anonymous. Book review: Circuit complexity and neural networks: By Ian Parberry. MIT Press, Cambridge, MA. (1994). 270 pages. $35.00. Computers and Mathematics with Applications, 29(3): 116, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195901078.

Anonymous:1995:BRCa

[Ano95r] Anonymous. Book review:

Computational learning theory and natural learning systems. Volume I: Constraints and prospects: Edited by Stephen José Hanson, George A. Drastal, and Ronald L. Rivest. MIT Press, Cambridge, MA. (1994). 565 pages. $49.95 (Volume I), $39.95 (Volume II). Computers and Mathematics with Applications, 29(3): 112, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900799.

Anonymous:1995:BRCb

[Ano95s]  Anonymous.  Book re-

view: Computational learning theory and natural learning systems. Volume II: Intersections between theory and experiment: Edited by Stephen José Hanson, George A. Drastal, and Ronald L. Rivest. MIT Press, Cambridge, MA. (1994). 449 pages. $49.95 (Volume I), $39.95 (Volume II). Computers and Mathematics with Applications, 29(3): 112, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900810.

Anonymous:1995:BRCf

[Ano95t] Anonymous. Book review:

Computer-aided verification of coordinating processes: the automata-theoretic approach: By Robert P. Kurshan. Princeton University Press, Princeton, NJ. (1994). 270 pages. $49.50 (cloth), £35. Computers and Mathematics with Applications, 29 (11):109, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902678.

Anonymous:1995:BRCj

[Ano95u] Anonymous. Book review: Computer algebra in industry 2: Problem solving in practice: Edited by Arjeh M. Cohen, Leendert van Gastel, and Sjoerd Verduyn Lunel. John Wiley & Sons, Chichester. (1995). 323 pages. $28.95. Computers and Mathematics with Applications, 30(9):122, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195901930.

Anonymous:1995:BRCh

[Ano95v] Anonymous. Book review: Computer-related risks: By Peter G. Neumann. ACM Press, New York/Addison-Wesley Publishing Company, Reading, MA. (1995). 367 pages. $22.95. Computers and Mathematics with Applications, 29(11):109, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902643.

Anonymous:1995:BRCm

[Ano95w]  Anonymous.  Book re-

view: Computer visualization: Edited by Richard S. Gallagher. CRC Press, Boca Raton. (1995). 312 pages. $69.95. Computers and Mathematics with Applications, 30(9):124, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902198.

Anonymous:1995:BRCk

[Ano95x] Anonymous. Book review:

Coordination and information: Historical perspectives on the organization of enterprise: Edited by Naomi R. Lamoreaux and Daniel M. G. Raff. University of Chicago Press, Chicago. (1995). 337 pages. $68.00 (cloth), $22.50 (paper). Computers and Mathematics with Applications, 30(9):123, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902058.

Anonymous:1995:BRCi

[Ano95y] Anonymous. Book review:

Cross-level inference: By Christopher H. Achen and W. Phillips Shively. University of Chicago Press, Chicago. (1995). 248 pages. $55.00 (cloth), $18.95 (paper). Computers and Mathematics with Applications, 30(9):121, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901914.

Anonymous:1995:BRCo

[Ano95z] Anonymous. Book review: Cryptography: Theory and practice: By Douglas R. Stinson. CRC Press, Boca Raton. (1995). 434 pages. $64.95. Computers and Mathematics with Applications, 30(9):125, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902252.

Anonymous:1995:BRDb

[Ano95-27] Anonymous. Book review: Decision making in economic, political, social and technological environments with the analytic hierarchy process: the analytic hierarchy process series, vol. VII: By Thomas L. Saaty and Luis G. Vargas. RWS Publications, Pittsburgh, PA. 330 pages. $15.00. Computers and Mathematics with Applications, 29(3):116, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901094.

Anonymous:1995:BRDa

[Ano95-28]  Anonymous.  Book re-

view: Dimensions of creativity: Edited by Margaret A. Boden. MIT Press, Cambridge, MA. (1994). 242 pages. $35.00. Computers and Mathematics with Applications, 29(3):111, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900772.

Anonymous:1995:BRDc

[Ano95-29] Anonymous. Book review:

Does technology drive history? The dilemma of technological determinism: Edited by Merritt Roe Smith and Leo Marx. MIT Press, Cambridge, MA. (1994). 280 pages. $35.00 (cloth), $16.95 (paper). Computers and Mathematics with Applications, 29(3):116, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901035.

Anonymous:1995:BRDf

[Ano95-30]  Anonymous.  Book re-

view: Dynamics of meaning: Anaphora, presupposition, and the theory of grammar: By Gennaro Chierchia. University of Chicago Press, Chicago. (1995). 270 pages. $75.00, £59.95 (cloth); $31.95, £25.50. Computers and Mathematics with Applications, 30(9):126, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902392.

Anonymous:1995:BREb

[Ano95-31] Anonymous. Book review: Existential cognition: Computational minds in the world: By Ron McClamrock. University of Chicago Press, Chicago. (1995). 205 pages. $28.95 (cloth), £23.25. Computers and Mathematics with Applications, 29 (11):109, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902694.

Anonymous:1995:BRFd

[Ano95-32]  Anonymous.  Book re-

view: Fields and rings, second edition (Chicago lectures in mathematics): By Irving Kaplansky. University of Chicago Press, Chicago. (1972). 206 pages. $20.00. Computers and Mathematics with Applications, 30(9): 123, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902112.

Anonymous:1995:BRFf

[Ano95-33] Anonymous. Book review:

Finite sums decompositions in mathematical analysis: By Themistocles M. Rassias and Jaromír Simsa. John Wiley & Sons, Chichester. (1995) 172 pages. $29.95. Computers and Mathematics with Applications, 30(9):125, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902317.

Anonymous:1995:BRFb

[Ano95-34] Anonymous. Book review:

Foundations of computing: System development with set theory and logic: By Thierry Scheurer. Addison-Wesley Publishing Company, Wokingham, England. (1994). 668 pages. $42.95. Computers and Mathematics with Applications, 29(11):109, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902651.

Anonymous:1995:BRFa

[Ano95-35] Anonymous. Book review:

Foundations of databases: By Serge Abiteboul, Richard Hull and Victor Vianu. Addison-Wesley Publishing Company, Reading, MA. (1995). 685 pages. $50.95. Computers and Mathematics with Applications, 29 (11):108, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902589.

Anonymous:1995:BRAb

[Ano95-36] Anonymous. Book review: From animals to animats 3: Proceedings of the third international conference on simulation of adaptive behavior: Edited by Dave Cliff, Philip Husbands, Jean-Arcady Meyer, and Stewart W. Wilson. MIT Press, Cambridge, MA. (1994). 508 pages. $55.00. Computers and Mathematics with Applications, 29(3):111, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900756.

Anonymous:1995:BRGa

[Ano95-37] Anonymous. Book review: Game theory and the social contract, volume 1: Playing fair: By Ken Binmore. MIT Press, Cambridge, MA. (1994). 364 pages. $39.95. Computers and Mathematics with Applications, 29(3): 113, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219590087X.

Anonymous:1995:BRGb

[Ano95-38] Anonymous. Book review:

Genetic programming II: Automatic discovery of reusable programs: By John R. Koza. MIT Press, Cambridge, MA. (1994). 746 pages. $45.00. Computers and Mathematics with Applications, 29(3): 115, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900993.

Anonymous:1995:BRGe

[Ano95-39] Anonymous. Book review:

Getting started with Lotus 1-2-3 5.0 for Windows: Barbara Farrell. John Wiley & Sons, New York. (1995). 176 pages. $13.95. Computers and Mathematics with Applications, 30(9):126, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902414.

Anonymous:1995:BRH

[Ano95-40]  Anonymous.  Book re-

view: Humour the computer: Edited by Andrew Davison. MIT Press, Cambridge, MA. (1995). 226 pages. $13.95.

Computers and Mathematics with Applications, 30(9): 120–121, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219590185X.

Anonymous:1995:BRLb

[Ano95-41] Anonymous. Book review: Large-scale neuronal theories of the brain: Edited by Christof Koch and Joel L. Davis. MIT Press, Cambridge, MA. (1994). 340 pages. $55.00. Computers and Mathematics with Applications, 29(3):115, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900977.

Anonymous:1995:BRLc

[Ano95-42] Anonymous. Book review: LaTEX: a document preparation system user’s guide and reference manual: By Leslie Lamport. Addison-Wesley Publishing Company, Reading, MA. (1994). 272 pages. $32.95. Computers and Mathematics with Applications, 29(11):108, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902619.

Anonymous:1995:BRLe

[Ano95-43] Anonymous. Book review:

Lie algebras and locally compact groups (Chicago lectures in mathematics): By Irving Kaplansky. University of Chicago Press, Chicago. (1971). 148 pages. $22.00. Computers and Mathematics with Applications, 30(9): 124, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902155.

Anonymous:1995:BRLd

[Ano95-44] Anonymous. Book review:

Linear algebra: (Modular Mathematics Series). By R. B. J. T. Allenby. Edward Arnold, London. (1995). 227 pages. $15.95. Computers and Mathematics with Applications, 29(11):110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902716.

Anonymous:1995:BRLa

[Ano95-45] Anonymous. Book review:

Logic programming: Proceedings of the eleventh international conference on logic programming: Edited by Pascal Van Hentenryck. MIT Press, Cambridge, MA. (1994). 752 pages. $75.00. Computers and Mathematics with Applications, 29 (3):112–113, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195900837.

Anonymous:1995:BRMd

[Ano95-46] Anonymous. Book review: Managing Internet information services: By Cricket Liu, Jerry Peek, Russ Jones, Bryan Buus, and Adrian Nye. O’Reilly & Associates, Sebastopol, CA. (1994). 630 pages. $29.95. Computers and Mathematics with Applications, 29(11):107, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902538.

Anonymous:1995:BRMc

[Ano95-47]  Anonymous.  Book re-

view: Massively parallel artificial intelligence: By Hiroaki Kitano and James A. Hendler. AAAI Press/MIT Press, Menlo Park, CA/ Cambridge, MA. (1994). 426 pages. $39.95. Computers and Mathematics with Applications, 29(3):115, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195900950.

Anonymous:1995:BRMg

[Ano95-48] Anonymous. Book review:

Mathematica(R) for scientists and engineers: By Thomas B. Bahder. Addison-Wesley Publishing Company, reading, MA. (1995). 846 pages. $42.95. Computers and Mathematics with Applications, 29 (11):108, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902597.

Anonymous:1995:BRMh

[Ano95-49] Anonymous. Book review:

Matrix algebra: Using MIN-Imal MATlab^{T M}: By Joel W. Robbin. A. K. Peters, Wellesley, MA. (1995). 544 pages. $59.95 (diskette included)ters, Wellesley, MA. (1995). 544 pages. Computers and Mathematics with Applications, 29(11):110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902740.

Anonymous:1995:BRMf

[Ano95-50]  Anonymous.  Book re-

view: Migrating to object technology: By Ian Graham. Addison-Wesley Publishing Company, Wokingham, England. (1995). 552 pages. $41.95 (diskette included). Computers and Mathematics with Applications, 29(11):108, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902600.

Anonymous:1995:BRMb

[Ano95-51] Anonymous. Book review: Modeling and simulation: By Hartmut Bossel. A. K. Peters, Ltd., Wellesley, MA. (1994). 484 pages. $59.95 (diskette included). Computers and Mathematics with Applications, 29(3):114, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195900896.

Anonymous:1995:BRMa

[Ano95-52] Anonymous. Book review: Motif tools: Streamlined GUI design and programming with the Xmt library: By David Flanagan. O’Reilley & Associates, Sebastopol, CA. (1994). 984 pages. $55.00. Computers and Mathematics with Applications, 29(3): 113, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195900853.

Anonymous:1995:BRNa

[Ano95-53] Anonymous. Book review: Neurocomputers: an overview of neural networks in VLSI: By Manfred Glesner and Werner Pöchmüller. Chapman & Hall, London. (1994). 281 pages. $59.95. Computers and Mathematics with Applications, 29(11):109, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219590266X.

Anonymous:1995:NQM

[Ano95-54] Anonymous. Book review:

Notes on quantum mechanics: a course given by Enrico Fermi (second edition): Compiled by Robert A. Schluter. University of Chicago Press, Chicago. (1995). 188 pages. $14.95. Computers and Mathematics with Applications, 30(9):122, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901957.

Anonymous:1995:BRNb

[Ano95-55] Anonymous. Book review:

NURB curves and surfaces: From projective geometry to practical use: By Gerald E. Farin. A. K. Peters, Wellesley, MA. (1995). 229 pages. $39.95. Computers and Mathematics with Applications, 29(11):110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902724.

Anonymous:1995:BROa

[Ano95-56]  Anonymous.  Book re-

view: Object engineering: the fourth dimension: By Philippe Desfray. Addison-Wesley Publishing Company, Wokingham, England. (1994). 342 pages. $36.95. Computers and Mathematics with Applications, 29(11):107, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902554.

Anonymous:1995:BROb

[Ano95-57] Anonymous. Book review: Object-oriented programming in Pascal: a graphical approach: By D. Brookshire Conner, David Niguidula and Andries van Dam. Addison-Wesley, Reading, MA. (1995). 616 pages. $44.06. Computers and Mathematics with Applications, 30(9):126, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902457.

Anonymous:1995:BRPa

[Ano95-58] Anonymous. Book review: PGP: Pretty good privacy: By Simson Garfinkel. O’Reilly & Associates, Sebastopol, CA. (1995). 393 pages. $19.95. Computers and Mathematics with Applications, 29 (11):108, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902562.

Anonymous:1995:BRPb

[Ano95-59] Anonymous. Book review:

PVM: Parallel virtual machine: a users’ guide and tutorial for networked parallel computing: By Al Geist, Adam Beguelin, Jack Dongarra, Weicheng Jiang, Robert Manchek and Vaidy Sunderam. MIT Press, Cambridge, MA. (1994). 279 pages. $19.95. Computers and Mathematics with Applications, 30(9):122, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901973.

Anonymous:1995:BRQ

[Ano95-60] Anonymous. Book review:

Qualitative reasoning: Modeling and simulation with incomplete knowledge: By Benjamin Kuipers. MIT Press, Cambridge, MA. (1994). 418 pages. $45.00. Computers and Mathematics with Applications, 29(3):116, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901051.

Anonymous:1995:BRRa

[Ano95-61] Anonymous. Book review:

Relational database writings 1991-1994: By C. J. Date. Addison-Wesley Publishing Company, Reading, MA. (1995). 542 pages. $45.95. Computers and Mathematics with Applications, 29 (11):106, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902503.

Anonymous:1995:BRRc

[Ano95-62] Anonymous. Book review: Repeated games with incomplete information: By Robert J. Aumann and Michael B. Maschler with the collaboration of Richard E. Stearns. MIT Press, Cambridge, MA. (1995). 342 pages. $34.95. Computers and Mathematics with Applications, 30(9): 126, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219590235X.

Anonymous:1995:BRRb

[Ano95-63] Anonymous. Book review: Running linux: By Matt Welsh and Lar Kaufman. O’Reilly & Associates, Inc., Sebastopol, CA. (1995). 575 pages. $24.95. Computers and Mathematics with Applications, 29(11):110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902708.

Anonymous:1995:BRSc

[Ano95-64] Anonymous. Book review:

Safeware: System safety and computers: By Nancy G. Leveson. Addison-Wesley, Reading, MA. (1995). 680 pages. $49.43. Computers and Mathematics with Applications, 30(9):123, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902031.

Anonymous:1995:BRSa

[Ano95-65] Anonymous. Book review:

Selected papers on automath: Edited by R. P. Nederpelt, J. H. Geuvers and R. C. De Vrijer. Elsevier, Amsterdam. (1994). 1024 pages. $214.25, Dfl. 375.00. Computers and Mathematics with Applications, 30(9):120, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901833.

Anonymous:1995:BRSb

[Ano95-66]  Anonymous.  Book re-

view: Several complex variables (Chicago lectures in mathematics): By Raghavan Narasimhan. University of Chicago Press, Chicago. (1971). 174 pages. $22.00.

Computers and Mathematics with Applications, 30(9): 123, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902090.

Anonymous:1995:BRSd

[Ano95-67] Anonymous. Book review: She does math!: Edited by Marla Parker. Mathematical Association of America, Washington, DC. (1995). 253 pages. $24.00. Computers and Mathematics with Applications, 30(9):125, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902295.

Anonymous:1995:SHG

[Ano95-68] Anonymous. Book review: Stable homotopy and generalised homology (Chicago lectures in mathematics): J. F. Adams. University of Chicago Press, Chicago (1974). 373 pages. $30.00. Computers and Mathematics with Applications, 30(9):124, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902171.

Anonymous:1995:BRTb

[Ano95-69] Anonymous. Book review: Technoscientific imaginaries: Conversations, profiles, and memoirs: Edited by George E. Marcus. University of Chicago Press, Chicago. (1995). 560 pages. $65.00 (cloth), $22.50 (paper). Computers and Mathematics with Applications, 30(9):121, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195901892.

Anonymous:1995:BRAf

[Ano95-70] Anonymous. Book review:

The antisymmetry of syntax: By Richard S. Kayne. MIT Press, Cambridge, MA. (1994). 195 pages. $17.95. Computers and Mathematics with Applications, 29 (11):108, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902627.

Anonymous:1995:BRCn

[Ano95-71]  Anonymous.  Book re-

view: The computer contradictionary (second edition): By Stan Kelly-Bootle. The MIT Press, Cambridge, MA. (1995). 239 pages. $14.95. Computers and Mathematics with Applications, 30(9): 125, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902333.

Anonymous:1995:BRCg

[Ano95-72] Anonymous. Book review: The cross-GUI handbook for multiplatform user interface design: By Aaron Marcus, Nick Smilonich and Lynne Thompson. Addison-Wesley Publishing Company, Reading, MA. (1995). 276 pages. $36.95. Computers and Mathematics with Applications, 29 (11):109, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902635.

Anonymous:1995:BRDe

[Ano95-73] Anonymous. Book review: The development of futureoriented processes: By Marshall M. Haith, Janette B. Benson, Ralph J. Roberts, Jr. and Bruce F. Pennington. University of Chicago Press, Chicago. (1994). 465 pages. $47.50 (cloth). Computers and Mathematics with Applications, 29(11):107, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219590252X.

Anonymous:1995:BRDd

[Ano95-74] Anonymous. Book review: The discourses of science: By Marcello Pera (translated by Clarissa Botsford). University of Chicago Press, Chicago. (1994). 250 pages. $29.95, £23.95. Computers and Mathematics with Applications, 29(11):106, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902511.

Anonymous:1995:BREc

[Ano95-75] Anonymous. Book review:

The electronic word, democracy, technology, and the arts: By Richard A. Lanham. University of Chicago Press, Chicago. (1993). 285 pages. $13.75, £9.50. Computers and Mathematics with Applications, 30(9):122, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902015.

Anonymous:1995:BRFe

[Ano95-76] Anonymous. Book review:

The future does not compute: Transcending the machines in our midst: By Stephen L. Talbott. O’Reilly & Associates, Inc., Sebastopol, CA. (1995). 481 pages. $22.95. Computers and Mathematics with Applications, 30(9):125, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902236.

Anonymous:1995:BRFc

[Ano95-77]  Anonymous.  Book re-

view: The future of software: Edited by Derek Leebaert. MIT Press, Cambridge, MA. (1995). 300 pages. $24.95. Computers and Mathematics with Applications, 29(11):109, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902686.

Anonymous:1995:BRGc

[Ano95-78] Anonymous. Book review: The Gödel programming language: By Patricia Hill and John Lloyd. MIT Press, Cambridge, MA. (1994). 348 pages. $45.00. Computers and Mathematics with Applications, 29(3):116, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195901132.

Anonymous:1995:BRIb

[Ano95-79] Anonymous. Book review: The ISDN literacy book: By Gerald L. Hopkins. Addison-Wesley, Reading, MA. (1995). 367 pages. $34.38. Computers and Mathematics with Applications, 30(9):126, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195902473.

Anonymous:1995:BRMe

[Ano95-80] Anonymous. Book review:

The mosaic handbook for the Macintosh: By Dale Dougherty and Richard Koman. O’Reilly & Associates, Sebastopol, CA. (1994). 171 pages. $29.95 (with diskette). Computers and Mathematics with Applications, 29 (11):108, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902570.

Anonymous:1995:BRTa

[Ano95-81] Anonymous. Book review:

Theoretical aspects of object oriented programming: Types, semantics, and language design: Edited by Carl A. Gunter and John C. Mitchell. MIT Press, Cambridge, MA. (1994). 546 pages. $45.00. Computers and Mathematics with Applications, 29(3): 114, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900934.

Anonymous:1995:BRU

[Ano95-82] Anonymous. Book review:

Using MPI: Portable parallel programming with the message-passing interface: By William Gropp, Ewing Lusk and Anthony Skjellum. MIT Press, Cambridge, MA.

(1994). 307 pages. $24.95. Computers and Mathematics with Applications, 30(9): 122, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219590199X.

Anonymous:1995:BRV

[Ano95-83] Anonymous. Book review: Variational inequalities and network equilibrium problems: Edited by F. Giannessi and A. Maugeri. Plenum Press, New York. (1995). 305 pages. $89.50. Computers and Mathematics with Applications, 30(9):121, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195901876.

Anonymous:1995:EBa

[Ano95-84] Anonymous. Editorial Board. Computers and Mathematics with Applications, 29(1): iii–v, January 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195900632.

Anonymous:1995:EBb

[Ano95-85] Anonymous. Editorial Board. Computers and Mathematics with Applications, 30 (1):iii–v, July 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195902481.

Anonymous:1995:LC

[Ano95-86] Anonymous. List of contents. Computers and Mathematics with Applications, 30 (12):iii–ix, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900535.

Anonymous:1995:LCA

[Ano95-87] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 29(12):iii–xi, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900659.

Anonymous:1996:BRCt

[Ano96a] Anonymous. Book review: A commentary on Kant’s critique of practical reason: By Lewis White Beck. University of Chicago Press, Chicago. (1960). 306 pages. $15.95, £12.75. Computers and Mathematics with Applications, 32(6):135, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902515.

Anonymous:1996:BRCu

[Ano96b] Anonymous. Book review: A compositional approach to performance modelling: By Jane Hillston. Cambridge University Press, Cambridge. (1996). 158 pages. $54.95, £37.50. Computers and Mathematics with Applications, 32(6):136, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902679.

Anonymous:1996:BRDi

[Ano96c] Anonymous. Book review: A dynamic systems approach to the development of cognition and action: By Esther Thelen and Linda Smith. MIT Press, Cambridge, MA. (1994). 376 pages. $25.00. Computers and Mathematics with Applications, 32(10):131, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900218.

Anonymous:1996:BRU

[Ano96d] Anonymous. Book review: A user’s guide to operator algebras: By Peter A. Fillmore. John Wiley & Sons, New York. (1996). 223 pages. $59.95. Computers and Mathematics with Applications, 32(6):134, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902370.

Anonymous:1996:BRAe

[Ano96e] Anonymous. Book review:

Achieving quality in software, proceedings of the third international conference on achieving quality in software, 1996: Edited by Sandro Bologna and Giacomo Bucci. Chapman & Hall, New York. (1996). 428 pages. $109.00. Computers and Mathematics with Applications, 32(10): 126, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900000.

Anonymous:1996:BRAh

[Ano96f] Anonymous. Book review:

Achieving quality software: Including its application to safety-related systems, (third edition): By David J. Smith. Chapman & Hall, New York. (1987). 285 pages. $69.95. Computers and Mathematics with Applications, 32(10): 129, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900115.

Anonymous:1996:BRAa

[Ano96g] Anonymous. Book review:

Advanced scientific Fortran:

By David R. Willé. John Wiley & Sons, Chichester. (1995). 234 pages. $24.95. Computers and Mathematics with Applications, 31(2): 128, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900772.

Anonymous:1996:BRAf

[Ano96h] Anonymous. Book review: Advances in knowledge discovery and data mining: Edited by Usama M. Fayyad, Gregory Piatetsky-Shapiro, Padhraic Smyth and Ramasamy Uthurusamy. AAAI Press/MIT Press (Menlo Park, CA/Cambridge, MA). (1996). 611 pages. $50.00. Computers and Mathematics with Applications, 32(10): 128, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900073.

Anonymous:1996:BRAc

[Ano96i] Anonymous. Book review: Advances in neural information processing systems 7: Edited by Gerald Tesauro, David Touretzky, and Todd Leen. MIT Press, Cambridge, MA. (1995). 1143 pages. $60.00. Computers and Mathematics with Applications, 31(7):144–146, April 1996.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904630.

Anonymous:1996:BRAj

[Ano96j] Anonymous. Book review:

Advances in the theory of computation and computational mathematics: Edited by Lee L. Keener. Ablex Publishing Corporation, Norwood, NJ. (1996). 243 pages. $65.00. Computers and Mathematics with Applications, 32(10):131, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690022X.

Anonymous:1996:BRAg

[Ano96k] Anonymous. Book review:

Algebraic semantics of imperative programs: By Joseph A. Goguen and Grant Malcolm. MIT Press, Cambridge, MA. (1996). 228 pages. $32.00. Computers and Mathematics with Applications, 32(10): 128, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900103.

Anonymous:1996:BRAk

[Ano96l] Anonymous. Book review:

American technological sublime: By David E. Nye.

MIT Press, CAmbridge, MA. (1995). 225 pages. $16.00. Computers and Mathematics with Applications, 32(10): 132, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900322.

Anonymous:1996:BRIh

[Ano96m] Anonymous. Book review: An information-theoretic approach to neural computing: By Gustavo Deco and Dragan Obradovic. Springer-Verlag, New York. (1996). 261 pages. $49.95. Computers and Mathematics with Applications, 32(6):134, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902412.

Anonymous:1996:BRIg

[Ano96n] Anonymous. Book review: An introduction to genetic algorithms: By Melanie Mitchell. MIT Press, Cambridge, MA. (1996). 205 pages. $30.00. Computers and Mathematics with Applications, 32(6): 133, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902278.

Anonymous:1996:BRIi

[Ano96o] Anonymous. Book review:

An introduction to highperformance scientific computing: By Lloyd D. Fosdick, Elizabeth R. Jessup, Carolyn J. C. Schauble and Gitta Domik. MIT Press, Cambridge, MA. (1996). 760 pages. $55.00. Computers and Mathematics with Applications, 32(6):135, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902576.

Anonymous:1996:BRIa

[Ano96p] Anonymous. Book review:

An introduction to neural networks: By James A. Anderson. MIT Press, Cambridge, MA. (1995). 650 pages. $55.00. Computers and Mathematics with Applications, 31(2):128, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900693.

Anonymous:1996:BRId

[Ano96q] Anonymous. Book review:

An introduction to programming with Mathematica(R) (second edition): By Richard Gaylord, Samuel Kamin and Paul Wellin. Springer-Verlag, New York. (1996). 452 pages. DM 72.00 (diskette included).

Computers and Mathematics with Applications, 31 (11):135, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873305.

Anonymous:1996:BRIf

[Ano96r] Anonymous. Book review: An introduction to symbolic dynamics and coding: By Douglas Lind and Brian Marcus. Cambridge University Press, Cambridge. (1995). 495 pages. $54.95 (cloth); $27.95 (paper). Computers and Mathematics with Applications, 31(11):139, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873457.

Anonymous:1996:BRAb

[Ano96s] Anonymous. Book review: Artificial minds: By Stan Franklin. MIT Press, Cambridge, MA. (1995). 449 pages. $30.00. Computers and Mathematics with Applications, 31(2):130, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900930.

Anonymous:1996:BRAi

[Ano96t] Anonymous. Book review: Axiomatic domain theory in categories of partial maps: By Marcelo P. Fiore. Cambridge University Press, Cambridge, U.K. (1996). 240 pages. $54.95. Computers and Mathematics with Applications, 32(10):130, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690019X.

Anonymous:1996:BRBc

[Ano96u] Anonymous. Book review:

Bandits on the information superhighway: By Daniel J. Barrett. O’Reilly & Associates, Sebastopol, CA. (1996). 229 pages. $17.95. Computers and Mathematics with Applications, 31 (11):137, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873330.

Anonymous:1996:BRBe

[Ano96v] Anonymous. Book review:

Bounding approaches to system identification: Edited by Mario Milanese, John Norton, Hélène Piet-Lahanier and Éric Walter. Plenum Press, New York. (1996). 565 pages. $135.00. Computers and Mathematics with Applications, 32(10):130, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900152.

Anonymous:1996:BRBa

[Ano96w] Anonymous. Book review: Building Internet firewalls: By D. Brent Chapman and Elizabeth D. Zwicky. O’Reilly & Associates, Sebastopol, CA. (1995). 517 pages. $29.95. Computers and Mathematics with Applications, 31 (7):141, April 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196904551.

Anonymous:1996:BRBd

[Ano96x] Anonymous. Book review: Business cycles: Theory, history, indicators, and forecasting: By Victor Zarnowitz. University of Chicago Press, Chicago. (1992). 593 pages. $32.50, £25.95. Computers and Mathematics with Applications, 32(6):132, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902199.

Anonymous:1996:BRBb

[Ano96y] Anonymous. Book review: Business data communications and networking, fifth edition: Jerry Fitzgerald and Alan Dennis. John Wiley & Sons, Inc., New York. (1996). 562 pages. $61.95. Computers and Mathematics with Applications, 31(11):137, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873342.

Anonymous:1996:BRCg

[Ano96z] Anonymous. Book review:

C++ and the object-oriented paradigm: an IS perspective: By Jan L. Harrington. John Wiley & Sons, New York. (1995). 458 pages. $42.95. Computers and Mathematics with Applications, 31(2): 131, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900991.

Anonymous:1996:BRCc

[Ano96-27] Anonymous. Book review:

C and Unix: Tools for software design: By Martin L. Barrett and Clifford H. Wagner. John Wiley & Sons, New York. (1996). 446 pages. $46.95. Computers and Mathematics with Applications, 31(2):128, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900759.

Anonymous:1996:BRCp

[Ano96-28] Anonymous. Book review:

CGI programming on the world wide web: By Shishir Gundavaram. O’Reilly & Associates, Sebastopol, CA. (1996). 433 pages. $29.95. Computers and Mathematics with Applications, 32(6): 132, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902217.

Anonymous:1996:BRCd

[Ano96-29] Anonymous. Book review: Circles: a mathematical view: By Dan Pedoe. Mathematical Association of America, Washington, DC. (1995). 102 pages. $18.95. Computers and Mathematics with Applications, 31(2):130, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900851.

Anonymous:1996:BRCw

[Ano96-30] Anonymous. Book review: City of bits: Space, place, and the infobahn: By William J. Mitchell. MIT Press, Cambridge, MA. (1995). 225 pages. $10.00. Computers and Mathematics with Applications, 32(10):127, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900061.

Anonymous:1996:BRCn

[Ano96-31] Anonymous. Book review:

COGNITION on cognition: Edited by Jacques Mehler and Susana Franck. MIT Press/ Elsevier, Cambridge, MA/ The Netherlands. (1995). 486 pages. $45.00. Computers and Mathematics with Applications, 31(11):136, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900383.

Anonymous:1996:BRCa

[Ano96-32] Anonymous. Book review:

Cognitive models of speech processing: Psycholinguistic and computational perspectives: Edited by Gerry T. M. Altmann. MIT Press, Cambridge, MA. (1990). 540 pages. $20.00. Computers and Mathematics with Applications, 31(2):125, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900516.

Anonymous:1996:BRCl

[Ano96-33] Anonymous. Book review:

Combinatorial geometry: By János Pach and Pankaj K. Agarwal. John Wiley & Sons, New York. (1995). 354 pages. $45.00. Computers and Mathematics with Applications, 31 (11):135, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873317.

Anonymous:1996:BRCm

[Ano96-34] Anonymous. Book review: Comparative approaches to cognitive science: Edited by Herbert L. Roitblat and Jean-Arcady Meyer. MIT Press, Cambridge, MA. (1995). 533 pages. $55.00. Computers and Mathematics with Applications, 31(11):136, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900413.

Anonymous:1996:BRCo

[Ano96-35] Anonymous. Book review: Computability, enumerability, unsolvability: Directions in recursion theory: Edited by S. B. Cooper, T. A. Slaman, and S. S. Wainer. Cambridge University Press, Cambridge. (1996). 347 pages. $39.95. Computers and Mathematics with Applications, 32(6):132, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902230.

Anonymous:1996:BRCi

[Ano96-36]  Anonymous.  Book re-

view: Computation and intelligence: Collected readings: Edited by George F. Luger. AAAI Press, Menlo Park, CA/MIT Press, Cambridge, MA. (1995). 735 pages. $35.00. Computers and Mathematics with Applications, 31(7):143, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904599.

Anonymous:1996:BRCaa

[Ano96-37] Anonymous. Book review:

Computation of special functions: By Shanjie Zhang and Jianming Jin. John Wiley & Sons, New York. (1996). 717 pages. $89.95. Computers and Mathematics with Applications, 32(10):131, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900206.

Anonymous:1996:BRCv

[Ano96-38]  Anonymous.  Book re-

view: Computational learning and probabilistic reasoning: Edited by A. Gammerman. John Wiley and Sons, New York. (1996). 312 pages. $40.00. Computers and Mathematics with Applications, 32(10):126, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900012.

Anonymous:1996:BRCb

[Ano96-39] Anonymous. Book review: Computational learning theory and natural learning systems; volume III: Selecting good models: Edited by Thomas Petsche, Stephen José Hanson, and Jude W. Shavlik. MIT Press, Cambridge, MA. (1995). 411 pages. $45.00. Computers and Mathematics with Applications, 31(2):127, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900590.

Anonymous:1996:BRCx

[Ano96-40] Anonymous. Book review: Computational theories of interaction and agency: Edited by Philip E. Agre and Stanley J. Rosenschein. MIT Press, Cambridge, MA. (1995). 767 pages. $50.00. Computers and Mathematics with Applications, 32(10):127, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900048.

Anonymous:1996:BRCq

[Ano96-41] Anonymous. Book review: Computer-aided analysis of difference schemes for partial differential equations: By Victor G. Ganzah and E. V. Vorozhtsov. John Wiley & Sons, New York. (1996). 458 pages. $65.00. Computers and Mathematics with Applications, 32(6):133, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902357.

Anonymous:1996:BRCe

[Ano96-42]  Anonymous.  Book re-

view: Computer crime: a crimefighter’s handbook: By David Icove, Karl Seger, and William VonStorch. O’Reilly & Associates, Sebastopol, CA. (1995). 437 pages. $24.95. Computers and Mathematics with Applications, 31(2): 131, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901054.

Anonymous:1996:BRCk

[Ano96-43] Anonymous. Book review:

Computer-integrated surgery: Technology and clinical applications: Edited by Russell H. Taylor, Stéphane Lavalléee, Grigore C. Burdea, and Ralph Mösges. MIT Press, Cambridge, MA. (1996). 736 pages. $95.00. Computers and Mathematics with Applications, 31(7):146–147, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904654.

Anonymous:1996:BRCh

[Ano96-44]  Anonymous.  Book re-

view: Computers and thought: Edited by Edward A. Feigenbaum and Julian Feldman. AAAI Press, Menlo Park, CA/MIT Press, Cambridge, MA. (1995). 535 pages. $18.00. Computers and Mathematics with Applications, 31(7):140, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196904459.

Anonymous:1996:BRCf

[Ano96-45] Anonymous. Book review: Computing in the information age, second edition: By Nancy Stern and Robert A. Stern. John Wiley & Sons, New York. (1996). 506 pages. $48.95. Computers and Mathematics with Applications, 31(2):131, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900978.

Anonymous:1996:BRCz

[Ano96-46] Anonymous. Book review: Control flow semantics: By Jaco de Bakker and Erik de Vink. MIT Press, Cambridge, MA. (1996). 564 pages. $50.00. Computers and Mathematics with Applications, 32(10):129, November 1996. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900139.

Anonymous:1996:BRCy

[Ano96-47] Anonymous. Book review:

Creative cognition: Theory, research, and applications: By Ronald A. Finke, Thomas B. Ward and Steven M. Smith. MIT Press, Cambridge, MA. (1992). 239 pages. $12.50. Computers and Mathematics with Applications, 32(10):128, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900097.

Anonymous:1996:BRDa

[Ano96-48] Anonymous. Book review:

Database security: By Silvana Castano, Maria Grazia Fugini, Giancarlo Martella and Pierangela Samarati. Addison-Wesley, Workingham, England. (1995). 456 pages. $49.44. Computers and Mathematics with Applications, 31(2):127, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900632.

Anonymous:1996:BRDe

[Ano96-49] Anonymous. Book review:

DCE security: Programming: By Wei Hu. O’Reilly & Associates, Sebastopol, CA. (1995). 364 pages. $29.95. Computers and Mathematics with Applications, 31(2): 131, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901091.

Anonymous:1996:BRDb

[Ano96-50] Anonymous. Book review: Descartes and his contemporaries: Meditations, objections, and replies: Edited by Roger Ariew and Marjorie Grene. University of Chicago Press, Chicago. (1995). 261 pages. $45, £35.95 (cloth); $17.95, £14.25 (paper). Computers and Mathematics with Applications, 31(2): 128, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900711.

Anonymous:1996:BRDh

[Ano96-51] Anonymous. Book review: Designing for the web: By Jennifer Niederst. O’Reilly & Associates, Sebastopol, CA. (1996). 165 pages. $24.95. Computers and Mathematics with Applications, 32(6): 135, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902497.

Anonymous:1996:BRDg

[Ano96-52]  Anonymous.  Book re-

view: Designing information technology in the postmodern age: From method to metaphor: By Richard Coyne. MIT Press, Cambridge, MA. (1995). 399 pages. $35.00. Computers and Mathematics with Applications, 31(7):148, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904770.

Anonymous:1996:BRDf

[Ano96-53]  Anonymous.  Book re-

view: Diagrammatic reasoning: Cognitive and computational perspectives: By Edited by B. Chandrasekaran, Janice Glasgow, and N. Hari Narayanan. AAAI Press, Menlo Park, CA/MIT Press, Cambridge, MA. (1995). 780 pages. $50.00. Computers and Mathematics with Applications, 31(7):140, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904472.

Anonymous:1996:BRDc

[Ano96-54] Anonymous. Book review:

Differential equations with MAPLE^{T M}: By Kevin R.

Coombes. Brian R. Hunt, Ronald L. Lipsman, John E. Osborn, and Garrett J. Stuck. John Wiley & Sons, New York. (1996). 232 pages. $20.95. Computers and Mathematics with Applications, 31 (2):129, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900838.

Anonymous:1996:BRDd

[Ano96-55] Anonymous. Book review: Digital mantras: the languages of abstract and virtual worlds: By Steven R. Holtzman. MIT Press, Cambridge, MA. (1994). 321 pages. $14.95. Computers and Mathematics with Applications, 31(2):130, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900954.

Anonymous:1996:BRDj

[Ano96-56] Anonymous. Book review: Dynamical systems: Differential equations, maps and chaotic behaviour: By D. K. Arrowsmith and C. Place. Chapman & Hall, New York. (1992). 330 pages. $32.95haotic behaviour. Computers and Mathematics with Applications, 32(10):132, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900358.

Anonymous:1996:BREe

[Ano96-57] Anonymous. Book review:

Einstein’s greatest blunder?: the cosmological constant and other fudge factors in the physics of the universe: By Donald Goldsmith. Harvard University Press, Cambridge, MA. (1995). 216 pages. $22.95. Computers and Mathematics with Applications, 31(2):131, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901078.

Anonymous:1996:BREl

[Ano96-58] Anonymous. Book review:

Electric words: Dictionaries, computers, and meanings: By Yorick Wilks, Brian Slator, and Louis Guthrie. MIT Press, Cambridge, MA. (1996). 289 pages. $32.50. Computers and Mathematics with Applications, 31 (11):140, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873536.

Anonymous:1996:BREo

[Ano96-59]  Anonymous.  Book re-

view: Electronic publishing on CD-ROM: By Steve Cunningham and Judson Rosebush. O’Reilly, Sebastopol, CA. (1996). 383 pages. $36.95 (CD-ROM included). Computers and Mathematics with Applications, 32(10): 130, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900176.

Anonymous:1996:BREi

[Ano96-60]  Anonymous.  Book re-

view: Elements of acoustic phoenetics, (2^{nd} edition): By Peter Ladefoged. University of Chicago Press, Chicago. (1996). 216 pages. $39.95, £31.95 (cloth); $14.95, £11.95 (paper). Computers and Mathematics with Applications, 31 (11):137, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873366.

Anonymous:1996:BREd

[Ano96-61] Anonymous. Book review: Empirical methods for artificial intelligence: By Paul R. Cohen. MIT Press, Cambridge, MA. (1995). 405 pages. $55.00. Computers and Mathematics with Applications, 31(2):130, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900875.

Anonymous:1996:BREh

[Ano96-62]  Anonymous.  Book re-

view: Enabling technologies for petaflops computing: By Thomas Sterling, Paul Messina, and Paul H. Smith. MIT Press, Cambridge, MA. (1995). 178 pages. $26.95. Computers and Mathematics with Applications, 31(7):147, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690471X.

Anonymous:1996:BREf

[Ano96-63]  Anonymous.  Book re-

view: Enrico Fermi, physicist: By Emilio Segrè. University of Chicago Press, Chicago. (1970). 276 pages. $13.95, £11.25. Computers and Mathematics with Applications, 31(2):132, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901157.

Anonymous:1996:BREp

[Ano96-64] Anonymous. Book review:

Error and the growth of experimental knowledge: By Deborah G. Mayo. University of Chicago Press, Chicago. (1996). 493 pages. $74.00, £59.25 (cloth); $29.95, £23.95 (paper). Computers and Mathematics with Applications, 32(10):132, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900280.

Anonymous:1996:BREc

[Ano96-65] Anonymous. Book review: Ethnomathematics: a multicultural view of mathematical ideas: By Marcia Ascher. Chapman & Hall, New York. (1991). 203 pages. $31.00. Computers and Mathematics with Applications, 31(2): 128, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900735.

Anonymous:1996:BREg

[Ano96-66]  Anonymous.  Book re-

view: Evolutionary game theory: By Jörgen W. Weibull. MIT Press, Cambridge, MA. (1995). 265 pages. $30.00. Computers and Mathematics with Applications, 31(2): 132, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196901236.

Anonymous:1996:BREb

[Ano96-67]  Anonymous.  Book re-

view: Evolutionary programming IV: Proceedings of the fourth annual conference on evolutionary programming: Edited by John R. McDonnell, Robert G. Reynolds, and David B. Fogel. MIT Press, Cambridge, MA. (1995). 805 pages. $65.00. Computers and Mathematics with Applications, 31(2):126, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900577.

Anonymous:1996:BREj

[Ano96-68] Anonymous. Book review:

Experiments in mathematics using Maple: By C. T. J. Dodson and E. A. Gonzalez. Springer, Berlin. (1995). 465 pages. DM 48.00. Computers and Mathematics with Applications, 31(11):138, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196900425.

Anonymous:1996:BREa

[Ano96-69] Anonymous. Book review:

Exploring general equilibrium: By Fischer Black. MIT Press, Cambridge, MA. (1995). 381 pages. $29.95. Computers and Mathematics with Applications, 31(2):125, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219690053X.

Anonymous:1996:BREn

[Ano96-70] Anonymous. Book review: Exploring Java: By Patrick Niemeyer and Joshua Peck. O’Reilly & Associates, Sebastopol, CA. (1996). 407 pages. $29.95. Computers and Mathematics with Applications, 32(6):135, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219690259X.

Anonymous:1996:BREk

[Ano96-71]  Anonymous.  Book re-

view: Exploring probability and statistics with spreadsheets: By R. Jackson and J. T. Callender. Prentice Hall, London. (1995). 236 pages. $36.00. Computers and Mathematics with Applications, 31(11):139, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873470.

Anonymous:1996:BREm

[Ano96-72] Anonymous. Book review: Extensions of first order logic: By María Manzano. Cambridge University Press, U.K. (1996). 388 pages. $59.95. Computers and Mathematics with Applications, 32(6): 134, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690245X.

Anonymous:1996:BRFb

[Ano96-73] Anonymous. Book review:

Foundations for programming languages: By John C. Mitchell. MIT Press, Cambridge, MA. (1996). 846 pages. $60.00. Computers and Mathematics with Applications, 32(10):130, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900164.

Anonymous:1996:BRFa

[Ano96-74] Anonymous. Book review:

Fundamentals of artificial neural networks: By Mohamad H. Hassoun. MIT Press, Cambridge, MA. (1995). 511 pages. $60.00. Computers and Mathematics with Applications, 31(2): 131, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901017.

Anonymous:1996:BRFc

[Ano96-75] Anonymous. Book review:

Fundamentals of numerical computing: By L. F. Shampine, R. C. Allen, Jr. and S. Pruess. John Wiley & Sons, New York. (1997). 268 pages. $72.95. Computers and Mathematics with Applications, 32(10):132, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900334.

Anonymous:1996:BRGa

[Ano96-76] Anonymous. Book review: Getting started with paradox 5.0 for windows: By Henry Gaylord, Babette Kronstadt, and David Sachs. John Wiley & Sons, New York. (1995). 196 pages. $13.95. Computers and Mathematics with Applications, 31(2):126, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900553.

Anonymous:1996:BRGc

[Ano96-77] Anonymous. Book review: Getting started with the Internet: By Joan Brady Lumpkin and Susan B. Durnbaugh. John Wiley & Sons, New York. (1995). 88 pages. $13.95. Computers and Mathematics with Applications, 31 (2):132, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901133.

Anonymous:1996:BRGb

[Ano96-78] Anonymous. Book review:

Goal-driven learning: Edited by Ashwin Ram and David B. Leake. MIT Press, Cambridge, MA. (1995). 507 pages. $55.00. Computers and Mathematics with Applications, 31(2):127, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900619.

Anonymous:1996:BRHb

[Ano96-79] Anonymous. Book review:

Handbook of algebraic topology: Edited by I. M. James. Elsevier, Amsterdam. (1995). 1324 pages. $205.75. Computers and Mathematics with Applications, 31(7):141, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904538.

Anonymous:1996:BRHf

[Ano96-80] Anonymous. Book review:

Handbook of combinatorics, volume 1: Edited by R. L. Graham, M. Grötschel, and L. Lovász. Elsevier/MIT Press, Amsterdam, Cambridge, MA. (1995). 1018 pages. $150.00. Computers and Mathematics with Applications, 31(11):138, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873391.

Anonymous:1996:BRHe

[Ano96-81] Anonymous. Book review: Handbook of combinatorics, volume 1: Edited by R. L. Graham, M. Grötschel, and L. Lovász. Elsevier/MIT Press, Amsterdam, Cambridge, MA. (1995). 1018 pages. $162.50. Computers and Mathematics with Applications, 31(11):138, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873408.

Anonymous:1996:BRHd

[Ano96-82] Anonymous. Book review: Handbook of incidence geometry: Buildings and foundations: Edited by F. Buekenhout. Elsevier, Amsterdam. (1995). 1420 pages. $229.50. Computers and Mathematics with Applications, 31 (7):148, April 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196904794.

Anonymous:1996:BRHh

[Ano96-83] Anonymous. Book review: Handbook of numerical analysis. volume IV: Finite element methods (part 2); numerical methods for solids (part 2): Edited by P. G. Ciarlet and J. L. Lions. Elsevier, Amsterdam, The Netherlands. (1996). 974 pages. $187.50, Dfl. 300.00. Computers and Mathematics with Applications, 32(6):136, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902618.

Anonymous:1996:HLV

[Ano96-84]  Anonymous.  Book re-

view: High-level vision: Object recognition and visual cognition: Shimon Ullman. MIT Press, Cambridge, MA. (1996). 412 pages. $40.00. Computers and Mathematics with Applications, 32(10): 132, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900346.

Anonymous:1996:BRHi

[Ano96-85] Anonymous. Book review:

High noon on the electronic frontier: Conceptual issues in cyberspace: By Peter Ludlow. MIT Press, Cambridge, MA. (1996). 536 pages. $30.00. Computers and Mathematics with Applications, 32(10): 129, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900127.

Anonymous:1996:BRHa

[Ano96-86]  Anonymous.  Book re-

view: High performance computing: Problem solving with parallel and vector architectures: Edited by Gary Sabot. Addison-Wesley, Reading, MA. (1995). 246 pages. $45.14. Computers and Mathematics with Applications, 31(2):127, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900656.

Anonymous:1996:BRHc

[Ano96-87] Anonymous. Book review: Holes and other superficialities: By Roberto Casati and Achille C. Varzi. MIT Press, Cambridge, MA. (1994). 256 pages. $15.00. Computers and Mathematics with Applications, 31(7):147, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196904678.

Anonymous:1996:BRHg

[Ano96-88] Anonymous. Book review: HTML: the definitive guide: By Chuck Musciano and Bill Kennedy. O’Reilly & Associates, Sebastopol, CA. (1996). 385 pages. $27.95. Computers and Mathematics with Applications, 32(6): 134, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902394.

Anonymous:1996:BRIb

[Ano96-89] Anonymous. Book review:

Improving science education: Edited by Barry J. Fraser and Herbert J. Walberg. University of Chicago Press, Chicago, IL. (1995). 233 pages. $28.00, £22.50. Computers and Mathematics with Applications, 31(2): 130, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900917.

Anonymous:1996:BRIe

[Ano96-90] Anonymous. Book review:

Inductive logic programming: From machine learning to software engineering: By Francesco Bergadano and Daniele Gunetti. MIT Press, Cambridge, MA. (1996). 240 pages. $37.50. Computers and Mathematics with Applications, 31(11):137, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873329.

Anonymous:1996:BRIc

[Ano96-91] Anonymous. Book review:

Introductory applications of partial differential equations:

With emphasis on wave propagation and diffusion: By G. L. Lamb, Jr. John Wiley & Sons, New York. (1995). 471 pages. $54.00. Computers and Mathematics with Applications, 31(2):132, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690111X.

Anonymous:1996:BRJa

[Ano96-92] Anonymous. Book review: Java in a nutshell: a desktop quick reference for Java programmers: By David Flanagan. O’Reilly & Associates, Sebastopol, CA. (1996). 438 pages. $19.95. Computers and Mathematics with Applications, 31(11):139, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873469.

Anonymous:1996:BRJb

[Ano96-93] Anonymous. Book review: JavaScript: the definitive guide, (beta Edition): By David Flanagan. O’Reilly, Sebastopol, CA. (1996). 454 pages. $29.95. Computers and Mathematics with Applications, 32(10):131, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900243.

Anonymous:1996:BRKb

[Ano96-94] Anonymous. Book review:

Knowing machines: Essays on technical change: By Donald MacKenzie. MIT Press, Cambridge, MA. (1996). 338 pages. $35.00. Computers and Mathematics with Applications, 31(11):138, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219687341X.

Anonymous:1996:BRKa

[Ano96-95] Anonymous. Book review:

Knowledge of meaning: an introduction to semantic theory: By Richard Larson and Gabriel Segal. MIT Press, Cambridge, MA. (1995). 639 pages. $40.00. Computers and Mathematics with Applications, 31(11):137, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873378.

Anonymous:1996:BRKc

[Ano96-96] Anonymous. Book review:

Krakatau: the destruction and reassembly of an island ecosystem: By Ian Thornton. Harvard University Press, Cambridge, MA. (1996). 346 pages. $39.95. Computers and Mathematics with Applications, 32(6):133, September 1996. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902333.

Anonymous:1996:BRLi

[Ano96-97] Anonymous. Book review: Languages of the mind: Essays on mental representation: By Ray Jackendoff. MIT Press, Cambridge, MA. (1992). 200 pages. $12.50. Computers and Mathematics with Applications, 32(6): 133, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902254.

Anonymous:1996:BRLc

[Ano96-98] Anonymous. Book review: Learning modelling with DE-RIVE: By M. Stewart Townend and David C. Pountney. Prentice Hall, London. (1995). 244 pages. $25.95. Computers and Mathematics with Applications, 31(2): 132, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196901194.

Anonymous:1996:BRLd

[Ano96-99] Anonymous. Book review: Learning networks: a field guide to teaching and learning online: By Linda Harasim, Starr Roxanne Hiltz, Lucio Teles, and Murray Turoff. MIT Press, Cambridge, MA. (1995). 329 pages. $35.00. Computers and Mathematics with Applications, 31(7):140, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904496.

Anonymous:1996:BRLb

[Ano96-100] Anonymous. Book review:

Lexico-logical form: a radically minimalist theory: By Michael Brody. MIT Press, Cambridge, MA. (1995). 155 pages. $20.00. Computers and Mathematics with Applications, 31(2):132, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690125X.

Anonymous:1996:BRLg

[Ano96-101] Anonymous. Book review:

Linear algebra: By Terry Lawson. John Wiley & Sons, New York. (1996). 408 pages. $67.95. Computers and Mathematics with Applications, 31 (11):139, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873421.

Anonymous:1996:BRLh

[Ano96-102] Anonymous. Book review:

Linear algebra labs with MATLAB(R) (second edition): By David R. Hill and David E. Zitarelli. Prentice Hall, Upper Saddle River, NJ. (1996). $22.33. Computers and Mathematics with Applications, 32(6):133, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690231X.

Anonymous:1996:BRLe

[Ano96-103] Anonymous. Book review: Logic programming: Proceedings of the 1995 international symposium: Edited by John Lloyd. MIT Press, Cambridge, MA. (1995). 646 pages. $75.00. Computers and Mathematics with Applications, 31(11):134, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219690036X.

Anonymous:1996:BRLa

[Ano96-104] Anonymous. Book review: Logic programming: Proceedings of the twelfth international conference on logic programming: Edited by Leon Sterling. MIT Press, Cambridge, MA. (1995). 834 pages. $80.00. Computers and Mathematics with Applications, 31(2):124–125, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900474.

Anonymous:1996:BRMd

[Ano96-105] Anonymous. Book review:

Making science: Between nature and society: By Stephen Cole. Harvard University Press, Cambridge, MA. (1992). 290 pages. $16.95. Computers and Mathematics with Applications, 31(7):148, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904812.

Anonymous:1996:BRMi

[Ano96-106] Anonymous. Book review:

Maple V: Learning guide: By K. M. Heal, M. L. Hansen and K. M. Richard. Springer-Verlag, New York. (1996). 269 pages. $24.00. Computers and Mathematics with Applications, 32(6):130, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902035.

Anonymous:1996:BRMj

[Ano96-107] Anonymous. Book review:

Maple V: Programming guide: By M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn and S. Vorkoetter. Springer-Verlag, New York. (1996). 379 pages. $34.00.

Computers and Mathematics with Applications, 32(6): 130, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902011.

Anonymous:1996:BRMm

[Ano96-108] Anonymous. Book review: Maps & civilization: Cartography in culture and society: By Norman J. W. Thrower. University of Chicago Press, Chicago. (1996). 326 pages. $55.00, £43.95 (cloth); $17.95, £14.25 (paper). Computers and Mathematics with Applications, 32(6):132, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902175.

Anonymous:1996:BRMb

[Ano96-109] Anonymous. Book review: Mastering C++. Second edition: an Introduction to C++ and object-oriented programming for C and Pascal programmers: Cay S. Horstmann. John Wiley & Sons, New York. (1996). 292 pages. $40.95. Computers and Mathematics with Applications, 31(2):131, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901030.

Anonymous:1996:BRMq

[Ano96-110] Anonymous. Book review:

Mastering excel: a problemsolving approach: By James Gips. John Wiley and Sons, New York. (1997). 324 pages. $35.95. Computers and Mathematics with Applications, 32(10):130, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900188.

Anonymous:1996:BRMg

[Ano96-111] Anonymous. Book review:

Math into LaTEX: an introduction to LaTEX and AMS-LaTEX: By G. Grätzer. Birkhäuser, Boston. (1996). 451 pages. $49.50. Computers and Mathematics with Applications, 31(11):137, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219687338X.

Anonymous:1996:BRMk

[Ano96-112] Anonymous. Book review:

Mathematical algorithms in visual basic for scientists & engineers: By Namir C. Shammas. McGraw-Hill, New York. (1996). 251 pages. $45.00 (including diskette). Computers and Mathematics with Applications, 32(6):

131, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902114.

Anonymous:1996:BRMf

[Ano96-113] Anonymous. Book review: Mathematical models for handling partial knowledge in artificial intelligence: Edited by Giulianella Coletti, Didier Dubois, and Romano Scozzafava. Plenum Press, New York. (1995). 308 pages. $89.50. Computers and Mathematics with Applications, 31 (11):136, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900401.

Anonymous:1996:BRMc

[Ano96-114] Anonymous. Book review: Meta-logics and logic programming: Edited by Krzysztof R. Apt and Franco Turini. MIT Press, Cambridge, MA. 339 pages. $32.50. Computers and Mathematics with Applications, 31(7):148, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219690485X.

Anonymous:1996:BRMp

[Ano96-115] Anonymous. Book review: Metacognition: Knowing about knowing: Edited by Janet Metcalfe and Arthur P. Shimamura. MIT Press, Cambridge, MA. (1994). 334 pages. $17.50. Computers and Mathematics with Applications, 32(10):128, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900085.

Anonymous:1996:BRMe

[Ano96-116] Anonymous. Book review:

Metamorphosis: a Guide to the World Wide Web & Electronic Commerce: Patrick G. McKeown and Richard T. Watson. John Wiley & Sons, New York. (1996). 159 pages. $17.95. Computers and Mathematics with Applications, 31(11):135, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873287.

Anonymous:1996:BRMn

[Ano96-117] Anonymous. Book review:

Modified Lagrangians and monotone maps in optimization: By E. G. Golshtein and N. V. Tretyakov. John Wiley & Sons, New York. (1996). 438 pages. $72.95. Computers and Mathematics with Applications, 32(6):136, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902710.

Anonymous:1996:BRMh

[Ano96-118] Anonymous. Book review: MPI: the compete reference: By Marc Snir, Steve Otto, Steven Huss-Lederman, David Walker, and Jack Dongarra. MIT Press, Cambridge, MA. (1996). 336 pages. $27.50. Computers and Mathematics with Applications, 31(11):140, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873494.

Anonymous:1996:BRMa

[Ano96-119] Anonymous. Book review: Multimedia programming: Objects, environments and frameworks: By Simon Gibbs and Dionysios Tsichritzis. Addison-Wesley, Wokingham, England. (1995). 323 pages. $35.45. Computers and Mathematics with Applications, 31(2):128, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219690067X.

Anonymous:1996:BRMo

[Ano96-120] Anonymous. Book review: MuPAD user’s manual: Multi-processing algebra data tool (MuPad version 1.2.2): B. Fuchssteiner, K. Drescher, A. Kemper, O. Kluge, K. Morisse, H. Naundorf, G. Oevel, F. Postel, T. Schulze, G. Siek, A. Sorgatz, W. Wiwianka and P. Zimmermann. John Wiley/ B. G. Teubner, Chichester, England/Stuttgart, Germany. (1996). 572 pages. £45.00 (2 CD-ROM disks included). Computers and Mathematics with Applications, 32(6): 136, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902655.

Anonymous:1996:BRNf

[Ano96-121] Anonymous. Book review:

Nanotechnology: Molecular speculations on global abundance: Edited by B. C. Crandall. MIT Press, Cambridge, MA. (1996). 214 pages. $17.00. Computers and Mathematics with Applications, 32(10):129, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900140.

Anonymous:1996:BRNa

[Ano96-122] Anonymous. Book review:

Neural networks and fuzzylogic control on personal computers and workstations: By Granino A. Korn. MIT Press, Cambridge, MA. (1995). 380 pages. $55.00 (with diskette). Computers and Mathematics with Applications, 31 (7):141, April 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196904514.

Anonymous:1996:BRNe

[Ano96-123] Anonymous. Book review: Nonlinear optimization and applications: Edited by G. Di Pillo and F. Giannessi. Plenum Press, New York. (1996). 367 pages. $110.00. Computers and Mathematics with Applications, 32(10): 127, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900036.

Anonymous:1996:BRNc

[Ano96-124] Anonymous. Book review: Notes on Fermat’s last theorem: By Alf van der Poorten. John Wiley & Sons, New York. (1996). 222 pages. $44.95. Computers and Mathematics with Applications, 31 (11):139, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873433.

Anonymous:1996:BRNb

[Ano96-125] Anonymous. Book review:

Numerical methods using Matlab: By John Penny and George Lindfield. Ellis Horwood, New York. (1995). 328 pages. $35.00. Computers and Mathematics with Applications, 31(7):147, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904733.

Anonymous:1996:BROa

[Ano96-126] Anonymous. Book review: Optimization and computational logic: Kenneth McAloon and Carol Trekoff. John Wiley & Sons, New York. (1996). 531 pages. $59.95. Computers and Mathematics with Applications, 32(10):131, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900267.

Anonymous:1996:BRPh

[Ano96-127] Anonymous. Book review:

Performance-based student assessment: Challenges and possibilities. Ninety-fifth yearbook of the National Society for the Study of Education, part 1: Edited by Joan Boykoff Baron and Dennie Palmer Wolf. University of Chicago Press, Chicago, IL. (1996). 318 pages. $31.00 (£24.75). Computers and Mathematics with Applications, 32(6):131, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902072.

Anonymous:1996:BRPb

[Ano96-128] Anonymous. Book review: Philosophy and AI: Essays at the interface: Edited by Robert Commins and John Pollock. MIT Press, Cambridge, MA. (1991). 304 pages. $15.00. Computers and Mathematics with Applications, 31(11):135, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873275.

Anonymous:1996:BRPc

[Ano96-129] Anonymous. Book review: Polynomial and matrix computations: Fundamental algorithms, volume 1: By Dario Bini and Victor Y. Pan. Birkhäuser, Boston. (1994). 415 pages. $64.50. Computers and Mathematics with Applications, 31(11):135, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873263.

Anonymous:1996:BRPj

[Ano96-130] Anonymous. Book review: Population dynamics in ecological space and time: Edited by Olin E. Rhodes, Jr., Ronald K. Chesser, and Michael H. Smith. University of Chicago Press, Chicago, IL. (1996). 388 pages. $50.00, £39.95 (cloth); $17.95, £14.25 (paper). Computers and Mathematics with Applications, 32(6):136, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902692.

Anonymous:1996:BRPd

[Ano96-131] Anonymous. Book review:

Practical numerical analysis: By Gwynne A. Evans. John Wiley & Sons, Chichester. (1995). 455 pages. $45.00. Computers and Mathematics with Applications, 31 (11):138, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196900437.

Anonymous:1996:BRPf

[Ano96-132] Anonymous. Book review:

Practical parallel programming: By Gregory V. Wilson. MIT Press, Cambridge, MA. (1995). 564 pages. $50.00. Computers and Mathematics with Applications, 31 (11):140, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873512.

Anonymous:1996:BRPi

[Ano96-133] Anonymous. Book review: Practical UNIX and Internet security, second edition: By Simson Garfinkel and Gene Spafford. O’Reilly & Associates, Sebastopol, CA. (1996). 971 pages. $39.95. Computers and Mathematics with Applications, 32(6): 135, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902552.

Anonymous:1996:BRPa

[Ano96-134] Anonymous. Book review: Principles and practice of constraint programming: the Newport papers: Edited by Vijay Saraswat and Pascal Van Hentenryck. MIT Press, Cambridge, MA. (1995). 475 pages. $49.95. Computers and Mathematics with Applications, 31(2):125, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900498.

Anonymous:1996:PFI

[Ano96-135] Anonymous. Book review: Proceedings: First international conference on multiagent systems, June 12–14, 1995, San Francisco, California: Victor Lesser, General Chair. AAAI Press, Menlo Park, CA/MIT Press, Cambridge, MA. (1995). 472 pages. $70.00. Computers and Mathematics with Applications, 31(7):142–143, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904575.

Anonymous:1996:BRPe

[Ano96-136] Anonymous. Book review:

Pseudorandomness and cryptographic applications: By Michael Luby. Princeton University Press, Princeton, NJ. (1996). 234 pages. $24.95, £20. Computers and Mathematics with Applications, 31 (11):139, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873482.

Anonymous:1996:BRRc

[Ano96-137] Anonymous. Book review:

Real computing made real: Preventing errors in scientific and engineering calculations: By Forman S. Acton. Princeton University Press, Princeton, NJ. (1996). 259 pages. $29.95. Computers and Mathematics with Applications, 31(7):147, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196904757.

Anonymous:1996:BRRa

[Ano96-138] Anonymous. Book review: Reasoning about knowledge: By Ronald Fagin, Joseph Y. Halpern, Yoram Moses, and Moshe Y. Vardi. MIT Press, Cambridge, MA. (1995). 477 pages. $45.00. Computers and Mathematics with Applications, 31(2):130, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900899.

Anonymous:1996:BRRd

[Ano96-139] Anonymous. Book review: Representations, targets, and attitudes: By Robert Cummins. MIT Press, Cambridge, MA. (1996). 153 pages. $25.00. Computers and Mathematics with Applications, 32(6):131, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902096.

Anonymous:1996:BRRb

[Ano96-140] Anonymous. Book review: Representing time in natural language: the dynamic interpretation of tense and aspect: By Alice G. B. ter Meulen. MIT Press, Cambridge, MA. (1995). 144 pages. $25.00. Computers and Mathematics with Applications, 31(2):132, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901170.

Anonymous:1996:BRSc

[Ano96-141] Anonymous. Book review: Science and culture; popular and philosophical essays: Hermann von Helmholtz: Edited by David Cahan. University of Chicago Press, Chicago. (1995). 418 pages. $52.00, £41.50 (cloth); $17.95, £14.25 (paper). Computers and Mathematics with Applications, 31(2): 132, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901212.

Anonymous:1996:BRSd

[Ano96-142] Anonymous. Book review:

Science at the bar: Law, science, and technology in America: By Sheila Jasanoff. Harvard University Press, Cambridge, MA. (1995). 285 pages. $29.95. Computers and Mathematics with Applications, 31(7):147, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904691.

Anonymous:1996:BRSi

[Ano96-143] Anonymous. Book review: Scientific knowledge: a sociological analysis: By Barry Barnes, David Bloor and John Henry. University of Chicago Press, Chicago, IL. (1996). 230 pages. $38.00 (cloth); $15.95 (paper). Computers and Mathematics with Applications, 32(6):135, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902539.

Anonymous:1996:BRSb

[Ano96-144] Anonymous. Book review: Scientific practice: Theories and stories of doing physics: Edited by Jed Z. Buchwald. University of Chicago Press, Chicago. (1995). 398 pages. $65.00, £51.95 (cloth); $24.95, £19.25 (paper). Computers and Mathematics with Applications, 31(2): 129, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900814.

Anonymous:1996:SPE

[Ano96-145] Anonymous. Book review: Solving problems in environmental engineering and geosciences with artificial neural networks: Farid Dowla and Leah L. Rogers. MIT Press, Cambridge, MA. (1995). 239 pages. $40.00. Computers and Mathematics with Applications, 31(11):137, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873354.

Anonymous:1996:BRSf

[Ano96-146] Anonymous. Book review:

Special functions: an introduction to the classical functions of mathematical physics: By Nico Temme. John Wiley & Sons, New York. (1996). 374 pages. $54.95. Computers and Mathematics with Applications, 31(11):134, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873251.

Anonymous:1996:BRSg

[Ano96-147] Anonymous. Book review: Speech coding: a computer laboratory textbook: By Thomas Barnwell III, Kambiz Nayebi, and Craig Richardson. John Wiley & Sons, New York. (1996). 184 pages. $37.95 (2 diskettes included). Computers and Mathematics with Applications, 31(11):139, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873445.

Anonymous:1996:BRSa

[Ano96-148] Anonymous. Book review: Standards policy for information infrastructure: Edited by Brian Kahin and Janet Abbate. MIT Press. Cambridge, MA. (1995). 653 pages. $35.00. Computers and Mathematics with Applications, 31(2):129, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196900796.

Anonymous:1996:BRSh

[Ano96-149] Anonymous. Book review: Syntactica, version 1.0 (for NEXTSTEP): By Richard K. Larson, David S. Warren, Juliana Freire de Lima e Silva, and Kostantinos Sagonas. MIT Press, Cambridge, MA. (1996). 130 pages. $25.00 (diskette included). Computers and Mathematics with Applications, 31(11):140, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873524.

Anonymous:1996:BRAd

[Ano96-150] Anonymous. Book review: The architecture of computer hardware systems software: an information technology approach: By Irv Englander. John Wiley & Sons, New York. (1996). 750 pages.

$67.95. Computers and Mathematics with Applications, 32(6):132, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902151.

Anonymous:1996:BRCj

[Ano96-151] Anonymous. Book review: The capability maturity model: Guidelines for improving the software process: Edited by Mark Paulk, Charles V. Weber, Bill Curtis, and Mary Beth Chrissis. Addison-Wesley, Reading, MA. (1995). 441 pages. $49.44. Computers and Mathematics with Applications, 31(7):143, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904617.

Anonymous:1996:BRCs

[Ano96-152] Anonymous. Book review:

The Challenger launch decision: Risky technology, culture, and deviance at NASA: By Diane Vaughan. University of Chicago Press, Chicago, IL. (1996). 575 pages. $24.95. Computers and Mathematics with Applications, 32(6):134, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902473.

Anonymous:1996:BRCr

[Ano96-153] Anonymous. Book review: The computer music tutorial: By Curtis Roads with John Strawn, Curtis Abbott, John Gordon, and Philip Greenspun. MIT Press, Cambridge, MA. (1996). 1234 pages. $50.00. Computers and Mathematics with Applications, 32(6):133, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196902291.

Anonymous:1996:BRLf

[Ano96-154] Anonymous. Book review: The little schemer (fourth edition): By Daniel P. Friedman and Matthias Felleisen. MIT Press, Cambridge, MA. (1996). 196 pages. $16.50. Computers and Mathematics with Applications, 31 (11):135, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196873299.

Anonymous:1996:BRMr

[Ano96-155] Anonymous. Book review: The map-building and exploration strategies of a simple sonar-equipped robot: an experimental, quantitative evaluation: By David Lee. Cambridge University Press, Cambridge, U.K. (1996). 228 pages. $54.95. Computers and Mathematics with Applications, 32(10):132, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900309.

Anonymous:1996:BRMl

[Ano96-156] Anonymous. Book review:

The Maple handbook, Maple V, release 4: By Darren Redfern. Springer-Verlag, New York. (1996). 495 pages. $29.00. Computers and Mathematics with Applications, 32(6):131, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902138.

Anonymous:1996:BRNd

[Ano96-157] Anonymous. Book review:

The new book of prime number records: By Paulo Ribenboim. Springer-Verlag, New York. (1996). 541 pages. $59.95. Computers and Mathematics with Applications, 32(6):134, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902436.

Anonymous:1996:BROb

[Ano96-158] Anonymous. Book review:

The origins of grammar: Evidence from early language comprehension: By Kathy Kirsh-Pasek and Roberta Michnick Golinkoff. MIT Press, Cambridge, MA. (1986). 230 pages. $30.00. Computers and Mathematics with Applications, 32(10): 132, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900310.

Anonymous:1996:BRPg

[Ano96-159] Anonymous. Book review: The practical approach^{T M} utilities for Maple^{T M}: Maple V, release 3: By Darren Redfern. Springer, New York. (1995). 312 pages. $69.00 (including diskette). Computers and Mathematics with Applications, 32(6):130, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196901996.

Anonymous:1996:SSD

[Ano96-160] Anonymous. Book review: The seasoned schemer: Daniel P. Friedman and Matthias Felleisen. MIT Press, Cambridge, MA. (1996). 210 pages. $18.50. Computers and Mathematics with Applications, 31 (11):135, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196900371.

Anonymous:1996:BRSe

[Ano96-161] Anonymous. Book review: The semantics of syntax: a minimalist approach to grammar: By Denis Bouchard. University of Chicago Press, Chicago. (1995). 525 pages. $95.00, £75.95 (cloth); $35.95, £28.75 (paper). Computers and Mathematics with Applications, 31(7):148, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904836.

Anonymous:1996:BRTd

[Ano96-162] Anonymous. Book review:

The Trouble with computers: Usefulness, usability, and productivity: By Thomas K. Landauer. MIT Press, Cambridge, MA. (1995). 425 pages. $15.00. Computers and Mathematics with Applications, 32(10):131, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900231.

Anonymous:1996:BRTb

[Ano96-163] Anonymous. Book review:

Theories of vision from Alkindi to Kepler: By David C. Lindberg. University of Chicago Press, Chicago. (1976). 324 pages. $16.95; £13.50. Computers and Mathematics with Applications, 32(6):136, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902631.

Anonymous:1996:BRTa

[Ano96-164] Anonymous. Book review: Time dependent problems and difference methods: By Bertil Gustafsson, Heinz-Otto Kreiss, and Joseph Oliger. John Wiley & Sons, New York. (1995). 642 pages. $79.95. Computers and Mathematics with Applications, 31 (11):140, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196873500.

Anonymous:1996:BRTc

[Ano96-165] Anonymous. Book review: Topics in advanced scientific computation: By R. E. Crandall. Springer-Verlag, Heidelberg, Germany. (1996). 340 pages. $49.00. Computers and Mathematics with Applications, 32(10):127, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219690005X.

Anonymous:1996:WWW

[Ano96-166] Anonymous. Book review: World Wide Web Journal: Fourth International World Wide Web Conference: O’Reilly & Associates, Sebastopol, CA. (1995). 735 pages. $39.95. Computers and Mathematics with Applications, 32(6): 130–131, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902059.

Anonymous:1996:BRW

[Ano96-167] Anonymous. Book review:

Would-be worlds: How simulation is changing the frontiers of science: By John L. Casti. John Wiley & Sons, New York. (1997). 242 pages. $24.95. Computers and Mathematics with Applications, 32(10):126, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196900024.

Anonymous:1996:EBa

[Ano96-168] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 31(1): iii–v, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196905015.

Anonymous:1996:EBb

[Ano96-169] Anonymous. Editorial Board. Computers and Mathematics with Applications, 32 (1):iii–v, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196901297.

Anonymous:1996:ENL

[Ano96-170] Anonymous. Errata to “68 new large amicable pairs”. Computers and Mathematics with Applications, 32 (7):123–127, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001617.

See [Yan94].

Anonymous:1996:EMT

[Ano96-171] Anonymous. Errata to “Modelling traffic queues at a signalized intersection with vehicle-actuated control”. Computers and Mathematics with Applications, 31 (7):137, April 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000235.

See [Alf95].

Anonymous:1996:LCAa

[Ano96-172] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 31(12):iii–xii, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196904435.

Anonymous:1996:LCAb

[Ano96-173] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 32 (12):iii–xii, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197867351.

Anonymous:1997:AIa

[Ano97a]  Anonymous.  Author in-

dex. Computers and Mathematics with Applications, 33 (12):x–xi, June 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197898902.

Anonymous:1997:AIb

[Ano97b]  Anonymous.  Author in-

dex. Computers and Mathematics with Applications, 34 (12):xi–xii, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197898094.

Anonymous:1997:BR

[Ano97c] Anonymous. Book reports. Computers and Mathematics with Applications, 34 (12):137–144, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197903072.

Anonymous:1997:BRLj

[Ano97d] Anonymous. Book review: A logical journey: From Gödel to philosophy: By Hao Wang. MIT Press, Cambridge, MA. (1996). 391 pages. $40.00. Computers and Mathematics with Applications, 33(9):148, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901553.

Anonymous:1997:BRAk

[Ano97e] Anonymous. Book review: Advanced topics in difference equations: By Ravi P. Agarwal and Patricia J. Y. Wong. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 507 pages. $245.00, Dfl. 395.00, £146.00. Computers and Mathematics with Applications, 34(10):139, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902418.

Anonymous:1997:BRAc

[Ano97f] Anonymous. Book review:

Advances in genetic programming, volume 2: Edited by Peter Angeline and Kenneth Kinnear, Jr. MIT Press, Cambridge, MA. (1996). 538 pages. $50.00. Computers and Mathematics with Applications, 33(5):129, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829331.

Anonymous:1997:BRAa

[Ano97g] Anonymous. Book review:

Advances in neural information processing systems 8: Proceedings of the 1995 conference: Edited by David S. Touretzky, Michael C. Mozer and Michael E. Hasselmo. MIT Press, Cambridge, MA. (1996). 1098 pages. $60.00. Computers and Mathematics with Applications, 33 (4):130–132, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900298.

Anonymous:1997:BRAi

[Ano97h]  Anonymous.  Book re-

view: American astronomy: Community, careers, and power, 1859-1940: By John Lankford. University of Chicago Press, Chicago. (1997). 447 pages. $65.00 (£51.95). Computers and Mathematics with Applications, 34(10):134, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901772.

Anonymous:1997:BRIb

[Ano97i] Anonymous. Book review: An introduction to distributed algorithms: By Valmir Barbosa. MIT Press, Cambridge, MA. (1996). 365 pages. $40.00. Computers and Mathematics with Applications, 33(4): 134, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900638.

Anonymous:1997:BRIj

[Ano97j] Anonymous. Book review: An introduction to Kolmogorov complexity and its applications. (second edition): By Ming Li and Paul Vitányi. Springer-Verlag, New York. (1997). 637 pages. DM 88.00, ös 642.40, sFr 77.50. Computers and Mathematics with Applications, 34(10): 137, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197902133.

Anonymous:1997:BRIh

[Ano97k] Anonymous. Book review:

An introduction to natural computation: By Dana H. Ballard. MIT Press, Cambridge, MA. (1997). 307 pages. $45.00. Computers and Mathematics with Applications, 34(10):136, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902017.

Anonymous:1997:BRAj

[Ano97l] Anonymous. Book review:

Analytical development economics: the less developed economy revisited: By Kaushik Basu. MIT Press, Cambridge, MA. (1997). 366 pages. $40.00. Computers and Mathematics with Applications, 34(10):139, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902376.

Anonymous:1997:BRAf

[Ano97m] Anonymous. Book review:

Animating calculus: Mathematica(R) notebooks for the laboratory: By Ed Packel and Stan Wagon. Springer-Verlag, New York. (1997). 292 pages. DM 54.00, öS 394.20, sFr 48.00. Computers and Mathematics with Applications, 33(7):140, April 1997. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197846111.

Anonymous:1997:BRAg

[Ano97n] Anonymous. Book review: Arrow logic and multi-modal logic: Edited by Maarten Marx, László Pólos and Michael Masuch. CSLI Publications/FoLLI, Stanford, CA. (1996). 247 pages. $59.95 (hardback), $22.95 (paperback). Computers and Mathematics with Applications, 33 (7):140, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197846123.

Anonymous:1997:BRAh

[Ano97o] Anonymous. Book review: Atlas for computing mathematical functions: an illustrated guide for practitioners, with programs in C and Mathematica: William J. Thompson. John Wiley & Sons, New York. (1997). 903 pages. $99.00 (CD included). Computers and Mathematics with Applications, 33(9):145, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901292.

Anonymous:1997:BRBa

[Ano97p] Anonymous. Book review:

Basic proof theory: By A. S. Troelstra and H. Schwichtenberg. Cambridge University Press, Cambridge, U.K. (1996). 343 pages. $49.95. Computers and Mathematics with Applications, 33(4): 129, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900274.

Anonymous:1997:BRBc

[Ano97q] Anonymous. Book review:

Building your own web conferences: By Susan B. Peck and Beverly Murray Scherf. O’Reilly, Sebastopol, CA. (1997). 246 pages. $59.95 (CD included)erf. O’Reilly, Sebastopol, CA. Computers and Mathematics with Applications, 33(9):148, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901656.

Anonymous:1997:BRCa

[Ano97r] Anonymous. Book review:

Case-based reasoning: Experiences, lessons, & future directions: Edited by David B. Leake. AAAI Press/MIT Press, Menlo Park, CA/ Cambridge, MA. (1996). 420 pages. $40.00. Computers and Mathematics with Applications, 33(4):128, February 1997. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900031.

Anonymous:1997:BRCc

[Ano97s] Anonymous. Book review: Cognition and the visual arts: By Robert L. Solso. MIT Press, Cambridge, MA. (1994). 294 pages. $17.50. Computers and Mathematics with Applications, 33(4): 134, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900651.

Anonymous:1997:BRCl

[Ano97t] Anonymous. Book review: Communication complexity and parallel computing: By Juraj Hromkovic. Springer-Verlag, Berlin. (1997). 336 pages. DM 58.00, öS 423.40, sFr 51.50. Computers and Mathematics with Applications, 34(10):137, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219790211X.

Anonymous:1997:BRCg

[Ano97u] Anonymous. Book review: Computability and complexity: From a programming perspective: By Neil D. Jones. MIT Press, Cambridge, MA. (1997). 466 pages. $45.00.

Computers and Mathematics with Applications, 33 (9):146, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901358.

Anonymous:1997:BRCn

[Ano97v]  Anonymous.  Book re-

view: Computational approaches to language acquisition: Edited by Michael R. Brent. MIT Press, Cambridge, MA. (1996). 199 pages. $20.00. Computers and Mathematics with Applications, 34(10):140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902479.

Anonymous:1997:BRCe

[Ano97w] Anonymous. Book review:

Computational economics and finance: Modeling and analysis with mathematica(R): Edited by Hal R. Varian. Springer-Verlag, New York. (1996). 468 pages. DM 88.00, öS 642.40, sFr 77.50 (diskette included). Computers and Mathematics with Applications, 33(5):129, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829367.

Anonymous:1997:BRCh

[Ano97x] Anonymous. Book review: Computer science logo style. Volume 1: Symbolic computing. (second edition): By Brian Harvey. MIT Press, Cambridge, MA. (1997). 318 pages. $85.00 (set of 3 books). Computers and Mathematics with Applications, 34(10): 134, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901711.

Anonymous:1997:BRCi

[Ano97y] Anonymous. Book review: Computer science logo style. Volume 2: Advanced techniques. (second edition): By Brian Harvey. MIT Press, Cambridge, MA. (1997). 324 pages. $85.00 (set of 3 books). Computers and Mathematics with Applications, 34(10): 134, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901735.

Anonymous:1997:BRCj

[Ano97z] Anonymous. Book review: Computer science logo style. Volume 3: Beyond programming. (second edition): By Brian Harvey. MIT Press, Cambridge, MA. (1997). 364 pages. $85.00 (set of 3 books).

Computers and Mathematics with Applications, 34(10): 134, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901759.

Anonymous:1997:BRCk

[Ano97-27] Anonymous. Book review:

Computing concepts with C++ essentials: By Cay S. Horstmann. John Wiley & Sons, New York. (1997). 669 pages. $55.95. Computers and Mathematics with Applications, 34(10):135, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901930.

Anonymous:1997:BRCd

[Ano97-28] Anonymous. Book review:

Computing tomorrow: Future research directions in computer science: Edited by Ian Wand and Robin Milner. Cambridge University Press, Cambridge, U.K. (1996). 373 pages. $39.95. Computers and Mathematics with Applications, 33(5):127, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829240.

Anonymous:1997:BRCo

[Ano97-29] Anonymous. Book review:

Conflict-controlled processes: By A. Chikrii. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 403 pages. $210.00, Dfl. 340.00, £126.00. Computers and Mathematics with Applications, 34(10):140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902492.

Anonymous:1997:BRCm

[Ano97-30] Anonymous. Book review: Conformal invariants, inequalities, and quasiconformal maps: By Glen D. Anderson, Mavina K. Vamanamurthy and Matti K. Vuorinen. John Wiley & Sons, New York. (1997). 505 pages. $69.95 (diskette included). Computers and Mathematics with Applications, 34(10): 139–140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902431.

Anonymous:1997:BRCb

[Ano97-31] Anonymous. Book review: Connectionism and the philosophy of psychology: By Terence Horgan and John Tienson. MIT Press, Cambridge, MA. (1996). 207 pages. $35.00. Computers and Mathematics with Applications, 33(4):134, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900535.

Anonymous:1997:BRCf

[Ano97-32] Anonymous. Book review:

Coordinating distributed objects: an actor-based approach to synchronization: By Svend Frølund. MIT Press, Cambridge, MA. (1996). 195 pages. $32.50. Computers and Mathematics with Applications, 33(5):132, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829549.

Anonymous:1997:BRDc

[Ano97-33] Anonymous. Book review:

Data structures and algorithms: a first course: By I. Adamson. Springer-Verlag, London. (1996). 419 pages. DM 38.00, öS 277.40, sFr 34.00. Computers and Mathematics with Applications, 33 (5):128, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829276.

Anonymous:1997:BRDe

[Ano97-34] Anonymous. Book review:

Data structures and algorithms: an object-oriented approach using Ada 95: By John Beidler. Springer-Verlag, New York. (1997). 364 pages. DM 78.00, öS 569.40, sFr 69.00. Computers and Mathematics with Applications, 33 (7):136, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197845981.

Anonymous:1997:BRDk

[Ano97-35] Anonymous. Book review: Database programming with JDBC and JAVA: By George Reese. O’Reilly, Sebastopol, CA. (1997). 224 pages. $29.95. Computers and Mathematics with Applications, 34(10): 141, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902595.

Anonymous:1997:BRDh

[Ano97-36] Anonymous. Book review: Design and engineering of intelligent communication systems: By Syed V. Ahamed and Victor B. Lawrence. Kluwer Academic Publishers, Boston. (1997). 675 pages. $135.00, Dfl. 250.00, £95.85. Computers and Mathematics with Applications, 34(10): 137–138, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902212.

Anonymous:1997:BRDg

[Ano97-37] Anonymous. Book review:

Designing incentive regulation for the telecommunications industry: By David E. M. Sappington and Dennis L. Weisman. MIT Press/ AIE Press, Cambridge, MA/ Washington, DC. (1996). 388 pages. $35.00. Computers and Mathematics with Applications, 34(10):136, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901991.

Anonymous:1997:BRDi

[Ano97-38] Anonymous. Book review:

Developing Java Beans: By Robert Englander. O’Reilly, Sebastopol, CA. (1997). 298 pages. $29.95. Computers and Mathematics with Applications, 34(10):140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902571.

Anonymous:1997:BRDd

[Ano97-39] Anonymous. Book review:

Deviant logic, fuzzy logic: Beyond the formalism: By Susan Haack. University of Chicago Press, Chicago. (1996). 291 pages. $55.00, £43.95 (cloth); $18.95, £15.25 (paper).

Computers and Mathematics with Applications, 33(5): 132, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829550.

Anonymous:1997:BRDa

[Ano97-40] Anonymous. Book review: Disciplining music: Musicology and its canons: Edited by Katherine Bergeron and Philip V. Bohlman. University of Chicago Press, Chicago. (1992). 220 pages. $15.95, £12.75. Computers and Mathematics with Applications, 33(4):128, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900079.

Anonymous:1997:BRDb

[Ano97-41] Anonymous. Book review: Domain decomposition: Parallel multilevel methods for elliptic partial differential equations: By Barry Smith, Petter E. Bjørstad and William D. Gropp. Cambridge University Press, Cambridge, U.K. (1996). 224 pages. $39.95. Computers and Mathematics with Applications, 33(4): 132, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900353.

Anonymous:1997:BRDj

[Ano97-42]  Anonymous.  Book re-

view: Dynamic macroeconomics: Instability, fluctuation, and growth in monetary economies: By Peter Flaschel, Reiner Franke and Willi Semmler. MIT Press, London England. (1997). 455 pages. $55.00. Computers and Mathematics with Applications, 34(10):141, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902637.

Anonymous:1997:BRDf

[Ano97-43] Anonymous. Book review:

Dynamic patterns: the selforganization of brain and behavior: By J. A. Scott Kelso. MIT Press, Cambridge, MA. (1995). 334 pages. $22.50. Computers and Mathematics with Applications, 33(9):147, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901516.

Anonymous:1997:BREb

[Ano97-44] Anonymous. Book review:

Elements of artificial neural networks: By Kishan Mehrotra, Chilukuri K. Mohan and Sanjay Ranka. MIT Press, Cambridge, MA. (1997). 344 pages. $55.00. Computers and Mathematics with Applications, 33(5):128, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829264.

Anonymous:1997:BREa

[Ano97-45] Anonymous. Book review: Encyclopedia of operations research and management science: Edited by Saul I. Gass and Carl M. Harris. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1996). 753 pages. $350.00, £248.50, Dfl. 650.00. Computers and Mathematics with Applications, 33(4): 129, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900134.

Anonymous:1997:BREd

[Ano97-46] Anonymous. Book review: Evolution of parallel cellular machines: the cellular programming approach: By Moshe Sipper. Springer-Verlag, Berlin. (1997). 198 pages. DM 58.00, öS 423.40, sFr 51.50. Computers and Mathematics with Applications, 34(10):136, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902091.

Anonymous:1997:BRFc

[Ano97-47] Anonymous. Book review:

Finite fields and applications: Proceedings of the third international conference, Glasgow, July 1995: Edited by S. Cohen and H. Niederreiter. Cambridge University Press, Cambridge, U.K. (1996). 401 pages. $42.95. Computers and Mathematics with Applications, 33(7):139, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846081.

Anonymous:1997:BRFb

[Ano97-48] Anonymous. Book review:

Finite fields: Encyclopedia of mathematics and its applications: By Rudolf Lidl and Harald Niederreiter. Cambridge University Press, Cambridge, U.K. (1997). 755 pages. $95.00. Computers and Mathematics with Applications, 33(7):136, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219784597X.

Anonymous:1997:BRFf

[Ano97-49] Anonymous. Book review:

Fiscal policy: Lessons from economic research: Edited by Alan J. Auerbach. MIT Press, Cambridge, MA. (1997). 475 pages. $50.00. Computers and Mathematics with Applications, 34(10):140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902455.

Anonymous:1997:BRFe

[Ano97-50] Anonymous. Book review: Foucault and his interlocutors: Edited by Arnold I. Davidson. University of Chicago Press, Chicago. (1997). 261 pages. $30.00. Computers and Mathematics with Applications, 33 (9):144, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901231.

Anonymous:1997:BRFd

[Ano97-51] Anonymous. Book review: Foundations of computational mathematics: Selected papers of a conference held at Rio de Janeiro, January 1997: Edited by Felipe Cucker and Michael Shub. Springer, Berlin. (1997). 441 pages. DM 128.00, öS 934.40, sFr 113.00. Computers and Mathematics with Applications, 33(9):144, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219790119X.

Anonymous:1997:FNN

[Ano97-52]  Anonymous.  Book re-

view: Foundations of neural networks, fuzzy systems, and knowledge engineering: Nikola K. Kasabov. MIT Press, Cambridge, MA. (1996). 550 pages. $60.00. Computers and Mathematics with Applications, 33 (7):136, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846007.

Anonymous:1997:BRAe

[Ano97-53] Anonymous. Book review:

From animals to animats 4: Proceedings of the fourth international conference on simulation of adaptive behavior: Edited by Pattie Maes, Maja J. Mataric, Jean-Arcady Meyer, Jordan Pollack and Stewart W. Wilson. MIT Press, Cambridge, MA. (1996). 644 pages. $65.00. Computers and Mathematics with Applications, 33(7): 138–139, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846044.

Anonymous:1997:BRFa

[Ano97-54] Anonymous. Book review:

Fuzzy logic: Edited by J. F. Baldwin. John Wiley and Sons, New York. (1996). 256 pages. $64.95. Computers and Mathematics with Applications, 33(4):133, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900432.

Anonymous:1997:BRGf

[Ano97-55] Anonymous. Book review: Gateways to knowledge: the role of academic libraries in teaching, learning, and research: Edited by Lawrence Dowler. MIT Press, Cambridge, MA. (1997). 240 pages. $35.00. Computers and Mathematics with Applications, 33(9):146, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901334.

Anonymous:1997:BRGi

[Ano97-56] Anonymous. Book review: Gauge theory and symplectic geometry: Edited by Jacques Hurtubise, François Lalonde and Gert Sabidussi. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 212 pages. $120.00, Dfl. 195.00, £72.00. Computers and Mathematics with Applications, 34(10): 142, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902753.

Anonymous:1997:BRGb

[Ano97-57] Anonymous. Book review:

Genetic programming: Proceedings of the first annual conference 1996: Edited by John R. Koza, David E. Goldberg, David B. Fogel and Rick L. Riolo. MIT Press, Cambridge, MA. (1996). 568 pages. $75.00. Computers and Mathematics with Applications, 33(5):126–127, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000254.

Anonymous:1997:BRGh

[Ano97-58] Anonymous. Book review:

Geometry of Lie groups: By Boris Rosenfeld. Kluwer Academic Publishers, Dordrecht. (1997). 393 pages. $195.00, Dfl. 295.00, £120.00. Computers and Mathematics with Applications, 34(10): 137, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902170.

Anonymous:1997:BRGg

[Ano97-59] Anonymous. Book review:

Geometry of nonpositively curved manifolds: By Patrick B. Eberlein. University of Chicago Press, Chicago. (1996). 449 pages. $90.00, £71.95 (cloth); $45.00, £35.95 (paper). Computers and Mathematics with Applications, 34(10):135, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901851.

Anonymous:1997:BRGe

[Ano97-60] Anonymous. Book review: Geometry of the quintic: By Jerry Shurman. John Wiley & Sons, New York. (1997). 200 pages. $39.95. Computers and Mathematics with Applications, 33(5):132, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829537.

Anonymous:1997:BRGd

[Ano97-61] Anonymous. Book review: Goodbye, Descartes: the end of logic and the search for a new cosmology of the mind: By K. Devlin. John Wiley & Sons, New York. (1997). 301 pages. Computers and Mathematics with Applications, 33(5):130, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829380.

Anonymous:1997:BRGa

[Ano97-62] Anonymous. Book review:

Graphs & digraphs, (third edition): By Gary Chartrand and Linda Lesniak. Chapman & Hall, New York. (1996). 422 pages. $59.95. Computers and Mathematics with Applications, 33(4):128, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900110.

Anonymous:1997:BRGc

[Ano97-63] Anonymous. Book review:

Growing artificial societies: Social science from the bottom up: By Joshua M. Epstein and Robert Axtell. MIT Press, Cambridge, MA. (1996). 208 pages. $18.95. Computers and Mathematics with Applications, 33(5): 127, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829239.

Anonymous:1997:BRHb

[Ano97-64] Anonymous. Book review:

Handbook of logic and language: Edited by Johan van Benthem and Alice ter Muelen. Elsevier/MIT Press, Amsterdam/Cambridge, MA. (1997). 1247 pages. $150.00. Computers and Mathematics with Applications, 33 (9):146, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901310.

Anonymous:1997:BRHc

[Ano97-65] Anonymous. Book review: Handbook of numerical analysis. Volume V: Techniques of scientific computing (Part 2): Edited by P. Ciarlet and J. L. Lions. Elsevier, Amsterdam. (1997). 818 pages. $187.50, NLG 300.00. Computers and Mathematics with Applications, 34(10):140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902510.

Anonymous:1997:BRIk

[Ano97-66]  Anonymous.  Book re-

view: Idempotent analysis and its applications: By Vassili N. Kolokoltsov and Victor P. Maslov. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 305 pages. $159.00, Dfl. 255.00, £95.00. Computers and Mathematics with Applications, 34(10):142, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902790.

Anonymous:1997:BRIf

[Ano97-67] Anonymous. Book review:

Interactive linear algebra: a laboratory course using Math-CAD(R): By Gerald Porter and David Hill. Springer-Verlag, New York. (1996). DM 69.00, öS 503.70, sFr 61.00 (2 diskettes included). Computers and Mathematics with Applications, 33 (7):136, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197845968.

Anonymous:1997:BRIg

[Ano97-68] Anonymous. Book review:

Interdisciplinary lively application projects (ILAPs): Edited by David C. Arney. Mathematical Association of America, Washington, DC. (1997). 222 pages. $27.50. Computers and Mathematics with Applications, 33(9):147, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901395.

Anonymous:1997:BRIi

[Ano97-69] Anonymous. Book review:

Internet economics: Edited by Lee W. McKnight and Joseph P. Bailey. MIT Press, Cambridge, MA. (1997). 525 pages. $35.00. Computers and Mathematics with Applications, 34(10):137, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197902194.

Anonymous:1997:BRId

[Ano97-70] Anonymous. Book review: Introduction to discrete mathematics with ISETL: By William E. Fenton and Ed Dubinsky. Springer-Verlag, New York. (1996). 194 pages. DM 64.00; öS 467.20; sFr 56.50. Computers and Mathematics with Applications, 33 (5):130, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829379.

Anonymous:1997:BRIc

[Ano97-71] Anonymous. Book review: Introduction to Maple, (second edition): By André Heck. Springer-Verlag, New York. (1996). 699 pages. DM 68.00; öS 496.40; sFr 60.00. Computers and Mathematics with Applications, 33(5):128, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829318.

Anonymous:1997:BRIa

[Ano97-72] Anonymous. Book review: Introduction to practical linear programming: By David J. Pannell. John Wiley & Sons, New York. (1997). 333 pages. $59.95 (including diskette). Computers and Mathematics with Applications, 33(4):133, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219790047X.

Anonymous:1997:BRIe

[Ano97-73] Anonymous. Book review:

Introduction to scientific programming: Computational problem solving using Maple and C: By Joseph Zachary. Springer-Verlag, New York. (1996). 380 pages. DM 78.00; öS 569.40; sFr 69.00 (diskette included). Computers and Mathematics with Applications, 33(5):131, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829446.

Anonymous:1997:BRJd

[Ano97-74] Anonymous. Book review:

Java AWT reference: By John Zukowski. O’Reilly, Sebastopol, CA. (1997). 1045 pages. $39.95. Computers and Mathematics with Applications, 34(10):135, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901875.

Anonymous:1997:BRJa

[Ano97-75] Anonymous. Book review: Java language reference: By Mark Grand. O’Reilly, Sebastopol, CA. (1997). 448 pages. $29.95. Computers and Mathematics with Applications, 33(7):139, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219784607X.

Anonymous:1997:BRJc

[Ano97-76] Anonymous. Book review: Java network programming: By Elliotte Rusty Harold. O’Reilly, Sebastopol, CA. (1997). 422 pages. $34.95. Computers and Mathematics with Applications, 33 (9):147, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901413.

Anonymous:1997:BRJb

[Ano97-77] Anonymous. Book review: Java threads: By Scott Oaks and Henry Wong. O’Reilly, Sebastopol, CA. (1997). 252 pages. $29.95. Computers and Mathematics with Applications, 33(7):140, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219784610X.

Anonymous:1997:BRJe

[Ano97-78]  Anonymous.  Book re-

view: Java virtual machine: By Jon Meyer and Troy Downing. O’Reilly, Sebastopol, CA. (1997). 426 pages. $32.95 (diskette included). Computers and Mathematics with Applications, 34(10):135, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901899.

Anonymous:1997:BRLe

[Ano97-79] Anonymous. Book review:

Laboratory experiences in group theory: a manual to be used with “exploring small groups”: By Ellen Maycock Parker. The Mathematical Association of America, Washington, DC. (1996). 81 pages. $22.00 (diskette included). Computers and Mathematics with Applications, 33(5):130, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829434.

Anonymous:1997:BRLi

[Ano97-80] Anonymous. Book review:

Language and reason: a study of Habermas’s pragmatics: By Maeve Cooke. MIT Press, Cambridge, MA. (1994). 207 pages. $15.00. Computers and Mathematics with Applications, 33(9):147, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901498.

Anonymous:1997:BRLb

[Ano97-81] Anonymous. Book review: Language and space: Edited by Paul Bloom, Mary A. Peterson, Lynn Nadel and Merrill F. Garrett. MIT Press, Cambridge, MA. (1996). 597 pages. $50.00. Computers and Mathematics with Applications, 33(4):132, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900316.

Anonymous:1997:BRLf

[Ano97-82]  Anonymous.  Book re-

view: Lectures on exceptional lie groups: Edited by Zafer Mahmud and Mamoru Mimura. University of Chicago Press, Chicago. (1996). 122 pages. $45.00, £35.95 (cloth); $19.95, £15.95 (paper). Computers and Mathematics with Applications, 33(7):136, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197845993.

Anonymous:1997:BRLa

[Ano97-83] Anonymous. Book review:

Lisp in small pieces: By Christian Queinnec (Translated by Kathleen Callaway). Cambridge University Press, Cambridge, U.K. (1996). 514 pages. $49.95. Computers and Mathematics with Applications, 33(4):129, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900237.

Anonymous:1997:BRLd

[Ano97-84] Anonymous. Book review:

Location, scheduling, design and integer programming: By Manfred Padberg and Minendra P. Rijal. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1996). 220 pages. $95.00, £67.50, Dfl. 175.00. Computers and Mathematics with Applications, 33(4):134, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900596.

Anonymous:1997:BRLc

[Ano97-85]  Anonymous.  Book re-

view: Logic-based 0–1 constraint programming: By Peter Barth. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1996). 253 pages. $98.50, £69.95, Dfl.

185.00. Computers and Mathematics with Applications, 33(4):134, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900614.

Anonymous:1997:BRLh

[Ano97-86]  Anonymous.  Book re-

view: Logic for applications: (Second edition). By Anil Nerode and Richard A. Shore. Springer, New York. (1997). 456 pages. DM 78.00, öS 569.40, sFr 69.00. Computers and Mathematics with Applications, 33(9):145, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901255.

Anonymous:1997:BRLl

[Ano97-87] Anonymous. Book review: Logic of mathematics: a modern course of classical logic: By Zofia Adamowicz and Pawel Zbierski. John Wiley & Sons, New York. (1997). 260 pages. $49.95. Computers and Mathematics with Applications, 34(10):139, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219790239X.

Anonymous:1997:BRLg

[Ano97-88] Anonymous. Book review:

Logic programming: Proceedings of ’96 joint international conference and symposium on logic programming: Edited by Michael Maher. MIT Press, Cambridge, MA. (1996). 554 pages. $85.00. Computers and Mathematics with Applications, 33(7):137, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846032.

Anonymous:1997:BRMf

[Ano97-89] Anonymous. Book review:

Mastering regular expressions: Powerful techniques for perl and other tools: By Jeffrey E. F. Friedl. O’Reilly, Sebastopol, CA. (1997). 342 pages. $29.95. Computers and Mathematics with Applications, 33(5):129, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829343.

Anonymous:1997:BRMh

[Ano97-90] Anonymous. Book review:

Mathematica 3.0: Standard add-on packages: Edited by Emily Martin. Wolfram Media/Cambridge University Press, Champaign, IL/ Cambridge, U.K. (1996). 516 pages. $59.95 (hardback); $29.95 (paperback). Computers and Mathematics with Applications, 33(5):130, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829392.

Anonymous:1997:BRMn

[Ano97-91] Anonymous. Book review: Mathematical methods for neural network analysis and design: By Richard M. Golden. MIT Press, Cambridge, MA. (1996). 419 pages. $65.00. Computers and Mathematics with Applications, 33(9):147, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901437.

Anonymous:1997:BRMb

[Ano97-92] Anonymous. Book review: Mathematica(R) in theoretical physics: Selected examples from classical mechanics to fractals: By Gerd Baumann. Springer-Verlag, Heidelberg, Germany. (1996). 348 pages. DM 88.00, sFr 77.50 (diskette included). Computers and Mathematics with Applications, 33(4):128, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900092.

Anonymous:1997:BRMd

[Ano97-93] Anonymous. Book review: Mathematics for engineers and scientists, (fifth edition): By Alan Jeffrey. Chapman & Hall, New York. (1996). 911 pages. $35.00. Computers and Mathematics with Applications, 33(4):132, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219790033X.

Anonymous:1997:BRMo

[Ano97-94] Anonymous. Book review:

Metamathematics, machines, and Gödel’s proof: By N. Shankar. Cambridge University Press, New York. (1994). 202 pages. $39.95 (hardback), $24.95 (paperback). Computers and Mathematics with Applications, 34(10):138, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902315.

Anonymous:1997:BRMc

[Ano97-95] Anonymous. Book review:

ML for the working programmer, (second edition): By Lawrence C. Paulson. Cambridge University Press, Cambridge, U.K. (1996). 476 pages. $80.00 (hardback); $32.95 (paper). Computers and Mathematics with Applications, 33(4):129, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900213.

Anonymous:1997:BRMg

[Ano97-96] Anonymous. Book review: Modeling nature: Cellular automata simulations with mathematica(R): By Richard Gaylord and Kazume Nishidate. Springer-Verlag, New York. (1996). 260 pages. DM 64.00; öS 467.20; sFr 56.50 (diskette included). Computers and Mathematics with Applications, 33(5):130, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829422.

Anonymous:1997:BRMl

[Ano97-97] Anonymous. Book review: Modern compiler implementation in C: Basic techniques: By Andrew W. Appel. Cambridge University Press, Cambridge, U.K. 398 pages. $64.95 (hardback), $29.95 (paper). Computers and Mathematics with Applications, 33(7):139, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197846056.

Anonymous:1997:BRMj

[Ano97-98] Anonymous. Book review: Modern compiler implementation in Java: Basic techniques: By Andrew W. Appel.

Cambridge University Press, Cambridge, U.K. (1997). 398 pages. $64.95 (hardback), $29.95 (paper). Computers and Mathematics with Applications, 33(7):139, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846093.

Anonymous:1997:BRMk

[Ano97-99] Anonymous. Book review:

Modern compiler implementation in ML: Basic techniques: By Andrew W. Appel. Cambridge University Press, Cambridge, U.K. 390 pages. $64.95 (hardback), $29.95 (paper). Computers and Mathematics with Applications, 33(7):139, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846068.

Anonymous:1997:BRMm

[Ano97-100] Anonymous. Book review:

Monetary policy: Edited by N. Gregory Mankiw. University of Chicago Press, Chicago. (1994). 345 pages. $50.00. Computers and Mathematics with Applications, 33(9):146, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901371.

Anonymous:1997:BRNb

[Ano97-101] Anonymous. Book review: NetActivism: How citizens use the Internet: By Edward Schwartz. O’Reilly & Associates, Sebastopol, CA. (1996). 212 pages. $24.95 (CD-ROM included). Computers and Mathematics with Applications, 33(4): 129, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900250.

Anonymous:1997:BRNd

[Ano97-102] Anonymous. Book review: NetSuccess: How real estate agents use the Internet: By Scott Kersnar. O’Reilly & Associates, Sebastopol, CA. (1996). 204 pages. $34.95 (CD included). Computers and Mathematics with Applications, 33(5):127, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829252.

Anonymous:1997:BRNh

[Ano97-103] Anonymous. Book review: NetTravel: How travelers use the Internet: By Michael Shapiro. O’Reilly, Sebastopol, CA. (1997). 284 pages. $24.95 (CD included). Computers and Mathematics with Applications, 34(10):

136, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902054.

Anonymous:1997:BRNm

[Ano97-104] Anonymous. Book review:

Networking security and standards: By Weidong Kou. Kluwer Academic Publishers, Boston, MA. (1997). 207 pages. $87.50, Dfl. 165.00, £62.25. Computers and Mathematics with Applications, 34(10):142, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219790273X.

Anonymous:1997:BRNl

[Ano97-105] Anonymous. Book review:

Networks and systems management: Platforms analysis and evaluation: By Iosif G. Ghetie. Kluwer Academic Publishers, Boston, MA. (1997). 512 pages. $125.00, Dlf. 230.00, £88.75. Computers and Mathematics with Applications, 34(10): 141–142, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902698.

Anonymous:1997:BRNf

[Ano97-106] Anonymous. Book review:

Neural networks and their applications: Edited by J. G. Taylor. John Wiley & Sons, New York. (1996). 293 pages. $69.95. Computers and Mathematics with Applications, 33(5):131, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829458.

Anonymous:1997:BRNk

[Ano97-107] Anonymous. Book review: Nonlinear dynamics, mathematical biology, and social science: Lecture notes volume IV: By Joshua M. Epstein. Addison-Wesley, Reading, MA. (1997). 164 pages. $34.38. Computers and Mathematics with Applications, 34(10):141, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902674.

Anonymous:1997:BRNj

[Ano97-108] Anonymous. Book review: Nonlinear wave dynamics: Complexity and simplicity: By Jüri Engelbrecht. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 183 pages, $109.00. Dlf. 175.00, £65.00. Computers and Mathematics with Applications, 34(10): 140, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902558.

Anonymous:1997:BRNg

[Ano97-109] Anonymous. Book review: Nonmonotonic reasoning: By Grigoris Antoniou. MIT Press, Cambridge, MA. (1997). 285 pages. $35.00. Computers and Mathematics with Applications, 34(10): 135, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901954.

Anonymous:1997:BRNi

[Ano97-110] Anonymous. Book review:

Numerica: a modeling language for global optimization: By Pascal Van Hentenryck, Laurent Michel and Yves Deville. MIT Press, Cambridge, MA. (1997). 210 pages. $25.00. Computers and Mathematics with Applications, 34(10):138, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902339.

Anonymous:1997:BRNa

[Ano97-111] Anonymous. Book review:

Numerical algorithms with C: By Gisela Engeln–Müllges and Frank Uhlig. Springer-Verlag, Heidelberg, Germany. (1996). 596 pages. DM 68.00 sFr 60.00 (CD-ROM included). Computers and Mathematics with Applications, 33(4):129, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900171.

Anonymous:1997:BRNc

[Ano97-112] Anonymous. Book review: Numerical algorithms with Fortran: By Gisela Engeln– Müllges and Frank Uhlig. Springer-Verlag, Heidelberg, Germany. (1996). 602 pages. DM 68.00, sFr 60.00 (CD-ROM included). Computers and Mathematics with Applications, 33(4):132, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900390.

Anonymous:1997:BROa

[Ano97-113] Anonymous. Book review: Operations research: an introduction: By Hamdy A. Taha. Prentice Hall, Upper Saddle River, NJ. (1997). 916 pages. $87.67 (diskette included). Computers and Mathematics with Applications, 33(5):132, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829483.

Anonymous:1997:BROb

[Ano97-114] Anonymous. Book review:

Operations research: an introduction (solutions manual): By Hamdy A. Taha. Prentice Hall, Upper Saddle River, NJ. (1997). 407 pages. Computers and Mathematics with Applications, 33(5):132, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829495.

Anonymous:1997:BRPb

[Ano97-115] Anonymous. Book review:

Parallel algorithms for regular architectures: Meshes and pyramids: By Russ Miller and Quentin F. Stout. MIT Press, Cambridge, MA. (1996). 310 pages, $40.00. Computers and Mathematics with Applications, 33(4):134, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900559.

Anonymous:1997:BRPa

[Ano97-116] Anonymous. Book review:

Partial differential equations: Methods and applications: By Robert C. McOwen. Prentice Hall, New Jersey. (1996). 419 pages. $62.00. Computers and Mathematics with Applications, 33(4):133, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900456.

Anonymous:1997:BRPc

[Ano97-117] Anonymous. Book review: Partiality, modality, and nonmonotonicity: Edited by Patrick Doherty. CSLI Publications, Stanford, CA. (1996). 300 pages. $64.95 (hardback); $22.95 (paper). Computers and Mathematics with Applications, 33(5):128, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829306.

Anonymous:1997:BRPj

[Ano97-118] Anonymous. Book review: Pattern matching algorithms: Edited by Alberto Apostolico and Zvi Galil. Oxford University Press, New York, (1997). 377 pages. $65.00. Computers and Mathematics with Applications, 34(10): 138–139, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197902352.

Anonymous:1997:BRPf

[Ano97-119] Anonymous. Book review: Perfect, amicable and sociable numbers: a computational approach: By Song Y. Yan. World Scientific, Singapore. (1996). 338 pages. $68.00. Computers and Mathematics with Applications, 33(9):147, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901450.

Anonymous:1997:BRPg

[Ano97-120] Anonymous. Book review:

Philosophy and cognitive science: By James H. Fetzer. Paragon House, New York. (1996). 191 pages. $13.95. Computers and Mathematics with Applications, 33 (9):148, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901590.

Anonymous:1997:BRPk

[Ano97-121] Anonymous. Book review:

Primes of the form x^{2} + ny^{2}: Fermat, class field theory, and complex multiplication: David A. Cox. John Wiley & Sons, New York. (1989). 351 pages. $49.95, £29.95. Computers and Mathematics with Applications, 34(10):141, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902650.

Anonymous:1997:BRPd

[Ano97-122] Anonymous. Book review:

Principles and practice of mathematics: By COMAP, Inc. Springer-Verlag, New York. (1997). 686 pages. DM 98.00; öS 715.40; sFr 86.50. Computers and Mathematics with Applications, 33(5): 128, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829288.

Anonymous:1997:BRPe

[Ano97-123] Anonymous. Book review: Programming Python: By Mark Lutz. O’Reilly, Sebastopol, CA. (1996). 880 pages. $44.95. Computers and Mathematics with Applications, 33(5):132, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829525.

Anonymous:1997:BRPh

[Ano97-124] Anonymous. Book review: Programming with GNU software: By Mike Loukides and Andy Oram. O’Reilly, Sebastopol, CA. (1997). 244 pages. $39.95 (CD included). Computers and Mathematics with Applications, 33 (9):148, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901577.

Anonymous:1997:BRQ

[Ano97-125] Anonymous. Book review:

Quaternionic and Clifford calculus for physicists and engineers: By Klaus Gürlebeck and Wolfgang Sprössig. John Wiley & Sons, Chichester, England. (1997). 371 pages. $29.95. Computers and Mathematics with Applications, 34(10):142, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902716.

Anonymous:1997:BRRb

[Ano97-126] Anonymous. Book review:

Real-time systems: Design principles for distributed embedded applications: By Hermann Kopetz. Kluwer Academic Publishers, Boston, MA. (1997). 338 pages. $72.50, Dfl. 145.00, £48.00. Computers and Mathematics with Applications, 34(10): 142, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902777.

Anonymous:1997:BRSo

[Ano97-127] Anonymous. Book review:

Sartre, Foucault, and historical reason. Volume one: Toward an existentialist theory of history: By Thomas R. Flynn. University of Chicago Press, Chicago, IL. (1997).

340 pages. $55.00, £43.95 (cloth); $18.95, £15.25 (paper). Computers and Mathematics with Applications, 34(10):141, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902613.

Anonymous:1997:BRSg

[Ano97-128] Anonymous. Book review: Science in the field: Edited by Henrika Kuklick and Robert E. Kohler. Osiris, a research journal, University of Chicago Press, Chicago, IL. (1996). 265 pages. $39.00 (cloth), $25.00 (paper). Computers and Mathematics with Applications, 33(7):136, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846019.

Anonymous:1997:SPCa

[Ano97-129] Anonymous. Book review: Selected papers on computer science: Donald E. Knuth. CSLI Publications/ Cambridge University Press/ Stanford, CA/Cambridge, U.K. (1996). 274 pages. $69.95 (hardback); $24.95 (paperback). Computers and Mathematics with Applications, 33(5):132, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829513.

Anonymous:1997:SPCb

[Ano97-130] Anonymous. Book review:

Selected papers: S. Chandrasekhar. Volume 7: The non-radial oscillations of stars in General Relativity and other writings: University of Chicago Press, Chicago. (1997). 294 pages. $99.00, £79.25 (cloth); $45.00, £35.95 (paper). Computers and Mathematics with Applications, 33(9):145, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901279.

Anonymous:1997:BRSj

[Ano97-131] Anonymous. Book review:

Shockwave studio: Designing multimedia for the web: By Bob Schmitt. O’Reilly, Sebastopol, CA. (1997). 183 pages. $39.95 (CD included). Computers and Mathematics with Applications, 33 (9):148, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901619.

Anonymous:1997:BRSm

[Ano97-132] Anonymous. Book review:

Simulation with visual SLAM and AweSim: By A. Alan B. Pritsker, Jean J. O’Reilly and David K. LaVal. John Wiley & Sonsg/Systems Publishing Corporation, New York/West Lafayette, IN. (1997). 811 pages. $89.95. Computers and Mathematics with Applications, 34(10):137, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902157.

Anonymous:1997:BRSn

[Ano97-133] Anonymous. Book review: Slaves of the machine: the quickening of computer technology: By Gregory J. E. Rawlins. MIT Press, Cambridge, MA. (1997). 135 pages. $25.00. Computers and Mathematics with Applications, 34(10):138, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902297.

Anonymous:1997:BRSl

[Ano97-134] Anonymous. Book review: Solving the frame problem: a mathematical investigation of the common sense law of inertia: By Murray Shanahan. MIT Press, Cambridge, MA. (1997). 407 pages. $45.00. Computers and Mathematics with Applications, 34(10): 135–136, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901978.

Anonymous:1997:BRSb

[Ano97-135] Anonymous. Book review:

Spaces, worlds, and grammar: Edited by Gilles Fauconnier and Eve Sweetser. University of Chicago Press, Chicago. (1996). 355 pages. $65.00, £51.95 (cloth); $24.95, £19.95 (paper). Computers and Mathematics with Applications, 33(4):133, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900419.

Anonymous:1997:BRSk

[Ano97-136] Anonymous. Book review: Specification of abstract data types: By Jacques Loeckx, Hans-Dieter Ehrich and Markus Wolf. Wiley/ Teubner, Chichester, U.K./ Stuttgart, Germany. 260 pages. $60.00. Computers and Mathematics with Applications, 34(10):134, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901814.

Anonymous:1997:BRSa

[Ano97-137] Anonymous. Book review:

Structure and interpretation of computer programs, (second edition): By Harold Abelson and Gerald Jay Sussman with Julie Sussman. MIT Press/McGraw-Hill, Cambridge, MA/New York. (1996). 657 pages. $55.00. Computers and Mathematics with Applications, 33(4):133, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900511.

Anonymous:1997:BRSh

[Ano97-138] Anonymous. Book review: Symbolic integration I: Transcendental functions: By Manuel Bronstein. Springer-Verlag, New York. (1997). 299 pages. DM 78.00, öS 569.40, sFr 69.00. Computers and Mathematics with Applications, 33(7):136, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197845956.

Anonymous:1997:BRTb

[Ano97-139] Anonymous. Book review: Telecommunications competition: the last ten miles: By Ingo Vogelsang and Bridger M. Mitchell. MIT Press/AEI Press, London/ Washington, DC. (1997). 364 pages. $35.00. Computers and Mathematics with Applications, 34(10):134, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901796.

Anonymous:1997:BRAb

[Ano97-140] Anonymous. Book review: The W_{3} algebra: Modules, semi-infinite cohomology and BV algebras: By Peter Bouwknegt, Jim McCarthy and Krzysztof Pilch. Springer-Verlag, Heidelberg, Germany. (1996). 204 pages. DM 62.00, sFr 55.00. Computers and Mathematics with Applications, 33(4):132, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900377.

Anonymous:1997:BRAd

[Ano97-141] Anonymous. Book review:

The ADI model problem: By E. Wachspress. (1995). Copies of this monograph may be purchased by sending $25.00 (plus $3.00 for overseas mailing) to the author at 9802 Montego Ct., Windsor, CA 95492, U.S.A. Computers and Mathematics with Applications, 33(5):131, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219782946X.

Anonymous:1997:BRBb

[Ano97-142] Anonymous. Book review:

The b-book: Assigning programs to meanings: By J.- R. Abrial. Cambridge University Press, Cambridge, U.K. (1996). 779 pages. $64.95. Computers and Mathematics with Applications, 33(4): 133, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900493.

Anonymous:1997:BREc

[Ano97-143] Anonymous. Book review: The educated mind: How cognitive tools shape our understanding: By Kieran Egan. University of Chicago Press, Chicago, IL. (1997). 299 pages. $24.95; £19.95. Computers and Mathematics with Applications, 33(9):148, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901632.

Anonymous:1997:BRHa

[Ano97-144] Anonymous. Book review: The human semantic potential: Spatial language and constrained connectionism: By T. Regier. MIT Press, Cambridge, MA. (1996). 220 pages. $37.50. Computers and Mathematics with Applications, 33(4):134, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900572.

Anonymous:1997:BRJf

[Ano97-145] Anonymous. Book review:

The joy of C: By Lawrence H. Miller and Alexander E. Quilici. John Wiley & Sons, New York. (1997). 788 pages. $58.95 (diskette included). Computers and Mathematics with Applications, 34(10):138, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219790225X.

Anonymous:1997:BRLk

[Ano97-146] Anonymous. Book review:

The limits of medicine: How science shapes our hope for the cure: By Edward Golub. University of Chicago Press, Chicago. (1994). 258 pages. $14.95. Computers and Mathematics with Applications, 34(10):135, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901917.

Anonymous:1997:BRMi

[Ano97-147] Anonymous. Book review:

The Maple(R) O.D.E. lab book: By Darren Redfern and Edgar Chandler. Springer-Verlag, New York. (1996). 160 pages. DM 44.00, öS 321.20, sFr 39.50 (diskette included). Computers and Mathematics with Applications, 33(7):137, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197846020.

Anonymous:1997:BRMe

[Ano97-148] Anonymous. Book review: The media equation: How people treat computers, television, & new media like real people & places: By Byron Reeves and Clifford Nass. CSLI Publications/Cambridge University Press, Stanford, CA/ Cambridge, U.K. (1996). 305 pages. $27.95. Computers and Mathematics with Applications, 33(5):128, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219782929X.

Anonymous:1997:BRMa

[Ano97-149] Anonymous. Book review: The mind-brain continuum: Sensory processes: Edited by Rodolfo Llinás and Patricia S. Churchland. MIT Press, Cambridge, MA. (1996). 315 pages. $50.00. Computers and Mathematics with Applications, 33(4):128, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900055.

Anonymous:1997:BRNe

[Ano97-150] Anonymous. Book review:

The new hacker’s dictionary, (third edition): Compiled by Eric S. Raymond. MIT Press, Cambridge, MA. (1996). 547 pages. $16.50. Computers and Mathematics with Applications, 33(5):128, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219782932X.

Anonymous:1997:BRPi

[Ano97-151] Anonymous. Book review: The practical imagination: the German sciences of state in the nineteenth century: By David F. Lindenfeld. University of Chicago Press, Chicago, IL. (1997). 382 pages. $57.00, £45.50 (cloth); $19.95, £15.95 (paper). Computers and Mathematics with Applications, 34(10):136, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902078.

Anonymous:1997:BRRa

[Ano97-152] Anonymous. Book review:

The resolution calculus: By Alexander Leitsch. Springer, Berlin. (1997). 300 pages. DM 58.00, öS 423.40, sFr 51.50. Computers and Mathematics with Applications, 34(10):134, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901838.

Anonymous:1997:BRSe

[Ano97-153] Anonymous. Book review: The sciences of the artificial, (third edition): By Herbert A. Simon. MIT Press, Cambridge, MA. (1996). 231 pages. $14.50. Computers and Mathematics with Applications, 33(5):130, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829410.

Anonymous:1997:BRSi

[Ano97-154] Anonymous. Book review: The scientific enterprise in early modern Europe: Readings from Isis: Edited by Peter Dear. University of Chicago Press, Chicago, IL (1997). 335 pages. $37.50 (cloth); $17.95 (paper). Computers and Mathematics with Applications, 33(9):144, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197901218.

Anonymous:1997:BRSf

[Ano97-155] Anonymous. Book review:

The scientific revolution: By Steven Shapin. University of Chicago Press, Chicago. (1996). 218 pages. $19.95. Computers and Mathematics with Applications, 33(5): 132, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829501.

Anonymous:1997:BRSc

[Ano97-156] Anonymous. Book review:

The six core theories of modern physics: By Charles Stevens. MIT Press, Cambridge, MA. (1995). 233 pages. $12.50. Computers and Mathematics with Applications, 33(4):134, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900675.

Anonymous:1997:BRSp

[Ano97-157] Anonymous. Book review:

The structure of classical diffeomorphism groups: By Augustin Banyaga. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 197 pages. $112.00, Dfl. 180.00, £67.00.95. Computers and Mathematics with Applications, 34(10):142, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197902832.

Anonymous:1997:BRSd

[Ano97-158] Anonymous. Book review: The structure of scientific revolutions, (third edition): By Thomas S. Kuhn. University of Chicago Press, Chicago. (1996). 212 pages. $10.95. Computers and Mathematics with Applications, 33 (5):129, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829355.

Anonymous:1997:BRWb

[Ano97-159] Anonymous. Book review: The way of Z: Practical programming with formal methods: By Jonathan Jacky. Cambridge University Press, Cambridge, U.K. (1997). 350 pages. $59.95 (hardback); $29.95 (paperback). Computers and Mathematics with Applications, 33(5):131, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829471.

Anonymous:1997:BRTa

[Ano97-160] Anonymous. Book review: Turtles, termites, and traffic jams: Explorations in massively parallel microworlds: By Mitchel Resnick. MIT Press, Cambridge, MA. (1994). 163 pages. $12.50. Computers and Mathematics with Applications, 33(9):147, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901474.

Anonymous:1997:BRU

[Ano97-161] Anonymous. Book review:

Universal service: Competition, interconnection, and monopoly in the making of the American Telephone System: By Milton L. Mueller, Jr. MIT Press/AEI Press, London/Washington, DC. (1997). 213 pages. $40.00. Computers and Mathematics with Applications, 33(9):148, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219790153X.

Anonymous:1997:BRVb

[Ano97-162] Anonymous. Book review:

Vicious circles: On the mathematics of non-wellfounded phenomena: By Jon Barwise and Lawrence Moss. CSLI Publications, Stanford, CA. (1996). 390 pages. $49.95 (hardback); $24.95 (paperback). Computers and Mathematics with Applications, 33 (5):130, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197829409.

Anonymous:1997:BRVc

[Ano97-163] Anonymous. Book review: Video database systems: Issues, products and applications: By Ahmed K. Elmagarmid, Haitao Jiang, Abdelsalam A. Helal, Anupam Joshi and Magdy Ahmed. Kluwer Academic Publishers, Boston, MA. (1997). 178 pages, $96.00, Dfl. 180.00, £68.25. Computers and Mathematics with Applications, 34(10):140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197902534.

Anonymous:1997:BRVa

[Ano97-164] Anonymous. Book review: Visualization and optimization: By Christopher V. Jones. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1996). 434 pages. $129.25, £92.50, Dfl. 240.00. Computers and Mathematics with Applications, 33(4):129, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197900158.

Anonymous:1997:BRWa

[Ano97-165] Anonymous. Book review: Weak and measure-valued solutions to evolutionary PDEs: By J. Málek, J. Nečas, M. Rokyta and M. Rokyta and M. Ružička. Chapman & Hall, New York. (1996). 317 pages. $55.00. Computers and Mathematics with Applications, 33(4):129, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900195.

Anonymous:1997:BRWc

[Ano97-166] Anonymous. Book review:

Web client programming with perl: By Clinton Wong. O’Reilly, Sebastopol, CA. (1997). 213 pages. $29.95. Computers and Mathematics with Applications, 34(10): 136, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902030.

Anonymous:1997:BRWe

[Ano97-167] Anonymous. Book review: Web security & commerce: By Simson Garfinkel. O’Reilly, Sebastopol, CA. (1997). 483 pages. $32.95. Computers and Mathematics with Applications, 34(10): 138, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197902236.

Anonymous:1997:BRWf

[Ano97-168] Anonymous. Book review: Windows annoyances: By David A. Karp. O’Reilly, Sebastopol, CA. (1997). 285 pages. $29.95. Computers and Mathematics with Applications, 34(10):142, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197902819.

Anonymous:1997:BRWd

[Ano97-169] Anonymous. Book review: Windows NT in a nutshell: a desktop quick reference for system administrators: By Eric Pearce. O’Reilly, Sebastopol, CA. (1997). 348 pages. $19.95. Computers and Mathematics with Applications, 34(10):138, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197902273.

Anonymous:1997:Ea

[Ano97-170] Anonymous. Editorial. Computers and Mathematics with Applications, 34(11): xiii, December 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002125.

Anonymous:1997:Ec

[Ano97-171] Anonymous. Editorial.

Computers and Mathematics with Applications, 34(11): xv, December 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002137.

Anonymous:1997:EBa

[Ano97-172] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 33(1– 2):iii–v, January 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197901693.

Anonymous:1997:EBb

[Ano97-173] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 34 (1):iii–v, July 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197900699.

Anonymous:1997:LCa

[Ano97-174] Anonymous. List of contents. Computers and Mathematics with Applications, 33 (12):iii–ix, June 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197898896.

Anonymous:1997:LCb

[Ano97-175] Anonymous. List of contents. Computers and Mathematics with Applications, 34 (12):iii–x, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197898082.

Anonymous:1998:AIa

[Ano98a]  Anonymous.  Author in-

dex. Computers and Mathematics with Applications, 35(12):xiii–xiv, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001023.

Anonymous:1998:AIb

[Ano98b] Anonymous. Author index. Computers and Mathematics with Applications, 36(10–12): xi–xii, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219880045X.

Anonymous:1998:BRBe

[Ano98c] Anonymous. Book report book report. Computers and Mathematics with Applications, 36(4):121–130, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001448.

Anonymous:1998:BRh

[Ano98d] Anonymous. Book reports.

Computers and Mathematics with Applications, 35 (4):131–134, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002964.

Anonymous:1998:BRj

[Ano98e] Anonymous. Book reports.

Computers and Mathematics with Applications, 36(2): 129–134, July 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001229.

Anonymous:1998:BRLd

[Ano98f] Anonymous. Book review:

A little Java, a few patterns: By Matthias Felleisen and Daniel P. Friedman. MIT Press, Cambridge, MA. (1998). $25.00. Computers and Mathematics with Applications, 35(9):142, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907039.

Anonymous:1998:BRAa

[Ano98g] Anonymous. Book review:

Advanced transaction models and architectures: Edited by Sushil Jajodia and Larry Kerschberg. Kluwer Academic Publishers, Boston. (1997).

381. pages. $135.00, NLG 285.00, £91.80. Computers and Mathematics with Applications, 35(5):132, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802222.

Anonymous:1998:BRAc

[Ano98h] Anonymous. Book review: Advances in neural information processing systems 9: Proceedings of the 1996 conference: Edited by Michael C. Mozer, Michael I. Jordan and Thomas Petsche. MIT Press, Cambridge, MA. (1997). 1098 pages. $70.00. Computers and Mathematics with Applications, 35(6):154–156, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198904990.

Anonymous:1998:BRAg

[Ano98i]  Anonymous.  Book re-

view: Algebraic model theory: Edited by Bradd T. Hart, Alistair H. Lachlan, and Matthew A. Valeriote. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 277 pages. NLG 210.00, $124.00, GBP 76.00. Computers and Mathematics with Applications, 36 (3):128–129, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890219X.

Anonymous:1998:BRAf

[Ano98j] Anonymous. Book review:

Algorithmic and computer methods for three-manifolds: By A. T. Fomenko and S. V. Matveev. Kluwer Academic Publishers, Dordrecht, The Netherlands. 334 pages. NLG 320.00, $182.00, GBP 109.00. Computers and Mathematics with Applications, 36 (3):127, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902127.

Anonymous:1998:BRAj

[Ano98k] Anonymous. Book review:

Algorithmic geometry: Jean-Daniel Boissonnat and Mariette Yvinec. Cambridge University Press, New York. (1998). 519 pages. $95.00 (hardback), $39.95 (paperback). Computers and Mathematics with Applications, 36 (8):123, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911476.

Anonymous:1998:BRAh

[Ano98l]  Anonymous.  Book re-

view: Analog and mixedsignal hardware description language: Edited by A. Vachoux, Jean-Michel Bergé, O. Levia, and Jacques Rouillard. Kluwer Academic Publishers, Boston, MA. (1997). 158 pages. NLG 285.00, $135.00, GBP 91.80. Computers and Mathematics with Applications, 36(3):129, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902218.

Anonymous:1998:BRAb

[Ano98m] Anonymous. Book review: Annotated bibliographies in combinatorial optimization: Edited by Mauro Dell’Amico, Francesco Maffioli, Silvano Martello John Wiley & Sons, Chichester, England. (1997). 471 pages. $60.00. Computers and Mathematics with Applications, 35(6):154, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198904977.

Anonymous:1998:BRAi

[Ano98n]  Anonymous.  Book re-

view: Applications of continuous mathematics to computer science: By Hung T. Nguyen and Vladik Kreinovich. Kluwer Academic Publishers, Dordrecht. (1997). 419 pages. $199.00, NLG 340.00, GBP 122.00.

Computers and Mathematics with Applications, 36(3): 130, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902292.

Anonymous:1998:BRAe

[Ano98o] Anonymous. Book review:

Applied asymptotic methods in nonlinear oscillations: By Yu. A. Mitropolskii and Nguyen Van Dao. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 341 pages. NLG 280,00, $165.00, GBP 99.00. Computers and Mathematics with Applications, 36(3): 127, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902115.

Anonymous:1998:BRAk

[Ano98p] Anonymous. Book review:

Atlas for computing mathematical functions: an illustrated guide for practitioners, with programs in Fortran 90 and Mathematica: By William J. Thompson. John Wiley & Sons, Inc., New York. (1997). $145.00 (CD-ROM included). Computers and Mathematics with Applications, 36(8):125, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911671.

Anonymous:1998:BRAd

[Ano98q] Anonymous. Book review: Automata and computability: By Dexter C. Kozen. Springer-Verlag, New York. (1997) 400 pages. $39.95, GDM 64.00, GBP 24.50. Computers and Mathematics with Applications, 35 (9):144, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198907258.

Anonymous:1998:BRBf

[Ano98r] Anonymous. Book review: Barriers and challenges in computational fluid dynamics: Edited by V. Venkatakrishnan, Manuel D. Salas and Sukumar R. Chakravarthy. Kluwer Academic Publishers, Dordrecht. (1998). 395 pages. $183.00, NLG 320.00, GBP 109.00. Computers and Mathematics with Applications, 36(8):119, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911178.

Anonymous:1998:BRBa

[Ano98s] Anonymous. Book review: Basic simple type theory: By J. Roger Hindley. Cambridge University Press, New York. (1997). 186 pages, $39.95. Computers and Mathematics with Applications, 35(5): 133, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802283.

Anonymous:1998:BRDa

[Ano98t] Anonymous. Book review: Be developer’s guide: By the Be Development Team. O’Reilly, Sebastopol, CA. (1997). 934 pages. $49.95 (CD-ROM included). Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802386.

Anonymous:1998:BRBb

[Ano98u] Anonymous. Book review:

Bioinformatics: the machine learning approach: By Pierre Baldi and Soren Brunak. MIT Press, Cambridge, MA. (1998). 351 pages. $40.00. Computers and Mathematics with Applications, 36(3): 124, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122198901948.

Anonymous:1998:BRBg

[Ano98v]  Anonymous.  Book re-

view: Black holes and relativistic stars: Edited by Robert M. Wald. University of Chicago Press, Chicago. (1998). 278 pages. $50.00, £39.95. Computers and Mathematics with Applications, 36(8):122, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911439.

Anonymous:1998:BRBc

[Ano98w] Anonymous. Book review: Brauer groups, Hopf algebras and Galois theory: By Stefaan Caenepeel. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1998). 488 pages. $197.00, NLG 345.00, GBP 118.00. Computers and Mathematics with Applications, 36(3):124–125, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198901985.

Anonymous:1998:BRBd

[Ano98x] Anonymous. Book review: Buffering techniques for delivery of compressed video in video-on-demand systems: By Wu-chi Feng. Kluwer Academic Publishers, Boston, MA. (1997). 132 pages. NLG 195.00, $88.50, GBP 58.50.

Computers and Mathematics with Applications, 36(3): 128, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902188.

Anonymous:1998:BRCf

[Ano98y] Anonymous. Book review:

Codes and algebraic curves: By Oliver Pretzel. Clarendon Press, Oxford, UK. (1998). 192 pages. $65.00. Computers and Mathematics with Applications, 36(3):127, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902103.

Anonymous:1998:BRCe

[Ano98z]  Anonymous.  Book re-

view: Codesign for real-time video applications: By Jörg Wilberg. Kluwer Academic Publishers, Boston, MA. (1997). 191 pages. $98.00, NLG 215.00, GBP 64.70. Computers and Mathematics with Applications, 36(3): 124, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122198901973.

Anonymous:1998:BRCg

[Ano98-27] Anonymous. Book review:

Cognitive ecology: the evolutionary ecology of information processing and decision making: Edited by Reuven Dukas. University of Chicago Press, Chicago, IL. (1998). 420 pages. $95.00, GBP 75.95 hardback, $30.00, GBP 23.95 paperback. Computers and Mathematics with Applications, 36(3):129, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902206.

Anonymous:1998:BRCc

[Ano98-28]  Anonymous.  Book re-

view: Cognitive neuroscience: Edited by Michael D. Rugg. MIT Press, Cambridge, MA. (1997). 390 pages. $25.00. Computers and Mathematics with Applications, 35 (9):141, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906976.

Anonymous:1998:BRCb

[Ano98-29] Anonymous. Book review: Combinatorial optimization: By William J. Cook, William H. Cunningham, William R. Pulleyblank and Alexander Schrijver. John Wiley & Sons, New York. (1998). 355 pages. $49.95. Computers and Mathematics with Applications, 35(9):139, May 1998. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906836.

Anonymous:1998:BRCn

[Ano98-30] Anonymous. Book review:

Combinatorial theory (second edition): By Marshall Hall, Jr. John Wiley & Sons, New York. (1983). 440 pages. $59.95. Computers and Mathematics with Applications, 36(8):124, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911579.

Anonymous:1998:BRCj

[Ano98-31] Anonymous. Book review:

Compiler technology: Tools, translators and language implementation: By Derek Beng Kee Kiong. Kluwer Academic Publishers, Boston, MA. (1997). 210 pages. $97.50, NLG 210.00, GBP 64.35. Computers and Mathematics with Applications, 36(6): 143, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901031.

Anonymous:1998:BRCk

[Ano98-32] Anonymous. Book review:

Computational complexity and feasibility of data processing and interval computations: By Vladik Kreinovich, Anatoly Lakeyev, JiríRohn and Patrick Kahl. Kluwer Academic Publishers, Dordrecht. (1998). 459 pages. $214.00, NLG 375.00, GBP 128.00. Computers and Mathematics with Applications, 36 (8):118, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001874.

Anonymous:1998:BRCi

[Ano98-33] Anonymous. Book review: Computational geometry: Algorithms and applications: By Mark de Berg, Marc van Kreveld, Mark Overmars and Otfried Schwarzkopf. Springer, Berlin. (1997). 365 pages. $32.95. Computers and Mathematics with Applications, 36(6):139, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198900955.

Anonymous:1998:BRCa

[Ano98-34] Anonymous. Book review: Computers and design in context: Edited by Morten Kyng and Lars Mathiassen. MIT Press, Cambridge, MA. (1997). 418 pages. $40.00. Computers and Mathematics with Applications, 35(5): 132, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802234.

Anonymous:1998:BRCl

[Ano98-35] Anonymous. Book review:

Concurrency in Ada (second edition): By Alan Burns and Andy Wellings. Cambridge University Press, New York. (1998). 390 pages. $32.95. Computers and Mathematics with Applications, 36(8): 123, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219891149X.

Anonymous:1998:BRCm

[Ano98-36] Anonymous. Book review:

Connectedness and necessary conditions for an extremum: By Alexander P. Abramov. Kluwer Academic Publishers, Dordrecht. (1996). 199 pages. $99.00, NLG 175.00, GBP 59.00. Computers and Mathematics with Applications, 36(8):124, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911592.

Anonymous:1998:BRCh

[Ano98-37] Anonymous. Book review:

Constraint-based design recovery for software reengineering: Theory and experiments: By Steven G. Woods, Alexander E. Quilici, and Qiang Yang. Kluwer Academic Publishers, Boston, MA. (1998). 188 pages. $120.00, NLG 260.00, GBP 79.50. Computers and Mathematics with Applications, 36 (3):130, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902243.

Anonymous:1998:BRCd

[Ano98-38]  Anonymous.  Book re-

view: Coordinating technology: Studies in the international standarization of telecommunications: By S. K. Schmidt and Raymund Werle. MIT Press, Cambridge, MA. (1998). 365 pages. $35.00. Computers and Mathematics with Applications, 35(9):143, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907155.

Anonymous:1998:BRDb

[Ano98-39] Anonymous. Book review: Dealing with risk: Why the public and the experts disagree on environmental issues: By H. Margolis. University of Chicago Press, Chicago. (1996). 227 pages. $15.95, £12.75. Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802362.

Anonymous:1998:BRDf

[Ano98-40] Anonymous. Book review:

Developing business objects: Edited by Andy Carmichael. Cambridge University Press, Cambridge, UK. (1998). 315 pages. $39.95. Computers and Mathematics with Applications, 36(3):129, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890222X.

Anonymous:1998:BRDe

[Ano98-41]  Anonymous.  Book re-

view: Developing business systems with CORBA: By Waqar Sadiq and Fred Cummins. Cambridge University Press, NY. (1998). 261 pages. $44.95 (Includes CD-ROM). Computers and Mathematics with Applications, 36(3): 126, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902048.

Anonymous:1998:BRDd

[Ano98-42] Anonymous. Book review:

Differential equations with Mathematica^{T M}: By Kevin R. Coombes, Brian R. Hunt, Ronald L. Lipsman, John E. Osborn and Garrett J. Stuck. John Wiley & Sons, Inc., New York (1998). 240 pages. $19.95. Computers and Mathematics with Applications, 35(9):144, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907192.

Anonymous:1998:BRDi

[Ano98-43] Anonymous. Book review: Digital Signal Processing Algorithms: Number Theory, Convolution, Fast Fourier Transforms, and Applications: By Hari Krishna Garg. CRC Press, Boca Raton, FL. (1998). 637 pages. $79.95; DM 166.00, öS 1.212.00, sFr 150.00, GBP 64.00. Computers and Mathematics with Applications, 36(8):123, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911452.

Anonymous:1998:BRDc

[Ano98-44] Anonymous. Book review: Discovering curves and surfaces with Maple(R): By Grazyna Klimek and Maciej Klimek. Springer, New York. (1997). 217 pages. $49.00, DM 86.00, öS 627.80, sFr 78.50, GBP 28.00 (diskette included). Computers and Mathematics with Applications, 35(5):135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890743X.

Anonymous:1998:BRDh

[Ano98-45]  Anonymous.  Book re-

view: Distributed and parallel databases: Edited by Jeffrey F. Naughton and Gerhard Weikum. Kluwer Academic Publishers, Boston, MA. (1998). 110 pages. $97.50, NLG 210.00, GBP 64.50. Computers and Mathematics with Applications, 36(8): 122, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911373.

Anonymous:1998:BRDg

[Ano98-46] Anonymous. Book review:

Dynamics: Numerical explorations: (Second, Revised and Enlarged Edition). By Helena E. Nusse and James A. Yorke. Springer, New York. (1998). 608 pages. $69.95 (diskette included). Computers and Mathematics with Applications, 36(6):138, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890092X.

Anonymous:1998:BREi

[Ano98-47] Anonymous. Book review: Electronic multimedia publishing: Enabling technologies and authoring issues: Edited by Fillia Makedon and Samuel A. Rebelsky. Kluwer Academic Publishers, Boston, MA. (1998). 197 pages. $94.00, NLG 205.00, GBP 62.25. Computers and Mathematics with Applications, 36(8):126, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219891183X.

Anonymous:1998:BREd

[Ano98-48] Anonymous. Book review: Elements of scientific inquiry: By Eric Martin and Daniel Osherson. MIT Press, Cambridge, MA. (1998). 270 pages. $30.00. Computers and Mathematics with Applications, 36(6):143, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198901080.

Anonymous:1998:BREe

[Ano98-49] Anonymous. Book review: Encyclopaedia of mathematics: Supplement volume I: Edited by M. Hazewinkel. Kluwer Academic Publishers, Dordrecht. (1997). 587 pages. $295.00, NLG 495.00, GBP 179.00. Computers and Mathematics with Applications, 36(6):144, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901109.

Anonymous:1998:BREh

[Ano98-50] Anonymous. Book review:

Engineering intelligent hybrid multi-agent systems: By Rajiv Khosla and Tharam Dillon. Kluwer Academic Publishers, Boston, MA. (1997). 410 pages. $126.00, NLG 275.00, GBP 83.20. Computers and Mathematics with Applications, 36(8):124, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911658.

Anonymous:1998:BREc

[Ano98-51] Anonymous. Book review:

Essential Windows NT system administration: By Æleen Frisch. O’Reilly, Sebastopol, CA. (1998). 467 pages. $34.95. Computers and Mathematics with Applications, 35(9):144, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907234.

Anonymous:1998:BREa

[Ano98-52] Anonymous. Book review:

Excel 97 annoyances: By Woody Leonhard, Lee Hudspeth and T. J. Lee. O’Reilly, Sebastopol, CA. (1997). 320 pages. $21.95. Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802350.

Anonymous:1998:BREg

[Ano98-53] Anonymous. Book review: Exercises in basic ring theory: By Grigore Calugareanu and Peter Hamburg. Kluwer Academic Publishers, Dordrecht. (1998). 198 pages. $90.00, NLG 160.00, GBP 55.00. Computers and Mathematics with Applications, 36(8):124, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911634.

Anonymous:1998:BREb

[Ano98-54] Anonymous. Book review: Exploring Java, (second edition): By Patrick Niemeyer and Joshua Peck. O’Reilly, Sebastopol, CA. (1997). 594 pages. $32.95. Computers and Mathematics with Applications, 35(9):144, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907179.

Anonymous:1998:BRFf

[Ano98-55] Anonymous. Book review:

Fast parallel algorithms for graph matching problems: By Marek Karpinski and Wojciech Rytter. Clarendon Press, Oxford. (1998). 212 pages. $65.00. Computers and Mathematics with Applications, 36(8):123, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911518.

Anonymous:1998:BRFe

[Ano98-56] Anonymous. Book review:

Fault-tolerant parallel and distributed systems: Edited by Dimiter R. Avresky and David R. Kaeli. Kluwer Academic Publishers, Boston, MA. (1998). 399 pages. $145.00, NLG 3155.00, GBP 95.70. Computers and Mathematics with Applications, 36 (8):120, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911233.

Anonymous:1998:BRFg

[Ano98-57] Anonymous. Book review:

Finite element analysis for composite structures: By Lazarus Teneketzis Tenek and John Argyris. Kluwer Academic Publishers, Dordrecht. (1998). 339 pages. $169.00, NLG 295.00, GBP 99.00 (diskette included). Computers and Mathematics with Applications, 36(8):124, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911610.

Anonymous:1998:BRFc

[Ano98-58] Anonymous. Book review: Finite-state language processing: Edited by Emmanuel Roche and Yves Schabes. MIT Press, Cambridge, MA. (1997). 464 pages. $45.00. Computers and Mathematics with Applications, 35 (9):142, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907015.

Anonymous:1998:BRFd

[Ano98-59] Anonymous. Book review: Fixed point theory and best approximation: the KKMmap principle: By Sankatha Singh, Bruce Watson and Pramila Sravastava. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 220 pages. NLG 180.00, $103.00, GBP 62.00. Computers and Mathematics with Applications, 36(3): 128, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902140.

Anonymous:1998:BRFa

[Ano98-60] Anonymous. Book review:

Flattening the earth: Two thousand years of map projections: By John P. Snyder. University of Chicago Press, Chicago. (1993). 365 pages. $19.95, £15.95. Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802325.

Anonymous:1998:BRFb

[Ano98-61] Anonymous. Book review:

Fortran 95 handbook: Complete ISO/ANSI reference: By Jeanne C. Adams, Walter S. Brainerd, Jeanne T. Martin, Brain T. Smith and Jerrold L. Wagener. MIT Press, Cambridge, MA. (1997). 711 pages. $50.00. Computers and Mathematics with Applications, 35(9):142, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907052.

Anonymous:1998:FSF

[Ano98-62] Anonymous. Book review:

Fractals and scaling in finance: Discontinuity, concentration, risk: Selecta Volume E. By Benoit B. Mandelbrot. Springer, Berlin. (1997). 551 pages. $39.95, DM 84.00, öS 613.20, sFr 76.50, GBP 32.50. Computers and Mathematics with Applications, 35(5):135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907453.

Anonymous:1998:BRGd

[Ano98-63] Anonymous. Book review: Game theory and the social contract II: Just playing: By Ken Binmore. MIT Press, Cambridge, MA. (1998). 589 pages. $50.00. Computers and Mathematics with Applications, 36(8):120, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911270.

Anonymous:1998:BRGa

[Ano98-64] Anonymous. Book review: Genetic algorithms and evolution strategy in engineering and computer science: Edited by D. Quagliarella, J. Périaux, C. Poloni and G. Winter. John Wiley and Sons, Chichester, England. (1998). 391 pages. GBP 75.00. Computers and Mathematics with Applications, 35(9):143, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890709X.

Anonymous:1998:BRGb

[Ano98-65] Anonymous. Book review:

Geometry, fields and cosmology: Techniques and applications: Edited by B. R. Iyer and C. V. Vishveshwara. Kluwer Academic Publishers, Dordrect, The Netherlands. (1997). 553 pages. NLG 420.00, $248.00, GBP 149.00. Computers and Mathematics with Applications, 36(3): 130, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902279.

Anonymous:1998:BRGc

[Ano98-66]  Anonymous.  Book re-

view: Graph symmetry: Algebraic methods and applications: Edited by Gena and Gert Sabidussi. Kluwer Academic Publishers, Dordrecht. (1997). 418 pages. $175.00, NLG 295.00, GBP 107.00. Computers and Mathematics with Applications, 36(6):139, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198900943.

Anonymous:1998:BRHb

[Ano98-67] Anonymous. Book review:

Handbook of multivalued analysis. Volume I: Theory: By Shouchuan Hu and Nikolas S. Papageorgiou. Kluwer Academic Publishers, Dordrecht. (1997). 964 pages. $395.00, NLG 695.00, GBP 249.00. Computers and Mathematics with Applications, 36(3): 130, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902280.

Anonymous:1998:BRHa

[Ano98-68] Anonymous. Book review: Hard real-time computing systems: Predictable scheduling algorithms and applications: By Giorgio C. Buttazzo. Kluwer Academic Publishers, Boston, MA. (1997). 379 pages. $87.50, NLG 190.00, GBP 57.75. Computers and Mathematics with Applications, 36(3):126, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890205X.

Anonymous:1998:BRHc

[Ano98-69]  Anonymous.  Book re-

view: Hardware/software codesign: Principles and practice: By Jørgen Staunstrup and Wayne Wolf. Kluwer Academic Publishers, Boston, MA. (1997). 395 pages. $98.00, NLG 215.00, GBP 64.70. Computers and Mathematics with Applications, 36(6):143, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901055.

Anonymous:1998:BRHd

[Ano98-70] Anonymous. Book review:

Human values and the design of computer technology: Edited by Batya Friedman. CSLI Publications/ Cambridge University Press, Standford, CA/New York. (1997). 320 pages. $64.95 (hardback); $24.95 (paperback). Computers and Mathematics with Applications, 36 (8):120, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219891121X.

Anonymous:1998:BRIa

[Ano98-71] Anonymous. Book review:

Information retrieval systems: Theory and implementation: By Gerald Kowalski. Kluwer Academic Publishers, Boston. (1997). 282 pages. $105.00, Dfl. 220.00, £71.50. Computers and Mathematics with Applications, 35(5):133, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219780226X.

Anonymous:1998:BRIc

[Ano98-72] Anonymous. Book review: Information systems and data compression: By Jerzy A. Seidler. Kluwer Academic Publishers, Boston. (1997). 472 pages. $160.00, NLG 335.00, GBP 108.80. Computers and Mathematics with Applications, 35(5):135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198907532.

Anonymous:1998:BRIb

[Ano98-73] Anonymous. Book review: Information theory: By Jan C. A. van der Lubbe. (Translated by Hendrik Jan Hoeve and Steve Gee). Cambridge University Press, New York. (English translation 1997). 350 pages. $74.95 (hardback), $29.95 (paperback). Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197802404.

Anonymous:1998:BRIj

[Ano98-74] Anonymous. Book review: Integral, measure, and ordering: By Beloslav Riečan and Tibor Neubrunn. Kluwer Academic Publishers/Ister Science, Dordrecht/Bratislava. (1997). 378 pages. $169.00, NLG 285.00, GBP 99.00.

Computers and Mathematics with Applications, 36(6): 143, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901067.

Anonymous:1998:BRIk

[Ano98-75] Anonymous. Book review:

Integral, probability, and fractal measures: By Gerald A. Edgar. Springer, New York. (1998). 286 pages. $39.95. Computers and Mathematics with Applications, 36(6): 144, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890116X.

Anonymous:1998:BRIh

[Ano98-76] Anonymous. Book review:

Intelligent image databases: Towards advanced image retrieval: By Yihong Gong. Kluwer Academic Publishers, Boston, MA. (1998). 134 pages. $95.00, NLG 205.00, GBP 62.75. Computers and Mathematics with Applications, 36(6):138, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198900918.

Anonymous:1998:BRIe

[Ano98-77] Anonymous. Book review:

Intelligent multimedia information retrieval: Edited by Mark T. Maybury. AAAI Press, Menlo Park, CA. (1997). 472 pages. $37.50. Computers and Mathematics with Applications, 35 (9):139, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906812.

Anonymous:1998:BRId

[Ano98-78] Anonymous. Book review: Interior point algorithms: Theory and analysis: By Yinyu Yee. John Wiley & Sons, New York. (1997). 418 pages. $74.95. Computers and Mathematics with Applications, 35(6):154, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198904953.

Anonymous:1998:BRIg

[Ano98-79] Anonymous. Book review: Internet in a nutshell: a desktop quick reference: By Valerie Quercia. O’Reilly, Sebastopol, CA. (1997). 431 pages. $19.95. Computers and Mathematics with Applications, 36(3):125–126, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902036.

Anonymous:1998:BRIm

[Ano98-80] Anonymous. Book review:

Introduction to scientific programming: Computational problem solving using Mathematica and C: By Joseph L. Zachary. Springer-Verlag, New York. (1998). $49.95 (diskette included). Computers and Mathematics with Applications, 36(8):122, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219891135X.

Anonymous:1998:BRIl

[Ano98-81] Anonymous. Book review:

Introduction to the theory of error-correcting codes (third edition): By Vera Pless. John Wiley & Sons, New York. (1998). 207 pages. $59.95. Computers and Mathematics with Applications, 36(8): 118, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122198911130.

Anonymous:1998:BRIf

[Ano98-82] Anonymous. Book review:

Investing in innovation: Creating a research and innovation policy that works: Edited by Lewis M. Branscomb and James H. Keller. MIT Press, Cambridge, MA. (1998). 516 pages. $35.00. Computers and Mathematics with Applications, 35(9):143, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907118.

Anonymous:1998:BRIn

[Ano98-83] Anonymous. Book review: IUTAM symposium on lubricated transport of viscous materials: Proceedings of the IUTAM symposium held in Tobago, West Indies, 7–10 January 1997: Edited by H. Ramkissoon. Kluwer Academic Publishers, Dordrecht. (1998). 239 pages. $114.00, NLG 200.00, GBP 68.00. Computers and Mathematics with Applications, 36(8): 125, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911695.

Anonymous:1998:BRIi

[Ano98-84] Anonymous. Book review: IUTAM symposium on variable density low-speed turbulent flows: Proceedings of the IUTAM symposium held in Marseille, France, 8–10 July 1996: Edited by Louis Fulachier, John L. Lumley and Fabien Anselmet. Kluwer Academic Publishers, Dordrecht. (1997). 368 pages. $165.00, NLG 280.00, GBP 99.00. Computers and Mathematics with Applications, 36(6):140, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198900967.

Anonymous:1998:BRJe

[Ano98-85] Anonymous. Book review:

JavaScript: the definitive guide (third edition): By D. Flanagan. O’Reilly, Sebastopol, CA. (1998). 776 pages. $39.95. Computers and Mathematics with Applications, 36(8):126, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911774.

Anonymous:1998:BRJd

[Ano98-86] Anonymous. Book review:

Java^{T M} cryptography: By Jonathan Knudsen. O’Reilly, Sebastopol, CA. (1998). 344 pages. $29.95. Computers and Mathematics with Applications, 36(6):141, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901006.

Anonymous:1998:BRJc

[Ano98-87] Anonymous. Book review:

Java^{T M} gems: Jewels from “Java Report”: Edited by Dwight Deugo. Cambridge University Press, Cambridge, UK. (1998). 230 pages. $29.95. Computers and Mathematics with Applications, 36 (3):126–127, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902097.

Anonymous:1998:BRJb

[Ano98-88] Anonymous. Book review: Java^{T M} security: By Scott Oaks. O’Reilly, Sebastopol, CA. (1998). 456 pages. $32.95. Computers and Mathematics with Applications, 36(3): 126, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902061.

Anonymous:1998:BRJa

[Ano98-89] Anonymous. Book review: Jordan, real and lie structures in operator algebras: By Shavkat Ayupov, Abdugafur Rakhimov, and Shukhrat Usmanov. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 225 pages. NLG 195.00, $115.00, GBP 70.00. Computers and Mathematics with Applications, 36(3):126, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902085.

Anonymous:1998:BRLe

[Ano98-90]  Anonymous.  Book re-

view: Language, truth and logic in mathematics: By Jaakko Hintikka. Kluwer Academic Publishers. Dordrecht, The Netherlands. (1998). 247 pages. $99.00, NLG 180.00, GBP 59.00. Computers and Mathematics with Applications, 36(3):124, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219890195X.

Anonymous:1998:BRLg

[Ano98-91] Anonymous. Book review:

Lectures on proof verification and approximation algorithms: Edited by Ernst W. Mayr, Hans Jürgen Prömel and Angelika Steger. Springer, Berlin. (1998). 344 pages. $45.00, DM 68.00, öS 497.00, sFr 62.00, GBP 26.00. Computers and Mathematics with Applications, 36(6):144, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901158.

Anonymous:1998:BRLh

[Ano98-92] Anonymous. Book review:

Lie groups and Lie algebras: Their representations, generalisations and applications: Edited by B. P. Komrakov, I. S. Krasil’shchik, G. L.

Litvinov and A. B. Sossinsky. Kluwer Academic Publishers, Dordrecht. (1998). 442 pages. $194.00, NLG 340.00, GBP 116.00. Computers and Mathematics with Applications, 36(8):119, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911154.

Anonymous:1998:BRLc

[Ano98-93]  Anonymous.  Book re-

view: Logic of computation: Nato ASI Series. Series F: Computer and Systems Sciences, Volume 157. Edited by Helmut Schwichtenberg. Springer-Verlag, Heidelberg, Germany. (1997). 396 pages. $99.00. Computers and Mathematics with Applications, 35(9):140, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906897.

Anonymous:1998:BRLb

[Ano98-94] Anonymous. Book review: Logic programming: Proceedings of the 1997 international symposium: Edited by Jan Maluszyński. MIT Press, Cambridge, MA. (1997). $85.00. Computers and Mathematics with Applications, 35(9):138–139, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906794.

Anonymous:1998:BRLf

[Ano98-95] Anonymous. Book review:

Logic programming: Proceedings of the 1998 joint international conference and symposium on logic programming: Edited by Joxan Jaffar. MIT Press, Cambridge, MA. (1998). 360 pages. $85.00. Computers and Mathematics with Applications, 36(6): 142, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890102X.

Anonymous:1998:BRLa

[Ano98-96] Anonymous. Book review:

Logic programming: Proceedings of the fourteenth international conference on logic programming: Edited by Lee Naish. MIT Press, Cambridge, MA. (1997) $85.00. Computers and Mathematics with Applications, 35 (9):138, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906770.

Anonymous:1998:BRMf

[Ano98-97] Anonymous. Book review:

Managing mailing lists: By Alan Schwartz. O’Reilly, Sebastopol, CA. (1998). 282 pages. $29.95. Computers and Mathematics with Applications, 36(3):125, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198902012.

Anonymous:1998:BRMg

[Ano98-98] Anonymous. Book review: Managing the windows NT registry: By Paul Robichaux. O’Reilly, Sebastopol, CA. (1998). 360 pages. $39.95. Computers and Mathematics with Applications, 36(3): 125, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198902000.

Anonymous:1998:BRMd

[Ano98-99] Anonymous. Book review: Managing usenet: By Henry Spencer and David Lawrence. O’Reilly, Sebastopol, CA. (1998). 492 pages. $32.95. Computers and Mathematics with Applications, 35 (9):144, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198907210.

Anonymous:1998:BRMi

[Ano98-100] Anonymous. Book review: Marcus contextual grammar:

By Gheorghe Paun. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 375 pages. NLG 210.00, $120.00, GBP 65.00. Computers and Mathematics with Applications, 36(3): 128, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902176.

Anonymous:1998:BRMj

[Ano98-101] Anonymous. Book review:

Material identification using mixed numerical experimental methods (proceedings of the EUROMECH colloquium held in Kerkrade, The Netherlands, 7–9 April 1997): Edited by H. Sol and C. W. J. Oomens. Kluwer Academic Publishers, Dordrecht. (1997). 240 pages. $110.00, NLG 195.00, GBP 67.00. Computers and Mathematics with Applications, 36(6):139, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198900931.

Anonymous:1998:BRMb

[Ano98-102] Anonymous. Book review:

Mathematical methods in program development: Edited by Manfred Broy and Birgit Schieder. Springer-Verlag, Berlin. (1997). 528 pages.

$119.00. Computers and Mathematics with Applications, 35(9):140, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219890685X.

Anonymous:1998:BRMk

[Ano98-103] Anonymous. Book review: MCSE the core exams in a nutshell: a desktop quick reference: By Michael Moncur. O’Reilly, Sebastopol, CA. (1998). 401 pages. $19.95. Computers and Mathematics with Applications, 36(6): 140, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198900979.

Anonymous:1998:BRMn

[Ano98-104] Anonymous. Book review: Methods in neuronal modeling: From ions to networks (second edition): Edited by Christof Koch and Idan Segev. MIT Press, Cambridge, MA. (1998). 671 pages. $60.00. Computers and Mathematics with Applications, 36(8):126, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911798.

Anonymous:1998:BRMc

[Ano98-105] Anonymous. Book review:

Modeling dynamic biological systems: By Bruce Hannon and Mathias Ruth. Springer-Verlag, New York. (1997). 399 pages. $59.95, GDM 88.00, CD ROM included. Computers and Mathematics with Applications, 35(9):144, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907295.

Anonymous:1998:BRMe

[Ano98-106] Anonymous. Book review:

Modern compiler implementation in Java: Revised and expanded edition: By Andrew W. Appel. Cambridge University Press, Cambridge, MA. (1998). 548 pages. $54.95. Computers and Mathematics with Applications, 36(3): 124, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122198901936.

Anonymous:1998:BRMm

[Ano98-107] Anonymous. Book review:

Modern differential geometry of curves and surfaces with mathematica (second edition): By Alfred Gray. CRC Press, Boca Raton, FL. (1998). 1053 pages. $84.95. Computers and Mathematics with Applications, 36(8):121, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911336.

Anonymous:1998:BRMh

[Ano98-108] Anonymous. Book review: Multilevel optimization: Algorithms and applications: Edited by Athanasios Migdalas, Panos M. Pardalos, and Peter Värbrand. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1998). 384 pages. NLG 320.00, $182.00, GBP 109.00. Computers and Mathematics with Applications, 36(3): 127, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198902139.

Anonymous:1998:BRMa

[Ano98-109] Anonymous. Book review: Multimedia information retrieval: Content-based information retrieval from large text and audio databases: By Peter Schäuble. Kluwer Academic Publishers, Boston. (1997). 190 pages. $87.75, NLG 175.00, GBP 57.75. Computers and Mathematics with Applications, 35(5): 135, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907519.

Anonymous:1998:BRMl

[Ano98-110] Anonymous. Book review:

Multimedia technologies and applications for the 21-st century: Visions of world experts: Edited by Borko Furht. Kluwer Academic Publishers, Boston, MA. (1998). 323 pages. $115.00, NLG 250.00, GBP 80.00. Computers and Mathematics with Applications, 36(6):141, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198900992.

Anonymous:1998:BRNe

[Ano98-111] Anonymous. Book review: Nearrings, nearfields, and K-loops: Proceedings of the conference on nearrings and nearfields, Hamburg, Germany, July 30– August 6, 1995: Edited by Gerhard Saad and Momme Johs Thomsen. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 444 pages. NLG 340.00, $194.00, GBP 116.00. Computers and Mathematics with Applications, 36(3):125, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902024.

Anonymous:1998:BRNg

[Ano98-112] Anonymous. Book review: Network and netplay: Virtual groups on the Internet: Edited by Fay Sudweeks, Margaret McLaughlin and Sheizaf Rafaeli. AAI Press/MIT Press Menlo Park, CA. (1998) 313 pages. $35.00. Computers and Mathematics with Applications, 36(3): 128, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198902152.

Anonymous:1998:BRNb

[Ano98-113] Anonymous. Book review: Neural organization: Structure, function, and dynamics: By Michael Arbib, Péter Érdi and János Szentágothai. MIT Press, Cambridge, MA. (1998). 407 pages. $60.00. Computers and Mathematics with Applications, 35(5): 133, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197802313.

Anonymous:1998:BRNh

[Ano98-114] Anonymous. Book review: Neurons, networks, and motor behavior: Edited by Paul S. G. Stein, Sten Grillner, Allen I. Selverston and Douglas G. Stuart. MIT Press, Cambridge, MA. (1997). 305 pages. $60.00. Computers and Mathematics with Applications, 36(6):141, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198900980.

Anonymous:1998:BRNa

[Ano98-115] Anonymous. Book review:

Non-commutative valuation rings and semi-hereditary orders: By Hidetoshi Marubayashi, Haruo Miyamoto and Akira Ueda. Kluwer Academic Publishers, Dordrecht. (1997). 191 pages. $99.00, Dfl. 160.00, £59.00. Computers and Mathematics with Applications, 35(5):132, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802258.

Anonymous:1998:BRNi

[Ano98-116] Anonymous. Book review: Nonconvex optimization in mechanics: Algorithms, heuristics and engineering applications by the F.E.M.: By E. S. Mistakidis and G. E. Stavroulakis. Kluwer Academic Publishers, Dordrecht. (1998). 285 pages. $149.00, NLG 260.00, GBP 89.00. Computers and Mathematics with Applications, 36(8):119, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911191.

Anonymous:1998:BRNc

[Ano98-117] Anonymous. Book review: Nonlinear system techniques and applications: By Julius S. Bendat. John Wiley & Sons, New York. (1998). 474 pages. $89.95. Computers and Mathematics with Applications, 35(9):140, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198906915.

Anonymous:1998:BRNf

[Ano98-118] Anonymous. Book review: Numerical computation in science and engineering: By C. Pozrikidis. Oxford University Press, N.Y. (1998). 627 pages. $75.00. Computers and Mathematics with Applications, 36(3):128, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198902164.

Anonymous:1998:BROc

[Ano98-119] Anonymous. Book review: Object modeling and design strategies: Tips and techniques: By Sanjiv Gossain. Cambridge University Press, Cambridge, UK. (1998). 310 pages. $39.95. Computers and Mathematics with Applications, 36(3):130, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902267.

Anonymous:1998:BROb

[Ano98-120] Anonymous. Book review:

Office 97 annoyances: By Woody Leonhard, Lee Hudspeth and T. J. Lee. O’Reilly, Sebastopol, CA. (1997). 379 pages. $21.95. Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802349.

Anonymous:1998:BROa

[Ano98-121] Anonymous. Book review: Operator algebras and applications: Edited by Aristides Katavolos. Kluwer Academic Publishers, Dordrecht. (1997). 467 pages. $198.00, Dfl. 335.00, £120.00. Computers and Mathematics with Applications, 35(5): 133, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802271.

Anonymous:1998:BRPh

[Ano98-122] Anonymous. Book review: PalmPilot: the ultimate guide: By David Pogue.

O’Reilly, Sebastopol, CA. (1998). 489 pages. $29.95 (CD ROM included). Computers and Mathematics with Applications, 36(8):124, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911555.

Anonymous:1998:BRPa

[Ano98-123] Anonymous. Book review: Parallel computing in optimization: Edited by Athanasios Migdalas, Panos M. Pardalos and Sverre Storøy. Kluwer Academic Publishers, Dordrecht. (1997). 585 pages. $260.00, Dfl. 440.00, £159.00. Computers and Mathematics with Applications, 35(5):133, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802301.

Anonymous:1998:BRPb

[Ano98-124] Anonymous. Book review: Parallel optimization: Theory, algorithms, and applications: By Yair Censor and Stavros A. Zenios. Oxford University Press, New York. (1997). $85.00. Computers and Mathematics with Applications, 35(9):143, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907131.

Anonymous:1998:BRPi

[Ano98-125] Anonymous. Book review:

Partial differential equations for computational science: With Maple(R) and vector analysis: By David Betounes. Springer, New York. (1998). $64.95, sFr 116.50, DM 128.00, öS 935.00 (CD-ROM included). Computers and Mathematics with Applications, 36(8):126, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911816.

Anonymous:1998:BRPc

[Ano98-126] Anonymous. Book review:

Partial differential equations through examples and exercises: By Endre Pap, Arpad Takaci, and Djurdjica Takaci. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 404 pages. $182.00, NLG 320.00, GBP 109.00. Computers and Mathematics with Applications, 36(3): 125, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901997.

Anonymous:1998:BRPd

[Ano98-127] Anonymous. Book review: Probabilistic methods for structural design: Edited by C. Guedes Soares. Kluwer Academic Publishers, Dordrecht, The Netherlands. (1997). 402 pages. NLG 325.00, $190.00, GBP 110.00. Computers and Mathematics with Applications, 36(3): 129, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902231.

Anonymous:1998:BRPe

[Ano98-128] Anonymous. Book review: Probabilistic modelling: By Isi Mitrani. Cambridge University Press, New York. (1998). 223 pages. $64.95 (hardback); $24.95 (paperback). Computers and Mathematics with Applications, 36(6):144, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901146.

Anonymous:1998:BRPf

[Ano98-129] Anonymous. Book review: Protecting networks with SA-TAN: By Martin Freiss. O’Reilly, Sebastopol, CA. (1997). 112 pages. $19.95. Computers and Mathematics with Applications, 36(6): 144, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901122.

Anonymous:1998:BRQ

[Ano98-130] Anonymous. Book review:

QuarkXPress in a nutshell: a desktop quick reference: By Donnie O’Quinn. O’Reilly, Sebastopol, CA. (1998). 526 pages. $19.95. Computers and Mathematics with Applications, 36(8):125, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911713.

Anonymous:1998:BRRd

[Ano98-131] Anonymous. Book review:

R&D and productivity: the econometric evidence: By Zvi Griliches. The University of Chicago Press, Chicago, IL. (1998). 382 pages. $56.00, GBP 44.75. Computers and Mathematics with Applications, 36(3):126, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902073.

Anonymous:1998:BRRe

[Ano98-132] Anonymous. Book review:

Real-time database and information systems: Research advances: Edited by Azer Bestavros and Victor Fay-Wolfe. Kluwer Academic Publishers, Boston, MA. (1997). 426 pages. $145.00, NLG 315.00, GBP 95.75. Computers and Mathematics with Applications, 36(6): 142, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901018.

Anonymous:1998:BRRa

[Ano98-133] Anonymous. Book review: Reasoning in boolean networks: Logic synthesis and verification using testing techniques: By Wolfgang Kunz and Dominik Stoffel. Kluwer Academic Publishers, Boston. (1997). 230 pages. $125.00, NLG 265.00, £85.00. Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802374.

Anonymous:1998:BRRg

[Ano98-134] Anonymous. Book review: Reasoning with the infinite: From the closed world to the mathematical universe: By Michel Blay (Translated by M. B. DeBevoise). University of Chicago Press, Chicago. (1998). 216 pages. $30.00; £25.95. Computers and Mathematics with Applications, 36(8):123, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911531.

Anonymous:1998:BRRf

[Ano98-135] Anonymous. Book review:

Recent progress in inequalities: Edited by G. V. Milovanović. Kluwer Academic Publishers, Dordrecht. (1998). 519 pages. $225.00, NLG 395.00, GBP 135.00. Computers and Mathematics with Applications, 36(8): 121, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911294.

Anonymous:1998:BRRc

[Ano98-136] Anonymous. Book review:

Regionalism versus multilateral trade arrangements: Edited by Takatoshi Ito and Anne O. Krueger. University of Chicago Press, Chicago. (1997). 419 pages. $59.00, GBP 47.25. Computers and Mathematics with Applications, 35(9):141, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906939.

Anonymous:1998:BRRb

[Ano98-137] Anonymous. Book review:

Retargetable code generation for digital signal processors: By Rainer Leupers. Kluwer Academic Publishers, Boston. (1997). 210 pages. $115.00, Dfl. 245.00, £78.20. Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802398.

Anonymous:1998:BRSe

[Ano98-138] Anonymous. Book review: Safety-critical real-time systems: Edited by Bernd Krämer and Norbert Völker. Kluwer Academic Publishers, Boston, MA. (1997). 306 pages. $85.00, NLG 185.00, GBP 56.25. Computers and Mathematics with Applications, 36(6):144, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901092.

Anonymous:1998:BRSg

[Ano98-139] Anonymous. Book review: Scalable high performance computing for knowledge discovery and data mining: Edited by Paul Stolorz and Ron Musick. Kluwer Academic Publishers, Boston, MA. (1998). 438 pages. $87.50, NLG 190.00, GBP 57.75. Computers and Mathematics with Applications, 36 (8):125, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911750.

Anonymous:1998:BRSh

[Ano98-140] Anonymous. Book review:

Set theory: Techniques and applications, Curaçao 1995 and Barcelona 1996 conferences: Edited by Carlos Augusto Di Prisco, Jean A. Larson, Joan Bagaria and A. R. D. Mathias. Kluwer Academic Publishers, Dordrecht. (1998). 226 pages. $99.00, NLG 175.00, GBP 59.00. Computers and Mathematics with Applications, 36 (8):126, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911853.

Anonymous:1998:BRSb

[Ano98-141] Anonymous. Book review:

State variables for engineers, (second edition): By Paul M. DeRusso, Rob J. Roy, Charles M. Close and Alan A. Desrochers. John Wiley & Sons, New York, (1998). 575 pages. $64.95. Computers and Mathematics with Applications, 35(9):142, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907076.

Anonymous:1998:BRSc

[Ano98-142] Anonymous. Book review: Statistical tests for mixed linear models: By André I. Khuri, Thomas Mathew and Bimal K. Sinha. John Wiley and Sons, New York. (1998). 352 pages. $69.95. Computers and Mathematics with Applications, 35(9):144, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198907271.

Anonymous:1998:BRSf

[Ano98-143] Anonymous. Book review: Structured COBOL programming: Eighth Edition, Year 2000 Update Edition. By Nancy Stern and Robert A. Stern. John Wiley & Sons, New York. (1999). 722 pages. $75.95 (diskette included). Computers and Mathematics with Applications, 36(8): 121, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911312.

Anonymous:1998:BRSd

[Ano98-144] Anonymous. Book review: System synthesis with VHDL: By Petru Eles, Krzysztof Kuchcinski and Zebo Peng. Kluwer Academic Publishers, Boston, MA. (1998). 370 pages. $115.00, NLG 250.00, GBP 75.90. Computers and Mathematics with Applications, 36(6):143, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901079.

Anonymous:1998:BRTb

[Ano98-145] Anonymous. Book review: Technology and American history: Edited by Stephen H. Cutcliffe and Terry S. Reynolds. University of Chicago Press, Chicago, IL. (1997). 444 pages. $37.95 (hardback), $18.95 (paperback). Computers and Mathematics with Applications, 35 (9):140, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906873.

Anonymous:1998:BRTd

[Ano98-146] Anonymous. Book review:

Technology and the west: Edited by Terry S. Reynolds and Stephen H. Cutcliffe. University of Chicago Press, Chicago, IL. (1997). 461 pages. $37.95 (hardback); $18.95 (paperback). Computers and Mathematics with Applications, 35(9):142, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219890699X.

Anonymous:1998:BRTg

[Ano98-147] Anonymous. Book review: Term rewriting and all that: By Franz Baader and Tobias Nipkow. Cambridge University Press, New York. 301 pages. $49.95. Computers and Mathematics with Applications, 36(6):144, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198901134.

Anonymous:1998:BREf

[Ano98-148] Anonymous. Book review: The economics of propertycasualty insurance: Edited by David F. Bradford. University of Chicago Press, Chicago. (1998). 203 pages. $37.50, £29.95. Computers and Mathematics with Applications, 36(8):122, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198911397.

Anonymous:1998:BRNd

[Ano98-149] Anonymous. Book review: The netscape programmer’s guide: Using OLE to build componentware apps: By Richard B. Lam. Cambridge University Press, New York. (1998). 394 pages. $39.95 (Includes CD ROM). Computers and Mathematics with Applications, 36(3):124, August 1998. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122198901961.

Anonymous:1998:BRPg

[Ano98-150] Anonymous. Book review:

The patterns handbook: Techniques, strategies, and applications: Edited by Linda Rising. Cambridge University Press, New York. (1998). 549 pages. $44.95. Computers and Mathematics with Applications, 36(8):118, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122198911117.

Anonymous:1998:BRSa

[Ano98-151] Anonymous. Book review: The syntax of nonfinite complementation: an economy approach: By Zeljko Bosković. MIT Press, Cambridge, MA. (1997). 247 pages. $25.00. Computers and Mathematics with Applications, 35(5):132, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802246.

Anonymous:1998:BRTf

[Ano98-152] Anonymous. Book review:

The theory of learning in games: By Drew Fudenberg and David K. Levine. MIT Press, Cambridge, MA.

(1998). 276 pages. $29.95. Computers and Mathematics with Applications, 36(6): 144, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901110.

Anonymous:1998:BRTa

[Ano98-153] Anonymous. Book review: The theory of partial algebraic operations: By E. S. Ljapin and A. E. Evseev. Kluwer Academic Publishers, Dordrecht. (1997). 235 pages. $133.00, Dfl. 225.00, £81.00. Computers and Mathematics with Applications, 35(5):134, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802337.

Anonymous:1998:BRTh

[Ano98-154] Anonymous. Book review: Theory of linear and integer programming: By Alexander Schrijver. John Wiley & Sons, Chichester, U.K. (1986). 471 pages. £29.95. Computers and Mathematics with Applications, 36(8):122, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911415.

Anonymous:1998:BRTe

[Ano98-155] Anonymous. Book review:

Thermal management of electronic systems II: Proceedings of EUROTHERM seminar 45, 20-22 September 1995, Leuven, Belgium: Edited by E. Beyne, C. J. M. Lasance and J. Berghmans. Kluwer Academic Publishers, Dordrecht. (1997). 359 pages. $169.00, NLG 285.00, GBP 103.00. Computers and Mathematics with Applications, 36(6):138, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001679.

Anonymous:1998:BRTi

[Ano98-156] Anonymous. Book review: Transaction management support for cooperative applications: Edited by Rolf A. de By, Wolfgang Klas and Jari Veijalainen. Kluwer Academic Publishers, Boston, MA. (1998). 222 pages. $125.00, NLG 270.00, GBP 82.50. Computers and Mathematics with Applications, 36(8):125, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911737.

Anonymous:1998:BRTc

[Ano98-157] Anonymous. Book review:

Trends in structural mechanics: Edited by J. Roorda and N. K. Srivastava. Kluwer Academic Publishers, Dordrecht, The Netherlands. 385 pages. (1997). $174.00. Computers and Mathematics with Applications, 35(9):141, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198906952.

Anonymous:1998:BRUb

[Ano98-158] Anonymous. Book review: Upgrading relational databases with objects: By Robert Vermeulen. SIGS Publications, Inc., NY. (1996). 212 pages. $39.95. Computers and Mathematics with Applications, 36(3):130, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198902255.

Anonymous:1998:BRUa

[Ano98-159] Anonymous. Book review: Urban travel demand modeling: From individual choices to general equilibrium: By N. Oppenheim. John Wiley & Sons, Inc., New York. (1995). 480 pages. $69.95. Computers and Mathematics with Applications, 35(5):135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907490.

Anonymous:1998:BRVa

[Ano98-160] Anonymous. Book review:

Verilog(R) quickstart: By James M. Lee. Kluwer Academic Publishers, Boston. (1997). 301 pages. $89.50, Dfl. 180.00, £59.10 (diskette included). Computers and Mathematics with Applications, 35(5):133, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197802295.

Anonymous:1998:BRVb

[Ano98-161] Anonymous. Book review:

Video on demand systems: Technology, interoperability and trials: Edited by Shih-Fu Chang, Dimitris Anastassiou, Alexandros Eleftheriadis and John V. Pavlik. Kluwer Academic Publishers, Boston, MA. (1997). 222 pages. $75.00, NLG 160.00, GBP 51.00. Computers and Mathematics with Applications, 36(6):143, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198901043.

Anonymous:1998:BRWb

[Ano98-162] Anonymous. Book review:

Web Board 3.0: Documentation by Susan B. Peck and Jay York. O’Reilly Software, Sebastopol, CA. (1998). $699.00 (includes CD ROM and 350-page book). Computers and Mathematics with Applications, 36(3):124, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001357.

Anonymous:1998:BRWa

[Ano98-163] Anonymous. Book review: Windows NT file system internals: a developer’s guide: By Rajeev Nagar. O’Reilly, Sebastopol, CA. (1997). 774 pages. $49.95. Computers and Mathematics with Applications, 35(5):135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198907477.

Anonymous:1998:BRWc

[Ano98-164] Anonymous. Book review: Workflow and process automation: Concepts and technology: By Andrzej Cichocki, Abdelsalam (Sumi) Helal, Marek Rusinkiewicz and Darrell Woelk. Kluwer Academic Publishers, Boston, MA. (1998). 121 pages. $79.95, NLG 175.00. GBP 53.00. Computers and Mathematics with Applications, 36(8): 120, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198911257.

Anonymous:1998:CSN

[Ano98-165] Anonymous. Corrigendum to “The solution of nonlinear Schrödinger equations using orthogonal spline collocation”. Computers and Mathematics with Applications, 35(6):151, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000261.

See [Rob97].

Anonymous:1998:EBa

[Ano98-166] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 35(1– 2):iii–v, January 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198905570.

Anonymous:1998:EBb

[Ano98-167] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 36 (1):v–vii, July 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198910318.

Anonymous:1998:LCa

[Ano98-168] Anonymous. List of contents. Computers and Mathematics with Applications, 35 (12):iii–xi, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800011.

Anonymous:1998:LCb

[Ano98-169] Anonymous. List of contents. Computers and Mathematics with Applications, 36(10–12): iii–x, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800448.

Anonymous:1999:AI

[Ano99a]  Anonymous.  Author in-

dex. Computers and Mathematics with Applications, 37 (11–12):XII–XIV, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001613.

Anonymous:1999:BRj

[Ano99b] Anonymous. Book report. Computers and Mathematics with Applications, 37 (2):121–128, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198002570.

Anonymous:1999:BRk

[Ano99c] Anonymous. Book reports. Computers and Mathematics with Applications, 37(7):

139–146, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199905011.

Anonymous:1999:BRCm

[Ano99d] Anonymous. Book review:

A concise course in algebraic topology: By J. P. May. University of Chicago Press, Chicago. (1999). $40.00, GBP 28.00 (cloth); $18.00, GBP 13.00 (paper). Computers and Mathematics with Applications, 38(11–12):283, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199911999.

Anonymous:1999:BRCj

[Ano99e] Anonymous. Book review: A course in mathematical modeling: By Douglas D. Mooney and Randall J. Swift. Mathematical Association of America, Washington, DC. (1999). 431 pages. $41.95. Computers and Mathematics with Applications, 38(1):128, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900512.

Anonymous:1999:BRHb

[Ano99f] Anonymous. Book review:

A history of modern computing: By Paul E. Ceruzzi. MIT Press, Cambridge, MA. (1998). 398 pages. $35.00. Computers and Mathematics with Applications, 37(3): 132, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903838.

Anonymous:1999:BRHg

[Ano99g] Anonymous. Book review: A history of modern computing: By Paul E. Ceruzzi. MIT Press, Cambridge, MA. (1998). 398 pages. $35.00. Computers and Mathematics with Applications, 37(6): 132, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990247X.

Anonymous:1999:BRHe

[Ano99h] Anonymous. Book review: A history of the modern fact: Problems of knowledge in the sciences of wealth and society: By Mary Poovey. University of Chicago Press, Chicago. (1998). 419 pages. $49.00, £39.25 (cloth); $17.00, £13.50 (paper). Computers and Mathematics with Applications, 37(3): 135, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990425X.

Anonymous:1999:BRHj

[Ano99i] Anonymous. Book review:

A history of the modern fact: Problems of knowledge in the sciences of wealth and society: By Mary Poovey. University of Chicago Press, Chicago. (1998). 419 pages. $49.00, £39.25 (cloth); $17.00, £13.50 (paper). Computers and Mathematics with Applications, 37(6): 135, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902894.

Anonymous:1999:BRMl

[Ano99j] Anonymous. Book review:

A mathematical mystery tour: Discovering the truth and beauty of the cosmos: By A. K. Dewdney. John Wiley & Sons, Inc., New York. (1999). 218 pages. $22.95. Computers and Mathematics with Applications, 37(10):173, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903395.

Anonymous:1999:BRPi

[Ano99k] Anonymous. Book review:

A programmer’s guide to ZPL: By Lawrence Snyder. MIT Press, Cambridge, MA.

(1999). 154 pages. $30.00. Computers and Mathematics with Applications, 38 (1):122, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990007X.

Anonymous:1999:BRRe

[Ano99l] Anonymous. Book review: A reformulation-linearization technique for solving discrete and continuous nonconvex problems: By Hanif D. Sherali and Warren P. Adams. Kluwer Academic Publishers, Dordrecht. (1999). 514 pages. $252.00, NLG 420.00, GBP 147.00. Computers and Mathematics with Applications, 38 (11–12):288, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912336.

Anonymous:1999:BRAt

[Ano99m] Anonymous. Book review: Access database design and programming, second edition: By Steven Roman. O’Reilly, Sebastopol, CA. (1999). 409 pages. $27.95. Computers and Mathematics with Applications, 38 (11–12):288, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912257.

Anonymous:1999:BRAk

[Ano99n] Anonymous. Book review:

Adaptive hypertext and hypermedia: Edited by Peter Brusilovsky, Alfred Kobsa and Julita Vassileva. Kluwer Academic Publishers, Dordrecht. (1998). 252 pages. $98.00, NLG 175.00, GBP 59.00. Computers and Mathematics with Applications, 37 (10):169, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903231.

Anonymous:1999:BRAh

[Ano99o] Anonymous. Book review:

Advanced techniques for embedded systems desion and test: Edited by Juan Carlos López, Román Hermida and Walter Geisselhardt. Kluwer Academic Publishers, Boston, MA. (1998). 290 pages. $125.00, NLG 285.00, GBP 85.00. Computers and Mathematics with Applications, 37(8):141, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901876.

Anonymous:1999:BRAu

[Ano99p] Anonymous. Book review:

Advances in genetic programming, volume III: Edited by Lee Spector, William B. Langdon, Una-May O’Reilly and Peter J. Angeline. MIT Press, Cambridge, MA. (1999). 476 pages. $55.00. Computers and Mathematics with Applications, 38 (11–12):291, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912671.

Anonymous:1999:BRAs

[Ano99q] Anonymous. Book review: Advances in neuro information processing systems 11: Proceedings of the 1998 conference: Edited by Michael S. Kearns, Sara A. Solla and David A. Cohn. MIT Press, Cambridge, MA. (1999). 1090 pages. $65.00. Computers and Mathematics with Applications, 38(11–12):285– 287, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912154.

Anonymous:1999:BRAi

[Ano99r] Anonymous. Book review: Advances in nonlinear programming: Proceedings of the 96 international conference on nonlinear programming: Edited by Ya-xiang Yuan. Kluwer Academic Publishers, Dordrecht. (1998). 351 pages. $149.00, NLG 280.00, GBP 96.00. Computers and Mathematics with Applications, 37(8):142–143, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901918.

Anonymous:1999:BRAj

[Ano99s] Anonymous. Book review:

Advances in robot kinematics: Analysis and control: Edited by Jadran Lenarčic and Manfred L. Husty. Kluwer Academic Publishers, Dordrecht. (1998). 581 pages. $254.00, NLG 470.00, GBP 160.00. Computers and Mathematics with Applications, 37(8): 143–144, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901955.

Anonymous:1999:BRAl

[Ano99t] Anonymous. Book review:

Advances in turbulence VII: Proceedings of the Seventh European Turbulence Conference, held in Saint-Jean Cap Ferrat, France, 30 June– 3 July 1998: Edited by Uriel Frisch. Kluwer Academic Publishers, Dordrecht. (1998). 613 pages. $268.00, NLG 495.00, GBP 169.00. Computers and Mathematics with Applications, 37(10): 170–172, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903279.

Anonymous:1999:BRAp

[Ano99u] Anonymous. Book review: Algebraic groups and their representations: Edited by R. W. Carter and J. Saxl. Kluwer Academic Publishers, Dordrecht. (1998). 374 pages. $173.00, NLG 320.00, GBP 109.00. Computers and Mathematics with Applications, 38(5–6):253, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901335.

Anonymous:1999:BRAc

[Ano99v] Anonymous. Book review: Algorithms and data structures in VLSI design: OBDD-Foundations and applications: By Christoph Meinel and Thorsten Theobald. Springer-Verlag, Berlin. (1998). 267 pages. $49.95, DM 79.00, öS 577.00, sFr 72.00, GBP 30.50. Computers and Mathematics with Applications, 37(3):134, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904078.

Anonymous:1999:BRAf

[Ano99w] Anonymous. Book review:

Algorithms and data structures in VLSI design: OBDD-Foundations and applications: By Christoph Meinel and Thorsten Theobald. Springer-Verlag, Berlin. (1998). 267 pages. $49.95, DM 79.00, öS 577.00, sFr 72.00, GBP 30.50. Computers and Mathematics with Applications, 37 (6):134, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902717.

Anonymous:1999:BRAn

[Ano99x] Anonymous. Book review:

Algorithms for parallel processing: Edited by Michael T. Heath, Abhiram Ranade and Robert S. Schreiber. Springer-Verlag, New York. (1999). 366 pages. $69.95, DM 129.00, öS 942.00, sFr 117.50, GBP 49.50. Computers and Mathematics with Applications, 38(1):126, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900275.

Anonymous:1999:BRAm

[Ano99y] Anonymous. Book review:

Apache: the definitive guide (second edition): By Ben Laurie and Peter Laurie. O’Reilly, Sebastopol, CA.

(1999). 369 pages. $34.95 (CD-ROM included). Computers and Mathematics with Applications, 38(1):122, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900093.

Anonymous:1999:BRAb

[Ano99z] Anonymous. Book review: Applications of point set theory in real analysis: By A. B. Kharazishvili. Kluwer Academic Publishers, Dordrecht. (1998). 236 pages. $105.00, NLG 195.00, GBP 66.00. Computers and Mathematics with Applications, 37(3):133, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903978.

Anonymous:1999:BRAe

[Ano99-27] Anonymous. Book review: Applications of point set theory in real analysis: By A. B. Kharazishvili. Kluwer Academic Publishers, Dordrecht. (1998). 236 pages. $105.00, NLG 195.00, GBP 66.00. Computers and Mathematics with Applications, 37 (6):133, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902614.

Anonymous:1999:BRAq

[Ano99-28] Anonymous. Book review:

Applied multivariate statistics with SAS software, second edition: By Ravindra Khattree and Dayanand Naik. SAS Institute, Inc./John Wiley & Sons, Cary, NC/ New York. (1999). 338 pages. $59.95. Computers and Mathematics with Applications, 38 (5–6):255, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901578.

Anonymous:1999:BRAa

[Ano99-29] Anonymous. Book review:

Artificial life VI: Proceedings of the sixth international conference on artificial life: Edited by Christoph Adami, Richard K. Belew, Hiroaki Kitano and Charles Taylor. MIT Press, Cambridge, MA. (1998). 498 pages. $65.00. Computers and Mathematics with Applications, 37 (3):131–132, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903711.

Anonymous:1999:BRAd

[Ano99-30] Anonymous. Book review:

Artificial life VI: Proceedings of the sixth international conference on artificial life: Edited by Christoph Adami, Richard K. Belew, Hiroaki Kitano and Charles Taylor. MIT Press, Cambridge, MA. (1998). 498 pages. $65.00. Computers and Mathematics with Applications, 37(6):131–132, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902353.

Anonymous:1999:BRAo

[Ano99-31] Anonymous. Book review: ASP in a nutshell: a desktop quick reference: By A. Keyton Weissinger. O’Reilly, Sebastopol, CA. (1999). 407 pages. $24.95. Computers and Mathematics with Applications, 38(1):128, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900536.

Anonymous:1999:BRAr

[Ano99-32] Anonymous. Book review: Associative learning for a robot intelligence: By John H. Andreae. Imperial College Press, London. (1998). 348 pages. $48.00. Computers and Mathematics with Applications, 38(5–6):256, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901633.

Anonymous:1999:BRBc

[Ano99-33] Anonymous. Book review:

Basic topological structures of ordinary differential equations: By V. V. Filippov. Kluwer Academic Publishers, Dordrecht. (1998). 516 pages. $225.00, NLG 395.00, GBP 135.00. Computers and Mathematics with Applications, 37(8):144, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902018.

Anonymous:1999:BRBd

[Ano99-34] Anonymous. Book review:

Beginning partial differential equations: By Peter V. O’Neil. John Wiley & Sons, Inc., New York. (1999). 500 pages. $51.95. Computers and Mathematics with Applications, 37(10):173, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903450.

Anonymous:1999:BRBa

[Ano99-35]  Anonymous.  Book re-

view: Be^{T M}: Advanced topics: By the Be Development Team. O’Reilly, Sebastopol, CA. (1998). 365 pages. $39.95 (CD-ROM included). Computers and Mathematics with Applications, 37(3):133, February 1999. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903875.

Anonymous:1999:BRBb

[Ano99-36]  Anonymous.  Book re-

view: Be&TM: Advanced topics: By the Be Development Team. O’Reilly, Sebastopol, CA. (1998). 365 pages. $39.95 (CD-ROM included). Computers and Mathematics with Applications, 37(6):133, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902511.

Anonymous:1999:BRBe

[Ano99-37] Anonymous. Book review: Business intelligence: the IBM solution: By Mark Whitehorn and Mary Whitehorn. Springer, London. (1999). 289 pages. $79.95, DM 115.00, öS 840.00, sF 105.00, GBP 37.50 (CD-ROM included). Computers and Mathematics with Applications, 38(11–12):284, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199912099.

Anonymous:1999:BRCq

[Ano99-38] Anonymous. Book review: Calculating risks? The spatial and political dimensions of hazardous waste policy: By James Hamilton and W. Kip Viscusi. MIT Press, Cambridge, MA. (1999). 326 pages. $37.50. Computers and Mathematics with Applications, 38 (11–12):292, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912774.

Anonymous:1999:BRCk

[Ano99-39]  Anonymous.  Book re-

view: Cartesian questions: Method and metaphysics: By Jean-Luc Marion. University of Chicago Press, Chicago. (1999). 215 pages. $25.00, £19.95. Computers and Mathematics with Applications, 38(5–6):254, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901475.

Anonymous:1999:BRCl

[Ano99-40]  Anonymous.  Book re-

view: Computation, causation, and discovery: Edited by Clark Glymour and Gregory F. Cooper. AAAI Press/ MIT Press, Menlo, CA/ Cambridge, MA. (1999). 552 pages. $45.00. Computers and Mathematics with Applications, 38(11–12):282, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199911872.

Anonymous:1999:BRCa

[Ano99-41]  Anonymous.  Book re-

view: Computational inelasticity: By J. C. Simo and T. J. R. Hughes. Springer-Verlag, New York. (1998). 392 pages. $59.95, DM 124.00, öS 906.00, sFr 113.00, GBP 47.50. Computers and Mathematics with Applications, 37(3):134, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904133.

Anonymous:1999:BRCc

[Ano99-42]  Anonymous.  Book re-

view: Computational inelasticity: By J. C. Simo and T. J. R. Hughes. Springer-Verlag, New York. (1998). 392 pages. $59.95, DM 124.00, öS 906.00, sFr 113.00, GBP 47.50. Computers and Mathematics with Applications, 37 (6):134, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902778.

Anonymous:1999:BRCn

[Ano99-43] Anonymous. Book review: Computational partial differential equations: Numerical methods and diffpack programming: By Hans Petter Langtangen. Springer-Verlag, Berlin. (1999). 682 pages. $69.95, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50. Computers and Mathematics with Applications, 38 (11–12):288, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912312.

Anonymous:1999:BRCg

[Ano99-44] Anonymous. Book review:

Computational visualization: Graphics, abstraction, and interactivity: By Thomas Strothotte. Springer-Verlag, Berlin. (1998). 459 pages. $64.95, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50. Computers and Mathematics with Applications, 37(8):146, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902213.

Anonymous:1999:BRCp

[Ano99-45] Anonymous. Book review:

Computer algebra systems: a practical guide: Edited by Michael J. Wester. John Wiley & Sons, Ltd, Chicester. (1999). 436 pages. £65.00. Computers and Mathematics with Applications, 38 (11–12):290, December 1999.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912579.

Anonymous:1999:BRCo

[Ano99-46]  Anonymous.  Book re-

view: Computer graphics and geometric modeling: By David Salomon. Springer-Verlag, New York. (1999). 851 pages. $64.95, DM 129.00, öS 942.00, sFr 117.50, GBP 49.50. Computers and Mathematics with Applications, 38 (11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912452.

Anonymous:1999:BRCb

[Ano99-47] Anonymous. Book review: Computing for scientists: Principles of programming with fortran 90 and C++: By R. J. Barlow and A. R. Barnett. John Wiley & Sons, Inc., Chichester. (1998). 276 pages. $22.50. Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904376.

Anonymous:1999:BRCd

[Ano99-48] Anonymous. Book review: Computing for scientists:

Principles of programming with Fortran 90 and C++: By R. J. Barlow and A. R. Barnett. John Wiley & Sons, Inc., Chichester. (1998). 276 pages. $22.50. Computers and Mathematics with Applications, 37(6):136, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903012.

Anonymous:1999:BRCi

[Ano99-49]  Anonymous.  Book re-

view: Continuous selections of multivalued mappings: By Dušan Repovš and Pavel Vladimirovič Semenov. Kluwer Academic Publishers, Dordrecht. (1998). 356 pages. $159.00, NLG 295.00, GBP 99.00. Computers and Mathematics with Applications, 38(1):126, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900299.

Anonymous:1999:BRCe

[Ano99-50] Anonymous. Book review:

Cross-Language information retrieval: Edited by Gregory Grefenstette. Kluwer Academic Publishers, Boston, MA. (1998). 182 pages. $115.00, NLG 260.00, GBP 78.25. Computers and Mathematics with Applications, 37 (8):145, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902092.

Anonymous:1999:BRCh

[Ano99-51] Anonymous. Book review: Cultural boundaries of science: Credibility on the line: By Thomas F. Gieryn. University of Chicago Press, Chicago. (1999). 398 pages. $21.00, £16.75. Computers and Mathematics with Applications, 37(10):169, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903255.

Anonymous:1999:BRDj

[Ano99-52]  Anonymous.  Book re-

view: Data mining: Methods for knowledge discovery: By Krzysztof J. Cios, Witold Pedrycz and Roman Swiniarski. Kluwer Academic Publishers, Boston, MA. (1998). 495 pages. $149.00, NLG 340.00, GBP 101.50. Computers and Mathematics with Applications, 37 (10):174, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903577.

Anonymous:1999:BRDc

[Ano99-53] Anonymous. Book review:

Data, statistics, and decision models with Excel: By Donald L. Harnett and James F. Horrell. John Wiley & Sons, Inc., New York. (1998). 605 pages. $93.95 (2 diskettes included). Computers and Mathematics with Applications, 37(3):135, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904236.

Anonymous:1999:BRDg

[Ano99-54] Anonymous. Book review:

Data, statistics, and decision models with Excel: By Donald L. Harnett and James F. Horrell. John Wiley & Sons, Inc., New York. (1998). 605 pages. $93.95 (2 diskettes included). Computers and Mathematics with Applications, 37(6):135, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902870.

Anonymous:1999:BRDd

[Ano99-55]  Anonymous.  Book re-

view: Data structure programming: With the standard template library in C++: By Joseph Bergin. Springer-Verlag, New York. (1998). 336 pages. $49.95, DM 89.00, öS 650.00, sFr 81.00, GBP 34.00. Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904339.

Anonymous:1999:BRDh

[Ano99-56]  Anonymous.  Book re-

view: Data structure programming: With the standard template library in C++: By Joseph Bergin. Springer-Verlag, New York. (1998). 336 pages. $49.95, DM 89.00, öS 650.00, sFr 81.00, GBP 34.00. Computers and Mathematics with Applications, 37 (6):136, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902973.

Anonymous:1999:BRDm

[Ano99-57] Anonymous. Book review: Database recovery: By Vijay Kumar and Sang H. Son. Kluwer Academic Publishers, Boston, MA. (1998). 94 pages. $110.00, NLG 250.00, GBP 75.00. Computers and Mathematics with Applications, 38(1):127, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199900470.

Anonymous:1999:BRDq

[Ano99-58] Anonymous. Book review:

Deadline scheduling for realtime systems: EDF and related algorithms: By John A. Stankovic, Marco Spuri, Krithi Ramamritham and Giorgio C. Buttazzo. Kluwer Academic Publishers, Boston, MA. (1998). 273 pages. $79.00, NLG 180.00, GBP 53.75. Computers and Mathematics with Applications, 38 (11–12):288, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219991235X.

Anonymous:1999:BRDp

[Ano99-59] Anonymous. Book review:

Delta functions: an introduction to generalised functions: By R. F. Hoskins. Horwood Publishing, Chichester. (1999). 262 pages. £25.00. Computers and Mathematics with Applications, 38 (11–12):288, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912294.

Anonymous:1999:BRDl

[Ano99-60] Anonymous. Book review:

Descriptive complexity: By Neil Immerman. Springer-Verlag, New York. (1999). 268 pages. $54.95, DM 99.00, öS 723.00, sFr 90.50, GBP 38.00. Computers and Mathematics with Applications, 38 (1):126, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199900354.

Anonymous:1999:BRDn

[Ano99-61] Anonymous. Book review: Developing ASP components: By Shelley Powers. O’Reilly, Sebastopol, CA. (1999). 490 pages. $29.95. Computers and Mathematics with Applications, 38(5–6):254, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901372.

Anonymous:1999:BRDo

[Ano99-62] Anonymous. Book review: Director in a nutshell: a desktop quick reference: By Bruce A. Epstein. O’Reilly, Sebastopol, CA. (1999). 628 pages. $24.95. Computers and Mathematics with Applications, 38(5–6):254, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901396.

Anonymous:1999:BRDa

[Ano99-63] Anonymous. Book review: Discrete and fractional programming techniques for location models: By Ana Isabel Barros. Kluwer Academic Publishers, Dordrecht.

(1998). 178 pages. $97.00, NLG 180.00, GBP 62.00. Computers and Mathematics with Applications, 37(3): 133, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903954.

Anonymous:1999:BRDe

[Ano99-64] Anonymous. Book review:

Discrete and fractional programming techniques for location models: By Ana Isabel Barros. Kluwer Academic Publishers, Dordrecht. (1998). 178 pages. $97.00, NLG 180.00, GBP 62.00. Computers and Mathematics with Applications, 37(6): 133, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902596.

Anonymous:1999:BRDk

[Ano99-65] Anonymous. Book review:

Discrete mathematics using Latin squares: By Charles F. Laywine and Gary L. Mullen. John Wiley & Sons, Inc., New York. (1998). 305 pages. $51.59. Computers and Mathematics with Applications, 37(10):174, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903516.

Anonymous:1999:BRDb

[Ano99-66]  Anonymous.  Book re-

view: DNA computing: New computing paradigms: By G. Paun, G. Rozenberg and A. Salomaa. Springer-Verlag, Berlin. (1998). 402 pages. $59.95, DM 89.00, öS 650.00, sFr 81.00, GBP 34.00. Computers and Mathematics with Applications, 37(3):134, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219990411X.

Anonymous:1999:BRDf

[Ano99-67]  Anonymous.  Book re-

view: DNA computing: New computing paradigms: By G. Paun, G. Rozenberg and A. Salomaa. Springer-Verlag, Berlin. (1998). 402 pages. $59.95, DM 89.00, öS 650.00, sFr 81.00, GBP 34.00. Computers and Mathematics with Applications, 37 (6):134, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902754.

Anonymous:1999:BRDr

[Ano99-68] Anonymous. Book review: Domains and lambda-calculi: By Roberto M. Amadio and Pierre-Louis Curien. Cambridge University Press, Cambridge. (1998). 484 pages. $74.95. Computers and Mathematics with Applications, 38 (11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912415.

Anonymous:1999:BRDi

[Ano99-69] Anonymous. Book review:

Dynamic modules: User’s manual and programming guide for MuPAD 1.4: By Andreas Sorgatz. Springer-Verlag, Berlin. (1999). 234 pages. $59.95, DM 89.00, öS 650.00, sFr 81.00, GBP 34.00 (CD-ROM included). Computers and Mathematics with Applications, 37(8):145, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902031.

Anonymous:1999:BREj

[Ano99-70] Anonymous. Book review:

Efficient solvers for incompressible flow problems: an algorithmic and computational approach: By Stefan Turek. Springer-Verlag, Berlin. (1999). 352 pages. $99.00, DM 149.00, öS 1,088.00, sFr 136.00, GBP 57.50 (CD-ROM included). Computers and Mathematics with Applications, 38 (11–12):292, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912853.

Anonymous:1999:BREf

[Ano99-71] Anonymous. Book review: Elementary lie group analysis and ordinary differential equations: By Nail H. Ibragimov. John Wiley & Sons, Chichester, England. (1999). 347 pages. £55.00. Computers and Mathematics with Applications, 38 (5–6):252, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901293.

Anonymous:1999:BREi

[Ano99-72] Anonymous. Book review: Enterprise Javabeans^{T M}: By Richard Monson-Haefel. O’Reilly, Sebastopol, CA. (1999). 317 pages. $32.95. Computers and Mathematics with Applications, 38(11–12):284, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199912130.

Anonymous:1999:BREe

[Ano99-73] Anonymous. Book review: Euler: the master of us all: By William Dunham.

Mathematical Association of America. (1999). 185 pages. $29.95. Computers and Mathematics with Applications, 38(1):126, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900330.

Anonymous:1999:BREg

[Ano99-74] Anonymous. Book review:

Evolutionary algorithms in engineering and computer science: Recent advances in genetic algorithms, evolution strategies, evolutionary programming, genetic programming and industrial applications: Edited by K. Miettinen, P. Neittaanmäki, M. M. Mäkelä and J. Périaux. John Wiley & Sons, Ltd., Chicester. (1999). £60.00. Computers and Mathematics with Applications, 38 (11–12):282, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199911896.

Anonymous:1999:BREa

[Ano99-75] Anonymous. Book review:

Evolutionary games and equilibrium selection: By Larry Samuelson. MIT Press, Cambridge, MA. (1997). 309 pages. $20.00. Computers and Mathematics with Applications, 37(3):132, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903814.

Anonymous:1999:BREc

[Ano99-76] Anonymous. Book review: Evolutionary games and equilibrium selection: By Larry Samuelson. MIT Press, Cambridge, MA. (1997). 309 pages. $20.00. Computers and Mathematics with Applications, 37(6):132, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902456.

Anonymous:1999:BREh

[Ano99-77] Anonymous. Book review: Exploring abstract algebra with mathematica(R): By Allen C. Hibbard and Kenneth M. Levasseur. Springer-Verlag, New York. (1999). 467 pages. $45.00, DM 84.00, öS 614.00, sFr 76.50, GBP 32.50. Computers and Mathematics with Applications, 38 (11–12):283, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199911938.

Anonymous:1999:BREb

[Ano99-78] Anonymous. Book review: Exploring mathematics with scientific notebook: By Wei-Chi Yang and Jonathan Lewin. Springer-Verlag, Singapore. (1998). 209 pages. $24.50, DM 49.00, öS 358.00, sFr 45.50, GBP 19.00 (includes CD-ROM). Computers and Mathematics with Applications, 37(3):134, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904054.

Anonymous:1999:BREd

[Ano99-79] Anonymous. Book review:

Exploring mathematics with scientific notebook: By Wei-Chi Yang and Jonathan Lewin. Springer-Verlag, Singapore. (1998). 209 pages. $24.50, DM 49.00, öS 358.00, sFr 45.50, GBP 19.00 (includes CD-ROM). Computers and Mathematics with Applications, 37(6):134, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902699.

Anonymous:1999:BRFh

[Ano99-80] Anonymous. Book review:

Feature extraction, construction and selection: a data mining perspective: Edited by Huan Liu and Hiroshi Motoda. Kluwer Academic Publishers, Boston, MA. (1998). 410 pages. $140.00, NLG 320.00, GBP 95.25. Computers and Mathematics with Applications, 38(1):125, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199900214.

Anonymous:1999:BRFf

[Ano99-81] Anonymous. Book review: Feynman and computation: Exploring the limits of computers: Edited by Anthony J. G. Hey. Perseus Books, Reading, MA. (1999). 438 pages. $50.00. Computers and Mathematics with Applications, 37(10):173, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903358.

Anonymous:1999:BRFm

[Ano99-82] Anonymous. Book review: Finiteness and regularity in semigroups and formal languages: By A. de Luca and Stefano Varricchio. Springer-Verlag, Berlin. (1999). 240 pages. $44.95, DM 79.00, öS 577.00, sFr 72.00, GBP 30.50. Computers and Mathematics with Applications, 38 (11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912518.

Anonymous:1999:BRFe

[Ano99-83] Anonymous. Book review:

Floating, flowing, flying: Pieter J. Zandbergen’s life las innovator, inspirator and instigator in numerical fluid dynamics: Edited by D. Dijkstra, B. J. Geurts, J. G. M. Kuerten and H. K. Kuiken. Kluwer Academic Publishers, Dordrecht. (1998). 253 pages. $105.00. Computers and Mathematics with Applications, 37(10):169, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903218.

Anonymous:1999:BRFa

[Ano99-84] Anonymous. Book review:

Focal boundary value problems for differential and difference equations: By Ravi P. Agarwal. Kluwer Academic Publishers, Dordrecht. (1998). 289 pages. $140.00, NLG 260.00, GBP 89.00. Computers and Mathematics with Applications, 37(3): 130, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990367X.

Anonymous:1999:BRFb

[Ano99-85] Anonymous. Book review:

Focal boundary value problems for differential and difference equations: By Ravi P. Agarwal. Kluwer Academic Publishers, Dordrecht.

(1998). 289 pages. $140.00, NLG 260.00, GBP 89.00. Computers and Mathematics with Applications, 37(6): 130, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902316.

Anonymous:1999:BRFc

[Ano99-86]  Anonymous.  Book re-

view: Formal equivalence checking and design debugging: By Shi-Yu Huang and Kwang-Ting (Tim) Cheng. Kluwer Academic Publishers, Boston, MA. (1998). 229 pages. $106.50, NLG 245.00, GBP 72.50. Computers and Mathematics with Applications, 37(8):141, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901852.

Anonymous:1999:BRFn

[Ano99-87] Anonymous. Book review: Formal semantics and proof techniques for optimizing VHDL models: By Kothanda Umamageswaran, Sheetanshu L. Pandey and Philip A. Wilsey. Kluwer Academic Publishers, Boston, MA. (1999). 158 pages. $105.00, NLG 240.00, GBP 72.00. Computers and Mathematics with Applications, 38 (11–12):289, December 1999.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912397.

Anonymous:1999:BRFg

[Ano99-88]  Anonymous.  Book re-

view: Foundations of knowledge systems: With applications to databases and agents: By Gerd Wagner. Kluwer Academic Publishers, Boston, MA. (1998). 296 pages. $120.00, NLG 275.00, GBP 81.75. Computers and Mathematics with Applications, 38(1):123, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900159.

Anonymous:1999:BRFl

[Ano99-89] Anonymous. Book review:

Fractals for the classroom: Strategic activities, volume three: By Heinz-Otto Peitgen, Hartmut Jürgens, Dietmar Saupe, Evan Maletsky and Terry Perciante. National Council of Teachers of Mathematics/Springer-Verlag, New York. (1999). 107 pages. $24.95, DM 49.00, öS 358.00, sFr 45.50, GBP 19.00. Computers and Mathematics with Applications, 38 (11–12):283, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199911975.

Anonymous:1999:BRAg

[Ano99-90]  Anonymous.  Book re-

view: From animals to animats 5: Proceedings of the fifth international conference on simulation of adaptive behavior: Edited by Rolf Pfeifer, Bruce Blumberg, Jean-Arcady Meyer and Stewart W. Wilson. MIT Press, Cambridge, MA. (1998). 561 pages. $65.00. Computers and Mathematics with Applications, 37(8): 139–141, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901839.

Anonymous:1999:BRWc

[Ano99-91]  Anonymous.  Book re-

view: From white dwarfs to black holes: the legacy of S. Chandrasekhar: Edited by G. Srinivasan. University of Chicago Press, Chicago. (1999). 240 pages. $40.00, £31.95. Computers and Mathematics with Applications, 38(5–6):252, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219990127X.

Anonymous:1999:BRFd

[Ano99-92] Anonymous. Book review:

Fundamentals of neural network modeling: Neuropsychology and cognitive neuroscience: Edited by Randolph W. Parks, Daniel S. Levine and Debra L. Long. MIT Press, Cambridge, MA. (1998). 428 pages. $60.00. Computers and Mathematics with Applications, 37 (10):168, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903139.

Anonymous:1999:BRFj

[Ano99-93] Anonymous. Book review:

Fuzzy control of industrial systems: Theory and Applications: By Ian S. Shaw. Kluwer Academic Publishers, boston, MA. (1998). 192 pages. $115.00, NLG 260.00, GBP 78.20. Computers and Mathematics with Applications, 38(1):128, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900573.

Anonymous:1999:BRFi

[Ano99-94] Anonymous. Book review:

Fuzzy logic in data modeling: Semantics, constraints, and database design: By Guoquin Chen. Kluwer Academic Publishers, Boston, MA. (1998). 224 pages. $115.00, NLG 260.00, GBP 78.25. Computers and Mathematics with Applications, 38(1):125, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199900238.

Anonymous:1999:BRGm

[Ano99-95] Anonymous. Book review: Generalized convexity, generalized monotonicity: Recent results: Edited by Jean-Pierre Crouzeix, Juan-Enrique Martinez-Legaz and Michel Volle. Kluwer Academic Publishers, Dordrecht. (1998). 467 pages. $208.00, NLG 385.00, GBP 130.00. Computers and Mathematics with Applications, 37 (10):168, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903115.

Anonymous:1999:BRGe

[Ano99-96] Anonymous. Book review: Generalized quasilinearization for nonlinear problems: By V. Lakshmikantham and A. S. Vatsala. Kluwer Academic Publishers, Dordrecht. (1998). 276 pages. $140.00, NLG 260.00, GBP 89.00. Computers and Mathematics with Applications, 37(3): 133, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903917.

Anonymous:1999:BRGk

[Ano99-97] Anonymous. Book review:

Generalized quasilinearization for nonlinear problems: By V. Lakshmikantham and A. S. Vatsala. Kluwer Academic Publishers, Dordrecht. (1998). 276 pages. $140.00, NLG 260.00, GBP 89.00. Computers and Mathematics with Applications, 37(6): 133, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902559.

Anonymous:1999:BRGn

[Ano99-98] Anonymous. Book review:

Generalized riccati theory and robust control: a Popov function approach: By Vlad Ionescu, Cristian Oară and Martin Weiss. John Wiley & Sons, Inc., Chichester. (1999). 380 pages. $65.00. Computers and Mathematics with Applications, 37(10):174, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903498.

Anonymous:1999:BRGc

[Ano99-99] Anonymous. Book review:

Genetic programming and data structures: Genetic programming + data structures = automatic programming!: By W. B. Langdon. Kluwer Academic Publishers, Boston, MA. (1998). 278 pages. $125.00. NLG 285.00, GBP 85.00. Computers and Mathematics with Applications, 37(3):132, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903759.

Anonymous:1999:BRGi

[Ano99-100] Anonymous. Book review: Genetic programming and data structures: Genetic programming + data structures = automatic programming: By W. B. Langdon. Kluwer Academic Publishers, Boston, MA. (1998). 278 pages. $125.00. NLG 285.00, GBP 85.00. Computers and Mathematics with Applications, 37(6):132, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902390.

Anonymous:1999:BRGo

[Ano99-101] Anonymous. Book review: Geometric methods and optimization problems: By V. Boltyanski, H. Martini and V. Soltan. Kluwer Academic Publishers, Dordrecht. (1999). 429 pages. $204.00, NLG 340.00, GBP 119.00. Computers and Mathematics with Applications, 38 (11–12):284, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912038.

Anonymous:1999:BRGg

[Ano99-102] Anonymous. Book review:

Geometry of vector sheaves: an axiomatic approach to differential geometry. Volume I: Vector sheaves. general theory: By Anastosios Mallios. Kluwer Academic Publishers, Dordrecht. (1998). 441 pages. $350.00, NLG 650.00, GBP 220.00 for 2 volume set. Computers and Mathematics with Applications, 37 (6):130, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902274.

Anonymous:1999:BRGb

[Ano99-103] Anonymous. Book review:

Geometry of vector sheaves: an axiomatic approach to differential geometry. Volume II: Geometry. Examples and applications: By Anastosios Mallios. Kluwer Academic Publishers, Dordrecht. (1998). 436 pages. $350.00, NLG 650.00, GBP 220.00 for 2 volume set. Computers and Mathematics with Applications, 37(3):130, February 1999. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903656.

Anonymous:1999:BRGh

[Ano99-104] Anonymous. Book review: Geometry of vector sheaves: an axiomatic approach to differential geometry. Volume II: Geometry. examples and applications: By Anastosios Mallios. Kluwer Academic Publishers, Dordrecht. (1998). 436 pages. $350.00, NLG 650.00, GBP 220.00 for 2 volume set. Computers and Mathematics with Applications, 37(6):130, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902298.

Anonymous:1999:BRGa

[Ano99-105] Anonymous. Book review: Geometry of vector sheaves: an exiomatic approach to differential geometry. Volume I: Vector sheaves. General theory: By Anastosios Mallios. Kluwer Academic Publishers, Dordrecht. (1998). 441 pages. $350.00, NLG 650.00, GBP 220.00 for 2 volume set. Computers and Mathematics with Applications, 37(3):130, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903632.

Anonymous:1999:BRGd

[Ano99-106] Anonymous. Book review:

Graphical models for machine learning and digital communication: By Brendan J. Frey. MIT Press, Cambridge, MA. (1998). 195 pages. $32.00. Computers and Mathematics with Applications, 37(3):133, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903899.

Anonymous:1999:BRGj

[Ano99-107] Anonymous. Book review:

Graphical models for machine learning and digital communication: By Brendan J. Frey. MIT Press, Cambridge, MA. (1998). 195 pages. $32.00. Computers and Mathematics with Applications, 37(6):133, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902535.

Anonymous:1999:BRGp

[Ano99-108] Anonymous. Book review: Graphics and GUIs with Matlab(R), second edition: By Patrick Marchand. CRC Press, Boca Raton, FL. (1999). 445 pages. $39.95, DM 70.00, öS 511.90, sFr 64.00, GBP 27.00. Computers and Mathematics with Applications, 38 (11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219991249X.

Anonymous:1999:BRGf

[Ano99-109] Anonymous. Book review: Graphics programming in C++: By Mark Walmsley. Springer-Verlag, London. (1998). 250 pages. $54.95, DM 89.00, öS 650.00, sFr 81.00, GBP 29.50. Computers and Mathematics with Applications, 37(3):135, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904212.

Anonymous:1999:BRGl

[Ano99-110] Anonymous. Book review: Graphics programming in C++: By Mark Walmsley. Springer-Verlag, London. (1998). 250 pages. $54.95, DM 89.00, öS 650.00, sFr 81.00, GBP 29.50. Computers and Mathematics with Applications, 37(6):135, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902857.

Anonymous:1999:BRHc

[Ano99-111] Anonymous. Book review: Handbook of mathematics and computational science: By John W. Harris and Horst Stocker. Springer-Verlag, New York. (1998). 1028 pages. $29.95, DM 40.00, öS 292.00, sFr 37.00, GBP 19.50. Computers and Mathematics with Applications, 37(3):133, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903851.

Anonymous:1999:BRHh

[Ano99-112] Anonymous. Book review: Handbook of mathematics and computational science: By John W. Harris and Horst Stocker. Springer-Verlag, New York. (1998). 1028 pages. $29.95, DM 40.00, öS 292.00, sFr 37.00, GBP 19.50. Computers and Mathematics with Applications, 37 (6):133, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902493.

Anonymous:1999:BRHk

[Ano99-113] Anonymous. Book review:

Handbook of numerical analysis: Volume VI, numerical methods for solids (Part 3), numerical methods for fluids (Part 1): Edited by G. G.

Ciarlet and J. L. Lions. Elsevier, Amsterdam. (1998). 688 pages. $164.00. Computers and Mathematics with Applications, 37(10):173, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903413.

Anonymous:1999:BRHn

[Ano99-114] Anonymous. Book review: Handbook of splines: By Gheorghe Micula and Sandra Micula. Kluwer Academic Publishers, Dordrecht. (1999). 604 pages. $105.00, NLG 240.00, GBP 72.00. Computers and Mathematics with Applications, 38 (11–12):284, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912117.

Anonymous:1999:BRHa

[Ano99-115] Anonymous. Book review: Harmonic analysis in hypercomplex systems: By Yu. M. Berezansky and A. A. Kalyuzhnyi. Kluwer Academic Publishers, Dordrecht. (1998). 483 pages. $238.00, NLG 440.00, GBP 149.00. Computers and Mathematics with Applications, 37(3): 132, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903735.

Anonymous:1999:BRHf

[Ano99-116] Anonymous. Book review:

Harmonic analysis in hypercomplex systems: By Yu. M. Berezansky and A. A. Kalyuzhnyi. Kluwer Academic Publishers, Dordrecht. (1998). 483 pages. $238.00, NLG 440.00, GBP 149.00. Computers and Mathematics with Applications, 37(6): 132, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902377.

Anonymous:1999:BRHl

[Ano99-117] Anonymous. Book review: Hierarchical annotated action diagrams: an interface-oriented specification and verification method: By E. Cerny, B. Berkane, P. Girodias and K. Khordoc. Kluwer Academic Publishers, Boston, MA. (1998). 211 pages. $112.00, NLG 255.00, GBP 76.20. Computers and Mathematics with Applications, 38(5–6):255, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901517.

Anonymous:1999:BRHd

[Ano99-118] Anonymous. Book review:

High level motion processing: Computational, neurobiological, and phychophysical perspectives: Edited by Takeo Watanabe. MIT Press, Cambridge, MA. (1998). 417 pages. $60.00. Computers and Mathematics with Applications, 37(3):135, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904157.

Anonymous:1999:BRHi

[Ano99-119] Anonymous. Book review: High level motion processing: Computational, neurobiological, and phychophysical perspectives: Edited by Takeo Watanabe. MIT Press, Cambridge, MA. (1998). 417 pages. $60.00. Computers and Mathematics with Applications, 37(6):135, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902791.

Anonymous:1999:BRHm

[Ano99-120] Anonymous. Book review: High performance networking: IFIP TC-6 eighth international conference on high performance networking (HPN’98) Vienna, Austria, September 21-25, 1998: Edited by Harmen R. van As. Kluwer Academic Publishers, Boston, MA. (1998).

699 pages. $195.00, NLG 440.00, GBP 132.00. Computers and Mathematics with Applications, 38(5–6):257– 258, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901736.

Anonymous:1999:BRIn

[Ano99-121] Anonymous. Book review: Independent component analysis: By Te-Won Lee. Kluwer Academic Publishers, Boston, MA. (1998). 210 pages. $118.00, NLG 270.00, GBP 80.25. Computers and Mathematics with Applications, 38(1):128, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990055X.

Anonymous:1999:BRIt

[Ano99-122] Anonymous. Book review: Information design: Edited by Robert Jacobson. MIT Press, Cambridge, MA. (1999). 357 pages. $35.00. Computers and Mathematics with Applications, 38 (11–12):291, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912713.

Anonymous:1999:BRIb

[Ano99-123] Anonymous. Book review: Initial approximations and root finding methods: By Nikolay Kyurkchiev. Wiley-VCH Verlag, Berlin. (1998). 180 pages. $45.00. Computers and Mathematics with Applications, 37(3):135, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904170.

Anonymous:1999:BRIg

[Ano99-124] Anonymous. Book review: Initial approximations and root finding methods: By Nikolay Kyurkchiev. Wiley-VCH Verlag, Berlin. (1998). 180 pages. $45.00. Computers and Mathematics with Applications, 37(6):135, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219990281X.

Anonymous:1999:BRIc

[Ano99-125] Anonymous. Book review: Inside the database object model: By Donald K. Burleson. CRC Press, Boca Raton, FL. (1999). 226 pages. $69.95, DM 135.00, öS 986.00, sFr 123.00, GBP 52.00. Computers and Mathematics with Applications, 37(3):135, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904194.

Anonymous:1999:BRIh

[Ano99-126] Anonymous. Book review: Inside the database object model: By Donald K. Burleson. CRC Press, Boca Raton, FL. (1999). 226 pages. $69.95, DM 135.00, öS 986.00, sFr 123.00, GBP 52.00. Computers and Mathematics with Applications, 37 (6):135, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902833.

Anonymous:1999:BRIr

[Ano99-127] Anonymous. Book review: Integer and combinatorial optimization: By George Nemhauser and Laurence Wolsey. John Wiley & Sons, Inc., New York. (1999). 763 pages. £48.50. Computers and Mathematics with Applications, 38 (11–12):290, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912592.

Anonymous:1999:BRIm

[Ano99-128] Anonymous. Book review: Intelligent systems for manufacturing: Multi-Agent systems and virtual organizations: (Proceedings of the BASYS’98-Third IEEE/ IFIP International conference on information technology for balanced automation systems in manufacturing, Prague, Czech Republic, August 1998). Edited by Luis M. Camarinha-Matos, Hamideh Afsarmanesh and Vladimir Marik. Kluwer Academic Publishers, Boston, MA. (1998). 632 pages. $210.00, NLG 475.00, GBP 143.00. Computers and Mathematics with Applications, 38(1):124–125, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900196.

Anonymous:1999:BRId

[Ano99-129] Anonymous. Book review: Interactive linear algebra with Maple V: By Elias Deeba and Ananda Gunawardena. Springer-Verlag, New York. (1998). 316 pages. $49.95, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50 (CD-ROM included). Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904315.

Anonymous:1999:BRIi

[Ano99-130] Anonymous. Book review:

Interactive linear algebra with Maple V: By Elias Deeba and Ananda Gunawardena. Springer-Verlag, New York. (1998). 316 pages. $49.95, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50 (CD-ROM included). Computers and Mathematics with Applications, 37(6):136, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990295X.

Anonymous:1999:BRIs

[Ano99-131] Anonymous. Book review:

Interactive video-on-demand systems: Resource management and scheduling strategies: By T. P. Jimmy To and Babak Hamidzadeh. Kluwer Academic Publishers, Boston, MA. (1998). 137 pages. $105.00. Computers and Mathematics with Applications, 38(11–12):290, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199912555.

Anonymous:1999:BRIq

[Ano99-132] Anonymous. Book review: Introduction to circuit complexity: a uniform approach: By Heribert Vollmer. Springer-Verlag, Berlin. (1999). 270 pages. $44.95, DM 68.00, öS 497.00, sFr 62.00, GBP 26.00. Computers and Mathematics with Applications, 38(11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199912476.

Anonymous:1999:BRIe

[Ano99-133] Anonymous. Book review: Introduction to parallel algorithms: By C. Xavier and S. S. Iyengar. John Wiley & Sons, Inc., New York. (1998). 365 pages. $74.95. Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904418.

Anonymous:1999:BRIj

[Ano99-134] Anonymous. Book review: Introduction to parallel algorithms: By C. Xavier and S. S. Iyengar. John Wiley & Sons, Inc., New York. (1998). 365 pages. $74.95. Computers and Mathematics with Applications, 37(6):136, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219990305X.

Anonymous:1999:BRIl

[Ano99-135] Anonymous. Book review: Invitation to discrete mathematics: By Jirf Matousek and Jaroslav Nesetril. Clarendon Press, Oxford. (1998). 410 pages. $79.00 (cloth), $35.00 (paper). Computers and Mathematics with Applications, 37(10):174, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903553.

Anonymous:1999:BRIo

[Ano99-136] Anonymous. Book review:

IUTAM symposium on nonlinear singularities in deformation and flow: Proceedings of the IUTAM symposium held in Haifa, Israel, 17-21 March 1997: Edited by D. Durban and J. R. A. Pearson. Kluwer Academic Publishers, Dordrecht. (1999). 360 pages. $170.00, NLG 315.00, GBP 108.00. Computers and Mathematics with Applications, 38(5–6):253, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901311.

Anonymous:1999:BRIp

[Ano99-137] Anonymous. Book review:

IUTAM symposium on rheology of bodies with defects: Proceedings of the IUTAM symposium held in Beijing, China, 2–5 September 1997: Edited by Ren Wang. Kluwer Academic Publishers, Dordrecht. (1999). 292 pages. $133.00, NLG 245.00, GBP 84.00. Computers and Mathematics with Applications, 38(5–6):255–256, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901591.

Anonymous:1999:BRJb

[Ano99-138] Anonymous. Book review: Java^{T M} 2D graphics: By Jonathan Knudsen. O’Reilly, Sebastopol, CA. (1999). 339 pages. $29.95. Computers and Mathematics with Applications, 38 (11–12):282, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199911914.

Anonymous:1999:BRJa

[Ano99-139] Anonymous. Book review: Java^{T M} servlet programming: By Jason Hunter. O’Reilly, Sebastopol, CA. (1998). 510 pages. $32.95. Computers and Mathematics with Applications, 37(8):145, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902079.

Anonymous:1999:BRKa

[Ano99-140] Anonymous. Book review: Kant’s conception of moral character: the “critical” link of morality, Anthropology, and reflective judgment: By G. Felicitas Munzel. University of Chicago Press, Chicago. (1999). 377 pages. $53.00, £37.50 (cloth); $24.00, £17.00 (paper). Computers and Mathematics with Applications, 38(11–12):284, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199912075.

Anonymous:1999:BRLc

[Ano99-141] Anonymous. Book review:

Learning Perl/Tk: By Nancy Walsh. O’Reilly, Sebastopol, CA. (1999). 358 pages. $32.95. Computers and Mathematics with Applications, 37 (10):173, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903437.

Anonymous:1999:BRLe

[Ano99-142] Anonymous. Book review:

Lexical semantics and knowledge representation in multilingual text generation: By Manfred Stede. Kluwer Academic Publishers, Boston, MA. (1999). 219 pages. $129.00, NLG 265.00, GBP 84.00. Computers and Mathematics with Applications, 38 (11–12):287, December 1999.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912178.

Anonymous:1999:BRLb

[Ano99-143] Anonymous. Book review: Lingo in a nutshell: a desktop quick reference: By Bruce A. Epstein. O’Reilly, Sebastopol, CA. (1998). 609 pages. $19.95. Computers and Mathematics with Applications, 37(8):146, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902195.

Anonymous:1999:BRLd

[Ano99-144] Anonymous. Book review: Linguistic diversity in space and time: By Johanna Nichols. University of Chicago Press, Chicago. (1992). 358 pages. $19.00, GBP 15.25. Computers and Mathematics with Applications, 38 (5–6):258, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901773.

Anonymous:1999:BRLa

[Ano99-145] Anonymous. Book review: Logics for databases and information systems: Edited by Jan Chomicki and Gunter Saake. Kluwer Academic Publishers, Boston, MA. (1998). 430 pages. $79.95, NLG 175.00, GBP 53.00. Computers and Mathematics with Applications, 37 (8):143, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901931.

Anonymous:1999:BRMp

[Ano99-146] Anonymous. Book review:

Making physics: a biography of Brookhaven National Laboratory, 1946-1972: By Robert P. Crease. University of Chicago Press, Chicago. (1999). 434 pages. $38.00, GBP 30.50. Computers and Mathematics with Applications, 38(1):128, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900494.

Anonymous:1999:BRMv

[Ano99-147] Anonymous. Book review:

Mastering algorithms with C: By Kyle Loudon. O’Reilly, Sebastopol, CA. (1999). 540 pages. $34.95 (diskette included). Computers and Mathematics with Applications, 38(11–12):292, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199912750.

Anonymous:1999:BRMt

[Ano99-148] Anonymous. Book review: Mastering algorithms with perl: By Jon Orwant, Jarkko Hietaniemi and John Macdonald. O’Reilly, Sebastopol, CA. (1999). 684 pages. $34.95. Computers and Mathematics with Applications, 38 (11–12):290, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912634.

Anonymous:1999:BRMo

[Ano99-149] Anonymous. Book review: Mathematical physics: a modern introduction to its foundations: By Sadri Hassani. Springer-Verlag, New York. (1999). 1025 pages. $89.95, NLG 179.00, öS 1,307.00, sFr 162.00, GBP 69.00. Computers and Mathematics with Applications, 38(1):123, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199900172.

Anonymous:1999:BRMu

[Ano99-150] Anonymous. Book review: Mathematics for computer graphics applications (second edition): By M. E. Mortenson. Industrial Press, New York. (1999). 354 pages.

$39.95. Computers and Mathematics with Applications, 38 (11–12):290, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912658.

Anonymous:1999:BRMm

[Ano99-151] Anonymous. Book review:

Maxima and minima with applications: Practical optimization and duality: By Wilfred Kaplan. John Wiley & Sons, Inc., New York. (1999). 284 pages. $48.50. Computers and Mathematics with Applications, 37(10):174, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903474.

Anonymous:1999:BRMk

[Ano99-152] Anonymous. Book review:

Mechanics of sandwich structures: Proceedings of the EUROMECH 360 colloquim held in Saint-Étienne, France, 13-15 May 1997: Edited by A. Vautrin. Kluwer Academic Publishers, Dordrecht. (1998). 430 pages. $189.00, NLG 350.00, GBP 119.00. Computers and Mathematics with Applications, 37(8): 141–142, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219990189X.

Anonymous:1999:BRMr

[Ano99-153] Anonymous. Book review: Meromorphic functions and projective curves: By Kichoon Yang. Kluwer Academic Publishers, Dordrecht. (1999). 201 pages. $105.00, NLG 175.00, GBP 62.00. Computers and Mathematics with Applications, 38 (11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912373.

Anonymous:1999:BRMe

[Ano99-154] Anonymous. Book review: Modelling systems: Practical tools and techniques in software development: By John Fitzgerald and Peter Larsen. Cambridge University Press, New York. (1998). 269 pages. $85.00 (hardback); $39.95 (paperback) (CD-ROM included). Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904273.

Anonymous:1999:BRMj

[Ano99-155] Anonymous. Book review: Modelling systems: Practical tools and techniques in software development: By John Fitzgerald and Peter Larsen. Cambridge University Press, New York. (1998). 269 pages. $85.00 (hardback); $39.95 (paperback) (CD-ROM included). Computers and Mathematics with Applications, 37(6):136, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902912.

Anonymous:1999:BRMd

[Ano99-156] Anonymous. Book review: Modern graph theory: By Béla Bollobás. Springer-Verlag, New York. (1998). 394 pages. $34.95, DM 68.00, öS 497.00, sFr 62.00, GBP 26.00. Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904297.

Anonymous:1999:BRMi

[Ano99-157] Anonymous. Book review: Modern graph theory: By Béla Bollobás. Springer-Verlag, New York. (1998). 394 pages. $34.95, DM 68.00, öS 497.00, sFr 62.00, GBP 26.00. Computers and Mathematics with Applications, 37 (6):136, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902936.

Anonymous:1999:BRMa

[Ano99-158] Anonymous. Book review: MPI — The complete reference: Volume 1, the MPI core, second edition: By Marc Snir, Steve Otto, Steven Huss-Lederman, David Walker and Jack Dongarra. MIT Press, Cambridge, MA. (1998). 426 pages. $35.00. Computers and Mathematics with Applications, 37(3):130, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903590.

Anonymous:1999:BRMf

[Ano99-159] Anonymous. Book review: MPI — The complete reference: Volume 1, the MPI core, second edition: By Marc Snir, Steve Otto, Steven Huss-Lederman, David Walker and Jack Dongarra. MIT Press, Cambridge, MA (1998). 426 pages. $35.00. Computers and Mathematics with Applications, 37(6): 130, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902237.

Anonymous:1999:BRMb

[Ano99-160] Anonymous. Book review:

MPI-The complete reference: Volume 2, the MPI-2 extensions: By William Gropp, Steven Huss-Lederman, Andrew Lumsdaine, Ewing Lusk, Bill Nitzberg, William Saphir and Marc Snir. MIT Press, Cambridge, MA. (1998). 344 pages. $35.00. Computers and Mathematics with Applications, 37(3): 130, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903619.

Anonymous:1999:BRMg

[Ano99-161] Anonymous. Book review:

MPI-The complete reference: Volume 2, the MPI-2 extensions: By William Gropp, Steven Huss-Lederman, Andrew Lumsdaine, Ewing Lusk, Bill Nitzberg, William Saphir and Marc Snir. MIT Press, Cambridge, MA. (1998). 344 pages. $35.00. Computers and Mathematics with Applications, 37(6): 130, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902250.

Anonymous:1999:BRMq

[Ano99-162] Anonymous. Book review:

Multifractals and 1/f noise:

Wild self-affinity in physics (1963–1976): By Benoit B. Mandelbrot. Springer-Verlag, New York. (1999). 442 pages. $42.95, DM 89.00, öS 650.00, sFr 81.00, GBP 34.00. Computers and Mathematics with Applications, 38 (11–12):283, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199911951.

Anonymous:1999:BRMn

[Ano99-163] Anonymous. Book review: Musical networks: Parallel distributed perception and performance: Edited by Niall Griffith and Peter M. Todd. MIT Press, Cambridge, MA. (1999). 385 pages. $37.50. Computers and Mathematics with Applications, 38 (1):122, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900056.

Anonymous:1999:BRMs

[Ano99-164] Anonymous. Book review: MySQL and mSQL: By Randy Jay Yarger, George Reese and Tim King. O’Reilly, Sebastopol, CA. (1999). 487 pages. $34.95. Computers and Mathematics with Applications, 38(11–12):289, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199912439.

Anonymous:1999:BRNe

[Ano99-165] Anonymous. Book review:

New trends in mathematical programming: Homage to Steven Vajda: Edited by Franco Giannessi, Sándor Komlósi and Tamás Rapcsák. Kluwer Academic Publishers, Boston, MA. (1998). 314 pages. $140.00, NLG 260.00, GBP 119.00. Computers and Mathematics with Applications, 37(8):144, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901979.

Anonymous:1999:BRNi

[Ano99-166] Anonymous. Book review:

Non-standard computation: Molecular computation, cellular automata, evolutionary algorithms, quantum computers: By Tino Gramß, Stefan Bornholdt, Michael Groß, Melanie Mitchell and Thomas Pellizzari. Wiley-VCH, Weinheim, Germany. (1998). 221 pages. $34.95. Computers and Mathematics with Applications, 37(10):174, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990353X.

Anonymous:1999:BRNa

[Ano99-167] Anonymous. Book review: Nondifferentiable optimization and polynomial problems: By Naum Z. Shor. Kluwer Academic Publishers, Boston, MA. (1998). 394 pages. $179.00, NLG 330.00, GBP 113.00. Computers and Mathematics with Applications, 37(3):130, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903693.

Anonymous:1999:BRNc

[Ano99-168] Anonymous. Book review: Nondifferentiable optimization and polynomial problems: By Naum Z. Shor. Kluwer Academic Publishers, Boston, MA. (1998). 394 pages. $179.00, NLG 330.00, GBP 113.00. Computers and Mathematics with Applications, 37(6):130, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219990233X.

Anonymous:1999:BRNf

[Ano99-169] Anonymous. Book review: Nonlinear modeling: Advanced black-box techniques: Edited by Johan A. K. Suykens and Joos Vandewalle. Kluwer Academic Publishers, Boston, MA. (1998). 256 pages. $125.00, NLG 285.00, GBP 85.00. Computers and Mathematics with Applications, 37(10):169, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903176.

Anonymous:1999:BRNl

[Ano99-170] Anonymous. Book review:

Nonlinear programming and variational inequality problems: a unified approach: By Michael Patriksson. Kluwer Academic Publishers, Dordrecht. (1999). 334 pages. $168.00, NLG 280.00, GBP 99.00. Computers and Mathematics with Applications, 38 (5–6):257, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901712.

Anonymous:1999:BRNg

[Ano99-171] Anonymous. Book review:

Nonsmooth approach to optimization problems with equilibrium constraints: Theory, applications and numerical results: By Jiri Outrata, Michal Kocvara and Jochem Zowe. Kluwer Academic Publishers, Dordrecht. (1998). 273 pages. $129.00, NLG 240.00, GBP 82.00. Computers and Mathematics with Applications, 37(10):172, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903292.

Anonymous:1999:BRNh

[Ano99-172] Anonymous. Book review: Numerical methods for wave propagation: Selected contributions from the workshop held in Manchester, U.K., containing the Harten Memorial Lecture: Edited by E. F. Toro and J. F. Clarke. Kluwer Academic Publishers, Dordrecht. (1998). 385 pages. $170.00, NLG 315.00, GBP 107.00. Computers and Mathematics with Applications, 37(10):172, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903310.

Anonymous:1999:BROa

[Ano99-173] Anonymous. Book review: OpenSources: Voices from the open source revolution: Edited by Chris DiBona, Sam Ockman and Mark Stone. O’Reilly, Sebastopol, CA. (1999). 272 pages. $24.95. Computers and Mathematics with Applications, 37 (10):173, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903371.

Anonymous:1999:BROb

[Ano99-174] Anonymous. Book review:

Optimal filtering, volume I: Filtering of stochastic processes: By Vladimir Fomin. Kluwer Academic Publishers, Dordrecht. (1999). 375 pages. $173.00, NLG 320.00, GBP 109.00. Computers and Mathematics with Applications, 38(1):127, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900457.

Anonymous:1999:BROg

[Ano99-175] Anonymous. Book review:

Optimization methods for logical inference: By Vijay Chandru and John Hooker. John Wiley & Sons, Inc., New York. (1999). GBP 48.50. Computers and Mathematics with Applications, 38 (5–6):256, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990161X.

Anonymous:1999:BROc

[Ano99-176] Anonymous. Book review:

Oracle database administration: the essential reference: By David C. Kreines and Brian Laskey. O’Reilly, Sebastopol, CA. (1999). 560 pages. $36.95. Computers and Mathematics with Applications, 38(5–6):252, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901256.

Anonymous:1999:BROd

[Ano99-177] Anonymous. Book review: Oracle distributed systems: By Charles Dye. O’Reilly, Sebastopol, CA. (1999). 530 pages. $39.95 (diskette included). Computers and Mathematics with Applications, 38(5–6):252, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901190.

Anonymous:1999:BROe

[Ano99-178] Anonymous. Book review: Oracle PL/SQL language pocket reference: By Steven Feuerstein, Bill Pribyl and Chip Dawes. O’Reilly, Sebastopol, CA. (1999). 94 pages. $7.95. Computers and Mathematics with Applications, 38(5–6):252, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901232.

Anonymous:1999:BROf

[Ano99-179] Anonymous. Book review: Oracle PL/SQL programming, second edition: By Steven Feuerstein with Bill Pribyl. O’Reilly, Sebastopol, CA. (1997). 987 pages. $46.95 (diskette included). Computers and Mathematics with Applications, 38(5–6):252, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901219.

Anonymous:1999:BRPg

[Ano99-180] Anonymous. Book review:

Palm programming: the developer’s guide: By Neil Rhodes and Julie McKeehan. O’Reilly, Sebastopol, CA. (1999). 457 pages. $32.95 (CD-ROM included). Computers and Mathematics with Applications, 37(8):145, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902055.

Anonymous:1999:BRPj

[Ano99-181] Anonymous. Book review:

Parametric lie group actions on global generalised solutions of nonlinear PDEs: Including a solution to Hilbert’s fifth problem: By Elmér E. Rosinger. Kluwer Academic Publishers, Dordrecht. (1998). 231 pages. $106.00, NLG 195.00, GBP 67.00. Computers and Mathematics with Applications, 38 (1):127, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199900378.

Anonymous:1999:BRPq

[Ano99-182] Anonymous. Book review: Partial differential and integral equations: Edited by Heinrich G. W. Begehr, Robert P. Gilbert and Guo-Chun Wen. Kluwer Academic Publishers, Dordrecht. (1999). 369 pages. $168.00, NLG 280.00, GBP 99.00. Computers and Mathematics with Applications, 38(11– 12):291–292, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912737.

Anonymous:1999:BRPa

[Ano99-183] Anonymous. Book review: Partial differential equations and boundary value problems: By Viorel Barbu. Kluwer Academic Publishers, Dordrecht. (1998). 277 pages. $129.00, NLG 240.00, GBP 82.00. Computers and Mathematics with Applications, 37(3):133, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199903930.

Anonymous:1999:BRPd

[Ano99-184] Anonymous. Book review: Partial differential equations and boundary value problems: By Viorel Barbu. Kluwer Academic Publishers, Dordrecht. (1998). 277 pages. $129.00, NLG 240.00, GBP 82.00. Computers and Mathematics with Applications, 37(6):133, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902572.

Anonymous:1999:BRPh

[Ano99-185] Anonymous. Book review:

Partial differential equations: Sources and solutions: By Arthur David Snider. Prentice Hall, Upper Saddle River, NJ. (1999). 658 pages. $74.67. Computers and Mathematics with Applications, 38 (1):122, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900111.

Anonymous:1999:BRPp

[Ano99-186] Anonymous. Book review: Particles and fields: Edited by Gordon Semenoff and Luc Vinet. Springer-Verlag, New York. (1999). 489 pages. $79.95, DM 158.00, öS 1,154.00, sFr 143.00, GBP 61.00. Computers and Mathematics with Applications, 38 (11–12):291, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912695.

Anonymous:1999:BRPl

[Ano99-187] Anonymous. Book review: Photoshop in a nutshell: a desktop quick reference, (second edition): By Donnie O’Quinn. O’Reilly, Sebastopol, CA. (1999). $24.95. Computers and Mathematics with Applications, 38 (5–6):255, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901499.

Anonymous:1999:BRPm

[Ano99-188] Anonymous. Book review: PNG: the definitive guide: By Greg Roelofs. O’Reilly, Sebastopol, CA. (1999). 321 pages. $32.95. Computers and Mathematics with Applications, 38 (11–12):287, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219991221X.

Anonymous:1999:BRPr

[Ano99-189] Anonymous. Book review: Polynomials: an algorithmic approach: By Maurice Mignotte and Doru Stefanescu. Springer-Verlag, Singapore. (1999). 306 pages. $49.00, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50. Computers and Mathematics with Applications, 38(11–12):292, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199912816.

Anonymous:1999:BRPc

[Ano99-190] Anonymous. Book review:

Probability and stochastic processes: a friendly introduction for electrical and computer engineers: By Roy D. Yates and David J. Goodman. John Wiley & Sons, Inc., New York. (1999). 454 pages. $90.95. Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904352.

Anonymous:1999:BRPf

[Ano99-191] Anonymous. Book review:

Probability and stochastic processes: a friendly introduction for electrical and computer engineers: By Roy D. Yates and David J. Goodman. John Wiley & Sons, Inc., New York. (1999). 454 pages. $90.95. Computers and Mathematics with Applications, 37(6):136, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199902997.

Anonymous:1999:BRPo

[Ano99-192] Anonymous. Book review: Programming the be operating system: By Dan Parks Sydow. O’Reilly, Sebastopol, CA. (1999). 390 pages. $34.95. Computers and Mathematics with Applications, 38 (11–12):290, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912531.

Anonymous:1999:BRPb

[Ano99-193] Anonymous. Book review: Programming with data: a guide to the S language: By John Chambers. Springer-Verlag, New York. (1998). 469 pages. $39.95, DM 84.00, öS 614.00, sFr 76.50, GBP 32.50. Computers and Mathematics with Applications, 37(3):134, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199904091.

Anonymous:1999:BRPe

[Ano99-194] Anonymous. Book review: Programming with data: a guide to the S language: By John Chambers. Springer-Verlag, New York. (1998). 469 pages. $39.95, DM 84.00, öS 614.00, sFr 76.50, GBP 32.50. Computers and Mathematics with Applications, 37 (6):134, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902730.

Anonymous:1999:BRPk

[Ano99-195] Anonymous. Book review:

Programming with Qt: By Matthias Kalle Dalheimer. O’Reilly, Sebastopol, CA. (1999). 361 pages. $32.95. Computers and Mathematics with Applications, 38 (5–6):254, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901438.

Anonymous:1999:BRQ

[Ano99-196] Anonymous. Book review: Quantum measures and spaces: By G. Kalmbach. Kluwer Academic Publishers, Dordrecht. (1998). 343 pages. $159.00, NLG 295.00, GBP 99.00. Computers and Mathematics with Applications, 38(1):126, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900317.

Anonymous:1999:BRRc

[Ano99-197] Anonymous. Book review:

Random fields and stochastic partial differential equations: By Yu. A. Rozanov. Kluwer Academic Publishers, Dordrecht. (1998). 229 pages. $105.00, NLG 195.00, GBP 67.00. Computers and Mathematics with Applications, 37(8):144, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901992.

Anonymous:1999:BRRf

[Ano99-198] Anonymous. Book review: Random walks in the quarterplane: Algebraic methods, boundary value problems and applications: By Guy Fayolle, Roudolf Iasnogorodski and Vadim Malyshev. Springer-Verlag, Berlin. (1999). 156 pages. $89.95, DM 139.00, öS 1,015.00, sFr 126.50, GBP 53.50. Computers and Mathematics with Applications, 38 (11–12):292, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912798.

Anonymous:1999:BRRa

[Ano99-199] Anonymous. Book review: Regression graphics: Ideas for studying regressions through graphics: By R. Dennis Cook. John Wiley & Sons, Inc., New York. (1998). 349 pages. $79.95. Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990439X.

Anonymous:1999:BRRb

[Ano99-200] Anonymous. Book review:

Regression graphics: Ideas for studying regressions through graphics: By R. Dennis Cook. John Wiley & Sons, Inc., New York. (1998). 349 pages. $79.95. Computers and Mathematics with Applications, 37(6):136, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903036.

Anonymous:1999:BRRd

[Ano99-201] Anonymous. Book review:

Reuse methodology manual for system-on-a-chip designs: By Michael Keating and Pierre Bricaud. Kluwer Academic Publishers, Boston, MA. (1998). 224 pages. $94.00, NLG 215.00, GBP 64.00. Computers and Mathematics with Applications, 37 (10):172, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903334.

Anonymous:1999:BRSg

[Ano99-202] Anonymous. Book review:

Science without laws: By Ronald N. Giere. University of Chicagp Press, Chicago. (1999). 285 pages. $25.00, £17.50. Computers and Mathematics with Applications, 38(5–6):257, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901694.

Anonymous:1999:BRSe

[Ano99-203] Anonymous. Book review: Secret origins of modern microeconomics: Dupuit and the engineers: By Robert B. Ekelund, Jr. and Robert F. Hébert. University of Chicago Press, Chicago. (1999). 468 pages. $40.00, £31.95 (cloth). Computers and Mathematics with Applications, 38 (1):127, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900433.

Anonymous:1999:BRSc

[Ano99-204] Anonymous. Book review: Semidistributive modules and rings: By Askar A. Tuganbaev. Kluwer Academic Publishers, Dordrecht. (1998). 352 pages. $157.00, NLG 290.00, GBP 99.00. Computers and Mathematics with Applications, 38(1):125, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900251.

Anonymous:1999:BRSh

[Ano99-205] Anonymous. Book review:

Spectral/hp element methods for CFD: By George Em Karniadakis and Spencer J. Sherwin. Oxford University Press, New York. (1999). 390 pages. $75.00. Computers and Mathematics with Applications, 38 (11–12):284, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912051.

Anonymous:1999:BRSa

[Ano99-206] Anonymous. Book review:

Steiner minimal trees: By Dietmar Cieslik. Kluwer Academic Publishers, Dordrecht. (1998). 319 pages. $140.00, NLG 260.00, GBP 89.00. Computers and Mathematics with Applications, 37(3): 134, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904030.

Anonymous:1999:BRSb

[Ano99-207] Anonymous. Book review:

Steiner minimal trees: By Dietmar Cieslik. Kluwer Academic Publishers, Dordrecht. (1998). 319 pages. $140.00, NLG 260.00, GBP 89.00.

Computers and Mathematics with Applications, 37(6): 134, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902675.

Anonymous:1999:BRSj

[Ano99-208] Anonymous. Book review: Supersymmetry and equivariant de Rham theory: By Victor W. Guillemin and Shlomo Sternberg. Springer-Verlag, Berlin. (1999). 228 pages. $59.95, DM 98.00, öS 716.00, sFr 89.50, GBP 37.50. Computers and Mathematics with Applications, 38 (11–12):292, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219991283X.

Anonymous:1999:BRSf

[Ano99-209] Anonymous. Book review: Survivable networks: Algorithms for diverse routing: By Ramesh Bhandari. Kluwer Academic Publishers, Boston, MA. (1999). 200 pages. $110.00, NLG 250.00, GBP 75.00. Computers and Mathematics with Applications, 38(5–6):255, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901554.

Anonymous:1999:BRTd

[Ano99-210] Anonymous. Book review:

TCL/TK in a nutshell: a desktop quick reference: By Paul Raines and Jeff Tranter. O’Reilly, Sebastopol, CA. (1999). 440 pages. $24.95. Computers and Mathematics with Applications, 38 (5–6):255, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901530.

Anonymous:1999:BRTe

[Ano99-211] Anonymous. Book review:

Telepresence: Edited by P. J. Sheppard and G. R. Walker. Kluwer Academic Publishers, Boston, MA. (1999). 383 pages. $135.00, NLG 305.00, GBP 92.00. Computers and Mathematics with Applications, 38(5–6):256, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901670.

Anonymous:1999:BRTb

[Ano99-212] Anonymous. Book review: Testing of communicating systems: Proceedings of the IFIP TC6 11th International Workshop on Testing of Communicating Systems (IWTCS’98) August 31–September 2, 1998, Tomsk, Russia: Edited by A. Petrenko and Nina Yevtushenko. Kluwer Academic Publishers, Boston, MA. (1998). 364 pages. $159.00, NLG 326.00, GBP 89.95. Computers and Mathematics with Applications, 38 (1):123, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900135.

Anonymous:1999:BRCf

[Ano99-213] Anonymous. Book review: The complete Verilog book: By Vivek Sagdeo. Kluwer Academic Publishers, Boston, MA. (1998). 464 pages. $135.00, NLG 310.00, GBP 91.80. Computers and Mathematics with Applications, 37(8):146, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902158.

Anonymous:1999:BRFk

[Ano99-214] Anonymous. Book review: The financing of catastrophe risk: Edited by Kenneth A. Froot. University of Chicago Press, Chicago. (1999). 477 pages. $68.00, GBP 54.50. Computers and Mathematics with Applications, 38(5–6):258, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990175X.

Anonymous:1999:BRIk

[Ano99-215] Anonymous. Book review:

The importance of being fuzzy and other insights from the border between math and computers: By Arturo Sangalli. Princeton University Press, Princeton, NJ. (1998). 173 pages. $24.95, £18.95. Computers and Mathematics with Applications, 37(10):169, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990319X.

Anonymous:1999:BRIa

[Ano99-216] Anonymous. Book review: The invisible computer: Why good products can fail, the personal computer is so complex and information appliances are the solution: By Donald A. Norman, MIT Press, Cambridge, MA. (1998). 302 pages. $25.00. Computers and Mathematics with Applications, 37(3): 132, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903796.

Anonymous:1999:BRIf

[Ano99-217] Anonymous. Book review: The invisible computer: Why good products can fail, the personal computer is so complex and information appliances are the solution: By Donald A. Norman, MIT Press, Cambridge, MA. (1998). 302 pages. $25.00. Computers and Mathematics with Applications, 37(6):132, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902432.

Anonymous:1999:BRMc

[Ano99-218] Anonymous. Book review: The Matlab(R) 5 handbook: By Darren Redfern and Colin Campbell. Springer-Verlag, New York. (1998). 488 pages. $34.95, DM 69.00, öS 504.00, sFr 63.00, GBP 19.50. Computers and Mathematics with Applications, 37(3):134, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199904017.

Anonymous:1999:BRMh

[Ano99-219] Anonymous. Book review: The Matlab(R) 5 handbook: By Darren Redfern and Colin Campbell. Springer-Verlag, New York. (1998). 488 pages. $34.95, DM 69.00, öS 504.00, sFr 63.00, GBP 19.50. Computers and Mathematics with Applications, 37 (6):134, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902651.

Anonymous:1999:BRNj

[Ano99-220] Anonymous. Book review: The nature of cognition: Edited by Robert J. Sternberg. MIT Press, Cambridge, MA. (1999). 736 pages. $35.00. Computers and Mathematics with Applications, 38(1):122, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900032.

Anonymous:1999:BRNk

[Ano99-221] Anonymous. Book review:

The networking CD bookshelf: 6 bestselling books on CD-ROM: O’Reilly, Sebastopol, CA. (1999). $79.95. Includes bonus book, DNS and BIND, 3 ^{r}d edition. By Paul Albitz and Cricket Liu. O’Reilly, Sebastopol, CA. (1998). 482 pages. Computers and Mathematics with Applications, 38(1):127, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900391.

Anonymous:1999:BRNb

[Ano99-222] Anonymous. Book review:

The new renaissance: Computers and the next level of civilization: By Douglas S. Robertson. Oxford University Press, New York. (1998). 200 pages. $25.00. Computers and Mathematics with Applications, 37(3):133, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903991.

Anonymous:1999:BRNd

[Ano99-223] Anonymous. Book review: The new renaissance: Computers and the next level of civilization: By Douglas S. Robertson. Oxford University Press, New York. (1998). 200 pages. $25.00. Computers and Mathematics with Applications, 37(6):133, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902638.

Anonymous:1999:BRPn

[Ano99-224] Anonymous. Book review: The politics of pure science, new edition: By Daniel S. Greenberg. University of Chicago Press, Chicago. (1999). 311 pages. $49.00, £34.50 (cloth); $15.00, £10.50 (paper). Computers and Mathematics with Applications, 38(11–12):287, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199912191.

Anonymous:1999:BRSd

[Ano99-225] Anonymous. Book review: The science of energy: a cultural history of energy physics in Victorian Britain: By Crosbie Smith. University of Chicago Press, Chicago. (1998). 404 pages. $60.00, £47.95 (cloth); $25.00, £19.95 (paper). Computers and Mathematics with Applications, 38(1):127, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219990041X.

Anonymous:1999:BRSi

[Ano99-226] Anonymous. Book review:

The simple genetic algorithm: By Michael D. Vose. MIT Press, Cambridge, MA. (1999). 251 pages. $37.50. Computers and Mathematics with Applications, 38 (11–12):290, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912610.

Anonymous:1999:BRVa

[Ano99-227] Anonymous. Book review: The Verilog(R) hardware description language (fourth edition): By Donald E. Thomas and Philip R. Moorby. Kluwer Academic Publishers, Boston, MA. (1998). 354 pages. $98.00, NLG 225.00, GBP 66.70 (CD-ROM included). Computers and Mathematics with Applications, 37(8):145, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902110.

Anonymous:1999:BRTg

[Ano99-228] Anonymous. Book review: Theories of programming languages: By John C. Reynolds. Cambridge University Press, Cambridge. (1998). 500 pages. $49.95. Computers and Mathematics with Applications, 38 (11–12):288, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912270.

Anonymous:1999:BRTa

[Ano99-229] Anonymous. Book review: Time-varying systems and computations: By Patrick Dewilde and Alle-Jan Van Der Veen. Kluwer Academic Publishers, Boston, MA. (1998). 459 pages. $122.00, NLG 280.00, GBP 83.00. Computers and Mathematics with Applications, 37 (8):146, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902171.

Anonymous:1999:BRTc

[Ano99-230] Anonymous. Book review: Transact-SQL programming: By Kevin Kline, Lee Gould and Andrew Zanevsky. O’Reilly, Sebastopol, CA. (1999). 807 pages. $44.95 (CD-ROM included). Computers and Mathematics with Applications, 38(5–6):253, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901359.

Anonymous:1999:BRTf

[Ano99-231] Anonymous. Book review:

Two- and three-dimensional patterns of the face: By Peter W. Hallinan, Gaile G. Gordon, A. L. Yuille. Peter Giblin and David Mumford. A. K. Peters, Natick, MA. (1999). 262 pages. $48.00. Computers and Mathematics with Applications, 38 (11–12):283, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199912014.

Anonymous:1999:BRUb

[Ano99-232] Anonymous. Book review:

Unsupervised learning: Foundations of neural computation: Edited by Geoffrey Hinton and Terrance J. Sejnowski. MIT Press, Cambridge, MA. (1999). 389 pages. $30.00. Computers and Mathematics with Applications, 38(5–6):256, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901657.

Anonymous:1999:BRUa

[Ano99-233] Anonymous. Book review: Untangling ecological complexity: the macroscopic perspective: By Brian A. Maurer. University of Chicago Press, Chicago. (1999). 251 pages. $50.00, £39.95 (cloth); $18.00, £14.50 (paper). Computers and Mathematics with Applications, 37(8):145, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902134.

Anonymous:1999:BRVb

[Ano99-234] Anonymous. Book review: Vector bundles and their applications: By Glenys Luke and Alexander S. Mishchenko. Kluwer Academic Publishers, Dordrecht. (1998). 254 pages. $130.00. Computers and Mathematics with Applications, 37 (10):168, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903152.

Anonymous:1999:BRWa

[Ano99-235] Anonymous. Book review:

Web performance tuning: By Patrick Killelea. O’Reilly, Sebastopol, CA. (1998). 351 pages. $32.95. Computers and Mathematics with Applications, 37(3):132, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199903772.

Anonymous:1999:BRWb

[Ano99-236] Anonymous. Book review:

Web performance tuning: By Patrick Killelea. O’Reilly, Sebastopol, CA. (1998). 351 pages. $32.95. Computers and Mathematics with Applications, 37(6):132, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199902419.

Anonymous:1999:BRWe

[Ano99-237] Anonymous. Book review:

Webmaster in a nutshell, second edition: By Stephen Spainhour and Robert Eckstein. O’Reilly, Sebastopol, CA. (1999). 523 pages. $24.95. Computers and Mathematics with Applications, 38 (11–12):287, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199912233.

Anonymous:1999:BRWd

[Ano99-238] Anonymous. Book review: Writing Apache modules with perl and C: By Lincoln Stein and Doug MacEachern. O’Reilly, Sebastopol, CA. (1999). 724 pages. $34.95. Computers and Mathematics with Applications, 38 (5–6):254, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901414.

Anonymous:1999:BRY

[Ano99-239] Anonymous. Book review: Yerkes observatory, 1892-1950: the birth, near death, and resurrection of a scientific research institution: By Donald E. Osterbrock. University of Chicago Press, Chicago. (1997). 384 pages. $25.00, £19.95. Computers and Mathematics with Applications, 38(5–6):254, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199901451.

Anonymous:1999:EBa

[Ano99-240] Anonymous. Editorial Board. Computers and Mathematics with Applications, 37(1): iii–v, January 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199906211.

Anonymous:1999:EBb

[Ano99-241] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 38 (1):iii–v, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199900019.

Anonymous:1999:EMM

[Ano99-242] Anonymous. Errata to “Mathematical modeling and computer simulation of transient flow in centrifuge cascade pipe network with optimizing techniques”. Computers and Mathematics with Applications, 37(7):137–138, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000942.

See [MATA98].

Anonymous:1999:EOC

[Ano99-243] Anonymous. Errata to “Oscillation criteria for a nonlinear difference equation”. Computers and Mathematics with Applications, 37(3): 127, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000528.

See [LC98b].

Anonymous:1999:EEC

[Ano99-244] Anonymous. Errata to “The errors in calculating the pseudospectral differentiation matrices for Čebyšev–Gauss– Lobatto points”. Computers and Mathematics with Applications, 38(1):119, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001741.

See [BB99].

Anonymous:1999:ESI

[Ano99-245] Anonymous. Errata to “The shift-inverted J-Lanczos algorithm for the numerical solutions of large sparse algebraic Riccati equations”. Computers and Mathematics with Applications, 38 (3–4):253–255, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001996.

See [FLW97].

Anonymous:1999:EPG

[Ano99-246] Anonymous. Erratum to “A practical guide to direct optimization for planar gridgeneration”. Computers and Mathematics with Applications, 38(3–4):257, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900200X.

See [CO99].

Anonymous:1999:LC

[Ano99-247] Anonymous. List of contents. Computers and Mathematics with Applications, 37 (11–12):III–XI, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199913093.

Anonymous:1999:LCA

[Ano99-248] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 38(11–12):iii– xiii, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199003077.

Anonymous:2000:CSA

[Ano00] Anonymous. Corrigendum to “Spectral analysis of the matrix multisplitting method for one-dimensional model problem”. Computers and Mathematics with Applications, 39(5–6):259, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/

science/article/pii/S0898122100000596.

See [BSS98].

Ashino:1994:CSS

[ANV94] R. Ashino, M. Nagase, and R. Vaillancourt. A computer search for simple multi-dimensional wavelets. Computers and Mathematics with Applications, 28 (5):57–69, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001413.

Ashino:1996:CM

[ANV96] R. Ashino, M. Nagase, and R. Vaillancourt. A construction of multiwavelets. Computers and Mathematics with Applications, 32(3): 23–37, August 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001113.

Adeniyi:1991:EEN

[AO91] R. B. Adeniyi and P. Onumanyi. Error estimation in the numerical solution of ODE with the Tau method. Computers and Mathematics with Applications, 21(9): 19–27, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190122K.

Ainsworth:1993:PEEb

[AO93a] M. Ainsworth and J. T.

Oden. A posteriori error estimators for second order elliptic systems part 2. an optimal order process for calculating self-equilibrating fluxes. Computers and Mathematics with Applications, 26 (9):75–87, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390007I.

Ainsworth:1993:PEEa

[AO93b] Mark Ainsworth and J. Tinsley Oden. A posteriori error estimators for second order elliptic systems:Part 1. theoretical foundations and a posteriori error analysis. Computers and Mathematics with Applications, 25 (2):101–113, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390227M.

Agarwal:1997:BVP

[AO97] R. P. Agarwal and D. O’Regan.

Boundary value problems for general discrete systems on infinite intervals. Computers and Mathematics with Applications, 33(7):85–99, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000448.

Agarwal:1998:DSI

[AO98a] R. P. Agarwal and D. O’Regan. Discrete systems on infinite intervals. Computers and Mathematics with Applications, 35(9):97–105, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000613.

Agarwal:1998:FPA

[AO98b] R. P. Agarwal and D. O’Regan. A fixed-point approach for nonlinear discrete boundary value problems. Computers and Mathematics with Applications, 36(10–12):115–121, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219880014X.

AliAbadi:1998:NTM

[AO98c] M. H. AliAbadi and E. L. Ortiz. Numerical treatment of moving and free boundary value problems with the tau method. Computers and Mathematics with Applications, 35(8):53–61, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000443.

Agarwal:1999:CMC

[AO99a] R. P. Agarwal and D. O’Regan.

Continuation methods for closed, weakly closed, DKT, and WDKT maps. Computers and Mathematics with Applications, 38(7–8):81– 88, October 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002400.

Agarwal:1999:EMA

[AO99b] R. P. Agarwal and D. O’Regan.

An essential map approach for multivalued A-proper, weakly inward, DKT and weakly closed maps. Computers and Mathematics with Applications, 38(7–8):89– 100, October 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002412.

Agarwal:1999:ESS

[AO99c] R. P. Agarwal and D. O’Regan.

Existence of solutions to singular integral equations. Computers and Mathematics with Applications, 37(9): 25–29, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001108.

Agarwal:1999:GEN

[AO99d] R. P. Agarwal and D. O’Regan. Global existence for nonlinear operator inclusions. Computers and Mathematics with Applications, 38(11– 12):131–139, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002928.

Agarwal:1999:MSH

[AO99e] R. P. Agarwal and D. O’Regan. Multiple solutions for higherorder difference equations. Computers and Mathematics with Applications, 37(9): 39–48, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001121.

Agarwal:1999:MSD

[AO99f] R. P. Agarwal and D. O’Regan. Multiple solutions for the Dirichlet second-order boundary value problem of nonsingular type. Computers and Mathematics with Applications, 37(8):99–106, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001030.

Ahlbrandt:1994:DTO

[AP94] C. D. Ahlbrandt and A. Peterson. The (n, n)-disconjugacy of a 2n-th order linear difference equation. Computers and Mathematics with Applications, 28(1–3):1–9, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194000883.

Agarwal:1998:CER

[AP98a] R. P. Agarwal and M. Pituk.

Convergence to equilibria in recurrence equations. Computers and Mathematics with Applications, 36(10–12):357– 368, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800369.

Atici:1998:BPS

[AP98b] F. Atici and A. Peterson.

Bounds for positive solutions for a focal boundary value problem. Computers and Mathematics with Applications, 36(10–12):99–107, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800126.

Agarwal:1997:III

[APB97] R. P. Agarwal, J. Pečarić, and I. Brnetić. Improved integral inequalities in n independent variables. Computers and Mathematics with Applications, 33(8):27–38, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000515.

Amodio:1996:SPD

[APP96] P. Amodio, M. Paprzycki, and T. Politi. A survey of parallel direct methods for block bidiagonal linear systems on distributed memory computers. Computers and Mathematics with Applications, 31(7):111–127, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000211.

Akishin:1997:HNM

[APV97] P. G. Akishin, I. V. Puzynin, and S. I. Vinitsky. A hybrid numerical method for analysis of dynamics of the classical Hamiltonian systems. Computers and Mathematics with Applications, 34(2– 4):45–73, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700117X.

Adomian:1990:EPS

[AR90a] G. Adomian and R. Rach. Equality of partial solutions in the decomposition method for linear or nonlinear partial differential equations.

Computers and Mathematics with Applications, 19(12): 9–12, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090246G.

Adomian:1990:PND

[AR90b] G. Adomian and R. Rach.

Purely nonlinear differential equations. Computers and Mathematics with Applications, 20(1):1–3, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090062O.

Arney:1990:ECF

[AR90c] D. C. Arney and B. T.

Robinson. Exhibiting chaos and fractals with a microcomputer. Computers and Mathematics with Applications, 19(3):1–11, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090037K.

Adomian:1992:FCP

[AR92a] G. Adomian and R. Rach. A further consideration of partial solutions in the decomposition method. Computers and Mathematics with Applications, 23(1):51–64, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192900802.

Adomian:1992:GAP

[AR92b] G. Adomian and R. Rach. Generalization of Adomian polynomials to functions of several variables. Computers and Mathematics with Applications, 24(5–6):11–24, September 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290037I.

Adomian:1992:NTD

[AR92c] G. Adomian and R. Rach. Noise terms in decomposition solution series. Computers and Mathematics with Applications, 24(11):61–64, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/089812219290031C.

Adomian:1992:NTS

[AR92d] G. Adomian and R. Rach. Nonlinear transformation of series-part II. Computers and Mathematics with Applications, 23(10):79–83, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290058P.

Anastassiou:1992:MPT

[AR92e] George Anastassiou and Svetlozar T. Rachev. Moment problems and their applications to the stability of queueing models. Computers and Mathematics with Applications, 24(8–9):229–246, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290201R.

Anastassiou:1998:MPR

[AR98] G. A. Anastassiou and T. Rychlik. Moment problems on random rounding rules subject to two moment conditions. Computers and Mathematics with Applications, 36(1):9–19, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001047.

Anastassiou:1999:RUP

[AR99a] G. A. Anastassiou and T. Rychlik. Rates of uniform Prokhorov convergence of probability measures with given three moments to a Dirac one. Computers and Mathematics with Applications, 38(7–8):101– 119, October 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002424.

Azhar:1999:EAL

[AR99b] S. Azhar and J. H. Reif. Efficient algorithmic learning of the structure of permutation groups by examples. Computers and Mathematics with Applications, 37(10):105–132, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001297.

Arato:1990:QFG

[Ara90] N. M. Arató. Quadratic functionals of generalized Gaussian random fields. Computers and Mathematics with Applications, 19(1):1–7, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090076V.

Arato:1995:EPS

[Ara95] N. M. Arató. The estimate of potential in stochastic Schrödinger’s equation. Computers and Mathematics with Applications, 29(6): 1–5, March 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500001F.

Arany:1996:PCG

[Ara96] I. Arany. The preconditioned conjugate gradient method with incomplete factorization preconditioners. Computers and Mathematics with Applications, 31(4–5): 1–5, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002103.

Arato:1997:MEB

[Ara97] N. M. Arató. Mean estimation of Brownian sheet. Computers and Mathematics with Applications, 33(8): 13–25, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000503.

Arany:1999:SSE

[Ara99a] I. Arany. Solving systems of elastic bar structures by preconditioned conjugate gradient method. Computers and Mathematics with Applications, 38(9–10):125– 134, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002680.

Arato:1999:MPE

[Ara99b] M. Arató. In memory of Paul Erdős: March 26, 1913– September 20, 1996. Computers and Mathematics with Applications, 37(8):1–2, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199901815.

Argyros:1998:NCT

[Arg98] I. K. Argyros. A new convergence theorem for the Jarratt method in Banach space. Computers and Mathematics with Applications, 36(8):13–18, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001795.

Argyros:1999:NCT

[Arg99] I. K. Argyros. A new convergence theorem for the inexact Newton methods based on assumptions involving the second Fréchet derivative. Computers and Mathematics with Applications, 37(7): 109–115, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000917.

Abdella:1996:IDL

[ARI96] K. Abdella, H. Rasmussen, and I. I. Inculet. Interfacial deformation of liquid drops by electric fields at zero gravity. Computers and Mathematics with Applications, 31(10):67–82, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000533.

Adomian:1991:NAD

[ARM91] G. Adomian, R. Rach, and R. Meyers. Numerical algorithms and decomposition. Computers and Mathematics with Applications, 22(8): 57–61, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190013T.

Arnow:1990:RPC

[Arn90] B. J. Arnow. Representation of permutations, combinations and dihedral elements as trees. Computers and Mathematics with Applications, 20(3):63–67, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090031E.

Ayala:1998:ODP

[ARS98] V. Ayala, O. Rojo, and R. Soto. Observability of the direct product of bilinear control systems on lie groups. Computers and Mathematics with Applications, 36 (3):107–112, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001321.

Arato:1991:P

[AS91] M. Arató and T. Szántai. Preface. Computers and Mathematics with Applications, 21(1):vii, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190225S.

Almasi:1993:QMN

[AS93] B. Almasi and J. Sztrik. A queueing model for a non-homogeneous terminal system subject to BR992. Computers and Mathematics with Applications, 25 (4):105–111, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390253R.

Antonoiu:1995:SSD

[AS95] G. Antonoiu and P. K. Srimani. A self-stabilizing distributed algorithm to construct an arbitrary spanning tree of a connected graph. Computers and Mathematics with Applications, 30(9): 1–7, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500143M.

Avdelas:1996:BRK

[AS96a] G. Avdelas and T. E. Simos.

Block Runge–Kutta methods for periodic initial-value problems. Computers and Mathematics with Applications, 31(1):69–83, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500183Y.

Avdelas:1996:EMN

[AS96b] G. Avdelas and T. E.

Simos. Embedded methods for the numerical solution of the Schrödinger equation. Computers and Mathematics with Applications, 31(2):85–102, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001964.

Antoniou:1997:QSD

[AS97] I. Antoniou and Z. Suchanecki. Quantum systems with diagonal singularity: Their logic and intrinsic irreversibility. Computers and Mathematics with Applications, 34(5–6):509–531, September 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219700151X.

Antonoiu:1998:SSD

[AS98] G. Antonoiu and P. K. Srimani. A self-stabilizing distributed algorithm for minimal spanning tree problem in a symmetric graph. Computers and Mathematics with Applications, 35(10):15–23, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000686.

Almasi:1999:OPP

[AS99] B. Almási and J. Sztrik. Optimization problems on the performance of a nonreliable terminal system. Computers and Mathematics with Applications, 38(3–4):13–21, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002023.

Aoki:1998:RGF

[ASW98] K. Aoki, Y. Sone, and Y. Waniguchi. A rarefied gas flow induced by a temperature field: Numerical analysis of the flow between two coaxial elliptic cylinders with different uniform temperatures. Computers and Mathematics with Applications, 35 (1–2):15–28, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002551.

Auger:1990:CLC

[Aug90] P. Auger. Cell lifetime and chemical kinetics. Computers and Mathematics with Applications, 20(4–6):79–85, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090317D.

Amato:1998:WSA

[AV98] U. Amato and D. T. Vuza.

Wavelet simultaneous approximation from samples affected by noise. Computers and Mathematics with Applications, 36(5):101–111, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001539.

Agarwal:1991:OEB

[AW91] Ravi P. Agarwal and Patricia J. Y. Wong. Optimal error bounds for the derivatives of two point Hermite interpolation. Computers and Mathematics with Applications, 21(8):21–35, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900489.

Aizawa:1994:SSP

[AW94] A. N. Aizawa and B. W.

Wah. A sequential sampling procedure for genetic algorithms. Computers and Mathematics with Applications, 27(9–10):77–82, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901260.

Agarwal:1997:PSI

[AWY97] R. P. Agarwal, Fu-Hsiang Wong, and Shiueh-Ling Yu. Positive solutions of integro-differential inequalities. Computers and Mathematics with Applications, 34 (10):61–69, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002071.

Anastassiou:1996:SAF

[AZ96] G. A. Anastassiou and M. Zhou. Simultaneous approximation of the Feller operator. Computers and Mathematics with Applications, 31 (10):31–44, May 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000508.

Azbelev:1999:ICT

[Azb99] N. V. Azbelev. Impact of certain traditions on development of the theory of differential equations. Computers and Mathematics with Applications, 37(4–5):1–8, February/March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000541.

Benouaz:1996:LSA

[BA96] T. Benouaz and O. Arino.

Least square approximation of a nonlinear ordinary differential equation. Computers and Mathematics with Applications, 31(8):69–84, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000326.

Benouaz:1998:OAI

[BA98] T. Benouaz and O. Arino.

Optimal approximation of the initial value problem. Computers and Mathematics with Applications, 36(1): 21–32, July 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001059.

Babovsky:1998:DNS

[Bab98] H. Babovsky. Discretization and numerical schemes for steady kinetic model equations. Computers and Mathematics with Applications, 35 (1–2):29–40, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002563.

Bashir-Ali:1998:LDC

[BACS98] Z. Bashir-Ali, J. R. Cash, and H. H. M. Silva. Lobatto deferred correction for stiff two-point boundary value problems. Computers and Mathematics with Applications, 36(10–12):59–69, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800096.

Badea:1995:FMS

[Bad95]  C. Badea.  K-functionals

and moduli of smoothness of functions defined on compact metric spaces. Computers and Mathematics with Applications, 30(3–6):23–31, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000836.

Bagchi:1990:BVP

[Bag90] Arunabha Bagchi. Boundary value processes: Estimation and identification. Computers and Mathematics with Applications, 19(1):9–19, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090077W.

Bai:1996:NDP

[Bai96a] S. X. Bai. A note on determining pure-strategy equilibrium points of bimatrix games. Computers and Mathematics with Applications, 32(7):29–35, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001538.

Bai:1996:CPI

[Bai96b] Zhong-Zhi Bai. The convergence of parallel iteration algorithms for linear complementarity problems. Computers and Mathematics with Applications, 32(9): 1–17, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001721.

Bai:1996:MCC

[Bai96c] Zhong-Zhi Bai. The monotone convergence of a class of parallel nonlinear relaxation methods for nonlinear complementarity problems. Computers and Mathematics with Applications, 31 (12):17–33, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000739.

Bai:1996:NCT

[Bai96d] Zhong-Zhi Bai. New comparison theorem for the nonlinear multisplitting relaxation method for the nonlinear complementarity problems. Computers and Mathematics with Applications, 32(4):41–48, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600123X.

Bai:1996:MCR

[Bai96e] Zhongzhi Bai. The monotone convergence rate of the parallel nonlinear AOR method. Computers and Mathematics with Applications, 31(7):1–8, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000132.

Bai:1996:PNA

[Bai96f] Zhongzhi Bai. Parallel nonlinear AOR method and its convergence. Computers and Mathematics with Applications, 31(2):21–31, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001905.

Bakry:1990:NMC

[Bak90] S. H. Bakry. A new method for computing Erlang-b formula. Computers and Mathematics with Applications, 19(2):73–74, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900099.

Baker:1992:MGC

[Bak92] Timothy Baker. Mesh generation for the computation of flowfields over complex aerodynamic shapes. Computers and Mathematics with Applications, 24(5– 6):103–127, September 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290044I.

Baldwin:1990:CMU

[Bal90]  J. F. Baldwin.  Compu-

tational models of uncertainty reasoning in expert systems. Computers and Mathematics with Applications, 19 (11):105–119, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090153B.

Balla:1996:BCT

[Bal96a] K. Balla. Boundary conditions and their transfer for differential-algebraic equations of index 1. Computers and Mathematics with Applications, 31(10):1–5, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000478.

Balla:1996:LSL

[Bal96b] K. Balla. Linear subspaces for linear DAEs of index

1. Computers and Mathematics with Applications, 31 (4–5):81–86, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002197.

Balla:1998:ABC

[Bal98] K. Balla. Asymptotic behavior of certain Riccati difference equations. Computers and Mathematics with Applications, 36(10–12):243–250, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800254.

Balla:1999:CSD

[Bal99] K. Balla. Characterization of solutions in the discretization of a parabolic equation on an infinite strip. Computers and Mathematics with Applications, 38(9– 10):101–106, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002655.

Banzhaf:1993:SRS

[Ban93] W. Banzhaf. Self-replicating sequences of binary numbers. Computers and Mathematics with Applications, 26(7): 1–8, October 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390046X.

Barnett:1991:SCP

[Bar91a] Michael P. Barnett. Summing P_{n}(cos θ)/p(n) for certain polynomials p(n). Computers and Mathematics with Applications, 21(10):79–86, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900629.

Barr:1991:AAS

[Bar91b] Thomas H. Barr. Approximation for age-structured population models using projection methods. Computers and Mathematics with Applications, 21(5):17–40, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190213N.

Barnett:1993:IRF

[Bar93] M. P. Barnett. Implicit rule formation in symbolic computation. Computers and Mathematics with Applications, 26 (10):35–50, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390024P.

Baritompa:1994:MBD

[Bar94]  W. Baritompa.  Multidi-

mensional bisection: a dual viewpoint. Computers and Mathematics with Applications, 27(4):11–22, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900515.

Bandtlow:1997:RDS

[BAS97] O. F. Bandtlow, I. Antoniou, and Z. Suchanecki. Resonances of dynamical systems and Fredholm–Riesz operators on rigged Hilbert spaces. Computers and Mathematics with Applications, 34(2– 4):95–102, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001193.

Bassett:1998:KS

[Bas98] B. Bassett. ∫ e^{−x^{2}} dx and the Kink soliton. Computers and Mathematics with Applications, 36(4):37–45, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001394.

Baxter:1992:ACF

[Bax92] B. J. C. Baxter. The asymptotic cardinal function of the multiquadratic φ(r) = (r^{2} + c^{2})^{12} as c → ∞. Computers and Mathematics with Applications, 24(12): 1–6, December 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290166F.

Bini:1993:FDT

[BB93] D. Bini and E. Bozzo.

Fast discrete transforms by means of eigenpolynomials. Computers and Mathematics with Applications, 26 (9):35–52, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390004F.

Baltensperger:1999:ECP

[BB99] R. Baltensperger and J.- P. Berrut. The errors in calculating the pseudospectral differentiation matrices for Čebyšev–Gauss–Lobatto points. Computers and Mathematics with Applications, 37(1):41–48, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198002405.

See errata [Ano99-244].

Basios:1997:MAO

[BBI97] V. Basios, A. Yu. Bonushkina, and V. V. Ivanov. A method for approximating one-dimensional functions. Computers and Mathematics with Applications, 34(7– 8):687–693, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001697.

Babloyantz:1997:AUC

[BBKN97] A. Babloyantz, N. A. Bobylev, S. K. Korovin, and A. P. Nosov. Approximating of unstable cycle in nonlinear autonomous systems. Computers and Mathematics with Applications, 34(2–4):333–354, July/August 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001314.

Barker:1995:BEK

[BBM95] A. L. Barker, D. E. Brown, and W. N. Martin. Bayesian estimation and the Kalman filter. Computers and Mathematics with Applications, 30 (10):55–77, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500156S.

Babuska:1999:DFE

[BBO99] I. Babuska, C. E. Baumann, and J. T. Oden. A discontinuous hp finite element method for diffusion problems: 1-D analysis. Computers and Mathematics with Applications, 37(9):103–122, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001170.

Baart:1997:TIC

[BC97] M. L. Baart and M. A.

Coetzee. Transfinite interpolation of conic triangles using quadric patches. Computers and Mathematics with Applications, 34 (9):75–85, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001909.

Brimberg:1998:NCS

[BC98] J. Brimberg and R. Chen. A note on convergence in the single facility minisum location problem. Computers and Mathematics with Applications, 35(9):25–31, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000546.

Boudiaf:1990:MOM

[BCB90] A. Boudiaf, Y. Cherruault, and R. Belle. Modelling and optimization of meiotic cell division of xenopus oocytes. Computers and Mathematics with Applications, 20(4–6):67–72, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090315B.

Buzano:1995:NPI

[BCC95] M. L. Buzano, S. E. Corno, and I. Cravero. A new procedure for integrating the point kinetic equations for fission reactors. Computers and Mathematics with Applications, 29(5):5–19, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400245G.

Buet:1998:RBO

[BCD98] C. Buet, S. Cordier, and P. Degond. Regularized Boltzmann operators. Computers and Mathematics with Applications, 35(1–2):55– 74, January 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002587.

Benedetto:1999:HMA

[BCGP99] D. Benedetto, E. Caglioti, F. Golse, and M. Pulvirenti. A hydrodynamic model arising in the context of granular media. Computers and Mathematics with Applications, 38 (7–8):121–131, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002436.

Barakahani:1990:ILT

[BD90] A. Barakahani and R. S.

Dahiya. Iterative Laplace transforms of two variables. Computers and Mathematics with Applications, 19(6): 71–93, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090048O.

Buttimore:1996:EAL

[BD96] N. H. Buttimore and D. A. Mac Dónaill. The exploitation of assembly language instructions in biological text manipulation: II. Amino acid sequences. Computers and Mathematics with Applications, 32(11):39–45, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122196001952.

Baszenski:1995:UAA

[BDT95] G. Baszenski, F.-J. Delvos, and M. Tasche. A united approach to accelerating trigonometric expansions. Computers and Mathematics with Applications, 30(3– 6):33–49, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000844.

Beckmann:1990:DPS

[Bec90] M. J. Beckmann. Dynamic programming and the secretary problem. Computers and Mathematics with Applications, 19(11):25–28, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090145A.

Beck:1996:VIP

[Bec96] S. D. Beck. The virtual instrument paradigm: an alternative approach to interactive computer music. Computers and Mathematics with Applications, 32(1):5–10, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600082X.

Bedivan:1995:GFE

[Bed95a] D. M. Bedivan. A Galerkin finite element method for an optimal shape design semiconductor problem. Computers and Mathematics with Applications, 30(7):37–46, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219500124H.

Bedivan:1995:TGM

[Bed95b] D. M. Bedivan. A twogrid method for solving elliptic problems with inhomogeneous boundary conditions. Computers and Mathematics with Applications, 29(6):59–66, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500007L.

Beineke:1997:BGS

[Bei97]  L. W. Beineke.  Bipla-

nar graphs: a survey. Computers and Mathematics with Applications, 34(11): 1–8, December 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002149.

Belenky:1994:PGP

[Bel94a] A. S. Belenky. A 3-person game on polyhedral sets. Computers and Mathematics with Applications, 28 (5):53–56, September 1994.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001405.

Bellman:1994:NLB

[Bel94b] Nina Bellman. Notes on learning Bellman style. Computers and Mathematics with Applications, 27(9–10):xiii, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901171.

Belenky:1997:PGP

[Bel97a]  A. S. Belenky.  A 2-

person game on a polyhedral set of connected strategies. Computers and Mathematics with Applications, 33(6):99–125, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000370.

Belenky:1997:NGP

[Bel97b] A. S. Belenky. A noncooperative game on polyhedral sets. Computers and Mathematics with Applications, 33(10):125–133, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000813.

Benders:1999:SPA

[Ben99] L. P. M. Benders. Specification and performance analysis of embedded systems with coloured Petri nets. Computers and Mathematics with Applications, 37 (11–12):177–190, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900156X.

Berry:1991:SAM

[Ber91] R. Stephen Berry. Symmetry in atomic and molecular systems. Computers and Mathematics with Applications, 21(10):39–52, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190060H.

Bernard:1992:MSI

[Ber92] Robert S. Bernard. A Mac-Cormack scheme for incompressible flow. Computers and Mathematics with Applications, 24(5–6):151– 168, September 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290046K.

Bertram:1993:NIB

[Ber93] B. Bertram. Numerical implementation of a brakhagelike technique for weakly singular integral equations using product integration. Computers and Mathematics with Applications, 26 (5):21–26, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900698.

Bertolazzi:1996:PCS

[Ber96]  E. Bertolazzi.  Positive

and conservative schemes for mass action kinetics. Computers and Mathematics with Applications, 32 (6):29–43, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001423.

Bera:1998:PWR

[Ber98]  R. K. Bera.  Propaga-

tion of waves in random rotating infinite magnetothermo-visco-elastic medium. Computers and Mathematics with Applications, 36 (9):85–102, November 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001941.

Bestor:1996:MOC

[Bes96] C. Bestor. MAX as an overall control mechanism for multidiscipline installation art. Computers and Mathematics with Applications, 32(1):11–16, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000831.

Benchekroun:1991:NPL

[BF91] Bahia Benchekroun and James E. Falk. A nonconvex, piecewise linear optimization problem. Computers and Mathematics with Applications, 21(6–7):77–85, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190162W.

Bedivan:1995:LSM

[BF95] D. M. Bedivan and G. J.

Fix. Least squares methods for optimal shape design problems. Computers and Mathematics with Applications, 30(2):17–25, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000749.

Belhaq:1997:HOA

[BF97] M. Belhaq and A. Fahsi.

Higher-order approximation of subharmonics close to strong resonances in the forced oscillators. Computers and Mathematics with Applications, 33(8):133–144, April 1997. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000618.

Bussa:1997:ACA

[BFF^{+}97] M. P. Bussa, L. Fava, L. Ferrero, A. Grasso, V. V. Ivanov, I. V. Kisel, E. V. Konotopskaya, and G. B. Pontecorvo. Application of a cellular automaton for recognition of straight tracks in the spectrometer DISTO. Computers and Mathematics with Applications, 34(7– 8):695–701, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001703.

Butzer:1994:SFW

[BFR94] P. L. Butzer, A. Fischer, and K. Rückforth. Scaling functions and wavelets with vanishing moments. Computers and Mathematics with Applications, 27(3): 33–39, February 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900442.

Brankin:1990:CAB

[BG90] R. W. Brankin and I. Gladwell. Codes for almost block diagonal systems. Computers and Mathematics with Applications, 19(7):1–6, ????

1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090188P.

Bunch:1991:SMG

[BG91a] K. J. Bunch and R. W.

Grow. A simple method to generate test matrices of known eigenvalue or singular value spectra. Computers and Mathematics with Applications, 22(7):65–67, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901835.

Bunch:1991:TSE

[BG91b] K. J. Bunch and R. W.

Grow. A technique to shift an eigenvalue of a complex matrix to accelerate convergence of the power and inverse power method. Computers and Mathematics with Applications, 21(4):11–16, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190242V.

Borno:1995:PAS

[BG95] I. Borno and Z. Gajic. Parallel algorithm for solving coupled algebraic Lyapunov equations of discrete-time jump linear systems. Computers and Mathematics with Applications, 30(7):1–4, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500119J.

Breuer:1990:MLM

[BGEZ90] S. Breuer, J. Gal-Ezer, and G. Zwas. Microcomputer laboratories in mathematics education. Computers and Mathematics with Applications, 19(3):13–34, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090038L.

Birkhoff:1990:DMS

[BGL90] G. Birkhoff, E. C. Gartland, Jr., and R. E. Lynch. Difference methods for solving convection–diffusion equations. Computers and Mathematics with Applications, 19 (11):147–160, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090158G.

Belyakov:1997:ASS

[BGL97] L. A. Belyakov, L. Yu. Glebsky, and L. M. Lerman. Abundance of stable stationary localized solutions to the generalized 1D Swift–Hohenberg equation. Computers and Mathematics with Applications, 34 (2–4):253–266, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001272.

Bonsignori:1990:LS

[BGLT90] A. Bonsignori, F. Giannotti, L. Lucchesi, and F. Turini. The logiform system. Computers and Mathematics with Applications, 20(9–10):83– 99, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090114Y.

Bharadwaj:1995:ELD

[BGM95] V. Bharadwaj, D. Ghose, and V. Mani. An efficient load distribution strategy for a distributed linear network of processors with communication delays. Computers and Mathematics with Applications, 29(9):95–112, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000392.

Bohm:1997:ESQ

[BGM97] A. Bohm, M. Gadella, and S. Maxson. Extending the stationary quantum mechanics of being to a nonstationary quantum theory of becoming and decaying. Computers and Mathematics with Applications, 34(5–6):427– 466, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001491.

Brink:1995:TAD

[BGO95] C. Brink, D. M. Gabbay, and H. J. Ohlbach. Towards automating duality. Computers and Mathematics with Applications, 29(2):73–90, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194002179.

Boisen:1992:AGT

[BGZ92] M. B. Boisen, Jr., G. V. Gibbs, and Z. G. Zhang. An application of graph theory to the estimation of bond numbers in crystals. Computers and Mathematics with Applications, 23(10):99–102, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290060U.

Buckwitz:1990:NMD

[BH90] D. Buckwitz and H.-G. Holzhütter. A new method to discriminate between enzymekinetic models. Computers and Mathematics with Applications, 20(4–6):117– 126, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090321A.

Benson:1991:BBO

[BH91] Harold P. Benson and Reiner Horst. A branch and boundouter approximation algorithm for concave minimization over a convex set. Computers and Mathematics with Applications, 21(6– 7):67–76, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190161V.

Bhattacharya:1994:RPT

[Bha94] P. Bhattacharya. The representation of permutations by trees. Computers and Mathematics with Applications, 28(9):67–71, November 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001782.

Benczur:1995:DEN

[BHK95] A. Benczúr, C. Hajas, and G. Kovács. Datalog extension for nested relations. Computers and Mathematics with Applications, 30 (12):51–79, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500174W.

Binder:1991:MIL

[Bin91] P. M. Binder. Machine iteration of a linear function: Local behaviour. Computers and Mathematics with Applications, 21(2–3):133–140, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190091H.

Bird:1998:RAC

[Bir98] G. A. Bird. Recent advances and current challenges for DSMC. Computers and Mathematics with Applications, 35(1–2):1–14, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700254X.

Bisoi:1991:MPG

[Bis91] A. K. Bisoi. A Maple program to generate orthonormal polynomials. Computers and Mathematics with Applications, 22(9):1–5, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190201E.

Bonushkina:1997:MDA

[BIZ97] A. Yu. Bonushkina, V. V. Ivanov, and P. V. Zrelov. Multivariate data analysis based on the ω_{n}^{k}-criteria and multilayer perceptron.

Computers and Mathematics with Applications, 34(7– 8):677–685, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001685.

Bronson:1990:MDS

[BJ90] R. Bronson and C. Jacobson. Modeling the dynamics of social systems. Computers and Mathematics with Applications, 19(3):35–42, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090039M.

Buckwitz:1990:KMP

[BJH90] D. Buckwitz, G. Jacobasch, and H.-G. Holzhütter. A kinetic model of phosphofructokinase from Plasmodium berghei — application of a novel procedure for model discrimination. Computers and Mathematics with Applications, 20(4–6):103–106, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090319F.

Balbas:1999:DTS

[BJH99] A. Balbas, P. Jimenez, and A. Heras. Duality theory and slackness conditions in multiobjective linear programming. Computers and Mathematics with Applications, 37(4– 5):101–109, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000620.

Borm:1993:PEB

[BJPT93] P. E. M. Borm, M. J. M. Jansen, J. A. M. Potters, and S. H. Tijs. Pareto equilibria for bimatrix games. Computers and Mathematics with Applications, 25(10– 11):19–25, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902773.

Belishev:1991:BCW

[BK91] Michael I. Belishev and Yaroslav V. Kurylev. Boundary control, wave field continuation and inverse problems for the wave equation. Computers and Mathematics with Applications, 22(4–5):27–52, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190130V.

Bavinck:1998:DOS

[BK98] H. Bavinck and R. Koekoek. Difference operators with Sobolev type Meixner polynomials as eigenfunctions. Computers and Mathematics with Applications, 36(10–12):163– 177, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800187.

Belyanin:1997:CCP

[BKK97] A. A. Belyanin, V. V.

Kocharovsky, and Vl. V. Kocharovsky. Cooperative coherent phenomena in annihilating electron-positron and electron-hole plasmas in a strong magnetic field. Computers and Mathematics with Applications, 34(7– 8):845–880, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001776.

Bateson:1999:STE

[BKK99] L. E. Bateson, M. A. Kelmanson, and C. Knudsen. Solution of a transcendental eigenvalue problem via interval analysis. Computers and Mathematics with Applications, 38(7–8):133– 142, October 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002448.

Benczur:1991:GCD

[BKM91] A. Benczúr, A. Kiss, and T. Márkus. On a general class of data dependencies in the relational model and its implication problems. Computers and Mathematics with Applications, 21(1):1–11, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190226T.

Boyer:1995:BMT

[BKM95] R. S. Boyer, M. Kaufmann, and J. S. Moore. The Boyer–Moore theorem prover and its interactive enhancement. Computers and Mathematics with Applications, 29(2):27–62, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194002157.

Babloyantz:1997:ASN

[BKN97] A. Babloyantz, A. P. Krishchenko, and A. Nosov. Analysis and stabilization of nonlinear chaotic systems. Computers and Mathematics with Applications, 34(2–4): 355–368, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001326.

Bales:1994:CFE

[BL94] L. Bales and I. Lasiecka. Continuous finite elements in space and time for the nonhomogeneous wave equation. Computers and Mathematics with Applications, 27 (3):91–102, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900485.

Bartelt:1995:EEL

[BL95] M. Bartelt and W. Li. Error estimates and Lipschitz constants for best approximation in continuous function spaces. Computers and Mathematics with Applications, 30(3– 6):255–268, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001042.

Balla:1996:SCB

[BL96] K. Balla and V. H. Linh.

The simultaneous computation of Bessel functions of first and second kind. Computers and Mathematics with Applications, 31(4–5):87–97, February/March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002200.

Bollobas:1997:EKR

[BL97] B. Bollobás and I. Leader. An Erdős–Ko–Rado theorem for signed sets. Computers and Mathematics with Applications, 34(11):9–13, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002150.

Brimberg:1999:LCR

[BL99] J. Brimberg and R. F. Love. Local convexity results in a generalized Fermat– Weber problem. Computers and Mathematics with Applications, 37(8):87–97, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001029.

Black:1995:PGS

[Bla95] K. Black. A Petrov–Galerkin spectral element technique for heterogeneous porous media flow. Computers and Mathematics with Applications, 29(1):49–65, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400206Z.

Bloom:1991:LTE

[Blo91] Frederick Bloom. Long time existence of solutions for ∂^{2}∂T^{2}u(x, t)+∂∂T α(u(x, t)) = ∂^{2}∂X^{2}β(u(x, t)). Computers and Mathematics with Applications, 21(5):41–56, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190214O.

Bellomo:1995:SNB

[BLP95] N. Bellomo, P. LeTallec, and B. Perthame. The solution of the nonlinear Boltzmann equation: a survey of analytic and computational methods. Computers and Mathematics with Applications, 30(7):21–30, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500122F.

Bevilacqua:1990:SCI

[BLR90] R. Bevilacqua, C. Lotti, and F. Romani. Storage compression of inverses of band matrices. Computers and Mathematics with Applications, 20(8):1–11, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090206Y.

Bernasconi:1994:SPS

[BLR94] A. Bernasconi, M. Leoncini, and G. Resta. Spectral properties of some matrices close to the Toeplitz triangular form. Computers and Mathematics with Applications, 27(6):79–92, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901120.

Bharadwaj:1996:SDL

[BLR96] V. Bharadwaj, H. F. Li, and T. Radhakrishnan. Scheduling divisible loads in bus networks with arbitrary processor release times. Computers and Mathematics with Applications, 32(7):57–77, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001563.

Busenberg:1989:CRD

[BM89] S. N. Busenberg and J. M. Mahaffy. A compartmental reaction–diffusion cell cycle model. Computers and Mathematics with Applications, 18 (10–11):883–892, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122189900060.

See [BM90c].

Bencsik:1990:IEV

[BM90a] I. Bencsik and Gy. Michaletzky. Identification of errorsin-variables models with faurre type realization algorithms. Computers and Mathematics with Applications, 20(12):69–74, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090166H.

Boland:1990:NMM

[BM90b] N. Boland and A. I. Mees.

New methods for multicommodity flows. Computers and Mathematics with Applications, 20(1):29–38, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090066S.

Busenberg:1990:ECR

[BM90c] S. N. Busenberg and J. M.

Mahaffy. Errata to “A compartmental reaction–diffusion cell cycle model”: Computers Math. Applic. 18 (10/ 11), 883–892 (1989). Computers and Mathematics with Applications, 20(7):73–74, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090350S.

See [BM89].

Baart:1991:CQP

[BM91] M. L. Baart and R. J. Y.

McLeod. Continuous quadric patch approximation of curved surfaces. Computers and Mathematics with Applications, 22(8):11–20, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190008R.

Buhmann:1992:MII

[BM92] Martin D. Buhmann and Charles A. Micchelli. Multiquadric interpolation improved. Computers and Mathematics with Applications, 24 (12):21–25, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290168H.

Bainov:1996:EIE

[BM96] D. Bainov and E. Minchev. An estimate of the interval of existence of solutions of the nonlinear Kirchhoff equation. Computers and Mathematics with Applications, 31(2):63–65, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500193X.

Berrut:1997:LCM

[BM97] J.-P. Berrut and H. D. Mittelmann. Lebesgue constant minimizing linear rational interpolation of continuous functions over the interval. Computers and Mathematics with Applications, 33(6):77–86, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000345.

Bocher:1994:NSV

[BMB94] P. Bocher, H. De Meyer, and G. Vanden Berghe. Numerical solution of Volterra equations based on mixed interpolation. Computers and Mathematics with Applications, 27(11):1–11, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900922.

Berghe:1992:PAM

[BMD92] G. Vanden Berghe, H. De Meyer, and M. Van Daele. A parallel approach to the modified Numerov-like eigenvalue determination for Sturm– Liouville problems. Computers and Mathematics with Applications, 23(12):69–74, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290092V.

Bender:1994:DTQ

[BMM94] C. M. Bender, L. R. Mead, and K. A. Milton. Discrete time quantum mechanics. Computers and Mathematics with Applications, 28(10–12):279–317, November/December 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001987.

Babin:1996:IRE

[BMN96] A. Babin, A. Mahalov, and B. Nicolaenko. Integrability and regularity of 3D Euler and equations for uniformly rotating fluids. Computers and Mathematics with Applications, 31(9):35–42, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000405.

Babu:1996:CBE

[BMR96] G. P. Babu, M. N. Murty, and B. R. Ram. Clan-based evolutionary approach for solving control problems. Computers and Mathematics with Applications, 31(6):41–60, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000053.

Beatson:1992:FER

[BN92] R. K. Beatson and G. N. Newsam. Fast evaluation of radial basis functions: I. Computers and Mathematics with Applications, 24 (12):7–19, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290167G.

Benczur:1996:CWK

[BNR96] A. Benczúr, Á. B. Novák, and P. Z. Revesz. Classical and weighted knowledgebase transformations. Computers and Mathematics with Applications, 32 (5):85–98, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600137X.

Bucy:1990:MC

[BNV90] R. S. Bucy, M. K. Namiri, and J. R. Velman. Minimax control. Computers and Mathematics with Applications, 19(4):51–63, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901379.

Bogacki:1996:FPR

[Bog96] P. Bogacki. A family of parallel Runge–Kutta pairs. Computers and Mathematics with Applications, 31(6): 23–31, March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600003X.

Boglaev:1998:FDD

[Bog98] I. P. Boglaev. Finite difference domain decomposition algorithms for a parabolic problem with boundary layers. Computers and Mathematics with Applications, 36 (6):25–40, September 1998.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800159X.

Boglaev:1999:IED

[Bog99] I. Boglaev. An implicitexplicit domain decomposition algorithm for a singularly perturbed parabolic problem. Computers and Mathematics with Applications, 38(5– 6):41–53, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002138.

Bohl:1996:DCD

[Boh96] E. Bohl. Discrete and continuous dynamical systems. Computers and Mathematics with Applications, 31(4– 5):99–104, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002219.

Bohner:1998:DLH

[Boh98] M. Bohner. Discrete linear Hamiltonian eigenvalue problems. Computers and Mathematics with Applications, 36(10–12):179–192, November/December 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800199.

Boley:1992:DON

[Bol92] Harold Boley. Declarative operations on nets. Computers and Mathematics with Applications, 23(6–9):601– 637, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901263.

Bonzani:1997:SNE

[Bon97]  I. Bonzani.  Solution of

nonlinear evolution problems by parallelized collocationinterpolation methods. Computers and Mathematics with Applications, 34(12):71–79, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122197002356.

Bonzani:1999:DDD

[Bon99] I. Bonzani. Domain decomposition and discretization of continuous mathematical models. Computers and Mathematics with Applications, 38(11–12):79– 88, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002874.

Borgers:1990:GDT

[Bör90] C. Börgers. Generalized Delaunay triangulations of nonconvex domains. Computers and Mathematics with Applications, 20(7):45–49, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090347M.

Bouchaffra:1993:TAA

[Bou93] D. Bouchaffra. Theory and algorithms for analysing the consistent region in probabilistic logic. Computers and Mathematics with Applications, 25(3):13–25, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390139M.

Boulbrachene:1998:NQV

[Bou98] M. Boulbrachene. The noncoercive quasi-variational inequalities related to impulse control problems. Computers and Mathematics with Applications, 35(12):101–108, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800100X.

Boumenir:1999:RVN

[Bou99]  A. Boumenir.  A rigor-

ous verification of a numerically computed eigenvalue.

Computers and Mathematics with Applications, 38 (7–8):39–41, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002369.

Boy:1996:ASC

[Boy96] A. Boy. Analysis for a system of coupled reaction-diffusion parabolic equations arising in biology. Computers and Mathematics with Applications, 32(4):15–21, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001204.

Boyd:1999:NCS

[Boy99] J. P. Boyd. A numerical comparison of seven grids for polynomial interpolation on the interval. Computers and Mathematics with Applications, 38(3–4):35–50, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002047.

Baldi:1990:GCB

[BP90] Pierre Baldi and Edward C.

Posner. Graph coloring bounds for cellular radio. Computers and Mathematics with Applications, 19(10): 91–97, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090361M.

Barnett:1994:HAS

[BP94a] M. P. Barnett and K. R. Perry. Hierarchical addressing in symbolic computation. Computers and Mathematics with Applications, 28(8):17–35, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001677.

Burrage:1994:IOC

[BP94b] K. Burrage and B. Pohl. Implementing an ODE code on distributed memory computers. Computers and Mathematics with Applications, 28(10–12):235–252, November/December 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001944.

Bramble:1997:ITT

[BP97] J. H. Bramble and J. E. Pasciak. Iterative techniques for time dependent Stokes problems. Computers and Mathematics with Applications, 33 (1–2):13–30, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002167.

Bukiet:1996:CBN

[BPLS96] B. Bukiet, J. Pelesko, X. L.

Li, and P. L. Sachdev. A characteristic based numerical method with tracking for nonlinear wave equations. Computers and Mathematics with Applications, 31(7): 75–99, April 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000193.

Benczur:1997:SCA

[BQN97] A. Benczúr, C. K. Quang, and T. Nikovits. Some constraints assuring serializability in multidatabases. Computers and Mathematics with Applications, 34(10):9–21, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002034.

Brodzik:1994:CSA

[BR94] M. L. Brodzik and W. C.

Rheinboldt. The computation of simplicial approximations of implicitly defined two-dimensional manifolds. Computers and Mathematics with Applications, 28(9):9–21, November 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001758.

Banerjee:1995:SST

[BR95a] S. Banerjee and S. K. Roychoudhuri. Spherically symmetric thermo-visco-elastic waves in a visco-elastic medium with a spherical cavity. Computers and Mathematics with Applications, 30 (1):91–98, July 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500070F.

Barkai:1995:IHV

[BR95b] S. M. Barkai and O. Rand. Implementation of the “Harmonic Variables” concept in the vibration analysis of a coupled rotor/fuselage system. Computers and Mathematics with Applications, 30 (9):87–104, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500149S.

Bellomo:1995:SNI

[BR95c] N. Bellomo and L. Ridolfi. Solution of nonlinear initialboundary value problems by sinc collocation-interpolation methods. Computers and Mathematics with Applications, 29(4):15–28, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400234C.

Braun:1992:SDM

[Bra92]  Phillip A. Braun.  State

dependent models of stock returns. Computers and Mathematics with Applications, 24(8–9):17–29, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290184J.

Bribiesca:1995:DCG

[BRGJ95] E. Bribiesca, D. A. Rosenblueth, and M. Garza-Jinich. Definite-clause grammars for 2D shape analysis. Computers and Mathematics with Applications, 30 (8):95–103, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500141K.

Bribiesca:1997:MDS

[Bri97] E. Bribiesca. Measuring 2-D shape compactness using the contact perimeter. Computers and Mathematics with Applications, 33(11):1–9, June 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000825.

Brown:1990:CRP

[Bro90] C. S. Brown. Capacity of the regular polyhedra. Computers and Mathematics with Applications, 20(1):43–56, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090068U.

Brooks:1994:ACP

[Bro94]  M. Brooks.  Approxima-

tion complexity for piecewise monotone functions and real data. Computers and Mathematics with Applications, 27 (8):47–58, April 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900655.

Broyden:1996:NTC

[Bro96] C. G. Broyden. A new taxonomy of conjugate gradient methods. Computers and Mathematics with Applications, 31(4–5):7–17, February/March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002111.

Brodzik:1998:CSA

[Bro98] M. L. Brodzik. The computation of simplicial approximations of implicitly defined p-dimensional manifolds. Computers and Mathematics with Applications, 36 (6):93–113, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001643.

Belingeri:1998:DLG

[BRV98] C. Belingeri, P. E. Ricci, and O. Villo’. A discretization of the Laplace and Gauss–Weierstrass operators. Computers and Mathematics with Applications, 35(8):127–140, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000510.

Brezinski:1997:BIL

[BRZS97] C. Brezinski, M. Redivo-Zaglia, and H. Sadok. Breakdowns in the implementation of the Lánczos method for solving linear systems. Computers and Mathematics with Applications, 33 (1–2):31–44, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002179.

Bogacki:1990:IHO

[BS90] P. Bogacki and L. F. Shampine.

Interpolating high-order Runge– Kutta formulas. Computers and Mathematics with Applications, 20(3):15–24, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090027H.

Betro:1991:STG

[BS91a] Bruno Betrò and Fabio Schoen. A stochastic technique for global optimization. Computers and Mathematics with Applications, 21 (6–7):127–133, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191901673.

Bhasker:1991:CPP

[BS91b] J. Bhasker and Tariq Samad. The clique-partitioning problem. Computers and Mathematics with Applications, 22 (6):1–11, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190001K.

Brink:1992:SCA

[BS92] Chris Brink and Renate A. Schmidt. Subsumption computed algebraically. Computers and Mathematics with Applications, 23(2–5):329– 342, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901469.

Boglaev:1993:SSP

[BS93] Igor P. Boglaev and Vadim V.

Sirotkin. The solution of singularity perturbed huge problems via domain decomposition. Computers and Mathematics with Applications, 25(9):31–42, May 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390130N.

Bryc:1994:SEB

[BS94] W. Bryc and W. Smoleński.

A sensitivity estimate for Boolean functions. Computers and Mathematics with Applications, 28(5):45–51, September 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001391.

Boglaev:1995:IDD

[BS95] I. P. Boglaev and V. V.

Sirotkin. Iterative domain decomposition algorithms for the solution of 2-D eddy current problem. Computers and Mathematics with Applications, 29(1):67–81, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002072.

Bogacki:1996:ERK

[BS96a] P. Bogacki and L. F. Shampine. An efficient Runge–Kutta (4,5) pair. Computers and Mathematics with Applications, 32(6):15–28, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001411.

Boglaev:1996:IDD

[BS96b] I. P. Boglaev and V. V. Sirotkin. Iterative domain decomposition algorithms for the solution of singularly perturbed parabolic problems. Computers and Mathematics with Applications, 31(10): 83–100, May 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000545.

Boglaev:1996:SSP

[BS96c] I. P. Boglaev and V. V. Sirotkin. Solution of a singularly perturbed parabolic problem by domain decomposition with time extrapolation. Computers and Mathematics with Applications, 32(2):75–90, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001058.

Boulbrachene:1996:FEA

[BS96d] M. Boulbrachene and H. Sissaoui. The finite element approximation of variational inequalities related to ergodic control problems. Computers and Mathematics with Applications, 31(4–5):137– 141, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002251.

Barrio:1999:PEC

[BS99a] R. Barrio and J. Sabadell.

Parallel evaluation of Chebyshev and trigonometric series. Computers and Mathematics with Applications, 38(11– 12):99–106, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002898.

Berner:1999:IJA

[BS99b] D. F. Berner and J. A.

Snyman. The influence of joint angle constraints on the optimum design of a planar robot manipulator following a complicated prescribed path. Computers and Mathematics with Applications, 37 (1):111–124, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002466.

Blaquiere:1994:SPU

[BSP94] A. Blaquière and M. Sigal-Pauchard. Schrödinger processes with unbounded or singular potentials, conditional Sanov property. Computers and Mathematics with Applications, 27(9–10):103– 120, May 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901295.

Bai:1998:SAM

[BSS98] Zhong-Zhi Bai, Jia-Chang Sun, and D. B. Szyld. The spectral analysis of the matrix multisplitting method for the one-dimensional model problem. Computers and Mathematics with Applications, 36(2):31–38, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800114X.

See corrigendum [Ano00].

Bai:1996:UFC

[BSW96] Zhong-Zhi Bai, Jia-Chang Sun, and De-Ren Wang. A unified framework for the construction of various matrix multisplitting iterative methods for large sparse system of linear equations. Computers and Mathematics with Applications, 32 (12):51–76, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002076.

Battiti:1994:SAT

[BT94] R. Battiti and G. Tecchiolli. Simulated annealing and tabu search in the long run: a comparison on QAP tasks. Computers and Mathematics with Applications, 28(6):1–8, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001472.

Bai:1995:PCP

[BT95] S. X. Bai and Y.-K. Tsai. A production control problem in competition. Computers and Mathematics with Applications, 29(5):65–80, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400249K.

Brutman:1997:EPE

[BT97] L. Brutman and D. Toledano.

An extremal problem of Erdős in interpolation theory. Computers and Mathematics with Applications, 34 (12):37–47, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002320.

Brugnano:1998:BVM

[BT98] L. Brugnano and D. Trigiante. Boundary value methods: the third way between linear multistep and Runge– Kutta methods. Computers and Mathematics with Applications, 36(10–12):269–284, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219880028X.

Bai:1997:PSP

[BTED97] S. X. Bai, Y.-K. Tsai, M. El-Hafsi, and K. Deng. Production scheduling in a price competition. Computers and Mathematics with Applications, 33(5):5–19, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000151.

Bokor:1990:FBM

[BTT90] J. Bokor, M. Tanyi, and Gy. Terdik. Forward and backward Markovian state space models of second order process. Computers and Mathematics with Applications, 19(1):21–29, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090078X.

Budinsky:1990:EHW

[BU90] N. Budinsky and A. A. Ungar. Electromagnetic head waves. Computers and Mathematics with Applications, 19(5):41–53, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090100X.

Burch:1992:VAP

[Bur92] Robert W. Burch. Valental aspects of Peircean algebraic logic. Computers and Mathematics with Applications, 23 (6–9):665–677, March/May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901285.

Butcher:1994:IVP

[But94] J. C. Butcher. Initial value problems: numerical methods and mathematics. Computers and Mathematics with Applications, 28(10–12):1–16, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001820.

Butcher:1996:GLM

[But96] J. C. Butcher. General linear methods. Computers and Mathematics with Applications, 31(4–5):105–112, February/March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002227.

Balakrishnarajan:1995:AIA

[BV95] M. M. Balakrishnarajan and P. Venuvanalingam. An artificial intelligence approach for the generation and enumeration of perfect matchings on graphs. Computers and Mathematics with Applications, 29(1):115–121, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002113.

Balakrishnarajan:1999:STA

[BV99] M. M. Balakrishnarajan and P. Venuvanalingam. A semantic tree algorithm for the generation of sextet polynomials of hexagonal systems. Computers and Mathematics with Applications, 37 (3):97–104, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000504.

Beckermann:1992:SDT

[BW92] Bernhard Beckermann and Jet Wimp. Some dynamically trivial mappings with applications to the improvement of simple iteration.

Computers and Mathematics with Applications, 24 (10):89–97, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900219.

Bucy:1993:MCR

[BW93] R. S. Bucy and B. G.

Williams. A matrix cross ratio theorem for the Riccati equation. Computers and Mathematics with Applications, 26(4):9–20, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390030Y.

Bai:1996:CPN

[BW96a] Zhong-Zhi Bai and De-Ren Wang. A class of parallel nonlinear multisplitting relaxation methods for the large sparse nonlinear complementarity problems. Computers and Mathematics with Applications, 32(8):79–95, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001691.

Bai:1996:ICT

[BW96b] Zhongzhi Bai and Deren Wang. Improved comparison theorem for the nonlinear multisplitting relaxation method. Computers and Mathematics with Applications, 31(8):23–30, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000272.

Barrlund:1997:IMT

[BWK97] A. Barrlund, H. Wallin, and J. Karlsson. Iteration of Möbius transformations and attractors on the real line. Computers and Mathematics with Applications, 33 (1–2):1–12, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002155.

Bykov:1997:OMA

[Byk97] V. V. Bykov. One mechanism of appearance of a spiral quasiattractor involving heteroclinic contours. Computers and Mathematics with Applications, 34(2–4): 229–244, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001259.

Cabos:1997:EMF

[Cab97] C. Cabos. Evaluation of matrix functions with the block Lanczos algorithm. Computers and Mathematics with Applications, 33 (1–2):45–57, January 1997.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002180.

Carroll:1992:SCU

[Car92] John Carroll. Sufficient conditions for uniformly secondorder convergent schemes for stiff initial-value problems. Computers and Mathematics with Applications, 24(10): 105–116, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290023B.

Carroll:1993:MEF

[Car93]  J. Carroll.  A matricial

exponentially fitted scheme for the numerical solution of stiff initial-value problems. Computers and Mathematics with Applications, 26(4):57–64, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390034S.

Carstensen:1994:NSR

[Car94]  C. Carstensen.  A note

on simultaneous rootfinding for algebraic, exponential, and trigonometric polynomials. Computers and Mathematics with Applications, 27(7):7–14, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901449.

Castillo:1991:ADV

[Cas91] JoséE. Castillo. An adaptive direct variational grid generation method. Computers and Mathematics with Applications, 21(5):57–64, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190215P.

Cash:1994:SCN

[Cas94] J. R. Cash. Stability concepts in the numerical solution of difference and differential equations. Computers and Mathematics with Applications, 28(1–3):45–53, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194000921.

Cash:1998:CSC

[Cas98] J. R. Cash. A comparison of some codes for the stiff oscillatory problem. Computers and Mathematics with Applications, 36(1):51–57, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001084.

Chawla:1999:ETF

[CAZ99] M. M. Chawla and M. A. Al-Zanaidi. An extended trapezoidal formula for the diffusion equation. Computers and Mathematics with Applications, 38(2):51–59, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001820.

Chawla:1995:CSE

[CAZAS95] M. M. Chawla, M. A. Al-Zanaidi, and M. S. Al-Sahhar. A class of stabilized extended one-step methods for the numerical solution of ODEs. Computers and Mathematics with Applications, 29(10):79–84, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000484.

Chang:1993:ACH

[CB93a] C. C. Chang and D. J.

Buehrer. Access control in a hierarchy using a oneway trap door function. Computers and Mathematics with Applications, 26 (5):71–76, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900757.

Cree:1993:ANE

[CB93b] M. J. Cree and P. J. Bones. Algorithms to numerically evaluate the Hankel transform. Computers and Mathematics with Applications, 26 (1):1–12, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193900816.

Chen:1992:MME

[CC92] Ye-Sho Chen and Pete Chong. Mathematical modeling of empirical laws in computer applications: a case study. Computers and Mathematics with Applications, 24(7): 77–87, October 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290156C.

Casulli:1994:SAE

[CC94a] V. Casulli and E. Cattani. Stability, accuracy and efficiency of a semiimplicit method for threedimensional shallow water flow. Computers and Mathematics with Applications, 27 (4):99–112, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900590.

Chang:1994:ASO

[CC94b] Y. F. Chang and G. Corliss.

ATOMFT: solving ODEs and DAEs using Taylor series. Computers and Mathematics with Applications, 28(10–12):209–233, November/December 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001936.

Chiou:1994:LFC

[CC94c] J. S. Chiou and T. B. Chang.

Laminar film condensation on a horizontal disk with suction at the wall. Computers and Mathematics with Applications, 27(12):61–68, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900868.

Chung:1995:SSD

[CC95] K. W. Chung and H. S. Y.

Chan. Spherical symmetries from dynamics. Computers and Mathematics with Applications, 29(7):67–81, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500019U.

Csallner:1996:CSI

[CC96] A. E. Csallner and T. Csendes.

The convergence speed of interval methods for global optimization. Computers and Mathematics with Applications, 31(4–5):173–178, February/March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002294.

Campo:1997:ASD

[CC97] A. Campo and L. Campo. An algebraic solution for a 2-d partial differential energy equation with a Robin boundary condition generated from the simpler solution for a Dirichlet boundary condition. Computers and Mathematics with Applications, 34(10): 101–114, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002095.

Chien:1998:MBR

[CC98a] C. S. Chien and M.-H. Chen. Multiple bifurcations in a reaction-diffusion problem. Computers and Mathematics with Applications, 35(8): 15–39, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800042X.

Choo:1998:CND

[CC98b] S. M. Choo and S. K. Chung.

Conservative nonlinear difference scheme for the Cahn– Hilliard equation. Computers and Mathematics with Applications, 36(7):31–39, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001710.

Cui:1998:EDT

[CC98c] Jingan Cui and Lansun Chen.

The effect of diffusion on the time varying logistic population growth. Computers and Mathematics with Applications, 36(3):1–9, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001242.

Cui:1999:EHF

[CC99] Jingan Cui and Lansun Chen.

The effects of habitat fragmentation and ecological invasion on population sizes. Computers and Mathematics with Applications, 38 (1):1–11, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001625.

Cochran:1998:CMS

[CCC^{+}98] H. D. Cochran, P. T. Cummings, S. T. Cui, S. A. Gupta, R. K. Bhupathiraju, and P. F. LoCascio.

Classical molecular simulations of complex, industriallyimportant systems on the Intel Paragon. Computers and Mathematics with Applications, 35(7):73–84, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000340.

Chen:1994:PCA

[CCCY94] M.-J. Chen, T.-Z. Chen, C.- S. Chien, and M.-S. Yu. A parallel continuation algorithm for nonlinear eigenvalue problems. Computers and Mathematics with Applications, 27(7):83–97, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901511.

Chattopadhyay:1999:CPH

[CCD99] P. Chattopadhyay, P. P. Choudhury, and K. Dihidar. Characterisation of a particular hybrid transformation of two-dimensional cellular automata. Computers and Mathematics with Applications, 38(5–6):207– 216, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002278.

Cairo:1999:CRC

[CCGL99] L. Cairó, J. Chavarriga, J. Giné, and J. Llibre. A class of reversible cubic systems with an isochronous center. Computers and Mathematics with Applications, 38(11–12):39–53, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002837.

Carrascosa:1990:STR

[CCH^{+}90] J. L. Carrascosa, J. M.

Carazo, L. Herranz, L. E. Donate, and J. P. Secilla. Study of two related configurations of the neck of bacteriophage o29. Computers and Mathematics with Applications, 20(4–6):57–65, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090314A.

Chen:1997:HAG

[CCH97] Deng-Jyi Chen, Ruey-Shun Chen, and Tien-Hsiang Huang. A heuristic approach to generating file spanning trees for reliability analysis of distributed computing systems. Computers and Mathematics with Applications, 34(10): 115–131, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002101.

Chang:1997:UIL

[CCK97] Chia-Ming Chang, Chien-Ming Chen, and Chung-Ta King. Using integer linear programming for instruction scheduling and register allocation in multi-issue processors. Computers and Mathematics with Applications, 34(9):1–14, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001843.

Cherruault:1990:SAB

[CCVV90] Y. Cherruault, M. Choubane, J. M. Valleton, and J. C. Vincent. Stability and asymptotic behavior of a numerical solution corresponding to a diffusion-reaction equation solved by a finite difference scheme (crank–nicolson). Computers and Mathematics with Applications, 20(11): 37–46, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190902178.

Chung:1996:HSD

[CCW96] K. W. Chung, H. S. Y. Chan, and B. N. Wang. Hyperbolic symmetries from dynamics. Computers and Mathematics with Applications, 31(2):33–47, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001913.

Chien:1993:AMM

[CCY93] Been-Chian Chien, Rong-Jaye Chen, and Wei-Pang Yang. Analysis of the multigroup method for manipulating multiple stacks. Computers and Mathematics with Applications, 25(3): 43–57, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390142I.

Chen:1995:NHA

[CCY95a] Ruey-Shun Chen, Deng-Jyi Chen, and Y. S. Yeh. A new heuristic approach for reliability optimization of distributed computing systems subject to capacity constraints. Computers and Mathematics with Applications, 29(3):37–47, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400227C.

Chen:1995:ROD

[CCY95b] Ruey-Shun Chen, Deng-Jyi Chen, and Y. S. Yeh. Reliability optimization of distributed computing systems subject to capacity constraints. Computers and Mathematics with Applications, 29(4):93–99, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400242D.

Cheng:1998:CSM

[CD98] T. C. E. Cheng and Q. Ding. The complexity of single machine scheduling with two distinct deadlines and identical decreasing rates of processing times. Computers and Mathematics with Applications, 35(12):95–100, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000996.

Chen:1991:API

[CDH91] Gen-Huey Chen, Dyi-Rong Duh, and Cheng-Chung Hsu. An algorithm paradigm for incomplete hypercubes. Computers and Mathematics with Applications, 22(6): 93–96, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190005O.

Chou:1995:PSM

[CDLW95] C.-C. Chou, Y.-F. Deng, G. Li, and Y. Wang. Parallelizing Strassen’s method for matrix multiplication on distributed-memory MIMD architectures. Computers and Mathematics with Applications, 30(2):49–69, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500077C.

Chen:1994:ORS

[CE94] S.-Z. Chen and L. Erbe. Oscillation results for second order scalar and matrix difference equations. Computers and Mathematics with Applications, 28(1–3):55–69, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400093X.

Carlson:1991:PMI

[CF91a] Ralph E. Carlson and Thomas A.

Foley. The parameter R^{2} in multiquadric interpolation. Computers and Mathematics with Applications, 21(9): 29–42, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190123L.

Coolsaet:1991:CSR

[CF91b] K. Coolsaet and V. Fack.

Classifying strongly regular graphs using lexical ordering of adjacency matrices. Computers and Mathematics with Applications, 21(2–3):15–21, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190076G.

Coolsaet:1991:TEA

[CF91c] K. Coolsaet and V. Fack. Total exchange algorithms on ‘sandwich graphs’. Computers and Mathematics with Applications, 22(10):45–48, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191901916.

Carlson:1992:ITD

[CF92] R. E. Carlson and T. A. Foley. Interpolation of track data with radial basis methods. Computers and Mathematics with Applications, 24 (12):27–34, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290169I.

Coolsaet:1994:ODE

[CF94] K. Coolsaet and V. Fack. Optimal data exchange algorithms on star graphs. Computers and Mathematics with Applications, 27(3): 21–25, February 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900426.

Chen:1998:SCP

[CF98] Y.-H. Chen and S.-C. Fang.

Solving convex programming problems with equality constraints by neural networks. Computers and Mathematics with Applications, 36(7): 41–68, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001722.

Chang:1995:CAC

[CFLC95] C. C. Chang, S. W. Fan, H. T. Liaw, and M. Y. Chiou. Cryptanalysis on an access control in a hierarchy. Computers and Mathematics with Applications, 29(4):69–72, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400239H.

Calio:1991:NDA

[CFM91] Franca Calio’, Marco Frontini, and Gradimir V. Milovanović. Numerical differentiation of analytic functions using quadratures on the semicircle. Computers and Mathematics with Applications, 22(10):99–106, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190196B.

Cooke:1994:NAS

[CG94] K. L. Cooke and I. Györi. Numerical approximation of the solutions of delay differential equations on an infinite interval using piecewise constant arguments. Computers and Mathematics with Applications, 28(1–3):81–92, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194000956.

Chou:1995:CSP

[CG95] S. C. Chou and X. S. Gao. The computer searches for Pascal conics. Computers and Mathematics with Applications, 29(2):63–71, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194002168.

Cai:1992:CAC

[CGH92] Wei Cai, David Gottlieb, and Ami Harten. Cell averaging Chebyshev methods for hyperbolic problems. Computers and Mathematics with Applications, 24(5– 6):37–49, September 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290039K.

Cercone:1992:EF

[CGHS92] Nick Cercone, Randy Goebel, John De Haan, and Stephanie Schaeffer. The ECO family. Computers and Mathematics with Applications, 23 (2–5):95–131, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901388.

Coll:1999:DCD

[CGS99] C. Coll, J. Gelonch, and E. Sanchez. Doubly coprime decompositions of a collection of rational matrices. Computers and Mathematics with Applications, 38(11– 12):89–97, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002886.

Chen:1991:HTT

[CH91] Cho-Kuang Chen and Tsan-Hui Hsu. Heat transfer of a thermomicropolar fluid past a porous stretching sheet. Computers and Mathematics with Applications, 21(8): 37–45, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190049A.

Chang:1993:USC

[CH93] C.-C. Chang and S.-J. Hwang. Using smart cards to authenticate remote passwords. Computers and Mathematics with Applications, 26(7): 19–27, October 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390048Z.

Cui:1999:CPC

[CH99] Yingchuan Cui and Bingsheng He. A class of projection and contraction methods for asymmetric linear variational inequalities and their relations to Fukushima’s descent method. Computers and Mathematics with Applications, 38(11–12):151– 164, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002941.

Chambers:1991:FDS

[Cha91] Marcus J. Chambers. Forecasting discrete stock and flow data generated by a second order continuous time system. Computers and Mathematics with Applications, 22 (10):107–114, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190197C.

Chambers:1993:NFC

[Cha93] Marcus J. Chambers. A note on forecasting in cointegrated systems. Computers and Mathematics with Applications, 25(2):93–99, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219390226L.

Chang:1994:PAP

[Cha94] Pei Chann Chang. A practical approach for parts scheduling with alternative process plans in automated manufacturing systems. Computers and Mathematics with Applications, 28(9):55–65, November 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001774.

Chalabi:1996:EBP

[Cha96] A. Chalabi. An error bound for the polygonal approximation of conservation laws with source terms. Computers and Mathematics with Applications, 32(8):59–63, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001678.

Chanane:1998:OCN

[Cha98] B. Chanane. Optimal control of nonlinear systems: a recursive approach. Computers and Mathematics with Applications, 35(3):29–33, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002769.

Chang:1999:BBA

[Cha99a] Pei Chann Chang. A branch and bound approach for single machine scheduling with earliness and tardiness penalties. Computers and Mathematics with Applications, 37(10):133–144, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001303.

Chang:1999:MII

[Cha99b]  Shih Sen Chang.  The

Mann and Ishikawa iterative approximation of solutions to variational inclusions with accretive type mappings. Computers and Mathematics with Applications, 37(9):17–24, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001091.

Chen:1995:HAR

[CHCC95] D.-J. Chen, W. C. Hol, R.- S. Chen, and D. T. K. Chen. A heuristic algorithm for the reliability-oriented file assignment in a distributed computing system. Computers and Mathematics with Applications, 29(10):85–104, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000495.

Chen:1990:HVF

[Che90a] Z. Chen. Hybrid variable finite elements for semiconductor devices. Computers and Mathematics with Applications, 19(4):65–73, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090138A.

Cheng:1990:DPA

[Che90b] T. C. E. Cheng. Dynamic programming approach to the single-machine sequencing problem with different due-dates. Computers and Mathematics with Applications, 19(2):1–7, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090001Z.

Cheng:1991:SFE

[Che91] Kwokming J. Cheng. Symbolic finite element analysis using computer algebra: heat transfer in rectangular duct flow. Computers and Mathematics with Applications, 22(12):15–22, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190143R.

Chen:1992:SLS

[Che92] Tsu-Fen Chen. Semidiscrete least squares methods for linear convection–diffusion problem. Computers and Mathematics with Applications, 24 (11):29–44, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290029H.

Chen:1993:HLT

[Che93a] Tzer-Ming Chen. Hybrid Laplace transform/weighting function scheme for transient multidimensional conservation equations. Computers and Mathematics with Applications, 26(12):25–34, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0898122193900562.

Chen:1993:PFE

[Che93b]  Zhangxin Chen.  Projec-

tion finite element methods for semiconductor device equations. Computers and Mathematics with Applications, 25(8):81–88, April 1993. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390173S.

Chen:1996:FEM

[Che96] Z. Chen. A finite element method for the quantum hydrodynamic model for semiconductor devices. Computers and Mathematics with Applications, 31(7):17–26, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000156.

Chen:1997:SDR

[Che97] Yixiang Chen. Stone duality and representation of stable domain. Computers and Mathematics with Applications, 34(1):27–41, July 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000965.

Chen:1999:OCS

[Che99a]  L. Chen.  Optimal com-

putation of shortest paths on doubly convex bipartite graphs. Computers and Mathematics with Applications, 38(3–4):1–12, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002011.

Chen:1999:EDS

[Che99b] Shaozhu Chen. Eventual disconjugacy of self-adjoint difference equations. Computers and Mathematics with Applications, 38(5–6):117– 128, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002199.

Chen:1999:FRQ

[Che99c] Yun-Shiow Chen. Fuzzy ranking and quadratic fuzzy regression. Computers and Mathematics with Applications, 38(11–12):265–279, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199003053.

Chichinadze:1991:SNN

[Chi91] V. K. Chichinadze. Solution of nonlinear nonconvex optimization problems by Ψ-transformation method. Computers and Mathematics with Applications, 21(6– 7):7–15, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190156X.

Chiam:1993:MBL

[Chi93] T. C. Chiam. Magnetohydrodynamic boundary layer flow due to a continuously moving flat plate. Computers and Mathematics with Applications, 26(4):1–7, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390029U.

Chou:1992:HBR

[Cho92] Mo-Hong Chou. Hopf bifurcation in a reaction-diffusion system governed by the fitz-hugh-nagumo equations. Computers and Mathematics with Applications, 23(10): 85–97, May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290059Q.

Choo:1994:SHO

[Cho94] J. Y. Choo. Stable high order methods for elliptic equations with large first order terms. Computers and Mathematics with Applications, 27(1):65–80, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219490006X.

Chou:1995:CAE

[Cho95] Mo-Hong Chou. Computeraided experiments on the Hopf bifurcation of the FitzHugh–Nagumo nerve model. Computers and Mathematics with Applications, 29(10): 19–33, May 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500041V.

Choi:1996:CSP

[Cho96] I. Choi. Computation and semiotic practice as compositional process. Computers and Mathematics with Applications, 32(1):17–35, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000843.

Choudhury:1999:DDR

[Cho99] P. P. Choudhury. Dynamic design of a reliable computer network on using information theory. Computers and Mathematics with Applications, 38(1):103–118, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900173X.

Chren:1990:NRN

[Chr90] W. A. Chren, Jr. A new residue number system division algorithm. Computers and Mathematics with Applications, 19(7):13–29, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090190U.

Castillo:1995:SAF

[CHSS95] J. E. Castillo, J. M. Hyman, M. J. Shashkov, and S. Steinberg. The sensitivity and accuracy of fourth order finite-difference schemes on nonuniform grids in one dimension. Computers and Mathematics with Applications, 30(8):41–55, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500136M.

Chung:1992:SHP

[Chu92]  Kuo-Liang Chung.  Solv-

ing hyperbolic PDE’s on hypercubes. Computers and Mathematics with Applications, 23(1):87–88, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290083T.

Chung:1999:ACB

[Chu99] Yun-Kung Chung. Application of a cascade BAM neural expert system to conceptual design for facility layout. Computers and Mathematics with Applications, 37(1):95–110, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198002454.

Chen:1993:ESS

[CHW93] Y.-Y. Chen, T.-W. Hwang, and C.-A. Wang. Existence of similarity solutions for surface-tension driven flows in floating rectangular cavities. Computers and Mathematics with Applications, 26 (1):35–52, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193900849.

Cowgill:1999:GAA

[CHW99] M. C. Cowgill, R. J. Harvey, and L. T. Watson. A genetic algorithm approach to cluster analysis. Computers and Mathematics with Applications, 37(7):99–108, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000905.

Cercignani:1998:P

[CI98] C. Cercignani and R. Illner. Preface. Computers and Mathematics with Applications, 35(1–2):xiii, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002538.

Ciesielski:1995:FFS

[Cie95] Z. Ciesielski. Fractal functions and Schauder bases. Computers and Mathematics with Applications, 30(3– 6):283–291, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001077.

Cimikowski:1990:CCP

[Cim90] R. J. Cimikowski. Coloring certain proximity graphs. Computers and Mathematics with Applications, 20(3): 69–82, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090032F.

Chen:1998:CSA

[CIS98] Ming-Po Chen, H. Irmak, and H. M. Srivastava. A certain subclass of analytic functions involving operators of fractional calculus. Computers and Mathematics with Applications, 35(5):83–91, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000078.

Chandrika:1990:SIA

[CJ90] B. Chandrika and R. R.

Joshi. Stochastic inventory analysis of antibody production. Computers and Mathematics with Applications, 20(8):37–49, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902082.

Chawla:1990:AEN

[CJS90] M. M. Chawla, M. K. Jain, and R. Subramanian. The application of explicit Nyström methods to singular second order differential equations. Computers and Mathematics with Applications, 19(12): 47–51, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090250N.

Choudury:1991:CSN

[CK91] Gilbert Choudury and Philip Korman. Computation of solutions of nonlinear boundary value problems. Computers and Mathematics with Applications, 22(8):49–55, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190012S.

Countryman:1994:CNB

[CK94a] M. Countryman and R. Kannan. A class of nonlinear boundary value problems on the half line. Computers and Mathematics with Applications, 28(4):121–130, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001332.

Countryman:1994:NBV

[CK94b] M. Countryman and R. Kannan. Nonlinear boundary value problems on semiinfinite intervals. Computers and Mathematics with Applications, 28(10–12):59–75, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001863.

Chen:1997:EDP

[CK97a] Huey-Ling Chen and Chung-Ta King. Efficient dynamic processor allocation for kary n-cube massively parallel processors. Computers and Mathematics with Applications, 33(8):59–73, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000540.

Choi:1997:DTG

[CK97b] B. D. Choi and J. W. Kim.

Discrete-time Geo_{1}, Geo_{2} /G/1 retrial queueing systems with two types of calls. Computers and Mathematics with Applications, 33(10): 79–88, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000783.

Chawla:1996:EDS

[CKAS96] M. M. Chawla, A. A. Karaballi, and M. S. Al-Sahhar. Extended double-stride Lstable methods for the numerical solution of ODEs. Computers and Mathematics with Applications, 31(2): 1–6, January 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001875.

Cho:1997:ESW

[CKK97a] C.-K. Cho, Y. Kwon, and S. Kang. Estimation of soil water distribution: Identifiability and observation design. Computers and Mathematics with Applications, 34 (12):105–120, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002381.

Cho:1997:PEI

[CKK97b] C.-K. Cho, Y.-H. Kwon, and S. Kang. Parameter estimation for an infiltration problem. Computers and Mathematics with Applications, 33(4):53–67, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000072.

Chai:1991:TTV

[CKL91] Wen Jian Chai, Gu Keqin, and Nan K. Loh. Transforming time-varying multivariable systems into block companion canonical forms. Computers and Mathematics with Applications, 21(2– 3):57–67, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190081E.

Choi:1998:MMC

[CKL98] B. D. Choi, Y. C. Kim, and Y. W. Lee. The M/M/c retrial queue with geometric loss and feedback. Computers and Mathematics with Applications, 36 (6):41–52, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001606.

Chien:1999:BKE

[CKM99] C. S. Chien, Y. J. Kuo, and Z. Mei. Bifurcations of the von Kármán equations with Robin boundary conditions. Computers and Mathematics with Applications, 38 (3–4):85–112, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002084.

Crisci:1998:SDV

[CKRV98] M. R. Crisci, V. B. Kolmanovskii, E. Russo, and A. Vecchio. Stability of difference Volterra equations: Direct Liapunov method and numerical procedure. Computers and Mathematics with Applications, 36(10–12): 77–97, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800114.

Codenotti:1990:PFS

[CL90] B. Codenotti and M. Leoncini. Parallelism and fast solution of linear systems. Computers and Mathematics with Applications, 19(10):1–18, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090354M.

Chiang:1991:FPNa

[CL91a] Jen-Shiun Chiang and Mi Lu. Floating-point numbers in residue number systems. Computers and Mathematics with Applications, 22(10): 127–140, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Chiang:1991:FPNb

[CL91b] Jen-Shiun Chiang and Mi Lu.

Floating-point numbers in residue number systems. Computers and Mathematics with Applications, 22 (10):127–140, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190200N.

Chang:1994:FLA

[CL94a] P.-T. Chang and E. S. Lee.

Fuzzy least absolute deviations regression and the conflicting trends in fuzzy parameters. Computers and Mathematics with Applications, 28 (5):89–101, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400143X.

Chang:1994:FLR

[CL94b] P.-T. Chang and E. S. Lee.

Fuzzy linear regression with spreads unrestricted in sign. Computers and Mathematics with Applications, 28(4): 61–70, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001278.

Chang:1994:RFS

[CL94c] P.-T. Chang and E. S. Lee.

Ranking of fuzzy sets based on the concept of existence. Computers and Mathematics with Applications, 27(9– 10):1–21, May 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219490118X.

Chang:1995:ENF

[CL95] P.-T. Chang and E. S. Lee. The estimation of normalized fuzzy weights. Computers and Mathematics with Applications, 29(5):21–42, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400246H.

Chen:1996:ABS

[CL96a] Ming-Po Chen and Bing Liu. Asymptotic behavior of solutions of first order nonlinear delay difference equations. Computers and Mathematics with Applications, 32 (4):9–13, August 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001198.

Chou:1996:EDB

[CL96b] Mo-Hong Chou and Yu-Tuan Lin. Exotic dynamic behavior of the forced FitzHugh–Nagumo equations. Computers and Mathematics with Applications, 32(10):

109–124, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001897.

Cheng:1998:GFS

[CL98] Sui Sun Cheng and Guo Hua Lin. Green’s function and stability of a linear partial difference scheme. Computers and Mathematics with Applications, 35(5):27–41, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000030.

Chang:1999:FDN

[CL99a] P.-T. Chang and E. S.

Lee. Fuzzy decision networks and deconvolution. Computers and Mathematics with Applications, 37 (11–12):53–63, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001431.

Cheng:1999:AFA

[CL99b] C.-B. Cheng and E. S. Lee.

Applying fuzzy adaptive network to fuzzy regression analysis. Computers and Mathematics with Applications, 38(2):123–140, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900187X.

Cheng:1999:NFR

[CL99c] C.-B. Cheng and E. S. Lee. Nonparametric fuzzy regression- k-NN and kernel smoothing techniques. Computers and Mathematics with Applications, 38 (3–4):239–251, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001984.

Cheng:1999:GST

[CL99d] Sui Sun Cheng and Yi-Feng Lu. General solutions of a three-level partial difference equation. Computers and Mathematics with Applications, 38(7–8):65–79, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002394.

Clark:1998:UBS

[Cla98]  D. A. Clark.  An up-

per bound of Σ1(a_{i} log a_{i}) for quasi-primitive sequences. Computers and Mathematics with Applications, 35 (4):105–109, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002939.

Cheung:1997:CGA

[CLD97] B. K.-S. Cheung, A. Langevin, and H. Delmaire. Coupling genetic algorithm with a grid search method to solve mixed integer nonlinear programming problems. Computers and Mathematics with Applications, 34 (12):13–23, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002290.

Chiccoli:1990:RRG

[CLM90] C. Chiccoli, S. Lorenzutta, and G. Maino. Recent results for generalized exponential integrals. Computers and Mathematics with Applications, 19(5):21–29, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900985.

Chiccoli:1992:CSI

[CLM92] C. Chiccoli, S. Lorenzutta, and G. Maino. Concerning some integrals of the generalized exponential-integral function. Computers and Mathematics with Applications, 23(11):13–21, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290065P.

Chien:1999:CGT

[CLM99] C.-S. Chien, W.-W. Lin, and Z. Mei. Conjugate gradient type methods for semilinear elliptic problems with symmetry. Computers and Mathematics with Applications, 37(8):3–22, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000978.

Camouzis:1994:RRS

[CLRN94] E. Camouzis, G. Ladas, I. W. Rodrigues, and S. Northshield. The rational recursive sequence x_{n+1} = (βx_{n}^{2})□(1 + x_{n−1}^{2}). Computers and Mathematics with Applications, 28 (1–3):37–43, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194000913.

Chung:1991:TPP

[CLY91] Kuo-Liang Chung, Ferng-Ching Lin, and Yeong-Nan Yeh. A three-phase parallel algorithm for solving linear recurrences. Computers and Mathematics with Applications, 21(2–3):187–195, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190098O.

Chen:1995:OND

[CLY95] M.-P. Chen, B. S. Lalli, and J. S. Yu. Oscillation in neutral delay difference equations with variable coefficients. Computers and Mathematics with Applications, 29(3):5–11, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002238.

Cubiotti:1997:GQV

[CLY97] P. Cubiotti, K. L. Lin, and J. C. Yao. Generalized quasivariational inequalities for fuzzy mappings. Computers and Mathematics with Applications, 33(7):121–134, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000473.

Chattopadhyay:1993:MOA

[CM93] Aditi Chattopadhyay and Thomas R. McCarthy. A multidisciplinary optimization approach for vibration reduction in helicopter rotor blades. Computers and Mathematics with Applications, 25(2):59–72, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390223I.

Castillo:1992:CAN

[CMMR92] JoséE. Castillo, Greg McDermott, Marc McEachern, and John Richardson. A comparative analysis of numerical techniques applied to a model of the otolith membrane. Computers and Mathematics with Applications, 24 (7):133–141, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290162B.

Campiti:1995:UCL

[CMP95] M. Campiti, G. Metafune, and D. Pallara. Uniformly convergent Lagrangetype approximation. Computers and Mathematics with Applications, 30(3–6):269– 276, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001050.

Calvo:1990:NEP

[CMR90] M. Calvo, J. I. Montijano, and L. Randez. A new embedded pair of Runge–Kutta formulas of orders 5 and 6. Computers and Mathematics with Applications, 20(1):15–24, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090064Q.

Carlson:1994:SAM

[CN94] R. E. Carlson and B. K.

Natarajan. Sparse approximate multiquadric interpolation. Computers and Mathematics with Applications, 27(6):99–108, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901147.

Castillo:1999:PGD

[CO99] J. E. Castillo and J. S. Otto.

A practical guide to direct optimization for planar gridgeneration. Computers and Mathematics with Applications, 37(9):123–156, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001182.

See erratum [Ano99-246].

Cohn:1992:CSH

[Coh92] Anthony G. Cohn. Completing sort hierarchies. Computers and Mathematics with Applications, 23(6–9):477– 491, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901193.

Cohen:1994:ACM

[Coh94] E. B. Cohen. Analysis of a class of multi-stage, multistep Runge–Kutta methods. Computers and Mathematics with Applications, 27 (3):103–116, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900493.

Coma:1992:ASS

[Com92] Josep RifàI Coma. All ssum sets of type 2 or 3 are triple-sum sets. Computers and Mathematics with Applications, 23(1):35–43, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290078V.

Conway:1990:SR

[Con90] G. Conway. Software review. Computers and Mathematics with Applications, 20 (4–6):341–343, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090342H.

Condon:1995:DIP

[Con95] D. J. Condon. Day’s implementation of the Pruess method for Sturm–Liouville eigenvalues. Computers and Mathematics with Applications, 29(4):9–14, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400233B.

Cong:1998:RTP

[Con98] Nguyen Huu Cong. RKNtype parallel block PC methods with Lagrange-type predictors. Computers and Mathematics with Applications, 35(9):45–57, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800056X.

Cooke:1997:HMA

[Coo97] C. H. Cooke. The Hadamard matroid and an anomaly in its single element extensions. Computers and Mathematics with Applications, 33(7):115–120, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000461.

Cooke:1998:AAM

[Coo98]  C. H. Cooke.  The ad-

joint alternative for matrix operators. Computers and Mathematics with Applications, 35(5):79–81, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000066.

Corless:1994:WGN

[Cor94] R. M. Corless. What good are numerical simulations of chaotic dynamical systems? Computers and Mathematics with Applications, 28(10–12):107–121, November/December 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400188X.

Corso:1995:RPS

[Cor95] G. M. Del Corso. Randomization and the parallel solution of linear algebra problems. Computers and Mathematics with Applications, 30 (11):59–72, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500164T.

Chalmers:1995:GMS

[CP95a] B. L. Chalmers and M. P. Prophet. The geometry of minimal shape-preserving projections onto Π_{n}. Computers and Mathematics with Applications, 30(3–6):277– 282, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001069.

Chattopadhyay:1995:MOU

[CP95b] A. Chattopadhyay and N. Pagaldipti. A multidisciplinary optimization using semianalytical sensitivity analysis procedure and multilevel decomposition. Computers and Mathematics with Applications, 29(7):55–66, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500018T.

Cash:1996:MST

[CP96] J. R. Cash and Y. Psihoyios.

The MOL solution of time dependent partial differential equations. Computers and Mathematics with Applications, 31(11):69–78, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000636.

Castillo:1998:ASS

[CP98a] S. Castillo and M. Pinto.

Asymptotic summation for second-order finite difference systems. Computers and Mathematics with Applications, 35(4):117–129, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122197002952.

Chou:1998:NRC

[CP98b] S. H. Chou and T. A. Porsching. A note on the restarted CG method and reduced space additive correction. Computers and Mathematics with Applications, 35(9):129–136, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000649.

Chattopadhyay:1993:DOP

[CPC93] A. Chattopadhyay, N. Pagaldipti, and K. T. Chang. A design optimization procedure for efficient turbine airfoil design. Computers and Mathematics with Applications, 26(4):21–31, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390031P.

Caruso:1998:IMC

[CQ98] C. Caruso and F. Quarta. Interpolation methods comparison. Computers and Mathematics with Applications, 35(12):109–126, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001011.

Choudhuri:1990:SST

[CR90a] S. K. Roy Choudhuri and G. Chatterjee (Roy). Spherically symmetric thermoelastic waves in a temperaturerate dependent medium with a spherical cavity. Computers and Mathematics with Applications, 20(11):1–12, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902134.

Choudhuri:1990:TRD

[CR90b] S. K. Roy Choudhuri and G. Chatterjee Roy. Temperaturerate dependent magnetothermoelastic waves in a finitely conducting elastic half-space. Computers and Mathematics with Applications, 19(5): 85–93, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090104R.

Cutler:1990:CSH

[CR90c] D. R. Cutler and I. L. Reilly.

A comparison of some Hausdorff notions in fuzzy topological spaces. Computers and Mathematics with Applications, 19(11):97–104, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090152A.

Craven:1998:OCI

[Cra98] B. D. Craven. Optimal control and invexity. Computers and Mathematics with Applications, 35(5):17–25, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000029.

Crisci:1998:TPR

[CRV98] M. R. Crisci, E. Russo, and A. Vecchio. Time point relaxation methods for Volterra integro-differential equations. Computers and Mathematics with Applications, 36 (9):59–70, November 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001928.

Craig:1990:ADI

[CS90] I. J. D. Craig and A. D. Sneyd. An alternating direction implicit scheme for parabolic systems of partial differential equations. Computers and Mathematics with Applications, 20(3): 53–62, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090030N.

Chawla:1993:SFD

[CS93a] M. M. Chawla and P. N. Shivakumar. A symmetric finite difference method for computing eigenvalues of Sturm– Liouville problems. Computers and Mathematics with Applications, 26(2):67–77, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390323N.

Choo:1993:SHO

[CS93b] Joshua Y. Choo and David H.

Schultz. Stable high order methods for differential equations with small coefficients for the second order terms. Computers and Mathematics with Applications, 25 (1):105–123, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390216I.

Chen:1994:SFB

[CS94a] M.-P Chen and H. M. Srivastava. Some families of bilinear and bilateral generating functions. Computers and Mathematics with Applications, 28(9):1–7, November 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400174X.

Choo:1994:HOD

[CS94b] J. Y. Choo and D. H.

Schultz. A high order difference method for the steady state Navier–Stokes equations. Computers and Mathematics with Applications, 27(11):105–119, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901015.

Carroll:1996:CSA

[CS96] J. Carroll and S. Stewart. A comparison of some adaptive space mesh solvers for the numerical solution of parabolic partial differential equations. Computers and Mathematics with Applications, 31(6):97–115, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000090.

Csenki:1995:DMM

[Cse95] A. Csenki. A dependability measure for Markov models of repairable systems: Solution by randomization and computational experience. Computers and Mathematics with Applications, 30(2):95–110, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500079E.

Cho:1996:SCF

[CSOL96] Y. Cho, V. I. Shin, M. Oh, and Y. Lee. Suboptimal continuous filtering based on the decomposition of observation vector. Computers and Mathematics with Applications, 32(4):23–31, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001216.

Cendra:1997:FNH

[CST97] H. Cendra, R. Salthu, and A. Torresi. Finding nondegenerate Hopf bifurcation points for four-dimensional twoparametric systems. Computers and Mathematics with Applications, 33(12):115– 124, June 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000928.

Cong:1999:GCE

[CSW99] Nguyen Huu Cong, K. Strehmel, and R. Weiner. A general class of explicit pseudo twostep RKN methods on parallel computers. Computers and Mathematics with Applications, 38(5–6):17–30, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002114.

Chen:1990:QNA

[CT90] Y. M. Chen and L. M.

Tseng. A queuing network approach to the topological optimization of linked cluster networks. Computers and Mathematics with Applications, 20(7):1–10, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090343I.

Corwin:1996:EES

[CT96] S. P. Corwin and S. Thompson. Error estimation and step size control for delay differential equation solvers based on continuously embedded Runge–Kutta–Sarafyan methods. Computers and Mathematics with Applications, 31(6):1–11, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000016.

Chartrand:1997:DD

[CT97] G. Chartrand and S. Tian. Distance in digraphs. Computers and Mathematics with Applications, 34(11):15–23, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122197002162.

Chaudhry:1996:ADT

[CTG96] M. L. Chaudhry, J. G. C. Templeton, and U. C. Gupta. Analysis of the discrete-time GI/Geom(n)/1/N queue. Computers and Mathematics with Applications, 31(1):59–68, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219500182X.

Chung:1995:PSC

[CTY95] K.-L. Chung, Y.-H. Tsai, and W.-M. Yan. A parallel solver for circulant block-tridiagonal systems. Computers and Mathematics with Applications, 29(1):109–113, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400210C.

Chang:1999:ETS

[CTY99] Shih-Sen Chang, H. B.

Thompson, and G. X.-Z. Yuan. The existence theorems of solutions for generalized vector-valued variational-like inequalities. Computers and Mathematics with Applications, 37(7):1–9, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000826.

Cubiotti:1995:DQL

[Cub95a]  P. Cubiotti.  Discontinu-

ous quasivariational-like inequalities. Computers and Mathematics with Applications, 29(12):9–12, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000532.

Cubiotti:1995:ESL

[Cub95b] P. Cubiotti. Existence of solutions for lower semicontinuous quasi-equilibrium problems. Computers and Mathematics with Applications, 30 (12):11–22, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500171T.

Cubiotti:1996:GTC

[Cub96]  P. Cubiotti.  A general

theorem of complementarity for the implicit GQVI problem. Computers and Mathematics with Applications, 31(11):97–104, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000661.

Casper:1990:SSL

[CV90] M. J. Casper and D. A. Voss. L_{0}-stable split linear multistep formulas for parabolic PDEs. Computers and Mathematics with Applications, 19(2):35–46, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190900055.

Chong:1995:CBU

[CW95] L. Chong and G. A. Watson.

Characterization of a best and a unique best approximation from constrained rationals. Computers and Mathematics with Applications, 30 (3–6):51–57, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000852.

Chui:1992:CID

[CWJ92] C. K. Chui, J. D. Ward, and K. Jetter. Cardinal interpolation with differences of tempered functions. Computers and Mathematics with Applications, 24 (12):35–48, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290170M.

Chung:1993:SST

[CY93] Kuo-Liang Chung and Wen-Ming Yan. Solving the symmetric tridiagonal eigenvalue problem on hypercubes. Computers and Mathematics with Applications, 25(9): 91–96, May 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390135I.

Cubiotti:1994:GQV

[CY94] P. Cubiotti and Je. Ch. Yao. The generalized quasivariational inequality problem over non-compact sets. Computers and Mathematics with Applications, 28(4): 93–97, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001308.

Cubiotti:1995:MOG

[CY95] P. Cubiotti and J.-C. Yao. Multivalued (S)_{+}^{1} operators and generalized variational inequalities. Computers and Mathematics with Applications, 29(12):49–56, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000576.

Chung:1996:VAS

[CY96] Kuo-Liang Chung and Wen-Ming Yan. Vectorized algorithms for solving special tridiagonal systems. Computers and Mathematics with Applications, 32 (12):1–14, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002039.

Chen:1994:PNS

[CYS94] Ming-Po Chen, J. S. Yu, and J. H. Shen. The persistence of nonoscillatory solutions of delay differential equations under impulsive perturbations. Computers and Mathematics with Applications, 27(8):1–6, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900612.

Cheng:1994:MSC

[CZ94] Sui Sun Cheng and Bing-Gen Zhang. Monotone solutions of a class of nonlinear difference equations. Computers and Mathematics with Applications, 28(1–3):71–79, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194000948.

Datta:1990:TAS

[Dat90]  B. K. Datta.  A tech-

nique for approximate solutions to Schrod̈inger-like equations. Computers and Mathematics with Applications, 20(1):61–65, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090070Z.

Datta:1993:AEC

[Dat93] Barun Kumar Datta. The approximate evaluation of certain integrals. Computers and Mathematics with Applications, 25(7):47–49, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390311I.

Davies:1996:PIB

[Dav96] A. J. Davies. Parallel implementations of the boundary element method. Computers and Mathematics with Applications, 31(6):33–40, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000041.

Donaill:1996:EAL

[DB96] D. A. Mac Dónaill and N. H. Buttimore. The exploitation of assembly language instructions in biological text manipulation: I. nucleotide sequences. Computers and Mathematics with Applications, 32(11):29–38, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122196001940.

Dauer:1998:NCN

[DBA98] J. P. Dauer, K. Balachandran, and S. M. Anthoni. Null controllability of nonlinear infinite neutral systems with delays in control. Computers and Mathematics with Applications, 36(2):39–50, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001151.

Daele:1992:FDF

[DBM92] M. Van Daele, G. Vanden Berghe, and H. De Meyer. Five-diagonal finite difference methods based on mixed-type interpolation for a certain fourth-order twopoint boundary-value problem. Computers and Mathematics with Applications, 24 (10):55–76, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290019E.

Daele:1994:NIB

[DBM94] M. Van Daele, G. Vanden Berghe, and H. De Meyer. Numerical integration based on mixed-type interpolation for a fourth-order boundaryvalue problem. Computers and Mathematics with Applications, 27(11):55–65, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900981.

Daele:1995:PIR

[DBM95] M. Van Daele, G. Vanden Berghe, and H. De Meyer. Properties and implementation of ρ-Adams methods based on mixed-type interpolation. Computers and Mathematics with Applications, 30 (10):37–54, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500155R.

Dautenhahn:1994:CSP

[DC94] K. Dautenhahn and H. Cruse. Computer simulations of path generation and path form modification with local rules working on a parallel cell-based architecture. Computers and Mathematics with Applications, 28 (5):75–88, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001464.

Dattoli:1993:GTT

[DCL^{+}93] G. Dattoli, C. Chiccoli, S. Lorenzutta, G. Maino, M. Richetta, and A. Torre. Graf-type theorem for Laguerre and Legendre functions. Computers and Mathematics with Applications, 25(12):99–105, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901893.

Dattoli:1994:TGH

[DCL^{+}94] G. Dattoli, C. Chiccoli, S. Lorenzutta, G. Maino, and A. Torre. Theory of generalized Hermite polynomials. Computers and Mathematics with Applications, 28(4):71–83, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001286.

Velho:1997:NAT

[dCVC97] H. F. de Campos Velho and J. C. R. Claeyssen. A nonmodal approach for timeintegration of a barotropic limited area model. Computers and Mathematics with Applications, 33(9):1–13, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219700062X.

Debnath:1999:IEW

[DD99] L. Debnath and S. Das. Interaction of elastic waves by an interfacial Griffith crack. Computers and Mathematics with Applications, 38 (3–4):151–158, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001893.

Dragomir:1999:ILA

[DDP99] S. S. Dragomir, N. M. Dragomir, and K. Pranesh. An inequality for logarithms and applications in information theory. Computers and Mathematics with Applications, 38(2):11–17, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001777.

Dubois:1993:EEP

[DdW93] N. Dubois and D. de Werra. Epcot: an efficient procedure for coloring optimally with Tabu Search. Computers and Mathematics with Applications, 25(10–11):35– 45, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902795.

DelaSen:1996:ROH

[De 96] M. De la Sen. The reachability and observability of hybrid multirate sampling linear systems. Computers and Mathematics with Applications, 31(1):109–122, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001852.

DelaSen:1999:ASF

[De 99] M. De la Sen. Adaptive stabilization of first-order systems using estimates modification based on a Sylvester determinant test. Computers and Mathematics with Applications, 37(10):51–62, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900125X.

Deane:1998:PCD

[Dea98] A. E. Deane. Parallel computations of 3-D Rayleigh– Taylor instability. Computers and Mathematics with Applications, 35(7):111–116, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000388.

Deissenberg:1991:ISG

[Dei91] Christophe Deissenberg. Interactive search for a global optimum in macroeconometric policy-making. Computers and Mathematics with Applications, 21(6–7):203– 213, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901743.

Deissenberg:1993:GFL

[Dei93] Christophe Deissenberg. The global failure of local preference optimization. Computers and Mathematics with Applications, 25(10–11):161– 172, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390290C.

Duenas:1999:MSD

[DELE99] E. Dueñas, R. England, and J. López-Estrada. Multiple shooting with dichotomically stable formulae for linear boundary-value problems. Computers and Mathematics with Applications, 38(9– 10):143–159, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002709.

Demkowicz:1994:ACFa

[Dem94a]  L. Demkowicz.  Asymp-

totic convergence in finite and boundary element methods: Part 1: theoretical results. Computers and Mathematics with Applications, 27 (12):69–84, June 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900876.

Demkowicz:1994:ACFb

[Dem94b] L. Demkowicz. Asymptotic convergence in finite and boundary element methods: part 2: the LBB constant for rigid and elastic scattering problems. Computers and Mathematics with Applications, 28(6):93–109, September 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001553.

Diniz-Ehrhardt:1994:PPM

[DEMS94] M. A. Diniz-Ehrhardt, J. M.

Martínez, and S. A. Santos. Parallel projection methods and the resolution of illposed problems. Computers and Mathematics with Applications, 27(1):11–24, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900027.

Dendrinos:1991:ASQ

[Den91a] Dimitrios S. Dendrinos. Asymptotic stability and quasiperiodicity in form and evolution of model socio-spatial structures. Computers and Mathematics with Applications, 22(6):29–45, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190003M.

Denman:1991:PSI

[Den91b] Eugene D. Denman. Projectors, solvents, interpolating polynomials and lambda matrices. Computers and Mathematics with Applications, 21 (11–12):145–151, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190116L.

Devroye:1993:RML

[Dev93] Luc Devroye. Records, the maximal layer, and uniform distributions in monotone sets. Computers and Mathematics with Applications, 25(5):19–31, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193901952.

Dey:1990:F

[Dey90] S. K. Dey. Foreward. Computers and Mathematics with Applications, 19(8–9):ix, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090259M.

Dixon:1994:PED

[DF94] A. E. Dixon and C. A. J. Fletcher. A preconditioner for equations with dominant convective terms. Computers and Mathematics with Applications, 27(7):59–66, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219490149X.

Ding:1998:GPC

[DF98] H. Q. Ding and R. D. Ferraro. An 18 GFLOPS parallel climate data assimilation PSAS package. Computers and Mathematics with Applications, 35(7):55–63, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000327.

Domingo-Ferrer:1999:RSC

[DFMS99] J. Domingo-Ferrer and J. M.

Mateo-Sanz. Resampling for statistical confidentiality in contingency tables. Computers and Mathematics with Applications, 38(11– 12):13–32, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002813.

DiBenedetto:1993:CGP

[DFS93] Fabio Di Benedetto, Giuseppe Fiorentino, and Stefano Serra. C.G. preconditioning for Toeplitz matrices. Computers and Mathematics with Applications, 25(6):35–45, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902979.

Dean:1991:PCB

[DG91] E. J. Dean and P. Gubernatis. Pointwise control of Burgers’ equation — a numerical approach. Computers and Mathematics with Applications, 22(7):93–100, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191901868.

Dean:1999:DDM

[DG99] E. J. Dean and R. Glowinski. Domain decomposition methods for mixed finite element approximations of wave problems. Computers and Mathematics with Applications, 38 (7–8):207–214, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002515.

Dubonos:1990:SIR

[DGU90] S. N. Dubonos, B. N. Gaifullin, and N. G. Ushakov. Statistical image restoration. Computers and Mathematics with Applications, 19(1): 39–45, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190900804.

Dattoli:1994:ANC

[DGV94] G. Dattoli, L. Giannessi, and G. Voykov. Analytical and numerical computation of undulator brightness: a comparison. Computers and Mathematics with Applications, 27(6):63–78, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901112.

Demetrovics:1991:CST

[DH91] János Demetrovics and Nguyen Xuan Huy. Closed sets and translations of relation schemes. Computers and Mathematics with Applications, 21(1): 13–23, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190227U.

Dutt:1997:LSD

[DH97] S. Dutt and J. P. Hayes. A local-sparing design methodology for fault-tolerant multiprocessorsa. Computers and Mathematics with Applications, 34(11):25–50, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122197002174.

Dai:1998:CSC

[DH98a] B. X. Dai and L. H. Huang.

Convergence of solutions for a class of neutral difference equations. Computers and Mathematics with Applications, 35(10):83–91, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000753.

Dai:1998:ABS

[DH98b] Binxiang Dai and Lihong Huang. Asymptotic behavior of solutions for a class of nonlinear difference equations. Computers and Mathematics with Applications, 36 (2):23–30, July 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001138.

Doual:1991:SST

[DHV91] Nasser Doual, James L. Howland, and Rémi Vaillancourt. Selective solutions to transcendental equations. Computers and Mathematics with Applications, 22(9): 61–76, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190207K.

Dongarra:1998:KCP

[DHW98] J. J. Dongarra, S. Hammarling, and D. W. Walker. Key concepts for parallel out-of-core LU factorization. Computers and Mathematics with Applications, 35(7):

13–31, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000297.

Diamond:1991:IMD

[Dia91] Phil Diamond. Iterated maps on discretized meshes of the unit interval. Computers and Mathematics with Applications, 21(8):57–63, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900515.

Ding:1996:IPE

[Din96] Xie Ping Ding. Iterative process with errors to locally strictly pseudocontractive maps in Banach spaces. Computers and Mathematics with Applications, 32 (10):91–97, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001873.

Ding:1997:GAS

[Din97a]  Xie Ping Ding.  Gen-

eral algorithm of solutions for nonlinear variational inequalities in Banach space. Computers and Mathematics with Applications, 34(9): 131–137, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001946.

Ding:1997:IPE

[Din97b] Xie Ping Ding. Iterative process with errors to nonlinear Φ-strongly accretive operator equations in arbitrary Banach spaces. Computers and Mathematics with Applications, 33(8):75–82, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000552.

Ding:1999:ASM

[Din99a] Xie Ping Ding. Algorithm of solutions for mixed implicit quasi-variational inequalities with fuzzy mappings. Computers and Mathematics with Applications, 38(5–6):231–241, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002291.

Ding:1999:GIQ

[Din99b] Xie Ping Ding. Generalized implicit quasivariational inclusions with fuzzy set-valued mappings. Computers and Mathematics with Applications, 38(1):71–79, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001704.

Ding:1999:GVI

[Din99c] Xie Ping Ding. Generalized variational inequalities and equilibrium problems in generalized convex spaces. Computers and Mathematics with Applications, 38(7– 8):189–197, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002497.

Ding:1999:GVQ

[Din99d] Xie Ping Ding. The generalized vector quasi-variational-like inequalities. Computers and Mathematics with Applications, 37(6):57–67, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000760.

Dorovsky:1997:HNMb

[DIR97] V. N. Dorovsky, Kh. Kh.

Imomnazarov, and E. I. Romensky. A hydrodynamic nonlinear model of bubble liquid (Part II). Computers and Mathematics with Applications, 33(6):13–15, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000278.

Dewson:1997:EVK

[DIT97] T. Dewson, A. D. Irving, and G. Terdik. Estimating of Volterra kernels in case of moment hierarchy and orthogonal representation. Computers and Mathematics with Applications, 33(4):5–14, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000023.

derJeugt:1991:PIN

[dJF91] J. Van der Jeugt and V. Fack. The pragacz identity and a new algorithm for Littlewood–Richardson coefficients. Computers and Mathematics with Applications, 21(2–3):39–47, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190079J.

Dimitrov:1998:FRR

[DJM98] V. S. Dimitrov, G. A. Jullien, and W. C. Miller. A fast and robust RNS algorithm for evaluating signs of determinants. Computers and Mathematics with Applications, 35(8):9–14, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000418.

deJager:1997:FPN

[dJPvZ97] N. G. de Jager, G. Pap, and M. C. A. van Zuijlen. Facts, phantasies, and a new proposal concerning the stringer bound. Computers and Mathematics with Applications, 33(5):37–54, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000187.

Dimov:1996:NAD

[DJV96] I. Dimov, U. Jaekel, and H. Vereecken. A numerical approach for determination of sources in transport equations. Computers and Mathematics with Applications, 32(5):31–42, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001332.

Diamond:1993:SDM

[DK93] Phil Diamond and Peter Kloeden. Spatial discretization of mappings. Computers and Mathematics with Applications, 25(6):85–94, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390302C.

Decorps:1997:ZBE

[DK97a] S. A. M. Decorps and M. A.

Kelmanson. A zonal boundary element method for analysing heat exchangers with thin extended surfaces. Computers and Mathematics with Applications, 33(8): 103–107, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000588.

Diamond:1997:EFL

[DK97b] P. Diamond and R. Körner. Extended fuzzy linear models and least squares estimates. Computers and Mathematics with Applications, 33(9):15–32, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000631.

Dmitrieva:1997:SCS

[DK97c] L. Dmitrieva and M. Khlabystova. Scheme of constructing solitontype solutions of the (2 + 1)- dimensional Harry Dym equation. Computers and Mathematics with Applications, 34(2–4):75–93, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001181.

Deshpande:1998:NDK

[DKG98] S. M. Deshpande, P. S. Kulkarni, and A. K. Ghosh. New developments in kinetic schemes. Computers and Mathematics with Applications, 35(1–2):75–93, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002599.

Dohi:1993:CTI

[DKO93] T. Dohi, N. Kaio, and S. Osaki. Continuous time inventory control for Wiener process demand. Computers and Mathematics with Applications, 26(9):11–22, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390002D.

Dmitrieva:1997:ECD

[DKR97] L. A. Dmitrieva, Yu. A. Kuperin, and G. E. Rudin. Extended class of Dubrovin’s equations related to the onedimensional quantum threebody problem. Computers and Mathematics with Applications, 34(5–6):571– 585, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001545.

deKlerk:1994:FDC

[dKSG94] E. de Klerk, J. A. Snyman, and Structural Optimization Research Group. A feasible descent cone method for linearly constrained minimization problems. Computers and Mathematics with Applications, 28 (6):33–44, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001502.

Delsanto:1994:TSP

[DKSI94] P. P. Delsanto, G. Kaniadakis, M. Scalerandi, and D. Iordache. Time scaling in the parallel processing simulation of diffusion processes. Computers and Mathematics with Applications, 27(6):51–61, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901104.

Dudas:1998:CCK

[DKW98] T. Dudás, B. Klinz, and G. J. Woeginger. The computational complexity of the κ-minimum spanning tree problem in graded matrices. Computers and Mathematics with Applications, 36 (5):61–67, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001503.

Ding:1999:EAS

[DL99] Xie Ping Ding and Chun Lin Luo. Existence and algorithm for solving some generalized mixed variational inequalities. Computers and Mathematics with Applications, 37(3):23–30, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000437.

Drakhlin:1999:VMC

[DLS99] M. E. Drakhlin, E. Litsyn, and E. Stepanov. Variational methods for a class of nonlocal functionals. Computers and Mathematics with Applications, 37(4–5):79–100, February/March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000619.

Dahlke:1995:AUP

[DM95] S. Dahlke and P. Maass. The affine uncertainty principle in one and two dimensions. Computers and Mathematics with Applications, 30(3– 6):293–305, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001085.

Dattoli:1995:UPV

[DMC^{+}95] G. Dattoli, G. Maino, C. Chiccoli, S. Lorenzutta, and A. Torre. A unified point of view on the theory of generalized Bessel functions. Computers and Mathematics with Applications, 30 (7):113–125, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500130Q.

Debinska-Nagorska:1998:ASG

[DNJS98] A. Dȩbińska-Nagórska, A. Just, and Z. Stempień. Analysis and semidiscrete Galerkin approximation of a class of nonlinear parabolic optimal control problems. Computers and Mathematics with Applications, 35(6):95–103, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000224.

Dikshit:1991:DSW

[DO91] H. P. Dikshit and A. Ojha. Dimensions of spaces of Wachspress type C^{1}-rational finite elements. Computers and Mathematics with Applications, 22(3):23–26, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190066D.

Dohi:1992:OTS

[DO92] Tadashi Dohi and Shunji Osaki. Optimal trading of stock options under alternative strategy. Computers and Mathematics with Applications, 24(1–2):127–134, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290237C.

Dohi:1995:OIP

[DO95] T. Dohi and S. Osaki. Optimal inventory policies under product obsolescent circumstance. Computers and Mathematics with Applications, 29(3):23–30, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400225A.

delOlmo:1991:SIM

[dOF91] V. del Olmo and R. Fuster.

Some iterative methods related to Frobenius norm minimization. Computers and Mathematics with Applications, 22(10):121–126, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190199E.

Doha:1990:ASP

[Doh90] E. H. Doha. An accurate solution of parabolic equations by expansion in ultraspherical polynomials. Computers and Mathematics with Applications, 19(4):75–88, ???? 1990. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090139B.

Doha:1991:CDE

[Doh91] E. H. Doha. The coefficients of differentiated expansions and derivatives of ultraspherical polynomials. Computers and Mathematics with Applications, 21 (2–3):115–122, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190089M.

Dorogovtsev:1990:NSC

[Dor90] A. Ya. Dorogovtsev. Necessary and sufficient conditions for existence of stationary and periodic solutions of a stochastic difference equation in Hilbert space. Computers and Mathematics with Applications, 19(1):31–37, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090079Y.

Dorovsky:1997:HNMa

[Dor97] V. N. Dorovsky. A hydrodynamic nonlinear model of bubble liquid. Computers and Mathematics with Applications, 33(6):1–12, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000266.

Dorovsky:1998:HPSa

[Dor98a] V. N. Dorovsky. The hydrodynamics of particles suspended in a melt with the self-consistent concentration field of an admixture (Part I). Computers and Mathematics with Applications, 35 (11):27–37, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000819.

Dorovsky:1998:HPSb

[Dor98b] V. N. Dorovsky. The hydrodynamics of particles suspended in a melt with the self-consistent concentration field of an admixture (Part II). Computers and Mathematics with Applications, 35 (11):39–47, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000820.

Dorovsky:1999:APM

[Dor99] V. N. Dorovsky. Absorption of particles from a melt by the surface of a bubble of variable size — Part III. Computers and Mathematics with Applications, 37(7):41–59, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000863.

Derr:1993:NMD

[DOS93] L. Derr, C. Outlaw, and D. Sarafyan. A new method for derivation of continuous Runge–Kutta formulas. Computers and Mathematics with Applications, 26(3): 7–13, August 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193901055.

Dosly:1998:FDH

[Doš98]  O. Došlý.  Factorization

of disconjugate higher-order Sturm–Liouville difference operators. Computers and Mathematics with Applications, 36(10–12):227–234, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800230.

Das:1998:SIF

[DPD98] S. Das, B. Patra, and L. Debnath. Stress intensity factors for an interfacial crack between an orthotropic halfplane bonded to a dissimilar orthotropic layer with a punch. Computers and Mathematics with Applications, 35(12):27–40, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000947.

Debnath:1996:EEG

[DR96] L. Debnath and T. Roy.

Energy eigenvalues of the Gellman–Feynman hydrogen atom. Computers and Mathematics with Applications, 32(7):9–11, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001502.

Dagnino:1997:PIS

[DR97] C. Dagnino and P. Rabinowitz. Product integration of singular integrands using quasi-interpolatory splines. Computers and Mathematics with Applications, 33 (1–2):59–67, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002192.

Dragomir:1999:OII

[Dra99] S. S. Dragomir. The Ostrowski’s integral inequality for Lipschitzian mappings and applications. Computers and Mathematics with Applications, 38(11–12):33– 37, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002825.

Dress:1990:ELS

[Dre90] W. B. Dress. Electronic life and synthetic intelligent systems. Computers and Mathematics with Applications, 20 (4–6):303–312, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090335H.

Driebe:1997:JBO

[Dri97] D. J. Driebe. Jordan blocks in a one-dimensional Markov map. Computers and Mathematics with Applications, 34 (2–4):103–109, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700120X.

Demetrovics:1991:DT

[DRS91] J. Demetrovics, L. Rónyai, and Hua Nam Son. Dependency types. Computers and Mathematics with Applications, 21(1):25–33, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190228V.

deSwart:1996:ESM

[dSB96] J. J. B. de Swart and J. G. Blom. Experiences with sparse matrix solvers in parallel ODE software. Computers and Mathematics with Applications, 31(9):43–55, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000417.

Diamond:1996:SPD

[DSKP96] P. Diamond, M. Suzuki, P. Kloeden, and P. Pokrovskii. Statistical properties of discretizations of a class of chaotic dynamical systems. Computers and Mathematics with Applications, 31(11): 83–95, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600065X.

Demetrovics:1993:AGA

[DT93] J. Demetrovics and V. D.

Thi. Algorithms for generating an Armstrong relation and inferring functional dependencies in the relational datamodel. Computers and Mathematics with Applications, 26(4):43–55, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390033R.

Demetrovics:1995:FFD

[DT95] J. Demetrovics and V. D.

Thi. Family of functional dependencies and its equivalent descriptions. Computers and Mathematics with Applications, 29(4):101–109, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/089812219400243E.

Dattoli:1998:GBF

[DTLM98] G. Dattoli, A. Torre, S. Lorenzutta, and G. Maino. Generalized Bessel functions and Kapteyn series. Computers and Mathematics with Applications, 35(8):117–125, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000509.

Du:1994:FEM

[Du94] Q. Du. Finite element methods for the time-dependent Ginzburg–Landau model of superconductivity. Computers and Mathematics with Applications, 27(12):119– 133, June 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900914.

Ducomet:1995:DEE

[Duc95a] B. Ducomet. Decay of electromagnetic energy in a perturbed half-space. Computers and Mathematics with Applications, 29(3):13–21, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122194002249.

Ducomet:1995:DSW

[Duc95b] B. Ducomet. Decay of solutions of the wave equation outside rough surfaces. Computers and Mathematics with Applications, 29 (1):89–108, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002094.

Dragomir:1997:IOG

[DW97] S. S. Dragomir and S. Wang.

An inequality of Ostrowski-Grüss’ type and its applications to the estimation of error bounds for some special means and for some numerical quadrature rules. Computers and Mathematics with Applications, 33(11):15–20, June 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000849.

Deng:1995:NWC

[DWC95] Young-Ching Deng, Yue-Li Wang, and Jou-Ming Chang. A new way of counting n^{m}. Computers and Mathematics with Applications, 29(12):

75–80, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500060C.

Dwyer:1993:ENF

[Dwy93] R. A. Dwyer. The expected number of k-faces of a Voronoi diagram. Computers and Mathematics with Applications, 26(5):13–19, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193900687.

Esogbue:1990:CEC

[EA90] A. O. Esogbue and C. W. Ahn. Computational experiments with a class of dynamic programming algorithms of higher dimensions. Computers and Mathematics with Applications, 19(11):3–23, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190901449.

El-Affendi:1992:ECP

[EA92] M. A. El-Affendi. Estimating computer performance metrics when the service and interarrival times are of the truncated normal type. Computers and Mathematics with Applications, 23(10): 35–40, May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290052J.

Erenguc:1991:AII

[EB91] S. S. Erenguc and H. P. Benson. An algorithm for indefinite integer quadratic programming. Computers and Mathematics with Applications, 21(6–7):99–106, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190164Y.

El-Banna:1993:SPM

[EB93]  A.-Z. H. El-Banna.  A

study on parametric multiobjective programming problems without differentiability. Computers and Mathematics with Applications, 26 (12):87–92, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390062Z.

Echevarria:1995:NSS

[Ech95] R. Echevarría. The numerical solution of some elliptic problems with nonlinear discontinuities using exact regularization. Computers and Mathematics with Applications, 29(12):57–66, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000587.

Echevarria:1996:NCS

[Ech96] R. Echevarría. The numerical computation of steady vortex pairs. Computers and Mathematics with Applications, 32 (5):117–130, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196856847.

Echevarria:1998:NSS

[Ech98]  R. Echevarría.  The nu-

merical solution of some elliptic nonscalar problems with nonlinear discontinuities. Computers and Mathematics with Applications, 35 (3):131–147, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002848.

El-Din:1991:EMS

[ED91a] M. M. Mohie El-Din. Estimation of the mean and standard deviation of Burr distribution by order statistics. Computers and Mathematics with Applications, 21(2–3):23–27, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190077H.

El-Din:1991:MLW

[ED91b] M. M. Mohie El-Din. Maximum likelihood weights of the location of the Burr distribution. Computers and Mathematics with Applications, 21(4):33–38, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190245Y.

Ebrahimpour:1991:ATN

[EDD91] M. Reza Ebrahimpour, John F.

Dorsey, and Stephen G. Demko. An acceleration technique for the Newton solution of quadratic equations. Computers and Mathematics with Applications, 22(12):79–87, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190151S.

El-Daou:1992:DEP

[EDNO92] M. K. El-Daou, S. Namasivayam, and E. L. Ortiz. Differential equations with piecewise approximate coefficients: Discrete and continuous estimation for initial and boundary value problems. Computers and Mathematics with Applications, 24(4):33–47, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192900053.

El-Daou:1993:EAT

[EDO93] M. K. El-Daou and E. L. Ortiz. Error analysis of the Tau method: Dependence of the error on the degree and on the length of the interval of approximation. Computers and Mathematics with Applications, 25(7):33–45, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390310R.

El-Daou:1996:DEL

[EDO96] M. K. El-Daou and E. L. Ortiz. Differential equations with locally perturbed coefficients: High order error estimates for function and derivative. Computers and Mathematics with Applications, 31(2):73–84, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001956.

El-Daou:1997:NEE

[EDO97] M. K. El-Daou and E. L. Ortiz. The number of extrema of the error function of a class of methods for differential equations. Computers and Mathematics with Applications, 33(1–2):69–79, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002209.

El-Daou:1998:RFG

[EDO98] M. K. El-Daou and E. L.

Ortiz. A recursive formulation of Galerkin’s method in terms of the tau method: Bounded and unbounded domains. Computers and Mathematics with Applications, 35 (12):83–94, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000984.

El-Daou:1993:UAT

[EDOS93] M. K. El-Daou, E. L. Ortiz, and H. Samara. A unified approach to the Tau method and Chebyshev series expansion techniques. Computers and Mathematics with Applications, 25(3):73–82, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390145L.

Edwards:1992:FCM

[Edw92] Thomas A. Edwards. Fluid/ chemistry modeling for hypersonic flight analysis. Computers and Mathematics with Applications, 24(5–6):25–36, September 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290038J.

Eidelman:1997:IFL

[EG97] Y. Eidelman and I. Gohberg. Inversion formulas and linear complexity algorithm for diagonal plus semiseparable matrices. Computers and Mathematics with Applications, 33(4):69–79, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000084.

Eidelman:1998:LAB

[EG98] Y. Eidelman and I. Gohberg. A look-ahead block Schur algorithm for diagonal plus semiseparable matrices. Computers and Mathematics with Applications, 35 (10):25–34, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000698.

Egert:1991:MRS

[Ege91] Eva Egert. The model RAS and its sensitivity analysis. Computers and Mathematics with Applications, 21(1): 35–47, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190229W.

El-Gindy:1995:CAS

[EGEHSEK95] T. M. El-Gindy, H. M. El-Hawary, M. S. Salim, and M. El-Kady. A Chebyshev approximation for solving optimal control problems. Computers and Mathematics with Applications, 29(6): 35–45, March 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500005J.

Ezquerro:1998:AIF

[EGHS98] J. A. Ezquerro, J. M.

Gutiérrez, M. A. Hernández, and M. A. Salanova. The application of an inverse-free Jarratt-type approximation to nonlinear integral equations of Hammerstein-type. Computers and Mathematics with Applications, 36(4): 9–20, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001370.

El-Gabali:1996:SGR

[EGK96] M. El-Gabali and S. Kalla.

Some generalised radiation field integrals. Computers and Mathematics with Applications, 32(12):121– 128, December 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219600212X.

Esser:1995:SEB

[EGLN95] H. Esser, St. J. Goebbels, G. Lüttgens, and R. J. Nessel. Sharp error bounds for the Crank–Nicolson and Saulyev difference scheme in connection with an initial boundary value problem for the inhomogeneous heat equation. Computers and Mathematics with Applications, 30(3– 6):59–68, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000860.

Ehrig:1992:IGG

[EHK92] Hartmut Ehrig, Annegret Habel, and Hans-Jörg Kreowski. Introduction to graph grammars with applications to semantic networks. Computers and Mathematics with Applications, 23(6–9):557– 572, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290124Z.

Eiseman:1992:CPG

[Eis92] Peter R. Eiseman. Control point grid generation. Computers and Mathematics with Applications, 24(5– 6):57–67, September 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290041F.

Egecioglu:1992:PPC

[EK92] Ömer Eğecioğlu and Çetin Kaya Koç. Parallel prefix computation with few processors. Computers and Mathematics with Applications, 24(4):77–84, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900097.

Eloe:1994:ULG

[EK94] P. W. Eloe and E. R. Kaufmann. A unique limiting Green’s function for a class of singular boundary value problems. Computers and Mathematics with Applications, 28(1–3):93–99, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194000964.

Emelyanov:1997:TNF

[EK97] S. V. Emelyanov and S. K.

Korovin. Theory of nonlinear feedback under uncertainty. Computers and Mathematics with Applications, 34 (2–4):273–323, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001296.

Egdbareciogdbarlu:1991:FCC

[EKC91] Omer Eğecioğlu, Çetin K. Koç, and Josep RifàI Coma. Fast computation of continued fractions. Computers and Mathematics with Applications, 21(2–3):167–169, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190095L.

Ekel:1999:ADM

[Eke99] P. Ya. Ekel. Approach to decision making in fuzzy environment. Computers and Mathematics with Applications, 37 (4–5):59–71, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000590.

Eck:1992:BSB

[EL92] Matthias Eck and Dieter Lasser. B-spline-Bézier representation of geometric spline curves: Quartics and quintics. Computers and Mathematics with Applications, 23 (11):23–39, June 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290066Q.

Ervin:1996:PEE

[EL96] V. J. Ervin and W. J. Layton. A posteriori error estimation for two level discretizations of flows of electrically conducting, incompressible fluids. Computers and Mathematics with Applications, 31(11):105–114, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000673.

Esogbue:1999:CVF

[EL99] A. O. Esogbue and Baoding Liu. Cluster validity for fuzzy criterion clustering. Computers and Mathematics with Applications, 37 (11–12):95–100, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001479.

Eloe:1998:GCF

[Elo98] P. W. Eloe. A generalization of concavity for finite differences. Computers and Mathematics with Applications, 36(10–12):109–113, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800138.

Esogbue:1994:AFA

[EM94] A. O. Esogbue and J. A. Murrell. Advances in fuzzy adaptive control. Computers and Mathematics with Applications, 27(9–10):29–35, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901201.

Escalante:1998:DAC

[ER98] R. Escalante and M. Raydan. Dykstra’s algorithm for constrained least-squares rectangular matrix problems. Computers and Mathematics with Applications, 35(6): 73–79, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000200.

Ermini:1992:SUI

[Erm92] Luigi Ermini. Some unresolved issues in the application of control theory to economic policy-making. Computers and Mathematics with Applications, 24(8–9): 125–136, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290192K.

El-Sheikh:1997:JTA

[ES97] M. G. El-Sheikh. The justification of the truncation applied to homogeneous integral equations with Cauchy’s kernel. Computers and Mathematics with Applications, 33(3):65–71, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002374.

El-Sayed:1998:NMR

[ESIP98] S. M. El-Sayed, I. G. Ivanov, and M. G. Petkov. A new modification of the Rojo method for solving symmetric circulant five-diagonal systems of linear equations. Computers and Mathematics with Applications, 35(10): 35–44, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000704.

El-Sheikh:1998:MIE

[ESKG98] M. G. El-Sheikh, M. E.

Khalifa, and V. Gavdzinski. The method of integral equation formulation and the unbounded solutions of elastic contact problems. Computers and Mathematics with Applications, 36(1):33–39, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001060.

Esogbue:1990:RCS

[Eso90]  A. O. Esogbue.  Reflec-

tions on the contributions to the second Bellman continuum. Computers and Mathematics with Applications, 19(11):1–2, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190901438.

Esogbue:1991:CAA

[Eso91] Augustine O. Esogbue. Computational aspects and applications of a branch and bound algorithms for fuzzy multistage decision processes. Computers and Mathematics with Applications, 21(11– 12):117–127, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190113I.

Esogbue:1999:CCS

[Eso99a] A. O. Esogbue. The computational complexity of some fuzzy dynamic programs. Computers and Mathematics with Applications, 37 (11–12):47–51, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900142X.

Esogbue:1999:FDP

[Eso99b] A. O. Esogbue. Fuzzy dynamic programming, fuzzy adaptive neuro control, and the general medical diagnosis problem. Computers and Mathematics with Applications, 37(11–12):37–45, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001418.

Esogbue:1999:MGA

[Eso99c] A. O. Esogbue. In memoriam: George Adomian (1922-1996). Computers and Mathematics with Applications, 37 (11–12):xiii–xiv, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219991307X.

Esogbue:1999:P

[Eso99d] A. O. Esogbue. Preface.

Computers and Mathematics with Applications, 37 (11–12):xv–xvi, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001352.

Etingof:1991:IPP

[Eti91] Pavel I. Etingof. Inverse problems of potential theory and flows in porous media with time-dependent free boundary. Computers and Mathematics with Applications, 22(4–5):93–99, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190134P.

Esham:1991:GMS

[EY91] Benjamin F. Esham, Jr. and Elizabeth Greenwell Yanik. Galerkin methods for a singularly perturbed hyperbolic problem with nonlocal nonlinearity. Computers and Mathematics with Applications, 22(2):1–22, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900386.

Erdos:1993:UBQ

[EZ93] P. Erdős and Zhenxiang Zhang. Upper bound of Σ1(a_{i} log a_{i}) for quasiprimitive sequences. Computers and Mathematics with Applications, 26(3):1–5, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901044.

Fabiano:1995:SPS

[Fab95] R. H. Fabiano. Stability preserving spline approximations for scalar functional differential equations. Computers and Mathematics with Applications, 29(8):87–94, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500032T.

Fabian:1997:SAS

[Fab97] V. Fabian. Simulated annealing simulated. Computers and Mathematics with Applications, 33(1–2):81–94, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002210.

Fairag:1998:TLF

[Fai98]  F. Fairag.  A two-level

finite-element discretization of the stream function form of the Navier–Stokes equations. Computers and Mathematics with Applications, 36(2):117–127, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001230.

Farago:1991:FEMa

[Far91a] I. Farago. Finite element method for solving nonlinear parabolic systems. Computers and Mathematics with Applications, 21(1):49–57, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191902302.

Farago:1991:FEMb

[Far91b] István Farago. Finite element method for solving nonlinear parabolic equations. Computers and Mathematics with Applications, 21(1):59–69, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190231R.

Farago:1996:QPN

[Far96a] I. Faragó. Qualitative properties of the numerical solution of linear parabolic problems with nonhomogeneous boundary conditions. Computers and Mathematics with Applications, 31(4– 5):143–150, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500226X.

Farrett:1996:IMI

[Far96b] P. W. Farrett. Intelligent musical instrument systems. Computers and Mathematics with Applications, 32(1): 37–48, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000855.

Farago:1999:REC

[Far99] I. Faragó. Regular and exponential convergence of difference schemes for the heatconduction equation. Computers and Mathematics with Applications, 38(9–10):71– 77, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900262X.

Faure:1994:NNM

[Fau94] R. Faure. Neural networks and mathematical distributions theory: periodical and almost periodical Dirac firings. Computers and Mathematics with Applications, 27(6):93–97, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901139.

Fazekas:1994:HTG

[Faz94a] I. Fazekas. Hellinger transform of Gaussian autoregressive processes. Computers and Mathematics with Applications, 27(7):15–21, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901457.

Fazekas:1994:MLE

[Faz94b] I. Fazekas. Maximum likelihood estimators of parameters of multidimensional stationary Gaussian AR processes. Computers and Mathematics with Applications, 27 (8):19–24, April 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900639.

Fazekas:1999:APE

[FBL99] I. Fazekas, S. Baran, and J. Lauridsen. Asymptotic properties of an estimator in errors-in-variables models in the presence of validation data. Computers and Mathematics with Applications, 38 (5–6):31–39, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002126.

Ferreira:1995:DMO

[FCP95] J. M. Ferreira, A. Chattopadhyay, and S. J. Pringnitz. Development of a multiobjective optimization procedure for reducing edge delamination stresses in composite plates. Computers and Mathematics with Applications, 29(6):81–97, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500010V.

Feuillebois:1990:PCN

[Feu90] F. Feuillebois. Precise calculation of numerical derivatives using Lagrange and Hermite interpolation. Computers and Mathematics with Applications, 19(5):1–11, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900963.

Feuillebois:1991:NCS

[Feu91] F. Feuillebois. Numerical calculation of singular integrals related to Hankel transform. Computers and Mathematics with Applications, 21(2–3):87–94, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190085I.

From:1992:CIE

[FF92] S. G. From and H. A.

Farhat. Confidence intervals for expected coverage from a beta testability model. Computers and Mathematics with Applications, 24 (7):97–107, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290158E.

Ferrari:1993:NEO

[FG93] André Ferrari and Efim A.

Galperin. Numerical experiments with one-dimensional adaptive cubic algorithm. Computers and Mathematics with Applications, 25(10– 11):47–56, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902809.

Frontini:1997:DCM

[FG97] M. Frontini and L. Gotusso. A discrete conservative model for the linear vibrating string and rod. Computers and Mathematics with Applications, 33(10):53–65, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700076X.

Frontini:1998:CLM

[FG98] M. Frontini and L. Gotusso. Conservative linear models for the vibrating string and rod free to slide, at one end point, on a vertical guide. Computers and Mathematics with Applications, 35(6): 41–50, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000169.

Friedrich:1990:GED

[FGN90] G. Friedrich, G. Gottlob, and W. Nejdl. Generating efficient diagnostic procedures from model-based knowledge using logic programming techniques. Computers and Mathematics with Applications, 20 (9–10):57–72, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090112W.

Farago:1999:OMC

[FH99] I. Faragó and R. Horváth.

An optimal mesh choice in the numerical solution of the heat equation. Computers and Mathematics with Applications, 38(9–10):79– 85, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002631.

Fang:1999:LPF

[FHWW99] S.-C. Fang, C.-F. Hu, H.-F.

Wang, and S.-Y. Wu. Linear programming with fuzzy coefficients in constraints. Computers and Mathematics with Applications, 37(10): 63–76, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001261.

Filippova:1993:NEP

[Fil93] T. F. Filippova. A note on the evolution property of the assembly of viable solutions to a differential inclusion. Computers and Mathematics with Applications, 25 (2):115–121, January 1993.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390228N.

Finta:1999:AOI

[Fin99] B. Finta. About Ostrowski’s iterative method in infinite dimensional spaces. Computers and Mathematics with Applications, 38(9–10):193– 196, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002746.

Fjelstad:1991:BTE

[Fje91a] P. Fjelstad. The binomial theorem and the extended Pascal’s triangle. Computers and Mathematics with Applications, 21(9):11–17, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190121J.

Fjelstad:1991:EPT

[Fje91b] P. Fjelstad. Extending Pascal’s triangle. Computers and Mathematics with Applications, 21(9):1–4, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190119O.

Fjelstad:1991:PTT

[Fje91c] P. Fjelstad. Pascal’s triangle on a torus and on a binary tree. Computers and Mathematics with Applications, 21(9):5–10, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190120S.

Fowler:1995:NFT

[FK95] A. C. Fowler and G. Kember. A nonlinear filtering technique for multi-oscillator systems. Computers and Mathematics with Applications, 29(4):55–67, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400238G.

Fazekas:1997:APE

[FK97] I. Fazekas and A. G. Kukush.

Asymptotic properties of an estimator in nonlinear functional errors-in-variables models with dependent error terms. Computers and Mathematics with Applications, 34 (10):23–39, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002046.

Frolov:1998:IAN

[FK98a] G. A. Frolov and M. Y. Kitaev. Improvement of accuracy in numerical methods for inverting Laplace transforms based on the Post–Widder formula. Computers and Mathematics with Applications, 36(5):23–34, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001473.

Frolov:1998:PNI

[FK98b] G. A. Frolov and M. Y. Kitaev. A problem of numerical inversion of implicitly defined Laplace transforms. Computers and Mathematics with Applications, 36 (5):35–44, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001485.

Fukumoto:1992:SCG

[FKO92] S. Fukumoto, N. Kaio, and S. Osaki. A study of checkpoint generations for a database recovery mechanism. Computers and Mathematics with Applications, 24(1–2):63–70, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290229B.

Favati:1991:PAM

[FL91] Paola Favati and Grazia Lotti. Parallel algorithms for matrix polynomial division. Computers and Mathematics with Applications, 22(7):37–42, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901798.

Fazekas:1996:SPH

[FL96] I. Fazekas and F. Liese. Some properties of the Hellinger transform and its application in classification problems. Computers and Mathematics with Applications, 31(8):107–116, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000351.

Flippen:1994:COS

[Fli94a] L. D. Flippen, Jr. Constitutiveoperator smoothing by condensation. Computers and Mathematics with Applications, 27(6):5–18, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901066.

Flippen:1994:CDS

[Fli94b] L. D. Flippen, Jr. Current dynamic substructuring methods as approximations to condensation model reduction. Computers and Mathematics with Applications, 27(12):17–29, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900825.

Flippen:1994:TCM

[Fli94c] L. D. Flippen, Jr. A theory of condensation model reduction. Computers and Mathematics with Applications, 27(2):9–40, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900345.

Flippen:1995:IBC

[Fli95] L. D. Flippen, Jr. Interpolationbased condensation of algebraic semi-discrete models with frequency response application. Computers and Mathematics with Applications, 29(9):39–52, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500036X.

Fairweather:1994:EEB

[FLM94] G. Fairweather and J. C. López-Marcos. An explicit extrapolated box scheme for the Gurtin–Maccamy equation. Computers and Mathematics with Applications, 27(2):41–53, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900353.

Fu:1998:HQC

[FLNW98] A. W. Fu, W. K. Lau, F. K.

Ng, and M. H. Wong. Hypercube quorum consensus for mutual exclusion and replicated data management. Computers and Mathematics with Applications, 36 (5):45–59, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001497.

Florchinger:1995:NDM

[Flo95]  P. Florchinger.  A new

decomposition method for stochastic dynamic stabilization. Computers and Mathematics with Applications, 29(3):31–36, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400226B.

Florchinger:1997:GSC

[Flo97] P. Florchinger. Global stabilization of composite stochastic systems. Computers and Mathematics with Applications, 33(6):127–135, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000382.

Favati:1992:AEE

[FLR92] P. Favati, G. Lotti, and F. Romani. Asymptotic expansion of error in interpolatory quadrature. Computers and Mathematics with Applications, 24(10):99–104, November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290022A.

Favati:1995:PKB

[FLR95] P. Favati, G. Lotti, and F. Romani. Peano kernel behaviour and error bounds for symmetric quadrature formulas. Computers and Mathematics with Applications, 29(6): 27–34, March 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500004I.

Ferng:1997:SIL

[FLW97] W. R. Ferng, Wen-Wei Lin, and Chern-Shuh Wang. The shift-inverted J-Lanczos algorithm for the numerical solutions of large sparse algebraic Riccati equations. Computers and Mathematics with Applications, 33(10):23–40, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000746.

See [Ano99-245].

Ferng:1999:NAU

[FLW99] W. R. Ferng, Wen-Wei Lin, and Chern-Shuh Wang. Numerical algorithms for undamped gyroscopic systems. Computers and Mathematics with Applications, 37(1): 49–66, January 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002417.

Ford:1996:MCM

[FM96] J. A. Ford and I. A. Moghrabi.

Minimum curvature multistep quasi-Newton methods. Computers and Mathematics with Applications, 31(4– 5):179–186, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002308.

Franconi:1992:PHL

[FMS92] Enrico Franconi, Bernardo Magnini, and Oliviero Stock. Prototypes in a hybrid language with primitive descriptions. Computers and Mathematics with Applications, 23 (6–9):543–555, March/May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290123Y.

Freilich:1991:ULE

[FO91] J. H. Freilich and E. L. Ortiz. Upper and lower error estimation for the Tau method: Some remarks on a problem of rational approximation. Computers and Mathematics with Applications, 22(10):89–97, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190195A.

Fogg:1993:VUF

[FO93] Iain Fogg and Maria Orlowska. Valid unbiased fragmentation for distributed database design. Computers and Mathematics with Applications, 25(9):97–106, May 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390136J.

Fogel:1994:AEP

[Fog94] D. B. Fogel. Applying evolutionary programming to selected control problems. Computers and Mathematics with Applications, 27(11): 89–104, July 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901007.

Foley:1992:MBS

[Fol92]  T. A. Foley.  The map

and blend scattered data interpolant on a sphere. Computers and Mathematics with Applications, 24 (12):49–60, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290171D.

Fox:1994:RRV

[Fox94] B. L. Fox. Random restarting versus simulated annealing. Computers and Mathematics with Applications, 27(6):33–35, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901082.

Frankel:1996:RML

[Fra96]  J. I. Frankel.  Residual-

minimization least-squares method for inverse heat conduction. Computers and Mathematics with Applications, 32(4):117–130, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001307.

Frezzotti:1998:MDE

[Fre98] A. Frezzotti. Molecular dynamics and Enskog theory calculation of shock profiles in a dense hard sphere gas. Computers and Mathematics with Applications, 35(1– 2):103–112, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002617.

Feher:1995:WTI

[FS95] F. Fehér and M. J. Strauss. Weak-type interpolation and orbits. Computers and Mathematics with Applications, 30(3–6):417–431, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001174.

Fiorentino:1996:FPS

[FS96] G. Fiorentino and S. Serra. Fast parallel solvers for elliptic problems. Computers and Mathematics with Applications, 32(2):61–68, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001034.

Feng:1997:AND

[FT97a] J. Feng and B. Tirozzi. An analysis on neural dynamics with saturated sigmoidal functions. Computers and Mathematics with Applications, 34(10):71–99, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002083.

Feng:1997:CTK

[FT97b] J. F. Feng and B. Tirozzi.

Convergence theorems for the Kohonen feature mapping algorithms with VL-RPs. Computers and Mathematics with Applications, 33(3):45–63, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002362.

Fu:1998:OND

[FT98] Sheng-Chen Fu and Long-Yi Tsai. Oscillation in nonlinear difference equations. Computers and Mathematics with Applications, 36(10–12):193– 201, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800205.

Funaro:1997:SRA

[Fun97] D. Funaro. Some remarks about the collocation method on a modified Legendre grid. Computers and Mathematics with Applications, 33 (1–2):95–103, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002222.

Fusaro:1990:I

[Fus90] B. A. Fusaro. Introduction. Computers and Mathematics with Applications, 19 (3):vii, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090036J.

Fang:1996:SMM

[FW96] S.-C. Fang and Soon-Yi Wu. Solving min-max problems and linear semi-infinite programs. Computers and Mathematics with Applications, 32 (6):87–93, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001459.

Fukao:1999:CPC

[FWI99] T. Fukao, J. Wu, and K. Ikeda. Critical phenomena in a collective computation algorithm for combinatorial optimization problems. Computers and Mathematics with Applications, 37 (11–12):213–223, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001595.

Fourie:1993:GAP

[FY93] J. P. Fourie and Y. Yavin.

The guidance accuracy of a projectile using a wind stabilized seeker. Computers and Mathematics with Applications, 25(3):27–36, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390140Q.

Frangos:1999:CSS

[FYS99] C. Frangos, Y. Yavin, and M. Sniedovich. Control strategies for a stochastic flexible manufacturing and assembly system model. Computers and Mathematics with Applications, 38(5–6):243– 249, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002308.

Ferreira:1993:BPS

[FZ93] Afonso G. Ferreira and Janez Zerovnik. Bounding the probability of success of stochastic methods for global optimization. Computers and Mathematics with Applications, 25 (10–11):1–8, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390275Z.

Gabet:1994:DMD

[Gab94a] L. Gabet. The decomposition method and distributions. Computers and Mathematics with Applications, 27(3):41–49, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900450.

Gabet:1994:TFA

[Gab94b] L. Gabet. The theoretical foundation of the Adomian method. Computers and Mathematics with Applications, 27(12):41–52, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900841.

Galperin:1991:ICA

[Gal91a] Efim A. Galperin. The integer cubic algorithm. Computers and Mathematics with Applications, 21(6–7):215– 224, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191901754.

Galperin:1991:PMC

[Gal91b] Efim A. Galperin. Problemmethod classification in optimization and control. Computers and Mathematics with Applications, 21(6–7):1–6,

????  1991.  CODEN

CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219190155W.

Galperin:1991:P

[Gal91c] Elfim A. Galperin. Preface. Computers and Mathematics with Applications, 21 (6–7):vii, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190154V.

Galperin:1993:AAA

[Gal93a] E. A. Galperin. The alpha algorithm and the application of the cubic algorithm in case of unknown Lipschitz constant. Computers and Mathematics with Applications, 25(10–11):71–78, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902832.

Galperin:1993:CAG

[Gal93b] E. A. Galperin. The cubic algorithm for global games with application to pursuit-evasion games. Computers and Mathematics with Applications, 26 (6):13–31, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390114B.

Galperin:1993:FCA

[Gal93c] E. A. Galperin. The fast cubic algorithm. Computers and Mathematics with Applications, 25(10–11):147– 160, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902898.

Galperin:1993:VFC

[Gal93d] E. A. Galperin. Validity of feedback controls depending on higher order derivatives. Computers and Mathematics with Applications, 25 (10–11):173–185, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902913.

Galantai:1999:PAP

[Gal99a] A. Galántai. Parallel ABS projection methods for linear and nonlinear systems with block arrowhead structure. Computers and Mathematics with Applications, 38(9– 10):11–17, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002588.

Galperin:1999:LTA

[Gal99b]  E. A. Galperin.  Linear

time algorithms for linear programming. Computers and Mathematics with Applications, 37(4–5):199–208, February/March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000693.

Galperin:1999:SOT

[Gal99c] E. A. Galperin. Some old traditions in mathematics and in mathematical education. Computers and Mathematics with Applications, 37(4–5): 9–17, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000553.

Galperin:1999:P

[Gal99d] Efin A. Galperin. Preface.

Computers and Mathematics with Applications, 37(4– 5):xiii–xiv, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900053X.

Gandolfo:1992:SAC

[Gan92] Giancarlo Gandolfo. Sensitivity analysis in continuous time econometric models. Computers and Mathematics with Applications, 24(8– 9):43–55, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290186L.

Garay:1996:VCC

[Gar96]  B. M. Garay.  Various

closeness concepts in numerical ODE’s. Computers and Mathematics with Applications, 31(4–5):113– 119, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002235.

Gaspar:1995:IMS

[Gás95] C. Gáspár. An interative and multigrid solution of boundary integral equations. Computers and Mathematics with Applications, 29(7): 89–101, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500021P.

Gautschi:1997:MQP

[Gau97] W. Gautschi. Moments in quadrature problems. Computers and Mathematics with Applications, 33(1–2):105– 118, January 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002234.

Gora:1990:HDP

[GB90] P. Góra and A. Boyarsky.

Higher-dimensional point transformations and asymptotic measures for cellular automata. Computers and Mathematics with Applications, 19(12):13–31, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090247H.

Gora:1992:MSH

[GB92] P. Góra and A. Boyarsky.

Modelling and simulating higher dimensional chaotic data. Computers and Mathematics with Applications, 24 (11):101–105, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290035G.

Gora:1996:ACC

[GB96] P. Góra and A. Boyarsky. An algorithm to control chaotic behavior in one-dimensional maps. Computers and Mathematics with Applications, 31(6):13–22, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000028.

Green:1991:HFS

[GBB91] H. S. Green, A. J. Bracken, and L. Bass. Harmonic functions satisfying a radiation boundary condition. Computers and Mathematics with Applications, 22(2):23–38, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900397.

Guo:1999:FRE

[GC99] Bao Zhu Guo and Jin Cheng Chen. The first real eigenvalue of a one-dimensional linear thermoelastic system. Computers and Mathematics with Applications, 38(11– 12):249–256, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900303X.

Guo:1998:DIR

[GCLT98] Jinn-Kwei Guo, Chin-Hsing Chen, Jiann-Der Lee, and Jia-Ming Tsai. 3-d image reconstruction of brain blood vessels from angiograms. Computers and Mathematics with Applications, 35(8): 79–94, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000467.

Gong:1998:SDN

[GCT98] Xiangyang Gong, Wanyi Chen, and Fengsheng Tu. The stability and design of nonlinear neural networks. Computers and Mathematics with Applications, 35(8):1–7, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000406.

Gemignani:1993:RIO

[Gem93] Luca Gemignani. Rational interpolation via orthogonal polynomials. Computers and Mathematics with Applications, 26(5):26–34, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390070C.

Gehrlein:1990:LEM

[GF90] W. V. Gehrlein and P. C.

Fishburn. Linear extension majority cycles for small (n ≤

9) partial orders. Computers and Mathematics with Applications, 20(2):41–44, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090239G.

Ge:1994:SOO

[GF94a] Y.-R. Ge and J. M. Fitzpatrick. A set of one-to-one two-dimensional biquadratic transformations. Computers and Mathematics with Applications, 27(11):13–25, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900930.

Gehrlein:1994:SGN

[GF94b] W. V. Gehrlein and P. C. Fishburn. The smallest graphs with niche number three. Computers and Mathematics with Applications, 27(4):53–57, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219490054X.

Gehrlein:1996:NNF

[GF96] W. V. Gehrlein and P. Fishburn. Niche number four. Computers and Mathematics with Applications, 32 (9):51–54, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001769.

Guan:1997:LPS

[GF97] S. Guan and S.-C. Fang. Linear programming with stochastic elements: an online approach. Computers and Mathematics with Applications, 33(9):61–82, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000667.

Greengard:1990:PVF

[GG90] L. Greengard and W. D.

Gropp. A parallel version of the fast multipole method. Computers and Mathematics with Applications, 20(7):63–71, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090349O.

Guillermo:1992:IIA

[GG92] L. Guillermo and M. Gómez.

Intertemporal issues associated with the control of macro-economic systems. Computers and Mathematics with Applications, 24(8– 9):77–98, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290189O.

Gelb:1997:RGP

[GG97] A. Gelb and D. Gottlieb. The resolution of the Gibbs phenomenon for “spliced” functions in one and two dimensions. Computers and Mathematics with Applications, 33(11):35–58, June 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000862.

Graesser:1992:QMQ

[GGB92] Arthur C. Graesser, Sallie E. Gordon, and Lawrence E. Brainerd. QUEST: a model of question answering. Computers and Mathematics with Applications, 23(6–9):733– 745, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901322.

Guellal:1997:NSL

[GGC97] S. Guellal, P. Grimalt, and Y. Cherruault. Numerical study of Lorenz’s equation by the Adomian method. Computers and Mathematics with Applications, 33(3): 25–29, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002349.

Glimm:1998:FTT

[GGG^{+}98] J. Glimm, M. J. Graham, J. Grove, X. L. Li, T. M. Smith, D. Tan, F. Tangerman, and Q. Zhang. Front tracking in two and three dimensions. Computers and Mathematics with Applications, 35(7):1–11, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000285.

Gavrea:1996:PLO

[GGK96] I. Gavrea, H. H. Gonska, and D. P. Kacsó. Positive linear operators with equidistant nodes. Computers and Mathematics with Applications, 32(8):23–32, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001630.

Gavrea:1997:VBP

[GGK97] I. Gavrea, H. H. Gonska, and D. P. Kacsó. Variation on Butzer’s problem: Characterization of the solutions. Computers and Mathematics with Applications, 34 (9):51–64, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001880.

Garey:1997:ASS

[GGS97] L. E. Garey, C. J. Gladwin, and R. E. Shaw. An approximate solution for a system with a symmetric matrix. Computers and Mathematics with Applications, 34 (12):25–31, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002307.

Ghosh:1990:DFB

[GH90] N. Ghosh and W. W. Hager. A derivative-free bracketing scheme for univariate minimization. Computers and Mathematics with Applications, 20(2):23–34, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090237E.

Graham:1997:SSS

[GH97] N. Graham and F. Harary. Some small sized spanning subgraphs of a hypercube. Computers and Mathematics with Applications, 34 (11):51–57, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002186.

Gutierrez:1998:RRS

[GH98] J. M. Gutiérrez and M. A. Hernández. Recurrence relations for the super-Halley method. Computers and Mathematics with Applications, 36(7):1–8, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001680.

Gunzburger:1996:AMC

[GHM96] M. D. Gunzburger, R. E. Hiromoto, and M. O. Mundt.

Analysis of a Monte Carlo boundary propagation method. Computers and Mathematics with Applications, 31(6): 61–70, March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000065.

Graf:1990:SHC

[GHPLZ90] T. Graf, P. Van Hentenryck, C. Pradelles-Lasserre, and L. Zimmer. Simulation of hybrid circuits in constraint logic programming. Computers and Mathematics with Applications, 20 (9–10):45–56, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090111V.

Gunzburger:1991:FEA

[GHS91] M. D. Gunzburger, L. Hou, and T. P. Svobodny. Finite element approximations of an optimal control problem associated with the scalar Ginzburg–Landau equation. Computers and Mathematics with Applications, 21 (2–3):123–131, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190090Q.

Guh:1996:FWA

[GHWL96] Yuh-Yuan Guh, Cheng-Chung Hon, Kuo-Ming Wang, and E. S. Lee. Fuzzy weighted average: a max-min paired elimination method. Computers and Mathematics with Applications, 32 (8):115–123, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600171X.

Gilbert:1991:CDN

[Gil91] William J. Gilbert. The complex dynamics of Newton’s method for a double root. Computers and Mathematics with Applications, 22 (10):115–119, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190198D.

Gilam:1993:TAP

[Gil93] A. Gilam. The technology of air power during the great war. Computers and Mathematics with Applications, 26 (6):149–152, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390124E.

Girosi:1992:SER

[Gir92] Federico Girosi. Some extensions of radial basis functions and their applications in artificial intelligence. Computers and Mathematics with Applications, 24 (12):61–80, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290172E.

Guram:1991:TDM

[GK91] G. S. Guram and M. Anwar Kamal. Three-dimensional micropolar flow near saddle and nodal points of attachment. Computers and Mathematics with Applications, 22(7):1–9, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901776.

Guo:1998:PEM

[GK98] X.-Z. Guo and I. N. Katz.

Performance enhancement of the multi- p preconditioner. Computers and Mathematics with Applications, 36(4): 1–8, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001369.

Galue:1999:SSG

[GKT99] L. Galué, S. Kalla, and V. K. Tuan. Some Sonine– Gegenbauer type integrals. Computers and Mathematics with Applications, 38 (3–4):73–83, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002072.

Grace:1993:OTS

[GL93] S. R. Grace and B. S. Lalli. Oscillation theorems for second order nonlinear functional differential equations with damping. Computers and Mathematics with Applications, 25(8):107–113, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390177W.

Guan:1994:SPN

[GL94] Zhi-Hong Guan and Yong-Qing Liu. The stability properties of nonlinear measure large scale systems with impulse effect. Computers and Mathematics with Applications, 28(9):89–99, November 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001812.

Gravvanis:1996:TDE

[GL96] G. A. Gravvanis and E. A. Lipitakis. A three-dimensional explicit preconditioned solver. Computers and Mathematics with Applications, 32(2): 111–131, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001083.

Glaister:1990:FDS

[Gla90a] P. Glaister. Flux-difference splitting for inviscid, real gases with non-equilibrium chemistry. Computers and Mathematics with Applications, 20(3):45–51, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090029J.

Glaister:1990:SSS

[Gla90b] P. Glaister. A similarity solution for a shock-reflection problem in isentropic gas dynamics. Computers and Mathematics with Applications, 20(1):67–75, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090071Q.

Glazebrook:1990:SHM

[Gla90c] J. F. Glazebrook. Strings, harmonic maps and hyperbolic systems. Computers and Mathematics with Applications, 19(8–9):117–125, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090269P.

Glaister:1991:EFD

[Gla91a] P. Glaister. An efficient finite difference scheme for threedimensional, non-equilibrium flows using operator splitting. Computers and Mathematics with Applications, 21(2–3):29–38, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190078I.

Glaister:1991:IFA

[Gla91b] P. Glaister. An isentropic flow algorithm with body fitted meshes. Computers and Mathematics with Applications, 22(12):47–53, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190146U.

Glaister:1991:NUM

[Gla91c] P. Glaister. A non-uniform mesh scheme for compressible flow. Computers and Mathematics with Applications, 21(4):39–62, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190246Z.

Glaister:1991:RSC

[Gla91d] P. Glaister. A Riemann solver for compressible flow of a real gas in a Lagrangian frame. Computers and Mathematics with Applications, 22(8):35–39, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900102.

Glaister:1991:TFD

[Gla91e] P. Glaister. A TVD finite difference scheme with nonuniform meshes and without upstream weighting. Computers and Mathematics with Applications, 22(3):45–58, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190069G.

Glaister:1992:ERS

[Gla92a] P. Glaister. An efficient Riemann solver for unsteady flows with non-ideal gases. Computers and Mathematics with Applications, 24(3): 77–93, August 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902165.

Glaister:1992:FDS

[Gla92b] P. Glaister. A finite difference scheme for steady, supersonic, two-dimensional, compressible flow of real gases. Computers and Mathematics with Applications, 24(4): 49–59, August 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192900064.

Glaister:1992:PSF

[Gla92c] P. Glaister. Prediction of supercritical flow in open channels. Computers and Mathematics with Applications, 24(7):69–75, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290155B.

Glaister:1992:RGF

[Gla92d] P. Glaister. Real gas flows in a duct. Computers and Mathematics with Applications, 24 (11):45–59, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290030L.

Glaister:1992:SSM

[Gla92e] P. Glaister. Similarity solutions for multi-component flows. Computers and Mathematics with Applications, 24 (10):77–88, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290020I.

Glaister:1993:ARS

[Gla93a] P. Glaister. An approximate Riemann solver for the H-system in generalized coordinates. Computers and Mathematics with Applications, 26(9):67–73, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390006H.

Glaister:1993:CSF

[Gla93b] P. Glaister. Cylindrically symmetric flows in a Lagrangian frame. Computers and Mathematics with Applications, 25(8):41–45, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390170Z.

Glaister:1993:FDSb

[Gla93c] P. Glaister. Flux difference splitting for hyperbolic systems of conservation laws with source terms. Computers and Mathematics with Applications, 26(7):79–96, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390052W.

Glaister:1993:FDSa

[Gla93d] P. Glaister. Flux difference splitting for the Euler equations in a general orthogonal curvilinear coordinate system with an application to axially symmetric flows. Computers and Mathematics with Applications, 26(1):75–89, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390087C.

Glaister:1993:SOA

[Gla93e]  P. Glaister.  Second or-

der accurate upwind difference schemes for scalar conservation laws with source terms. Computers and Mathematics with Applications, 25(4):65–73, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390249U.

Glaister:1994:AAA

[Gla94a]  P. Glaister.  An analy-

sis of arithmetic averaging in approximate Riemann solvers with an application to steady, supercritical flows. Computers and Mathematics with Applications, 28(8): 77–85, October 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001715.

Glaister:1994:ENM

[Gla94b] P. Glaister. An efficient numerical method for compressible flows of a real gas using arithmetic averaging. Computers and Mathematics with Applications, 28 (7):97–113, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001634.

Glaister:1994:ENSb

[Gla94c] P. Glaister. An efficient numerical scheme for the Euler equations. Computers and Mathematics with Applications, 27(8):91–100, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900698.

Glaister:1994:ENSa

[Gla94d] P. Glaister. An efficient numerical scheme for the twodimensional shallow water equations using arithmetic averaging. Computers and Mathematics with Applications, 27(1):97–117, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900086.

Glaister:1994:ESC

[Gla94e] P. Glaister. An efficient shock capturing scheme for the pseudo-unsteady equations representing steady inviscid flow. Computers and Mathematics with Applications, 28(7):71–76, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001618.

Glaister:1994:WFR

[Gla94f] P. Glaister. A weak formulation of Roe’s approximate Riemann solver applied to ‘barotropic’ flows. Computers and Mathematics with Applications, 27(8):87–90, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219490068X.

Glaister:1995:AAA

[Gla95a] P. Glaister. An analysis of arithmetic averaging in a Riemann solver for the one-dimensional Euler equations. Computers and Mathematics with Applications, 30 (7):103–112, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500129M.

Glaister:1995:CCE

[Gla95b]  P. Glaister.  A compari-

son of the computational efficiency of three averaging procedures in a numerical scheme for the shallow water equations. Computers and Mathematics with Applications, 30(7):31–35, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500123G.

Glaister:1995:CDE

[Gla95c]  P. Glaister.  A compari-

son of the different extensions of a weak formulation of an approximate Riemann solver for supercritical flows and their relationship to existing schemes. Computers and Mathematics with Applications, 29(12):27–38, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000554.

Glaister:1995:ESC

[Gla95d] P. Glaister. An efficient shock capturing scheme for two-dimensional, open channel, unsteady flows in a generalised coordinate system. Computers and Mathematics with Applications, 29(1):83–88, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002083.

Glaister:1995:SCS

[Gla95e] P. Glaister. Shock capturing for supercritical, free-surface flows in curvilinear channel geometries. Computers and Mathematics with Applications, 29(1):27–37, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400204X.

Glaister:1995:WFR

[Gla95f] P. Glaister. A weak formulation of Roe’s scheme for twodimensional, unsteady, compressible flows and steady, supersonic flows. Computers and Mathematics with Applications, 30(8):85–93, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500140T.

Glaister:1996:EDA

[Gla96a] P. Glaister. The efficiency of different averaging procedures in an algorithm for the Euler equations. Computers and Mathematics with Applications, 31(2):67–71, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001948.

Glaister:1996:EPS

[Gla96b] P. Glaister. The efficient prediction of shallow water flows. Part 1: Theory. Computers and Mathematics with Applications, 32 (12):129–143, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002131.

Glaister:1996:RBD

[Gla96c] P. Glaister. The relationship between the direct and weak formulations of a linearised Riemann solver for systems of conservation laws. Computers and Mathematics with Applications, 31(7):65–73, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000181.

Glaister:1997:AAP

[Gla97a] P. Glaister. An analysis of averaging procedures in a Riemann solver for compressible flows of a real gas. Computers and Mathematics with Applications, 33 (3):105–119, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002416.

Glaister:1997:EPS

[Gla97b] P. Glaister. The efficient prediction of shallow water flows — Part II: Application. Computers and Mathematics with Applications, 33(9):115–141, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000709.

Glaister:1998:AAP

[Gla98] P. Glaister. An analysis of averaging procedures in a Riemann solver for the two-dimensional Euler equations. Computers and Mathematics with Applications, 35 (8):63–77, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000455.

Glaister:1999:RSP

[Gla99] P. Glaister. Riemann solvers with primitive parameter vectors for two-dimensional compressible flows of a real gas. Computers and Mathematics with Applications, 37(2): 75–92, January 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198002545.

Gonska:1995:SCI

[GlZ95] H. H. Gonska and X. l. Zhou. The strong converse inequality for Bernstein–Kantorovich operators. Computers and Mathematics with Applications, 30(3–6):103–128, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000895.

Grad:1996:SLQ

[GM96] J. R. Grad and K. A. Morris.

Solving the linear quadratic optimal control problem for infinite-dimensional systems. Computers and Mathematics with Applications, 32 (9):99–119, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001800.

Gleyse:1999:ECN

[GM99] B. Gleyse and M. Moflih. Exact computation of the number of zeros of a real polynomial in the open unit disk by a determinant representation. Computers and Mathematics with Applications, 38 (11–12):257–263, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199003041.

Gupta:1994:GCT

[GMBS94] S. Sen Gupta, K. Mukhopadhyaya, B. B. Bhattacharya, and B. P. Sinha. Geometric classification of triangulations and their enumeration in a convex polygon. Computers and Mathematics with Applications, 27(7):99–115, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219490152X.

Guo:1990:MSM

[GMR90] L. Guo, D. A. Murio, and C. Roth. A mollified space marching finite differences algorithm for the inverse heat conduction problem with slab symmetry. Computers and Mathematics with Applications, 19(7):75–89, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090196Q.

Grote:1998:ELD

[GMS98] K. Grote and R. Meyer-Spasche. Euler-like discrete models of the logistic differential equation. Computers and Mathematics with Applications, 36(10–12):211–225, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800229.

Grotschel:1996:OPP

[GMW96] M. Grötschel, A. Martin, and R. Weismantel. Optimum path packing on wheels: the consecutive case. Computers and Mathematics with Applications, 31(11):23–35, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000582.

Giraldo:1999:SAE

[GN99] F. X. Giraldo and B. Neta.

Stability analysis for Eulerian and semi-Lagrangian finiteelement formulation of the advection-diffusion equation. Computers and Mathematics with Applications, 38(2): 97–112, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001856.

Germano:1995:CMD

[GNR95] B. Germano, P. Natalini, and P. E. Ricci. Computing the moments of the density of zeros for orthogonal polynomials. Computers and Mathematics with Applications, 30 (3–6):69–81, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000879.

Goh:1994:GAS

[Goh94] B. S. Goh. Global attractivity and stability of a scalar nonlinear difference equation. Computers and Mathematics with Applications, 28 (1–3):101–107, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194000972.

Goldstein:1990:PIM

[Gol90] C. I. Goldstein. Preconditioned iterative methods for convection diffusion and related boundary value problems. Computers and Mathematics with Applications, 19(8–9):11–29, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090261H.

Golik:1994:OIG

[Gol94] W. L. Golik. Order-increasing grid adaption for Runge– Kutta methods applied to two-point boundary value problems. Computers and Mathematics with Applications, 27(4):59–75, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900558.

Garcia-Ortiz:1993:ASC

[GOWR^{+}93] A. Garcia-Ortiz, J. Wootton, E. Y. Rodin, S. M. Amin, M. Meusey, C. Ruan, A. Y. Wu, P. De, and J. Revetta. Application of semantic control to a class of pursuer-evader problems. Computers and Mathematics with Applications, 26 (6):97–124, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390121B.

Gyongy:1990:ASD

[GP90] I. Gyöngy and T. Pröhle. On the approximation of stochastic differential equation and on Stroock–Varadhan’s support theorem. Computers and Mathematics with Applications, 19(1):65–70, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090082U.

Gunzburger:1991:PSS

[GP91] Max D. Gunzburger and Janet S. Peterson. Predictor and steplength selection in continuation methods for the Navier–Stokes equations. Computers and Mathematics with Applications, 22(8):73–81, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190015V.

Ghose:1993:DRS

[GP93] D. Ghose and U. R. Prasad.

Determination of rational strategies for players in twotarget games. Computers and Mathematics with Applications, 26(6):1–11, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390113A.

Gonzalez-Pinto:1992:TUS

[GPCGV92] S. González-Pinto, L. Casasús, and P. González-Vera. Two uniform schemes for the singularly perturbed Riccati equation. Computers and Mathematics with Applications, 23(1):75–85, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290082S.

Guellal:1990:MMA

[GPCSL90] S. Guellal, C. J. Papapanayotou, Y. Cherruault, and J. Sainte-Laudy. A mathematical model of the allergentriggered human basophil receptivity. Computers and Mathematics with Applications, 20(4–6):73–77, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090316C.

Gonzalez-Pinto:1994:ISG

[GPGCM94] S. González-Pinto, C. González-Concepción, and J. I. Montijano. Iterative schemes for Gauss methods. Computers and Mathematics with Applications, 27(7):67–81, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901503.

Gunzburger:1999:OBD

[GPK99] M. D. Gunzburger, J. S. Peterson, and H. Kwon. An optimization based domain decomposition method for partial differential equations. Computers and Mathematics with Applications, 37(10): 77–93, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001273.

Galperin:1993:AGO

[GPZ93] Efim A. Galperin, Zhongxiong Pan, and Quan Zheng. Application of global optimization to implicit solution of partial differential equations. Computers and Mathematics with Applications, 25 (10–11):119–124, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902876.

Grace:1994:OCF

[Gra94] S. R. Grace. Oscillation of certain functional differential equations. Computers and Mathematics with Applications, 27(4):91–98, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900582.

Grigorescu:1996:TFH

[GRC^{+}96] C. E. A. Grigorescu, K. B. Radev, V. Chesaru, T. Necsoiu, and I. Pricop. Thermal fluxes from the human body. Computers and Mathematics with Applications, 32 (11):47–55, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196001964.

Greenspan:1990:CDF

[Gre90a] D. Greenspan. Conservative difference formulations of Calogero and Toda Hamiltonian systems. Computers and Mathematics with Applications, 19(2):91–95, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190900129.

Greenspan:1990:SSC

[Gre90b] D. Greenspan. Supercomputer simulation of colliding microdrops of water. Computers and Mathematics with Applications, 19(7):91–97, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090197R.

Greenspan:1990:SSA

[Gre90c] D. Greenspan. Supercomputer studies in absolute minimization. Computers and Mathematics with Applications, 20(11):47–52, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902189.

Greenspan:1991:EQQ

[Gre91a] Donald Greenspan. The error in quasi-quantum mechanical calculations. Computers and Mathematics with Applications, 22(2):73–79, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900434.

Greenspan:1991:TDF

[Gre91b] Donald Greenspan. Threedimensional fluid microdrop interaction in the nucleation of raindrops. Computers and Mathematics with Applications, 22(8):83–89, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190016W.

Greenspan:1992:PMG

[Gre92a] Donald Greenspan. A particle model of ground state H_{2}. Computers and Mathematics with Applications, 24(3):45–48, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902132.

Greenspan:1992:SRK

[Gre92b] Donald Greenspan. Studies in rapid kinetic reactions by quasi-quantum mechanical, conservative methodology. Computers and Mathematics with Applications, 24 (11):11–15, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290027F.

Greenspan:1993:CSS

[Gre93a] D. Greenspan. Computer study of a semiclassical model of the dynamical water molecule. Computers and Mathematics with Applications, 26(8):13–20, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390327R.

Greenspan:1993:EAN

[Gre93b] Donald Greenspan. Electron attraction and Newtonian methodology for approximating quantum mechanical phenomena. Computers and Mathematics with Applications, 25(6):75–83, March 1993. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390301B.

Greenspan:1994:CDE

[Gre94]  D. Greenspan.  Conser-

vative difference equations. Computers and Mathematics with Applications, 28 (1–3):109–114, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194000980.

Greenspan:1995:CCC

[Gre95] D. Greenspan. Completely conservative, covariant numerical methodology. Computers and Mathematics with Applications, 29(4):37–43, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219400236E.

Greenspan:1998:DDE

[Gre98a] D. Greenspan. Dynamical difference equations. Computers and Mathematics with Applications, 36(10–12):49–58, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800084.

Greenspan:1998:DSS

[Gre98b]  D. Greenspan.  Dynami-

cal simulation of the simplest hydrides. Computers and Mathematics with Applications, 35(10):55–62, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000728.

Greenspan:1999:EAC

[Gre99] D. Greenspan. Electron attraction and chemical bonding. Computers and Mathematics with Applications, 38 (11–12):217–227, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199003004.

Gomes-Ruggiero:1996:NSL

[GRKM96] M. A. Gomes-Ruggiero, D. N. Kozakevich, and J. M. Martínez. A numerical study on large-scale nonlinear solvers. Computers and Mathematics with Applications, 32(3):1–13, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001095.

Groetsch:1995:RPI

[Gro95] C. W. Groetsch. Regularized product integration for Hadamard finite part integrals. Computers and Mathematics with Applications, 30(3–6):129–135, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000909.

Grycko:1994:EVT

[Gry94] E. Grycko. Extreme values and tests for randomness. Computers and Mathematics with Applications, 28(7):47–50, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001596.

Galligani:1996:SMC

[GRZ96] E. Galligani, V. Ruggiero, and L. Zanni. Splitting methods for constrained quadratic programs in data analysis. Computers and Mathematics with Applications, 32(5): 1–9, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001319.

Garey:1991:ASS

[GS91] L. E. Garey and R. E. Shaw.

Algorithms for the solution of second order Volterra integro-differential equations. Computers and Mathematics with Applications, 22(3): 27–34, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190067E.

Garey:1993:NMS

[GS93] L. E. Garey and R. E. Shaw. Numerical methods for singular nonlinear Volterra type equations with two point boundary conditions. Computers and Mathematics with Applications, 26 (1):53–61, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390085A.

Genuit:1995:PCE

[GS95] M. Genuit and G. Schöttler. A problem of L. L. Campbell on the equivalence of the Kramer and Shannon sampling theorems. Computers and Mathematics with Applications, 30(3–6):433– 443, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001182.

Gockenbach:1996:HCL

[GS96a] M. S. Gockenbach and W. W. Symes. Hilbert class library: a library of abstract C++ classes for optimization and inversion. Computers and Mathematics with Applications, 32(6):1–13, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600140X.

Greenberg:1996:SLC

[GS96b] R. I. Greenberg and Jau-Der Shih. Single-layer channel routing and placement with single-sided nets. Computers and Mathematics with Applications, 32(4):1–7, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001186.

Gupta:1996:GLD

[GS96c] A. K. Gupta and D. Song.

Generalized Liouville distribution. Computers and Mathematics with Applications, 32(2):103–109, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001071.

Gergel:1999:SPA

[GS99] V. P. Gergel and Ya. D.

Sergeyev. Sequential and parallel algorithms for global minimizing functions with Lipschitzian derivatives. Computers and Mathematics with Applications, 37(4–5):163– 179, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900067X.

Gori:1998:PSN

[GSC98] L. Gori, E. Santi, and M. G. Cimoroni. Projector-splines in the numerical solution of inetgro-differential equations. Computers and Mathematics with Applications, 35(5):107–116, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000091.

Ganesan:1991:SDL

[GSM91] Sekar Ganesan, Leang S. Shieh, and Mohamad M. Mehio. Sequential design of linear quadratic state regulators with prescribed eigenvalues and specified relative stability. Computers and Mathematics with Applications, 21(4):1–10, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190241U.

Gonchenko:1997:BSS

[GSST97] S. V. Gonchenko, L. P. Shil’nikov, O. V. Sten’kin, and D. V. Turaev. Bifurcations of systems with structurally unstable homoclinic orbits and moduli of ω-equivalence. Computers and Mathematics with Applications, 34(2–4):111–142, July/August 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001211.

Gonchenko:1997:QHT

[GST97] S. V. Gonchenko, L. P.

Shil’nikov, and D. V. Turaev. Quasiattractors and homoclinic tangencies. Computers and Mathematics with Applications, 34(2–4):195–227, July/August 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001247.

Garnett:1992:ERA

[GT92] Lucy Garnett and Abdullah U. Tansel. Equivalence of the relational algebra and calculus for nested relations. Computers and Mathematics with Applications, 23(10):3–25, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290050R.

Gao:1995:NFO

[GT95] Yi-Tian Gao and Bo Tian.

New family of overturning soliton solutions for a typical breaking soliton equation. Computers and Mathematics with Applications, 30 (12):97–100, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500176Y.

Gao:1997:GTM

[GT97a] Yi-Tian Gao and Bo Tian. Generalized tanh method with symbolic computation and generalized shallow water wave equation. Computers and Mathematics with Applications, 33 (4):115–118, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000114.

Gao:1997:USC

[GT97b] Yi-Tian Gao and Bo Tian. Using symbolic computation to exactly solve for the Bogoyavlenskii’s generalized breaking soliton equation. Computers and Mathematics with Applications, 33(4):35–37, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000047.

Graef:1999:OTD

[GT99] J. R. Graef and E. Thandapani. Oscillation of twodimensional difference systems. Computers and Mathematics with Applications, 38 (7–8):157–165, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002461.

Guelman:1993:CSP

[Gue93] M. Guelman. Control strategies in a planar pursuit evasion game with energy constraints. Computers and Mathematics with Applications, 26(6):33–41, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390115C.

Gundem:1999:NOM

[Gün99] T. I. Gündem. Near optimal multiple choice index selection for relational databases. Computers and Mathematics with Applications, 37 (2):111–120, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002569.

Gupta:1994:FSE

[Gup94]  S. M. Gupta.  Finite

source Erlang based queueing systems: complementarity, equivalence and their implications. Computers and Mathematics with Applications, 28(6):57–74, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001529.

Gurarie:1991:MAI

[Gur91]  David Gurarie.  Multipa-

rameter averaging and inverse spectral problem for anharmonic oscillators in R^{n}. Computers and Mathematics with Applications, 22(4– 5):81–92, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190133O.

Gustafson:1997:OSS

[Gus97]  K. Gustafson.  Operator

spectral states. Computers and Mathematics with Applications, 34(5–6):467– 508, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001508.

Gutman:1991:EDC

[GW91] S. Gutman and L. W. White. The estimation of L^{∞} diffusion coefficients in the parabolic equations. Computers and Mathematics with Applications, 22(4–5):125– 136, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190138T.

Goh:1997:ZDT

[GW97] W. M. Y. Goh and J. Wimp.

The zero distribution of the Tricomi–Carlitz polynomials. Computers and Mathematics with Applications, 33(1– 2):119–127, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002246.

Galperin:1999:RUB

[GW99] E. A. Galperin and M. M.

Wiecek. Retrieval and use of the balance set in multiobjective global optimization. Computers and Mathematics with Applications, 37(4– 5):111–123, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000632.

Guh:1996:HCA

[GWL96] Yuh-Yuan Guh, Kuo-Ming Wang, and E. S. Lee. Hierarchies consistency analysis for nontransitive problems. Computers and Mathematics with Applications, 32 (9):67–77, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001782.

Guo:1993:VI

[GY93] Jong-Shenq Guo and Jen-Chih Yao. The variational inequalities. Computers and Mathematics with Applications, 25(3):99–105, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390148O.

Gyires:1996:EPV

[Gyi96] B. Gyires. Elementary proof for a Van der Waerden’s conjecture and related theorems. Computers and Mathematics with Applications, 31 (10):7–21, May 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600048X.

Gyongy:1990:ASP

[Gyö90a] I. Gyöngy. The approximation of stochastic partial differential equations and applications in nonlinear filtering. Computers and Mathematics with Applications, 19(1):47–63, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090081T.

Gyori:1990:SMA

[Gyö90b]  I. Györi.  Some math-

ematical aspects of modelling cell population dynamics. Computers and Mathematics with Applications, 20 (4–6):127–138, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090322B.

Galperin:1991:IGO

[GZ91] Efim A. Galperin and Quan Zheng. Integral global optimization method for nonlinear games. Computers and Mathematics with Applications, 21(6–7):145–159, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901695.

Galperin:1993:SCP

[GZ93] Efim A. Galperin and Quan Zheng. Solution and control of PDE via global optimization methods. Computers and Mathematics with Applications, 25(10–11):103– 118, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902865.

Gonska:1995:LSF

[GZ95] H. H. Gonska and D.-X.

Zhou. Local smoothness of functions and Bernstein– Durrmeyer operators. Computers and Mathematics with Applications, 30(3–6):83–101, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000887.

Ha:1990:ENS

[Ha90] S. N. Ha. Experimental numerical studies on a supercomputer of natural convection in an enclosure with localized heating. Computers and Mathematics with Applications, 20(12):1–17, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090159H.

Hakim:1999:ESS

[Hak99] A. Hakim. Existence of slow steady flows of viscoelastic fluids of White–Metzner type. Computers and Mathematics with Applications, 38(11– 12):239–247, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199003028.

Halilov:1990:ANL

[Hal90] A. Halilov. Asymptotic normality of lag window estimates. Computers and Mathematics with Applications, 19(1):71–74, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090083V.

Hanson:1992:CGS

[Han92] James N. Hanson. The computer generated symbolic approximations to systems of nonlinear ODE’s by matrix annihilation and the Newton– Kantorovich method. Computers and Mathematics with Applications, 24(3):29–35, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219290211Y.

Hosseini:1991:TMB

[HAO91] M. Hosseini, Ali Abadi, and E. L. Ortiz. A Tau method based on non-uniform spacetime elements for the numerical simulation of solitons. Computers and Mathematics with Applications, 22 (9):7–19, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190202F.

Hardy:1990:TAM

[Har90] R. L. Hardy. Theory and applications of the multiquadricbiharmonic method 20 years of discovery 1968-1988. Computers and Mathematics with Applications, 19(8–9):163–

208, ????  1990.  CO-

DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090272L.

Hardy:1992:CDE

[Har92a] Rolland L. Hardy. Comments on the discovery and evolution of the multiquadric method. Computers and Mathematics with Applications, 24(12):xi–xii, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/089812219290165E.

Hardy:1992:CMM

[Har92b] Rolland L. Hardy. A contribution of the multiquadric method: Interpolation of potential inside the earth. Computers and Mathematics with Applications, 24 (12):81–97, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290173F.

Hartley:1992:URT

[Har92c] Roger T. Hartley. A uniform representation for time and space and their mutual constraints. Computers and Mathematics with Applications, 23(6–9):441– 457, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290117Z.

Harison:1996:DEC

[Har96] V. Harison. Drift estimation of a certain class of diffusion processes from discrete observation. Computers and Mathematics with Applications, 31(6):121–133, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000119.

Hashimoto:1991:NAS

[Has91] Naoki Hashimoto. Numerical algebraic surfaces. Computers and Mathematics with Applications, 21(2–3):111– 114, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190088L.

Hassenpflug:1993:CN

[Has93a] W. C. Hassenpflug. Convolution number. Computers and Mathematics with Applications, 26(11):3–135, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122193900665.

Hassenpflug:1993:MTN

[Has93b] W. C. Hassenpflug. Matrix tensor notation. Part I. Rectilinear orthogonal coordinates. Computers and Mathematics with Applications, 26(3):55–93, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390110H.

Hassenpflug:1993:RNP

[Has93c] W. C. Hassenpflug. Revised notation for partial derivatives. Computers and Mathematics with Applications, 26(3):95–105, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193901118.

Hassenpflug:1995:MTN

[Has95] W. C. Hassenpflug. Matrix tensor notation. Part II. Skew and curved coordinates. Computers and Mathematics with Applications, 29 (11):1–103, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000509.

Hassenpflug:1997:EIS

[Has97] W. C. Hassenpflug. Elliptic integrals and the Schwarz– Christoffel transformation.

Computers and Mathematics with Applications, 33(12): 15–114, June 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000916.

Hassenpelug:1998:FS

[Has98] W. C. Hassenpflug. Freestreamlines. Computers and Mathematics with Applications, 36(1):69–129, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001102.

Hautamaki:1992:CSA

[Hau92] Antti Hautamäki. A conceptual space approach to semantic networks. Computers and Mathematics with Applications, 23(6–9):517– 525, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290121W.

Haws:1991:SGF

[Haw91] LaDawn Haws. Symmetric Green’s functions for certain hyperbolic problems. Computers and Mathematics with Applications, 21(5): 65–78, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190216Q.

Huang:1999:PIP

[HB99] Nan-Jing Huang and Min-Ru Bai. A perturbed iterative procedure for multivalued pseudo-contractive mappings and multivalued accretive mappings in Banach spaces. Computers and Mathematics with Applications, 37(6):7–15, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000723.

Hanson:1990:CGS

[HBB90] J. H. Hanson, A. C. Benander, and B. A. Benander. The computer generated symbolic solution of a system of linear first order differential equations. Computers and Mathematics with Applications, 19(7):7–12, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090189Q.

hCong:1996:EPT

[hC96a] N. h. Cong. Explicit parallel two-step Runge–Kutta– Nyström methods. Computers and Mathematics with Applications, 32(3):119–130, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122196001174.

hCong:1996:ESR

[hC96b] N. h. Cong. Explicit symmetric Runge–Kutta–Nyström methods for parallel computers. Computers and Mathematics with Applications, 31 (2):111–121, 123, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001980.

Hwang:1998:PMS

[HC98] H.-C. Hwang and S. Y.

Chang. Parallel machines scheduling with machine shutdowns. Computers and Mathematics with Applications, 36(3):21–31, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001266.

hCong:1998:NES

[hCPW98] N. h. Cong, H. Podhaisky, and R. Weiner. Numerical experiments with some explicit pseudo two-step RK methods on a shared memory computer. Computers and Mathematics with Applications, 36(2):107–116, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001217.

He:1995:FPC

[He95a] M. He. The Faber polynomials for circular lunes. Computers and Mathematics with Applications, 30(3– 6):307–315, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001093.

He:1995:SCB

[He95b] T.-X. He. Shape criteria of Bernstein–Bézier polynomials over simplexes. Computers and Mathematics with Applications, 30(3–6):317– 333, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001107.

Hegedus:1991:GCDa

[Heg91a] Cs. J. Hegedüs. Generating conjugate directions for arbitrary matrices by matrix equations i. Computers and Mathematics with Applications, 21(1):71–85, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190232S.

Hegedus:1991:GCDb

[Heg91b] Cs. J. Hegedüs. Generating conjugate directions for arbitrary matrices by matrix equations II. Computers and Mathematics with Applications, 21(1):87–94, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190233T.

Helmuth:1996:GSC

[Hel96]  M. Helmuth.  Granular

synthesis composition with StochGran and max. Computers and Mathematics with Applications, 32(1):57–74, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000879.

Hendler:1992:MPM

[Hen92] James A. Hendler. Massivelyparallel marker-passing in semantic networks. Computers and Mathematics with Applications, 23(2–5):277– 291, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901447.

Henderson:1994:SSD

[Hen94]  J. Henderson.  Smooth-

ness of solutions for delaydifference equations. Computers and Mathematics with Applications, 28(1–3):127– 129, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001006.

Hennefeld:1997:PPO

[Hen97] J. Hennefeld. Parallel processing and OOP as an analogy for the discovery of certain mathematical proofs. Computers and Mathematics with Applications, 33(6): 91–97, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000369.

Hevia:1997:TVG

[Hev97] H. Hevia. Transitional values of graphs. Computers and Mathematics with Applications, 34(11):65–79, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122197002204.

Hsieh:1998:ASS

[HG98] T.-H. Hsieh and Z. Gajic. An algorithm for solving the singularly perturbed H_{∞} algebraic Riccati equation. Computers and Mathematics with Applications, 36 (5):69–77, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001515.

Hwang:1999:FBA

[HGG99] Chyi Hwang, Jia-Chyu Guo, and Tong-Yi Guo. An FFTbased algorithm for 2D power series expansions. Computers and Mathematics with Applications, 37(10):19–27, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001224.

He:1993:VGC

[HGK93] M. He, D. B. Goldgof, and C. Kambhamettu. Variation of Gaussian curvature under conformal mapping and its application. Computers and Mathematics with Applications, 26(1):63–74, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390086B.

Hon:1996:FMA

[HGWL96] Cheng-Chuang Hon, Yuh-Yuan Guh, Kou-Ming Wang, and E. S. Lee. Fuzzy multiple attributes and multiple hierarchical decision making. Computers and Mathematics with Applications, 32 (12):109–119, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002118.

Hwang:1993:IAD

[HH93] T. Hwang and S. Y. Hwang. An improved algorithm for davida and Cowles’s decoding method. Computers and Mathematics with Applications, 26(7):59–66, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193900506.

Hill:1998:NMP

[HH98a] S. Hill and L. D. Humphreys. Numerical mountain pass periodic solutions of a nonlinear spring equation. Computers and Mathematics with Applications, 35(12):59–67, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000960.

How:1998:TMC

[HH98b] Sey-Ping How and Tsan-Hui Hsu. Transient mixed convection in a partially divided enclosure. Computers and Mathematics with Applications, 36(8):95–115, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001862.

Halanay:1994:ASS

[HI94] A. Halanay and V. Ionescu.

Anticausal stabilizing solution to discrete reverse-time Riccati equation. Computers and Mathematics with Applications, 28(1–3):115– 126, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194000999.

Halanay:1998:TVD

[HI98] A. Halanay and V. Ionescu.

Time-varying discrete Hamiltonian systems. Computers and Mathematics with Applications, 36(10–12):307–326, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800321.

Hirabayashi:1999:DTP

[HIIN99] R. Hirabayashi, Y. T. Ikebe, K. Iwamura, and T. Nakayama. Drawing a tree on parallel lines. Computers and Mathematics with Applications, 37 (11–12):171–176, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001558.

Hinestroza:1995:ITC

[Hin95] D. Hinestroza. Identification of transmissivity coefficients by mollification techniques. Part II: n-dimensional elliptic problems. Computers and Mathematics with Applications, 30(1):11–30, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000646.

Hentzel:1991:DPM

[HJ91] Irvin Roy Hentzel and David Pokrass Jacobs. A dynamic programming method for building free algebras. Computers and Mathematics with Applications, 22(12):61–66, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190148W.

Herceg:1996:CRF

[HK96] D. Herceg and N. Krejić. Convergence results for fixed point iterations in R. Computers and Mathematics with Applications, 31(2):7–10, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001883.

Henderson:1998:FBV

[HK98] J. Henderson and E. R. Kaufmann. Focal boundary value problems for singular difference equations. Computers and Mathematics with Applications, 36(10–12):1–10, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800047.

Huynh:1990:EDS

[HL90] T. Huynh and C. Lassez. An expert decision-support system for option-based investment strategies. Computers and Mathematics with Applications, 20(9–10):1–14, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090108V.

Haase:1996:EPS

[HL96] G. Haase and U. Langer.

The efficient parallel solution of PDEs. Computers and Mathematics with Applications, 31(4–5):151– 159, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002278.

Hicks:1999:SHC

[HL99] D. L. Hicks and L. M.

Liebrock. SPH hydrocodes can be stabilized with shapeshifting. Computers and Mathematics with Applications, 38(5–6):1–16, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002102.

Hwang:1991:IFB

[HLS91] Chyi Hwang, Ming-Jeng Lu, and Leang S. Shieh. Improved FFT-based numerical inversion of Laplace transforms via fast Hartley transform algorithm. Computers and Mathematics with Applications, 22(1):13–24, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190021U.

Huang:1999:GSV

[HLTB99] Nan-Jing Huang, Ya-Ping Liu, Ya-Yong Tang, and Min-Ru Bai. The generalized set-valued strongly nonlinear implicit variational inequalities. Computers and Mathematics with Applications, 37 (10):29–36, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001236.

Hou:1991:SNS

[HM91] Lisheng Hou and Hassan Manouzi. Solving the Navier– Stokes equations by Laplacians: an optimal control approach. Computers and Mathematics with Applications, 22(11):59–67, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900342.

Hinestroza:1993:ITC

[HM93a] Doris Hinestroza and Diego A.

Murio. Identification of transmissivity coefficients by mollification techniques. Part I: One-dimensional elliptic and parabolic problems. Computers and Mathematics with Applications, 25(8):59–79, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390172R.

Hinestroza:1993:RTH

[HM93b] Doris Hinestroza and Diego A.

Murio. Recovery of the transient heat transfer coefficient in the nonlinear boundary value problem for the heat equation. Computers and Mathematics with Applications, 25(5):101–111, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902038.

Hu:1994:DPG

[HM94] T. C. Hu and J. D. Morgenthaler. Dynamic programming and graph optimization problems. Computers and Mathematics with Applications, 27(9–10):53–58, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901236.

Hurley:1995:NSU

[HM95] W. Hurley and L. Mc-Donough. A note on simulating unbiased heterogeneous expectations. Computers and Mathematics with Applications, 30(8):29–32, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500134K.

Hayashi:1996:FAS

[HM96] H. Hayashi and T. Mitsui. Further analysis of stability for Lambert’s method based on Euler’s rule. Computers and Mathematics with Applications, 31(8):31–36, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000284.

He:1999:CBI

[HM99] Yinnian He and R. M. M. Mattheij. The coupling of boundary integral and finite element methods for the exterior steady Oseen problem. Computers and Mathematics with Applications, 37(2):27–40, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800251X.

Hinestroza:1999:RTN

[HMZ99] D. Hinestroza, D. A. Murio, and S. Zhan. Regularization techniques for nonlinear problems. Computers and Mathematics with Applications, 37(10):145–159, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001315.

Hogarth:1990:CSF

[HNS^{+}90] W. L. Hogarth, B. J. Noye, J. Stagnitti, J. Y. Parlange, and G. Bolt. A comparative study of finite difference methods for solving the onedimensional transport equation with an initial-boundary value discontinuity. Computers and Mathematics with Applications, 20(11):67–82, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090220E.

Harary:1997:ICG

[HO97] F. Harary and O. R. Oellermann. The irregularity cost of a graph. Computers and Mathematics with Applications, 34(11):59–63, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122197002198.

Holcombe:1990:TFD

[Hol90] W. M. L. Holcombe. Towards a formal description of intracellular biochemical organization. Computers and Mathematics with Applications, 20(4–6):107–115, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090320J.

Honner:1999:HWS

[Hon99] M. Honner. Heat waves simulation. Computers and Mathematics with Applications, 38(9–10):233–243, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002783.

Henderson:1994:DTO

[HP94a] J. Henderson and A. Peterson. Disconjugacy for a third order linear difference equation. Computers and Mathematics with Applications, 28 (1–3):131–139, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001014.

Hofmann:1994:SWN

[HP94b] N. Hofmann and E. Platen. Stability of weak numerical schemes for stochastic differential equations. Computers and Mathematics with Applications, 28(10–12):45–57, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001855.

Hung:1994:ASD

[HP94c] C. Y. Hung and B. Parhami.

An approximate sign detection method for residue numbers and its application to RNS division. Computers and Mathematics with Applications, 27(4):23–35, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900523.

Hamalainen:1996:DLC

[HP96] T. Hämäläinen and S. Pohjolainen. The design of linear controllers with symbolic algebra. Computers and Mathematics with Applications, 32 (7):105–116, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001599.

Henderson:1999:CEL

[HP99] J. Henderson and K. R.

Prasad. Comparison of eigenvalues for Lidstone boundary value problems on a measure chain (preprint). Computers and Mathematics with Applications, 38(11–12):55– 62, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002849.

Hoover:1994:MDS

[HPH^{+}94] W. G. Hoover, T. G. Pierce, C. G. Hoover, J. O. Shugart, C. M. Stein, and A. L. Edwards. Molecular dynamics, smoothed-particle applied mechanics, and irreversibility. Computers and Mathematics with Applications, 28(10–12):155–174, November/December 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400191X.

Hansen:1999:NSU

[HR99] E. B. Hansen and H. Rasmussen. A numerical study of unstable Hele–Shaw flow. Computers and Mathematics with Applications, 38(5– 6):217–230, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900228X.

Hosea:1994:ECM

[HS94] M. E. Hosea and L. F. Shampine. Efficiency comparisons of methods for integrating ODEs. Computers and Mathematics with Applications, 28(6):45–55, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001510.

Hai:1995:CPC

[HS95] N. T. Hai and H. M. Srivastava. The convergence problem of certain multiple Mellin–Barnes contour integrals representing Hfunctions in several variables. Computers and Mathematics with Applications, 29(6): 17–25, March 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500003H.

Hyman:1997:NDD

[HS97] J. M. Hyman and M. Shashkov.

Natural discretizations for the divergence, gradient, and curl on logically rectangular grids. Computers and Mathematics with Applications, 33(4):81–104, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000096.

Hyman:1998:ABC

[HS98] J. M. Hyman and M. Shashkov.

Approximation of boundary conditions for mimetic finitedifference methods. Computers and Mathematics with Applications, 36(5):79–99, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001527.

Harris:1990:CGI

[HSH90a] L. F. Harris, M. R. Sullivan, and D. F. Hickok. Conservation of genetic information between regulatory protein DNA binding alpha helices and their cognate operator sites: a simple code for sitespecific recognition. Computers and Mathematics with Applications, 20(4–6):1–23, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190903117.

Harris:1990:GSH

[HSH90b] L. F. Harris, M. R. Sullivan, and D. F. Hickok. Genetic sequences of hormone response elements share similarity with predicted alpha helices within DNA binding domains of steroid receptor proteins: a basis for site-specific recognition. Computers and Mathematics with Applications, 20(4–6):25–48, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190903128.

Hsueh:1991:UAS

[Hsu91] Yuang-Cheh Hsueh. A unifying approach to the structures of the stable matching problems. Computers and Mathematics with Applications, 22(6):13–27, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190002L.

Hsueh:1996:RBV

[Hsu96] Yuang-Chen Hsueh. Relationbased variations of the discrete radon transform. Computers and Mathematics with Applications, 31(3):119–131, February 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122195002081.

Hou:1992:AFT

[HTT92] Ting-Wei Hou, S. R. Tsai, and L. M. Tseng. Adaptive and fault-tolerant routing algorithms for high performance 2D torus interconnection network. Computers and Mathematics with Applications, 23(1):3–15, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290075S.

Hu:1995:PCS

[Hu95]  Y.-K. Hu.  Positive and

copositive spline approximation in L_{p}[0, 1]. Computers and Mathematics with Applications, 30(3–6):137– 146, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000917.

Huang:1991:ESS

[Hua91] Cheng-Gui Huang. The explicit solutions of the Sturm– Liouville problem. Computers and Mathematics with Applications, 22(11):29–37, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190031X.

Huang:1993:EGV

[Hua93] T. C. Huang. Extended generalized variational inequalities. Computers and Mathematics with Applications, 26(8):21–30, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390328S.

Huang:1995:CBU

[Hua95] Lan Chieh Huang. Conservative bicharacteristic upwind schemes for hyperbolic conservation laws II.

Computers and Mathematics with Applications, 29 (3):91–107, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400230I.

Huang:1997:NCR

[Hua97] Zhenyu Huang. A new convergence result for fixed-point iteration in bounded interval of R^{n}. Computers and Mathematics with Applications, 34 (12):33–36, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002319.

Huang:1998:MIT

[Hua98a] Nan-Jing Huang. Mann and Ishikawa type perturbed iterative algorithms for generalized nonlinear implicit quasivariational inclusions. Computers and Mathematics with Applications, 35(10):1–7, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000662.

Huang:1998:NCG

[Hua98b] Nan-Jing Huang. A new completely general class of variational inclusions with noncompact valued mappings. Computers and Mathematics with Applications, 35 (10):9–14, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000674.

Huang:1998:AFP

[Hua98c] Zhenyu Huang. Approximating fixed points of Φhemicontractive mappings by the Ishikawa iteration process with errors in uniformly smooth Banach spaces. Computers and Mathematics with Applications, 36(2):13–21, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001126.

Huang:1999:MII

[Hua99] Zhenyu Huang. Mann and Ishikawa iterations with errors for asymptotically nonexpansive mappings. Computers and Mathematics with Applications, 37(3):1–7, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000401.

Hujter:1996:IMS

[Huj96] M. Hujter. Improving a method of search for solving polynomial equations. Computers and Mathematics with Applications, 31(4– 5):187–189, February/March 1996. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002316.

Hujter:1999:SNP

[Huj99] M. Hujter. Some numerical problems in discrete geometry. Computers and Mathematics with Applications, 38(9–10):175–178, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002722.

Huntingford:1995:ESO

[Hun95]  C. Huntingford.  An ex-

act solution to the one-phase zero-surface-tension Hele– Shaw free-boundary problem. Computers and Mathematics with Applications, 29 (10):45–50, May 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500044Y.

Howland:1992:RTS

[HV92] James L. Howland and Rémi Vaillancourt. Rational transformation from Schur to Jordan form. Computers and Mathematics with Applications, 23(11):91–101, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290072P.

Hwang:1990:ECS

[HW90] T.-W. Hwang and C.-A. Wang. Existence and classification of solutions for a problem of surface-tension driven flows in a slot with an insulated bottom. Computers and Mathematics with Applications, 19(12):1–8, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090245F.

Hansen:1993:BLM

[HW93a] E. R. Hansen and G. W. Walster. Bounds for Lagrange multipliers and optimal points. Computers and Mathematics with Applications, 25(10–11):59– 69, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390282Z.

Hansen:1993:NEO

[HW93b] E. R. Hansen and G. W. Walster. Nonlinear equations and optimization. Computers and Mathematics with Applications, 25(10–11):125– 145, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902887.

Huang:1996:CCP

[HW96] Lan Chieh Huang and Ya Dan Wu. The componentconsistent pressure correction projection method for the incompressible Navier– Stokes equations. Computers and Mathematics with Applications, 31(11):1–21, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000570.

Hwang:1999:IDC

[Hwa99]  Min-Shiang Hwang.  An

improvement of a dynamic cryptographic key assignment scheme in a tree hierarchy. Computers and Mathematics with Applications, 37(3):19–22, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000425.

Hwang:1994:TIA

[HWL94] Chyi Hwang, Rong-Yuang Wu, and Ming-Jeng Lu. A technique for increasing the accuracy of the FFT-based method of numerical inversion of Laplace transforms. Computers and Mathematics with Applications, 27(7): 23–29, April 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901465.

Han:1998:DLV

[HY98] Jen M. Han and Maria C. Yuang. A delay and loss versatile scheduling discipline in ATM switches. Computers and Mathematics with Applications, 35(5):93–106, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800008X.

Ioakimidis:1994:AGB

[IA94] N. I. Ioakimidis and E. G. Anastasselou. Application of Gröbner bases to problems of movement of a particle. Computers and Mathematics with Applications, 27(3):51–57, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900469.

Ihlenburg:1995:FES

[IB95] F. Ihlenburg and I. Babuska. Finite element solution of the Helmholtz equation with high wave number. Part I: the h-version of the FEM. Computers and Mathematics with Applications, 30(9): 9–37, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500144N.

Ibrahim:1995:SSM

[IESAH95] M. A.-K. Ibrahim, A. El-Safty, and S. M. Abo-Hasha. 2 h-step spline method for the solution of delay differential equations. Computers and Mathematics with Applications, 29(8):1–6, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500024S.

Ivanov:1998:SST

[IGM98] M. S. Ivanov, S. F. Gimelshein, and G. N. Markelov. Statistical simulation of the transition between regular and Mach reflection in steady flows. Computers and Mathematics with Applications, 35 (1–2):113–125, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002629.

Imado:1993:PEG

[II93] F. Imado and T. Ishihara. Pursuit-evasion geometry analysis between two missiles and an aircraft. Computers and Mathematics with Applications, 26(6): 125–139, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390122C.

Ivanov:1997:DRN

[IIP^{+}97] V. V. Ivanov, Yu. M. Itkis, B. Purevdorj, I. V. Puzynin, and S. I. Vinitsky. Dynamics of realistic neural network with time-dependent external signal. Computers and Mathematics with Applications, 34 (7–8):667–675, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001673.

Ishida:1995:SAS

[IK95] Y. Ishida and K. Kotovsky. Symmetry analysis on symmetry cognition on multilevel figures. Computers and Mathematics with Applications, 30(7):93–102, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500128L.

Ilin:1997:SPS

[IK97] V. A. Il’in and L. V. Kritskov. Some properties of spectral expansions related to the one-dimensional Stark effect Hamiltonian. Computers and Mathematics with Applications, 34(5–6):633– 640, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001594.

Ikehata:1991:SGF

[Ike91] Masaru Ikehata. A special Green’s function for the biharmonic operator and its application to an inverse boundary value problem. Computers and Mathematics with Applications, 22(4–5):53–66, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190131M.

Ilin:1997:HEB

[Il’97] V. A. Il’in. How to express basis conditions and conditions for the equiconvergence with trigonometric series of expansions related to nonself-adjoint differential operators. Computers and Mathematics with Applications, 34(5–6):641–647, September 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001600.

Illes:1991:OCG

[Ill91] Tibor Illés. Optimality criteria for general unconstrained geometric programming problems. Computers and Mathematics with Applications, 21(1):103–110, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190235V.

Ivanov:1993:PAR

[IM93] M. N. Ivanov and E. P. Maslov. A problem of avoidance of a rotating segment. Computers and Mathematics with Applications, 26 (6):67–75, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390118F.

Ito:1992:OIP

[IN92] Kodo Ito and Toshio Nakagawa. Optimal inspection policies for a system in storage. Computers and Mathematics with Applications, 24(1–2):87–90, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192902327.

Iserles:1997:BPZ

[IN97] A. Iserles and S. P. Nørsett. Biorthogonal polynomials and zero-mapping transformations. Computers and Mathematics with Applications, 33(1–2):129–143, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122196002258.

Ioakimidis:1991:LDI

[Ioa91] N. I. Ioakimidis. The location of discontinuity intervals of sectionally analytic functions: Application to the interface crack problem. Computers and Mathematics with Applications, 21(2–3):69–74, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190082F.

Iraniparast:1993:SFS

[Ira93] Nezam Iraniparast. Shaking a forced string to rest. Computers and Mathematics with Applications, 25(12):49–54, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390185X.

Iri:1998:TUF

[Iri98] M. Iri. Theory of uncontrollable flows — a new type of network-flow theory as a model for the 21^{st} century of multiple values. Computers and Mathematics with Applications, 35(10):107–123, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000777.

Iserles:1994:ABC

[Ise94] A. Iserles. The asymptotic behaviour of certain difference equations with proportional delays. Computers and Mathematics with Applications, 28 (1–3):141–152, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001022.

Inoue:1993:GIC

[ISM93] Hidehiko Inoue, Yukio Shibata, and Haruo Miyazaki. Generation of interpolation curves with the least energy using dynamic programming. Computers and Mathematics with Applications, 25(6): 25–34, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902968.

Ismail:1997:FPD

[Ism97] M. E. H. Ismail. Fractional powers of a difference operator. Computers and Mathematics with Applications, 33 (1–2):145–150, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219600226X.

Ispany:1998:SSO

[Isp98] M. Ispány. Stationarity and second-order properties of a scalar-valued nonlinear time series with Gaussian residuals. Computers and Mathematics with Applications, 35 (4):95–104, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002927.

Itakura:1993:SML

[Ita93] Hidekiyo Itakura. A solution to multiple linear regression problems with ordered attributes. Computers and Mathematics with Applications, 25(2):47–57, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390222H.

Iwamoto:1991:IIV

[Iwa91] Seiichi Iwamoto. Iterative integral versus dynamic programming. Computers and Mathematics with Applications, 21(11–12):23–39, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190104C.

Ivanov:1997:NIS

[IZ97] V. V. Ivanov and P. V.

Zrelov. Nonparametric integral statistics ω_{n}^{k} = n^{k2} ∫_{−∞}^{∞}[S_{n}(x)−F (x)]^{k} dF (x): Main properties and applications. Computers and Mathematics with Applications, 34 (7–8):703–726, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001715.

Jodar:1996:ACN

[JA96] L. Jódar and P. Almenar. Accurate continuous numerical solutions of time dependent mixed partial differential problems. Computers and Mathematics with Applications, 32(2):5–19, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000995.

Jodar:1997:USM

[JA97] L. Jódar and P. Almenar. Uniqueness of solution for mixed problems related to the generalized wave equation. Computers and Mathematics with Applications, 33(4):49–51, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000060.

Jabri:1996:SGS

[Jab96] A. Kh. Al Jabri. Shrinking generators and statistical leakage. Computers and Mathematics with Applications, 32(4):33–39, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001228.

Jha:1996:EEH

[JAK96] P. K. Jha, N. Agnihotri, and R. Kumar. Edge exchanges in Hamiltonian decompositions of Kroneckerproduct graphs. Computers and Mathematics with Applications, 31(2):11–19, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001891.

Janikow:1996:MPP

[Jan96] C. Z. Janikow. A methodology for processing problem constraints in genetic programming. Computers and Mathematics with Applications, 32(8):97–113, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001708.

Jarrett:1997:ERE

[Jar97] E. B. Jarrett. Edge rotation and edge slide distance graphs. Computers and Mathematics with Applications, 34 (11):81–87, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002216.

Jayaraman:1990:SEP

[Jay90]  B. Jayaraman.  Subset-

equational programming in intelligent decision systems. Computers and Mathematics with Applications, 20 (9–10):73–82, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090113X.

Jaynes:1992:CTA

[Jay92] E. T. Jaynes. Commentary on two articles by C. A. Los. Computers and Mathematics with Applications, 24(8–9): 267–273, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290204U.

See [Los89b, Los89a].

Janikow:1993:AAE

[JCL93] Cezary Z. Janikow, Haiyan Cai, and Xiaolong Luo. An approximate algorithm for estimating treatment lags from right censored data. Computers and Mathematics with Applications, 25(12):73–85, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390187Z.

Jodar:1996:CCV

[JCM96] L. Jódar, J. L. Cortés, and J. L. Morera. Construction and computation of variable coefficient Sylvester differential problems. Computers and Mathematics with Applications, 32(8):41–50, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001654.

Jodar:1997:ANS

[JD97] L. Jódar and E. Defez.

Analytic-numerical solutions with a priori error bounds for time-dependent mixed partial differential problems. Computers and Mathematics with Applications, 34 (12):49–59, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002332.

Jekeli:1994:HMB

[Jek94] C. Jekeli. Hardy’s multiquadricbiharmonic method for gravity field predictions. Computers and Mathematics with Applications, 28(7):43–46, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122194001588.

Jodar:1995:ANS

[JG95] L. Jódar and D. Goberna. Analytic numerical solution of coupled semi-infinite diffusion problems. Computers and Mathematics with Applications, 30(12):1–10, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001704.

Jodar:1996:EAN

[JG96] L. Jódar and D. Goberna. Exact and analytic numerical solution of coupled diffusion problems in a semiinfinite medium. Computers and Mathematics with Applications, 31(9):17–24, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000387.

Jodar:1998:MDF

[JG98] L. Jódar and D. Goberna. A matrix d’Alembert formula for coupled wave initial value problems. Computers and Mathematics with Applications, 35(9):1–15, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000522.

Jablonski:1995:IMG

[JGB95] M. Jabloński, P. Góra, and A. Boyarsky. Invariant measures generated by sequences of approximating transformations. Computers and Mathematics with Applications, 30(7):75–91, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500127K.

Jha:1990:NCS

[Jha90] N. K. Jha. A new class of symmetric error correcting/unidirectional error detecting codes. Computers and Mathematics with Applications, 19(5):95–104, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090105S.

Jones:1994:ARI

[JI94] B. Jones and S. Iyengar. Approximate root isolation for nonlinear systems by Monte-Carlo. Computers and Mathematics with Applications, 27 (7):1–5, April 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901430.

Janicki:1996:WOA

[JK96] R. Janicki and W. W.

Koczkodaj. A weak order approach to group ranking. Computers and Mathematics with Applications, 32 (2):51–59, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001022.

Jun:1998:VSM

[JKK98] S.-R. Jun, Y.-H. Kwon, and S. Kang. A variational spectral method for the twodimensional Stokes problem. Computers and Mathematics with Applications, 35(4): 1–17, February 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700285X.

Jaimini:1994:SMS

[JKS94] B. B. Jaimini, C. L. Koul, and H. M. Srivastava. Some multiple series identities. Computers and Mathematics with Applications, 28(4): 19–24, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001235.

Jeurissen:1994:LBG

[JL94] R. Jeurissen and W. Layton. Load balancing by graph coloring, an algorithm. Computers and Mathematics with Applications, 27(3): 27–32, February 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900434.

Jwo:1992:CPT

[JLDL92] Jung-Sing Jwo, S. Lakshmivarahan, Sudarshan K. Dhall, and John M. Lewis. Comparison of performance of three parallel versions of the block cyclic reduction algorithm for solving linear elliptic partial differential equations. Computers and Mathematics with Applications, 24(5–6):83–101, September 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290043H.

Jin:1993:JSG

[JLL^{+}93] De-Jun Jin, Si-Zhang Liu, Sang-Ho Lee, Hai-Lou Liu, Xi-Guan Lu, and Dazhi Zhang. The joint sum of graceful trees. Computers and Mathematics with Applications, 26(10):83–87, November 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390027S.

Jacobs:1991:ARS

[JM91] David Jacobs and Semyon M.

Meerkov. Asymptotically reliable serial production lines with a quality control system. Computers and Mathematics with Applications, 21 (11–12):85–90, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190110P.

Jansen:1994:SEW

[JMPS94] J. K. M. Jansen, R. M. M. Mattheij, M. T. M. Penders, and W. H. A. Schilders. Stability and efficiency of waveform relaxation methods. Computers and Mathematics with Applications, 28 (1–3):153–166, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001030.

Johnson:1995:SR

[Joh95] J. A. Johnson. Semantic relatedness. Computers and Mathematics with Applications, 29(5):51–63, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400248J.

Johnston:1999:CSL

[Joh99] P. R. Johnston. Calculating surface Laplacians as a boundary element method postprocess. Computers and Mathematics with Applications, 37(10):95–103, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001285.

Jones:1996:UUG

[Jon96] M. Jones. Universal unfoldings of group invariant equations which model second and third harmonic resonant capillary-gravity waves. Computers and Mathematics with Applications, 32 (10):59–89, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001861.

Joshi:1990:ISD

[Jos90] R. R. Joshi. An intelligent system for DNA repair. Computers and Mathematics with Applications, 19(6):1–7, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090042I.

Jackiewicz:1998:PWR

[JOW98] Z. Jackiewicz, B. Owren, and B. Welfert. Pseudospectra of waveform relaxation operators. Computers and Mathematics with Applications, 36(8):67–85, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001849.

Johnstone:1993:RNA

[JP93] Paul Johnstone and Frederick E. Petry. Rational number approximation in higher radix floating point systems. Computers and Mathematics with Applications, 25(6): 103–108, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390304E.

Jodar:1995:CCN

[JP95a] L. Jodar and E. Ponsoda. Computing continuous numerical solutions of matrix differential equations. Computers and Mathematics with Applications, 29(4): 73–84, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400240L.

Jodar:1995:CNS

[JP95b] L. Jódar and E. Ponsoda. Continuous numerical solutions and error bounds for time dependent systems of partial differential equations: Mixed problems. Computers and Mathematics with Applications, 29(8):63–71, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000303.

Jodar:1999:CPA

[JPC99] L. Jódar, A. E. Posso, and H. Castejón. Constructive polynomial approximation with a priori error bounds for nonlinear initial value differential problems. Computers and Mathematics with Applications, 38(7–8):9–19, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002333.

Jodar:1996:WCV

[JPF96] L. Jódar, E. Ponsoda, and M. Legua Fernández. Wellconditioning for variable coefficient second order boundary value difference systems. Computers and Mathematics with Applications, 31(12): 35–48, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000740.

Jayakumar:1994:NMS

[JR94] J. Jayakumar and N. Ramanujam. A numerical method for singular perturbation problems arising in chemical reactor theory. Computers and Mathematics with Applications, 27(5): 83–99, March 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900787.

Jana:1995:FPA

[JS95] P. K. Jana and B. P. Sinha. Fast parallel algorithm for polynomial interpolation. Computers and Mathematics with Applications, 29(4):85–92, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400241C.

Jana:1997:FPA

[JS97a] P. K. Jana and B. P. Sinha. Fast parallel algorithms for forecasting. Computers and Mathematics with Applications, 34(9):39–49, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001879.

Joshi:1997:NLN

[JS97b] R. R. Joshi and D. K. Satyanarayana. A nested layered network model for parallel solutions of discrete SPPs. Computers and Mathematics with Applications, 33(5):111–123, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000242.

Jana:1998:FPA

[JS98] P. K. Jana and B. P. Sinha. Fast parallel algorithms for Graeffe’s root squaring technique. Computers and Mathematics with Applications, 35(3):71–80, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002800.

Joshi:1999:CFM

[JS99] S. B. Joshi and H. M. Srivastava. A certain family of meromorphically multivalent functions. Computers and Mathematics with Applications, 38(3–4):201– 211, August 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001947.

Jacobs:1999:LTL

[JT99] D. P. Jacobs and V. Trevisan. Linear-time LUP decomposition of forest-like matrices. Computers and Mathematics with Applications, 37 (10):37–50, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001248.

Jiang:1997:NCM

[JTY97] D. C. Jiang, K. L. Teo, and W. Y. Yan. A new computational method for the functional inequality constrained minimax optimization problem. Computers and Mathematics with Applications, 33(6):53–63, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700031X.

Jodar:1996:ESI

[JV96] L. Jódar and R. J. Villanueva. Explicit solutions of implicit second-order difference systems in unbounded bilateral domains. Computers and Mathematics with Applications, 32(9):19–28, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001733.

Jan:1993:DAC

[JW93] Jinn-Ke Jan and Shiuh-Jeng Wang. A dynamic access control scheme based upon the knapsack problem. Computers and Mathematics with Applications, 26 (12):75–86, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390061Y.

Jou:1993:AGM

[JY93] Chin-Rong Jou and Jen-Chih Yao. Algorithm for generalized multivalued variational inequalities in Hilbert spaces. Computers and Mathematics with Applications, 25 (9):7–16, May 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390127H.

Khusainov:1999:SEC

[KAD99] D. Khusainov, R. P. Agarwal, and V. Davidov. Stability and estimates for the convergence of solutions for systems involving quadratic terms with constant deviating arguments. Computers and Mathematics with Applications, 38 (11–12):141–149, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900293X.

Kaiser:1994:BSP

[Kai94] M. J. Kaiser. The Blaschke– Steinhardt point of a planar convex set. Computers and Mathematics with Applications, 28(7):1–11, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001561.

Kaiser:1996:CPR

[Kai96] M. J. Kaiser. The critical point and related symmetry measures of a planar convex set. Computers and Mathematics with Applications, 32(9):79–98, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001794.

Kalamboukis:1990:TBS

[Kal90] T. Z. Kalamboukis. Tridiagonalization of band symmetric matrices for vector computers. Computers and Mathematics with Applications, 19(2):29–34, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190900044.

Kampke:1994:EVM

[Käm94a] T. Kämpke. Efficient versioning for matrix structures. Computers and Mathematics with Applications, 27(3):9–19, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900418.

Kampke:1994:TBC

[Käm94b] T. Kämpke. Tight bounds for capacities. Computers and Mathematics with Applications, 27(8):67–86, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900671.

Kansa:1990:MSDa

[Kan90a] E. J. Kansa. Multiquadrics — a scattered data approximation scheme with applications to computational fluid-dynamics — I: surface approximations and partial derivative estimates. Computers and Mathematics with Applications, 19(8–9):127– 145, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090270T.

Kansa:1990:MSDb

[Kan90b] E. J. Kansa. Multiquadrics — a scattered data approximation scheme with applications to computational fluiddynamics — II. Solutions to parabolic, hyperbolic and elliptic partial differential equations. Computers and Mathematics with Applications, 19 (8–9):147–161, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090271K.

Kansa:1992:SCS

[Kan92a] E. J. Kansa. A strictly conservative spatial approximation scheme for the governing engineering and physics equations over irregular regions and inhomogeneously scattered nodes. Computers and Mathematics with Applications, 24(5–6):169–190, September 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290047L.

Kansa:1992:P

[Kan92b] Edward J. Kansa. Preface. Computers and Mathematics with Applications, 24(12):vii, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290163C.

Karageorghis:1994:SBC

[Kar94]  A. Karageorghis.  Satis-

faction of boundary conditions for Chebyshev collocation methods in cuboidal domains. Computers and Mathematics with Applications, 27(12):85–90, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900884.

Karafiat:1996:AIT

[Kar96a] A. Karafiat. Adaptive integration techniques for almost singular functions in the boundary element method. Computers and Mathematics with Applications, 32 (5):11–30, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001320.

Karasozen:1996:CIB

[Kar96b] B. Karasözen. Composite integrators for bi-Hamiltonian systems. Computers and Mathematics with Applications, 32(7):79–86, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001575.

Kiryakova:1999:ESH

[KAS99] V. Kiryakova and B. Al-Saqabi. Explicit solutions to hyper-Bessel integral equations of second kind. Computers and Mathematics with Applications, 37(1): 75–86, January 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002430.

Katayama:1992:PAO

[Kat92] Tsuyoshi Katayama. Performance analysis and optimization of a cyclic-service tandem queueing system with multiclass customers. Computers and Mathematics with Applications, 24(1–2):25–33, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902246.

Kulev:1990:GSS

[KB90] G. K. Kulev and D. D. Bainov. Global stability of sets for impulsive differential systems by Lyapunov’s direct method. Computers and Mathematics with Applications, 19(2):17–28, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900033.

Krieger:1990:ITC

[KBLS90] D. Krieger, T. W. Berger, S. Levitan, and R. J. Sclabassi. An interactive toolset for characterizing complex neural systems. Computers and Mathematics with Applications, 20 (4–6):231–246, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090331D.

Kostova:1991:MDI

[KC91] T. V. Kostova and N. H. Chipev. A model of the dynamics of intramolluscan trematode populations: Some problems concerning oscillatory behavior. Computers and Mathematics with Applications, 21(5):1–15, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190212M.

Kansa:1992:IAM

[KC92a] E. J. Kansa and R. E. Carlson. Improved accuracy of multiquadric interpolation using variable shape parameters. Computers and Mathematics with Applications, 24 (12):99–120, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290174G.

King:1992:ISQ

[KC92b] R. B. King and E. R. Canfield. Icosahedral symmetry and the quintic equation. Computers and Mathematics with Applications, 24(3):13–28, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902109.

Kwak:1990:AGM

[KCB90] Do Y. Kwak, U. Jin Choi, and Sung H. Bae. An algebraic geometric method to calculate branches near a bifurcation point. Computers and Mathematics with Applications, 19(10):87–90, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090360V.

Khan:1997:VAC

[KCD^{+}97] A. R. Khan, P. P. Choudhury, K. Dihidar, S. Mitra, and P. Sarkar. VLSI architecture of a cellular automata machine. Computers and Mathematics with Applications, 33(5):79–94, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000217.

Khan:1999:TCU

[KCDV99] A. R. Khan, P. P. Choudhury, K. Dihidar, and R. Verma. Text compression using twodimensional cellular automata. Computers and Mathematics with Applications, 37(6): 115–127, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000802.

Kelly:1991:PME

[Kel91] Theresa D. Kelly. Particle modelling of an elastic arch in three dimensions. Computers and Mathematics with Applications, 22(6):47–91, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190004N.

Kelley:1994:AAS

[Kel94] W. Kelley. Asymptotic analysis of solutions in the “double root” case. Computers and Mathematics with Applications, 28(1–3):167– 173, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001049.

Kertesz:1991:CES

[Ker91a] V. Kertész. Conditions for the existence of solutions converging to zero in the case of homogeneous linear differential equations. Computers and Mathematics with Applications, 21(1):111–118, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190236W.

Kertesz:1991:PSI

[Ker91b] V. Kertész. Partial stability investigations of differential equations of second order. Computers and Mathematics with Applications, 21(1):95–102, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190234U.

Khater:1990:ASR

[KESH90] A. H. Khater, A. E. El-Sharif, and N. A. Hussien. Analytical solutions of the Rayleigh flow problem for one component charged rarefied gas. Computers and Mathematics with Applications, 19(5):67–73, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090102P.

Kharab:1994:NMW

[KG94] A. Kharab and R. B. Guenther. A numerical model of water, heat and chemical transport in porous media. Computers and Mathematics with Applications, 27(2):55–64, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900361.

Korzeniowski:1995:NSM

[KG95a] A. Korzeniowski and D. Greenspan. Numerical studies of microturbulence in water. Computers and Mathematics with Applications, 29(6):7–15, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500002G.

Kraut:1995:CIP

[KG95b] G. L. Kraut and I. Gladwell. Convergence and instability in PCG methods for bordered systems. Computers and Mathematics with Applications, 30 (12):101–109, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500177Z.

Korlie:1997:PSC

[KG97] M. S. Korlie and D. Greenspan.

Particle simulation of carbon dioxide bubbles in water. Computers and Mathematics with Applications, 34 (9):23–37, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001867.

Katti:1998:PMC

[KG98] C. P. Katti and S. Goel. A parallel mesh chopping algorithm for a class of twopoint boundary value problems. Computers and Mathematics with Applications, 35(9):121–128, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000637.

Koc:1992:AAS

[KH92]  Çetin K. Koç and Ching-

Yu Hung. Adaptive mary segmentation and canonical recoding algorithms for multiplication of large binary numbers. Computers and Mathematics with Applications, 24(3):3–12, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290209Z.

Kharab:1990:PFP

[Kha90] A. Kharab. A Poiseuille flow past a permeable hill. Computers and Mathematics with Applications, 19(2):9–15, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190900022.

Kharab:1997:UMS

[Kha97] A. Kharab. Use of multiple sheets for the solution of a three-dimensional transient heat conduction problem. Computers and Mathematics with Applications, 34 (1):71–79, July 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000989.

Kuttner:1992:PNL

[KHC92] Eliza Kuttner, William Havens, and Nick Cercone. Processing natural language with schema constraint networks. Computers and Mathematics with Applications, 24 (11):3–10, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290026E.

Kondoh:1994:KON

[KHI94] Y. Kondoh, Y. Hosaka, and K. Ishii. Kernel optimum nearly-analytical discretization (KOND) algorithm applied to parabolic and hyperbolic equations. Computers and Mathematics with Applications, 27(3): 59–90, February 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900477.

Kaykobad:1998:MPA

[KIAM98] M. Kaykobad, Md. M. Islam, M. E. Amyeen, and M. M. Murshed. 3 is a more promising algorithmic parameter than 2. Computers and Mathematics with Applications, 36(6):19–24, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001588.

Kijima:1992:TSC

[Kij92] Masaaki Kijima. The transient solution to a class of Markovian queues. Computers and Mathematics with Applications, 24(1–2):17– 24, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122192902235.

Kimura:1992:DBM

[Kim92] Toshikazu Kimura. Duality between the M/G/s/s and GI/M/1/s/s queues. Computers and Mathematics with Applications, 24(1– 2):35–40, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192902257.

Kim:1998:SSI

[Kim98] S. Kim. Solving singular integral equations using Gaussian quadrature and overdetermined system. Computers and Mathematics with Applications, 35(10):63–71, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800073X.

Kim:1999:NSC

[Kim99] S. Kim. Numerical solutions of Cauchy singular integral equations using generalized inverses. Computers and Mathematics with Applications, 38(5–6):183– 195, September 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002254.

Kiss:1991:ASP

[Kis91] A. Kiss. The application of the separation principle for the linear continuous systems with coloured noise. Computers and Mathematics with Applications, 21(1): 119–128, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190237X.

Kostreva:1991:DHA

[KK91] M. M. Kostreva and L. A.

Kinard. A differentiable homotopy approach for solving polynomial optimization problems and noncooperative games. Computers and Mathematics with Applications, 21 (6–7):135–143, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901684.

Kawai:1992:OMP

[KK92] Hajime Kawai and Junji Koyanagi. An optimal maintenance policy of a discrete time Markovian deterioration system. Computers and Mathematics with Applications, 24(1–2):103–108, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902349.

Kurano:1994:EOS

[KK94] M. Kurano and M. Kawai.

Existence of optimal stationary policies in discounted Markov decision processes: approaches by occupation measures. Computers and Mathematics with Applications, 27(9–10):95–101, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901287.

Kalla:1997:TMA

[KK97a] S. L. Kalla and H. G. Khajah. Tau method approximation of some integrals related to radiation field problems. Computers and Mathematics with Applications, 33(5):21–27, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000163.

Kang:1997:CEM

[KK97b] K. S. Kang and D. Y. Kwak. Convergence estimates for multigrid algorithms. Computers and Mathematics with Applications, 34 (9):15–22, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001855.

Kocharovsky:1997:ISE

[KKB97] V. V. Kocharovsky, Vl. V. Kocharovsky, and A. A. Belyanin. Inhibited spontaneous emission and electromagnetic instability of an atom in the near zone from the surface of an active medium. Computers and Mathematics with Applications, 34(7–8):795– 805, October 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001752.

Kocharovsky:1997:NDS

[KKD^{+}97] V. V. Kocharovsky, Vl. V.

Kocharovsky, E. A. Derishev, S. A. Litvak, I. A. Shereshevsky, and S. Tasaki. Nonstationary dressed states and effects of decay in nonadiabatic crossing of decaying levels. Computers and Mathematics with Applications, 34 (7–8):727–750, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001727.

Kocharovsky:1997:MIN

[KKG97] V. V. Kocharovsky, Vl. V.

Kocharovsky, and E. R. Golubyatnikova. Mode instability and nonlinear superradiance phenomena in open Fabry–Perot cavity. Computers and Mathematics with Applications, 34(7–8):773– 793, October 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001740.

Kagiwada:1991:AMT

[KKK91] Harriet H. Kagiwada, Julia K. Kagiwada, and Robert E. Kalaba. Associative memories for third-generation passive ranging. Computers and Mathematics with Applications, 21(2–3):81–85, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190084H.

Kagiwada:1993:PPR

[KKK93] H. H. Kagiwada, J. K. Kagiwada, and R. E. Kalaba. Precision passive ranging. Computers and Mathematics with Applications, 26(4): 89–96, August 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390036U.

Kocharovsky:1997:ESD

[KKK^{+}97] V. V. Kocharovsky, Vl. V. Kocharovsky, V. A. Kukushkin, Yu. M. Mironov, and V. Yu. Zaitsev. Effects of the spatial dispersion and instabilities in the wave electrodynamics of weakly ionized gases. Computers and Mathematics with Applications, 34 (7–8):807–844, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001764.

Kim:1996:DRB

[KKM96] H. O. Kim, J. H. Kim, and S. Y. Moon. Degree reduction of Bézier curves and filter banks. Computers and Mathematics with Applications, 31(10):23–30, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000491.

Katsumata:1996:SFD

[KKOS96] M. Katsumata, M. Kurihara, A. Ohuchi, and Y. Sugasawa. Serial failure diagnosis of a distributed processing system by Petri nets. Computers and Mathematics with Applications, 31(9):57–62, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000429.

Kim:1998:EAD

[KKS98] H. S. Kim, H. O. Kim, and S. Y. Shin. An efficient algorithm for determining the extreme vertices of a moving 3D convex polyhedron with respect to a plane. Computers and Mathematics with Applications, 36(3):55–61, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800128X.

Kelmanson:1995:ABS

[KL95] M. A. Kelmanson and B. Lonsdale. Annihilation of boundary singularities via suitable Green’s functions. Computers and Mathematics with Applications, 29(4):1–7, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400232A.

Kim:1996:ADT

[KL96a] S. Kim and M. Lee. Artificial damping techniques for scalar waves in the frequency domain. Computers and Mathematics with Applications, 31(8):1–12, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000259.

Kostova:1996:OSD

[KL96b] T. Kostova and J. Li. Oscillations and stability due to juvenile competitive effects on adult fertility. Computers and Mathematics with Applications, 32 (11):57–70, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196001976.

Klar:1998:AAC

[Kla98] A. Klar. Asymptotic analysis and coupling conditions for kinetic and hydrodynamic equations. Computers and Mathematics with Applications, 35(1–2):127– 137, January 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002630.

Klementiev:1993:DDP

[Kle93] D. A. Klementiev. The 3-D detection problem of an evader moving in a fixed plane. Computers and Mathematics with Applications, 26 (6):77–85, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390119G.

Kuperin:1997:QLG

[KLMP97] Yu. A. Kuperin, S. B. Levin, Yu. B. Melnikov, and B. S. Pavlov. Quantum Lorentz gas: Effective equations and spectral analysis. Computers and Mathematics with Applications, 34(5–6):599– 612, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001569.

Kuperin:1997:AET

[KLMY97] Yu. A. Kuperin, S. B. Levin, Yu. B. Melnikov, and E. A. Yarevsky. Application of extension theory to antiprotonnucleon systems with continuous set of resonances in annihilation channel. Computers and Mathematics with Applications, 34(5–6):559– 570, September 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001533.

Kurkal:1990:CSU

[KM90] K. R. Kurkal and S. S. Munukutla. Computational study of an unsteady heat transfer problem related to high power lasers. Computers and Mathematics with Applications, 19(8–9):77–83, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090265L.

Kamiyama:1992:DTM

[KM92] Shin-Ichi Kamiyama and Yasuo Murata. The discretetime model of Bagchi’s regional investment allocation problem. Computers and Mathematics with Applications, 24(8–9):181–194, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290197P.

Klimov:1993:KIB

[KM93] G. P. Klimov and G. R. Marshall. Kinetics of interactions between drugs and receptors: stochastic conception. Computers and Mathematics with Applications, 26 (10):1–25, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390022N.

Klimov:1996:MSA

[KM96] G. P. Klimov and G. R. Marshall. Measure of similarity of amino acids or how to obtain the genetic code table from “nothing”. Computers and Mathematics with Applications, 31(8):13–21, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000260.

Khamis:1997:CGM

[KM97a] H. J. Khamis and T. A. Mc-Kee. Chordal graph models of contingency tables. Computers and Mathematics with Applications, 34 (11):89–97, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002228.

Kim:1997:DRB

[KM97b] H. O. Kim and S. Y. Moon. Degree reduction of Bézier curves by L ^{1}-approximation with endpoint interpolation. Computers and Mathematics with Applications, 33(5): 67–77, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000205.

Kocic:1997:SPA

[KM97c] L. M. Kocić and G. V. Milovanović. Shape preserving approximations by polynomials and splines. Computers and Mathematics with Applications, 33(11):59–97, June 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000874.

Kumar:1999:RDI

[KMD99] R. Kumar, A. Miglani, and L. Debnath. Radial displacements of an infinite liquid saturated porous medium with cylindrical cavity. Computers and Mathematics with Applications, 37(7):117–123, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000929.

Kim:1990:RPC

[KMR90] S. Kim, S. M. Meerkov, and T. Runolfsson. Residence probability control. Computers and Mathematics with Applications, 19 (11):121–125, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090154C.

Kaczynski:1998:HCR

[KMŚ98] T. Kaczyński, M. Mrozek, and M. Ślusarek. Homology computation by reduction of chain complexes. Computers and Mathematics with Applications, 35(4): 59–70, February 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002897.

Kwapisz:1994:DEL

[KN94] M. Kwapisz and K. Nowicka. Difference equations and local convergence of inexact Newton methods. Computers and Mathematics with Applications, 28(1–3):175– 183, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001057.

Kulkarni:1999:AFC

[KNS99] S. R. Kulkarni, U. H. Naik, and H. M. Srivastava. An application of fractional calculus to a new class of multivalent functions with negative coefficients. Computers and Mathematics with Applications, 38(5–6):169–182, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002242.

Kalaba:1999:EPT

[KNU99] R. E. Kalaba, H. H. Natsuyama, and S. Ueno. The estimation of parameters in time-dependent transport problems: Dynamic programming and associative memories. Computers and Mathematics with Applications, 37(2):41–45, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002521.

Kalaba:1999:BGP

[KNUX99] R. Kalaba, H. Natsuyama, S. Ueno, and R. Xu. The Bellman–Gauss principle for constrained motion. Computers and Mathematics with Applications, 37(11–12):1– 7, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199001364.

Kaio:1990:OOP

[KO90] N. Kaio and S. Osaki. Optimal ordering policies with two types of randomized lead times. Computers and Mathematics with Applications, 19(11):43–52, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090147C.

Khajah:1991:ULE

[KO91] H. G. Khajah and E. L. Ortiz. Upper and lower error estimation for the Tau method and related polynomial techniques. Computers and Mathematics with Applications, 22(3):81–87, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190072C.

Kolar:1993:FPR

[KO93] Miroslav Kolár and Seamus F.

O’Shea. Fast, portable, and reliable algorithm for the calculation of Halton numbers. Computers and Mathematics with Applications, 25 (7):3–13, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390307H.

Khajah:1994:UHP

[KO94] H. G. Khajah and E. L. Ortiz. Ultra-high precision computations. Computers and Mathematics with Applications, 27(7):41–57, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901481.

Katsurada:1996:CPF

[KO96] M. Katsurada and H. Okamoto. The collocation points of the fundamental solution method for the potential problem. Computers and Mathematics with Applications, 31 (1):123–137, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001863.

Koczkodaj:1997:OBC

[KO97] W. W. Koczkodaj and M. Orlowski. An orthogonal basis for computing a consistent approximation to a pairwise comparisons matrix. Computers and Mathematics with Applications, 34 (10):41–47, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002058.

Koczkodaj:1999:CCA

[KO99] W. W. Koczkodaj and M. Orlowski. Computing a consistent approximation to a generalized pairwise comparisons matrix. Computers and Mathematics with Applications, 37(3):79–85, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000486.

Koc:1991:IAM

[Koç91]  Çetin K. Koç.  An im-

proved algorithm for mixedradix conversion of residue numbers. Computers and Mathematics with Applications, 22(8):63–71, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190014U.

Koc:1992:PAE

[Koç92]  Çetin K. Koç. A parallel al-

gorithm for exact solution of linear equations via congruence technique. Computers and Mathematics with Applications, 23(12):13–24, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290089Z.

Koc:1995:ASW

[Koç95] C. K. Koç. Analysis of sliding window techniques for exponentiation. Computers and Mathematics with Applications, 30(10):17–24, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500153P.

Kohler:1991:FSC

[Köh91]  Adelhard E. Köhler.  A

fuzzy symmetry concept for forms with imperfect symmetries. Computers and Mathematics with Applications, 22(9):35–50, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190205I.

Kohler:1993:IMF

[Köh93a] Adelhard E. Köhler. Isosymmetric manifolds in form spaces and the normal deformations of polygonal forms. Computers and Mathematics with Applications, 25(9): 67–89, May 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390134H.

Kohler:1993:SAF

[Köh93b] Adelhard E. Köhler. A second approach to fuzzy symmetries: Fuzzy symmetry requirements. Computers and Mathematics with Applications, 25(1):17–34, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390208D.

Kohler:1994:CFM

[Köh94] A. E. Köhler. Crisp and fuzzy motif and arrangement symmetries in composite geometric figures. Computers and Mathematics with Applications, 27(11):67–87, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219490099X.

Kohler:1995:DIM

[Köh95] A. E. Köhler. Deformations, isosymmetric manifolds, and higher-dimensional form space symmetries for point ensembles (polygonal forms) under o(2) symmetry i. two and three points. Computers and Mathematics with Applications, 29(3): 63–90, February 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400229E.

Kohler:1996:DIMa

[Köh96a] A. E. Köhler. Deformations, isosymmetric manifolds, and higher dimensional form space symmetries for point ensembles (polygonal forms) under O(2) symmetry II. Four points. Computers and Mathematics with Applications, 31(9):89–116, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000454.

Kohler:1996:DIMb

[Köh96b] A. E. Köhler. Deformations, isosymmetric manifolds, and higher dimensional form space symmetries for point ensembles (polygonal forms) under O(2) symmetry III. Motifs in geometric figures and constraints in form spaces. Computers and Mathematics with Applications, 31(9):117–140, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000466.

Kohler:1996:PFT

[Köh96c]  A. E. Köhler.  Polygo-

nal figures and their form spaces i. the structure of isodiastemic manifolds (manifolds of equilateral polygons). Computers and Mathematics with Applications, 32(6): 101–128, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001472.

Kohler:1997:TQS

[Köh97] A. E. Köhler. De triangularitate quadranguli or symmetry regions and symmetry centroids of geometric figures. Computers and Mathematics with Applications, 34 (9):95–121, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001922.

Kohler:1999:FSP

[Köh99] A. E. Köhler. Fuzzy symmetries of point ensembles. Computers and Mathematics with Applications, 37(6): 89–113, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000796.

Kormos:1990:HTN

[Kor90a] J. Kormos. Hypothesis testing for nearly nonstationary autoregressive models. Computers and Mathematics with Applications, 19(1): 75–82, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090084W.

Koroljuk:1990:CLT

[Kor90b] V. S. Koroljuk. Central limit theorem for semi-Markov random evolutions. Computers and Mathematics with Applications, 19(1):83–88, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090085X.

Kormos:1991:HTN

[Kor91] J. Kormos. Hypothesis testing for nearly nonstationary AR(1) model with Gaussian autoregressive innovation. Computers and Mathematics with Applications, 21(1):129–136, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190238Y.

Korzeniowski:1992:QMS

[Kor92] Andrzej Korzeniowski. Quantum mechanical simulation of the hydrogen molecule. Computers and Mathematics with Applications, 24(4): 99–102, August 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900127.

Kornhuber:1996:PEE

[Kor96] R. Kornhuber. A posteriori error estimates for elliptic variational inequalities. Computers and Mathematics with Applications, 31(8): 49–60, April 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000302.

Korlie:1997:PML

[Kor97] M. S. Korlie. Particle modeling of liquid drop formation on a solid surface in 3-D. Computers and Mathematics with Applications, 33 (9):97–114, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000692.

Kostova:1990:NSE

[Kos90] T. V. Kostova. Numerical solutions to equations modelling nonlinearly interacting age-dependent populations. Computers and Mathematics with Applications, 19 (8–9):95–103, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090267N.

Kostreva:1993:CON

[Kos93] Michael M. Kostreva. Combinatorial optimization in Nash games. Computers and Mathematics with Applications, 25(10–11):27–34, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902784.

Kostreva:1999:RPA

[KOT99] M. M. Kostreva, W. Ogryczak, and D. W. Tonkyn. Relocation problems arising in conservation biology. Computers and Mathematics with Applications, 37(4–5):135– 150, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000656.

Kouyoumdjian:1989:STA

[Kou89] E. R. Kouyoumdjian. A set theoretic approach to the symmetry analysis of hexcadecahedrane. Computers and Mathematics with Applications, 17(1–3):375– 395, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122189901697.

See errata [Kou90b].

Koukhaleishvili:1990:ADQ

[Kou90a] Rati E. Koukhaleishvili. Asymptotic distribution of the quadratic deviation of the regression curve recursive estimator. Computers and Mathematics with Applications, 19(1):89–96, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090086Y.

Kouyoumdjian:1990:EST

[Kou90b] E. R. Kouyoumdjian. Errata to “A set theoretic approach to the symmetry analysis of hexadecahedrane”: Computers math. applic. 17 (1–3), 375–395 (1989). Computers and Mathematics with Applications, 19(2):97, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090013A.

See [Kou89].

Kounchev:1995:ORL

[Kou95] O. Kounchev. Optimal recovery of linear functionals of Peano type through data on manifolds. Computers and Mathematics with Applications, 30(3–6):335– 351, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001115.

Koura:1998:INC

[Kou98] K. Koura. Improved nullcollision technique in the direct simulation Monte Carlo method: Application to vibrational relaxation of nitrogen. Computers and Mathematics with Applications, 35 (1–2):139–154, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002642.

Kovacheva:1995:NLD

[Kov95] R. K. Kovacheva. A note on the limiting distribution of the zeros of Padé approximants for some classes of functions. Computers and Mathematics with Applications, 30(3–6):147– 153, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000925.

Kim:1995:MAP

[KP95a] M.-Y. Kim and E.-J. Park. Mixed approximation of a population diffusion equation. Computers and Mathematics with Applications, 30 (12):23–33, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500172U.

Kim:1995:USN

[KP95b] M.-Y Kim and E.-J Park. An upwind scheme for a nonlinear model in age-structured population dynamics. Computers and Mathematics with Applications, 30(8):5–17, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500132I.

Kjurkchiev:1996:BAS

[KP96] N. Kjurkchiev and M. S.

Petković. The behaviour of approximations of the SOR Weierstrass method. Computers and Mathematics with Applications, 32 (7):117–121, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001605.

Kordonis:1997:OND

[KP97a] I.-G. E. Kordonis and Ch. G.

Philos. Oscillation of neutral difference equations with periodic coefficients. Computers and Mathematics with Applications, 33(7):11–27, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000400.

Kormos:1997:NUM

[KP97b] J. Kormos and G. Pap. Nearly unstable multidimensional AR processes. Computers and Mathematics with Applications, 34(10):1–7, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002022.

Karageorghis:1998:DMS

[KP98] A. Karageorghis and M. Paprzycki. Direct methods for spectral approximations in nonconforming domain decompositions. Computers and Mathematics with Applications, 35(11):75–82, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000856.

Kim:1999:FDM

[KP99a] M.-Y. Kim and E.-J. Park. Fully discrete mixed finite element approximations for non-Darcy flows in porous media. Computers and Mathematics with Applications, 38 (11–12):113–129, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002916.

Kordonis:1999:BSL

[KP99b] I.-G. E. Kordonis and Ch. G. Philos. The behavior of solutions of linear integrodifferential equations with unbounded delay. Computers and Mathematics with Applications, 38(2):45–50, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001819.

Kukolj:1997:AUL

[KPB97] D. Kukolj, D. Popovic, and M. Borota. Applied unsupervised learning in model reduction of linear dynamic systems. Computers and Mathematics with Applications, 33 (3):95–103, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002404.

Kozak:1996:RMC

[KPS96] I. Kozák, I. Páczelt, and Gy. Szeidl. Relative motion of continua with applications to the analysis of nonlinear problems by the finite element method. Computers and Mathematics with Applications, 31(4–5):191– 199, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002324.

Krass:1990:SMC

[Kra90]  Iosif A. Krass.  Shadow

method for convex programming with application for Navy credit sea/shore rotation problem. Computers and Mathematics with Applications, 20(2):67–80, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090242C.

Krawczyk:1992:PRC

[Kra92] Jacek B. Krawczyk. Properties of repetitive control of partially observed processes. Computers and Mathematics with Applications, 24(8–9): 111–124, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290191J.

Kreith:1991:SAP

[Kre91] Kurt Kreith. A self-adjoint problem for the wave equation in higher dimensions. Computers and Mathematics with Applications, 21(5): 129–132, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190221O.

Krishchenko:1997:EDC

[Kri97] A. P. Krishchenko. Estimations of domains with cycles. Computers and Mathematics with Applications, 34 (2–4):325–332, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001302.

Kropinski:1999:IEM

[Kro99] M. C. A. Kropinski. Integral equation methods for particle simulations in creeping flows. Computers and Mathematics with Applications, 38 (5–6):67–87, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002151.

Kwak:1991:PSS

[KS91] Do Y. Kwak and S. Son.

Prime sample scheme for almost sure convergence of a Galerkin approximation. Computers and Mathematics with Applications, 22(1): 45–48, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190024X.

Kaplan:1992:PMI

[KS92] Wilfred Kaplan and Frederick B. Sleator. Procedures for multiple integration. Computers and Mathematics with Applications, 24 (7):125–126, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290160J.

Kimmel:1995:STD

[KS95] R. Kimmel and G. Sapiro.

Shortening three-dimensional curves via two-dimensional flows. Computers and Mathematics with Applications, 29(3):49–62, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400228D.

Krattenthaler:1996:SBH

[KS96] C. Krattenthaler and H. M. Srivastava. Summations for basic hypergeometric series involving a q-analogue of the digamma function. Computers and Mathematics with Applications, 32(3): 73–91, August 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001149.

Kalpakam:1997:LSI

[KS97a] S. Kalpakam and K. P. Sapna. A lost sales inventory system with supply uncertainty. Computers and Mathematics with Applications, 33(3):81–93, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002398.

Kuchkina:1997:OCP

[KS97b] N. Kuchkina and L. Shaikhet. Optimal control problem for nonlinear stochastic difference second kind volterra equations. Computers and Mathematics with Applications, 34(9):65–73, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001892.

Khabibrakhmanov:1998:UGL

[KS98a] I. K. Khabibrakhmanov and D. Summers. The use of generalized Laguerre polynomials in spectral methods for nonlinear differential equations. Computers and Mathematics with Applications, 36(2):65–70, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001175.

Kuchkina:1998:OCV

[KS98b] N. Kuchkina and L. Shaikhet.

Optimal control of Volterra type stochastic difference equations. Computers and Mathematics with Applications, 36(10–12):251–259, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800266.

Kamei:1999:PNM

[KS99] M. Kamei and Y. Sugasawa. Petri net model and performance evaluation of an autonomous decentralized network system. Computers and Mathematics with Applications, 38(11–12):165– 173, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002953.

Kersten:1990:MDM

[KSK90] G. E. Kersten, S. Szpakowicz, and Z. Koperczak. Modelling of decision making for discrete processes in dynamic environments. Computers and Mathematics with Applications, 20(9–10):29–43, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190901106.

Komori:1994:SID

[KSM94] Y. Komori, Y. Saito, and T. Mitsui. Some issues in discrete approximate solution for stochastic differential equations. Computers and Mathematics with Applications, 28(10–12):269–278, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001979.

Kalaba:1990:NAS

[KT90] R. Kalaba and L. Tesfatsion. Nonlocal automated sensitivity analysis. Computers and Mathematics with Applications, 20(2):53–65, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090241B.

Kim:1992:QNM

[KT92] S. Kim and R. P. Tewarson. A quasi-Newton method for solving nonlinear algebraic equations. Computers and Mathematics with Applications, 24(4):93–97, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900116.

Kim:1995:CQG

[KT95] S. Kim and R. P. Tewarson.

The convergence of quasi-Gauss–Newton methods for nonlinear problems. Computers and Mathematics with Applications, 29(8):27–38, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500027V.

Kalla:1996:AFG

[KT96] S. L. Kalla and Vu Kim Tuan. Asymptotic formulas for generalized elliptictype integrals. Computers and Mathematics with Applications, 32(4):49–55, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001241.

Kao:1997:UBS

[KT97] Chiang Kao and Huey-Chin Tang. Upper bounds in spectral test for multiple recursive random number generators with missing terms. Computers and Mathematics with Applications, 33 (4):119–125, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000126.

Kao:1998:SET

[KT98] Chiang Kao and Hui-Chin Tang. Several extensively tested multiple recursive random number generators. Computers and Mathematics with Applications, 36(6): 129–136, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001667.

Kida:1997:MEH

[KTO97] S. Kida, E. Trimandalawati, and S. Ogawa. Matrix expression of Hermite interpolation polynomials. Computers and Mathematics with Applications, 33(11):11–13, June 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000837.

Kalaba:1991:ALA

[KU91] Robert E. Kalaba and Firdaus E. Udwadia. An adaptive learning approach to the identification of structural and mechanical systems. Computers and Mathematics with Applications, 22(1):67–75, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190026Z.

Kawata:1999:ANS

[KU99] Y. Kawata and S. Ueno. Analytical and numerical simulation of satellite image from space. Computers and Mathematics with Applications, 37 (11–12):123–131, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001510.

Kun:1990:ICS

[Kun90] I. Kun. Iteration convergence speed estimation using the LIL. Computers and Mathematics with Applications, 19(1):97–100, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090087Z.

Kunen:1995:SSA

[Kun95] K. Kunen. The shortest single axioms for groups of exponent 4. Computers and Mathematics with Applications, 29(2):1–12, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002124.

Kupershmidt:1991:FIU

[Kup91] B. A. Kupershmidt. Forminvariance upon relativisation of the Hamiltonian structures of fluids is not universal. Computers and Mathematics with Applications, 21(5):95–99, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190219T.

Kurano:1991:FCA

[Kur91] Masami Kurano. Functional characterization for average cost Markov decision processes with Doeblin’s conditions. Computers and Mathematics with Applications, 21 (11–12):57–63, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190107F.

Kuusela:1994:ADD

[Kuu94]  T. Kuusela.  AC-driven

damped Toda lattice. Computers and Mathematics with Applications, 28(10–12):327– 351, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002002.

Karoui:1994:CSS

[KV94a] A. Karoui and R. Vaillancourt. Computer solutions of state-dependent delay differential equations. Computers and Mathematics with Applications, 27(4): 37–51, February 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900531.

Karoui:1994:FBW

[KV94b] A. Karoui and R. Vaillancourt. Families of biorthogonal wavelets. Computers and Mathematics with Applications, 28(4):25–39, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001243.

Karoui:1995:MTC

[KV95] A. Karoui and R. Vaillancourt. McClellan transformation and the construction of biorthogonal wavelet bases of L^{2}R^{2}. Computers and Mathematics with Applications, 29(12):13–26, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000543.

Kovacevic-Vujcic:1999:TSM

[KVCA^{+}99] V. V. Kovacević-Vujcić, M. M. Cangalović, M. D. Asić, L. Ivanović, and M. Drazić. TABU search methodology in global optimization. Computers and Mathematics with Applications, 37(4–5):125– 133, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000644.

Kao:1998:RNG

[KW98] Chiang Kao and J. Y. Wong. Random number generators with long period and sound statistical properties. Computers and Mathematics with Applications, 36 (3):113–121, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001333.

Kwok:1992:SAS

[Kwo92] Yue-Kuen Kwok. Stability analysis of six-point finite difference schemes for the constant coefficient convectivediffusion equation. Computers and Mathematics with Applications, 23(12):3–11, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290088Y.

Kurano:1999:TAR

[KYNY99] M. Kurano, M. Yasuda, J. Nakagami, and Y. Yoshida. The time average reward for some dynamic fuzzy systems. Computers and Mathematics with Applications, 37 (11–12):77–86, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001455.

Kimura:1992:SRA

[KYO92] Mitsuhiro Kimura, Shigeru Yamada, and Shunji Osaki. Software reliability assessment for an exponential-S-shaped reliability growth phenomenon. Computers and Mathematics with Applications, 24(1–2):71–78, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290230F.

Kapur:1995:ORR

[KZ95] D. Kapur and H. Zhang. An overview of rewrite rule laboratory (RRL). Computers and Mathematics with Applications, 29(2):91–114, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400218A.

Lee:1996:CFD

[LA96] H. Y. Lee and M. J. Ahn. The convergence of the fully discrete solution for the roseneau equation. Computers and Mathematics with Applications, 32(3):15–22, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001101.

Lagarias:1991:EML

[Lag91]  Jeffrey C. Lagarias.  Ef-

fects of misspecification of lag structure in certain twovariable distributed lag models. Computers and Mathematics with Applications, 22(10):3–23, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190188A.

Lai:1997:EPV

[Lai97] C.-H. Lai. An extension of Purcell’s vector method with applications to panel element equations. Computers and Mathematics with Applications, 33(7):101–114, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700045X.

Lalli:1994:OTN

[Lal94]  B. S. Lalli.  Oscillation

theorems for neutral difference equations. Computers and Mathematics with Applications, 28(1–3):191– 202, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001073.

Lardner:1990:SNS

[Lar90]  R. W. Lardner.  Stabil-

ity of the numerical solution of a parabolic system with integral subsidiary conditions. Computers and Mathematics with Applications, 19(12):41–46, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090249J.

Lasser:1994:RTP

[Las94] D. Lasser. Rational tensor product Bézier volumes. Computers and Mathematics with Applications, 28(8): 49–62, October 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001693.

Lasser:1995:RTP

[Las95] D. Lasser. Rational tensor product Bézier volumes. Computers and Mathematics with Applications, 29(8): 95–108, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500033U.

Laske:1996:KTA

[Las96] O. Laske. Knowledge technology and the arts: a personal view. Computers and Mathematics with Applications, 32(1):85–88, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000909.

Layton:1993:TLD

[Lay93] W. Layton. A two-level discretization method for the Navier–Stokes equations. Computers and Mathematics with Applications, 26(2): 33–38, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390318P.

Layton:1999:WIN

[Lay99] W. Layton. Weak imposition of “no-slip” conditions in finite element methods. Computers and Mathematics with Applications, 38(5– 6):129–142, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002205.

Li:1994:MDP

[LB94] D. Li and J. A. Burger. Multilevel dynamic programming for general multiple linearquadratic control in discretetime systems. Computers and Mathematics with Applications, 27(9–10):59–66, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901244.

Lin:1994:SAN

[LC94] J.-F. Lin and S.-J. Chen.

Scheduling algorithm for nonpreemptive multiprocessor tasks. Computers and Mathematics with Applications, 28(4):85–92, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001294.

Lin:1996:SCT

[LC96a] Yi-Zhong Lin and Sui Sun Cheng. Stability criteria for two partial difference equations. Computers and Mathematics with Applications, 32 (7):87–103, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001587.

Lu:1996:FPL

[LC96b] Tzon Tzer Lu and Sui Sun Cheng. Form preserving linear integral operators. Computers and Mathematics with Applications, 31(8):117–134, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000363.

Lau:1997:EWF

[LC97a] P. Lau and C. W. Chan. An enhanced weighted fuzzy reasoning algorithm for engineering applications. Computers and Mathematics with Applications, 34(1):81–87, July 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000990.

Liu:1997:EPA

[LC97b] Zhiyong Liu and D. W. Cheung. Efficient parallel algorithm for dense matrix LU decomposition with pivoting on hypercubes. Computers and Mathematics with Applications, 33(8):39–50, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000527.

Li:1998:ATP

[LC98a] Wan Tong Li and Sui Sun Cheng. Asymptotic trichotomy for positive solutions of a class of odd order nonlinear neutral difference equations. Computers and Mathematics with Applications, 35(8):101–108, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000480.

Li:1998:OCN

[LC98b] Wan Tong Li and Sui Sun Cheng. Oscillation criteria for a nonlinear difference equation. Computers and Mathematics with Applications, 36(8):87–94, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001850.

See errata [Ano99-243].

Liu:1999:DAI

[LCC99] P. Liu, Lansun Chen, and X. Cui. A discrete analogue of an integrodifferential equation. Computers and Mathematics with Applications, 37(3):41–55, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000450.

Lee:1994:DRL

[LCL94] Jiann-Der Lee, Chin-Hsing Chen, and Jau-Yien Lee. Determination of robot location using generalized cylindrical object shapes. Computers and Mathematics with Applications, 28(8):87–98, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001723.

Lai:1998:DEA

[LCP98] C.-H. Lai, A. M. Cuffe, and K. A. Pericleous. A defect equation approach for the coupling of subdomains in domain decomposition methods. Computers and Mathematics with Applications, 35(6):81–94, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000212.

Li:1998:URB

[LCSN98] D. Li, D. Cheung, Xuemei Shi, and V. Ng. Uncertainty reasoning based on cloud models in controllers. Computers and Mathematics with Applications, 35 (3):99–123, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002824.

Lanza:1990:AMG

[LD90] Antonio Lanza and Mark R. Dubal. An application of the multi-grid method to the construction of initial data for brill waves. Computers and Mathematics with Applications, 19(10):77–85, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090359R.

Long:1993:PSN

[LD93] Nguyen Thanh Long and Alain Pham Ngoc Dinh. Periodic solution of a nonlinear parabolic equation involving Bessel’s operator. Computers and Mathematics with Applications, 25(5):11–18, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390194Z.

Long:1995:PSN

[LD95] N. T. Long and A. P. N.

Dinh. Periodic solutions of a nonlinear parabolic equation associated with the penetration of a magnetic field into a substance. Computers and Mathematics with Applications, 30(1):63–78, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500068A.

Lellouche:1994:BCB

[LDD94] J.-M. Lellouche, J.-L. Devenon, and I. Dekeyser. Boundary control of Burgers’ equation — a numerical approach. Computers and Mathematics with Applications, 28(5):33–44, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001383.

Leble:1998:EBD

[Leb98] S. B. Leble. Elementary and binary Darboux transformations at rings. Computers and Mathematics with Applications, 35(10):73–81, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000741.

Lee:1991:F

[Lee91] E. S. Lee. Foreword. Computers and Mathematics with Applications, 21(11–12):vii– viii, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190100I.

Lee:1993:RSS

[Lee93] Min Ho Lee. The reduction of solutions of some integrable partial differential equations. Computers and Mathematics with Applications, 25(8):95–100, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390175U.

Lee:1994:FDM

[Lee94]  H. Y. Lee.  Fully dis-

crete methods for the nonlinear Schrödinger equation. Computers and Mathematics with Applications, 28 (6):9–24, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001480.

Lee:1996:FGA

[Lee96] Jiann-Der Lee. A fast geometrical approach to camera extrinsic parameters. Computers and Mathematics with Applications, 32 (12):93–100, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219600209X.

Lehmann:1992:SN

[Leh92] Fritz Lehmann. Semantic networks. Computers and Mathematics with Applications, 23 (2–5):1–50, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901355.

Leinbach:1990:CAM

[Lei90] L. C. Leinbach. Computer algebra and the microcomputer. Computers and Mathematics with Applications, 19(3):43–51, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090040Q.

Leibowitz:1996:RAN

[Lei96] G. M. Leibowitz. Remarks on the article “Norm inequalities which yield inclusions for Euler sequence spaces”. Computers and Mathematics with Applications, 31(11):79–81, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000648.

See [MSR95].

Leonenko:1990:SMG

[Leo90] Nick N. Leonenko. Sojourns of multidimensional Gaussian random fields with dependent components. Computers and Mathematics with Applications, 19(1):101–108, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900882.

Leonard:1996:LSL

[Leo96] N. Leonard III. “Legacy: San Lazaro” the integration of composition, performance, and computer programming. Computers and Mathematics with Applications, 32(1):89–92, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000910.

Levinson:1992:PAR

[Lev92] Robert Levinson. Pattern associativity and the retrieval of semantic networks. Computers and Mathematics with Applications, 23(6–9): 573–600, March/May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901252.

Lew:1991:FDT

[Lew91a] Art Lew. Fuzzy decision tables for expert systems. Computers and Mathematics with Applications, 21(11– 12):111–116, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190112H.

Lew:1991:RDO

[Lew91b] Art Lew. Reduction of dimensionality of an optimal decision table conversion algorithm. Computers and Mathematics with Applications, 21 (11–12):41–49, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190105D.

Lew:1994:DCS

[Lew94] A. Lew. A discrete and continuous systems simulator: applications of PNT-BLSIM to manufacturing systems. Computers and Mathematics with Applications, 27(9–10):37–43, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219490121X.

Luo:1991:SDM

[LF91] Jenn-Ching Luo and Morton B. Friedman. A study on decomposition methods. Computers and Mathematics with Applications, 21(8): 79–84, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900537.

Liu:1998:SCT

[LF98] X. Liu and X. Fu. Stability criteria in terms of two measures for discrete systems. Computers and Mathematics with Applications, 36(10–12):327–337, November/December 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800333.

Li:1999:OCS

[LF99] Wan-Tong Li and Xian-Ling Fan. Oscillation criteria for second-order nonlinear difference equations with damped term. Computers and Mathematics with Applications, 37(6):17–30, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000735.

Lopez-Fenner:1997:TAN

[LFP97] J. López-Fenner and M. Pinto.

(h, k)-trichotomies and asymptotics of nonautonomous difference systems. Computers and Mathematics with Applications, 33(10):105–124, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000801.

Liu:1999:MUL

[LG99] Xinzhi Liu and Dajun Guo.

Method of upper and lower solutions for second-order impulsive integro-differential equations in a Banach space. Computers and Mathematics with Applications, 38 (3–4):213–223, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001960.

Lisio:1998:CSM

[LGP98] R. Di Lisio, E. Grenier, and M. Pulvirenti. The convergence of the SPH method. Computers and Mathematics with Applications, 35 (1–2):95–102, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002605.

Li:1991:EDP

[LH91] Duan Li and Yacov Y. Haimes. Extension of dynamic programming to nonseparable dynamic optimization problems. Computers and Mathematics with Applications, 21(11–12):51– 56, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190106E.

Lew:1994:DPD

[LH94] A. Lew and R. Halverson, Jr. Dynamic programming, decision tables, and the Hawaii parallel computer. Computers and Mathematics with Applications, 27 (9–10):121–127, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901309.

Lee:1997:RCB

[LH97] Narn-Yih Lee and Tzonelih Hwang. Remarks on Change– Buehrer access control scheme in a hierarchy. Computers and Mathematics with Applications, 34(12):11–12, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122197002289.

Lew:1999:DPF

[LH99] A. Lew and R. Halverson, Jr. Dynamic programming on a functional memory computer. Computers and Mathematics with Applications, 37 (11–12):17–22, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001388.

Lu:1996:PPG

[LHC96] Tan-Chun Lu, Yu-Song Hou, and Rong-Jaye Chen. A parallel Poisson generator using parallel prefix. Computers and Mathematics with Applications, 31(3):33–42, February 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122195002049.

Li:1996:NCB

[Li96a] L. Li. A new complexity bound for the least-squares problem. Computers and Mathematics with Applications, 31(12):15–16, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000727.

Li:1996:NII

[Li96b] Z. Li. A note on immersed interface method for threedimensional elliptic equations. Computers and Mathematics with Applications, 31(3):9–17, February 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002022.

Li:1996:SIM

[Li96c] Zi-Cai Li. Splitting-integrating methods for the image transformations T and T^{−1}T. Computers and Mathematics with Applications, 32(3): 39–60, August 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001125.

Li:1997:QUC

[Li97] J. Li. Quasioptimal uniformly convergent finite element methods for the elliptic boundary layer problem. Computers and Mathematics with Applications, 33 (10):11–22, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000734.

Li:1998:GPE

[Li98a] J. Li. Global pointwise error estimates for uniformly convergent finite element methods for the elliptic boundary layer problem. Computers and Mathematics with Applications, 36(1):59–67, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001096.

Li:1998:RFE

[Li98b] J. Li. A robust finite element method for a singularly perturbed elliptic problem with two small parameters. Computers and Mathematics with Applications, 36 (7):91–110, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001758.

Li:1998:ACP

[Li98c] K. Li. Average-case performance analysis of an approximation algorithm for maximum subset sum using recurrence relations. Computers and Mathematics with Applications, 36(6):63–75, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800162X.

Li:1998:NDT

[Li98d] Zi-Cai Li. New discrete techniques for 3D image transformations. Computers and Mathematics with Applications, 36(4):77–109, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001424.

Li:1999:SPN

[Li99a] Longtu Li. Stability properties of nonlinear difference equations and conditions for boundedness. Computers and Mathematics with Applications, 38(2):29–35, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001790.

Li:1999:EMS

[Li99b] Luoluo Li. Estimation for matrix singular values. Computers and Mathematics with Applications, 37(9):9–15, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900108X.

Li:1999:FMG

[Li99c] Rong-Jun Li. Fuzzy method in group decision making. Computers and Mathematics with Applications, 38(1): 91–101, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001728.

Li:2010:CUC

[Li10]  Jichun Li.  Corrigendum

to “Uniformly convergent finite element methods for singularly perturbed elliptic boundary value problems I: Reaction–diffusion type” [Comput. Math. Appl. 35 (3) (1998) 57–70]. Computers and Mathematics with Applications, 59(10):3374–3376, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122110002026.

Liaw:1995:PAU

[Lia95]  Horng-Twu Liaw.  Pass-

word authentications using triangles and straight lines. Computers and Mathematics with Applications, 30 (9):63–71, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500147Q.

Liaw:1999:BCC

[Lia99a] Horng-Twu Liaw. Broadcasting cryptosystem in computer networks. Computers and Mathematics with Applications, 37(6):85–87, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000784.

Liaw:1999:DAC

[Lia99b] Horng-Twu Liaw. The design of an arithmetic coding mechanism to determine relationships in a hierarchy. Computers and Mathematics with Applications, 37(9):61–71, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001145.

Lin:1991:GSS

[Lin91a] Chau-Jy Lin. Generating subsets on a systolic array. Computers and Mathematics with Applications, 21 (2–3):103–109, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190087K.

Lin:1991:DTV

[Lin91b] Lai-Jiu Lin. Duality theorems of vector-valued n-set functions. Computers and Mathematics with Applications, 21 (11–12):165–175, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190118N.

Lin:1991:TST

[Lin91c] Shu-Kun Lin. Time symmetry and thermodynamics. Computers and Mathematics with Applications, 22(12):67–76, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190149X.

Lin:1994:SAD

[Lin94] Chau-Jy Lin. A systolic algorithm for dynamic programming. Computers and Mathematics with Applications, 27(1):1–10, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900019.

Lin:1995:MDW

[Lin95] E.-B. Lin. Multi-dimensional wavelet Stieltjes transforms. Computers and Mathematics with Applications, 30(3– 6):353–359, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001123.

Lin:1996:SAS

[Lin96]  Chau-Jy Lin.  A sys-

tolic algorithm for solving dense linear systems. Computers and Mathematics with Applications, 32 (12):77–91, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002088.

Liu:1996:DNS

[Liu96a]  Jen-Chwan Liu.  Duality

for nondifferentiable static multiobjective variational problems involving generalized (F, ρ)-convex functions. Computers and Mathematics with Applications, 31(12): 77–89, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000788.

Liu:1996:ODG

[Liu96b] Jen-Chwan Liu. Optimality and duality for generalized fractional programming involving nonsmooth (F, ρ)- convex functions. Computers and Mathematics with Applications, 32(2):91–102, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600106X.

Liu:1997:DCP

[Liu97] Baoding Liu. Dependentchance programming: a class of stochastic optimization. Computers and Mathematics with Applications, 34 (12):89–104, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219700237X.

Liu:1998:SNE

[Liu98] Baoding Liu. Stackelberg– Nash equilibrium for multilevel programming with multiple followers using genetic algorithms. Computers and Mathematics with Applications, 36(7):79–89, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001746.

Lee:1998:PPS

[LJ98] J.-Y. Lee and K. Jeong. A parallel Poisson solver using the fast multipole method on networks of workstations. Computers and Mathematics with Applications, 36(4): 47–61, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001400.

Liu:1993:RPG

[LJL^{+}93] Si-Zhang Liu, De-Jun Jin, Sang-Ho Lee, Hai-Lou Liu, Xi-Guan Lu, and Dazhi Zhang. The radical product of graceful trees. Computers and Mathematics with Applications, 26(10):89–93, November 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390028T.

Lakshmikantham:1994:MFF

[LK94] V. Lakshmikantham and B. Kaymakçalan. Monotone flows and fixed points for dynamic systems on time scales. Computers and Mathematics with Applications, 28 (1–3):185–189, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001065.

Laih:1997:SCE

[LK97] Chi-Sung Laih and Wen-Chung Kuo. Speeding up the computations of elliptic curves cryptoschemes. Computers and Mathematics with Applications, 33(5): 29–36, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000175.

Liang:1999:CCD

[LK99] Jihua Liang and K. Keimel.

Compact continuous L-domains. Computers and Mathematics with Applications, 38(1): 81–89, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001716.

Lee:1995:CSS

[LKC95] E.-Y Lee, K. J. Kim, and U. J. Choi. A construction of the simplest super pseudorandom permutation generator. Computers and Mathematics with Applications, 29(8):19–25, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500026U.

Lo:1997:ACP

[LKE97] S. Lo, C. Kee, and P. Enge.

Algorithms for crossover point determination. Computers and Mathematics with Applications, 33(11):117– 131, June 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000898.

Liu:1999:TTD

[LKT99] J. C. Liu, Y. Kimura, and K. Tanaka. Three types dual model for minimax fractional programming. Computers and Mathematics with Applications, 38(7–8):143– 155, October 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900245X.

Li:1990:MCN

[LL90a] R. J. Li and E. S. Lee. Multi-criteria de Novo programming with fuzzy parameters. Computers and Mathematics with Applications, 19(5):13–20, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900974.

Liebermann:1990:SQA

[LL90b] P. Liebermann and R. Liebermann. Symmetry in question and answer sequences in music. Computers and Mathematics with Applications, 19(7):59–66, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090194O.

Lirov:1990:SBL

[LL90c] V. Lirov and Y. Lirov. A subject bibliography of logic programming applications in control and decision support systems. Computers and Mathematics with Applications, 20(9–10):141–179, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901184.

Li:1991:EMF

[LL91a] R. J. Li and E. Stanley Lee. An exponential membership function for fuzzy multiple objective linear programming. Computers and Mathematics with Applications, 22(12):55–60, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190147V.

Lin:1991:LAC

[LL91b] S. S. Lin and F. C. Lin.

An o(log n) algorithm for computing periodic continued fractions and its applications. Computers and Mathematics with Applications, 21(2–3):1–6, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190074E.

Liaw:1992:EHS

[LL92] Horng-Twu Liaw and Chin-Laung Lei. An efficient hierarchy scheme for determining relationships in information systems. Computers and Mathematics with Applications, 24(11):71–78, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/089812219290033E.

Lee:1994:SHS

[LL94a] C. S. Lee and G. Leitmann. A stabilizing harvesting strategy for an uncertain model of an ecological system. Computers and Mathematics with Applications, 27(9–10):199–212, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901384.

Lew:1994:F

[LL94b] Art Lew and E. Stanley Lee. Foreword. Computers and Mathematics with Applications, 27(9–10):xi, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901163.

Leitmann:1996:SUC

[LL96] G. Leitmann and C. S. Lee. Stabilization of an uncertain competing species system. Computers and Mathematics with Applications, 31(10):101–129, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000557.

Lee:1999:BHS

[LL99] C. S. Lee and G. Leitmann.

A bounded harvest strategy for an ecological system in the presence of uncertain disturbances. Computers and Mathematics with Applications, 37 (11–12):141–157, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001534.

Layton:1994:TLN

[LLP94] W. Layton, H. W. J.

Lenferink, and J. S. Peterson. A two-level newton, finite element algorithm for approximating electrically conducting incompressible fluid flows. Computers and Mathematics with Applications, 28 (5):21–31, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001375.

Li:1990:ETM

[LLR^{+}90] Baoxin Li, A. G. Law, H. Raafat, P. H. Nguyen, and Y.-F. Yan. Eigenvalues of tridiagonal matrices: an alternative to Givens’ method. Computers and Mathematics with Applications, 19(4):89–94, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090140F.

Lopez-Marcos:1991:USN

[LM91] J. C. Lopez-Marcos. An upwind scheme for a nonlinear hyperbolic integro-differential equation with integral boundary condition. Computers and Mathematics with Applications, 22(11):15–28, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900308.

Liu:1996:NED

[LM96a] Y. Liu and D. A. Murio. Numerical experiments in 2-D IHCP on bounded domains. Part I: the “interior” cube problem. Computers and Mathematics with Applications, 31(1):15–32, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001807.

Llanas:1996:FPP

[LM96b] B. Llanas and C. Moreno. Finding the projection on a polytope: an iterative method. Computers and Mathematics with Applications, 32(8):33–39, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001642.

Liska:1995:NTL

[LMW95] R. Liska, L. Margolin, and B. Wendroff. Nonhydrostatic two-layer models of incompressible flow. Computers and Mathematics with Applications, 29(9):25–37, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500035W.

Li:1998:UCF

[LN98] J. Li and I. M. Navon.

Uniformly convergent finite element methods for singularly perturbed elliptic boundary value problems i: Reaction–diffusion type. Computers and Mathematics with Applications, 35(3): 57–70, February 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002794.

See corrigendum [Li10].

Li:1999:GUC

[LN99] Jichun Li and I. M. Navon.

A global uniformly convergent finite element method for a quasi-linear singularly perturbed elliptic problem. Computers and Mathematics with Applications, 38(5– 6):197–206, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002266.

Lustman:1992:SOD

[LNG92] L. Lustman, B. Neta, and W. Gragg. Solution of ordinary differential initial value problems on an Intel Hypercube. Computers and Mathematics with Applications, 23(10):65–72, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290056N.

Le:1990:SLS

[LNM90] S. Le, R. Nussinov, and J. V. Maizel, Jr. Studies of local stability in histone, UsnRNA and globin precursor mRNAs around transcription termination sites. Computers and Mathematics with Applications, 20(4–6):49–56, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190903139.

Looges:1993:OGA

[LO93] Peter J. Looges and Stephan Olariu. Optimal greedy algorithms for indifference graphs. Computers and Mathematics with Applications, 25(7): 15–25, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390308I.

Liu:1999:DFC

[LO99] Baoding Liu and T. Odanaka.

Dynamic fuzzy criterion model for reservoir operations and a case study. Computers and Mathematics with Applications, 37 (11–12):65–75, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001443.

Long:1996:NBD

[LOD96] N. T. Long, E. L. Ortiz, and A. Pham Ngoc Dinh. A nonlinear Bessel differential equation associated with Cauchy conditions. Computers and Mathematics with Applications, 31(10):131–139, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000569.

Logemann:1996:SMS

[Log96] G. W. Logemann. Software for the Mathews Stanford radio drum. Computers and Mathematics with Applications, 32(1):93–96, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000922.

Long:1995:TMT

[Lon95] Yong-Hong Long. Thinning and multinomial thinning of point processes. Computers and Mathematics with Applications, 30(8):1–4, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500131H.

Lopez:1993:TSB

[Lop93] L. Lopez. Two-step boundary value methods in the solution of ODEs. Computers and Mathematics with Applications, 26(1):91–100, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390088D.

Lin:1995:OPC

[LOP95] R. Lin, S. Olariu, and G. Pruesse. An optimal path cover algorithm for cographs. Computers and Mathematics with Applications, 30(8):75–83, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500139P.

Los:1989:ILS

[Los89a] C. A. Los. Identification of a linear system from inexact data: a three-variable example. Computers and Mathematics with Applications, 17 (8–9):1285–1304, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122189900965.

See commentary [Jay92, Zel92] and reply [Los92].

Los:1989:PLS

[Los89b]  C. A. Los.  The preju-

dices of least squares, principal components and common factors schemes. Computers and Mathematics with Applications, 17(8–9):1269– 1283, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122189900953.

See commentary [Jay92, Zel92] and reply [Los92].

Los:1992:RJZ

[Los92] Cornelis A. Los. Reply to E.

T. Jaynes’ and A. Zellner’s comments on my two articles. Computers and Mathematics with Applications, 24(8–9): 277–288, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290206W.

Lee:1999:PDS

[LOS99] Poh-Aun Lee, Seng-Huat Ong, and H. M. Srivastava. Probabilistic derivation of some generating functions for the Laguerre polynomials. Computers and Mathematics with Applications, 38 (7–8):21–29, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002345.

Liu:1999:ASS

[LP99] K. M. Liu and C. K. Pan. The automatic solution to systems of ordinary differential equations by the tau method. Computers and Mathematics with Applications, 38(9– 10):197–210, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002758.

Lloyd:1996:CIC

[LPR96] N. G. Lloyd, J. M. Pearson, and V. A. Romanovsky. Computing integrability conditions for a cubic differential system. Computers and Mathematics with Applications, 32(10):99–107, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196001885.

Li:1995:LPI

[LQ95a] S. Li and Mengzhao Qin. Lie–Poisson integration for rigid body dynamics. Computers and Mathematics with Applications, 30(9):105–118, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500150W.

Li:1995:NLP

[LQ95b] S.-T. Li and Mengzhao Qin. A note for Lie–Poisson Hamilton–Jacobi equation and Lie–Poisson integrator. Computers and Mathematics with Applications, 30(7): 67–74, October 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500126J.

Lirov:1990:IID

[LR90] Y. Lirov and S. Ravikumar. Intelligent interaction in diagnostic expert systems. Computers and Mathematics with Applications, 20 (9–10):111–123, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901162.

Leyva-Ramos:1993:NLP

[LR93a] J. Leyva-Ramos. A new look at partial fraction expansion of transfer function matrices from a computational viewpoint. Computers and Mathematics with Applications, 26(3):27–35, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193901077.

Lin:1993:FCF

[LR93b] Qun Lin and J. G. Rokne. A family of centered forms for rational functions. Computers and Mathematics with Applications, 25(4):87–93, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/089812219390251P.

Lancaster:1996:SSV

[LR96a] P. Lancaster and P. Rózsa. The spectrum and stability of a vibrating rail supported by sleepers. Computers and Mathematics with Applications, 31(4–5):201– 213, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002332.

Lin:1996:IAH

[LR96b] Qun Lin and J. G. Rokne. Interval approximation of higher order to the ranges of functions. Computers and Mathematics with Applications, 31(7):101–109, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600020X.

Laita:1995:FMV

[LRLR95] L. M. Laita, B. Ramírez, L. De Ledesma, and A. Riscos. A formal model for verification of dynamic consistency of KBSs. Computers and Mathematics with Applications, 29(5):81–96, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400250O.

Lister:1994:ACN

[LS94] L. Lister and J. Schaeffer. An analysis of the conspiracy numbers algorithm. Computers and Mathematics with Applications, 27(1): 41–64, January 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900051.

Luchko:1995:ESC

[LS95] Yu. F. Luchko and H. M. Srivastava. The exact solution of certain differential equations of fractional order by using operational calculus. Computers and Mathematics with Applications, 29(8):73–85, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500031S.

Lasser:1996:BIP

[LS96a] D. Lasser and T. Stüttgen. Boundary improvement of piecewise linear interpolants defined over Delaunay triangulations. Computers and Mathematics with Applications, 32(10):43–58, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219600185X.

Latifi:1996:SIJ

[LS96b] S. Latifi and P. K. Srimani. Scheduling of independent jobs in star graph networks. Computers and Mathematics with Applications, 32(6):45–55, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001435.

Lakshmikantham:1998:SMI

[LS98a] V. Lakshmikantham and S. Sivasundaram. Stability of moving invariant sets and uncertain dynamic systems on time scales. Computers and Mathematics with Applications, 36(10–12):339–346, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800345.

Lasser:1998:IIP

[LS98b] D. Lasser and T. Stüttgen. Interior improvement of piecewise linear interpolants defined over Delaunay triangulations. Computers and Mathematics with Applications, 36(4):21–36, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001382.

Li:1999:SMD

[LS99] Jian Li and W. W. Sun.

Spectral method of decoupling the vorticity and stream function for the incompressible fluid flows. Computers and Mathematics with Applications, 37(1):35–40, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002399.

Lam:1995:CBR

[LSH95] K.-Y. Lam, K. Sung, and L. C.-K. Hui. A cardinalised binary representation for exponentiation. Computers and Mathematics with Applications, 30(8):33–39, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500135L.

Lepper:1998:PSC

[LSH98] J. Lepper, U. Schnell, and K. R. G. Hein. Parallelization of a simulation code for reactive flows on the intel paragon. Computers and Mathematics with Applications, 35(7):101–109, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000376.

Liu:1994:RCO

[LT94] C. H. Liu and W. H. Tsai. Recognizing cylindrical objects by single camera views. Computers and Mathematics with Applications, 27(8): 101–117, April 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900701.

Li:1996:SVF

[LT96] M. Li and T. Tang. Steady viscous flow in a triangular cavity by efficient numerical techniques. Computers and Mathematics with Applications, 31(10):55–65, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000521.

Liu:1999:SDM

[LT99] Chin-Sung Liu and Ching-Hung Tseng. Space-decomposition minimization method for large-scale minimization problems. Computers and Mathematics with Applications, 37 (7):73–88, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000887.

Longo:1996:DNM

[LTB96] E. Longo, G. Teppati, and N. Bellomo. Discretization of nonlinear models by sinc collocation-interpolation methods. Computers and Mathematics with Applications, 32(4):65–81, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001265.

Lee:1998:OCA

[LTC98] H. W. J. Lee, K. L. Teo, and X. Q. Cai. An optimal control approach to nonlinear mixed integer programming problems. Computers and Mathematics with Applications, 36(3):87–105, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800131X.

Lee:1998:SLF

[LTR98] H. W. J. Lee, K. L. Teo, and V. Rehbock. Suboptimal local feedback control for a class of constrained discrete time nonlinear control problems. Computers and Mathematics with Applications, 36(10–12):133–148, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800163.

Lu:1996:CAP

[Lu96] Zhengyi Lu. Computer aided proof for the global stability of Lotka–Volterra systems. Computers and Mathematics with Applications, 31 (12):49–59, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000752.

Lubinsky:1997:WNO

[Lub97]  D. S. Lubinsky.  The

weighted L_{p}-norms of orthonormal polynomials for Erdős weights. Computers and Mathematics with Applications, 33(1–2):151– 163, January 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002271.

Luo:1993:ECS

[Luo93a] Jenn-Ching Luo. Extended concept of stair-shape sparsity for the inverse of an asymmetric matrix. Computers and Mathematics with Applications, 25(5):113–122, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902049.

Luo:1993:IIP

[Luo93b] Jenn-Ching Luo. An incomplete inverse as a preconditioner for the conjugate gradient method. Computers and Mathematics with Applications, 25(2):73–79, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390224J.

Luo:1993:NCD

[Luo93c] Jenn-Ching Luo. A new class of decomposition for inverting asymmetric and indefinite matrices. Computers and Mathematics with Applications, 25(4):95–104, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390252Q.

Luo:1993:SDS

[Luo93d] Jenn-Ching Luo. A sparsity for decomposing a symmetric matrix. Computers and Mathematics with Applications, 25(5):83–90, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902016.

Luo:1998:SIC

[Luo98]  Zhong-Xuan Luo.  C^{1},

C^{2}-smooth interpolants on curved sides element. Computers and Mathematics with Applications, 35(5):125–130, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S089812219800011X.

Luo:1998:AAE

[LURL98] F.-L. Luo, R. Unbehauen, K. Reif, and Y.-D. Li. Analog approach for the eigendecomposition of positive definite matrices. Computers and Mathematics with Applications, 35(11):49–61, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000832.

Leavens:1992:SP

[LV92] Gary T. Leavens and Mike Vermeulen. 3x + 1 search programs. Computers and Mathematics with Applications, 24 (11):79–99, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290034F.

Lucka:1996:MPP

[LVV96] M. Lucká, M. Vajtersic, and E. Viktorinová. Massively parallel Poisson and QR factorization solvers. Computers and Mathematics with Applications, 31(4–5):19–26, February/March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500212X.

Lu:1991:SLE

[LW91] An Lu and E. L. Wachspress.

Solution of Lyapunov equations by alternating direction implicit iteration. Computers and Mathematics with Applications, 21(9):43–58, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190124M.

Li:1996:CBS

[LW96] Chong Li and G. A. Watson. A class of best simultaneous approximation problems. Computers and Mathematics with Applications, 31 (10):45–53, May 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600051X.

Lewinter:1997:HHL

[LW97] M. Lewinter and W. Widulski. Hyper-Hamilton laceable and caterpillar-spannable product graphs. Computers and Mathematics with Applications, 34(11):99–104, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812219700223X.

Li:1999:BSA

[LW99a] Chong Li and G. A. Watson. Best simultaneous approximation of an infinite set of functions. Computers and Mathematics with Applications, 37(2):1–9, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800248X.

Lu:1999:PDC

[LW99b] Zhengyi Lu and G. Wang. The positive definiteness of a class of polynomials from the global stability analysis of Lotka–Volterra systems. Computers and Mathematics with Applications, 38 (2):19–27, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001789.

Liaw:1993:DCK

[LWL93] H. T. Liaw, S. J. Wang, and C. L. Lei. A dynamic cryptographic key assignment scheme in a tree structure. Computers and Mathematics with Applications, 25(6):109–114, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390305F.

Li:1996:FS

[LWL96] Q. Li, P. Wang, and E. S.

Lee. r-fuzzy sets. Computers and Mathematics with Applications, 31(2):49–61, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001921.

Liang:1998:CNC

[LX98] Jin Liang and Tijun Xiao.

A characterization of norm continuity of propagators for second order abstract differential equations. Computers and Mathematics with Applications, 36(2):87–94, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001199.

Liu:1992:GDT

[LY92] Jen-Yu Liu and Hong-Sen Yan. Geometry design of trapezoidal threaded variable pitch lead screws. Computers and Mathematics with Applications, 23(1):65–73, January 1992. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290081R.

Li:1994:NND

[LY94] Horng Jaan Li and Cheh Chih Yeh. Nonoscillation in nonlinear difference equations. Computers and Mathematics with Applications, 28 (1–3):203–208, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001081.

Li:1998:OCS

[LY98] Horng-Jaan Li and Cheh-Chih Yeh. Oscillation criteria for second-order neutral delay difference equations. Computers and Mathematics with Applications, 36(10–12):123– 132, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800151.

Lin:1999:SPP

[LY99a] Rong S. Lin and Maria C. Yuang. Synchronization paradigm for protocol testing under multiparty configuration. Computers and Mathematics with Applications, 37(7):125–135, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000930.

Liu:1999:EOD

[LY99b] Jen-Chwan Liu and K. Yokoyama.

ϵ-Optimality and duality for multiobjective fractional programming. Computers and Mathematics with Applications, 37(8):119–128, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001054.

Lin:1998:EMH

[LYH98] Rong S. Lin, Maria C. Yuang, and Steen J. Hsu. Efficient minimization of homogeneous FSMs for fault diagnosis. Computers and Mathematics with Applications, 35(5):117–124, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000108.

Lee:1991:FMC

[LYL^{+}91] Jiann-Der Lee, Yuan-Chiuan You, Jau-Yien Lee, Chin-Hsing Chen, and Chia-Nian Shyi. A fast method for cylindrical object location determination. Computers and Mathematics with Applications, 22(2):49–57, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900412.

Lian:1995:DBZ

[LYL95] Wei-Cheng Lian, Cheh-Chih Yeh, and Horng-Jaan Li. The distance between zeros of an oscillatory solution to a half-linear differential equation. Computers and Mathematics with Applications, 29(8):39–43, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500028W.

Lytinen:1992:CDD

[Lyt92a] Steven L. Lytinen. Conceptual dependency and its descendants. Computers and Mathematics with Applications, 23(2–5):51–73, January/March 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901366.

Lytinen:1992:UBI

[Lyt92b]  Steven L. Lytinen.  A

unification-based, integrated natural language processing system. Computers and Mathematics with Applications, 23(6–9):403–418, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290115X.

Lee:1992:IDS

[LZ92] E. S. Lee and Qing Zhu.

An interval Dempster–Shafer approach. Computers and Mathematics with Applications, 24(7):89–95, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290157D.

Lee:1994:DFS

[LZ94] S.-H. Lee and D. Zhang.

Dual fuzzy similarity relation equations. Computers and Mathematics with Applications, 27(11):49–53, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900973.

Liu:1998:CPM

[LZ98] X. Liu and Shenghai Zhang.

A cell population model described by impulsive PDEs

— existence and numerical approximation. Computers and Mathematics with Applications, 36(8):1–11, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001783.

MacLeod:1993:HAN

[Mac93]  A. J. MacLeod.  High-

accuracy numerical values in the Gauss–Kuzmin continued fraction problem. Computers and Mathematics with Applications, 26(3):37–44, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901088.

Madych:1992:MEB

[Mad92] W. R. Madych. Miscellaneous error bounds for multiquadric and related interpolators. Computers and Mathematics with Applications, 24 (12):121–138, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290175H.

Magnitskii:1997:SUP

[Mag97]  N. A. Magnitskii.  Sta-

bilization of unstable periodic orbits of chaotic maps. Computers and Mathematics with Applications, 34(2–4): 369–372, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001338.

Mahmoudi:1990:PMF

[Mah90]  M. Mahmoudi.  Particle

modelling of fluid phenomena in 3-D. Computers and Mathematics with Applications, 20(3):25–44, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090028I.

Maini:1996:F

[Mai96a]  P. K. Maini.  Foreword.

Computers and Mathematics with Applications, 32(11): xiii, December 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197866692.

Maini:1996:SSP

[Mai96b] P. K. Maini. Spatial and spatiotemporal pattern formation in generalised turing systems. Computers and Mathematics with Applications, 32 (11):71–77, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001988.

Makridakis:1990:CMC

[Mak90] Ch. G. Makridakis. Cosine methods for a class of semilinear second-order wave equations. Computers and Mathematics with Applications, 19(10):19–34, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090355N.

Mauser:1991:SCB

[MAM91] Norbert Mauser, Anton Arnold, and Peter A. Markowich. The self-consistent Blochequation in one dimension. Computers and Mathematics with Applications, 21 (2–3):197–202, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190099P.

Manzini:1996:PSR

[Man96]  G. Manzini.  Perimeter

search in restricted memory. Computers and Mathematics with Applications, 32(7):37–45, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600154X.

Martin:1990:SIL

[Mar90] R. Martin. Single-interval learning by simile within a simulated Hebbian neural network. Computers and Mathematics with Applications, 20(4–6):217–226, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090329I.

Marshak:1991:SDM

[Mar91] Robert E. Marshak. From symmetry to dynamics in modern particle physics. Computers and Mathematics with Applications, 22(10): 33–43, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190190F.

Marty:1992:FSN

[Mar92] Robert Marty. Foliated semantic networks: Concepts, facts, qualities. Computers and Mathematics with Applications, 23(6–9):679– 696, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901296.

Marciniak:1994:AON

[Mar94] A. Marciniak. Arbitrary order numerical methods conserving integrals for solving dynamic equations. Computers and Mathematics with Applications, 28(10–12):33–43, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001847.

Marti:1995:DPF

[Mar95]  K. Marti.  Differentia-

tion of probability functions: the transformation method. Computers and Mathematics with Applications, 30(3– 6):361–382, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001131.

Marz:1996:CPL

[Mär96] R. März. Canonical projectors for linear differential algebraic equations. Computers and Mathematics with Applications, 31(4–5):121– 135, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002243.

Mastin:1992:LVM

[Mas92]  C. Wayne Mastin.  Lin-

ear variational methods and adaptive grids. Computers and Mathematics with Applications, 24(5–6):51–56, September 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290040O.

Matsumoto:1999:ANI

[Mat99] M. Matsumoto. The accuracy of numerical inversion of the Laplace transform applied to the timedependent reflection function from a semi-infinite atmosphere. Computers and Mathematics with Applications, 37 (11–12):113–122, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001509.

Malik:1998:MMC

[MATA98] M. N. Malik, M. Afzal, G. F.

Tariq, and N. Ahmed. Mathematical modeling and computer simulation of transient flow in centrifuge cascade pipe network with optimizing techniques. Computers and Mathematics with Applications, 36(4):63–76, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001412.

See errata [Ano99-242].

Mavronicolas:1993:ESS

[Mav93] Marios Mavronicolas. Efficiency of semi-synchronous versus asynchronous systems: Atomic shared memory. Computers and Mathematics with Applications, 25(2):81–91, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390225K.

Mayer:1999:NPG

[May99] M. K. Mayer. A network parallel genetic algorithm for the one machine sequencing problem. Computers and Mathematics with Applications, 37(3):71–78, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000474.

Mazumdar:1990:ESO

[Maz90] T. Mazumdar. Existence of solution of the operator equation AX — XB = Q with possibly unbounded a, B, Q. Computers and Mathematics with Applications, 19 (8–9):105–116, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090268O.

Milev:1990:SLI

[MB90] N. V. Milev and D. D. Bainov. Stability of linear impulsive differential equations. Computers and Mathematics with Applications, 20(12):35–41, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090162D.

Mukhopadhyay:1992:ETE

[MB92] B. Mukhopadhyay and R. K. Bera. Effect of temperature on the edge crack in orthotropic elastic half-plane. Computers and Mathematics with Applications, 24 (10):3–10, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290015A.

McIntosh:1993:NRR

[MB93] Aubrey McIntosh and Jennifer Brodbelt. The natural representation of real numbers. Computers and Mathematics with Applications, 25 (6):3–14, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902946.

McNabb:1994:MES

[MB94a] A. McNabb and L. Bass. Moment expressions for some diffusion-reaction problems. Computers and Mathematics with Applications, 27 (9–10):163–168, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901341.

Mukhopadhyay:1994:NAT

[MB94b] B. Mukhopadhyay and R. K.

Bera. Nonlinear analysis of thin homogeneous orthotropic elastic plates under large deflection and thermal loading. Computers and Mathematics with Applications, 28 (9):81–88, November 1994.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001804.

Misra:1992:RDS

[MBMS92] J. C. Misra, G. C. Bera, Subhas C. Misra, and S. Samanta. Remodelling of diaphysial surfaces of a long bone induced due to force-fitting of a medullary pin: an analytical study. Computers and Mathematics with Applications, 24(7):3–15, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290150G.

Meyer:1991:MBD

[MBV91] H. De Meyer, G. Vanden Berghe, and J. Vanthournout. Modified backward differentiation methods of the Adamstype based on exponential interpolation. Computers and Mathematics with Applications, 21(2–3):171–179, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190096M.

Margavio:1993:SAQ

[MC93] T. M. Margavio and J. M. Chin. A statistical approximation for queueing measures in a decision support system. Computers and Mathematics with Applications, 26(2):39–44, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390319Q.

Mora:1997:CGA

[MC97] G. Mora and Y. Cherruault. Characterization and generation of α-dense curves. Computers and Mathematics with Applications, 33(9): 83–91, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000679.

Marszalek:1999:DAT

[MC99] W. Marszalek and S. L.

Campbell. DAEs arising from traveling wave solutions of PDEs II. Computers and Mathematics with Applications, 37(1):15–34, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002387.

Mandal:1998:CFO

[MCG98a] C. A. Mandal, P. P. Chakrabarti, and S. Ghose. Complexity of fragmentable object bin packing and an application. Computers and Mathematics with Applications, 35 (11):91–97, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800087X.

Mandal:1998:PEV

[MCG98b] C. A. Mandal, P. P. Chakrabarti, and S. Ghose. A probabilistic estimator for the vertex deletion problem. Computers and Mathematics with Applications, 35(6):1–4, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000121.

Mandal:1998:SNR

[MCG98c] C. A. Mandal, P. P. Chakrabarti, and S. Ghose. Some new results in the complexity of allocation and binding in data path synthesis. Computers and Mathematics with Applications, 35(10):93–105, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000765.

McKenney:1996:AFM

[McK96] A. McKenney. An adaptation of the fast multipole method for evaluating layer potentials in two dimensions. Computers and Mathematics with Applications, 31(1):33–57, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500181W.

McLachlan:1995:CPS

[McL95] R. McLachlan. Comment on “Poisson schemes for Hamiltonian systems on Poisson manifolds”. Computers and Mathematics with Applications, 29(3):1, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002216.

See [ZQ94] and reply [ZQ95].

Melnikov:1997:NSB

[MDVK97] L. A. Melnikov, V. L. Derbov, I. V. Veshneva, and A. I. Konukhov. Numerical studies of beam and pulse propagation in lasers and nonlinear media: Transverse pattern dynamics and nonparaxial effects. Computers and Mathematics with Applications, 34 (7–8):881–909, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001788.

Medhin:1990:MSD

[Med90] N. G. Medhin. Minimizing sequences for differential inclusion problems in the presence of state constraints. Computers and Mathematics with Applications, 19(11):127–134, ???? 1990. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090155D.

Medina:1997:ABS

[Med97] R. Medina. The asymptotic behavior of the solutions of a Volterra difference equation. Computers and Mathematics with Applications, 34(1):19–26, July 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000953.

Meir:1993:ESI

[Mei93]  A. J. Meir.  The equa-

tions of stationary, incompressible magnetohydrodynamics with mixed boundary conditions. Computers and Mathematics with Applications, 25(12):13–29, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390182U.

Meinardus:1997:SRS

[Mei97] G. Meinardus. Some results in segmented approximation. Computers and Mathematics with Applications, 33(1– 2):165–180, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002283.

Mentzer:1991:LFI

[Men91] Stuart G. Mentzer. LP-form inclusion functions for global optimization. Computers and Mathematics with Applications, 21(6–7):51–65, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901606.

Meszaros:1996:FCF

[Més96] C. Mészáros. Fast Cholesky factorization for interior point methods of linear programming. Computers and Mathematics with Applications, 31 (4–5):49–54, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002154.

Meywerk:1999:NMC

[Mey99] M. Meywerk. A numerical method to classify zeros of implicit eigenvalue problems. Computers and Mathematics with Applications, 38(9– 10):39–47, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002606.

Mezey:1997:CMG

[Mez97] P. G. Mezey. Chirality measures and graph representations. Computers and Mathematics with Applications, 34 (11):105–112, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002241.

Murio:1990:DSA

[MG90a] D. A. Murio and L. Guo. Discrete stability analysis of the mollification method for numerical differentiation. Computers and Mathematics with Applications, 19(6): 15–26, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090044K.

See errata [MG90b].

Murio:1990:EDS

[MG90b] D. A. Murio and L. Guo. Errata to “Discrete stability analysis of the modification method for numerical diffentiation”: Computers Math. Applic. 19 (6), 15– 26 (1990). Computers and Mathematics with Applications, 20(12):75, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090167I.

See [MG90a].

Murio:1990:SSM

[MG90c] D. A. Murio and L. Guo. A stable space marching finite differences algorithm for the inverse heat conduction problem with no initial filtering procedure. Computers and Mathematics with Applications, 19(10):35–50, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090356O.

Marciniak:1991:AOH

[MG91] Andrzej Marciniak and Donald Greenspan. Arbitrary order, Hamiltonian conserving numerical solutions of Calogero and Toda systems. Computers and Mathematics with Applications, 22(7): 11–35, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901787.

Mendivil:1993:MDD

[MG93] Jose R. González De Mendívil and Jose R. Garitagoitia. A model for deadlock detection based on automata and languages theory. Computers and Mathematics with Applications, 25(6):47–55, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390298A.

Manoranjan:1995:ADI

[MG95] V. S. Manoranjan and M. O.

Gomez. Alternating direction implicit method with exponential upwinding. Computers and Mathematics with Applications, 30(11):47–58, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/089812219500163S.

Manoranjan:1997:TSJ

[MG97] V. S. Manoranjan and M. Olmos Gomez. A two-step Jacobi-type iterative method. Computers and Mathematics with Applications, 34 (1):1–9, July 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700093X.

McCalla:1992:GH

[MGBP92] Gordon McCalla, Jim Greer, Bryce Barrie, and Paul Pospisil. Granularity hierarchies. Computers and Mathematics with Applications, 23(2–5):363–375, January/March 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290148B.

Murio:1992:NSN

[MGM92] Diego A. Murio, Doris Hinestroza G., and Carlos E. Mejía. New stable numerical inversion of Abel’s integral equation. Computers and Mathematics with Applications, 23(11):3–11, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290064O.

Mo-Hong:1993:MCC

[MH93a] Chou Mo-Hong. A modified cell-to-cell mapping method for nonlinear systems. Computers and Mathematics with Applications, 25 (8):47–57, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390171Q.

Murio:1993:SMS

[MH93b] Diego A. Murio and Doris Hinestroza. The space marching solution of the inverse heat conduction problem and the identification of the initial temperature distribution. Computers and Mathematics with Applications, 25(4):55–63, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390248T.

Maleknejad:1999:NCM

[MH99] K. Maleknejad and M. Hadizadeh.

A new computational method for Volterra–Fredholm integral equations. Computers and Mathematics with Applications, 37(9):1–8, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001078.

Meyer:1996:MSF

[MHBD96] H. De Meyer, T. Van Hecke, G. Vanden Berghe, and M. Van Daele. Multistride L-stable fourth-order methods for the numerical solution of ODEs. Computers and Mathematics with Applications, 32(7):1–8, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001496.

Michelson:1990:SMC

[Mic90] S. Michelson. A system for Monte Carlo simulation of heterogeneous tumor cell populations. Computers and Mathematics with Applications, 20(4–6):139–148, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090323C.

Mickens:1994:CAS

[Mic94] R. E. Mickens. Construction of asymptotic solutions to discrete Bessel equations. Computers and Mathematics with Applications, 28 (1–3):219–226, August 1994.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001103.

Mickens:1998:APS

[Mic98]  R. E. Mickens.  Asymp-

totic properties of solutions to discrete Coulomb equations. Computers and Mathematics with Applications, 36(10–12):285–289, November/December 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800291.

Milner:1990:NMM

[Mil90] F. A. Milner. A numerical method for a model of population dynamics with spatial diffusion. Computers and Mathematics with Applications, 19(4):31–43, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901357.

Miller:1991:CPC

[Mil91] F. R. Miller. Characterizing principal curvature data as functions of the normal vectors and determining the surfaces. Computers and Mathematics with Applications, 22 (4–5):119–124, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190137S.

Mills:1992:MII

[Mil92] Jeffrey A. Mills. Making inflexible investment decisions with incomplete information. Computers and Mathematics with Applications, 24(8–9): 247–258, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290202S.

Milaszewicz:1995:MDN

[Mil95] J. P. Milaszewicz. Monotone discrete Newton iterations and elimination. Computers and Mathematics with Applications, 30(1):79–90, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500069B.

Milovanovic:1998:NCI

[Mil98] G. V. Milovanović. Numerical calculation of integrals involving oscillatory and singular kernels and some applications of quadratures. Computers and Mathematics with Applications, 36(8):19–39, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001801.

Mishkoy:1990:PQI

[Mis90] G. K. Mishkoy. Priority queueing involving orientation and the problems of their software implementation. Computers and Mathematics with Applications, 19(1):109–113, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900893.

Mitra:1990:MBS

[Mit90] A. K. Mitra. Minimal Boolean sum-of-product forms in prolog. Computers and Mathematics with Applications, 19(2):75–81, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090010H.

Mittnik:1992:FIG

[Mit92a] S. Mittnik. Forecasting international growth rates with leading indicators: a systemtheoretic approach. Computers and Mathematics with Applications, 24(8–9):31–41, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290185K.

Mittnik:1992:P

[Mit92b]  Stefan Mittnik.  Preface.

Computers and Mathematics with Applications, 24(8–9): ix, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290182H.

Michalewicz:1992:MGA

[MJK92] Zbigniew Michalewicz, Cezary Z. Janikow, and Jacek B. Krawczyk. A modified genetic algorithm for optimal control problems. Computers and Mathematics with Applications, 23(12):83–94, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290094X.

Mandal:1991:NTO

[MK91] B. N. Mandal and P. K. Kundu. A note on the transmission of oblique waves through small apertures. Computers and Mathematics with Applications, 21(4): 17–22, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190243W.

Mika:1993:FOA

[MK93] Janusz R. Mika and Jan M. Kozakiewicz. First-order asymptotic expansion method for singularly perturbed systems of second-order ordinary differential equations.

Computers and Mathematics with Applications, 25(3): 3–11, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390138L.

Morokoff:1998:CSA

[MK98] W. J. Morokoff and A. Kersch. A comparison of scattering angle models. Computers and Mathematics with Applications, 35(1–2):155– 164, January 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002654.

Moore:1991:LPR

[MKS^{+}91] James E. Moore II, Moon Kim, Jong-Gook Seo, Ying Wu, and Robert Kalaba. Linear programming, recurrent associative memories, and feed-forward neural networks. Computers and Mathematics with Applications, 22(11):71–90, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900364.

McCahon:1990:JSF

[ML90a] C. S. McCahon and E. S. Lee.

Job sequencing with fuzzy processing times. Computers and Mathematics with Applications, 19(7):31–41, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090191L.

Michelson:1990:EDC

[ML90b] S. Michelson and J. T. Leith. Effects of differential cell kill on the dynamic composition of heterogeneous tumors. Computers and Mathematics with Applications, 20 (4–6):149–159, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090324D.

Misra:1990:OES

[ML90c] P. K. Misra and E. S. Lee. Optimal estimation in stream quality modeling. Computers and Mathematics with Applications, 19(11):81–96, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190901519.

Marquina:1993:SMS

[MM93a] A. Marquina and V. Martínez. Shooting methods for 1D steady-state free boundary problems. Computers and Mathematics with Applications, 25(2):39–46, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390221G.

Mejia:1993:MHM

[MM93b] C. E. Mejía and D. A. Murio.

Mollified hyperbolic method for coefficient identification problems. Computers and Mathematics with Applications, 26(5):1–12, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900676.

Milaszewicz:1993:EFP

[MM93c] J. P. Milaszewicz and S. Abdel Masih. Elimination and fixed point iterations. Computers and Mathematics with Applications, 25(5): 43–53, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901974.

See [MM94].

Milaszewicz:1994:EEF

[MM94] J. P. Milaszewicz and S. Abdel Masih. Errata to “Elimination and fixed point iterations”: Computers and mathematics with applications, vol. 25, no. 5, pp. 43–54, 1993. Computers and Mathematics with Applications, 27 (4):113–115, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900604.

See [MM93c].

Mejia:1995:NID

[MM95a] C. E. Mejía and D. A. Murio. Numerical identification of diffusivity coefficient and initial condition by discrete mollification. Computers and Mathematics with Applications, 30 (12):35–50, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500173V.

Melville:1995:NAT

[MM95b] R. Melville and D. S. Mackey. A new algorithm for twodimensional numerical continuation. Computers and Mathematics with Applications, 30(1):31–46, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000657.

Miller:1995:RCF

[MM95c] A. R. Miller and I. S. Moskowitz. Reduction of a class of Fox–Wright psi functions for certain rational parameters. Computers and Mathematics with Applications, 30(11):73–82, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219500165U.

Murthy:1995:NGE

[MM95d] K. N. Balasubramanya Murthy and C. Siva Ram Murthy. A new Gaussian eliminationbased algorithm for parallel solution of linear equations. Computers and Mathematics with Applications, 29 (7):39–54, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500017S.

Manzini:1996:MRA

[MM96a] G. Manzini and L. Margara. Minimal residual algorithm and matrix–vector information. Computers and Mathematics with Applications, 32(4):57–63, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001253.

Marek:1996:IAD

[MM96b] I. Marek and P. Mayer. Iterative aggregation/disaggregation method for computing stationary probability vectors of Markov type operators. Computers and Mathematics with Applications, 31(4–5): 27–40, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002138.

Mejia:1996:NSG

[MM96c] C. E. Mejía and D. A. Murio. Numerical solution of generalized IHCP by discrete mollification. Computers and Mathematics with Applications, 32(2):33–50, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001010.

Miley:1996:SID

[MM96d] T. Miley and T. McFadden. A sonar, interactive dance and music system. Computers and Mathematics with Applications, 32(1):97–107, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000934.

Miller:1996:GCK

[MM96e] A. R. Miller and I. S. Moskowitz. Generalizations of the Carlton–Kimball distribution for a target’s future location. Computers and Mathematics with Applications, 31(8):61–68, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000314.

Milovanovic:1996:DPT

[MMMS96] I. S. Milovanović, E. I. Milovanović, I. Z. Milentijević, and M. K. Stojcev. Designing of processor-time optimal systolic arrays for band matrix–vector multiplication. Computers and Mathematics with Applications, 32(2): 21–31, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001009.

Milentijevic:1997:DOP

[MMMS97] I. Z. Milentijević, I. Ž. Milovanović, E. I. Milovanović, and M. K. Stojčev. The design of optimal planar systolic arrays for matrix multiplication. Computers and Mathematics with Applications, 33(6):17–35, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219700028X.

Manickam:1998:SOS

[MMP98] A. V. Manickam, K. M.

Moudgalya, and A. K. Pani. Second-order splitting combined with orthogonal cubic spline collocation method for the Kuramoto–Sivashinsky equation. Computers and Mathematics with Applications, 35(6):5–25, March 1998. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000133.

Maponi:1991:IPT

[MMZ91] Pierluigi Maponi, Luciano Misici, and Francesco Zirilli. An inverse problem for the three dimensional vector Helmholtz equation for a perfectly conducting obstacle. Computers and Mathematics with Applications, 22 (4–5):137–146, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190139U.

Murio:1998:DMA

[MMZ98] D. A. Murio, C. E. Mejía, and S. Zhan. Discrete mollification and automatic numerical differentiation. Computers and Mathematics with Applications, 35(5):1–16, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000017.

Manohar:1991:AAM

[MN91] Ram P. Manohar and Tat Y. Ngai. Automated allocation of mesh points for stiff twopoint boundary value problems. Computers and Mathematics with Applications, 22(3):75–79, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190071B.

Marsaglia:1993:SIS

[MN93] G. Marsaglia and B. Narasimhan.

Simulating interpolation search. Computers and Mathematics with Applications, 26(8): 31–42, October 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390329T.

Ma:1997:DSN

[MN97] Shilong Ma and Shucheng Ning. Deriving some new conditions on the existence of eight limit cycles for a cubic system. Computers and Mathematics with Applications, 33(7):59–84, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000436.

Meyer:1993:IPG

[MNB93] H. De Meyer, C. Niyokindi, and G. Vanden Berghe. The implementation of parallel Gram–Schmidt orthogonalisation algorithms on a ring of transputers. Computers and Mathematics with Applications, 25(1):65–72, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390212E.

Markowitz:1992:BPW

[MNE92] Judith A. Markowitz, J. Terry Nutter, and Martha W. Evens. Beyond is-a and part-whole: More semantic network links. Computers and Mathematics with Applications, 23(6–9):377– 390, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290113V.

Mastroianni:1995:IPZ

[MO95] G. Mastroianni and D. Occorsio. Interlacing properties of the zeros of the orthogonal polynomials and approximation of the Hilbert transform. Computers and Mathematics with Applications, 30(3– 6):155–168, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000933.

Meehan:1998:EPNb

[MO98a] M. Meehan and D. O’Regan. Existence principles for nonlinear resonant operator and integral equations. Computers and Mathematics with Applications, 36(8):41–52, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001813.

Meehan:1998:EPNa

[MO98b] M. Meehan and D. O’Regan.

Existence principles for nonresonant operator and integral equations. Computers and Mathematics with Applications, 35(9):79–87, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000595.

Moiseev:1997:SCS

[Moi97] E. Moiseev. Spectral characteristics of some nonlocal boundary-value problems. Computers and Mathematics with Applications, 34(5– 6):649–655, September 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001612.

Moll:1993:PAF

[Mol93] V. Moll. Polygonal approximation to the flow on the critical surface for the bistable equation. Computers and Mathematics with Applications, 25(1):45–51, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390210M.

Molenaar:1996:SCC

[Mol96] J. Molenaar. A simple cellcentered multigrid method for 3D interface problems. Computers and Mathematics with Applications, 31(9): 25–33, May 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000399.

Molodtsov:1999:SST

[Mol99] D. Molodtsov. Soft set theory — first results. Computers and Mathematics with Applications, 37(4–5):19–31, February/March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000565.

Moore:1991:GOP

[Moo91] Ramon E. Moore. Global optimization to prescribed accuracy. Computers and Mathematics with Applications, 21(6–7):25–39, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190158Z.

See erratum [Moo93b].

Moore:1993:RCM

[Moo93a] R. E. Moore. The resolution of close minima. Computers and Mathematics with Applications, 25(10–11):57– 58, May/June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390281Y.

Moore:1993:EGO

[Moo93b]  Ramon E. Moore.  Er-

ratum to: “Global optimization to prescribed accuracy” [Comput. Math. Appl. 21 (1991), no. 6–7, 25– 39; MR1096131 (91m:90159)]. Computers and Mathematics with Applications, 25 (10–11):187, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902924.

See [Moo91].

Moore:1994:NSD

[Moo94] R. E. Moore. Numerical solution of differential equations to prescribed accuracy. Computers and Mathematics with Applications, 28(10–12):253– 261, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001952.

Moree:1993:BPP

[Mor93] P. Moree. Bertrand’s Postulate for primes in arithmetical progressions. Computers and Mathematics with Applications, 26(5):35–43, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193900713.

Mika:1991:AAS

[MP91] Janus R. Mika and Andrzej Palczewski. Asymptotic analysis of singularly perturbed systems of ordinary differential equations. Computers and Mathematics with Applications, 21(10):13–32, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190058C.

Misra:1993:OFV

[MP93] J. C. Misra and M. K. Patra. Oscillatory flow of a viscous fluid through convergingdiverging orthotropic tethered tubes. Computers and Mathematics with Applications, 26(2):87–106, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390325P.

Misra:1994:PTP

[MP94] J. C. Misra and S. K. Pandey. Peristaltic transport of a particle-fluid suspension in a cylindrical tube. Computers and Mathematics with Applications, 28 (4):131–145, August 1994.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001340.

Maroulas:1996:GPN

[MP96] J. Maroulas and P. Psarrakos. Geometrical properties of numerical range of matrix polynomials. Computers and Mathematics with Applications, 31(4–5):41–47, February/March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002146.

Milner:1999:NAM

[MP99] F. A. Milner and C. A.

Patton. A new approach to mathematical modeling of host-parasite systems. Computers and Mathematics with Applications, 37 (2):93–110, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002557.

Misra:1993:BAA

[MPM93] J. C. Misra, M. K. Patra, and S. C. Misra. Biorheological aspects of arterial flow near bifurcations. Computers and Mathematics with Applications, 26(9):53–65, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390005G.

Mardanov:1997:GPR

[MPP97] K. Mardanov, Yu. Pis’mak, and A. Potyagailo. General properties of realistic neural network dynamics. Computers and Mathematics with Applications, 34(7– 8):657–666, October 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001661.

Mardanov:1997:MMP

[MPR97] K. A. Mardanov, Yu. M. Pismak, and V. L. Rapoport. Mathematical modelling of pirimidine photodimerization in polynucleotides. Computers and Mathematics with Applications, 34(7–8):911– 922, October 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700179X.

Moore:1998:RID

[MPR98] P. K. Moore, L. R. Petzold, and Y. Ren. Regularization of index-1 differential-algebraic equations with rank-deficient constraints. Computers and Mathematics with Applications, 35(5):43–61, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000042.

Misra:1992:UFB

[MPS92] J. C. Misra, M. K. Patra, and B. K. Sahu. Unsteady flow of blood through narrow blood vessels- a mathematical analysis. Computers and Mathematics with Applications, 24 (10):19–31, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290017C.

Miloh:1993:EFR

[MPS93] T. Miloh, M. Pachter, and A. Segal. The effect of a finite roll rate on the missdistance of a bank-to-turn missile. Computers and Mathematics with Applications, 26 (6):43–54, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390116D.

McCammon:1994:ODE

[MPS94] J. A. McCammon, B. M. Pettitt, and L. R. Scott. Ordinary differential equations of molecular dynamics. Computers and Mathematics with Applications, 28(10–12):319– 326, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001995.

Misra:1991:MAS

[MR91] J. C. Misra and S. Roy. Mathematical analysis of a specimen of tabular bone subjected to torsional loading. Computers and Mathematics with Applications, 21 (2–3):141–147, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190092I.

Mili:1992:MHB

[MR92] Hafedh Mili and Roy Rada. A model of hierarchies based on graph homomorphisms. Computers and Mathematics with Applications, 23(2– 5):343–361, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290147A.

Mickens:1994:FDS

[MR94] R. E. Mickens and I. Ramadhani. Finite-difference schemes having the correct linear stability properties for all finite step-sizes III. Computers and Mathematics with Applications, 27(4):77–84, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900566.

Mikolajczak:1999:DDP

[MR99] B. Mikolajczak and J. T.

Rumbut, Jr. Distributed dynamic programming using concurrent object-orientedness with actors visualized by high-level Petri nets. Computers and Mathematics with Applications, 37(11–12):23– 34, June 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900139X.

Morera:1997:HOI

[MRJ97] J. L. Morera, G. Rubio, and L. Jódar. Higher order implicit multistep methods for matrix differential equations. Computers and Mathematics with Applications, 33(4):39–48, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000059.

Mrozek:1996:TIM

[Mro96] M. Mrozek. Topological invariants, multivalued maps and computer assisted proofs in dynamics. Computers and Mathematics with Applications, 32(4):83–104, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001277.

McQuain:1994:PIM

[MRWM94] W. D. McQuain, C. J. Ribbens, L. T. Watson, and R. C. Melville. Preconditioned iterative methods for sparse linear algebra problems arising in circuit simulation. Computers and Mathematics with Applications, 27 (8):25–45, April 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900647.

Maponi:1996:TDT

[MRZ96] P. Maponi, M. C. Recchioni, and F. Zirilli. Threedimensional time harmonic electromagnetic inverse scattering: the reconstruction of the shape and the impedance of an obstacle. Computers and Mathematics with Applications, 31(3):1–7, February 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002014.

Melton:1991:FRF

[MS91] Austin Melton and Sujeet Shenoi. Fuzzy relations and fuzzy relational databases. Computers and Mathematics with Applications, 21(11– 12):129–138, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190114J.

Mahmoud:1993:RMS

[MS93] F. A. Mahmoud and M. A.

Saleem. Regularization method for solving radon integral equations. Computers and Mathematics with Applications, 25(4):3–13, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390243O.

Madden:1996:LES

[MS96] N. Madden and M. Stynes.

Linear enhancements of the streamline diffusion method for convection–diffusion problems. Computers and Mathematics with Applications, 32 (10):29–42, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001848.

Mukai:1999:AAR

[MS99] S. Mukai and I. Sano. Algorithms for aerosol retrieval based on scattering calculations. Computers and Mathematics with Applications, 37 (11–12):133–139, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001522.

Meaux:1990:CMC

[MSB90] L. M. Meaux, J. W. Seaman, Jr., and T. L. Boullion. Calculation of multivariate Chebyshev-type inequalities. Computers and Mathematics with Applications, 20(12):55–60, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090164F.

Mohapatra:1995:NIW

[MSR95] R. N. Mohapatra, F. Salzmann, and D. Ross. Norm inequalities which yield inclusion for Euler sequence spaces. Computers and Mathematics with Applications, 30(3–6):383–387, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500114X.

See remarks [Lei96].

Mazzia:1992:NMS

[MT92] F. Mazzia and D. Trigiante. Numerical methods for second order singular perturbation problems. Computers and Mathematics with Applications, 23(11):81–89, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290071O.

Marz:1994:SMG

[MT94a] R. März and C. Tischendorf. Solving more general index-2 differential algebraic equations. Computers and Mathematics with Applications, 28(10–12):77–105, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001871.

Mazzia:1994:RDE

[MT94b] F. Mazzia and D. Trigiante.

The role of difference equations in numerical analysis. Computers and Mathematics with Applications, 28 (1–3):209–217, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400109X.

Moon:1998:NSD

[MT98] I. H. Moon and R. P. Tewarson. Numerical solutions of differential equations for renal concentrating mechanism in inner medullary vasa recta models. Computers and Mathematics with Applications, 36(7):69–78, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001734.

Martinez-Torres:1993:TGMa

[MTGFG93] E. Martínez-Torres, J. J. López González, and M. Fernández-Gómez. A topological geometric method for the obtention of symmetry-adapted functions for point groups i. general theory. Computers and Mathematics with Applications, 26(2):79–85, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390324O.

Martinez-Torres:1993:TGMb

[MTLGFG93] E. Martínez-Torres, J. J. López-González, and M. Fernández-Gómez. A topological geometric method for the obtention of symmetry-adapted functions for point groups II. The icosahedral group. Computers and Mathematics with Applications, 26(7): 67–77, October 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390051V.

Martinez-Torres:1995:TGMa

[MTLGFG95a] E. Martinez-Torres, J. J. Lopez-Gonzalez, and M. Fernandez-Gomez. A topological geometric method for the obtention of symmetry-adapted functions for point groups III.

The cubic group. Computers and Mathematics with Applications, 29(10):35–39, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500042W.

Martinez-Torres:1995:TGMb

[MTLGFG95b] E. Martinez-Torres, J. J.

Lopez-Gonzalez, and M. Fernandez-Gomez. A topological geometric method for the obtention of symmetry-adapted functions for point groups IV. The dihedral groups. Computers and Mathematics with Applications, 29(10): 41–44, May 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500043X.

Markov:1996:P

[MU96] Svetoslav Markov and Christian Ullrich. Preface. Computers and Mathematics with Applications, 32(11): xv, December 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197866709.

Muraskin:1991:DAF

[Mur91a] M. Muraskin. The determinant of g_{ij} in Aesthetic Field Theory. Computers and Mathematics with Applications, 22(7):43–53, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190180C.

Murthy:1991:SHP

[Mur91b] C. Siva Ram Murthy. Solving hyperbolic PDE’s on hypercubes. Computers and Mathematics with Applications, 21(5):79–82, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190217R.

Murio:1992:NII

[Mur92a] Diego A. Murio. Numerical identification of interface source functions in heat transfer problems under nonlinear boundary conditions. Computers and Mathematics with Applications, 24(4): 65–76, August 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192900086.

Murio:1992:SIH

[Mur92b]  Diego A. Murio.  Solu-

tion of inverse heat conduction problems with phase changes by the mollification method. Computers and Mathematics with Applications, 24(7):45–57, October 1992. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901539.

Muraskin:1993:CA

[Mur93a] M. Muraskin. Chaos and aesthetics. Computers and Mathematics with Applications, 26 (5):93–103, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900779.

Muraskin:1993:INA

[Mur93b] M. Muraskin. Instantons in nonintegrable aesthetic field theory. Computers and Mathematics with Applications, 26(8):81–96, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390332P.

Muraskin:1994:DLS

[Mur94] M. Muraskin. Dynamical lattice systems. Computers and Mathematics with Applications, 28(7):77–95, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001626.

Malek:1995:CPZ

[MV95a] F. Malek and R. Vaillancourt. A composite polynomial zerofinding matrix algorithm. Computers and Mathematics with Applications, 30(2):37–47, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500076B.

Malek:1995:PZI

[MV95b] F. Malek and R. Vaillancourt. Polynomial zerofinding iterative matrix algorithms. Computers and Mathematics with Applications, 29(1):1–13, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400202V.

Mistiri:1999:DFD

[MW99] F. S. Mistiri and A. P. Wang. Disconjugacy of functional and differential systems (associated with input-output scattering model). Computers and Mathematics with Applications, 38(7–8):43–55, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002370.

Myers:1992:KCR

[Mye92] Donald E. Myers. Kriging, cokriging, radial basis functions and the role of positive definiteness. Computers and Mathematics with Applications, 24(12):139– 148, December 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290176I.

Mylopoulos:1992:PT

[Myl92]  John Mylopoulos.  The

PSN tribe. Computers and Mathematics with Applications, 23(2–5):223–241, January/March 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901425.

Marsaglia:1993:MTR

[MZ93] George Marsaglia and Arif Zaman. Monkey tests for random number generators. Computers and Mathematics with Applications, 26(9): 1–10, November 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390001C.

See also [PW95].

Murio:1995:SAI

[MZ95] D. A. Murio and H. C.

Zheng. A stable algorithm for 3D-IHCP. Computers and Mathematics with Applications, 29(5):97–110, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400251F.

Malajovich:1997:FSA

[MZ97] G. Malajovich and J. P. Zubelli. A fast and stable algorithm for splitting polynomials. Computers and Mathematics with Applications, 33(3):1–23, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002337.

Meng-Zhao:1990:MSS

[MZMQ90] Qin Meng-Zhao and Zhang Mei-Qing. Multi-stage symplectic schemes of two kinds of Hamiltonian systems for wave equations. Computers and Mathematics with Applications, 19(10):51–62, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090357P.

Nabli:1998:PMA

[Nab98]  H. Nabli.  Performability

measure for acyclic Markovian models. Computers and Mathematics with Applications, 35(8):41–51, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000431.

Nairay:1992:DPC

[Nai92] Alain Nairay. Dynamic portfolio choice under asset price lognormality. Computers and Mathematics with Applications, 24(8–9):157–166, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290195N.

Nakagami:1999:TPN

[Nak99] J.-I. Nakagami. A two-person noncooperative game for assets selling problem: Independent case. Computers and Mathematics with Applications, 37(11–12):207– 212, June 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001583.

Nazeer:1998:MMC

[NAKA98] M. M. Nazeer, M. Afzal, A. F. Khan, and N. Ahmed. Mathematical modeling and computer simulation of two orthogonally forced transverse vibrations in multisegment hanging cantilever. Computers and Mathematics with Applications, 36(6): 115–127, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001655.

Nambiar:1996:NN

[Nam96a]  K. K. Nambiar.  Nu-

Machine and NuAlgebra. Computers and Mathematics with Applications, 32 (4):105–107, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001289.

Nambiar:1996:SAG

[Nam96b] K. K. Nambiar. Sentient arithmetic and Gödel’s theorems. Computers and Mathematics with Applications, 32 (2):69–73, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001046.

Nambiar:1997:JCH

[Nam97] K. K. Nambiar. Justification of the continuum hypothesis. Computers and Mathematics with Applications, 33(5):1–3, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700014X.

Nambiar:1999:ADC

[Nam99a] K. K. Nambiar. Axiomatic derivation of the continuum hypothesis. Computers and Mathematics with Applications, 38(3–4):225– 227, August 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001959.

Nambiar:1999:RST

[Nam99b] K. K. Nambiar. Real set theory. Computers and Mathematics with Applications, 38 (7–8):167–171, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002473.

Nanda:1990:IME

[Nan90]  T. Nanda.  An interac-

tive method for the eigenvalue problem for matrices. Computers and Mathematics with Applications, 19(7): 43–51, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090192M.

Narayanan:1992:NPE

[Nar92] A. Narayanan. A note on parameter estimation in the multivariate beta distribution. Computers and Mathematics with Applications, 24 (10):11–17, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290016B.

Naschie:1992:PDM

[Nas92] M. S. El Naschie. Peano dynamics as a model for ergodic and strange nonchaotic behaviour. Computers and Mathematics with Applications, 23(1):25–33, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290077U.

Naschie:1995:SGC

[Nas95] M. S. El Naschie. Statistical geometry of a Cantor discretum and semiconductors. Computers and Mathematics with Applications, 29(12):103–110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000624.

Noor:1997:IMG

[NAS97] M. Aslam Noor and E. A. Al-Said. Iterative methods for generalized nonlinear variational inequalities. Computers and Mathematics with Applications, 33(8):1–11, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000497.

Nandy:1995:UAF

[NB95] S. C. Nandy and B. B. Bhattacharya. A unified algorithm for finding maximum and minimum object enclosing rectangles and cuboids. Computers and Mathematics with Applications, 29(8): 45–61, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500029X.

Nandy:1998:MEC

[NB98] S. C. Nandy and B. B. Bhattacharya. Maximal empty cuboids among points and blocks. Computers and Mathematics with Applications, 36(3):11–20, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001254.

Nikolopoulos:1995:PCP

[ND95] S. D. Nikolopoulos and S. D.

Danielopoulos. Parallel computation of perfect elimination schemes using partition techniques on triangulated graphs. Computers and Mathematics with Applications, 29(6):47–57, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500006K.

Nelson:1996:RTT

[Nel96] G. L. Nelson. Real time transformation of musical material with fractal algorithms.

Computers and Mathematics with Applications, 32(1): 109–116, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000946.

Neumann:1991:IA

[Neu91] Gustaf Neumann. Interpretational abstraction. Computers and Mathematics with Applications, 21(8):65–78, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900526.

Nado:1992:SMF

[NF92] Robert Nado and Richard Fikes. Saying more with frames: Slots as classes. Computers and Mathematics with Applications, 23(6–9): 719–731, March/May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290131Z.

Navarro:1996:ASP

[NFJ96] E. Navarro, M. V. Ferrer, and L. Jódar. Analytic solution with a priori error bounds for a class of mixed coupled partial differential equations. Computers and Mathematics with Applications, 31(3):19–31, February 1996. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002030.

Newbery:1991:IMC

[NG91] A. C. R. Newbery and Tylene S. Garrett. Interpolation with minimized curvature. Computers and Mathematics with Applications, 22(1):37–43, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190023W.

Nambiar:1997:BCN

[NGNM97] K. K. Nambiar, B. Gopinath, T. Nagaraj, and S. Manjunath. Boyce–Codd normal form decomposition. Computers and Mathematics with Applications, 33(4): 1–3, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000011.

Nguyen:1997:PCP

[Ngu97] G. D. Nguyen. A polynomial construction of perfect codes. Computers and Mathematics with Applications, 33(8):127–131, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000606.

Nishishiraho:1995:APR

[Nis95] T. Nishishiraho. Approximation processes with respect to positive multiplication operators. Computers and Mathematics with Applications, 30(3–6):389–408, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001158.

Nishikawa:1997:NPS

[NIS97a] T. Nishikawa, M. Imai, and S. Shimizu. Nonlinear phenomena in a self-organizing model. Computers and Mathematics with Applications, 33(3):73–79, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002386.

Nishizawa:1997:MFE

[Nis97b] K. Nishizawa. A method to find elements of cycles in an incomplete directed graph and its applications-binary AHP and Petri nets. Computers and Mathematics with Applications, 33(9):33–46, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000643.

Niwa:1996:MMS

[Niw96]  H.-S. Niwa.  Mathemati-

cal model for the size distribution of fish schools. Computers and Mathematics with Applications, 32 (11):79–88, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219600199X.

Nishimura:1992:QMC

[NJ92] Shoichi Nishimura and Yong Jiang. Queueing models of concurrency control in database with Poisson arrivals. Computers and Mathematics with Applications, 24(1–2):41–50, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902268.

Nketsa:1994:CDS

[Nke94] A. Nketsa. Conservative distributed simulation of Petri nets. Computers and Mathematics with Applications, 27(9–10):45–51, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901228.

Naadimuthu:1991:IIF

[NL91] G. Naadimuthu and E. S.

Lee. Invariant imbedding filter in the modeling of water resources. Computers and Mathematics with Applications, 21(11–12):9–22, ????

1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190103B.

Negi:1993:PPC

[NL93] Devendra S. Negi and E. Stanley Lee. Possibility programming by the comparison of fuzzy numbers. Computers and Mathematics with Applications, 25(9):43–50, May 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390131E.

Nakashima:1992:ODR

[NMUY92] K. Nakashima, H. Matznaga, Y. Ueda, and K. Yamato. Optimal design of redundant systems for improving fail-safe characteristics. Computers and Mathematics with Applications, 24(1–2):109–117, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290235A.

Nixon:1991:ETP

[NN91] J. B. Nixon and B. J. Noye. Explicit three-point three-level FDMS for the one-dimensional constantcoefficient advection-diffusion equation. Computers and Mathematics with Applications, 22(11):47–58, ????

1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190033Z.

Nandy:1997:EAS

[NNB97] S. C. Nandy, G. N. Nandakumar, and B. B. Bhattacharya. Efficient algorithms for single- and two-layer linear placement of parallel graphs. Computers and Mathematics with Applications, 34 (12):121–135, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002393.

Namasivayam:1993:EAT

[NO93] S. Namasivayam and E. L.

Ortiz. Error analysis of the Tau Method: dependence of the approximation error on the choice of perturbation term. Computers and Mathematics with Applications, 25(1):89–104, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390215H.

Noakes:1991:AFE

[Noa91] Lyle Noakes. Adapting functions to experimental measurements. Computers and Mathematics with Applications, 22(3):35–44, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190068F.

Noor:1995:GMV

[Noo95] M. A. Noor. Generalized multivalued variational inequalities II. Computers and Mathematics with Applications, 30 (11):83–89, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500166V.

Noor:1996:GMQ

[Noo96] M. A. Noor. Generalized multivalued quasi-variational inequalities. Computers and Mathematics with Applications, 31(12):1–13, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000715.

Noor:1998:GMQ

[Noo98] M. A. Noor. Generalized multivalued quasi-variational inequalities (II). Computers and Mathematics with Applications, 35(5):63–78, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000054.

Noor:1999:MEM

[Noo99] M. A. Noor. A modified extragradient method for general monotone variational inequalities. Computers and Mathematics with Applications, 38(1):19–24, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001649.

Novakovic:1990:SSN

[Nov90] Z. R. Novaković. Solving systems of non-linear equations using the Lyapunov direct method. Computers and Mathematics with Applications, 20(12):19–23, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090160L.

Newman:1996:FPB

[NP96] M. Newman and Y. C. Pao.

Factorization of polynomials based on the multiplicities of their roots and applications. Computers and Mathematics with Applications, 32(9):29–40, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001745.

Navarro:1995:MAA

[NPJ95] E. Navarro, E. Ponsoda, and L. Jódar. A matrix approach to the analytic-numerical solution of mixed partial differential systems. Computers and Mathematics with Applications, 30(1):99–109, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000716.

Natalini:1996:MDZ

[NR96a] P. Natalini and P. E. Ricci. The moments of the density of zeros for the relativistic Hermite and Laguerre polynomials. Computers and Mathematics with Applications, 31(6):87–96, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000089.

Nyland:1996:ATG

[NR96b] L. S. Nyland and J. H. Reif. An algebraic technique for generating optimal CMOS circuitry in linear time. Computers and Mathematics with Applications, 31(1):85–108, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500184Z.

Nath:1998:MTS

[NSD98] S. Nath, P. R. Sengupta, and L. Debnath. Magnetothermoelastic surface waves in micropolar elastic media. Computers and Mathematics with Applications, 35(3):47–55, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002782.

Nowicki:1992:RSA

[NSS92] Ryszard Nowicki, Roman Slowiński, and Jerzy Stefanowski. Rough sets analysis of diagnostic capacity of vibroacoustic symptoms. Computers and Mathematics with Applications, 24 (7):109–123, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290159F.

Noronha:1992:CSF

[NWB92] W. P. Noronha, E. L. Wachspress, and A. J. Baker. Consistent sparse factorization. Computers and Mathematics with Applications, 24(7):25–44, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901528.

Naadimuthu:1991:APM

[NWL91] G. Naadimuthu, K. W. Wang, and E. S. Lee. Air pollution modeling by quasilinearization. Computers and Mathematics with Applications, 21 (11–12):153–164, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190117M.

Nyiri:1999:SFN

[Nyí99] A. Nyíri. Surface fitting and a new direct method for solving block band linear system. Computers and Mathematics with Applications, 38(9– 10):161–173, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002710.

Odell:1997:SDR

[OB97] P. L. Odell and T. L. Boullion. Simultaneous diagonalization of rectangular matrices. Computers and Mathematics with Applications, 33 (9):93–96, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000680.

Oh:1996:LOS

[OCS96] M. Oh, U. J. Choi, and V. I. Shin. The lower-order suboptimal filter for discrete nonlinear systems. Computers and Mathematics with Applications, 31(7):9–16, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000144.

Odanaka:1991:CCS

[Oda91] Toshio Odanaka. Control chart and stochastic control processes. Computers and Mathematics with Applications, 21(11–12):65–78, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190108G.

Odanaka:1994:DPO

[Oda94] T. Odanaka. Dynamic programming and optimal inventory processes. Computers and Mathematics with Applications, 27(9–10):213– 217, May 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901392.

Onana:1993:DSC

[OH93] A. Onana and A. Mbala Hikolo. A direct solver for a class of symmetric linear systems. Computers and Mathematics with Applications, 26(2):45–50, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390320U.

Ojika:1991:NTO

[OHI91] Takeo Ojika, Takayuki Hashimoto, and Yoshiro Izumoto. Nonlinear time optimal control of a rotating rigid body. Computers and Mathematics with Applications, 21 (11–12):79–84, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190109H.

Ohnishi:1992:PIN

[Ohn92] Masamitsu Ohnishi. Policy iteration and Newton–Raphson methods for Markov decision processes under average cost criterion. Computers and Mathematics with Applications, 24(1–2):147–155, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290240I.

Olaf:1990:QSS

[Ola90] G. Olaf. Quadrature solution of singular integral equations

— I: a uniform treatment of Fredholm and Volterra equations. Computers and Mathematics with Applications, 19 (11):141–146, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090157F.

Olariu:1992:MCP

[Ola92] Stephan Olariu. The morphology of convex polygons. Computers and Mathematics with Applications, 24(7): 59–68, October 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290154A.

Oliker:1993:RTE

[Oli93] V. Oliker. The ray tracing and energy conservation equations for mirror systems with two reflecting surfaces. Computers and Mathematics with Applications, 26(7):9–18, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390047Y.

Oliveira:1998:AMM

[Oli98] S. Oliveira. Analysis of a multigrid method for a transport equation by numerical Fourier analysis. Computers and Mathematics with Applications, 35(12):7–12, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000923.

Ollerton:1990:DSA

[Oll90] R. L. Ollerton. A discrete segments approach to the optimization of insulin infusion algorithms. Computers and Mathematics with Applications, 20(4–6):207–215, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090328H.

Ober:1998:SII

[OOWD98] C. C. Ober, R. A. Oldfield, D. E. Womble, and J. Van Dyke. Seismic imaging on the intel paragon. Computers and Mathematics with Applications, 35(7):65–72, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000339.

Ortiz:1997:SRS

[OPD97] E. L. Ortiz, A. Pham, and Ngoc Dinh. Some remarks on structural relations between the tau method and the finite element method. Computers and Mathematics with Applications, 33 (4):105–113, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000102.

Oppenheim:1996:DME

[Opp96] D. V. Oppenheim. DMIX: a multifaceted environment for composing and performing computer music. Computers and Mathematics with Applications, 32(1):117–135, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000958.

ORegan:1995:FPT

[O’R95]  D. O’Regan.  A fixed

point theorem for condensing operators and applications to Hammerstein integral equations in Banach spaces. Computers and Mathematics with Applications, 30 (9):39–49, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500145O.

Ortiz:1997:PCP

[OR97a] E. L. Ortiz and T. J. Rivlin.

Preface: a collection of papers to commemorate the Cornelius Lanczos centennial. Computers and Mathematics with Applications, 33 (1–2):xiii–xiv, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002143.

Ortiz:1997:ATA

[OR97b] Eduardo L. Ortiz and T. J.

Rivlin. Approximation theory and applications. Preface: a collection of papers to commemorate the Cornelius Lanczos centennial.

Computers and Mathematics with Applications, 33(1–2): xiii–xiv, January 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). Dedicated to Cornelius Lanczos (1893–1974).

ORegan:1998:ASN

[O’R98a] D. O’Regan. Approximation of solutions of nonlinear operator equation on the half line. Computers and Mathematics with Applications, 35(9):65–77, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000583.

ORegan:1998:NFP

[O’R98b] D. O’Regan. New fixedpoints results for 1-set contractive set valued maps. Computers and Mathematics with Applications, 35(4): 27–34, February 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002873.

ORegan:1998:SDE

[O’R98c] D. O’Regan. Singular differential equations with linear and nonlinear boundary conditions. Computers and Mathematics with Applications, 35(3):81–97, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002812.

ORegan:1999:MDE

[O’R99] D. O’Regan. Multivalued differential equations in Banach spaces. Computers and Mathematics with Applications, 38(5–6):109–116, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002187.

Oh:1998:SDF

[OSLC98] M. Oh, V. I. Shin, Y. Lee, and U. J. Choi. Suboptimal discrete filters for stochastic systems with different types of observations. Computers and Mathematics with Applications, 35(3):17–27, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002757.

Ozawa:1994:SGC

[Oza94]  K. Ozawa.  Some glob-

ally convergent iterative method based on the bisection iteration for solving nonlinear scalar equations. Computers and Mathematics with Applications, 28 (6):83–91, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001545.

Ozyildirim:1996:TCT

[Özy96] S. Özyildirim. Three-country trade relations: a discrete dynamic game approach. Computers and Mathematics with Applications, 32 (5):43–56, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001344.

Pang:1994:MIM

[PA94] P. Y. H. Pang and R. P. Agarwal. Monotone iterative methods for a general class of discrete boundary value problems. Computers and Mathematics with Applications, 28 (1–3):243–254, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400112X.

Pang:1998:SGC

[PA98] P. Y. H. Pang and R. P. Agarwal. Stability of a general class of difference systems. Computers and Mathematics with Applications, 36(10–12):423–429, November/December 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800436.

Popenda:1999:DGI

[PA99] J. Popenda and R. P. Agarwal. Discrete Gronwall inequalities in many variables. Computers and Mathematics with Applications, 38(1): 63–70, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001698.

Pachpatte:1994:SNF

[Pac94]  B. G. Pachpatte.  Some

new finite difference inequalities. Computers and Mathematics with Applications, 28 (1–3):227–241, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001111.

Padgham:1992:DIL

[Pad92a] Lin Padgham. Defeasible inheritance: a lattice based approach. Computers and Mathematics with Applications, 23 (6–9):527–541, March/May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290122X.

Padoan:1992:NSA

[Pad92b] Pier Carlo Padoan. Nonlinear simulation analysis in continuous time econometric models. Computers and Mathematics with Applications, 24(8–9):57–65, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290187M.

Page:1990:CAS

[Pag90]  W. Page.  Computer al-

gebra systems: Issues and inquiries. Computers and Mathematics with Applications, 19(6):51–69, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090047N.

Palmer:1997:HAO

[Pal97] E. M. Palmer. The hidden algorithm of Ore’s theorem on Hamiltonian cycles. Computers and Mathematics with Applications, 34 (11):113–119, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002253.

Pan:1990:EEE

[Pan90a] V. Pan. Estimating the extremal eigenvalues of a symmetric matrix. Computers and Mathematics with Applications, 20(2):17–22, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090236D.

Pan:1990:MBF

[Pan90b] V. Pan. The modified barrier function method for linear programming and its extensions. Computers and Mathematics with Applications, 20(3):1–14, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090026G.

Pan:1992:PNI

[Pan92a] Victor Pan. Parametrization of Newton’s iteration for computations with structured matrices and applications. Computers and Mathematics with Applications, 24(3):61–75, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902154.

Panek:1992:OTM

[Pan92b] Emil Panek. Optimal trajectories in a multisectoral model of economic growth. Computers and Mathematics with Applications, 24(8–9): 99–109, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290190S.

Pan:1993:BSM

[Pan93a] V. Pan. Binary segmentation for matrix and vector operations. Computers and Mathematics with Applications, 25(3):69–71, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390144K.

Panton:1993:DFV

[Pan93b] Don B. Panton. Distribution function values for logstable distributions. Computers and Mathematics with Applications, 25(9):17–24, May 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390128I.

Pan:1994:IPS

[Pan94] V. Y. Pan. Improved parallel solution of a triangular linear system. Computers and Mathematics with Applications, 27(11):41–43, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900957.

Pan:1995:DIC

[Pan95] V. Y. Pan. Deterministic improvement of complex polynomial factorization based on the properties of the associated resultant. Computers and Mathematics with Applications, 30(2):71–94, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500078D.

Pan:1996:ONO

[Pan96] V. Y. Pan. Optimal and nearly optimal algorithms for approximating polynomial zeros. Computers and Mathematics with Applications, 31(12):97–138, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000806.

Pan:1998:BDA

[Pan98a] P.-Q. Pan. A basis-deficiency-allowing variation of the simplex method for linear programming. Computers and Mathematics with Applications, 36(3):33–53, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001278.

Pan:1998:DPS

[Pan98b] Ping-Qi Pan. A dual projective simplex method for linear programming. Computers and Mathematics with Applications, 35(6):119–135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000248.

Pan:1998:NFA

[Pan98c] V. Y. Pan. New fast algorithms for polynomial interpolation and evaluation on the Chebyshev node set. Computers and Mathematics with Applications, 35 (3):125–129, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002836.

Pao:1998:MMF

[Pao98] C. V. Pao. Monotone methods for a finite difference system of reaction diffusion equation with time delay. Computers and Mathematics with Applications, 36(10–12): 37–47, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800072.

Papakyriazis:1992:OEP

[Pap92] Panagiotis A. Papakyriazis. Optimal estimation policies of stochastic linear systems with time-varying parameters. Computers and Mathematics with Applications, 24(8–9):67–75, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290188N.

Pap:1994:DEP

[Pap94a] G. Pap. The distribution of estimates of parameters of multidimensional stationary AR processes. Computers and Mathematics with Applications, 27(2):1–8, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900337.

Papaschinopoulos:1994:CKS

[Pap94b] G. Papaschinopoulos. Conditions for kinematical similarity of a linear difference equation to a block upper triangular equation. Computers and Mathematics with Applications, 28 (1–3):255–257, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001138.

Pardalos:1991:GOA

[Par91a] Panos M. Pardalos. Global optimization algorithms for linearly constrained indefinite quadratic problems. Computers and Mathematics with Applications, 21(6– 7):87–97, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190163X.

Parisi:1991:GSA

[Par91b] Jürgen Parisi. Global symmetry aspects of a compartmentalized reaction-diffusion system. Computers and Mathematics with Applications, 22(12):23–31, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190144S.

Pascal:1992:SSA

[Pas92] J. P. Pascal. Similarity solutions for axisymmetric plane radial power law fluid flows through a porous medium. Computers and Mathematics with Applications, 23(12): 25–41, June 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900905.

Paszkowski:1997:FCQ

[Pas97] S. Paszkowski. Fast convergent quasipower series for some elementary and special functions. Computers and Mathematics with Applications, 33(1–2):181– 191, January 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002295.

Paternoster:1998:OBF

[Pat98]  B. Paternoster.  Order

bound for a family of parallel Runge–Kutta–Nyström methods through computer algebra. Computers and Mathematics with Applications, 35(9):107–119, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000625.

Pavani:1997:NAR

[Pav97]  R. Pavani.  The numer-

ical approximation of the rotation number of planar maps. Computers and Mathematics with Applications, 33(5):103–110, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000230.

Pazzani:1990:LOA

[Paz90] M. J. Pazzani. Learning in order to avoid search in logic programming. Computers and Mathematics with Applications, 20(9–10):101– 110, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090115Z.

Pozzi:1995:APA

[PB95] A. Pozzi and E. Bassano.

Application of Padé approximants to the study of plane jets. Computers and Mathematics with Applications, 30 (11):107–123, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500168X.

Pap:1990:AMM

[PBG90] Gy. Pap, J. Bokor, and P. Gáspár. Analysis of multiple model method for change detection of AR processes. Computers and Mathematics with Applications, 19(1):115–121, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900907.

Petkovic:1991:HMP

[PC91] M. S. Petković and Lj. Cvetković. A hybrid method for a polynomial complex zero. Computers and Mathematics with Applications, 21 (2–3):181–186, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190097N.

Pan:1992:SVD

[PC92] Yi Pan and Henry Y. H. Chuang. Singular value decomposition on SIMD hypercube and shuffle-exchange computers. Computers and Mathematics with Applications, 24(4):23–32, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900042.

Petkovic:1993:SII

[PC93] Miodrag S. Petković and Carsten Carstensen. Some improved inclusion methods for polynomial roots with Weierstrass’ corrections. Computers and Mathematics with Applications, 25(3): 59–67, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390143J.

Park:1994:EAD

[PC94a] Y. Park and U. J. Choi.

The error analysis for degree reduction of Bézier curves. Computers and Mathematics with Applications, 27 (12):1–6, June 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900809.

Peng:1994:ABN

[PC94b] Qiu-Liang Peng and Lan-Sun Chen. Asymptotic behavior of the nonautonomous twospecies Lotka–Volterra competition models. Computers and Mathematics with Applications, 27(12):53–60, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219490085X.

Pagaldipti:1996:DSP

[PC96a] N. Pagaldipti and A. Chattopadhyay. A discrete semianalytical procedure for aerodynamic sensitivity analysis including grid sensitivity. Computers and Mathematics with Applications, 32(3):61–71, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001137.

Paprzycki:1996:USM

[PC96b] M. Paprzycki and C. Cyphers. Using Strassen’s matrix multiplication in high performance solution of linear systems. Computers and Mathematics with Applications, 31 (4–5):55–61, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002162.

Papapanayotou:1990:MMC

[PCR90] C. J. Papapanayotou, Y. Cherruault, and B. De La Rochefoucauld. A mathematical model of the circle of Willis in the presence of an arteriovenous anomaly.

Computers and Mathematics with Applications, 20 (4–6):199–206, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090327G.

Panteliou:1996:DII

[PDK96] S. D. Panteliou, A. D. Dimarogonas, and I. N. Katz. Direct and inverse interpolation for Jacobian elliptic functions, zeta function of Jacobi and complete elliptic integrals of the second kind. Computers and Mathematics with Applications, 32(8):51–57, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001666.

Pengilly:1996:IPL

[Pen96] S. Pengilly. Integrating performance, live electronics, and interactive video. Computers and Mathematics with Applications, 32(1):75–77, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000880.

Petkovski:1992:KBS

[Pet92] Djordjija B. Petkovski. A knowledge-based system for robustness analysis of largescale economic systems. Computers and Mathematics with Applications, 24(8–9):137– 148, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290193L.

Pfiffer:1998:MPO

[Pfi98]  A. K. Pfiffer.  Moving

Paragon^{T M} operating system to a new hardware platform. Computers and Mathematics with Applications, 35 (7):33–43, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000303.

Piccioni:1991:CMA

[Pic91] Mauro Piccioni. A combined multistart-annealing algorithm for continuous global optimization. Computers and Mathematics with Applications, 21(6–7):173–179, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190171Y.

Pinto:1994:DD

[Pin94]  M. Pinto.  Discrete di-

chotomies. Computers and Mathematics with Applications, 28(1–3):259–270, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122194001146.

Pinto:1998:WCB

[Pin98] M. Pinto. Weighted convergent and bounded solutions of difference systems. Computers and Mathematics with Applications, 36(10–12):391– 400, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800394.

Piotrowski:1994:SDF

[Pio94] J. A. Piotrowski. Sequential and data flow models of processing. Computers and Mathematics with Applications, 27(1):81–96, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900078.

Petkovic:1997:FSZ

[PIT97] M. Petković, S. Ilić, and S. Tricković. A family of simultaneous zerofinding methods. Computers and Mathematics with Applications, 34(10):49–59, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219700206X.

Pardalos:1991:GST

[PJ91] Panos M. Pardalos and Somesh Jha. Graph separation techniques for quadratic zero-one programming. Computers and Mathematics with Applications, 21(6–7):107– 113, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190165Z.

Powell:1990:BDO

[PL90] P. Powell and H. C. Lau. Branch distance optimization of structured programs. Computers and Mathematics with Applications, 19(12): 53–64, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090251E.

Parthasarathy:1999:IPB

[PL99] P. R. Parthasarathy and R. B. Lenin. An inverse problem in birth and death processes. Computers and Mathematics with Applications, 38(1):33–40, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001662.

Pardalos:1995:LPA

[PLH95] P. M. Pardalos, Y. Li, and W. W. Hager. Linear programming approaches to the convex hull problem in R^{m}. Computers and Mathematics with Applications, 29(7): 23–29, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500015Q.

Petkovic:1991:CSM

[PM91] M. S. Petković and Z. M. Marjanović. A class of simultaneous methods for the zeros of analytic functions. Computers and Mathematics with Applications, 22(10):79–87, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901949.

Padmanabhan:1995:ARA

[PM95a] R. Padmanabhan and W. Mc-Cune. Automated reasoning about cubic curves. Computers and Mathematics with Applications, 29(2): 17–26, January 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194002146.

Padmanabhan:1995:SIT

[PM95b] R. Padmanabhan and W. Mc-Cune. Single identities for ternary Boolean algebras. Computers and Mathematics with Applications, 29(2):

13–16, January 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194002135.

Pao:1998:DAE

[PN98] Y. C. Pao and M. Newman. A direct algorithm for extracting a transfer function’s multiple zeros or poles. Computers and Mathematics with Applications, 35(9):33–44, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000558.

Punnen:1999:CBO

[PN99] A. P. Punnen and K. P. K. Nair. Constrained balanced optimization problems. Computers and Mathematics with Applications, 37(9): 157–163, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001194.

Piziak:1999:CPS

[POH99] R. Piziak, P. L. Odell, and R. Hahn. Constructing projections on sums and intersections. Computers and Mathematics with Applications, 37(1):67–74, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002429.

Pokrovski:1997:LPT

[Pok97] A. A. Pokrovski. Lax–Phillips theory for sound waves scattering by thin infinite elastic planes. Computers and Mathematics with Applications, 34 (2–4):143–171, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001223.

Pol:1990:LEI

[Pol90] A. T. Pol. Life: Energyinformation relationship within material systems — I. General outline of the concept. Computers and Mathematics with Applications, 20 (4–6):269–285, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090333F.

Pomeranz:1991:CMN

[Pom91] Shirley B. Pomeranz. Computational methods for nonlinear elliptic eigenvalue problems. Computers and Mathematics with Applications, 21 (8):9–20, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900478.

Pomeranz:1992:SWA

[Pom92] S. B. Pomeranz. A solitary wave application of an iterative method for nonlinear elliptic eigenvalue problems. Computers and Mathematics with Applications, 23(1):45–49, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290079W.

Popenda:1994:OSD

[Pop94] J. Popenda. The oscillation of solutions of difference equations. Computers and Mathematics with Applications, 28 (1–3):271–279, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001154.

Pourbabai:1990:AOP

[Pou90a] B. Pourbabai. Approximation of the overflow process of a D/G/K loss system. Computers and Mathematics with Applications, 19(7):53–58, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090193N.

Pourbabai:1990:ACA

[Pou90b] B. Pourbabai. An assembly cell with an automated quality control station: Exponentially distributed processing times. Computers and Mathematics with Applications, 19(4):17–30, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901346.

Petrosky:1997:ECD

[PP97] T. Petrosky and I. Prigogine. The extension of classical dynamics for unstable Hamiltonian systems. Computers and Mathematics with Applications, 34(2– 4):1–44, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001168.

Park:1999:AVL

[PPC99] H. Park, K. Park, and Y. Cho.

Analysis of the variable length nonzero window method for exponentiation. Computers and Mathematics with Applications, 37(7):21–29, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900084X.

Petkovic:1998:PEF

[PPH98] M. S. Petkovic, L. D.

Petković, and D. D. Herceg. Point estimation of a family of simultaneous zero-finding methods. Computers and Mathematics with Applications, 36(2):1–12, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001114.

Pan:1990:BCD

[PR90] V. Pan and J. Reif. The bit-complexity of discrete solutions of partial differential equations: Compact multigrid. Computers and Mathematics with Applications, 20 (2):9–16, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090235C.

Powar:1991:CCR

[PR91] P. L. Powar and R. Rao. A counterexample of the construction of C^{1}-rational finite element due to Wachspress. Computers and Mathematics with Applications, 22(3):17–22, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190065C.

Pan:1993:GCM

[PR93] Victor Pan and John Reif. Generalized compact multigrid. Computers and Mathematics with Applications, 25 (9):3–5, May 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390126G.

Pettigrew:1996:CMS

[PR96] M. F. Pettigrew and H. Rasmussen. A compact method for second-order boundary value problems on nonuniform grids. Computers and Mathematics with Applications, 31(9):1–16, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000375.

primak:1994:ASL

[pri94]  M. primak.  Algorithm

for solving a linear exchange model with rationing. Computers and Mathematics with Applications, 27(4): 85–90, February 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900574.

Prnic:1997:SOS

[Prn97]  Ž. Prnić.  Second-order

scheme for the scalar nonlinear conservation laws with flux depending on the space variable. Computers and Mathematics with Applications, 33(8):51–57, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000539.

Prolla:1995:SLB

[Pro95] J. B. Prolla. Semilattices and lattices of bounded continuous functions. Computers and Mathematics with Applications, 30(3–6):183– 190, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500095X.

Poguda:1990:DMR

[PS90] A. L. Poguda and E. K. Shpilewski. Dynamical methods for random processes recognition. Computers and Mathematics with Applications, 19(1):123–132, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090091W.

Pan:1992:FPC

[PS92] Victor Pan and Robert Schreiber. A fast, preconditioned conjugate gradient Toeplitz solver. Computers and Mathematics with Applications, 24(7):17–24, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901517.

Parisi:1993:PSC

[PS93a] J. Parisi and R. Stoop. A possible scheme for classifying phase transitions in the generalized entropy spectrum of a nonhyperbolic system. Computers and Mathematics with Applications, 25(5): 33–41, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901963.

Price:1993:OTA

[PS93b] C. C. Price and M. Salama.

Optimal task allocation in hypercube multiprocessor ensembles. Computers and Mathematics with Applications, 26(12):17–24, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219390055Z.

Prestin:1995:MTW

[PS95] J. Prestin and F. Sprengel. Multivariate trigonometric wavelet decompositions. Computers and Mathematics with Applications, 30(3– 6):169–182, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000941.

Papakaliatakis:1996:NMN

[PS96] G. Papakaliatakis and T. E. Simos. A new method for the numerical solution of fourthorder BVP’s with oscillating solutions. Computers and Mathematics with Applications, 32(10):1–6, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001812.

Panja:1996:HFR

[PSD96] S. Panja, P. R. Sengupta, and L. Debnath. Hydromagnetic flow of reiner-rivlin fluid between two coaxial circular cylinders with porous walls. Computers and Mathematics with Applications, 32 (2):1–4, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000983.

Pan:1993:NAF

[PSLT93] Victor Pan, Akimou Sadikou, Elliott Landowne, and Olen Tiga. A new approach to fast polynomial interpolation and multipoint evaluation. Computers and Mathematics with Applications, 25 (9):25–30, May 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390129J.

Paczelt:1999:SCP

[PSS99] I. Páczelt, B. A. Szabó, and T. Szabó. Solution of contact problem using the hp-version of the finite element method. Computers and Mathematics with Applications, 38(9– 10):49–69, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002618.

Piterbarg:1991:EAL

[PT91] V. Piterbarg and V. Taran.

Exact asymptotics for large deviation probabilities of normalized Brownian bridge with applications to empirical processes. Computers and Mathematics with Applications, 22(7):85–91, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901857.

Petkovic:1996:TLM

[PT96] M. S. Petković and S. B.

Tricković. Tchebychef-like method for the simultaneous finding zeros of analytic functions. Computers and Mathematics with Applications, 31(8):85–93, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000338.

Pan:1998:NTC

[PTC^{+}98] V. Y. Pan, M. A. Tabanjeh, Z. Q. Chen, E. I. Landowne, and A. Sadikou. New transformations of Cauchy matrices and Trummer’s problem. Computers and Mathematics with Applications, 35 (12):1–5, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000911.

Pap:1995:SBC

[PvZ95] G. Pap and M. C. A. van Zuijlen. The Stringer bound in case of uniform taintings. Computers and Mathematics with Applications, 29 (10):51–59, May 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500045Z.

Percus:1995:TAM

[PW95] Ora E. Percus and Paula A. Whitlock. Theory and application of Marsaglia’s monkey test for pseudorandom number generators. ACM Transactions on Modeling and Computer Simulation, 5(2): 87–100, April 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [MZ93].

Pavarino:1997:ISM

[PW97] L. F. Pavarino and O. B. Widlund. Iterative substructuring methods for spectral elements: Problems in three dimensions based on numerical quadrature. Computers and Mathematics with Applications, 33(1–2):193– 209, January 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002301.

Peng:1999:AOB

[PXHG99] Mingshu Peng, Qianli Xu, Lihong Huang, and Weigao Ge. Asymptotic and oscillatory behavior of solutions of certain second-order nonlinear difference equations. Computers and Mathematics with Applications, 37(3): 9–18, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000413.

Potts:1994:DEM

[PY94] R. B. Potts and X.-H. Yu.

Difference equation modelling of a variable structure system. Computers and Mathematics with Applications, 28 (1–3):281–289, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001162.

Pan:1997:ANT

[PYS97] V. Y. Pan, Y. Yu, and C. Stewart. Algebraic and numerical techniques for the computation of matrix determinants. Computers and Mathematics with Applications, 34(1):43–70, July 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000977.

Pan:1999:GOS

[PZ99a] Xing-Chen Pan and Quan Zheng. Global optimum shape design. Computers and Mathematics with Applications, 37(4–5):151–162, February/March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000668.

Pap:1999:APN

[PZ99b] G. Pap and M. C. A. Van Zuijlen. Asymptotic properties of nearly unstable multivariate AR processes. Computers and Mathematics with Applications, 37(2):11–19, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198002491.

Pan:1995:FPC

[PZDH95] V. Pan, A. L. Zheng, O. Dias, and X. H. Huang. A fast, preconditioned conjugate gradient Toeplitz and Toeplitzlike solvers. Computers and Mathematics with Applications, 30(8):57–63, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500137N.

Qu:1996:SSI

[QA96] R. Qu and R. P. Agarwal.

Smooth surface interpolation to scattered data using interpolatory subdivision algorithms. Computers and Mathematics with Applications, 32(3):93–110, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001150.

Qu:1998:SAC

[QA98] R. Qu and R. P. Agarwal.

A subdivision approach to the construction of approximate solutions of boundaryvalue problems with deviating arguments. Computers and Mathematics with Applications, 35(11):121–135, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000893.

Qu:1996:SCI

[Qu96] R. Qu. Smooth curve interpolation with generalized conics. Computers and Mathematics with Applications, 31 (7):37–64, April 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600017X.

Qin:1991:CRK

[QZ91] Meng-Zhao Qin and Wen-Jie Zhu. Canonical Runge– Kutta–Nyström (RKN) methods for second order ordinary differential equations. Computers and Mathematics with Applications, 22(9): 85–95, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190209M.

Qin:1993:VPS

[QZ93a] M.-Z. Qin and W.-J. Zhu. Volume-preserving schemes and numerical experiments. Computers and Mathematics with Applications, 26(4): 33–42, August 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390032Q.

Qin:1993:NST

[QZ93b] Meng-Zhao Qin and Wen-Jei Zhu. A note on stability of a three-stage difference scheme for ordinary differential equations. Computers and Mathematics with Applications, 25(1):35–44, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390209E.

Qin:1993:CSS

[QZ93c] Mengzhao Qin and Wenjie Zhu. Construction of symplectic schemes for wave equations via hyperbolic functions sinh(x), cosh(x) and tanh(x). Computers and Mathematics with Applications, 26(8):1–11, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390326Q.

Rodin:1992:MPA

[RA92a] Ervin Y. Rodin and S. Massoud Amin. Maneuver prediction in air combat via artificial neural networks. Computers and Mathematics with Applications, 24(3): 95–112, August 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902176.

Rucker:1992:DAC

[RA92b] Rob Rucker and Tario A.

Aldowaisan. A design approach for constructing engineering scenario maps. Computers and Mathematics with Applications, 23(6–9):419– 440, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290116Y.

Rabut:1992:ISA

[Rab92] Christophe Rabut. An introduction to Schoenberg’s approximation. Computers and Mathematics with Applications, 24(12):149– 175, December 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290177J.

Rachowicz:1999:APG

[Rac99]  W. Rachowicz.  Analysis

of a preconditioned GMRES solver for a nonsymmetric SUPG discretization of the compressible Navier–Stokes equations. Computers and Mathematics with Applications, 38(2):73–96, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001844.

Ramachandran:1990:BEM

[Ram90a] P. A. Ramachandran. A boundary element method based solution procedure for non-linear boundary value problems. Computers and Mathematics with Applications, 19(10):63–76, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090358Q.

Ramos:1990:KTF

[Ram90b] J. A. Ramos. A Kalmantracking filter approach to nonlinear programming. Computers and Mathematics with Applications, 19(11):63–74, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219090149E.

Ramachandran:1991:BIS

[Ram91a] P. A. Ramachandran. Boundary integral solution method for diffusion-reaction problems with both material and rate nonlinearities. Computers and Mathematics with Applications, 21(2–3):7–13, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190075F.

Ramm:1991:FCS

[Ram91b] A. G. Ramm. Finding conductivity from surface measurements. Computers and Mathematics with Applications, 21(8):85–91, ???? 1991.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900548.

Ramm:1991:ILA

[Ram91c] A. G. Ramm. Inversion of limited-angle tomographic data. Computers and Mathematics with Applications, 22 (4–5):101–111, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190135Q.

Ramm:1991:P

[Ram91d]  A. G. Ramm.  Preface.

Computers and Mathematics with Applications, 22(4– 5):vii, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190128Q.

Ramm:1991:PIG

[Ram91e]  A. G. Ramm.  A prob-

lem of integral geometry. Computers and Mathematics with Applications, 22 (4–5):113–117, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190136R.

Rach:1992:MD

[RAM92] R. Rach, G. Adomian, and R. E. Meyers. A modified decomposition. Computers and Mathematics with Applications, 23(1):17–23, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290076T.

Ramachandran:1993:NSI

[Ram93] M. P. Ramachandran. Numerical solution of an integral equation with logarithmic singularity. Computers and Mathematics with Applications, 26(10):51–57, November 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390025Q.

Rampone:1995:LCI

[Ram95] S. Rampone. Linear codes interpolation from noisy patterns by means of a vector quantization process. Computers and Mathematics with Applications, 30 (11):91–106, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500167W.

Randez:1992:IEM

[Rán92] L. Rández. Improving the efficiency of the multiple shooting technique. Computers and Mathematics with Applications, 24(7):127–132, October 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290161A.

Rand:1993:LCS

[Ran93] O. Rand. Limit cycles and stability of highly nonlinear periodic systems by harmonic variables. Computers and Mathematics with Applications, 25(4):15–25, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390244P.

Raqab:1998:OSB

[Raq98] M. Z. Raqab. Order statistics from the Burr type X model. Computers and Mathematics with Applications, 36 (4):111–120, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001436.

Raina:1997:DSG

[RB97] R. K. Raina and M. Bolia. The decompositional structure of a generalized hypergeometric transformation of convolution type. Computers and Mathematics with Applications, 34 (9):87–93, November 1997.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001910.

Randell:1992:ELT

[RC92] David A. Randell and Anthony G. Cohn. Exploiting lattices in a theory of space and time. Computers and Mathematics with Applications, 23(6–9):459– 476, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901182.

Radwan:1996:HIG

[RD96a] A. E. Radwan and L. Debnath. Hydromagnetic instability of a gravitating finite resistive fluid layer sandwiched in a different fluid. Computers and Mathematics with Applications, 31(11):61–67, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000624.

Richardson:1996:SCH

[RD96b] A. D. Richardson and J. R.

Dormand. The simulated cooling of the hot-rolled structural steel sections. Computers and Mathematics with Applications, 31(8):37–47, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000296.

Ruffino:1999:PAR

[RD99] E. Ruffino and P. P. Delsanto. Problems of accuracy and reliability in 2-d LISA simulations. Computers and Mathematics with Applications, 38 (5–6):89–97, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002163.

Rachev:1992:PAO

[RDK92] S. T. Rachev, B. Dimitrov, and Z. Khalil. A probabilistic approach to optimal quality usage. Computers and Mathematics with Applications, 24(8–9):219–227, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192902002.

Read:1997:GTA

[Rea97] R. C. Read. Graph theory and the amateur cryptographer. Computers and Mathematics with Applications, 34 (11):121–127, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002265.

Reggini:1994:TGR

[Reg94] H. C. Reggini. Tree graph representation of Hamiltonian paths. Computers and Mathematics with Applications, 27(5):101–114, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900795.

Reif:1993:PPP

[Rei93a] J. H. Reif. Probabilistic parallel prefix computation. Computers and Mathematics with Applications, 26(1): 101–110, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390089E.

Reif:1993:EVF

[Rei93b] John H. Reif. Efficient VLSI fault simulation. Computers and Mathematics with Applications, 25(2):15–32, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390219L.

Reid:1994:ODE

[Rei94] D. J. Reid. Optimal distributed execution of join queries. Computers and Mathematics with Applications, 27(11):27–40, July 1994. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900949.

Reif:1998:EAS

[Rei98] J. H. Reif. Efficient approximate solution of sparse linear systems. Computers and Mathematics with Applications, 36(9):37–58, November 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001916.

See errata [Rei99b].

Reid:1999:ODE

[Rei99a]  D. J. Reid.  Optimising

the distributed execution of join queries in polynomial time. Computers and Mathematics with Applications, 37 (3):105–126, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000516.

Reif:1999:EEA

[Rei99b] J. H. Reif. Errata to “Efficient approximate solution of sparse linear systems”. Computers and Mathematics with Applications, 38(2):141, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001881.

See [Rei98].

Repaci:1991:NIH

[Rèp91] Antonino Rèpaci. A nonlinear inverse heat transfer problem. Computers and Mathematics with Applications, 21 (11–12):139–143, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190115K.

Resta:1991:APC

[Res91] G. Resta. The average parallel complexity of Cholesky factorization. Computers and Mathematics with Applications, 22(9):27–33, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190204H.

Ramm:1991:JFM

[RF91] A. G. Ramm and V. I.

Fabrikant. Justification of Fabrikant’s method for solving mixed problems of potential theory. Computers and Mathematics with Applications, 22(6):97–103, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190006P.

Roman-Flores:1999:LCF

[RFRM99] H. Román-Flores and M. Rojas-Medar. Level-continuity of functions and applications.

Computers and Mathematics with Applications, 38 (3–4):143–149, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900190X.

Read:1990:DDP

[RG90a] E. G. Read and J. A. George. Dual dynamic programming for linear production/inventory systems. Computers and Mathematics with Applications, 19(11): 29–42, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090146B.

Rodin:1990:FFC

[RG90b] E. Y. Rodin and D. Geist. Flight and fire control with logic programming. Computers and Mathematics with Applications, 20(9–10):15– 27, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090109W.

Ruggiero:1990:IML

[RG90c] V. Ruggiero and E. Galligani. An iterative method for large sparse linear systems on a vector computer. Computers and Mathematics with Applications, 20(1):25–28, ????

1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090065R.

Ruggiero:1992:PAS

[RG92] V. Ruggiero and E. Galligani. A parallel algorithm for solving block tridiagonal linear systems. Computers and Mathematics with Applications, 24(4):15–21, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290003Z.

Read:1991:ENE

[RGBK91] E. G. Read, J. A. George, C. H. Bannister, and R. J. Kaye. Explanatory note to the Editors. Computers and Mathematics with Applications, 22(12):77, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901503.

Radev:1996:NSP

[RGBL96] K. B. Radev, C. E. A. Grigorescu, K. Berovski, and B. Logofatu. Nonlinear signal processing in bioenergetics. Computers and Mathematics with Applications, 32 (11):89–91, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002003.

Rheinboldt:1996:MSA

[Rhe96] W. C. Rheinboldt. MAN-PAK: a set of algorithms for computations on implicitly defined manifolds. Computers and Mathematics with Applications, 32 (12):15–28, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002040.

Rheinboldt:1997:SAE

[Rhe97] W. C. Rheinboldt. Solving algebraically explicit DAEs with the MANPAK-manifold-algorithms. Computers and Mathematics with Applications, 33(3):31–43, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002350.

Rosa-Hatko:1991:SMQ

[RHG91] Wanda Rosa-Hatko and Eldon Gunn. Some models of queueing control with switchover. Computers and Mathematics with Applications, 21(11–12):91–110, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190111G.

Rojo:1990:PBF

[RHRZ90] H. J. Rojo, P. L. Huerta, J. B. Rojo, and F. Zamorano. Polynomial basis functions for curved elements using hyperbolas. Computers and Mathematics with Applications, 19(2):47–54, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900066.

Rao:1994:PSA

[RI94] R. L. Rao and S. S. Iyengar. Bin-packing by simulated annealing. Computers and Mathematics with Applications, 27(5):71–82, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900779.

Ricci:1995:IAG

[Ric95] P. E. Ricci. Improving the asymptotics for the greatest zeros of Hermite polynomials. Computers and Mathematics with Applications, 30(3– 6):409–416, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001166.

Rivlin:1997:SRM

[Riv97]  T. J. Rivlin.  Some re-

marks on multivariate Chebyshev polynomials. Computers and Mathematics with Applications, 33(1–2):223– 228, January 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002325.

Rizk:1995:ASI

[Riz95] M. M. Rizk. Approximate solution of initial-value problem for retarded differential equations on complex domain. Computers and Mathematics with Applications, 30(8):19–27, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500133J.

Rojo:1994:DSU

[RJ94] O. Rojo and R. Jiménez. A decreasing sequence of upper bounds for the Perron root. Computers and Mathematics with Applications, 28(8):9–15, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001669.

Ramachandran:1998:MIU

[RK98] P. A. Ramachandran and S. R. Karur. Multidimensional interpolation using osculatory radial basis functions. Computers and Mathematics with Applications, 35 (11):63–73, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000844.

Ramm:1991:NSS

[RL91] A. G. Ramm and Peiqing Li. Numerical solution of some integral equations in distributions. Computers and Mathematics with Applications, 22(1):1–11, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900205.

Raghavan:1998:ACA

[RM98a] J. Raghavan and V. S.

Manoranjan. Arnoldi–Chebyshev approach for convection– diffusion computations. Computers and Mathematics with Applications, 36(2):95–105, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001205.

Rojo:1998:SRH

[RM98b] O. Rojo and E. Montaño.

Some results on h-hypertournament matrices. Computers and Mathematics with Applications, 35(4):89–93, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002915.

Roberts:1992:EG

[Rob92] Don D. Roberts. The existential graphs. Computers and Mathematics with Applications, 23(6–9):639– 663, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901274.

Robinson:1997:SNS

[Rob97] M. P. Robinson. The solution of nonlinear Schrödinger equations using orthogonal spline collocation. Computers and Mathematics with Applications, 33(7):39–57, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000424.

See corrigendum [Ano98-165].

Rodin:1990:SSS

[Rod90] Erwin Y. Rodin. Software survey section. Computers and Mathematics with Applications, 19(5):I–II, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090106T.

Rodin:1991:SPEa

[Rod91a]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 22 (11):1–2, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900283.

Rodin:1991:SPEb

[Rod91b]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 22 (12):1–2, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190141P.

Rodin:1992:SPEa

[Rod92a] Ervin Y. Rodin. Speed of publication — an editorial. Computers and Mathematics with Applications, 23(1): 1–2, January 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290074R.

Rodin:1992:SPEb

[Rod92b]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 23 (10):1–2, May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290049N.

Rodin:1992:SPEc

[Rod92c]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 23 (12):1–2, June 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290087X.

Rodin:1992:SPEd

[Rod92d] Ervin Y. Rodin. Speed of publication — an editorial. Computers and Mathematics with Applications, 24(4):1–2, 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Rodin:1992:SPEe

[Rod92e] Ervin Y. Rodin. Speed of publication — an editorial. Computers and Mathematics with Applications, 24(4): 1–2, August 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290001X.

Rodin:1992:SPEf

[Rod92f]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 24(10):1–2, November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900149.

Rodin:1992:SPEg

[Rod92g]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 24(11):1–2, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290025D.

Rodin:1993:BR

[Rod93a] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 26 (9):89–95, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390008J.

Rodin:1993:E

[Rod93b]  Ervin Y. Rodin.  Edito-

rial. Computers and Mathematics with Applications, 26(11):1, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900654.

Rodin:1993:SPEa

[Rod93c] Ervin Y. Rodin. Speed of publication — an editorial.

Computers and Mathematics with Applications, 25(1): 1–2, January 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390206B.

Rodin:1993:SPEb

[Rod93d] Ervin Y. Rodin. Speed of publication — an editorial. Computers and Mathematics with Applications, 25(2): 1–2, January 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390217J.

Rodin:1993:SPEc

[Rod93e]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 25(3):1–2, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390137K.

Rodin:1993:SPEd

[Rod93f]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 25(4):1–2, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390242N.

Rodin:1993:SPEe

[Rod93g]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 25 (5):1–2, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390192X.

Rodin:1993:SPEf

[Rod93h]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 25 (6):1–2, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902935.

Rodin:1993:SPEg

[Rod93i]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 25 (7):1–2, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390306G.

Rodin:1993:SPEh

[Rod93j]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 25 (9):1–2, May 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390125F.

Rodin:1993:SPEi

[Rod93k]  Ervin Y. Rodin.  Speed

of publication — an editorial. Computers and Mathematics with Applications, 25 (12):1–2, June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901804.

Rodin:1994:BRa

[Rod94a] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 27 (2):107–112, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900396.

Rodin:1994:BRb

[Rod94b] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 27(11):121–126, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901023.

Rodin:1994:BRc

[Rod94c] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 27(11):121–126, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901023.

Rodin:1994:BRd

[Rod94d] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 28 (5):111–114, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900108.

Rodin:1994:BRe

[Rod94e] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 28 (7):115–120, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900116.

Rodin:1994:SPA

[Rod94f] Ervin Y. Rodin. Széchenyi prize for Arató Mátyás. Computers and Mathematics with Applications, 28(4):1, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001219.

Rojo:1990:NMS

[Roj90] O. Rojo. A new method for solving symmetric circulant tridiagonal systems of linear equations. Computers and Mathematics with Applications, 20(12):61–67, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090165G.

Rojo:1999:FBS

[Roj99] O. Rojo. Further bounds for the smallest singular value and the spectral condition number. Computers and Mathematics with Applications, 38(7–8):215– 228, October 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002527.

Rokhlin:1990:EPC

[Rok90] V. Rokhlin. End-point corrected trapezoidal quadrature rules for singular functions. Computers and Mathematics with Applications, 20(7):51–62, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090348N.

Romanenko:1998:ACD

[Rom98] E. Romanenko. Attractors of continuous difference equations. Computers and Mathematics with Applications, 36(10–12):377–390, November/December 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800382.

Roth:1996:MAT

[Rot96] R. Roth. Music and animation tool kit: Modules for computer multimedia composition. Computers and Mathematics with Applications, 32(1):137–144, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600096X.

Rouff:1992:CSF

[Rou92] Marc Rouff. The computation of C^{k} spline functions. Computers and Mathematics with Applications, 23 (1):103–110, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290085V.

Rozsa:1996:P

[Róz96] Pál Rózsa. Preface. Computers and Mathematics with Applications, 31(4–5):xiii, February/March 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196904411.

Rozsa:1999:KPT

[Róz99] P. Rózsa. Kronecker polynomials and their applications. Computers and Mathematics with Applications, 38 (9–10):1–10, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002576.

Rosenberg:1997:MNS

[RP97] M. Rosenberg and T. Petrosky. Microscopic nonequilibrium structure in quantum fields and H-functions. Computers and Mathematics with Applications, 34(5– 6):533–557, September 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001521.

Rand:1990:AMU

[RR90] O. Rand and A. Rosen. Approximate model of the unsteady lift of thin airfoils performing arbitrary motion in the presence of an arbitrary downwash distribution. Computers and Mathematics with Applications, 20(7):31–43, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090346L.

Ratschek:1991:ITG

[RR91] H. Ratschek and J. Rokne.

Interval tools for global optimization. Computers and Mathematics with Applications, 21(6–7):41–50, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901592.

Romanov:1997:SAG

[RR97a] R. V. Romanov and G. E.

Rudin. Scattering on p-adic graphs. Computers and Mathematics with Applications, 34(5–6):587–597, September 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001557.

Ruland:1997:PDP

[RR97b] K. S. Ruland and E. Y.

Rodin. The pickup and delivery problem: Faces and branch-and-cut algorithm. Computers and Mathematics with Applications, 33 (12):1–13, June 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000904.

Rojo:1999:RBE

[RRS99] H. Rojo, O. Rojo, and R. Soto. Related bounds for the extreme eigenvalues. Computers and Mathematics with Applications, 38(7– 8):229–242, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002539.

Ruymgaart:1990:IES

[RS90] P. A. Ruymgaart and T. T. Soong. An integral equation in systems theory. Computers and Mathematics with Applications, 19(11):135–140, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090156E.

Rojo:1992:NCA

[RS92] O. Rojo and R. Soto. New conditions for the additive inverse eigenvalue problem for matrices. Computers and Mathematics with Applications, 23(11):41–46, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290067R.

Reyna:1993:DTL

[RS93] L. G. Reyna and J. R. Sobehart. Dynamics of twolevel ions in a laser beam.

Computers and Mathematics with Applications, 25(3): 37–42, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390141H.

Ravindran:1994:TSI

[RS94] R. Ravindran and S. Sundar. Two shock interaction using new theory of shock dynamics. Computers and Mathematics with Applications, 28(8):37–47, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001685.

Raghavarao:1995:NPS

[RS95] C. V. Raghavarao and S. T.

P. T. Srinivas. A note on parametric spline function approximation. Computers and Mathematics with Applications, 29(12):67–73, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000598.

Raina:1996:CSA

[RS96a] R. K. Raina and H. M. Srivastava. A certain subclass of analytic functions associated with operators of fractional calculus. Computers and Mathematics with Applications, 32(7):13–19, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001514.

Rojo:1996:PRB

[RS96b] O. Rojo and R. Soto. Perron root bounding for nonnegative persymmetric matrices. Computers and Mathematics with Applications, 31 (12):69–76, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000776.

Ravindran:1997:MLE

[RS97] R. Ravindran and S. Sundar. Mode locking for an externally excited droplet. Computers and Mathematics with Applications, 33(11): 21–33, June 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000850.

Raina:1999:SSA

[RS99a] R. K. Raina and H. M. Srivastava. Some subclasses of analytic functions associated with fractional calculus operators. Computers and Mathematics with Applications, 37(9):73–84, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001157.

Raina:1999:UPC

[RS99b] R. K. Raina and H. M. Srivastava. A unified presentation of certain subclasses of prestarlike functions with negative coefficients. Computers and Mathematics with Applications, 38(11–12):71– 78, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002862.

Rheinboldt:1999:CSS

[RS99c] W. C. Rheinboldt and B. Simeon.

Computing smooth solutions of DAEs for elastic multibody systems. Computers and Mathematics with Applications, 37(6):69–83, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000772.

Rojo:1995:LEE

[RSAR95] O. Rojo, R. L. Soto, T. Avila, and H. Rojo. Localization of eigenvalues in elliptic regions. Computers and Mathematics with Applications, 29(7):3–11, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500013O.

Raduenz:1993:AAB

[RSC93] B. D. Raduenz, B. W. Suter, and E. R. Christensen. Analysis of an Ada based version of Glassman’s general N-point Fast Fourier Transform. Computers and Mathematics with Applications, 26 (2):61–65, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390322M.

Ravindran:1994:LTB

[RSP94] R. Ravindran, S. Sundar, and P. Prasad. Long time behaviour of the solution of a system of equations from new theory of shock dynamics. Computers and Mathematics with Applications, 27(12):91–104, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900892.

Rojo:1993:NEE

[RSR93] Oscar Rojo, Ricardo L. Soto, and Hector Rojo. New eigenvalue estimates for complex matrices. Computers and Mathematics with Applications, 25(3):91–97, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390147N.

Rojo:1998:BSR

[RSR98] O. Rojo, R. Soto, and H. Rojo. Bounds for the spectral radius and the largest singular value. Computers and Mathematics with Applications, 36(1):41–50, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001072.

Raghavarao:1994:NAC

[RSS94a] C. V. Raghavarao, Y. V. S. S.

Sanyasiraju, and S. Suresh. A note on application of cubic splines to two point boundary value problems. Computers and Mathematics with Applications, 27(11):45–48, July 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900965.

Rassias:1994:SGF

[RSS94b] Th. M. Rassias, S. N. Singh, and H. M. Srivastava. Some q-generating functions associated with basic multiple hypergeometric series. Computers and Mathematics with Applications, 27(1): 33–39, January 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900043.

Ruedin:1998:CSS

[RSV98] A. M. C. Ruedin, D. C. Skigin, and R. Vaillancourt. Computer solution of the scattering problem for a groove in a metallic plane using the modal method. Computers and Mathematics with Applications, 35(11): 98–119, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000881.

Reif:1994:AKP

[RT94] J. H. Reif and S. R. Tate. Approximate kinodynamic planning using L_{2}-norm dynamic bounds. Computers and Mathematics with Applications, 27(5):29–44, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900744.

Rigas:1996:SAT

[RT96] A. G. Rigas and D. S. Tsitsis. Spectral analysis techniques of stationary point processes: Extensions and applications to neurophysiological problems. Computers and Mathematics with Applications, 32 (11):93–99, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002015.

Reutskiy:1999:QTS

[RT99a] S. Yu. Reutskiy and B. Tirozzi.

Quasi-Trefftz spectral method for separable linear elliptic equations. Computers and Mathematics with Applications, 37(2):47–74, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002533.

Ronto:1999:MTC

[RT99b] M. Rontó and Á. Tuzson.

Modification of trigonometric collocation method for impulsive periodic BVP. Computers and Mathematics with Applications, 38(9–10):117– 123, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002679.

Rehbock:1996:OSF

[RTJ96] V. Rehbock, K. L. Teo, and L. S. Jennings. Optimal and suboptimal feedback controls for a class of nonlinear systems. Computers and Mathematics with Applications, 31(6):71–86, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000077.

Ruan:1991:ONO

[Rua91] Shigui Ruan. Oscillations of nth-order functional differential equations. Computers and Mathematics with Applications, 21(2–3):95–102, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190086J.

Ruzevic:1990:LPF

[Ruz90] N. A. Ruzevic. Limit properties of functionals on irregular semi-Markovian processes. Computers and Mathematics with Applications, 19(1):133–139, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090092X.

Ramos:1994:QMA

[RV94] J. I. Ramos and F. R. Villatoro. A quantum mechanics analogy for the nonlinear Schrödinger equation in the finite line. Computers and Mathematics with Applications, 28(4):3–17, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001227.

Ramm:1991:IS

[RW91] A. G. Ramm and O. L. Weaver. 3D inverse scattering. Computers and Mathematics with Applications, 22(4–5):1–25, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190129R.

Rjasanow:1998:GCM

[RW98] S. Rjasanow and W. Wagner. A generalized collision mechanism for stochastic particle schemes approximating Boltzmann-type equations. Computers and Mathematics with Applications, 35 (1–2):165–178, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002666.

Rouff:1991:ARE

[RZ91] Marc Rouff and Wen Li Zhang. An application of RE-DUCE: the elimination process for bilinear or linear systems. Computers and Mathematics with Applications, 21 (2–3):149–157, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190093J.

Rouff:1994:SFL

[RZ94] M. Rouff and W.-L. Zhang.

C^{k} spline functions and linear operators. Computers and Mathematics with Applications, 28(4):51–59, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400126X.

Rizk:1995:ASN

[RZ95] M. M. Rizk and S. L. Zaher. Approximate solutions of nonlinear integrodifferential equations on complex domain. Computers and Mathematics with Applications, 29(10):69–78, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000473.

Riganti:1996:IPE

[RZ96] R. Riganti and M. G. Zavattaro. Invariant probability and entropy measures for multidimensional maps of chaotic attractors. Computers and Mathematics with Applications, 32 (10):7–16, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196001824.

Sheng:1994:MMH

[SA94] Q. Sheng and R. P. Agarwal. Monotone methods for higher-order partial difference equations. Computers and Mathematics with Applications, 28(1–3):291– 307, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001170.

Srivastava:1995:SAF

[SA95] H. M. Srivastava and M. K.

Aouf. Some applications of fractional calculus operators to certain subclasses of prestarlike functions with negative coefficients. Computers and Mathematics with Applications, 30(1):53–61, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000679.

Sadeh:1996:ODC

[Sad96a] I. Sadeh. Optimal data compression algorithm. Computers and Mathematics with Applications, 32(5):57–72, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001356.

Sadeh:1996:PAI

[Sad96b] I. Sadeh. Polynomial approximation of images. Computers and Mathematics with Applications, 32(5):99–115, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001381.

Sakell:1990:ULT

[Sak90]  L. Sakell.  Use of local

time stepping with full pseudospectral solutions to the Navier–Stokes equations of motion for flows with shock waves. Computers and Mathematics with Applications, 19(8–9):85–94, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090266M.

Salzer:1991:CPA

[Sal91] Herbert E. Salzer. Confluent Prony approximation. Computers and Mathematics with Applications, 22(10): 25–31, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190189B.

See erratum [Sal92b].

Salkauskas:1992:MLS

[Sal92a] K. Salkauskas. Moving least squares interpolation with thin-plate splines and radial basis functions. Computers and Mathematics with Applications, 24(12):177– 185, December 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290178K.

Salzer:1992:ECP

[Sal92b] H. E. Salzer. Erratum to “Confluent Prony approximation”: Computers Math. Applic., Vol. 22, No. 10, pp. 25–31, 1991. Computers and Mathematics with Applications, 24(3):113, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902187.

See [Sal91].

Sanchez:1998:AFO

[Sán98] F. J. Sánchez. Application of a first-order operator splitting method to Bingham fluid flow simulation. Computers and Mathematics with Applications, 36(3):71–86, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001308.

Sarafyan:1990:NAC

[Sar90] D. Sarafyan. New algorithms for the continuous approximate solutions of ordinary differential equations and the upgrading of the order of the processes. Computers and Mathematics with Applications, 20(1):77–100, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090072R.

Sarafyan:1994:ASO

[Sar94] D. Sarafyan. Approximate solution of ordinary differential equations and their systems through discrete and continuous embedded Runge– Kutta formulae and upgrading of their order. Computers and Mathematics with Applications, 28(10–12):353–384, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194002010.

Samareh-Abolhassani:1992:PAA

[SAS92] Jamshid Samareh-Abolhassani and Robert E. Smith. A practical approach to algebraic grid adaption. Computers and Mathematics with Applications, 24(5–6):69–81, September 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290042G.

Syam:1999:CCT

[SAS99] M. I. Syam and S. M. Al-Sharo. Collocationcontinuation technique for solving nonlinear ordinary boundary value problems. Computers and Mathematics with Applications, 37(10): 11–17, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001212.

Saunderson:1992:MIF

[Sau92] Houston C. Saunderson. Multiquadric interpolation of fluid speeds in a natural river channel. Computers and Mathematics with Applications, 24(12):187– 193, December 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290179L.

Saunders:1995:BCG

[Sau95] B. V. Saunders. A boundary conforming grid generation system for interface tracking. Computers and Mathematics with Applications, 29(10):1–17, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000406.

Sawaki:1992:OEP

[Saw92]  Katsushige Sawaki.  Op-

timal exercise policies for call options and their valuation. Computers and Mathematics with Applications, 24(1–2):141–146, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290239E.

Sayas:1999:NSS

[Say99] F.-J. Sayas. The numerical solution of Symm’s equation on smooth open arcs by spline Galerkin methods. Computers and Mathematics with Applications, 38(9– 10):87–99, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002643.

Steppe:1997:FSM

[SB97] J. M. Steppe and K. W. Bauer, Jr. Feature saliency measures. Computers and Mathematics with Applications, 33(8):109–126, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219700059X.

Sundar:1999:CKS

[SB99] S. Sundar and B. K. Bhagavan. Comparison of Krylov subspace methods with preconditioning techniques for solving boundary value problems. Computers and Mathematics with Applications, 38 (11–12):197–206, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002989.

Sundar:1999:CEL

[SBD99] S. Sundar, B. K. Bhagavan, and A. Datta. Computing eigenvalues: Lanczos algorithm with a new recursive partitioning method. Computers and Mathematics with Applications, 38(5– 6):99–107, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002175.

Sundar:1999:CLC

[SBS99] S. Sundar, B. K. Bhagavan, and K. S. Sastri. Comparison of Lanczos and CGS solvers for solving numerical heat transfer problems. Computers and Mathematics with Applications, 37(8):107–117, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001042.

Song:1998:PGS

[SC98a] Xinyu Song and Lansun Chen. Persistence and global stability for nonautonomous predator–prey system with diffusion and time delay. Computers and Mathematics with Applications, 35(6): 33–40, March 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000157.

Stefanov:1998:MCS

[SC98b] S. Stefanov and C. Cercignani. Monte Carlo simulation of the propagation of a disturbance in the channel flow of a rarefied gas. Computers and Mathematics with Applications, 35(1–2):41–53, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002575.

Soto:1993:PDP

[SCB^{+}93] J. L. Cruz Soto, M. C. Calzada Canalejo, M. Marín Beltrán, I. Luque Ruiz, and M. A. Gómez-Nieto. Parallelization of differential problems by partitioning method (synchronized algorithm). Computers and Mathematics with Applications, 26(2): 25–31, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390317O.

Soto:1993:PAS

[SCBFC93] J. L. Cruz Soto, M. C. Calzada Canalejo, M. Marín Beltrán, and E. Fernández-Cara. A parallel algorithm for solving the incompressible Navier– Stokes problems. Computers and Mathematics with Applications, 25(9):51–58, May 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390132F.

Soto:1998:CRP

[SCBFC98] J. L. Cruz Soto, M. C. Calzada Canalejo, M. Marín Beltrán, and E. Fernández-Cara. A convergence result for a parallel algorithm for solving the Navier–Stokes equations. Computers and Mathematics with Applications, 35(4): 71–88, February 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002903.

Soto:1994:NAM

[SCBGN94] J. L. Cruz Soto, M. C. Calzada Canalejo, M. Marín Beltrán, and M. A. Gómez-Nieto. Numerical analysis of the M.P.P. model of turbulence. Computers and Mathematics with Applications, 27(5): 17–28, March 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900736.

Sun:1997:NCT

[SCC97] Hai-Wei Sun, R. H. Chan, and Qian-Shun Chang. A note on the convergence of the two-grid method for Toeplitz systems. Computers and Mathematics with Applications, 34(1):11–18, July 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000941.

Song:1997:AFT

[SCG^{+}97] Y. Song, W. Chui, J. Glimm, B. Lindquist, and F. Tangerman. Applications of front tracking to the simulation of resin transfer molding. Computers and Mathematics with Applications, 33(9): 47–60, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000655.

Schwabe:1990:GPH

[Sch90] C. Schwabe. The genomic potential hypothesis and phasestate mathematics. Computers and Mathematics with Applications, 20(4–6):287– 301, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090334G.

Schafer:1992:REP

[Sch92a] Martin Schäfer. Resource extraction and production of a substitute. Computers and Mathematics with Applications, 24(8–9):195–207, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290198Q.

Schneider:1992:SSU

[Sch92b] Wolfgang Schneider. Systems of seemingly unrelated regression equations with time varying coefficients — an interplay of Kalman filtering, scoring, EM- and MINQUE-method. Computers and Mathematics with Applications, 24(8– 9):1–16, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290183I.

Schiesser:1994:MLS

[Sch94] W. E. Schiesser. Method of lines solution of the Korteweg–de Vries equation. Computers and Mathematics with Applications, 28(10–12):147–154, November/December 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001901.

Schmidt:1996:SAC

[Sch96] J. W. Schmidt. Staircase algorithm and construction of convex spline interpolants up to the continuity C^{3}. Computers and Mathematics with Applications, 31(4–5): 67–79, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195002189.

Scheiber:1999:NST

[Sch99] E. Scheiber. The numerical solution of Theodorsen integral equation. Computers and Mathematics with Applications, 38(9–10):221– 231, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002771.

Scott:1996:IEP

[Sco96]  C. L. Scott.  Interactive

electroacoustic performance transactional composition. Computers and Mathematics with Applications, 32(1): 79–83, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000892.

Scroggs:1991:IMS

[Scr91] Jeffrey S. Scroggs. An iterative method for systems of nonlinear hyperbolic equations. Computers and Mathematics with Applications, 21(5):137–144, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190223Q.

Sherif:1993:DTN

[SD93] Yosef S. Sherif and Roger G.

Dear. The development and testing of a new composite random number generator. Computers and Mathematics with Applications, 25(1):3–16, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390207C.

See remarks and improvements [Sez95].

Schmidt:1995:DTB

[SD95] F. Schmidt and P. Deuflhard. Discrete transparent boundary conditions for the numerical solution of Fresnel’s equation. Computers and Mathematics with Applications, 29(9):53–76, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500037Y.

Sebo:1991:PTP

[Seb91] András Sebö. A particular timetable problem: Terminal scheduling. Computers and Mathematics with Applications, 21(1):137–156, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190239Z.

Semper:1993:FEM

[Sem93]  B. Semper.  Finite ele-

ment methods for suspension bridge models. Computers and Mathematics with Applications, 26(5):77–91, September 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193900768.

Sengupta:1992:FSA

[Sen92] Jati K. Sengupta. A fuzzy systems approach in data envelopment analysis. Computers and Mathematics with Applications, 24(8–9):259–266, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290203T.

Serra:1993:MIM

[Ser93]  S. Serra.  Multi-iterative

methods. Computers and Mathematics with Applications, 26(4):65–87, August 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390035T.

Sezgin:1995:SRN

[Sez95]  F. Sezgin.  Some re-

marks on a new composite random number generator. Computers and Mathematics with Applications, 30 (11):125–130, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500169Y.

See [SD93].

Slaney:1995:ARE

[SFS95] J. Slaney, M. Fujita, and M. Stickel. Automated reasoning and exhaustive search: Quasigroup existence problems. Computers and Mathematics with Applications, 29 (2):115–132, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400219B.

Sheu:1999:AEE

[SFT99] T. W. H. Sheu, C. C. Fang, and S. F. Tsai. Application of an element-by-element BiCGSTAB iterative solver to a monotonic finite element model. Computers and Mathematics with Applications, 37(3):57–70, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000462.

Snyman:1992:PFS

[SFY92] J. A. Snyman, C. Frangos, and Y. Yavin. Penalty function solutions to optimal control problems with general constraints via a dynamic optimisation method. Computers and Mathematics with Applications, 23(11):47–55, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290068S.

Sztrik:1990:QMT

[SG90] J. Sztrik and T. Gál. A queueing model for a terminal system subject to breakdowns. Computers and Mathematics with Applications, 19(1):143–147, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090094Z.

Shibata:1995:EBG

[SG95a] Y. Shibata and Y. Gonda. Extension of de Bruijn graph and Kautz graph. Computers and Mathematics with Applications, 30 (9):51–61, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500146P.

Srivastava:1995:SFG

[SG95b] H. M. Srivastava and L. C. Gupta. Some families of generating functions for the Jacobi polynomials. Computers and Mathematics with Applications, 29(4):29–35, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219400235D.

Santis:1993:KFA

[SGS93] A. De Santis, A. Germani, and C. Scoglio. Kalman filter approach to solution of rational expectations models. Computers and Mathematics with Applications, 25 (12):39–47, June 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390184W.

Szabo:1999:P

[SGS99] Barna Szabó, Aurél Galántai, and György Szeidl. Preface. Computers and Mathematics with Applications, 38 (9–10):xiii, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002564.

Sun:1994:KGA

[SH94] Hung-Min Min Sun and Tzonelih Hwang. Key generation of algebraic-code cryptosystems. Computers and Mathematics with Applications, 27(2):99–106, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900388.

Shi:1995:SAC

[SH95a] J. Shi and R. Huotari. Simultaneous approximation from convex sets. Computers and Mathematics with Applications, 30(3–6):197– 206, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195000976.

Srivastava:1995:FIF

[SH95b] H. M. Srivastava and M. A. Hussain. Fractional integration of the H-function of several variables. Computers and Mathematics with Applications, 30(9):73–85, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500148R.

Skrien:1996:CTL

[SH96] D. Skrien and J. Hallstrom. Composers toolbox, a Lispbased precomposition environment: general system overview and an introduction to the package’s new Krayola MIDI paint module. Computers and Mathematics with Applications, 32(1):49–55, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000867.

Shamardan:1990:CFD

[Sha90a] A. B. Shamardan. Central finite difference schemes for nonlinear dispersive waves. Computers and Mathematics with Applications, 19(4): 9–15, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901335.

Shamardan:1990:NTN

[Sha90b] A. B. Shamardan. The numerical treatment of the nonlinear Schrödinger equation. Computers and Mathematics with Applications, 19(7): 67–73, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090195P.

Shao:1990:ILA

[Sha90c] P.-L. Shao. An improved Lanczos algorithm for solving ill-conditioned linear equations. Computers and Mathematics with Applications, 20(12):25–33, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090161C.

Sharma:1990:NAQ

[Sha90d] R. K. Sharma. Numerical analysis of a quasi linear differential equation with weak diffusion. Computers and Mathematics with Applications, 19(2):55–60, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190900077.

Shastri:1992:SCM

[Sha92] Lokendra Shastri. Structured connectionist models of semantic networks. Computers and Mathematics with Applications, 23(2–5):293– 328, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901458.

Shariff:1995:CCG

[Sha95] M. H. B. M. Shariff. A constrained conjugate gradient method and the solution of linear equations. Computers and Mathematics with Applications, 30 (11):25–37, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500161Q.

Shaidurov:1996:SER

[Sha96a] V. V. Shaidurov. Some estimates of the rate of convergence for the cascadic conjugate-gradient method. Computers and Mathematics with Applications, 31(4– 5):161–171, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002286.

Shawagfeh:1996:AAS

[Sha96b]  N. T. Shawagfeh.  Ana-

lytic approximate solution for a nonlinear oscillator equation. Computers and Mathematics with Applications, 31(6):135–141, March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000120.

Shampine:1998:LCL

[Sha98]  L. F. Shampine.  Linear

conservation laws for ODEs. Computers and Mathematics with Applications, 35(10): 45–53, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000716.

Shampine:1999:CLN

[Sha99a] L. F. Shampine. Conservation laws and the numerical solution of ODEs, II. Computers and Mathematics with Applications, 38(2):61–72, July 1999. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001832.

Srivastava:1999:UPS

[SHA99b] H. M. Srivastava, H. M. Hossen, and M. K. Aouf. A unified presentation of some classes of meromorphically multivalent functions. Computers and Mathematics with Applications, 38(11– 12):63–70, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002850.

Sheikh:1992:IEF

[She92] M. G. El Sheikh. The integral equation formulation of mixed finite time-dependent elastic problems. Computers and Mathematics with Applications, 24(4):3–14, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290002Y.

Sheikh:1993:FMS

[She93] M. G. El Sheikh. The finite mixed Sturm–Liouville problems of the Dirichlet– Newton type. Computers and Mathematics with Applications, 25(5):3–10, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390193Y.

Shen:1994:MMM

[She94] Z. Shen. A mathematical model for mesh’s dynamic behavior. Computers and Mathematics with Applications, 28 (6):25–32, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001499.

Shekhtman:1995:IIF

[She95] B. Shekhtman. Interpolation of individual functions. Computers and Mathematics with Applications, 30(3– 6):191–196, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000968.

Shen:1998:ADG

[She98] Z. Shen. The average diameter of general tree structures. Computers and Mathematics with Applications, 36 (7):111–130, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800176X.

Shen:1999:NSI

[She99] Shih-Yu Shen. A numerical study of inverse heat conduction problems. Computers and Mathematics with Applications, 38(7–8):173– 188, October 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002485.

Shibberu:1994:TDH

[Shi94] Y. Shibberu. Time-discretization of Hamiltonian dynamical systems. Computers and Mathematics with Applications, 28(10–12):123–145, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001898.

Shi:1995:SOP

[Shi95a] Y. Shi. Studies on optimumpath ratios in multicriteria De Novo programming problems. Computers and Mathematics with Applications, 29(5):43–50, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400247I.

Shisha:1995:GGR

[Shi95b] O. Shisha. The genesis of the generalized Riemann integral. Computers and Mathematics with Applications, 30(3– 6):207–211, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000984.

Shilnikov:1997:HPL

[Shi97] A. L. Shil’nikov. Homoclinic phenomena in laser models. Computers and Mathematics with Applications, 34(2–4): 245–251, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001260.

Shin:1998:FEA

[Shi98] J. Y. Shin. Finite-element approximation of a fourthorder differential equation. Computers and Mathematics with Applications, 35(8): 95–100, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000479.

Shkoller:1997:AHS

[Shk97] S. Shkoller. An approximate homogenization scheme for nonperiodic materials. Computers and Mathematics with Applications, 33(4): 15–34, February 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000035.

Shohdohji:1999:OQT

[SHK99] T. Shohdohji, Y. Hoshino, and N. Kutsuwada. Optimization of quantization table based on visual characteristics in DCT image coding. Computers and Mathematics with Applications, 37 (11–12):225–232, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001601.

Seither:1990:SNT

[SHT^{+}90] R. L. Seither, D. Hearne, D. Trent, D. C. Mikulecky, and I. D. Goldman. SPICE2 network thermodynamic simulation of antifolate effects on purine and pyrimidine biosynthesis: Exploring the role of tetrahydrofolate cofactor depletion versus dihydrofolate feedback inhibition. Computers and Mathematics with Applications, 20 (4–6):87–101, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090318E.

Shu:1993:AII

[Shu93a]  H.-B. Shu.  Application

of the integral invariant of Poincarè-Cartan to contact systems. Computers and Mathematics with Applications, 26(2):51–59, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390321L.

Shu:1993:NAG

[Shu93b] Hai-Bin Shu. A new approach to generating functions for contact systems. Computers and Mathematics with Applications, 25(8):101–106, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390176V.

Shvachko:1995:SSG

[Shv95] K. V. Shvachko. Space saving generalization of B-trees with 23 utilization. Computers and Mathematics with Applications, 30(7):47–66, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500125I.

Sugisaka:1994:MPC

[SI94] M. Sugisaka and M. Ino.

Model predictive control based on neural identification method. Computers and Mathematics with Applications, 27(9–10):83–93, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901279.

Siboni:1998:CGA

[Sib98] S. Siboni. Complex/generalized arithmetico-geometric maps and elliptic integrals. Computers and Mathematics with Applications, 36(5):1–10, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800145X.

Sieburg:1990:CDM

[Sie90] H. B. Sieburg. The cellular device machine: Point of departure for large-scale simulations of complex biological systems. Computers and Mathematics with Applications, 20(4–6):247–267, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090332E.

Simos:1993:RKN

[Sim93a]  T. E. Simos.  Runge–

Kutta–Nyström interpolants for the numerical integration of special second-order periodic initial-value problems. Computers and Mathematics with Applications, 26 (12):7–15, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390054Y.

Simos:1993:RKF

[Sim93b] T. E. Simos. A Runge–Kutta Fehlberg method with phaselag of order infinity for initialvalue problems with oscillating solution. Computers and Mathematics with Applications, 25(6):95–101, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390303D.

Simos:1993:RKI

[Sim93c] T. E. Simos. Runge–Kutta interpolants with minimal phase-lag. Computers and Mathematics with Applications, 26(8):43–49, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390330X.

Simos:1997:ENT

[Sim97]  T. E. Simos.  An ex-

tended numerov-type method for the numerical solution of the Schrödinger equation. Computers and Mathematics with Applications, 33(10): 67–78, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000771.

Simos:1998:NHI

[Sim98] T. E. Simos. A new hybrid imbedded variable-step procedure for the numerical integration of the Schrödinger equation. Computers and Mathematics with Applications, 36(2):51–63, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001163.

Sirotkin:1997:SSP

[Sir97] V. V. Sirotkin. The solution of singularly perturbed semilinear elliptic problems by iterative domain decomposition algorithms. Computers and Mathematics with Applications, 33(11):99–116, June 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000886.

Sirotkin:1999:SSP

[Sir99] V. V. Sirotkin. The solution of singularly perturbed parabolic problems by parallel algorithms combining Crank–Nicolson scheme and overlapping Schwarz methods. Computers and Mathematics with Applications, 38 (3–4):119–142, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001911.

Savioli:1997:SAN

[SJB97] G. B. Savioli, P. M. Jacovkis, and M. S. Bidner. Stability analysis and numerical simulation of 1-D and 2d radial flow towards an oil well. Computers and Mathematics with Applications, 33 (3):121–135, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002428.

Sjogren:1992:CFS

[Sjo92] Jon A. Sjogren. Closed-form solution of decomposable stochastic models. Computers and Mathematics with Applications, 23(12):43–67, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290091U.

Sugasawa:1992:ESP

[SJS92] Yoshio Sugasawa, Qun Jin, and Koichiro Seya. Extended stochastic Petri net models for systems with parallel and cooperative motions. Computers and Mathematics with Applications, 24(1–2):119–126, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290236B.

Santiago:1991:UCR

[SK91] Marilyn Santiago and David R.

Kincaid. Using cyclic reduction on a parallel computer to improve the performance of an underwater sound implicit finite difference model. Computers and Mathematics with Applications, 21(5):83–94, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190218S.

Sandoh:1992:OBP

[SKI92] Hiroaki Sandoh, Hajime Kawai, and Toshihide Ibaraki. An optimal backup policy for a hard computer disk depending on age under availability criterion. Computers and Mathematics with Applications, 24(1–2):57–62, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290228A.

Sobol:1999:PRN

[SL99] I. M. Sobol’ and Yu. L. Levitan. A pseudo-random number generator for personal computers. Computers and Mathematics with Applications, 37(4–5):33–40, February/March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000577.

Translation of the Russian original, O datchike psevdosluchainykh chisel dlya personalnykh kompyuterov, Matematicheskoe Modelirovanie, 2 (8) (1990), pp. 119–126, with the authors reversed.

Slator:1992:SP

[Sla92]  Brian M. Slator.  Sense

and preference. Computers and Mathematics with Applications, 23(6–9):391– 402, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290114W.

Si:1997:ASI

[SLC97] Jian-Guo Si, Wen-Rong Li, and Sui Sun Cheng. Analytic solutions of an iterative functional differential equation. Computers and Mathematics with Applications, 33(6):47–51, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000308.

Sharma:1990:HSN

[SM90] R. K. Sharma and P. Majumdar. High speed numerical technique to solve coupled moving boundary value heat problems. Computers and Mathematics with Applications, 19(4):1–8, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190901324.

Simos:1994:FFS

[SM94] T. E. Simos and G. V. Mitsou. A family of four-step exponential fitted methods for the numerical integration of the radial Schrödinger equation. Computers and Mathematics with Applications, 28(4):41–50, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001251.

Simos:1995:TSM

[SM95] T. E. Simos and G. Mousadis. A two-step method for the numerical solution of the radial Schrödinger equation. Computers and Mathematics with Applications, 29 (7):31–37, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500016R.

Saied:1998:EPM

[SM98] F. Saied and G. Mahinthakumar. Efficient parallel multigrid based solvers for large scale groundwater flow simulations. Computers and Mathematics with Applications, 35 (7):45–54, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000315.

Satratzemi:1998:APM

[SMT98] M. Satratzemi, K. G. Margaritis, and C. Tsouros. An algorithm for prescribed multiple domination in arbitrary graphs. Computers and Mathematics with Applications, 35(8):109–115, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000492.

Sun:1997:PAM

[SMTS97] K. Sun, I. H. Moon, R. P.

Tewarson, and J. L. Stephenson. Parallel algorithms for multinephron renal medullary models. Computers and Mathematics with Applications, 33(6):37–45, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000291.

Stoyan:1998:CNN

[SMU98] G. Stoyan, C. Mihálykó, and Z. Ulbert. Convergence and nonnegativity of numerical methods for an integrodifferential equation describing batch grinding. Computers and Mathematics with Applications, 35(12):69–81, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000972.

Sniedovich:1990:AME

[Sni90] M. Sniedovich. An analysis of a machine executable generalized Σ-notation. Computers and Mathematics with Applications, 20(11):53–65, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090219A.

Sniedovich:1991:MEG

[Sni91a] Moshe Sniedovich. A machine executable generalized inner product and its applications. Computers and Mathematics with Applications, 21(2–3):159–165, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190094K.

Sniedovich:1991:SSV

[Sni91b]  Moshe Sniedovich.  So-

lution strategies for variance minimization problems. Computers and Mathematics with Applications, 21(2– 3):49–56, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190080N.

Sakawa:1998:IFP

[SNU98] M. Sakawa, I. Nishizaki, and Y. Uemura. Interactive fuzzy programming for multilevel linear programming problems. Computers and Mathematics with Applications, 36(2):71–86, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001187.

Snyman:1990:IFD

[Sny90] J. A. Snyman. An interior feasible direction method with constraint projections for linear programming. Computers and Mathematics with Applications, 20(12):43–54, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090163E.

Sod:1990:RCM

[Sod90]  G. A. Sod.  A random

choice method for the Stefan problem. Computers and Mathematics with Applications, 19(8–9):1–9, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090260Q.

Sod:1992:NMT

[Sod92] Gary A. Sod. A numerical model of a thermal boundary layer. Computers and Mathematics with Applications, 24 (5–6):1–10, September 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290036H.

Segawa:1992:OMR

[SOI92] Yoshiyuki Segawa, Masamitsu Ohnishi, and Toshihide Ibaraki. Optimal minimal-repair and replacement problem with age dependent cost structure. Computers and Mathematics with Applications, 24(1–2):91–101, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902338.

Sokolov:1991:NHS

[Sok91] Viatcheslav I. Sokolov. The “Natural” hierarchy of symmetry elements based on the intrinsic property of dimensionality and the corresponding tree of the symmetry groups. Computers and Mathematics with Applications, 22(7):63–64, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901824.

Solntsev:1990:ASS

[Sol90]  S. Solntsev.  On almost

sure stability of the solutions of stochastic recurrent equations. Computers and Mathematics with Applications, 19(1):141–142, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090093Y.

Soloveichik:1997:IMS

[Sol97]  Y. Soloveichik.  Iterative

method for solving finite element systems of algebraic equations. Computers and Mathematics with Applications, 33(6):87–90, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000357.

Sommeijer:1990:IRS

[Som90] B. P. Sommeijer. Increasing the real stability boundary of explicit methods. Computers and Mathematics with Applications, 19(6):37–49, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090046M.

Soni:1992:GGI

[Son92]  Bharat K. Soni.  Grid

generation for internal flow configurations. Computers and Mathematics with Applications, 24(5–6):191– 201, September 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290048M.

Souidi:1990:NST

[Sou90] R. Souidi. A new sequential test for detection of a point of change in ARMA parameters. Computers and Mathematics with Applications, 19(5):31–39, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122190900996.

Sovegjarto:1999:CSM

[Söv99] A. Sövegjártó. Conservative spline methods for secondorder IVPs of ODEs. Computers and Mathematics with Applications, 38(9–10):135– 141, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002692.

Sowa:1992:CGU

[Sow92]  John F. Sowa.  Concep-

tual graphs as a universal knowledge representation. Computers and Mathematics with Applications, 23(2–5): 75–93, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901377.

Shapiro:1992:SF

[SR92] Stuart C. Shapiro and William J.

Rapaport. The SNePS family. Computers and Mathematics with Applications, 23(2– 5):243–275, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901436.

Sarkar:1998:PDU

[SRCB98] P. Sarkar, B. K. Roy, P. P.

Choudhury, and R. Barua. Polynomial division using left shift register. Computers and Mathematics with Applications, 35(6):27–31, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000145.

Srivastava:1995:SRI

[Sri95] R. Srivastava. Some results on integrals involving generalized Jacobi and related functions. Computers and Mathematics with Applications, 30(1):47–52, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000668.

Sondak:1990:NDM

[SS90] V. K. Sondak and N. E.

Sondak. New directions for medical artificial intelligence. Computers and Mathematics with Applications, 20 (4–6):313–319, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090336I.

Sculli:1991:ASP

[SS91a] D. Sculli and Y. W. Shum. An approximate solution to the PERT problem. Computers and Mathematics with Applications, 21(8):1–7, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900467.

Stoyan:1991:CDG

[SS91b] Gisbert Stoyan and Robert Stoyan. Colouring the discretization graphs arising in the multigrid method. Computers and Mathematics with Applications, 22(7): 55–62, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901813.

Sur:1993:TPS

[SS93] Sumit Sur and Pradip K. Srimani. Topological properties of star graphs. Computers and Mathematics with Applications, 25(12):87–98, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901882.

Shahruz:1994:ASH

[SS94] S. M. Shahruz and A. L.

Schwartz. An approximate solution for homogeneous boundary-value problems with slowly-varying coefficient matrices. Computers and Mathematics with Applications, 28(6):75–82, September 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001537.

Sun:1996:ECP

[SS96] Hung-Min Sun and Shiuh-Pyng Shieh. An efficient construction of perfect secret sharing schemes for graphbased structures. Computers and Mathematics with Applications, 31(7):129–135, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000223.

Sacco:1998:FEM

[SS98a] R. Sacco and M. Stynes.

Finite element methods for convection–diffusion problems using exponential splines on triangles. Computers and Mathematics with Applications, 35(3):35–45, February 1998. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002770.

Sun:1998:SSS

[SS98b] Hung-Min Sun and Shiuh-Pyng Shieh. Secret sharing schemes for graphbased prohibited structures. Computers and Mathematics with Applications, 36 (7):131–140, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001771.

Sun:1999:RCP

[SS99] Hung-Min Sun and Shiuh-Pyng Shieh. Recursive constructions for perfect secret sharing schemes. Computers and Mathematics with Applications, 37(3):87–96, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199000498.

Srimani:1991:PCT

[SSBG91] Pradip K. Srimani, Bhabani P. Sinha, Bhargab B. Bhattacharya, and Suranjan Ghosh. Properties of a class of trivalent network graphs and optimal routing. Computers and Mathematics with Applications, 22(2):39–47, ???? 1991. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190040B.

Srivastava:1992:SGF

[SSH92] H. M. Srivastava, R. K. Saxena, and Z. Hussain. Some generating functions for the Jacobi polynomials. Computers and Mathematics with Applications, 23(10):51–56, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290054L.

Stone:1994:PSO

[SSN94] L. C. Stone, S. B. Shukla, and B. Neta. Parallel satellite orbit prediction using a workstation cluster. Computers and Mathematics with Applications, 28(8):1–8, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001650.

Shi:1991:NSO

[SSZ91] Peter Shi, Meir Shillor, and Xiu-Lin Zou. Numerical solutions to one-dimensional problems of thermoelastic contact. Computers and Mathematics with Applications, 22(10):65–78, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191901938.

Shraga:1990:PEU

[ST90] Y. Shraga and R. P. Tewarson. Parameter estimation in underdetermined problems. Computers and Mathematics with Applications, 20 (4–6):325–337, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090338K.

Shores:1994:SPV

[ST94] T. Shores and C. Tiahrt. Subspace projection variants on Newton’s method. Computers and Mathematics with Applications, 27(8):7–17, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900620.

Stynes:1995:NUC

[ST95] M. Stynes and L. Tobiska. Necessary L^{2}-uniform convergence conditions for difference schemes for two-dimensional convection–diffusion problems. Computers and Mathematics with Applications, 29(4):45–53, February 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400237F.

Shilnikov:1997:SBL

[ST97] L. P. Shil’nikov and D. V.

Turaev. Simple bifurcations leading to hyperbolic attractors. Computers and Mathematics with Applications, 34 (2–4):173–193, July/August 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001235.

Shen:1998:NOC

[ST98] Jianhua Shen and Xianhua Tang. New oscillation criteria for linear delay differential equations. Computers and Mathematics with Applications, 36(6):53–61, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001618.

Stavroulakis:1995:ODD

[Sta95] I. P. Stavroulakis. Oscillations of delay difference equations. Computers and Mathematics with Applications, 29 (7):83–88, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500020Y.

Steiner:1990:CBB

[Ste90] G. Steiner. The connection between the bump number problem and flow-shop scheduling with precedence constraints. Computers and Mathematics with Applications, 20(11):13–18, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902145.

Stilman:1993:LAG

[Sti93a] B. Stilman. A linguistic approach to geometric reasoning. Computers and Mathematics with Applications, 26(7):29–57, October 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900492.

Stilman:1993:NLC

[Sti93b] B. Stilman. Network languages for complex systems. Computers and Mathematics with Applications, 26(8): 51–79, October 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390331O.

Stilman:1994:TNL

[Sti94] B. Stilman. Translations of network languages. Computers and Mathematics with Applications, 27(2):65–98, January 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219490037X.

Stilman:1996:NLI

[Sti96] B. Stilman. Network languages for intelligent control. Computers and Mathematics with Applications, 31 (3):91–118, February 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002073.

Stilman:1997:NLC

[Sti97] B. Stilman. Network languages for concurrent multiagent systems. Computers and Mathematics with Applications, 34(1):103–136, July 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001028.

Stojmenovic:1992:SSA

[Sto92] Ivan Stojmenovic. A simple systolic algorithm for generating combinations in lexicographic order. Computers and Mathematics with Applications, 24(4):61–64, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192900075.

Stoyan:1999:TDV

[Sto99] G. Stoyan. Towards discrete velte decompositions and narrow bounds for inf-sup constants. Computers and Mathematics with Applications, 38 (7–8):243–261, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002540.

Struckmeier:1998:SBV

[Str98] J. Struckmeier. Simulation of boundary value problems for the Boltzmann equation. Computers and Mathematics with Applications, 35(1– 2):179–192, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002678.

Sugiyama:1994:SDP

[Sug94] H. Sugiyama. Solving deterministic problems via stochastic approximation — a simple yet powerful numerical method. Computers and Mathematics with Applications, 27(9–10):129–144, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901317.

Sun:1991:UFD

[Sun91] Ziqi Sun. Uniqueness for formally determined inverse boundary value problems. Computers and Mathematics with Applications, 22(4– 5):67–80, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190132N.

Sun:1992:CSC

[Sun92]  Xingping Sun.  Cardinal

and scattered-cardinal interpolation by functions having non-compact support. Computers and Mathematics with Applications, 24 (12):195–200, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290180P.

Srinivasan:1991:PMDa

[SV91a] S. K. Srinivasan and R. Vasudevan. Particle multiplicity distribution a la invariant imbedding and natural scaling. Computers and Mathematics with Applications, 21(11–12):7, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190102A.

Srinivasan:1991:PMDb

[SV91b] S. K. Srinivasan and R. Vasudevan. Particle multiplicity distribution a la invariant imbedding and natural scaling. Computers and Mathematics with Applications, 22(2):59–71, ???? 1991.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900423.

Stepanov:1997:MPE

[SV97] E. Stepanov and S. A. Vavilov. The main problem of external ballistics. Computers and Mathematics with Applications, 33(5):95–101, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000229.

Saucez:1998:AML

[SVS98] P. Saucez, A. Vande Wouwer, and W. E. Schiesser. An adaptive method of lines solution of the Korteweg–de Vries equation. Computers and Mathematics with Applications, 35(12):13–25, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000935.

Sheu:1995:PDI

[SWF95] R.-L. Sheu, S.-Y. Wu, and S.-C. Fang. A primal-dual infeasible-interior-point algorithm for linear semi-infinite programming. Computers and Mathematics with Applications, 29(8):7–18, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500025T.

Shi:1996:ACS

[SWY96a] B. Shi, Z. C. Wang, and J. S. Yu. Asymptotic constancy of solutions of linear parabolic Volterra difference equations. Computers and Mathematics with Applications, 32(8):65–77, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219600168X.

See notes and corrections [Zha98].

Shi:1996:ONP

[SWY96b] B. Shi, Z. C. Wang, and J. S. Yu. Oscillation of nonlinear partial difference equations with delays. Computers and Mathematics with Applications, 32 (12):29–39, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002052.

Shi:1997:GAS

[SWY97] B. Shi, Z. C. Wang, and J. S.

Yu. Global asymptotic stability in a nonlinear nonautonomous difference equation with delays. Computers and Mathematics with Applications, 33(8):93–102, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000576.

Shi:1998:SBD

[SWY98] B. Shi, Z. C. Wang, and J. S. Yu. Stability and boundedness in difference systems with finite delay. Computers and Mathematics with Applications, 36(10–12):369–376, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800370.

Steiner:1996:LTA

[SY96] G. Steiner and J. S. Yeomans. A linear time algorithm for maximum matchings in convex, bipartite graphs. Computers and Mathematics with Applications, 31 (12):91–96, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600079X.

Syau:1997:SFM

[Sya97]  Yu-Ru Syau.  Sequences

in a fuzzy metric space. Computers and Mathematics with Applications, 33(6): 73–76, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000333.

Syau:1999:PFM

[Sya99] Yu-Ru Syau. Preinvex fuzzy mappings. Computers and Mathematics with Applications, 37(3):31–39, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000449.

Symes:1991:DSA

[Sym91]  William W. Symes.  A

differential semblance algorithm for the inverse problem of reflection seismology. Computers and Mathematics with Applications, 22 (4–5):147–178, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190140Y.

Shi:1994:GND

[SYZ94] Y. Shi, P.-L. Yu, and D. Zhang. Generating new designs using union operations. Computers and Mathematics with Applications, 27(12):105–117, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900906.

Srivastav:1991:SCS

[SZ91] R. P. Srivastav and Fenggang Zhang. Solving Cauchy singular integral equations by using general quadraturecollocation nodes. Computers and Mathematics with Applications, 21(9):59–71, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190125N.

Shao:1998:TOC

[SZ98a] Fang-Ming Shao and Lian-Chang Zhao. Topological optimization of computer network expansion with reliability constraint. Computers and Mathematics with Applications, 35(11):17–26, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000807.

Shen:1998:QSS

[SZ98b] S. S. Shen and Shunian Zhang. A quantitative study on the stability of quadratic delay difference systems. Computers and Mathematics with Applications, 35(11):1–16, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000790.

Szablowski:1990:ETA

[Sza90a] P. J. Szablowski. Expansions of E(XZ|Y + ϵX) and their applications to the analysis of elliptically contoured measures. Computers and Mathematics with Applications, 19(5):75–83, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090103Q.

Szabowski:1990:ECR

[Sza90b] P. J. Szabowski. Elliptically contoured random variables and their application to the extension of the Kalman filter. Computers and Mathematics with Applications, 19(2):61–72, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900088.

Szablowski:1991:MQF

[Sza91] Paweł J. Szabłowski. Minimization of quadratic functionals on cones in Hilbert spaces. Computers and Mathematics with Applications, 22(8):21–34, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190009S.

Shieh:1991:ODR

[SZC91] Leang S. Shieh, Jian L.

Zhang, and Norman P. Coleman. Optimal digital redesign of continuous-time controllers. Computers and Mathematics with Applications, 22(1):25–35, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190022V.

Szkodny:1995:MKI

[Szk95] T. Szkodny. Modelling of kinematics of the IRb-6 manipulator. Computers and Mathematics with Applications, 29(9):77–94, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500038Z.

Szmolyan:1990:ITS

[Szm90] P. Szmolyan. Initial transients of solutions of the semiconductors device equations. Computers and Mathematics with Applications, 19(8–9):43–58, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090263J.

Sarma:1998:UHP

[SZM98] G. Sarma, T. Zacharia, and D. Miles. Using High Performance Fortran for parallel programming. Computers and Mathematics with Applications, 35(12):41–57, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000959.

Sztrik:1992:MSB

[Szt92] J. Sztrik. Modelling of a single bus multiprocessor system operating in Markovian environments. Computers and Mathematics with Applications, 23(11):57–65, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290069T.

Szulc:1996:TSP

[Szu96]  T. Szulc.  Testing some

properties of real matrices. Computers and Mathematics with Applications, 31(4–5): 63–65, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002170.

Tait:1991:ARS

[TAD91] R. J. Tait, K. Abdella, and D. B. Duncan. Approximate Riemann solvers and waves on a nonlinear elastic string. Computers and Mathematics with Applications, 21(4):77–89, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191902483.

Tadi:1993:FEM

[Tad93] M. Tadi. Finite elements for modeling and control of a Mindlin plate. Computers and Mathematics with Applications, 25(7):81–99, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390314L.

Tadi:1994:OID

[Tad94] M. Tadi. Optimal infinitedimensional estimator and compensator for a Timoshenko beam. Computers and Mathematics with Applications, 27(6):19–32, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901074.

Taha:1991:NSN

[Tah91]  Thiab R. Taha.  A nu-

merical scheme for the nonlinear Schrödinger equation. Computers and Mathematics with Applications, 22(9): 77–84, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190208L.

Tang:1992:SDA

[Tan92] Hui-Chin Tang. Simulated division with approximate factoring for the multiple recursive generator with both unrestricted multiplier and non-Mersenne prime modulus. Computers and Mathematics with Applications, 40 (3):1173–1181, July 24, 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122103902094.

Tang:1993:GFC

[Tan93a] Y.-F. Tang. Geodesic flows on compact surfaces-as an application of Hamiltonian formalism. Computers and Mathematics with Applications, 26(1):21–33, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900838.

Tang:1993:NCR

[Tan93b]  Y.-F. Tang.  The neces-

sary condition for a Runge– Kutta scheme to be symplectic for Hamiltonian systems. Computers and Mathematics with Applications, 26 (1):13–20, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900827.

Tang:1993:NCT

[Tan93c] Yi-Fa Tang. Non-conservativity of traditional schemes for Liouville and contact systems. Computers and Mathematics with Applications, 25(8): 89–94, April 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390174T.

Tang:1993:SMS

[Tan93d] Yi-Fa Tang. The symplecticity of multi-step methods. Computers and Mathematics with Applications, 25(3):83–90, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390146M.

Tang:1994:FESa

[Tan94] Y.-F. Tang. Formal energy of a symplectic scheme for Hamiltonian systems and its applications (I). Computers and Mathematics with Applications, 27(7):31–39, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901473.

Tanikawa:1996:RCH

[Tan96]  A. Tanikawa.  The rate

of convergence of a homogeneous Markov chain arising from two-queue networks. Computers and Mathematics with Applications, 31 (2):103–109, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001972.

Tauber:1990:CRM

[Tau90a] S. Tauber. A chemical robot model. Computers and Mathematics with Applications, 20 (4–6):321–324, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090337J.

Tauber:1990:SMN

[Tau90b] S. Tauber. A systematic model for nerve transmission and memory. Computers and Mathematics with Applications, 20(4–6):227–230, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090330M.

Tadj:1999:HQS

[TBA99] L. Tadj, L. Benkherouf, and L. Aggoun. A hysteretic queueing system with random server capacity. Computers and Mathematics with Applications, 38(1):51–61, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001686.

tenCate:1995:AAF

[tC95] H. H. ten Cate. Applying abstraction and formal specification in numerical software design. Computers and Mathematics with Applications, 29(12):81–102, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000613.

Teng:1999:UPE

[TC99] Zhidong Teng and Lansun Chen. Uniform persistence and existence of strictly positive solutions in nonautonomous Lotka– Volterra competitive systems with delays. Computers and Mathematics with Applications, 37(7):61–71, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000875.

Tcheremissine:1998:CEB

[Tch98] F. G. Tcheremissine. Conservative evaluation of Boltzmann collision integral in discrete ordinates approximation. Computers and Mathematics with Applications, 35 (1–2):215–221, January 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002691.

Tandon:1991:AMF

[TCJG91] P. N. Tandon, Amita Chaurasia, V. K. Jain, and T. Gupta.

Application of magnetic fields to synovial joints. Computers and Mathematics with Applications, 22(12):33–45, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190145T.

Tsai:1993:GME

[TCS93] J. S. H. Tsai, S. S. Chen, and L. S. Shieh. Generalized matrix Euclidean algorithms for solving Diophantine equations and associated problems. Computers and Mathematics with Applications, 25(12):55–72, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390186Y.

Tsai:1994:RDI

[TCS94] J. S. H. Tsai, W. S. Chen, and L. S. Shieh. Realization, design and implementation of row-pseudoproper left matrix fraction descriptions with impulsive modes. Computers and Mathematics with Applications, 28(7):13–41, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400157X.

Torok:1999:ADP

[TD99] Cs. Török and N. D. Dikoussar. Approximation with discrete projective transformation. Computers and Mathematics with Applications, 38(9–10):211–220, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900276X.

Tanaka:1992:CDS

[TDF^{+}92] H. Tanaka, T. Dohi, H. Fujiwara, S. Osaki, and N. Kaio. Construction of a decisionsupport system for a combination of options. Computers and Mathematics with Applications, 24(1–2):135– 140, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290238D.

Tsubakitani:1993:TDM

[TE93] S. Tsubakitani and J. R. Evans. A two-dimensional mapping for the traveling salesman problem. Computers and Mathematics with Applications, 26(12):65–73, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0898122193900609.

Telegdi:1991:ISB

[Tel91] L. Telegdi. Investigation of the structure of binary variables. Computers and Mathematics with Applications, 21(1):157–168, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191902405.

Temme:1993:ANA

[Tem93]  N. M. Temme.  Asymp-

totic and numerical aspects of the noncentral chi-square distribution. Computers and Mathematics with Applications, 25(5):55–63, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901985.

Terdik:1991:BSS

[Ter91] György Terdik. Bilinear state space realization for polynomial stochastic systems. Computers and Mathematics with Applications, 22(7): 69–83, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901846.

Terceiro:2000:CKF

[Ter00]  J. Terceiro.  Comments

on “Kalman-filtering methods for computing information matrices for timeinvariant, periodic, and generally time-varying VARMA models and samples”. Computers and Mathematics with Applications, 40(2–3):405– 411, July/August 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122100001681.

See [ZM94].

Tang:1995:EPD

[TG95] D. Tang and G. Gupta. An efficient parallel dynamic programming algorithm. Computers and Mathematics with Applications, 30(8):65–74, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500138O.

Tian:1996:CAP

[TG96] Bo Tian and Yi-Tian Gao. Computer algebra, Painlevé analysis and the time-dependent-coefficient nonlinear Schrödinger equation. Computers and Mathematics with Applications, 31(11):115–119, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000685.

Tsai:1995:TTC

[TH95] So-Yenn Tsai and Tsan-Hui Hsu. Thermal transport of a continuous moving plate in a non-Newtonian fluid. Computers and Mathematics with Applications, 29(6):

99–108, March 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500011M.

Thandapani:1994:OTP

[Tha94] E. Thandapani. Oscillation theorems for perturbed nonlinear second order difference equations. Computers and Mathematics with Applications, 28(1–3):309– 316, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001189.

Thomason:1992:NSP

[Tho92] Richmond H. Thomason.

NETL and subsequent pathbased inheritance theories. Computers and Mathematics with Applications, 23(2– 5):179–204, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290140D.

Tada:1996:IFT

[TI96] M. Tada and H. Ishii. An integer fuzzy transportation problem. Computers and Mathematics with Applications, 31(9):71–87, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000442.

Tang:1994:FESb

[TL94] Yi-Fa Tang and Yong-Hong Long. Formal energy of symplectic scheme for Hamiltonian systems and its applications (II). Computers and Mathematics with Applications, 27(12):31–39, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900833.

Thissell:1991:SAC

[TM91] William R. Thissell and Patrick L. Mills. Some applications of computerized symbolic manipulation in the analysis of chemical engineering systems. Computers and Mathematics with Applications, 21(10):53–78, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900618.

Thandapani:1998:MPC

[TMGS98] E. Thandapani, M. M. S. Manuel, J. R. Graef, and P. W. Spikes. Monotone properties of certain classes of solutions of second-order difference equations. Computers and Mathematics with Applications, 36(10–12):291–297, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800308.

Tuzun:1998:CSC

[TNS98] R. E. Tuzun, D. W. Noid, and B. G. Sumpter. Computer simulation of complex, strongly coupled nanometerscale systems: Breaking the billion atom barrier. Computers and Mathematics with Applications, 35(7): 93–100, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000364.

Tandon:1991:ANT

[TNTR91] P. N. Tandon, P. Nirmala, M. Tewari, and U. S. Rana. Analysis of nutritional transport through a capillary: Normal and stenosed. Computers and Mathematics with Applications, 22(12):3–13, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190142Q.

Torabi:1990:DBC

[Tor90] M. Torabi. Decomposed block Cholesky factorization in the Karmarkar algorithm: Solving a class of super large LP problems. Computers and Mathematics with Applications, 20(2):1–7, ???? 1990.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090234B.

Torok:1994:NSL

[Tör94] C. Török. Numerical solution of linear stochastic differential equations. Computers and Mathematics with Applications, 27(4):1–10, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900507.

Toyoizumi:1992:EMS

[Toy92] Hiroshi Toyoizumi. Evaluating mean sojourn time estimates for the M/M/1 queue. Computers and Mathematics with Applications, 24(1– 2):7–15, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192902224.

Thandapani:1998:AOB

[TP98] E. Thandapani and S. Pandian. Asymptotic and oscillatory behavior of solutions of general nonlinear difference equations of second order. Computers and Mathematics with Applications, 36(10–12):413–421, November/December 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800424.

Telek:1998:ENM

[TPF98] M. Telek, A. Pfening, and G. Fodor. An effective numerical method to compute the moments of the completion time of Markov reward models. Computers and Mathematics with Applications, 36(8):59–65, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001837.

Thandapani:1999:BMP

[TR99] E. Thandapani and K. Ravi.

Bounded and monotone properties of solutions of secondorder quasilinear forced difference equations. Computers and Mathematics with Applications, 38(2):113–121, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001868.

Trazaska:1997:SSD

[Tra97] Z. W. Trazaska. State-space/ descriptor models and asymptotic behavior of continuoustime positive control systems. Computers and Mathematics with Applications, 34 (12):1–10, December 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002277.

Trench:1994:ABS

[Tre94]  W. F. Trench.  Asymp-

totic behavior of solutions of Poincaré recurrence systems. Computers and Mathematics with Applications, 28 (1–3):317–324, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001197.

Trench:1995:ABS

[Tre95a]  W. F. Trench.  Asymp-

totic behavior of solutions of asymptotically constant coefficient systems of linear differential equations. Computers and Mathematics with Applications, 30 (12):111–117, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001782.

Trench:1995:ICI

[Tre95b] W. F. Trench. Invertibly convergent infinite products of matrices, with applications to difference equations. Computers and Mathematics with Applications, 30 (11):39–46, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500162R.

Trench:1998:LAE

[Tre98]  W. F. Trench.  Linear

asymptotic equilibrium and uniform, exponential, and strict stability of linear difference systems. Computers and Mathematics with Applications, 36(10–12):261–267, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800278.

Trzaska:1992:ACN

[Trz92] Z. W. Trzaska. The algorithm for computing nonmonic second order multivariable systems. Computers and Mathematics with Applications, 24 (11):17–27, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290028G.

Tsao:1990:APL

[Tsa90] N. K. Tsao. The accuracy of a parallel LU decomposition algorithm. Computers and Mathematics with Applications, 20(7):25–30, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090345K.

Tselnik:1995:BRH

[Tse95] D. S. Tselnik. A bound for the remainder of the Hilbert– Schmidt series and other results on representation of solutions to the functional equation of the second kind with a self-adjoint compact operator as an infinite series. Computers and Mathematics with Applications, 29(10):61–68, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000462.

Tsiotras:1991:EAJ

[Tsi91] George D. Tsiotras. Exact analysis of the joint production quantity-quality of a twostation stable tandem manufacturing system with restricted capacity. Computers and Mathematics with Applications, 21(4):23–31, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190244X.

Tsitouras:1999:HOE

[Tsi99] C. Tsitouras. A high-order explicit Runge–Kutta pair for initial value problems with oscillating solutions. Computers and Mathematics with Applications, 37(6):31–36, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000747.

Thandapani:1996:OTH

[TSL96] E. Thandapani, P. Sundaram, and B. S. Lalli. Oscillation theorems for higher order nonlinear delay difference equations. Computers and Mathematics with Applications, 32(3):111–117, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001162.

Takata:1998:ECS

[TSLW98] S. Takata, Y. Sone, D. Lhuillier, and M. Wakabayashi. Evaporation from or condensation onto a sphere: Numerical analysis of the Boltzmann equation for hardsphere molecules. Computers and Mathematics with Applications, 35(1–2):193– 214, January 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219700268X.

Tsonis:1991:ETR

[Tso91] Anastasios A. Tsonis. The effect of truncation and roundoff on computer simulated chaotic trajectories. Computers and Mathematics with Applications, 21(8):93–94, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900559.

Tsiotras:1992:WCQ

[TT92] George Tsiotras and Charles S. Tapiero. WIP and CSP-1 quality control in a tandem queueing production system. Computers and Mathematics with Applications, 23(1):89–101, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290084U.

Tsai:1996:ACA

[TTC96] Ying Teh Tsai, Chuan Yi Tang, and Yunn Yen Chen. An average case analysis of a greedy algorithm for the on-line Steiner tree problem. Computers and Mathematics with Applications, 31(11):121–131, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000697.

Tang:1996:SMN

[TVZPG96] Y.-F. Tang, L. Vázquez, F. Zhang, and V. M. Pérez-García. Symplectic methods for the nonlinear Schrödinger equation. Computers and Mathematics with Applications, 32(5):73–83, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001368.

Tang:1998:HGA

[TWIF98] Jiafu Tang, Dingwei Wang, A. Ip, and R. Y. K. Fung. A hybrid genetic algorithm for a type of nonlinear programming problem. Computers and Mathematics with Applications, 36 (5):11–21, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001461.

Tian:1998:MOS

[TXC98] Chuan-Jun Tian, Sheng-Li Xie, and Sui Sun Cheng. Measures for oscillatory sequences. Computers and Mathematics with Applications, 36(10–12):149–161, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198800175.

Tang:1999:FRO

[TY99a] X. H. Tang and J. S. Yu. A further result on the oscillation of delay difference equations. Computers and Mathematics with Applications, 38 (11–12):229–237, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199003016.

Tang:1999:ODD

[TY99b] X. H. Tang and J. S. Yu. Oscillation of delay difference equation. Computers and Mathematics with Applications, 37(7):11–20, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000838.

Tseng:1993:FCM

[TYC93] Frank S. C. Tseng, Wei-Pang Yang, and Arbee L. P. Chen. Finding a complete matching with the maximum product on weighted bipartite graphs. Computers and Mathematics with Applications, 25(5):65–71, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193901996.

Tu:1993:IGO

[TZ93] Renjin Tu and Quan Zheng. Integral global optimization method in statistical applications. Computers and Mathematics with Applications, 25(10–11):9–17, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902762.

Uebe:1992:MLN

[Ueb92] Götz Uebe. The model of lang- national income accounts in 1807. Computers and Mathematics with Applications, 24(8–9):167–179, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290196O.

Udwadia:1999:UAR

[UK99] F. E. Udwadia and F. E.

Kalaba. A unified approach for the recursive determination of generalized inverses. Computers and Mathematics with Applications, 37 (1):125–130, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002478.

Ungar:1990:UAS

[Ung90] A. A. Ungar. A unified approach for solving quadratic, cubic and quartic equations by radicals. Computers and Mathematics with Applications, 19(12):33–39, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090248I.

See errata [Ung91].

Ungar:1991:EUA

[Ung91] Abraham A. Ungar. Errata to “A unified approach for solving quadratic, cubic and quartic equations by radicals”. Computers and Mathematics with Applications, 22(2):81, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191900445.

See [Ung90].

Ugrin-Sparac:1991:TSM

[US91] Gordana Ugrin-Sparac. Two simulation models of the birthday problem and certain consequences. Computers and Mathematics with Applications, 22(9):21–25, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190203G.

Usher:1994:OUG

[Ush94a] J. R. Usher. Observations on a useful general nonlinear equation. Computers and Mathematics with Applications, 28(4):99–105, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001316.

Usher:1994:SMM

[Ush94b] J. R. Usher. Some mathematical models for cancer chemotherapy. Computers and Mathematics with Applications, 28(9):73–80, November 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001790.

Usmani:1994:IJT

[Usm94] R. A. Usmani. Inversion of Jacobi’s tridiagonal matrix. Computers and Mathematics with Applications, 27 (8):59–66, April 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194900663.

Ueno:1991:IIS

[UW91] S. Ueno and Alan P. Wang.

Invariant imbedding and searchlight problem in turbid slab. Computers and Mathematics with Applications, 21(11–12):1–6, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901019.

Ueno:1994:IIO

[UW94] S. Ueno and A. P. Wang. Invariant imbedding and orderof-scattering theory in radiation field. Computers and Mathematics with Applications, 27(9–10):175–183, May 1994. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901368.

Ueno:1999:IIC

[UWT99] S. Ueno, A. P. Wang, and M. Toho. Invariant imbedding and Cauchy system of scattering function in searchlight problem in rugged terrain. Computers and Mathematics with Applications, 37 (11–12):101–106, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001480.

Vukina:1994:PFS

[VA94] T. Vukina and J. L. Anderson. Price forecasting with state-space models of nonstationary time series: case of the Japanese salmon market. Computers and Mathematics with Applications, 27(5):45–62, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900752.

Voss:1997:BPC

[VA97] D. Voss and S. Abbas. Block predictor–corrector schemes for the parallel solution of ODEs. Computers and Mathematics with Applications, 33(6):65–72, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000321.

Venkatesan:1991:NTC

[VAB91] R. Venkatesan and S. Al-Bassam. New techniques for constructing t- ECd — EDAUED and t- ECd — EDs

— UED codes. Computers and Mathematics with Applications, 22(1):49–66, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190025Y.

Van:1992:LUT

[Van92] Cuong Le Van. Lagging and unlagging theorems: a quasiequivalence. Computers and Mathematics with Applications, 24(8–9):149–156, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290194M.

Varadarajan:1992:SM

[Var92] V. S. Varadarajan. Symmetry in mathematics. Computers and Mathematics with Applications, 24(3):37–44, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290212Z.

Varga:1998:NUA

[Var98] K. Varga. Nearly unstable AR models with coefficient matrices in Jordan normal form. Computers and Mathematics with Applications, 36(9):1–10, November 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001886.

Vanthournout:1990:FBD

[VBM90] J. Vanthournout, G. Vanden Berghe, and H. De Meyer. Families of backward differentiation methods based on a new type of mixed interpolation. Computers and Mathematics with Applications, 20(11):19–30, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902156.

vanDuin:1995:INS

[vDHO^{+}95] A. C. N. van Duin, P. C. Hansen, Tz. Ostromsky, H. Wijshoff, and Z. Zlatev. Improving the numerical stability and the performance of a parallel sparse solver. Computers and Mathematics with Applications, 30 (12):81–96, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500175X.

vanderHouwen:1994:PPR

[vdHS94] P. J. van der Houwen and B. P. Sommeijer. Preconditioning in parallel Runge– Kutta methods for stiff initial value problems. Computers and Mathematics with Applications, 28(10–12):17–31, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001839.

Vandermeulen:1995:TBD

[VDLM95] E. Vandermeulen, H. A.

Donegan, M. Larnac, and J. Magnier. The temporal Boolean derivative applied to verification of extended finite state machines. Computers and Mathematics with Applications, 30(2):27–36, July 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500075A.

vanderVeen:1995:SPA

[vdV95] W. A. van der Veen. Stepparallel algorithms for stiff initial value problems. Computers and Mathematics with Applications, 30(11):9–23, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219500160Z.

vanderVeen:1999:EAN

[vdVVdB99a] H. van der Veen, C. Vuik, and R. de Borst. An eigenvalue analysis of nonassociated plasticity. Computers and Mathematics with Applications, 38(9–10):107– 115, November 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002667.

vanderVeen:1999:RBN

[vdVVdB99b] H. van der Veen, K. Vuik, and R. de Borst. The relation between numerical and material stress states. Computers and Mathematics with Applications, 38(9– 10):245–249, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002795.

Verma:1999:GHK

[Ver99a] R. U. Verma. G-h-KKM type theorems and their applications to a new class of minimax inequalities. Computers and Mathematics with Applications, 37(8):45–48, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000991.

Verma:1999:GGH

[Ver99b]  R. U. Verma.  General-

ized G–H-convexity and minimax theorems in G–H-spaces. Computers and Mathematics with Applications, 38(1): 13–18, July 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001637.

Vesel:1998:INS

[Ves98] A. Vesel. The independence number of the strong product of cycles. Computers and Mathematics with Applications, 36(7):9–21, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001692.

Vlieg-Hulstman:1995:ESK

[VHH95] M. Vlieg-Hulstman and W. D.

Halford. Exact solutions to KdV equations with variable coefficients and/or nonuniformities. Computers and Mathematics with Applications, 29(1):39–47, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400205Y.

Vichnevetsky:1990:QLT

[Vic90] R. Vichnevetsky. A quantumlike theory of wave propagation in periodic structures. Computers and Mathematics with Applications, 19(8–9):59–75, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090264K.

Voss:1995:PLM

[VK95] D. A. Voss and A. Q. M. Khaliq. Parallel LOD methods for second order time dependent PDEs. Computers and Mathematics with Applications, 30(10):25–35, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500154Q.

Voss:1999:LIP

[VK99] D. A. Voss and A. Q. M. Khaliq. A linearly implicit predictor–corrector method for reaction-diffusion equations. Computers and Mathematics with Applications, 38 (11–12):207–216, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002990.

Vulanovic:1992:NSQ

[VL92] Relja Vulanović and Ping Lin. Numerical solution of quasilinear attractive turning point problems. Computers and Mathematics with Applications, 23(12):75–82, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290093W.

Varma:1995:NQF

[VL95] A. K. Varma and E. Landau. New quadrature formulas based on the zeros of Jacobi polynomials. Computers and Mathematics with Applications, 30(3– 6):213–220, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195000992.

Voudouri:1999:SDM

[VMA99] A. P. Voudouri, J. I. Moshakis, and P. T. Artikis. A stochastic discounting model arising in competing risks management. Computers and Mathematics with Applications, 38(3–4):51–59, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002059.

vanNiekerk:1990:EFE

[vNvN90a] A. van Niekerk and F. D.

van Niekerk. An explicit finite element method for convection–diffusion equations using rational basis functions. Computers and Mathematics with Applications, 19(12):71–77, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090253G.

vanNiekerk:1990:GMR

[vNvN90b] A. van Niekerk and F. D. van Niekerk. A Galerkin method with rational basis functions for Burgers’ equation. Computers and Mathematics with Applications, 20(2):45–51, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090240K.

Voit:1990:SME

[Voi90] E. O. Voit. S-system modelling of endemic infections. Computers and Mathematics with Applications, 20 (4–6):161–173, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090325E.

Vose:1991:PLM

[Vos91] Michael D. Vose. Piecewise linear models of processor utilization. Computers and Mathematics with Applications, 22(8):1–9, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190007Q.

Vyridis:1999:ICA

[VPK99] D. G. Vyridis, S. D. Panteliou, and I. N. Katz. An inverse convergence approach for arguments of Jacobian elliptic functions. Computers and Mathematics with Applications, 37(2):21–26, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198002508.

Venkatarangan:1995:MAD

[VR95a] S. N. Venkatarangan and K. Rajalakshmi. Modification of Adomian’s decomposition method to solve equations containing radicals. Computers and Mathematics with Applications, 29(6): 75–80, March 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500009N.

Venkatarangan:1995:MAS

[VR95b] S. N. Venkatarangan and K. Rajalakshmi. A modification of Adomian’s solution for nonlinear oscillatory systems. Computers and Mathematics with Applications, 29(6):67–73, March 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500008M.

Vasudevan:1999:DPQ

[VV99] R. Vasudevan and L. S. Varadharajan. Dynamic programming and quantum mechanical motion. Computers and Mathematics with Applications, 37(11–12):9–15, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001376.

Wong:1994:EEE

[WA94a] P. J. Y. Wong and R. P. Agarwal. Explicit error estimates for quintic and biquintic spline interpolation II. Computers and Mathematics with Applications, 28(7): 51–69, October 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219400160X.

Wong:1994:SEB

[WA94b] P. J. Y. Wong and R. P. Agarwal. Sharp error bounds for the derivatives of Lidstone-spline interpolation. Computers and Mathematics with Applications, 28 (9):23–53, November 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001766.

Wong:1996:DPS

[WA96a] P. J. Y. Wong and R. P. Agarwal. Double positive solutions of (n, p) boundary value problems for higher order difference equations. Computers and Mathematics with Applications, 32(8):1–21, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001629.

Wong:1996:OCN

[WA96b] P. J. Y. Wong and R. P.

Agarwal. Oscillation criteria for nonlinear partial difference equations with delays. Computers and Mathematics with Applications, 32 (6):57–86, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001447.

Wong:1996:SEB

[WA96c] P. J. Y. Wong and R. P.

Agarwal. Sharp error bounds for the derivatives of Lidstone-spline interpolation II. Computers and Mathematics with Applications, 31(3):61–90, February 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002065.

Wong:1998:ONH

[WA98] P. J. Y. Wong and R. P. Agarwal. Oscillations and nonoscillations of half-linear difference equations generated by deviating arguments. Computers and Mathematics with Applications, 36(10–12): 11–26, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800059.

Wachspress:1991:RFE

[Wac91a] E. L. Wachspress. C^{1} rational finite elements. Computers and Mathematics with Applications, 22(3):11–16, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190064B.

Wachspress:1991:PBA

[Wac91b] E. L. Wachspress. Polynomial bases for algebraic elements. Computers and Mathematics with Applications, 22(3):1–9, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190063A.

Wachspress:1994:TVA

[Wac94] E. L. Wachspress. Threevariable alternating-direction-implicit iteration. Computers and Mathematics with Applications, 27(3):1–7, February 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219490040X.

Wade:1995:WSP

[Wad95] W. R. Wade. A Walsh system for polar coordinates. Computers and Mathematics with Applications, 30(3– 6):221–227, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500100X.

Wagstaff:1990:SUM

[Wag90] S. S. Wagstaff, Jr. Some uses of microcomputers in number theory research. Computers and Mathematics with Applications, 19(3):53–58, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090041H.

Willox:1997:DCC

[WAL97] R. Willox, I. Antoniou, and J. Levitan. Decay of correlations and control of chaotic billiards. Computers and Mathematics with Applications, 34(2–4):391–398, July/August 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001351.

Wang:1990:IIN

[Wan90] A. P. Wang. Invariant imbedding and nonpredictive operators. Computers and Mathematics with Applications, 19(11):75–80, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090150I.

Wang:1991:SDF

[Wan91] Daoliu Wang. Semi-discrete Fourier spectral approximations of infinite dimensional Hamiltonian systems and conservation laws. Computers and Mathematics with Applications, 21(4):63–75, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191902472.

Wang:1993:IPR

[Wan93a]  Alan P. Wang.  An in-

verse problem of reflection. Computers and Mathematics with Applications, 25(9): 59–65, May 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390133G.

Wang:1993:RNN

[Wan93b] J. Wang. Recurrent neural networks for solving linear matrix equations. Computers and Mathematics with Applications, 26(9):23–34, November 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390003E.

Wang:1996:CIS

[Wan96] C. L.-Y. Wang. Computer implementation of a series solution for constant coefficient ordinary differential equations. Computers and Mathematics with Applications, 31(3):43–60, February 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195002057.

Wang:1998:MEC

[Wan98a] Yuan-Ming Wang. Monotone enclosure for a class of discrete boundary value problems without monotone nonlinearities. Computers and Mathematics with Applications, 35(6):51–60, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000170.

Wang:1998:MMB

[Wan98b] Yuan-Ming Wang. Monotone methods for a boundary value problem of secondorder discrete equation. Computers and Mathematics with Applications, 36(6):77–92, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001631.

Wang:1999:PME

[Wan99] Yuan-Ming Wang. Parallel multisplitting explicit AOR methods for numerical solutions of semilinear elliptic boundary value problems. Computers and Mathematics with Applications, 38(5– 6):55–66, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900214X.

Wang:1997:PPS

[WC97] Wendi Wang and Lansun Chen. A predator–prey system with stage-structure for predator. Computers and Mathematics with Applications, 33(8):83–91, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000564.

Wan:1999:RPB

[WCC99] Zhuang Wan, Yubo Chen, and Jie Chen. Remarks on the periodic boundary value problems for first-order differential equations. Computers and Mathematics with Applications, 37(8):49–55, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001005.

Wu:1996:CNF

[WCHL96] Congxin Wu, Lixin Cheng, Minghu Ha, and E. S. Lee. Convexification of nonconvex functions and application to minimum and maximum principles for nonconvex sets. Computers and Mathematics with Applications, 31(7):27–36, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000168.

Wang:1999:MCL

[WCL99] Ching-Te Wang, Chin-Chen Chang, and Chu-Hsing Lin. A method for computing Lucas sequences. Computers and Mathematics with Applications, 38(11–12):187– 196, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002977.

Wu:1994:ACA

[WCY94] T.-C. Wu, C.-C. Chang, and Y.-S. Yeh. An authenticationcombined access control scheme using a one-way function. Computers and Mathematics with Applications, 27(5):63–70, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900760.

Weatherill:1992:DTC

[Wea92] N. P. Weatherill. Delaunay triangulation in computational fluid dynamics. Computers and Mathematics with Applications, 24(5– 6):129–150, September 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290045J.

Weba:1995:CRF

[Web95] M. Weba. Cubature of random fields by product-type integration rules. Computers and Mathematics with Applications, 30(3–6):229– 234, September 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001018.

Wei:1999:TFA

[Wei99] Yuchuan Wei. Triangular function analysis. Computers and Mathematics with Applications, 37(6):37–56, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000759.

Wendroff:1999:TDH

[Wen99]  B. Wendroff.  A two-

dimensional HLLE Riemann solver and associated Godunovtype difference scheme for gas dynamics. Computers and Mathematics with Applications, 38(11–12):175– 185, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002965.

Wolff:1990:VUG

[WF90] J. M. Wolff and S. Fleeter.

Viscous unsteady gust aerodynamics of a flat plate airfoil by a locally analytical method. Computers and Mathematics with Applications, 19(6):27–36, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090045L.

Wolff:1991:AVA

[WF91] James M. Wolff and Sanford Fleeter. Analysis of the viscous aerodynamics of a thin airfoil cascade by a locally analytic method. Computers and Mathematics with Applications, 21(10):1–12, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219190057B.

Wilks:1992:PSF

[WF92] Yorick Wilks and Dann Fass. The preference semantics family. Computers and Mathematics with Applications, 23(2–5):205– 221, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901414.

Wang:1996:SIP

[WF96] Dingwei Wang and S.-C. Fang. A semi-infinite programming model for earliness/tardiness production planning with a genetic algorithm. Computers and Mathematics with Applications, 31(8):95–106, April 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219600034X.

Wang:1999:PMF

[WF99a] P. Wang and R. D. Ferraro. Parallel multigrid finite volume computation of threedimensional thermal convection. Computers and Mathematics with Applications, 37 (9):49–60, May 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001133.

Wu:1999:SCP

[WF99b] Soon-Yi Wu and S.-C. Fang.

Solving convex programs with infinitely many linear constraints by a relaxed cutting plane method. Computers and Mathematics with Applications, 38 (3–4):23–33, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002035.

Wang:1990:ESB

[WHC90] C.-A. Wang, T.-W. Hwang, and Y.-Y. Chen. Existence of solutions for Berman’s equation from laminar flows in a porous channel with suction. Computers and Mathematics with Applications, 20(2):35–40, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090238F.

Wong:1999:MDS

[WHL^{+}99] A. S. M. Wong, Y. C. Hon, T. S. Li, S. L. Chung, and E. J. Kansa. Multizone decomposition for simulation of time-dependent problems using the multiquadric scheme. Computers and Mathematics with Applications, 37(8):23–43, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219900098X.

Wille:1992:CLC

[Wil92] Rudolf Wille. Concept lattices and conceptual knowledge systems. Computers and Mathematics with Applications, 23(6–9):493–515, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192901207.

Winnie:1991:CSS

[Win91] John A. Winnie. Computing structure: Sets, structures, and invariants in LISP. Computers and Mathematics with Applications, 22(10): 49–63, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122191901927.

Witten:1990:MMC

[Wit90]  M. Witten.  Mathemati-

cal modeling and computers in medicine: Editor’s remarks. Computers and Mathematics with Applications, 20 (4–6):xiii–xv, ???? 1990.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090310G.

Witten:1996:P

[Wit96a] M. Witten. Preface. Computers and Mathematics with Applications, 32(1):xiii, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196887074.

Witten:1996:SSI

[Wit96b] M. Witten. The sounds of science: II. Listening to dynamical systems-towards a musical exploration of complexity. Computers and Mathematics with Applications, 32(1):145–173, July 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000971.

Wang:1990:SEC

[WJ90] S.-J. Wang and N. K. Jha.

Systematic t-error correcting/ all unidirectional error detecting codes with easy encoding/decoding. Computers and Mathematics with Applications, 20(1):5–13, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090063P.

Weng:1999:EPS

[WJ99] Peixuan Weng and Daqing Jiang. Existence of positive solutions for boundary value problem of secondorder FDE. Computers and Mathematics with Applications, 37(10):1–9, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001200.

Wimp:1994:NRR

[WK94] J. Wimp and H. Kiesel. Nonlinear recurrence relations and induced orthogonal polynomials. Computers and Mathematics with Applications, 28(1–3):325– 332, August 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001200.

Wang:1999:PMS

[WK99] M.-H. Wang and Y.-E. Kuo. A perturbation method for solving linear semi-infinite programming problems. Computers and Mathematics with Applications, 37(4–5):181– 198, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000681.

Wong:1996:PSS

[WL96] Fu-Hsiang Wong and Wei-Cheng Lian. Positive solution for singular boundary value problems. Computers and Mathematics with Applications, 32(9):41–49, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001757.

Wong:1998:EPM

[Won98a]  P. J. Y. Wong.  Even-

tually positive and monotonely decreasing solutions of partial difference equations. Computers and Mathematics with Applications, 35(4):35–58, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002885.

Wong:1998:TPS

[Won98b]  P. J. Y. Wong.  Triple

positive solutions of conjugate boundary value problems. Computers and Mathematics with Applications, 36 (9):19–35, November 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001904.

Wood:1991:MBA

[Woo91] G. R. Wood. Multidimensional bisection applied to global optimisation. Computers and Mathematics with Applications, 21(6–7):161– 172, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901709.

Wos:1995:RS

[Wos95a] L. Wos. The resonance strategy. Computers and Mathematics with Applications, 29 (2):133–178, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400220F.

Wos:1995:PFA

[Wos95b] Larry Wos. Preface: the field of automated reasoning. Computers and Mathematics with Applications, 29(2):xi–xiv, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195900691.

Wu:1993:CSN

[WR93] Y. Wu and E. Y. Rodin. Control systems in neural networks with local memory. Computers and Mathematics with Applications, 26 (2):1–24, July 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390316N.

Woods:1992:KOF

[WS92] William A. Woods and James G. Schmolze. The KL–ONE family. Computers and Mathematics with Applications, 23(2–5):133– 177, January/March 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192901399.

Wang:1994:EPA

[WSDT94] H. Wang, J. L. Stephenson, Y.-F. Deng, and R. P. Tewarson. An efficient parallel algorithm for solving n-nephron models of the renal inner medulla. Computers and Mathematics with Applications, 28(5):1–12, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001359.

Wang:1998:FPL

[WSNS98] Y. Wang, G. M. Stocks, D. M. C. Nicholson, and W. A. Shelton. The first principles o[ N ] LSMS method and its applications to magnetic structure of alloys. Computers and Mathematics with Applications, 35 (7):85–92, April 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000352.

Wang:1992:RSR

[WSS92] Yeih J. Wang, Leang S. Shieh, and John W. Sunkel. Robust stabilization, robust performance, and disturbance attenuation for uncertain linear systems. Computers and Mathematics with Applications, 23(11):67–80, June 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290070X.

Warsi:1990:AVM

[WT90a] Z. U. A. Warsi and J. F. Thompson. Application of variational methods in the fixed and adaptive grid generation. Computers and Mathematics with Applications, 19(8–9):31–41, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090262I.

Wong:1990:EPF

[WT90b] K. H. Wong and K. L. Teo. An exact penalty function algorithm for time-lag control problems with control and terminal equality constraints. Computers and Mathematics with Applications, 19(12): 79–94, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219090254H.

Wang:1993:QGN

[WT93] H. Wang and R. P. Tewarson. A quasi-Gauss–Newton method for solving nonlinear algebraic equations. Computers and Mathematics with Applications, 25(1): 53–63, January 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390211D.

Wilson:1997:ICP

[WT97] S. J. Wilson and T. S. Tan.

Ignition of carbon particle suspension. Computers and Mathematics with Applications, 33(7):1–10, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000394.

Wu:1990:RDA

[Wu90a] J.-S. Wu. Refining the diffusion approximation for the G/G/c queue. Computers and Mathematics with Applications, 20(11):31–36, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190902167.

Wu:1990:DVA

[Wu90b] Y. Wu. The discrete variational approach to the Euler– Lagrange equation. Computers and Mathematics with Applications, 20(8):61–75, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090210B.

Wang:1994:STD

[WU94] A. P. Wang and S. Ueno. Searchlight on a target with diffuse background II. Computers and Mathematics with Applications, 27(9–10):169– 173, May 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219490135X.

Wu:1998:SOS

[Wu98]  Xin-Yuan Wu.  A sixth-

order A-stable explicit onestep method for stiff systems. Computers and Mathematics with Applications, 35 (9):59–64, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000571.

Wang:1999:III

[WU99a] A. P. Wang and S. Ueno. Invariant imbedding and inverse problems with applications to radiative transfer. Computers and Mathematics with Applications, 37 (11–12):107–112, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001492.

Wu:1999:ETS

[Wu99b] Xian Wu. Existence theorems of solutions for a kind of quasi-variational inequalities with monotone type multivalued mapping. Computers and Mathematics with Applications, 37(10):161–166, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001327.

Watson:1991:LDR

[WW91] Layne T. Watson and C.- Y. Wang. Large deformations of rotating polygonal space structures. Computers and Mathematics with Applications, 22(9):51–59, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190206J.

White:1993:ACF

[WW93] David White and Alan P.

Wang. An algebraic characterization of fixed modes in a decentralized control (a simplified existence proof).

Computers and Mathematics with Applications, 25(6): 57–59, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390299B.

Wang:1995:SSS

[WW95] Ching-An Wang and Tu-Cheng Wu. Similarity solutions of steady flows in a channel with accelerating walls. Computers and Mathematics with Applications, 30 (10):1–16, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500152O.

Wang:1998:FCM

[WW98] Hsiao-Fan Wang and C. H. Wang. A fixed-charge model with fuzzy inequality constraints composed by maxproduct operator. Computers and Mathematics with Applications, 36(7):23–29, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001709.

Wu:1997:NGB

[WX97] Xian Wu and Yu-Guang Xu. New generalizations of Browder’s variational inequalities and the Ky fan minimax inequality. Computers and Mathematics with Applications, 34(1):89–96, July 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197001004.

Wu:1993:CKD

[WY93] Tzong-Chen Wu and Yi-Shiung Yeh. A conference key distribution system based on cross-product. Computers and Mathematics with Applications, 25(4): 39–46, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390246R.

Wu:1999:NVI

[WY99] Xian Wu and Xian-Zhi Yuan.

Nonlinear variational inequalities and implicit variational inequality of Ky fan type in H-spaces. Computers and Mathematics with Applications, 38(7–8):1–8, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002321.

Wyatt:1994:CMW

[Wya94] B. Wyatt. Collisions of microdrops of water. Computers and Mathematics with Applications, 28(10–12):175–208, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001928.

Xu:1992:IPR

[XC92] Yuan Xu and E. W. Cheney. Interpolation by periodic radial functions. Computers and Mathematics with Applications, 24 (12):201–215, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290181G.

Xiao:1997:SAE

[XL97] Tijun Xiao and Jin Liang. Semigroups arising from elastic systems with dissipation. Computers and Mathematics with Applications, 33 (10):1–9, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000722.

Xu:1995:OID

[XLG95] C.-Z. Xu, P. Ligarius, and J.-P. Gauthier. An observer for infinite-dimensional dissipative bilinear systems. Computers and Mathematics with Applications, 29(7): 13–21, April 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500014P.

Luo:1996:RSF

[xLhW96] Zhong xuan Luo and Ren hong Wang. Rational shape functions for C^{1}- interpolation on quadrilaterals. Computers and Mathematics with Applications, 32 (9):55–66, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001770.

Xiao:1998:MLP

[XLJS98] Wei Xiao, Zhibin Liu, Maoze Jiang, and Y. Shi. Multiobjective linear programming model on injection oilfield recovery system. Computers and Mathematics with Applications, 36(5):127–135, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001552.

Xu:1999:DAN

[XLPG99] Daoyi Xu, Shuyong Li, Zhiling Pu, and Qingyi Guo. Domain of attraction of nonlinear discrete systems with delays. Computers and Mathematics with Applications, 38(5–6):155–162, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002229.

Xu:1990:BVP

[XS90] Y.-H. Xu and E. P. Salathe. Boundary value problems in microcirculatory physiology. Computers and Mathematics with Applications, 20 (4–6):175–197, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090326F.

Xu:1997:CFB

[Xu97] Y. Xu. Commuting family of block Jacobi matrices. Computers and Mathematics with Applications, 33 (1–2):211–222, January 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002313.

Xue:1994:GCA

[Xue94] G.-L. Xue. A globally convergent algorithm for facility location on a sphere. Computers and Mathematics with Applications, 27(6):37–50, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901090.

Yamaguti:1994:DMS

[Yam94] M. Yamaguti. Discrete models in social sciences. Computers and Mathematics with Applications, 28(10–12):263– 267, November/December 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001960.

Yan:1991:DTL

[Yan91] Song Yuan Yan. Declarative testing of logic databases. Computers and Mathematics with Applications, 22(11): 39–45, ???? 1991. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190032Y.

Yang:1993:OGO

[Yan93a] H. H. Yang. A one-grid-overlapped spectral multidomain method for the PDEs. Computers and Mathematics with Applications, 26 (10):27–33, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390023O.

Yang:1993:CPG

[Yan93b] M. S. Yang. Convergence properties of the generalized fuzzy c-means clustering algorithms. Computers and Mathematics with Applications, 25(12):3–11, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390181T.

Yang:1993:PDE

[Yan93c] Yi Xian Yang. Period distribution for error-correcting codes. Computers and Mathematics with Applications, 26(12):1–6, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390053X.

Yanushevsky:1993:LFR

[Yan93d] R. T. Yanushevsky. Lyapunov’s functionals and related optimal problems for differential-difference systems. Computers and Mathematics with Applications, 25(10–11):89–101, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122193902854.

Yan:1994:NLAb

[Yan94] S. Y. Yan. 68 new large amicable pairs. Computers and Mathematics with Applications, 28(5):71–74, September 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194001421.

See errata [Ano96-170].

Yan:1995:PTL

[Yan95] S. Y. Yan. Primality testing of large numbers in Maple. Computers and Mathematics with Applications, 29 (12):1–8, June 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500052Z.

Yang:1997:DDD

[Yan97] D. Yang. Dynamic domain decomposition and grid modification for parabolic problems. Computers and Mathematics with Applications, 33(10):89–103, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000795.

Yang:1999:TEB

[Yan99a]  Tianruo Yang.  Theoret-

ical error bounds on the convergence of the Lanczos and block-Lanczos methods. Computers and Mathematics with Applications, 38(9– 10):19–38, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900259X.

Yanushevsky:1999:CNT

[Yan99b] R. T. Yanushevsky. A class of nonlinear time-delay systems and related optimal problems. Computers and Mathematics with Applications, 37 (4–5):73–78, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000607.

Yao:1992:NIB

[Yao92] Jen-Chih Yao. Nonlinear inequalities in Banach spaces. Computers and Mathematics with Applications, 23(12): 95–98, June 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290095Y.

Yao:1993:AVI

[Yao93] Jen-Chih Yao. Abstract variational inequality problems and a basic theorem of complementarity. Computers and Mathematics with Applications, 25(1):73–79, January 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390213F.

Yasuda:1994:LSC

[Yas94] M. Yasuda. A linear structure and convexity for relations in dynamic fuzzy systems. Computers and Mathematics with Applications, 27(9–10):23–27, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194901198.

Yasar:1998:P

[Yas98a] O. Yasar. Preface. Computers and Mathematics with Applications, 35(7):xiii, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000273.

Yasar:1998:SMC

[Yas98b] O. Yasar. A scalable model for complex flows. Computers and Mathematics with Applications, 35(7):117–128, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219800039X.

Yavin:1991:OLE

[Yav91a] Y. Yavin. Open-loop evasion strategies in a pursuitevasion problem in a reduced state space. Computers and Mathematics with Applications, 22(11):3–13, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900294.

Yavin:1991:OLC

[Yav91b] Y. Yavin. Optimal launch conditions: an optimal stochastic control problem.

Computers and Mathematics with Applications, 21 (6–7):115–126, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191901662.

Yavin:1992:SRM

[Yav92] Yaakov Yavin. A search for a randomly moving object: a numerical study. Computers and Mathematics with Applications, 23(10):41–49, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290053K.

Yavin:1993:ASD

[Yav93] Y. Yavin. Applications of stochastic differential games to the suboptimal design of pulse motors. Computers and Mathematics with Applications, 26(6):87–95, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390120K.

Yavin:1996:NCT

[Yav96] Y. Yavin. Navigation and control of a trolley by using a dynamical model. Computers and Mathematics with Applications, 32(7):47–55, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001551.

Yang:1992:MEU

[YC92] Sheng-An Yang and Cha’o-Kuang Chen. Magnetohydrodynamic effects upon momentum transfer on a continuously moving flat plane. Computers and Mathematics with Applications, 23(10): 27–33, May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290051I.

Yang:1994:CRF

[YC94] Miin-Shen Yang and Chih-Tien Chen. Convergence rate of the fuzzy generalized nearest neighbor rule. Computers and Mathematics with Applications, 27(5):1–8, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219490071X.

Yang:1996:PPP

[YCC96] Xia Yang, Lansun Chen, and Jufang Chen. Permanence and positive periodic solution for the single-species nonautonomous delay diffusive models. Computers and Mathematics with Applications, 32(4):109–116, August 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196001290.

Yu:1998:ONN

[YCZ98] J. S. Yu, Ming-Po Chen, and H. Zhang. Oscillation and nonoscillation in neutral equations with integrable coefficients. Computers and Mathematics with Applications, 35(6):65–71, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000194.

Ye:1999:LSF

[Ye99]  X. Ye.  A least-squares

finite-element method for the Stokes equations with improved mass balances. Computers and Mathematics with Applications, 38(3–4):229– 237, August 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001972.

Yavin:1992:OSC

[YF92a] Y. Yavin and C. Frangos. Optimal selection and control strategies for a flexible manufacturing and assembly system. Computers and Mathematics with Applications, 23(10):73–78, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290057O.

Yavin:1992:OSP

[YF92b] Y. Yavin and C. Frangos.

An optimal selection policy for a flexible manufacturing cell feeding several production lines. Computers and Mathematics with Applications, 23(10):57–63, May 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290055M.

Yavin:1993:PSM

[YF93] Y. Yavin and C. Frangos.

The performance of some manufacturing queueing networks when some of the machines are in the failure mode. Computers and Mathematics with Applications, 25(4):47–54, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390247S.

Yavin:1992:PPWb

[YFF92a] Y. Yavin, C. Frangos, and J. P. Fourie. The performance of a projectile which uses a bang-bang type guidance law: Part 2 — a semirigid body model. Computers and Mathematics with Applications, 24(4):85–92, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290010F.

Yavin:1992:PPWa

[YFF92b] Y. Yavin, C. Frangos, and J. P. Fourie. The performance of a projectile which uses a bang-bang type guidance law:Part 1. Computers and Mathematics with Applications, 23(1):111–118, January 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290086W.

Yavin:1993:PPW

[YFF93] Y. Yavin, C. Frangos, and J. P. Fourie. The performance of a projectile which uses a bang-bang type guidance law. part 3 — the point-mass model revisited. Computers and Mathematics with Applications, 25(4): 27–37, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390245Q.

Yavin:1995:CFCa

[YFZM95a] Y. Yavin, C. Frangos, G. Zilman, and T. Miloh. Computation of feasible command strategies for the navigation of a ship in a narrow zigzag channel. Computers and Mathematics with Applications, 29(9):1–23, May 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500034V.

Yavin:1995:CFCb

[YFZM95b] Y. Yavin, C. Frangos, G. Zilman, and T. Miloh. Computation of feasible command strategies for the navigation of a ship in a narrow zigzag channel. Computers and Mathematics with Applications, 30(10):79–101, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500157T.

Yoneyama:1997:TRI

[YI97] K. Yoneyama and H. Ishii.

The throughput rate of interchangeable parallel two-stage tandem queue with correlated service times. Computers and Mathematics with Applications, 33(7):29–37, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000412.

Yan:1994:NLAa

[YJ94] S. Y. Yan and T. H. Jackson. A new large amicable pair. Computers and Mathematics with Applications, 27(6):1–3, March 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901058.

Yan:1999:FGO

[YJ99] S. Y. Yan and G. James. Fast group operations on elliptic curves in Maple. Computers and Mathematics with Applications, 37(8):129–138, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001066.

Yon:1995:NMM

[YK95] Y. J. Yon and D. Y. Kwak. Nonconforming multigrid method for nonsymmetric and indefinite problems. Computers and Mathematics with Applications, 30(11):1–7, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500159V.

Yen:1993:SAH

[YL93a] S.-M. Yen and C.-S. Laih. Server-aided honest computation for cryptographic applications. Computers and Mathematics with Applications, 26(12):61–64, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122193900595.

Yen:1993:DDA

[YL93b] Sung-Ming Yen and Chi-Sung Laih. The design of dynamic access control scheme with user authentication. Computers and Mathematics with Applications, 25(7):27–32, April 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390309J.

Yun:1999:DLP

[YL99] B. I. Yun and S. Lee. Double layer potential scheme for Dirichlet problems on smooth open arcs. Computers and Mathematics with Applications, 37(7):31–40, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000851.

Yun:1996:MBI

[YLC96] B. I. Yun, S. Lee, and U. J.

Choi. A modified boundary integral method on open arcs in the plane. Computers and Mathematics with Applications, 31(11):37–43, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000594.

Yu:1996:GNV

[YLY96] S. J. Yu, T. C. Lai, and J. C. Yao. Generalized nonlinear variational inequalities. Computers and Mathematics with Applications, 32(7):21–27, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196001526.

Yasui:1992:RCE

[YNK92] Kazumi Yasui, Toshio Nakagawa, and Shin-Ichi Koike. Reliability consideration on error control policies for a data communication system. Computers and Mathematics with Applications, 24(1– 2):51–55, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902279.

Yamada:1992:TDD

[YOO92] Shigeru Yamada, Hiroshi Ohtera, and Mitsuru Ohba. Testing-domain dependent software reliability models. Computers and Mathematics with Applications, 24(1– 2):79–86, July 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902316.

Yoon:1993:CFO

[Yoo93] D. H. H. Yoon. The categorical framework of objectoriented concurrent systems. Computers and Mathematics with Applications, 25(2): 33–38, January 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390220P.

Yoshihara:1990:CEP

[Yos90] Ken-Ichi Yoshihara. Conditional empirical processes defined by ϕ — mixing sequences. Computers and Mathematics with Applications, 19(1):149–158, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900952.

Yoshida:1993:OSP

[Yos93] Y. Yoshida. Optimal stopping problems for multiarmed bandit processes with arms’ independence. Computers and Mathematics with Applications, 26 (12):47–60, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900584.

Yoshida:1996:OSP

[Yos96] Y. Yoshida. An optimal stopping problem in dynamic fuzzy systems with fuzzy rewards. Computers and Mathematics with Applications, 32 (10):17–28, November 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196001836.

Yoshida:1999:SFS

[Yos99] Y. Yoshida. Superharmonic fuzzy sets on recurrent sets in dynamic fuzzy systems. Computers and Mathematics with Applications, 37 (11–12):87–93, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001467.

Yan:1998:UIC

[YP98a] G. Yan and P. Y. H. Pang. Uniqueness of the inverse conductive scattering problem. Computers and Mathematics with Applications, 36(8):53–57, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001825.

Yan:1998:UIO

[YP98b] G. Yan and P. Y. H. Pang. The uniqueness of the inverse obstacle scattering problem with transmission boundary conditions. Computers and Mathematics with Applications, 36(3):63–69, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198001291.

Yang:1999:GSR

[YQM99] Biao Yang, L. Qiu, and T. Mitsui. GP _{G}-stability of Runge–Kutta methods for generalized delay differential systems. Computers and Mathematics with Applications, 37(7):89–97, April 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000899.

Yalcinalp:1990:DJE

[YS90] L. Ü. Yalçinalp and L. Sterling. Diagnosing jaundice expert system. Computers and Mathematics with Applications, 20(9–10):125–140, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190901173.

Yamazaki:1992:MIC

[YSM92] Genji Yamazaki, Karl Sigman, and Masakiyo Miyazawa. Moments in infinite channel queues. Computers and Mathematics with Applications, 24(1–2):1–6, July 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122192902213.

Yamada:1998:ODP

[YT98] Y. Yamada and Y. Teraoka. An optimal design of piping route in a CAD system for power plant. Computers and Mathematics with Applications, 35(6):137–149, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S089812219800025X.

Yamashiro:1994:ASB

[YTLK94] S. M. Yamashiro, G. A. Taylor, S. K. M. Luk, and R. E. Kalaba. Automated solutions of breathing pattern optimizations. Computers and Mathematics with Applications, 27(9–10):155–161, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901333.

Yu:1998:ASL

[Yu98] J. S. Yu. Asymptotic stability for a linear difference equation with variable delay. Computers and Mathematics with Applications, 36(10–12):203– 210, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800217.

Yabe:1995:DCS

[YX95] T. Yabe and F. Xiao. Description of complex and sharp interface with fixed grids in incompressible and compressible fluid. Computers and Mathematics with Applications, 29(1):15–25, January 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219400203W.

Yu:1998:SPE

[YY98] J. Yu and G. X.-Z. Yuan. The study of Pareto equilibria for multiobjective games by fixed point and Ky fan minimax inequality methods. Computers and Mathematics with Applications, 35(9):17–24, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000534.

Yavin:1994:FSS

[YZM94] Y. Yavin, G. Zilman, and T. Miloh. A feasibility study of ship maneuverability in the vicinity of an obstacle: a stochastic control approach. Computers and Mathematics with Applications, 28(8):63–76, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122194001707.

Yavin:1995:SMF

[YZM95] Y. Yavin, G. Zilman, and T. Miloh. Ship maneuverability in finite depth water in the vicinity of an obstacle: a stochastic control approach. Computers and Mathematics with Applications, 30(10): 103–117, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500158U.

Zhou:1995:SSP

[ZA95] M. Zhou and G. A. Anastassiou. Some saturation properties of the representation formulae for (C_{0})mparameter operator semigroups. Computers and Mathematics with Applications, 30(3–6):243–253, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122195001034.

Zadrozny:1989:ADE

[Zad89]  P. A. Zadrozny.  Ana-

lytic derivatives for estimation of linear dynamic models. Computers and Mathematics with Applications, 18 (6–7):539–553, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122189901065.

See errata [Zad92].

Zadrozny:1992:EAD

[Zad92] Peter A. Zadrozny. Errata to “Analytic derivatives for estimation of linear dynamic models”: Computers and Mathematics with Applications, Vol. 18, No. 6/7, pp. 539–553, 1989. Computers and Mathematics with Applications, 24(8–9):289–290, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290207X.

See [Zad89].

Zadeh:1999:FLC

[Zad99] L. A. Zadeh. Fuzzy logic and the calculi of fuzzy rules, fuzzy graphs, and fuzzy probabilities. Computers and Mathematics with Applications, 37(11–12):35, June 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001406.

Zafer:1998:EPS

[Zaf98a] A. Zafer. The existence of positive solutions and oscillation of solutions of higherorder difference equations with forcing terms. Computers and Mathematics with Applications, 36(10–12):27–35, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800060.

Zafer:1998:NSC

[Zaf98b] A. Zafer. Necessary and sufficient condition for oscillation of higher order nonlinear delay difference equations. Computers and Mathematics with Applications, 35(10):125–130, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198000789.

Zalik:1995:SRS

[Zal95] R. A. Zalik. Some remarks on spectral approximation. Computers and Mathematics with Applications, 30(3– 6):235–241, September 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001026.

Zarri:1992:DCH

[Zar92] Gian Piero Zarri. The “descriptive” component of a hybrid knowledge representation language. Computers and Mathematics with Applications, 23(6–9):697– 718, March/May 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219290130A.

Zhang:1995:TMAa

[ZB95] J. Zhang and J. A. Belward. Tau-method approximations for the Bessel function Y_{0}(z). Computers and Mathematics with Applications, 30(7):5–14, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500120N.

Zappe:1991:AGB

[ZC91] Christopher J. Zappe and A. Victor Cabot. The application of generalized Benders decomposition to certain nonconcave programs. Computers and Mathematics with Applications, 21 (6–7):181–190, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190172Z.

Zhang:1996:PGS

[ZC96] Jingru Zhang and Lansun Chen. Permanence and global stability for a two-species cooperative system with time delays in a two-patch environment. Computers and Mathematics with Applications, 32 (12):101–108, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002106.

Zhang:1998:NRT

[ZC98] Shunian Zhang and Ming-Po Chen. A new Razumikhin theorem for delay difference equations. Computers and Mathematics with Applications, 36(10–12):405–412, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800412.

Zhang:1999:PSC

[ZC99] Xin-An Zhang and Lansun Chen. The periodic solution of a class of epidemic models. Computers and Mathematics with Applications, 38 (3–4):61–71, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002060.

Zhuang:1996:MMD

[ZCC96] Wan Zhuang, Yubo Chen, and Sui Sun Cheng. Monotone methods for a discrete boundary problem. Computers and Mathematics with Applications, 32 (12):41–49, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122196002064.

Zedan:1990:AEJ

[Zed90] H. Zedan. Avoiding the exactness of the Jacobian matrix in Rosenbrock formulae. Computers and Mathematics with Applications, 19(2):83–89, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122190900118.

Zedan:1999:ETF

[Zed99] H. A. Zedan. Equivalence transformation for the filtration problem. Computers and Mathematics with Applications, 37(6):1–5, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199000711.

Zellner:1992:BCC

[Zel92] Arnold Zellner. Brief comment on C. A. Los’ papers. Computers and Mathematics with Applications, 24(8– 9):275, October/November 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290205V.

See [Los89b, Los89a].

Zlatev:1996:RBE

[ZFM96] Z. Zlatev, J. Fenger, and L. Mortensen. Relationships between emission sources and excess ozone concentrations. Computers and Mathematics with Applications, 32 (11):101–123, December 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196002027.

Zhang:1997:PML

[ZG97] A. X. Zhang and Sy-Ming Guu. Properties of the multiple lot-sizing problem with rigid demands and general yield distributions. Computers and Mathematics with Applications, 33(5): 55–65, March 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197000199.

Zhang:1991:HTQ

[Zha91] Fenggang Zhang. A hyperbolic tangent quadrature rule for solving singular integral equations with Hadamard finite part integrals. Computers and Mathematics with Applications, 22(3):59–73, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190070K.

Zhang:1993:IBA

[Zha93a] C. N. Zhang. An improved binary algorithm for RSA. Computers and Mathematics with Applications, 25(6): 15–24, March 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902957.

Zhang:1993:PPT

[Zha93b] Liqing Zhang. Phase path tracing methods for solving the high order zeros of nonlinear systems. Computers and Mathematics with Applications, 26(5):45–50, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900724.

Zhang:1993:PIM

[Zha93c]  Zhenyue Zhang.  A par-

allel iterative method for solving symmetric tridiagonal extreme eigenpair problems. Computers and Mathematics with Applications, 26 (12):35–45, December 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900573.

Zhang:1993:SMS

[Zha93d] Zhenyue Zhang. A shifted method for solving symmetric eigenvalue problems. Computers and Mathematics with Applications, 25(4): 75–86, February 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390250Y.

Zhang:1995:TMAb

[Zha95] J. Zhang. Tau-method approximations for the Bessel function Y_{1}(z). Computers and Mathematics with Applications, 30(7):15–19, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219500121E.

Zhang:1996:NTM

[Zha96] J. Zhang. A note on the tau-method approximations for the Bessel functions Y_{0}(z) and Y_{1}(z). Computers and Mathematics with Applications, 31(9):63–70, May 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122196000430.

Zhang:1997:SAD

[Zha97]  Shunian Zhang.  Stabil-

ity analysis of delay difference systems. Computers and Mathematics with Applications, 33(10):41–52, May 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197000758.

Zhang:1998:NCA

[Zha98] Z. Q. Zhang. Notes and corrections on “Asymptotic constancy criteria of solution of linear parabolic Volterra difference equations”. Computers and Mathematics with Applications, 35(6):61–63, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122198000182.

See [SWY96a].

Zhang:1999:ETS

[Zha99] Shunian Zhang. Estimate of total stability of delay difference systems. Computers and Mathematics with Applications, 37(9):31–38, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219900111X.

Zheng:1991:RAG

[Zhe91] Quan Zheng. Robust analysis and global optimization. Computers and Mathematics with Applications, 21(6–7):17–24, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190157Y.

Zheng:1993:DMR

[Zhe93] Quan Zheng. Discontinuity and measurability of robust functions in the integral global minimization. Computers and Mathematics with Applications, 25(10– 11):79–88, May/June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193902843.

Zhou:1990:PGA

[Zho90] C. Zhou. A permutationgenerating algorithm. Computers and Mathematics with Applications, 20(1):39–42, ???? 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219090067T.

Zhu:1992:NCE

[Zhu92]  Li-Xing Zhu.  A note

on the consistent estimator of nonparametric regression constructed by splines. Computers and Mathematics with Applications, 24 (11):65–70, December 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219290032D.

Zheleznov:1993:ADD

[ZIKM93] V. S. Zheleznov, M. N. Ivanov, E. A. Kurskii, and E. P. Maslov. Avoidance of detection in 3-D. Computers and Mathematics with Applications, 26 (6):55–66, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390117E.

Zuberek:1999:TPN

[ZK99] W. M. Zuberek and W. Kubiak. Timed Petri nets in modeling and analysis of simple schedules for manufacturing cells. Computers and Mathematics with Applications, 37(11–12):191– 206, June 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001571.

Zafer:1998:ABH

[ZKÖ98] A. Zafer, B. Kaymakçalan, and S. A. Özgün. Asymptotic behavior of higher-order nonlinear equations on time scales. Computers and Mathematics with Applications, 36(10–12):299–306, November/December 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812219880031X.

Zhang:1998:OPD

[ZL98a] B. G. Zhang and S. T. Liu.

Oscillation of partial difference equations with variable coefficients. Computers and Mathematics with Applications, 36(10–12):235–242, November/December 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198800242.

Zhang:1998:OTS

[ZL98b] Zhenguo Zhang and Qiaoluan Li. Oscillation theorems for second-order advanced functional difference equations. Computers and Mathematics with Applications, 36 (6):11–18, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001576.

Zhang:1999:NSC

[ZL99a] B. G. Zhang and B. M. Liu. Necessary and sufficient conditions for oscillations of partial difference equations with continuous variables. Computers and Mathematics with Applications, 38(5– 6):163–167, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002230.

Zhang:1999:OCC

[ZL99b] B. G. Zhang and B. M. Liu. Oscillation criteria of certain nonlinear partial difference equations. Computers and Mathematics with Applications, 38(11–12):107– 112, December 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002904.

Zhi:1999:IBP

[ZL99c] Lihong Zhi and Zhuojun Liu.

P -irreducibility of binding polynomials. Computers and Mathematics with Applications, 38(2):1–10, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001765.

Zadrozny:1994:KFM

[ZM94] P. A. Zadrozny and S. Mittnik. Kalman-filtering methods for computing information matrices for timeinvariant, periodic, and generally time-varying VARMA models and samples. Computers and Mathematics with Applications, 28(4):107–119, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122194001324.

See comments [Ter00].

Zheng:1996:IFC

[ZM96] H. Zheng and D. A. Murio. 3D-IHCP on a finite cube. Computers and Mathematics with Applications, 31(1):1–14, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122195001793.

Zhan:1998:IPO

[ZM98] S. Zhan and D. A. Murio. Identification of parameters in one-dimensional IHCP. Computers and Mathematics with Applications, 35(3):1–16, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002745.

Zhan:1999:SFN

[ZM99] S. Zhan and D. A. Murio. Surface fitting and numerical gradient computations by discrete mollification. Computers and Mathematics with Applications, 37(9):85–102, May 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199001169.

Zhang:1992:EUI

[ZO92] Yanchun Zhang and Maria E. Orlowska. The effect of unary inclusion dependencies on relational database design. Computers and Mathematics with Applications, 24(3):49–59, August 1992. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122192902143.

Zou:1991:OGE

[Zou91] Qisu Zou. An observation on Gauss elimination. Computers and Mathematics with Applications, 22(11):69–70, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122191900353.

Zhang:1993:ESS

[ZQ93a] Mei-Qing Zhang and Meng-Zhao Qin. Explicit symplectic schemes to solve vortex systems. Computers and Mathematics with Applications, 26(5):51–56, September 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193900735.

Zhu:1993:AHO

[ZQ93b] W.-J. Zhu and M.-Z. Qin.

Applications of higher order self-adjoint schemes to partial differential equations. Computers and Mathematics with Applications, 26(3): 15–26, August 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122193901066.

Zhu:1993:CHO

[ZQ93c] Wen-Jie Zhu and Meng-Zhao Qin. Constructing higher order schemes by formal power series. Computers and Mathematics with Applications, 25(12):31–38, June 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390183V.

Zhu:1993:OCT

[ZQ93d] Wen-Jie Zhu and Meng-Zhao Qin. Order conditions of two kinds of canonical difference schemes. Computers and Mathematics with Applications, 25(6):61–74, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219390300K.

Zhu:1994:PSH

[ZQ94] Wenjie Zhu and Mengzhao Qin. Poisson schemes for Hamiltonian systems on Poisson manifolds. Computers and Mathematics with Applications, 27(12):7–16, June 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194900817.

See [McL95] and reply [ZQ95].

Zhu:1995:RCP

[ZQ95] W.-J. Zhu and M.-Z. Qin. Reply to “Comment on ‘Poisson schemes for Hamiltonian systems on Poisson manifolds”’. Computers and Mathematics with Applications, 29(7):1, April 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219500012N.

See [ZQ94].

Zou:1991:NSS

[ZR91] Q. Zou and A. G. Ramm.

Numerical solution of some inverse scattering problems of geophysics. Computers and Mathematics with Applications, 21(2–3):75–80, ???? 1991. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219190083G.

Zamani:1993:ECF

[ZS93] Nader G. Zamani and Weiwei Sun. Enriched collocation finite element method, a model problem. Computers and Mathematics with Applications, 25(5):73–81, March 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/089812219390200F.

Zhou:1997:OMS

[ZS97] Zongfang Zhou and Y. Shi.

An ODE method of solving nonlinear programming. Computers and Mathematics with Applications, 34(1): 97–102, July 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197001016.

Zhou:1999:NST

[ZSDB99] X. Zhou, Z. Sun, F. Durst, and G. Brenner. Numerical simulation of turbulent jet flow and combustion. Computers and Mathematics with Applications, 38(9– 10):179–191, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002734.

Zhang:1998:NEP

[ZT98a] B. G. Zhang and Chuan Jun Tian. Nonexistence and existence of positive solutions for difference equations with unbounded delay. Computers and Mathematics with Applications, 36(1):1–8, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001035.

Zhang:1998:OCD

[ZT98b] B. G. Zhang and Chuan Jun Tian. Oscillation criteria for difference equations with unbounded delay. Computers and Mathematics with Applications, 35(4):19–26, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197002861.

Zhang:1999:SCC

[ZT99] B. G. Zhang and Chuan Jun Tian. Stability criteria for a class of linear delay partial difference equations. Computers and Mathematics with Applications, 38(2):37–43, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001807.

Zhou:1996:OPH

[ZY96] Xiaoliang Zhou and Jurang Yan. Oscillatory property of higher order nonlinear difference equations. Computers and Mathematics with Applications, 31(12):61–68, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122196000764.

Zhang:1998:LOT

[ZY98] B. G. Zhang and J. S. Yu.

Linearized oscillation theorems for certain nonlinear delay partial difference equations. Computers and Mathematics with Applications, 35 (4):111–116, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122197002940.

Zhang:1998:ODE

[ZYC98] B. G. Zhang, Jun Yan, and S. K. Choi. Oscillation for difference equations with continuous variable. Computers and Mathematics with Applications, 36(9):11–18, November 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001898.

Zhou:1998:GAN

[ZYW98] Zhan Zhou, Jianshe Yu, and Zhicheng Wang. Global attractivity of neutral difference equations. Computers and Mathematics with Applications, 36(6):1–10, September 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122198001564.

Zak:1994:UDC

[ZZ94] M. Zak and A. P. Zak. Unpredictable dynamics and collective brain. Computers and Mathematics with Applications, 27(9–10):185–197, May 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0898122194901376.

Zhang:1999:SSD

[ZZ99a] Bing Gen Zhang and Yong Zhou. The semicycles of solutions of delay difference equations. Computers and Mathematics with Applications, 38 (7–8):31–38, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002357.

Zhang:1999:SAL

[ZZ99b] C. J. Zhang and S. Z.

Zhou. Stability analysis of LMMs for systems of neutral multidelay-differential equations. Computers and Mathematics with Applications, 38 (3–4):113–117, August 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199002096.

Zhang:1999:OCS

[ZZ99c] Zhenguo Zhang and Jinlian Zhang. Oscillation criteria for second-order functional difference equations with “summation small” coefficient. Computers and Mathematics with Applications, 38(1):25–31, July 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122199001650.

Zheng:1999:GMC

[ZZ99d] Quan Zheng and Liansheng Zhang. Global minimization of constrained problems with discontinuous penalty functions. Computers and Mathematics with Applications, 37 (4–5):41–58, February/March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199000589.

Zhou:1999:GAN

[ZZ99e] Zhan Zhou and Qinqin Zhang. Global attractivity of a nonautonomous logistic difference equation with delay. Computers and Mathematics with Applications, 38 (7–8):57–64, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002382.

Zhou:1999:GSA

[ZZT99] S. Zhou, J. Zeng, and X. Tang. Generalized Schwarz algorithm for obstacle problems. Computers and Mathematics with Applications, 38 (7–8):263–271, October 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122199002552.